A Complete Bibliography of Publications in Probability Theory and Related Fields

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

19 October 2024 Version 1.42

Title word cross-reference

(1 + 1) [DGZ05]. (2 + 1) [CMT15]. (2 + 1)d [BCF18]. (Z/NZ)^{2} [Rod19]. (t, κ) [FTY21]. + [Gia95]. 0 [Dow88, Rod17]. 0 < α < 1 [DD88]. 0 < d < 1 [Ber90b]. 0 < p ≤ 1 [Wei90]. 1 [Ald86, DD16, Dow88, DL20, FJL09, Rod17, SS15]. 1/15 [CGN14]. 2 [AST19, ALS20, BF11, BCF15, Bur89, CL07, CM11a, CV12, DS11, FL19, Hry98, Kor04, LZ10, Leb17, Mél01, PV97, TW20, Yao19, vdHL10]. 2L × 2L [LT15]. 2M − X [CC21]. 2 × 2 [BD23]. 3 [DS20, DR14a, FR08, FG19, Han12, LZ10, LS97b, Ndu91, RZ09]. 4 + 1 [vdHdHS07]. ^{∗} [BG97b]. ^{d} [BP89, Tan97]. ^{∞} [Nis97]. _{6} [CN07]. A [Kat15]. Airy_{1} [QR13]. α [BY05, BKK08, DD88, GVA86, GM06]. α = 0 [BGL19]. B [Ein88]. BES_{0}(d) [Ber90b]. β [AH20, Ahn22, ABZ23, GZ18, HL19]. mod1 [Sch88a]. C [HNY99, Iwa87]. C[0, 1] [vV89]. C[0, 1]^{Z^{d}} [Deu87]. C^{1} [Kim04, RV96]. C^{1,λ} [ERV02]. C^{2} [DT12]. C_{0} [Blu05]. d [CM13, LLT^{+}19, TY03]. D(R) [SK86]. δ [IV00]. Δ_{2} [KN98]. Δϕ [FN01, Fun03, Nis03, Zam04]. Δu = 4u^{2} [AD04]. Δu = u^{2} [Le 95a]. Δφ [DD05, DGI00]. d ≥ 3 [MS17b]. d ≥ 3 [CLO01, HTV12, JR08, Pov99, QRV99, Sab11]. d ≥ 5 [AC07]. E [CSK99]. exp(Φ)_{2} [HKK23]. f(n_{k}x) [AB10]. F_{Dt} [Mai89]. \frac{∞_{x≥y}−∞_{x≤y}}{4} − − − \frac{3(x−y)(1+xy)}{8} [SS24]. G [She93, CHMP02, JZ10, JP10]. G(n, d/n) [MS10]. Γ [Le99]. GL(N, C) [Ahn23]. h [KP89, Mik04, Von92]. H^{p} [RS94]. ∞ [AB11]. K [Tal01a, BL19, CM88, Deh86, Deh88, Fal90, HHT21, Hou16, Kol03]. κ [FS17, VR14]. k ≤ c log_{2} n [Deh86]. L′ [BG86b, HJS88, Wat96]. L^{1} [HKK23]. L^{2} [BN13, FT04]. L^{p} [DMS05, KKT10, Li08, Li11, Li14b, RS94, Li14a]. L_{0} [Jur88]. l_{1} [BMN12]. L_{2} [Gri95, HP18, Li92, NN04, Por95]. L_{d} [Kry21]. L_{∞} [Kaz88, Kon90]. L_{p} [CH88b, GT98b, Hit90, Kim04, Kry10, DJ94]. L_{r} [HJM20, LNS99]. Λ [GPW09]. M [Cla86, DM16b, LBE18, WZ99, HTV89, Pet90]. D^{1} [JMP89]. L_{2} [Bir14]. L_{p} [GL09]. P^{k} [Dup10]. R^{2} [ATT18]. R^{d} [Bre05, CT11a, DCRT19, GLZ11, Pov99, Tan96]. R^{N} [Fra05, DeB87a, DeB88, DS88a]. Z [BSS14]. Z^{2} [BS17, CGN14, CDCIV14]. Z^{3} [Ass08]. Z^{d} [BKU10, Bur15, Der22, FK16, Mor11, XZ13]. Z_{∗}^{d} [Hig86]. Z_{n}^{d} [MS17b]. Z/mZ [NS23b]. R^{3} [Wei86b]. W_{∈} [Zha18a]. ∫↕_{∋} [Cer24]. SLE_{16/3} [Li19]. N [HL21, IW94, Mai16, Pru98, Bia91, HS95b]. N/V [GKR07]. ∇ϕ [DR24]. O(N) [LT21]. p [ABXY24, AMP05, BS24b, CP13, CDP17, CHL18, Hou90a, JT17, KS88b, KS88c, LM17, Li94b, PX88, SZ17, Sub23, Tal00b]. Φ [BC12, Yu23]. ±1 [Hig86]. ±J [Maz99]. ψ [DD17]. p → ∞ [Por86]. Q [CV16, BCF18, BW05, DM03, FM12a, IS19, IMS20, Kor22, MP02]. q ≥ 1 [BDC12, DCM16]. R [Eva00b]. R^{2} [vdBS90, CF19, CL90]. R^{3} [Oss05]. R^{d} [Bug91, GL14b, Mas88, IW94]. R^{n} [Bel86]. R^{p} [Por86]. R_{+}^{n} [JY03]. ρ [Bor20, FL19, KS00]. Sch_{τ} [HV15]. σ [Art01, BPR10, Jaw95, NUZ90, OT20]. Sine_{β} [HV15]. SL(n, R) [Lia93]. SL_{2} [CC21]. SL_{2}(C) [DKW23]. SLE_{κ} [Aru15, VR14]. SLE_{κ}(ρ) [Sch20]. ⋆ [ARV13]. T_{c} [CIvV03, Iof95]. t_{g} [Cha19a]. RCD^{∗}(K, N) [GT20]. θ ≥ 2 [FS08]. U [CK20, DDP86, GJV88, IR89, JN91, Maj06, Szu92]. U_{q}(sl_{2}) [CGRS16]. V [Hou90a, Tal16]. φ [Sam87]. W [LL24]. W_{p}^{1} [Kry10]. W_{2}^{n} [Kry94]. X_{1} [Sas86]. X_{2} [Sas86]. X_{3} [Sas86]. Z [BdT10, BdTR19]. Z/nZ [Hil94]. Z^{d} [FM06, Wan01]. Z_{+}^{p} [Joh86]. Z_{2}^{d} [Wil97]. |ρ| = 3 [Bor20].

* [HNY99]. *-probability [HNY99].

-algebras [Jaw95]. -avoiding [Bor20]. -balls [DJ94]. -boundedness [Hou90a, Hou90a]. -Brownian [Von92]. -capacities [Yos93b]. -Catalan [MP02]. -coalescent [GPW09]. -completeness [Li94b]. -convex [JP10]. -currents [KKT10]. -D [DL20]. -dependent [Pet90, HTV89]. -dimension [DGZ05]. -dimensional [Ald86, AST19, Bur89, CMT15, CM13, Kor04, LS97b, TW20]. -entropy [LL24]. -error [DJ94]. -estimates [LBE18]. -estimation [DM16b, WZ99]. -expectation [JP10]. -expectations [CHMP02]. -fields [Art01, BPR10, NUZ90, OT20]. -friendly [TY03]. -functionals [Cla86]. -functions [ERV02, KN98, RV96]. -Hahn [Kor22]. -harmonic [BY05, BKK08]. -harmonizable [RS94]. -Hodge [Li11]. -invariant [BdT10, BdTR19]. -Jacobi [Ahn22, GZ18]. -Laplacian [AB11, LM17]. -limit [GKR07]. -loss [Bir14]. -lower [Por95]. -mappings [AMP05]. -martingales [CSK99, Le99]. -means [CM88]. -measures [DK04]. -Meixner [BW05]. -mixing [Sam87]. -moment [BC12]. -nary [Kol03]. -norm [Li14a, FT04, GL09, Gri95, Kry10, Li92, Li08]. -Norms [CH88b, HJM20, Hit90]. -parameter [HS95b, IW94]. -particle [Mai16]. -path [KP89, Mik04]. -percolation [KS00]. -pinning [IV00]. -Plancherel [FM12a]. -polar [She93]. -probabilities [JZ10]. -process [CV16, CK20]. -quantum [HKK23]. -recurrence [Cha19a]. -regular [DS20, LLT^{+}19]. -regularized [BMN12]. -sat [Tal01a]. -scale [ARV13]. -self-similar [GVA86]. -sided [HL21]. -skeleton [HHT21]. -small [NN04]. -sphere [Ndu91]. -spin [ABXY24, BS24b, CP13, CDP17, CHL18, JT17, SZ17, Sub23]. -spins [Tal00b]. -stability [Yu23]. -Stable [GM06, KS88b, KS88c]. -statistics [DDP86, GJV88, HJS88, IR89, JN91, Maj06, Szu92, Tal16]. -Super [DD17]. -tacnode [BL19]. -TASEP [IS19]. -th [Deh86, Deh88]. -theory [Kim04, Kry94]. -trees [Eva00b]. -tuplewise [Pru98]. -valued [Nis97, BG86b, Ein88, Hig86, Kry10, vV89]. -variable [HS95b]. -variation [PX88]. -White [FL19]. -Whittaker [BCF18, IMS20].

129 [DeB05]. 136 [Fon09].

2d [HMO23, BRM24, CM24, DLTV24, vdBN21]. 2D-discrete [CM24]. 2f [HMT23].

3-regular [ABS21]. 3D [FL21, HLP24].

455 [Ano98a].

89 [Ano98a].

a.s. [BMP04]. abc [HHW94]. abc-gaskets [HHW94]. abelian [DM22b, Fel05, DSC96, Fel03, JR08]. above [CIvV03, CMT15, vdHS05, vdHdHS07]. Absence [Reg23, Bor03, Pap87]. absolument [Rio98]. Absolute [MM90b, OS16, Wat00, Bel86, Bor88a, KS89, LO24, Sim11, ÜZ93, OS22]. absolute-value-FKG [LO24]. absolutely [Bra89, Hen07, Imk96, ÜZ94a, Vie97, Rio98]. absorbed [BJL17, QSY24]. absorption [BBS14]. abstract [Har04]. Accelerating [DM22b]. Accuracy [BB91, CL97b]. Acknowledgement [Fon09, Xia07]. across [CD13b, DM05].

Action [KL93]. actions [Ros88, SCZ21]. Adam [CCL^{+}24]. adaptation [CT02, GL09, GS15b]. adapté [Thi91]. adapted [BVBE20, BJP22, ÜZ07, Yon97]. Adaptive [Bru16, BN13, EL94, GP05, GP00, RB03, Ruk86, BM19, CST24, Chi12, GN09, GL14b, KT04, Kle06, KSvdVvZ16, MW90, NM89, Str16]. Addendum [LS87b, Ano87a]. addition [CL19]. Additive [Eba86, ADN22, AP00, Bel08, BG02a, BEN23, BM16, Che00, Eri86, GHM23, GRS94, HS94, KSX08, KX09, KSX09, Mit86, Mor92, Oda08, Ove94, SY01, Yan09]. adjacency [BR21, Coo17]. adjacent [GZ18]. adjointness [DT00]. Admissible [Joh86]. advection [BBPS21]. affects [Ber20, Ber21a]. Affine [Def16, KRST11, DM09, Mir13, Urb90]. against [ABK08, Roh11]. ageing [BGS13a]. Aggregate [TZ01]. Aggregation [BB07b, LM09, BPP97, Luc14, Men17, NST23, RS23]. Aging [ADG01, ACM06, CW17, Gay19, MM15, BKdS18]. Ahlfors [Li14b, Li11]. Aire [Ber86b]. Airy [Gro89]. Aizenman [CV12]. Akaike [Hal90a]. Aldous [AP14]. aléatoire [KPyWmW88]. aléatoires [DD88, DLP87, Fer91, FS88]. algebra [ELP08, JP91]. Algebraic [DSS07]. algebras [Def16, Eig86, Jaw95, Mło99]. algorithm [BCT12, BCZ^{+}10, Cat98]. Algorithmic [HPR20]. algorithms [Bér05, MP87, Sta04a]. algorithms- [Sta04a]. Allen [HLR23, TW20]. allocation [Hec86]. Almost [BBPS21, BD89, BMP97, Boy19, Ers19, FHS12, GKM00, LWR92, LDS02, Mör00, Sch89a, Ber90a, Ber95a, BMS08, CG07a, CFR93, CS94, DD88, Dup10, EHM88, LP19, Oll87, RAS05, Rio00, Rol90, Sun98, Thr87, MP87]. Almost-sure [BBPS21, BMP97, CG07a, DD88, Rio00]. along [BG24, Bre08, BNS14, MR03, Sam22]. alternative [AS22, Thu86]. always [BMP04]. amenable [EM21, Rau88, Ros88]. among [ACS21, DHZ19, DFSX20, NUZ90, Pov99, SS04, vdBBdH05]. Ampère [FÜ04]. analogue [BL91]. analogues [Zai87]. analyse [Bar99]. analysis [Bar99, BS98b, BK01, BZ07, BG97b, CM11a, CT86, Che90, Cla86, CQ02, DG93a, DT10, DSS07, DJ12, FT04, Hou90b, HT24, HS87, KX09, MN93, NU91]. analytic [Arn97, Bar02, Bor88b, DW90, Zam00]. analytical [GT98b]. analyticity [CZ16]. Anchored [BKP21]. Anderson [BKdS18, CM95, CMS05, CM07, CGM10, DDD19, DL20, EdHM15, GM98, GdH99, GKM00, MPS18]. Andrei [Pen90, Pen90]. angle [Le96, LM22]. angles [Le99]. angular [Mar08]. anharmonic [BH13, OS13]. Anisotropic [BCF18, HL00, KY17, DCvEH18, KLP01, SS87a]. annals [Ano89a]. Annealed [ANV15, CM24, MO15, BBT16, CGZ00, CGZ03, CM07, Lac14]. Annealing [Cat98, Mon18, Rai09, Tro95]. annihilating [CRS18]. Announcement [Ano86a]. Anomalous [Ber01a, BKM24, FN15, GH20, IN16]. Anosov [Lov95]. ant [BCF19b]. anti [CCK15]. anti-concentration [CCK15]. anticipating [BF93, CMN92, NP88]. Anticipative [Buc91a, MN93, Pri96a, Ust88, ÜZ92]. anticipatory [AP91]. Anticoncentration [DT23]. any [EHJT13]. appearances [CP88]. Appell [GT97, GTT98, TT91]. Application [BD11, DNS90, Wan94a, Bon20, CR95, Cos92, CFG96, CH88a, DT00, DS18, DDP86, DD16, Duf92, Flo90, FRS23, Fuk91, GS90, GTS84, GTS87, Guy87, Har04, Hor00, Kry03, Lap90, LP11b, MMS14, Mat94, Rio94, Sch88a, Stu91, Vie97, CR88]. Applications [Ber89, Com89, Lig87, ÜZ93, AK16, ADS16, AC13, BC11a, Bar00b, BS22, Ber90c, BRS96, BR01, BM16, BMS08, BQR00, CK20, Con19, CS94, DK99a, DP05, DD05, DSS14, DK04, Eld20, ELY99, Fin02, FMS11, GKN92, GN09, Gor92, GPW13, HMS11, IPS13, Ino08, Ino23, JN91, JZ15, Kal07a, Kal07b, Kan90, Kry02, LM89, Le 93a, LS02, MW14, Ndu91, Nis97, Oli16, Pin07, RR92, RR00, Stu95, Wu00, Zam02, dA00, Ber89]. applied [Lov95, Rei89a]. approach [AMS94, AST19, Aug21, BL15, BK93, Bor88b, BF93, BDG21, CRZ96, CGM08, Cha12, CDR86, CL97b, CP13, Chi12, CIK99, CC94, COT19, Cra93, Div22, Dyn91b, FvdH17, FL95, GR10, GL07, GHS20, HLØ^{+}93, LRV23, MS00, MW16, MMW15, Son93, Sta04a, Sub23, Thu86, Uek89, Zam00, vdHdHS98, vdH01, vdHS02]. approche [CRZ96]. Approximate [BZ12, DS18, DM16b]. approximating [RD13]. Approximation [AFZ87, BJP22, BW93, BDG21, Epl86, Fer89, GNSS95, Jia09a, Owh03, RW05, Zha07, Aug21, BK04, BL12a, Ber86a, Bon20, BB91, CF94a, Cha16, CR95, CKK13, CL99, DS87, Ebe97, ER96, EM97a, Eld20, GM12, GT05b, HPU04, Kal03, Kal09, Kom01, Kuw09, LPS16, Mam87, MN98b, MP91, Nol14, NW91, Pfe87, Wei93, Zai87, AFZ87, BW93]. Approximations [GS89, Ass01, Bar90, Bes91, Bol86, Bol87, BM05b, BCH81, BCH88, Cek98, CK20, Che95, Deh88, DK87, EM13, Ein88, FSX19a, FSX19b, Fon09, GK96b, JW24, KR14, Kry00, KL00, LS89, Mél01, Ver03, GS89]. Arbitrage [DS95]. arbitrary [Blu05, Uch94]. arbre [Has88]. arc [CPY94]. arcsine [Ber95b]. Area [vdBB04, Ber86b, BH98, CR88, Dri13, FV08, GVV93, IVW18, LQ97, Maj92, Xu91]. area-like [Dri13]. areas [BW17a, DH96]. arises [Rol03]. arising [BDS93, Col05, EP19, HLN13, LS87a, MT95, War05]. arithmetic [HNY08, vdB00]. arm [DM23]. armed [DS11, Man86]. arrangement [Nes19]. arrays [Kal95, Pel12, Sam87, Str86a]. Arrow [Mos12]. article [Fon09]. ASEP [BN22, Yan22]. aspect [Kaz88]. aspects [BCS11, DL05]. assembly [DL23]. assignment [Ald92, LW04b]. associate [XZ13]. associated [AD08, AR89, AB11, Bar21, BC11b, BC11c, BL03, Bob10, BP99, CT86, Day88, DO05, DM09, FS10, HKK23, Léa87a, LU03, Léo95, LM00, Leu10, LQ19, RW05, Shi94a, Shi94b, Tan97, Tzv10, Wat87, Yat13, Yu93]. Association [Eva90]. associé [Leu10]. associée [Léa87a]. asymmetric [Ale01, ABL88, Ass05, BMRS20, CGRS16, CLO01, Dit90, Fer92, FF94, LY97, MM15, MS02, Pin07, Wee90]. Asymptotic [Alt89, AG23, Bab91, BBM00, BM08, Bar86, Bax86, Ber86a, Bry08, Cas93, CG13, CC94, CMS00, CMS03, DR06, DR07, DR13, Don94, DK87, DR92, Dzi91, Fil89, FT04, GW04, GHK22, GH89, GN98, GK01, HMS11, Höp88, Hös96, Kee88, Ker89, KS00, Klo13, Lah99, MR13, MST14, Ohl89, Par91, Roz86, Sha92a, Wen96, Wen99, Wen04, Wer97, Yos92, Yos93a, ABF23, CF03, Don89, DM16b, EK87, GT98b, Hai02, HJL90, HR95, IW94, JN91, Jaw95, KKR88, Kho92, LLR10, LS86, Lus92b, Mam92, Myk95, PR08, Tao13a, Wei88, WZ13, Yos97, Zes88, DR16, GHK23, LY86, LY87, Wei86b]. Asymptotical [Tro95, GS90]. Asymptotically [HI86, HHW94, LT00, Bal98, Cla87, CHM86, SCC19, Wac12]. Asymptotics [AS92, Ald92, EL87, FM12a, HM24, KM93, LBE18, OV19, Pet14, WHZY07, Yao19, Yuk95, vdBB94, CY14, CY15, CF19, CCM16, Dur02, FM90, GKM00, GM05b, Ham00, HvdH11, HvdH19, KSW16, MMF92, Mer03, MN04, NN04, Nut12, Szn93a]. Asymptotique [Wei86b, LY86, LY87]. Atiyah [Léa88]. Atomic [BPR10, BC04, Hil87b, Wei93]. atoms [Var07]. attachment [CCL^{+}24]. attack [LPW13]. attracted [AT86, BD89, Chr86, HS88, KM88, LPR97]. attracting [HH19, Rai97a]. attraction [Ber19, Ber95a, Cli86, Ein88, GH91, Wei86a]. attractive [Mou92a, Uel02]. Attractors [CF94b, BCF93, CC99, Tea08]. augmented [BBW14]. autocorrelation [Ino08]. autocovariance [HP94]. automata [DvdW01, HMM19]. automatic [NM89]. automorphisms [LS87a]. autonomous [Mac21]. autoregression [MN98b, MSW09]. Autoregressive [Pou88, DJ00]. autour [LY87, Wer94b]. aux [Ber89, Leu10]. avalanche [JRS20]. Average [HT24, BDN17, GN22, HH90, HS90b, Mor91b]. Average-case [HT24]. Averaged [Gui97, BCK21, Yil11]. averages [CH93, PT02, SRMC93]. Averaging [CF09, EF93, Wat88, BK04, DGPZ02, YY21]. aversion [Nut12]. avoiding [Bor20, HH19, HHK90, HH91, HK92, HRS17, HTV12, Kön93, NRW87, Uel02, YK88, vdHdHS98, vdH01, vdHS05]. ax [BV22]. Aztec [ACJvM15, BD23].

B [CS94, BV22]. B-valued [CS94]. back [PS22]. backtracking [FvdH17]. Backward [Bla05, DHB11, PP94, RZ16, Roz03, BP99, BQR00, DT10, DT12, HP95, HMY02, MPY94, MY99, MZ02, MZ11, PT99, RV93, STZ12, Yon97, Yon08, Zha10]. Bahadur [CK08b]. balanced [BD14, DG22, GZ12b, Hal90c]. ball [CQ23, FJK23, GPS22, Luc14, NN04, SW95]. ballistic [Bau16, CR14, Szn02, vdH01]. Ballisticity [DR11, GVV22]. balloons [ARS23]. balls [Dar92, DJ94, Dwy90, KLL94b, LT88, Tan96, Tan97]. Banach [BR90, BG93, Ber90a, Cer99, CM88, MWZ05, MWZ08, MT87, MN87, Shi94a, Thu86]. Banach-valued [MT87, Shi94a]. band [AZ06, BE17a, BW18, FM21, YY21]. band-limited [BW18]. bandits [Man86]. bands [HR02, KNP12, Lah99]. bandwidth [HM91, Hös96, Rob94]. bandwidths [WT04]. barely [vdHHN20]. Barlow [Mat95]. barrier [GOZ21]. Barriers [BGR15, CLM^{+}12, For09, HH05]. barycenters [ACLP20, CDM24, LL17]. based [Bar86, FIKP13, HR02, Hil87b, Hsi93, IPS13, IKS13, Imk94, Lah99, MMF92, PT17, RS17, Tal16, WZ99]. basic [Kal00a, SV86a]. basis [Kla09]. Bass [LST20]. Baxter [BHSY23, CP93]. Bayes [Bar24, CKP14, KSvdVvZ16, Yos93a]. Bayes-like [Bar24]. Bayesian [BM19, Chi12, MST14]. BBP [BW22]. be [CL17, CDR86]. bead [FFN12, NV21]. beads [Vir03]. beam [BMS05]. beaucoup [Wer94b]. behavior [AB15, AM95, BCL20, BY86, BG99a, BBC^{+}04, CR14, CP10, CS08b, CS09, CV98, CG90, DHR21, DK87, DR92, EG95, EK87, Erd01, FJL09, FZ99, Hsi93, KS00, Lal09, Leb17, NN04, QR13, Sav09, Sch86, Szn02, Tri95, Wat96]. behaviour [And08, Bax86, Ber86a, Ber01a, Ber95b, CMS00, Dit90, Don89, Don12, DR13, DR16, Dzi91, Gri88b, GM05c, Har90, Hry98, HV04, HS07, IW94, KKR88, LWR92, Maj88a, Mou90, Sta02, Tro95, Wer97, Zes88, vdH01]. being [Mou92c]. Belavkin [AGM03]. Bellmans [Kry00]. below [CM01, Hut21]. Benamou [GT20]. Benedicks [CMS03]. Benford [Sch88a]. Benign [CST24]. Benjamin [Tzv10]. Bergman [BFQ18]. Bernoulli [AGKW09, CG09, DFSX20, Dup10, HTV89, UA03]. Bernstein [Cas12, FNS97, MPR11, Zai87, vdGL13]. Berry [BCG14a, Bob18, CS01, DT23, DH23, Kla09, Neu24, Rio96, Tan86]. Berry-Esseen [Rio96]. Bessel [AZ20, DS95, DMRYZ04, HM24, Lee90, MR07a, RW19]. Best [FNS87, CM88]. Best-possible [FNS87]. Beta [BC17, LL04, NV21, Kor22]. Beta-distributed [BC17]. Bethe [DMR11]. between [Bel86, BD20, Cha10a, CLM94, DM22a, Dok04, EHM88, FR89, FI19, FI07, GIM22, GIM23, GRS24, GH91, Häg95, Hos91a, LM22, MS87, Mas88, Mic99, Mue91a, NT87, NTV17, Szn89, dHHKR91]. Beurling [Li14b, Li11]. beyond [DSS07, MWW09]. bi [HHW18]. bi-free [HHW18]. bias [AB15, Hös96, Zho88]. Biased [BD23, CFK13, LPP96, Tak98, Aïd14, AdR17, BGP03, Bow18, FHS12, GG11, Hec86, PZ08, PPY92]. Bienaymé [ADN22, Kor17]. bifurcation [Bax94, BEN23]. bifurcations [BG02a]. bilinear [Avr88, KN98, Zha20]. binary [AS88, ACCH23, Gan94]. binomial [GIM22, GIM23]. biometric [BCH81, BCH88]. bipartite [TY23]. Birth [Dyn93a, FKL04, CCM16, CMMM11, DLP10, FM86, Joh86]. birth-and-death [CCM16, DLP10]. birth-death [Joh86]. Bismut [FSX19a, FSX19b]. bispectrum [Mar08]. bivariate [GT97, GTT98, Kal88, TT91]. BK [vdBJ12, vdBG13]. BK-type [vdBG13]. Black [Rév92, Gei05]. block [BE17a, BCZ^{+}10, DMB20, Lah01, NO15]. blow [FL21, Rom14]. blow-up [FL21, Rom14]. Blowup [MS93, FKN24]. BM [Son04]. BMO [Gei05, Jia09b, Kaz88]. BMO-martingales [Kaz88]. bodies [Bob10, CY14, Kla09]. body [BR97]. Bolthausen [BL00]. Boltzmann [BF11, DR14b, Léo95, Uch89]. Boltzmann-equation [DR14b]. Bond [GM14, Der22]. Bonnet [Nic16]. Boolean [BGS13b, Cha16, FM19b]. Bootstrap [BB06, Mam92, NSU19, BBS16, BB91, CL97b, CK20, Düm93, DCvEH18, DCH24, FR89, GHM12, GH19b, HM88b, Hal89, Hal90c, Hol03, JW24, LS86, Mou92c, Lo89]. Bootstrapping [FR89, GH23]. borderline [Erd01]. Borel [SS87b]. Borell [vH18]. born [HP99]. Bottlenecks [FR07]. bottom [Yur99]. Bouchaud [CM17, ACM06]. boules [LT88]. bound [AB15, CS01, KN03, MR07a, Szn11, Zho88]. boundaries [CN16, Cra93, DFP15, DM05, DG21b, EK87, GH19a, GPS22, Lia93]. Boundary [AS17, BCR04, SZ10, AS11, AP00, Bai08, BE17b, BLQ23, CJN21, CWZ95, CK04, Gal22, GOZ21, GM89, HLØ^{+}93, IR10, Kan00, KMM98, LRV23, Las19, MP19, NN04, Nis03, OP89, Pal12, Pap90, PZ98, RW88, RW89, Ric18, Sai87, Sch20, UA03, ZD90, vdBS90, vdB94]. Bounded [AB15, AW15a, AF12, Ber87b, BBW14, BM01, BH98, DDP86, Ein92, Fuk11, Kal00b, Kön94, Las19, Pow19, Pru98, Tao13b, Tao13c, Tru03]. boundedness [Hou90a]. Bounds [CMS15, PT04, AGKW09, AG14, BBM99, BCG14a, Bob18, BCT12, Bre08, CCK15, CSC93, CS12, DM23, DT23, FNS87, FMS11, GT16, Gri87a, GS15b, HM91, Hit90, Hut21, Iof95, Jir16, Kan90, KN98, KH03, Kor17, Kül03, Lou97, MP05, NTV17, OTZ24, Por95, SSSY17, Son04, UA03, Wan01]. bouts [Has88]. box [DCTT18]. box-crossing [DCTT18]. braided [FNS97]. Branching [ABBS13, AS05, AS09, DKS94, GdH92, Lig96, Rei89a, ABB11, AHJ21, ABK13, BBS14, BFF18, BL00, CM11b, CR88, CR90, CH19, CGS23, DF91, DH91, DvdW01, DK04, FV00, GHK22, GHK23, GRW92, Gre91, GK12, HL00, Hut20a, Isc86, KM12a, Kyp00, LS97a, LZ10, LS15a, Lam02, LPW13, Lee90, Liu96, Mai16, Myt98, NZ17, NZ21, NW90, Per92, Pow19, Rei89b, Sal90, Wan97c, Win91, vdHHN20, HS21]. Branching-coalescing [AS05, AS09]. breaking [Tal00a]. Brenier [GT20]. Breuer [NN20]. bridge [BJ99, Gol96, Léa01, PR12]. bridges [AK08, CV18, Con19, Con24, DW21, Hsu90, Sam22]. Broadwell [Uch89]. broken [BH05]. Brown [DHK22, HH23]. Brownian [DeB88, RW89, AL94, AD02b, AD04, AE14, AK08, AH20, AL23, Ahn23, ABBS13, AD22, ABK13, AC03, BR92, Ban16, BB24, BS97a, BPP04, BP88, BB90, BB95, BB96, BH00, BH07, BCR04, BBS14, Ber89, Ber90c, Ber00, BM17b, BS05, BS98b, BM05a, BG24, Bor91, Bor86a, Bor86b, BJ99, BM11b, Bud18, Bur89, BL90, BL01, BN02, BM05b, BS14b, BNS14, CPY94, CRZ96, CP10, CHM22, CD99, CR88, Che92b, CGS23, CMS00, CDK18, CQ02, CKM93, CCFS00, CL16, DW93a, DW93b, DM97, Dav99, DeB87a, DeB87b, DT93, DeB12, Def16, Del99, DF01, DL97, DM91a, DM03, Dri13, DHK22, DD17, DR92, DP99, DvZ04, Eis90, EF98, Eva00b, EMS13, FT04, FW20, FL95, FM97, FR92, FP00, FL96a, GHM23, Gan94]. Brownian [Gol96, GdH14, GS15a, GVV93, Gro89, Gru98, HH23, Ham92, HS13, Has89, HRS17, HTV12, Hsu90, HS95c, HL19, Jeg21, KP89, Kho92, Kni01, KMP22, Kre95, Kry03, Kum93, Kwo92, Kwo95, Lac90, LS97a, LW04a, Le96, LY86, LY87, Le 87, LM92, Le 93b, Le 95a, Le 15, Le 18, Le 95b, LQ13, LS87b, Léa01, Leu10, LS01b, Lio05, Mai16, Mal95, Mat95, McG86, McG89, MW95, MS24a, MS21, MN02, Mör00, Mör01, MS09, Mou90, MS24b, MP19, NRW87, OB04, Osa98, Ove93, Ove94, PPS96, PS16, PT87, PW97, Pes02, PT00, Pit96, PR12, Pov99, PR08, RS12, RW88, Rog87, RW91, RS93, SV99, Sal90, SS18, Sch96, Ser95, Shi92, SV03, Sto86, Swa07, Szn93a, Szn93b, Szn96a, Szn96b, Tal98a, Tan96, Tan97, TW93, Tha96, TTV03, Tri91]. Brownian [Tro21, Val91, Ver97, Vir03, Von92, Von95, Wan97c, Wan03, Wer94a, Wer94b, Wil87, Wüt98, XZ02, Xu91, vdBS90, vdB94, vdBBdH05]. brownien [AL94, Gol96, LY87, Le 87, LM92, Leu10, Wer94b, Ber89, Mal95]. brownienne [Wer94a]. Browniennes [CRZ96]. Browniens [Mat95, Ber90c, DM91a, Eis90, LY86, Val91]. bruit [Bié98]. Brunn [Bor08]. Brusselator [Sch86]. BSDE [FMS11, HT10, Pen99]. BSDEs [BH08, HH05, PZ98]. bubble [DW93b]. Bulk [Agg19, EYY12]. bumping [MMŚ21]. bundle [Léa01]. Burdzy [CM91, LST20]. Burgers [HW13a, AMS94, GP20, HW13b, Kif97, Szn86]. Burgers-like [HW13a, HW13b]. Burkholder [CSK99]. Busemann [CP12, GRAS17b]. BV [Tre15].

C. [Kry97]. cônes [Le 87, LM92]. cactus [Le 15]. Càdlàg [ERV02]. Cahn [HLR23, TW20]. Calcul [Thi91]. Calculus [Ust88, AJJ22, BS98b, DNS90, EM90, ERV02, Flo90, Imk96, KY00, NZ86, NP88, PP87, PS87, Pri96b, Sch23, Thi91, TZ97, Yos97, Zäh98]. calibration [Roh11]. calls [LMU03]. Cam [PV20]. Can [EHJT13, BP89, CDR86, LT15]. Cantelli [HJS88, Yu93, vH13]. Cantor [DG88]. capacities [HS95a, Li03, Nor86, Stu95, Yos93b]. Capacity [GS86, Iha90, ASS19, DPR22, Haw98, PPS96, vdBB04]. capacity-equivalent [PPS96]. Capture [LS01b]. CAR [ELP08]. caracterisation [DS88a]. Carleson [Jia09b]. Carlo [DJP01, EM13, Mél01]. carpet [BB90, Shi03]. carpets [CQ24, Osa90, Osa01]. carré [Asp89]. cas [HS90a, LM92]. cascades [BPW09, BJ14, LS97b]. Case [Rai97a, AA88, Arb91, BRR11, BB96, Bax86, Boi90, BYY14b, BdT10, CGM08, CV98, DR07, Don97, Erd01, GT98a, Gho15, Häg03, HS90a, HT24, Kol13, LM92, MS90, MP11, Oda08, Qia18, Tal01a, Wu04, Yin14, Zha20]. cases [BM16, Por96, Sch05]. Catalan [MP02]. catalysts [DS94]. catalytic [DF01, EF98, FL95, FM97, GKW01, GK12, KM12a, Myt98]. Cauchy [AW15b, Ber19, DS88a, Has88, Roz03]. Cauchy-compliant [DS88a]. Cauchy-conforme [DS88a]. Cayley [HP99]. cellular [DvdW01, HMM19, IN16]. Censored [BBC03, CK02a, CKS10, LMW89, TA96, WHZY07]. censorship [Bab91, BCH81, BCH88]. center [Box89]. Centered [Ale02]. Central [Ale87a, BLM11, CS89, CD13b, CH88b, FT87, GR10, GT97, Gou04, GLZ11, HLLS23, HLW24, MN04, Mor87, Pel12, Rei05, Wit87a, AB10, AGOZ88, Ans99, ABBN04, Bar86, BSC02, BV95, Ber95a, BHS10, BS05, BCG14a, Bob18, BMP97, Bon20, Boy19, Bra89, CES24, CFR93, DS18, Ded98, Ded10, Ded13, DG98, DL01, DL03b, DG22, Dun24, Fra05, Fuk91, Fuk11, GN08, GS90, GTT98, GVV89, Gol07, Gra05, Häg05, HW86, HHT21, HL19, JB04, KZ97, Kol10, KL87a, LT88, Mar08, Nis97, PPZ14, PZ08, Pet92, Por86, Sch88b, Shi06, Tho14, Wat87, WY93, dJ87, dHHKR91, vW97, BCG14b, Fuk94, LT88, Pet92, WY93]. certain [CT86, CF03, DeB87b, FKL04, Fer94, Gol96, HNY99, Kol87, Le96, LY86, Man91, Mar87, MU12, Tal09]. certaines [Fer94, Gol96]. certains [LY86]. Cesaro [DD88]. chaînes [CSC93, Mic99]. Chafee [BEN23]. chain [Alh04, BH13, BV22, CGH24, Dar88, EM13, LM00, OTZ24, OS13, RR03, TY23]. Chains [Ber87a, ABBP02, BP89, BKU10, BC04, Bla88, BCS19, CD20, CD21, Che00, CSC93, DDMP19, DL01, DL03b, DLP10, DH23, EP97, Ful09, GL14a, GT98a, GLL20, Gre87a, Gre87b, Häg05, Höp88, Joh86, KM00, KM09, KM12b, Lev88, Lov95, Mic99, MW14, MR16, NOT22, NN93, Pel12, Ros89, Tho88]. chaleur [AL91a, AL91b, FL93]. challenge [Tal98c]. chambers [BJ02, Def16, Nes19]. champ [Pra93]. champs [Guy87]. change [Hös91b, Kal90, Mit86, PS93, Thi91, ZD90]. change-of-variables [Thi91]. change-of-variance [Hös91b]. changement [Thi91]. channels [Iha90, Liu22a]. Chaos [DK99a, GM94, AN92, AJJ22, BLM24, BWW18, Bor88a, CF24, CP13, Der03, Gra95, Gui97, Imk92, Imk94, Jeg21, KLZ23, LL23, LP11a, MMW15, NT86, NT87, NP09, Szn86, Won20]. Chaotic [FIN99, AB07, Che95, CJKS13]. chaotic-representation [AB07]. Character [GM05b, Bor86a, Bor86b, CJKS13]. characterisation [BPR20, GOZ21, McG86]. characteristic [ABZ23, BCGZ22, Mit87, Shc13]. characteristics [FIKP13]. Characterization [GTS84, GTS87, DS88a, Fag90a, Fel03, Fel05, HP18, KM12a, Kry02, Lac16, Led00, Liu17, RS91, RT02, Von92]. Characterizations [BLS87, HNY99, Klü89, Lin87]. characterize [Tay89]. characterized [Xia04, Xia07]. Characterizing [Sas86]. charge [RSX13]. charged [BCFN89]. charges [FL96a, Mat92]. chart [Ndu91]. chemical [Dit88, Kot88]. Chen [Léa01]. Chern [Nic16]. Chernoff [HS87]. Chi [SCC19]. Chi-square [SCC19]. Choquet [Bia92, Rau88]. chordal [RZ18]. chosen [BR97]. Chung [BM11b, EV21]. circle [Cha92, CC94, LR04]. Circular [BCC12, Rid04, BYY14a, BYY14b, GT03, GT16, Yin14]. claims [Kra96]. class [AS88, AB07, BE17a, BL07, BLM11, BT03, Bau16, Ber86a, BGS16, BCW21, Bru13, Buc21, BMS08, BDG10, Cer05, CF09, CMN92, CV98, DP98, DNS90, Dyn91b, FZ05, FG19, GRS94, Hös96, HV04, HLN13, Isc86, Jur88, Kot92, Kot95, KV18, Lal86, Le 93a, Led00, LS87a, LWXZ12, MN98a, MiS09, Maj06, Man91, Mar98, MP87, Mou92b, Por95, Rez96, Wan97c, Wat96, Wil88, WZ93, Zhe95]. classe [MP87]. classes [BvdHS18, BS24a, Gri91b, Jur88, Klü89, MiS09, Men17, NUZ90, Ort89, Rio01, Tal88b]. Classical [AR89, CG07a, DZ95, Erd01, Fit87, FR05, Gri88b]. Classification [Uma06, BG07a, FKL04, She93, Wan97c]. CLE [MSW21]. close [CC94, HMZ20]. closed [Tri91, Yat13]. closeness [Zai88]. closure [GG96, MiS09]. CLT [AZ06, ABZ23, BG93, BG97a, CMS07, CMS23, CG13, DKP22, Gol08, GJMM24, Mou14, Pru98, Sta04b, Zha18a]. Cluster [DPR22, Ale01, BDC12, BM03, BS17, BGJ96, DG93b, DCM16, DCMT21, GKZ93, Häg96, Kes86, Lac14, LS24, MW14, DM16a, Tei11, vdBG13]. cluster-formation [DG93b]. Clustering [EG95, Swa00, EM93, FG94, GK91]. clusters [BGP03, BB07a, CS12, DS11, HP99, HS11a, Hut24, PS98, PRS16, Sch99, SS04]. Co [BBH24]. Co-evolving [BBH24]. Coagulation [LS20, Kol10]. Coagulation-transport [LS20]. Coalescence [BC02]. coalescent [AP00, BL00, BL03, BBW14, BvdHS18, BM16, DK17, GPW09, PWW11]. coalescents [ABL12, LS20]. Coalescing [EMS13, AS05, AS09, CK19]. cocycle [Jou87]. cocycles [AS95, GRAS17b, Jou87]. coding [Dup10]. coefficients [ALN99, BP87, Bou88, Cer24, Cha24, CG07b, DP05, Fag90b, FZ05, Flo90, GM12, GM16, Hau07, IKS13, Kol11, Kry00, Mit87, Mou14, MP11, Wat88]. Coexistence [DS94, BK07, CP07b, Erd01]. cohomology [IL01, Léa01]. coin [HK97, Hec86]. coincidence [Lac14]. Collapse [Tea08]. collision [PR08]. collisions [DGL87, NT87, Szn89]. colonies [KM12a]. colorings [MS10]. coloured [NT87]. columnar [BSV10]. combinatorial [GT24, Sch88b]. coming [LP19]. Community [GV16]. Commutation [Ful09]. commutative [Ans13, AG97, GVV89, Par22, Sch05]. Compact [Art01, BM17b, Dun07, AUY08, BSC02, Bia92, BLV12, BGV07, Fel05, HLW24, Jaw95, KNP12, LR21, Li03, MP92, Rai97b]. Compactness [BR23, Hen07]. compared [Asp89, Cat98]. Comparison [BBT16, CCK15, CFG96, Hit88, JR88, Kot92, Lus92a, QRZ03, Ale01, LS02, Tro12]. compensation [SK86]. compensator [McG86]. competition [BFF18, DM91b, GRAS17a]. competitive [GLPP17]. complement [GHM20, Wer94a]. complémentaire [Wer94a]. complementary [MPV01]. Complete [DM91b, Fab90, BGJ96, Bon90, DF16, Led00, Li08, Li11, Li14a, Li14b]. Completely [DW90, FQ88]. Completeness [Mat92, Gho15, Li94b, MS95]. completion [KLT17]. Complex [ADP16, KK14, Por96, Bry08, FI19, LY86, Lio05, MM17, Mou90, Tho01]. complexes [BKP21, LY86, YSA17]. complexities [RST15]. complexity [RS17]. compliant [DS88a]. component [BM11b, GKN92, HP19, Jan95, JPRR18]. components [AKLP23, BKYY16, CF03, Häg03, Mor03, Tri91, Wer94a]. composantes [Wer94a]. composition [ÜZ94a]. compositions [KM22]. Compound [CR95, Coc18, DIT14, HV09, KR14]. compressible [BFHM19]. compression [YS92]. Computable [Key87]. computation [Sub23]. computational [EM21]. concave [ASZ09, BGP24, Har04]. Concentration [AW15a, AS22, BGM18, BG07b, BGV07, BFO09, Bru19, CCKR07, GG11, KNP12, MM13a, Pal12, AGKW09, BKdS18, BL97, Cha07, CCK15, DGI00, FHS90, Hua24, Kar12, MMS14, Mar98, MR07a, Men18, Mil12, NST24, NO15, RB03, Rio01, Rio01]. concentration-variance [FHS90]. concerning [Fon09, Xia07]. concrete [CLW11]. condensed [BL12b]. condensing [AGL17, BJL17]. condition [AK08, Don95, EKZ22, Léo22, Mas92, Mij88a, RST^{+}13]. Conditional [BW05, CS03, Art01, BJT11, Bry88, BF93, Com89, Der03, Kry16, LPR97, MP91, NU91, Röm96, Zie88, vW97]. Conditioned [BB95, GLP17, GLL20, Ove93, ADN22, Car05, CN20, CFK13, DeB87b, Def16, DFSX20, GVV93, KP89, Kor17, SV99, VW09, Xu91]. conditioning [DZ96, GQZ88]. conditions [AD04, Ale87a, AGOZ88, AP91, BL21, BG93, BH08, CJN21, CL17, Cha24, CWZ95, DR11, GVV22, KR16b, LLR10, LS21, Nis03, OP89, Ros89, Sch98, Stu93, ÜZ07, Wit87b, Xia97, Zai87]. conductance [ABDH13, ADS16, BBT16, BS22, BCW21, DNS18]. conductances [BC11b, DHZ19, FJL09, FM06, SS04, XZ13, BC11c]. conducted [Bié98]. conductivity [Owh03]. conduite [Bié98]. Cone [Ver97, AD13, DeB12, Kwo92, Kwo95]. Cones [KMTT19, BS97a, DeB87a, DeB88, Le 87, MW95]. confidence [AM88, BN13, Hal89, HR02, KNP12, NSU19]. configuration [CQ20]. configurations [COP93]. Confinement [Pov99, Szn96a, Szn96b]. Conformal [Li19, Qia18, GHM20, KW16, MWW15, MS21, MSW21, vdBCL16]. conforme [DS88a]. Conga [Ban16]. Conic [LM92]. conjecture [DMR11, DR14b, Le 88, LW04b]. conjectures [Yu23]. conjugacy [BS19]. conjugacy-invariant [BS19]. connected [Ber00, Der22, Izy17, Tri91, Wer94a, Zha04]. Connection [FPW24, Arn97, Ber95b, Cha10a, DCRT19, Mue91a]. connections [Ber87a, BGS16, BBCK18, CG09, DF16, MS87]. connective [Lac14, SSSY17]. connectivities [CCC91]. Connectivity [CCD88, MPV01]. connexes [Wer94a]. consequences [BPR10, CPY94, Li94a]. conservation [CD13a, DV15, FW04, FW07, KR16b, Kot95, Mar10, Oua22]. conservative [CR93, LO96]. consistency [LL17, Lie89, NM89]. Consistent [AS91, CHMP02, HKR18, WZ99]. Consolidation [BCG86]. constant [BP87, FS17, GN22, KY92, Kwo92, MRT23, Rai97a, SSSY17, Wan97b, Web89]. constants [ESZ09, GT98b, KT04, Lac14, Osa90]. constrained [BEES18, BHR01, CMRT08, CMRT15, DK87, HMT23]. constraint [Gam04]. constraints [FK97]. Construction [Bal98, Chi01, DL03a, FT96, Sal86a, Shi94b, Tal10, Tzv10, Yos96, AR89, Arb91, BLM24, KM12a, Ove94, Sto86]. constructions [Ber03]. contact [AS10, Cha16, CDS91, DS88b, DP99, FS08, GS89, LS01a, Lig96, MT94, MT95, Pen96, SV17, SV86a, Sta01, Swa18, Swi90, vdHS05]. Contacts [Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano90n, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano91k, Ano91l, Ano91m, Ano91n, Ano92a, Ano92b, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j]. Contacts [Ano92k, Ano92l, Ano92m, Ano92c, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano94k, Ano94l, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95l, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k]. Contacts [Ano99l, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e]. Contacts [Ano08f, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f]. content [Zha19, vdB94]. contestants [HL18]. contingent [Kra96]. continuation [Yon97]. continuité [Bar99]. Continuity [AK16, Rog87, Bar99, BGN91, Bel86, BK01, Bor88a, Ebe97, HLN13, KS89, LWXZ12, MM90b, MS92, OS16, OS22, Sim11, ÜZ93, VR14, Wat00, Zha19, MW89, RR00]. Continuous [DV02, Eri90, AL00, Arn97, ADP16, BGM18, BL00, Bru13, DS14, DL20, Epl86, GKM00, GM05a, GHV90, GS07, GKR07, Hen07, Höp88, Imk96, JS86, Jen91, Kol11, Lam02, LS87a, Mor87, MP11, NZ90, Pap87, Tho94, ÜZ94a]. continuous-state [BL00, Lam02]. continuous-time [ADP16, GM05a, GS07, Höp88]. Continuum [BSW17, ABBG12, ATT18, AP00, AMP04, AMP05, ALS20, BvdHS18, BR23, CSZ16, CGH24, Cro08, Kre95, Pen96, PV97, SS13, Tro21]. contour [SS13]. contoured [Gor87]. contours [KSX08, KSX09, MS99]. contraction [DLM03, DFM24, Ebe16, Sug94]. contractivity [LW00]. control [Bal98, FL21, Zha20]. controls [Bal98]. Convection [LOY98, Kom01]. Convection-diffusion [LOY98, Kom01]. conventional [KR14]. converge [DP99]. Convergence [AKTW18, ACLP20, And86, BC11b, BKU10, Ber87b, BCFN89, BDS93, BCGZ22, BGS13a, BDG10, DZ95, FRS23, GPW09, HW86, HH90, Hoo91, IPS13, JMP89, KMP22, MRT19, MRT20, Pic89, TT91, AMP05, Art01, ABBN04, Ass01, BMRS20, Bel98, BG93, BG97a, BK93, Bor20, Bor86a, Bor86b, Bou90, BNPS18, BNS14, CRZ13, CN07, CGG07, CT18, CV16, CP93, CH88a, DS87, DD88, DR14b, Dun24, DM91b, Ein92, FHS12, FW89, FGM09, FG15, Gam14, GKS04, GT03, GT16, Gri87b, GZ12a, HM88b, Hal90b, Hal91, HJS88, IR89, Jac87, JS86, Jan88, JM90, KM12a, Kol08, Kry00, KS18, KU97, Kyp00, LR86, LP11b, Li19, MM13a, MP87, Mij88b, Mir11, NW22, Ogu01, Rob94, SK86, SK87, Sch89b, Sta04a, Swa07, Tei88, Tha91, Tho01, Thr87, Wan03]. convergence [Yu93, Zha18a, dBM13, BC11c, BM05b, JMP89]. Convergent [Cla87, KM09, Sch89a]. converging [Dub02, KM00]. Convex [Bob10, CFV23, DR99a, DR99b, BF88, Bor91, Bor88b, BH98, BH08, Bru16, BR97, CG94, CY14, CQ23, CGM08, CLL20, CL90, Dob88, Dwy90, GGLS20, Gol96, Gro88, GS15b, Har04, HLO23, JP10, KMTT19, KM22, Kho92, Kla09, LR88, MMS14, PR12, Zam02, dA00]. convex/log [Har04]. convex/log-concave [Har04]. convexe [Gol96]. Convexity [FMS11, ACLP20, GRST14, RT86]. Convolution [Cli86, Wat08, Bar89, BL91, Bel08, BG07a, BG09, BSC02, FNS97, KS89]. convolutions [BM08, Bon88, Bon90, BH09, MSS22, Zai87, Zai88]. cookie [BS08, KMP22]. cooperative [ABBL24]. copies [CLM94]. copolymer [BGL19]. copolymers [GT05a]. core [HMM19, Mas92, Osa98, Pap87]. corner [GRAS17a, GRAS17b]. corners [Cos92, GZ18]. Correction [Ano87a, BRZ20a, Ber21a, BGdH23, CMS23, CD21, CS23, DeB05, FSX19a, FW07, GIM23, GTS87, GHK23, HvdH19, Kim23, MRT20, OS22, Osa22, RW89, BCH88, LB95]. corrections [Ber01a, CFV23, DCvEH18]. correctors [Gu17]. correlated [CFV23, DZ18, HL18, LP19, Mar87, MP02, PRS16]. Correlation [GdH99, Cer24, CS09, DNS22, DSS23, Har90, Har04, LLR10, MRT23, NST24, Shc13]. Correlations [FR04, AEK19, CW93, DM17, Eld20, Gal22, GKdV89, HK20, HP91, IV00, LT21, OV19]. correspondence [Kal00a, MMŚ21]. corresponding [EF93, FW12]. cost [Con19, GT20, GL07, HPU04, Mik04, Mor91b]. Coulomb [Cha19b, GS20, Leb17]. count [Slu91]. countable [Fic91, VW24]. countably [BL12a, LW90]. counter [Eig86]. counter-example [Eig86]. counterexample [Pru98]. counting [DS18, Ham00, RST^{+}13]. counts [GW04]. coupled [BK04, Bar02, Gre91, Yos99]. couples [Kwa24]. coupling [AB15, CCKR07, DW21, Hai02, HMS11, Man91, Tho94, Wan94a]. Couplings [CV18, Gre87a, Gre87b, BK17a, BJP22, BBC07, BC02, Ebe16, FW24, GG11, IK13, KS18]. courbe [Wer94a]. covariance [Au22, BKYY16, Bru13, CRZ13, CMW15, CT18, LP11a, LP11c, LV12, NSU19, Péc09, Shc13, YBK88]. covariances [LP11b]. covariant [AS11]. covariate [MU90]. cover [AB22, Bel13, BK17b, BRZ20a, BRZ20b, BGGM13, CQ20, Liu22a]. coverage [Pen23, PSSS13]. covers [DP15]. Cramér [BV95]. Cranston [Bañ87]. critère [LT88]. criteria [CR14, CR93, Kou03]. criterion [BD88, Den00, Hal90a, LT88, Shi90, Szn02, WZ99]. Critical [BBS14, CN07, CS08b, CS09, DF91, HS11a, Izy17, Jeg21, KW04, Rei89a, Wan94b, vdHS05, ABBG12, AV24, BW22, BGN91, BB96, BDC12, BM03, BNP11, BGL19, BBW14, BSW17, BvdHS18, BBC^{+}04, BdT10, CHS22, CDS91, CFK13, DCTT18, DCMT21, FV00, FR95, Gal22, GS24, GL99, Gho15, GRW92, Har90, HMT20, HV04, Hut21, KZ97, KS00, Kol13, Lal09, LZ10, LS15a, LM92, Lee90, LLP10, Mou92a, Naj18, NZ21, SZ17, Szn11, Tal01b, Wan01, YK88, Yao19, Zha20, vdBBdH05, vdBN21, vdHdHS07]. criticality [BBC^{+}04, CMRT15, CM11a, For21, GM14]. critically [Kaj13]. critique [GS89, LM92]. croissance [Fou98]. cross [HM87, HMP92]. cross-validation [HM87, HMP92]. Crossing [CFS01, Dub06, DCTT18, Slu91, Wüt98]. crushed [vdBB04]. crystallization [SS15]. cube [ELP08, EI20, FM19b, Mas88, PPZ14]. cubes [Mou92c]. cubic [BNS14, Fab90, Lam22, OT17]. Cumulants [Leh03, Mat04, Ber01a]. cumulative [Lah99]. current [DCL19, KW15]. currents [KKT10, WZ89]. Curvature [Zäh86, ACLP20, CC21, DHR21, Dar92, HRvR17, HD21, Sam22]. curve [BD23, GHM20, Zha16]. Curves [FI07, BK92, Bre08, BR97, RZ18]. cusp [DT93]. cusps [FT96]. customers [Gra05]. Cut [BV22, DDMP19, Fre19, CKM93, KV19, LLP10, MW17]. Cut-off [BV22, DDMP19, Fre19, LLP10]. Cutoff [BS19, BCS19, BN22, PTZ20, BF11, BCL23, DLP10]. cycle [BU11, Hou16, Nac03]. cycles [Lou22]. Cyclically [Pit96, HMO23]. Cylinder [HST15]. cylinders [CD13b, Dok04, Szn09, TW12]. cylindrical [BM02, PZ11].

d [FG19, ALS20, BF11, BCF15, CL07, CM11a, CV12, DS11, DR14a, DD16, DL20, FJL09, FR08, FL19, Hry98, Leb17, Mél01, PV97, RZ09, SS15, Yao19, vdHL10]. D-Ising [CM11a]. d’algorithmes [MP87]. damping [BL24a]. dans [HS90a, Léa87a, LT88, NW91]. Darling [Ein89a]. darning [CF08]. Data [YS92, AL02, BSV24, GH10, GS04, Hal88, HH90, Hös91b, KV08, LMW89, Mél01, Oua22, TA96, Wil87, Yur11]. data-driven [Yur11]. d’Atiyah [Léa88]. datum [HLR23]. Davis [CSK99]. death [CCM16, CMMM11, DLP10, FKL04, FM86, Joh86]. decay [AL91a, AL91b, BLM24, CRS18, CX96, DCRT19, FL93, IV00, LT21, Mön24, NST24, Rob10, Shi97, Swa18]. decimations [FR04]. decision [Alt89, Bor88b]. decomposable [Ger86, Thu86]. Decomposition [Lia00, PT02, BNT05, Bau13, Bel08, DK99a, DS18, FRS23, FT96, Hos91a, Jan95, Kra96, Li11, Li14b, MP88, Pen99, Tru03, Uma06]. decompositions [AGKW09, CHW88, Che93, Dub06, Eld20, FK97]. Decoupling [MT87]. decreasing [MP87]. Décroissance [AL91a, AL91b, FL93]. décroissant [MP87]. deep [LLT23]. defect [BSV10]. defined [GLL20, KLS86, LW18, SCZ21]. definite [O’C21]. deformation [LL24]. deformations [BGGM12]. deformed [CP16, CES23, DŚ19, LS16]. Degenerate [ABBP02, ACS21, BRR11, BL07, BL24a, Bru13, DNS18, EP19, HMY02, Lap90, Léa87a, Léa87b, LWX04, MY99, Wan18a]. dégénérée [Lap90, Léa87a, Léa87b]. Degree [ÜZ97, DMX21, GN22, JPRR18, ÜZ93]. degrees [Jon09, KSS86]. Delaunay [BKP21]. delay [Dyn93a, HMS11]. Delocalisation [LO24, Lam22]. Delocalising [MPS18]. Delocalization [BE17a, BG13a, GT05a, LT20]. delocalized [YY21]. delta [NM89]. delta-sequence [NM89]. denaturation [BL24b]. denoising [KLP01]. dense [GIM22, GIM23]. densité [Léa87b]. densities [Ale02, BGN91, Day88, DR14a, Dri13, Dzi91, GRT20, HM88a, HM92, Imk92, Kal14, KH03, KM00, Kum93, NRS23, Pic96, Pic10, QRZ03, Sch23, Wit86]. Density [AJJ22, GJ18, LMW89, Bir86, Bir14, BCZ^{+}10, Bru16, CRS18, CMS15, CR86a, CH88b, Den00, DP96, Gaj89, GKS04, GN08, GN09, GL14b, HM87, Hal90a, HH90, HM91, HI86, HK20, Hös96, KNP12, Kon90, Kot88, Léa87b, LV86, MMF92, Pes02, PR05, PP91, Rio94, Son04, Vie97, WT04]. denumerable [CR93]. Deny [Bia92, Rau88]. dépendantes [Rio96]. dependence [CT18, DP05, FIN99, FT87, GGLS20, Kül98, Rob94, TT91, Yas02]. dependency [RR03]. dependency-neighborhoods [RR03]. Dependent [BBS00, ACS21, BBG12, BB24, BJMW11, CM96, CK08b, DL03a, DD16, DDG12, DG22, DM23, DÜ23, ES03, Flo90, FR05, FR95, GHM23, GKdV89, GS90, Hös91b, HTV89, Jen91, KO01, KM93, KR05, MPR11, MM13b, MW13, Mor87, Pet90, Rio96, Wan94b, Wan18a, Wit87a]. depinned [CLM^{+}12]. depth [BGP24, Bru19]. Derivation [Fun89, Oel89]. derivative [LS01a]. derivatives [AW15a, DD05, Gaj89, Nor93]. derived [Mie03, Mie05]. Derrida [CHS22]. description [BG02a]. design [Bar00a, Hal90b, LBE18]. desquels [Wer94b]. Detecting [PV20, BBN20]. detection [BDN17, CMW15, GV16, PSSS13]. Determinantal [Gho15, BFQ18, Kat15]. determine [JPRR18, LQ13]. determined [Bry88, GW04, Hu94, MS21]. determining [HW86]. Deterministic [ABR08, CD20, CD21, CFM20, CF03, DS18, Mal12, Par22, Ver11]. deux [Car86, LR90, Mat95, Thi91, WY93]. deviation [BGdH10, BGdH23, BGP95, BW17b, CH19, CK88, Com89, Don97, Fil89, Fin02, FL96b, FS00, Gan94, GL07, HK01, HLLS23, Iof95, JZ15, KKT10, LV86, O’C98, Pes94, PR05, PPZ99, QSY24, SY01, Wat87, Yos93b, Zhe95, dA00, vdGL13]. Deviations [EL04, FKT88, Rid04, AK08, AN92, AKLP23, AG95a, AG95b, AG97, AC03, Ass08, AS21, Bal86, BS88, BGGM12, BR90, BC15, Bor03, BDE00, BCG88, Bre08, CX96, CT11a, CR95, CL04, CGZ00, CP03, CGZ03, CD90, CGM10, DG94, DGPZ02, DP96, DGI00, DMRYZ04, DK87, DE92, DS88b, ES03, EK96, FLT19, Fen94, FT04, Fin02, FW08, FL96c, FN01, GN22, GZ93, Geo94, HK03, HV15, HM24, IR06, JLLV93, Kif95, KV14, LS88, Léo95, LS15b, Lip96, Mar10, MPR11, MW13, MW16, MS02, Müc94, Num90, Oll87, Oll88, PV97, Pis96, QRV99, Roz86, Sch96, Sep93, Sep98, Sow92, Ver03, Yat13, Yil11, Zha14]. DF [HJS88]. Diaconis [EV21]. Diagonal [Lie98, AL91a, AL91b, Kaj13, Kor22]. diagonale [AL91a, AL91b]. diagram [JT17, MT11]. diagrams [DŚ19]. diameter [HvdH11, HvdH19, MNS21]. diamond [BD23]. diamonds [ACJvM15]. dice [HMZ20, SS24]. dichotomy [BSC02]. Dickman [AB15]. Dickman-function [AB15]. dictionary [ASS00]. diffeomorphisms [Bax86, BS88, Bax89, LQ98]. difference [Kry00, Str86a]. different [BNS14, HV16, Pri96b, PR08]. Differentiability [AD13, Bel86, Cla86]. Differentiable [Arn97]. differential [AR91, AS95, AI98, AM95, ABR08, ABBP02, BT96, BC01, BL24a, Bla05, Buc91b, Buc92, BN94, BP99, BQR00, CJ94, CIK99, Cos92, CH88a, DP98, DT10, DT12, DNS90, Dyn91b, DK04, Fag90b, FZ05, FS00, GMM17, Gyö86, GP93, HP95, HMY02, IL01, Iwa87, Ker89, KKM99, Kim04, KS88a, Kot92, Kot95, KK10, Kry14, Lac16, LM01, Lee90, Li94b, Li94a, LDS02, LWX04, Li16, MPY94, MY99, MZ02, MN93, Myt02, OP89, Osa12, OT20, Osa22, PP94, PT99, Rei89b, Rut90, Sai87, Sch23, Sch90, Tak87, Tsa16, Wan18b, WZ87, Yon97]. differentially [Ose08]. differintegrals [KT07]. diffraction [Kül03]. diffusian [CRZ96]. Diffusing [CL97a, NT86]. Diffusion [BK04, BPP97, BB07b, FP96, FS00, HK03, HP91, Kom01, MN98b, NT87, Rei89a, Rho09, AK87a, AR89, AS88, AL02, Bai08, Bal86, Bar90, BBPS21, BJL17, BKM24, BCFN89, BS95a, Buc92, CRS18, CF94a, CD13a, Cer03, Cer05, CF09, CW93, Che93, CZ16, Deu87, DDL90, Dit88, DS09, DGL87, FLT19, FS10, FV00, FL96b, FL96c, FW04, FW07, FW12, FG19, GJMM24, GM89, Gra95, GH20, Hab18, HSWZ18, IN16, JLLV93, Kan00, Kol08, Kot88, Kry15, Kry21, KY00, KU97, LV96, LOY98, Lap90, LPW13, Léa87a, Léa87b, LR90, Luc14, MMW15, Mou92b, MT11, Mur24, Nag89, Oel87, Oel89, OS13, Osa98, Osa01, Pic02, QRZ03, Rei89b, RS23, Sow92, Stu93, Wan95, Yos96]. Diffusion-limited [BPP97, CRS18]. diffusion-transmutation [FL96b, FL96c]. Diffusions [Aeb93, DG21a, Ebe97, BK17a, Bañ87, BP87, BBZQC10, BLR02, BG17, BRS14, BO09, BDE00, Bru13, CRZ96, CM02, Cha92, CGS95, CFG96, DRZ02, DR06, DR07, DG93b, DG93a, DG94, Day88, DD05, Ebe16, EHJT13, EP19, Eri90, Eva89a, Feh17, FIKP13, FG94, Fra05, FKW19, Fuk08, FT96, GQZ88, GKW01, HHW94, HR91, IVW18, KR16a, KTT89, KM00, KS88a, KL93, Kre97, LRW23, LR90, LS22, Lim99, Lip96, Mat94, MNS89, O’C21, Off14, OS95, Pal12, Per92, Pow19, QSY24, Rai97a, Rai09, Ram88, Ram90, Ram98, Str16, Stu95, Swa00, Swa04, Szn89, Tay89, Tru03, Uch94, Wan97c, Yos92, Yos93a, vV89]. Diffusive [CM01, FG94, HTV12, BMP04, BS95b]. diffusivities [GK24]. diffusivity [BBPS21, FKW19]. digraphs [BCS18, Coo17]. dilatation [AM96]. dilations [FM89]. dilute [BGW13, CM01]. diluted [PT04, Tal16]. Dimension [BBG12, GHM20, PS16, ABBL24, ASS19, Ber03, CLO01, DDMP19, DGZ05, EK23, GG96, Hin13, Hut21, KS91, KSX08, KSX09, LS00, Lal09, Liu96, LST20, MW17, MP19, PR20, Roh11, Sab11, Ser95, Szn11, Urb90, YK88, Zha19, Zyg09]. dimension-free [KS91]. dimensional [AB22, AHR01, Ald86, AST19, AC87, ACM06, Ass01, Bal89, BCK21, BLM11, BDC12, BK17b, BRZ20a, BRZ20b, BCF19b, Ble93, BR01, Bol90, BO09, BM11a, Bur89, CMT15, COP93, CRZ96, CM13, Cha19b, CGZ00, CGZ03, CSC93, CS23, DT00, DRZ02, DL03a, Deu87, Din13, DM16b, DG10, DCTT18, DCH24, DGL87, DS88b, Ebe97, El 18, EZ22, EF98, Fon09, FR95, Fuk04, Fun89, Gol07, GM13, Gyö86, Has89, HHW94, HvdH11, HvdH19, Hig93, HRvR17, Hol03, HT10, HLN13, Hut20b, Iwa87, JLS09, Ji89, KM22, KR16a, KTT89, Kes86, Kön93, Kön94, Kor04, LS97b, LBE18, LW00, Luc91, MM90b, MR09, MNS89, Osa12, OT20, Osa22, PV20, Pap87, PPZ99, Ram98, Rei89b, Ros89, Sai87, SV86a, Sch09, Shi92, Shi12, SCC19, Tsa16, TW20]. dimensional [Xia04, Xia07, YBK88, Yos96, Zha18a, dHNOS03, vdH01, vdHdHK03]. dimensionnelles [CRZ96]. dimensions [AR91, AK16, AC07, BBS00, BBS16, BK12, BR23, CJN21, CCGS91, DGRZ20, DP99, GS20, GHM12, GH19b, Har90, HST15, HP20, HP19, Iof95, JR08, KZ97, KS00, Kis15, LZ10, MNS21, MS92, PS93, QRV99, RR00, Tel90, XZ13, Yil11, Zam00, vdHS05, vdHdHS07, vdHH10]. dimer [dT07a]. dimers [BLQ23, BdT10, BdTR19, KSW16]. d’intégrales [JMP89]. d’intersection [Ber89, Ber90c, Mat95]. d’invariance [DLP87]. diophantiennes [Bre05]. Diophantine [Bre05]. directed [CQ20, CH02, Dau23, Kif97, Liu22b, Tak98, Var07]. direction [Hal88, PV20]. directions [HMT20, HV16]. Dirichlet [AR89, AR91, Ass05, BC01, Ber89, CQ24, Che92a, EF93, FS10, FL96b, GL14a, Joh86, Kry94, KY92, Lig87, LQ19, Mat94, MW01, PW20, Sab11, Son93, Tan97, Wan95, Yos96, Zhe95]. Disappearance [Yas02]. disastrous [Shi97]. Disconnection [Szn17, CN20, Qia21, Wer97]. discontinuity [ABR08]. discontinuous [BDE00, DE92, KL93, Mit86]. discrepancy [DFGGR19]. Discrete [BB07b, BG99b, CKK13, Kem05, Man86, Sei00, XZ13, AUY08, BKN14, Bel13, Bre08, CM24, DJP01, Din13, ELP08, GP05, Gei05, Hil87a, Hor00, Jen89, KV08, Lia22, MS87, Mar87, Mor86, MW05, Sch09, Szn09, Wu04, Yor92]. discrete-time [DJP01]. discretely [Coc18]. Discrimination [BG86a, Kan90]. discussions [KKR88]. disintegration [Kal14]. disintegrations [Kal07a, Kal07b, Kal11]. disks [BM17b]. Disorder [BGL19, BMRS20, CGH24, CK15a, GLT10, HNY08, Lac10, Zyg09]. disordered [BL24b, BH13, BEGK01, CSZ16, CD01, DG24, FM03, FIN99, Sep95]. Displacement [GRST14, DH91, HK92, LS15a, NZ17, NZ21, PSS15]. displacements [CH19]. dissipation [BBPS21, CP07a, LY97]. Dissipative [CSS23, DR02, DDG02, DR09, Hol96]. distance [Bel98, BPR11, Bir86, DGS21, FR89, FSX19a, FSX19b, FG15, Gaj89, GRS24, GHS19, HS90b, Hor00, KLS86, Kan90, KV19, Mas88, Mou14, Zha18a]. distance-regular [Hor00]. distances [Bob18, Bon20, DM22a, Dok04, DV90, GHS20]. distinct [Nán11]. Distoriton [GVV93]. distribuées [DD88]. distributed [BC17, DD88, Jia05, Jia09a, Pru98, Ruc88]. Distribution [CG04, GHM23, KR16a, Tei88, AW15b, AW01, BW22, BM12, BEES18, BW18, BL97, BC15, BD20, BG97b, CQ20, CV16, CCM16, CS08b, CS09, CH19, Com89, DG93c, EHJT13, FR89, Fel03, FS10, FNS87, GKS04, GN09, GPW09, GNW23, Hal90c, HS90a, HV09, Hil87a, Hoo91, Jen89, Joh19, KM22, KP89, KN03, KY13, Kry11, Kry16, Lah99, LS01a, Léa87a, LLR10, Lia22, LV86, Liu22b, LT86, Maj88a, Mas88, MM13b, MS90, MSW09, PR08, RT86, Sas86, Swa00, Tao13a, TT91, Urb90, Wei88]. distribution-constrained [BEES18]. distribution-valued [Kry11]. distributional [DD16].

Distributions [Bre05, BB24, Bar24, BM08, BBZQC10, Bel08, BG07a, BM15, BZ07, Bry88, CMMM11, DHS15, DH91, DR13, DR16, Dyn93a, EHM88, Ers19, Fal90, FF91, FW89, GTS84, GTS87, Gor87, GT16, Gyö86, Hor00, HTV89, JN91, Jia09a, Jur88, Kal03, Klü89, Lin87, MN98a, MiS09, Nie89, Nor86, Off14, Pan90, Pop09, QSY24, Rai03, Res88, Sas86, Szu92, TW01, Uch94, Von92, Von95, Wat00, Wat08, Wei86a, Wie02, Wil88, Xia04, Xia07, Zai88]. divergence [ABL12, Kim04, Lej01, MWZ05, MWZ08, Roz03]. divergence-form [Lej01]. divisibility [Eis03]. divisible [AT86, BJ14, BOR16, BG07a, CHR18, FQ88, GHV90, MiS09, RR89, Sam87]. DLA [AG14, LS19]. DNA [BL24b]. Do [Tay89, JRdlR08]. d’occupation [DM91a]. Does [HHT21, Aru15, LS19]. domain [BM12, Ber19, CT11a, CL90, Ein88, GM89, GH91, KM02, Le 95a, MS95, Sai87, SV03, Wei86a, Wil87]. domain-valued [KM02]. domains [Ass05, Bañ87, BB95, BH00, BY05, BW18, Ber95a, Che92b, Cli86, CMS00, CMS03, Cos92, DeB87b, DT12, Eba86, FT96, GH19a, Han12, Hua24, Izy17, KP89, Kim04, KMM98, Kry94, Kry14, Mur24, OS95, Pow19, Tru03, Xu91, Zha04]. domains-II [Eba86]. Dominating [EK96]. domino [LT15]. Donsker [Nic07, Ber20, Ber21a, Mue91a, Nic08, NRST16, Tal88b]. Donsker-type [Nic07]. Doob [Pen99]. d’opérateurs [FS88]. d’ordre [DD88]. dot [Qui10]. double [ACJvM15, Din13, DCL19, MW17, MS09]. Doubly [ST88, BKdS18, Fug88, PP94, Zha04]. doubly-exponential [BKdS18]. drawn [Rév92, vdBJ12]. Drift [BS95a, CL07, Aeb93, BBZQC10, Ber90b, Bor91, DD16, EL87, EZ22, Gro89, KLZ23, Kom01, Kön93, Kön94, KR05, Kry21, LY86, MMW15, NRW87, PS16, Rut90, Str16, Stu93, Swa04]. drifted [Luc14]. drifts [BB24, FZ99, GRT20, PR08, Tak87, Wan18a]. driven [AJV18, BC06, BD11, BFHM19, BH09, BCF15, CD13a, CHKN18, DR14a, Den00, DMP93, Fer89, GHM23, GMW24, Hau07, HKK23, HN09, KM02, Myt02, NP92, PZ11, Tre15, Yur11]. drivers [FS17]. droites [LY87]. Droplet [CL07, dHNOS03]. droplets [UA03]. dual [BDC12, Bia91, Bia92, LW20]. duality [BB03, CGRS16, DR99a, DR99b, Fag22, Fit87, GS86, GT20, Mit86, RR95, Sun98]. d’un [Ber89, Bia92]. d’une [Bié98, DLP87, Léa87b]. Dunkl [HM24]. duplication [MPS18]. dyadic [LW20, Rom14]. Dye [Eig86]. Dynamic [BC22, BBH24, BG02a, EdHM15, Kry14, SS87b, dlR95]. Dynamical [Kho08, ANV15, BMP97, BCF93, CG07a, Che95, CFM20, CMS02, CF94b, DHHL23, FW04, FW07, FK10, FM89, GHS21, HV09, Hof86, KLMH06, LS24, LRW94, PS98, PSS15, PSS20]. Dynamics [BPW09, AG95a, AG95b, BKMP05, BGP14, BBW14, BS17, BEGK01, CM01, CP07a, CLM^{+}12, CW17, DS20, DFM24, DM22b, Gay19, GPW13, GKR07, Gru96, GNW23, Gui97, LRW23, LT15, LLP10, LS21, MM15, Mor11, Nac03, Oel89, QSY24, Sta17, dHNOS03]. dynamique [dlR95]. Dynkin [Eis94]. Dyson [AH20, AEK19, BCL23, GS15a, HL19, IVW18, Tsa16].

E. [Kry97]. each [Mór91a]. easy [BCS11]. Edge [Ben10, CES21, Hua24, KV18, MR09, AH20, AD13, BYY14b, CS20, CDK18, CDS91, DKL02, MR07b, NRS23, RR17, Rol03]. Edge-reinforced [MR09, DKL02, MR07b, Rol03]. edges [CP16, GS04, Yuk95]. Edgeworth [BC04, Bob18, DH23, Fuk08, Jen89, KM09, RR03, Yos04]. Edgeworth-type [KM09]. Editorial [Ano17a, Ano17b, Ano21a]. EDPS [Bié98]. effect [CJN21, dBD02]. Effective [GK24, KZ98, Owh03, Szn02]. Effects [Hal90b, Lah01, IR89, MU90]. efficiency [Kou03]. Efficient [Coc18, DMX21, Par90, Eld20, HI86]. Ehrhard [vH18]. eigen [Jir16]. Eigenfunction [AC87]. eigenfunctions [Mic08]. eigengaps [NST24]. Eigenvalue [Wie02, EK87, FM12b, GN22, Ham00, HK20, LLS24, MW01, Ndu91, Péc06a, Péc06b, PR05, Rai03, Tao13a, Wan94a, Wan95, YBK88]. eigenvalues [Alh04, BGGM12, CT18, Cha92, Cha09, FI19, HK21, Jia09a, Jia12, Klo13, LS16, NTV17, Péc09]. Eigenvector [BW22, KY13, Ben21, BLZ20]. Eigenvectors [LP11c, Alh04, BG13a, BD20, LS16, LT20]. Einstein [BS97b]. elections [HMZ20]. electric [Tel89]. electrostatic [CL97a]. Elementary [vEH88]. elements [JP91, Mor87, Rai97b]. elimination [HT24]. Elliptic [BR01, Kat15, AM96, BK17a, Bañ87, BL07, BD14, BD23, CWZ95, DT00, DH23, FK16, KH03, Nol14, Pap90]. Elliptically [Gor87]. Ellipticity [CR14]. embedded [AKTW18, AD22]. Embedding [BCH20, Myk95, Ber97, COT19, FNS97]. embeddings [BS14a, BSS18, Cha12, MR16, Tro21]. embraces [LY87]. emergent [Mac21]. Empirical [BGV07, ACLP20, Ale87b, AGOZ88, Ber20, Ber21a, BC15, Bru19, CT18, CH19, CM96, Ded10, Ded13, DLP87, ES03, EHM88, EM97a, FG15, GT98a, Gol04, HJS88, JW24, Lah99, LP11b, LM09, Mas88, MP86, PR05, Rio98, Rio01, Yu93, Yuk87]. empirique [DLP87]. empiriques [Rio98, Rio01]. encoding [BM16]. end [DS13]. endedness [BvE24]. endomorphisms [Dup10]. ends [Cra93]. energies [LS88]. Energy [BDN17, BCD21, BS24b, BHW19, Cat98, CW17, CDP17, CHL18, FW24, FS17, GRS94, Hin05, KK14, Kül98, KZ98, Le 95b, MM15, Mer03, Ngu88, OS13, RAS14, SS15, Sub23, Tal06, Wan01]. Enhanced [BG02b, BBPS21]. enjoy [GY05]. enlacements [LY87]. enlarged [CLM94]. Enlargement [Imk96]. ensemble [AG22, CH16, FG02, MWW15, RSX13]. ensembles [BCF19a, Ben10, Col05, FR04, FR05, Geo94, Jia09a, KW16, LP11c, MSW21, NV21, Péc09, Yat13, vdBCL16]. entrance [AP00]. entre [Mic99]. entries [Au22, Jia05]. Entropic [CN20, NW22, Sam22, Vel04, ABBN04, BCG14a, BCS19, CGG07, Con19, GT20]. entropie [Rio98]. entropies [Eba86]. Entropy [CP07a, DM22a, LY88, Ale87a, AG97, BCL20, Bax89, Ces01, Eld20, GZ93, GRST14, Léo22, LL24, MMS14, Ram98, Rio98, Ros88, Sep95, Stu93, Mil12]. entropy-efficient [Eld20]. Entropy-Transport [DM22a]. entry [LLR10]. enveloppe [Gol96]. Environment [DGPS07, Aid08, BC17, Bau16, BD14, BMP97, BMP04, CR14, CH02, CV10, CGZ00, CP03, CGZ03, DR86, DD05, DG22, DR11, EdHM15, Feh17, FGP10, GI20, Gol07, Gol08, Gre90, GdH92, GZ12b, HS00, HS07, LN05, MR07b, PT17, PPZ99, RAS05, Rho09, Sab11, Shi97, Sta04b, Szn93a, Szn99, Szn02, Tót86]. environments [Bér05, BZ07, BGS13a, Fag22, FK16, GVV22, KTT89, Kif97, Off14, Yil11]. epidemics [Lal09, LZ10]. equal [BVBE20]. Equality [Yil11, BGN91]. Equation [FÜ04, Zha07, AEK19, AMS94, AHR01, AGM03, ALN99, BF11, BRR11, BG99a, Bia92, BEN23, BKM24, Bra86, BKWW23, BHW19, BQR00, CR88, CHKN18, Con19, CJKS13, Coo23, CL90, CL99, DLTV24, Dey16, DR14b, DGRZ20, Fon09, FJK23, Fun89, GR10, GP20, HLR23, HP91, HN09, Iwa87, Kif97, Kom01, LLT23, LY97, LOY98, LWX04, MS87, Mue91b, MP92, MS93, Mue98, Myt02, Nol14, OT17, Pra93, Sow92, Szn86, Tri95, TW20, Tzv10, Yan22, Yor92, dBD02, vdB00, Pra93, Bia92]. equations [AV24, AUY08, AR91, ASZ09, AL02, AS95, AI98, Ass01, ABBP02, Bai86, BT96, BC11a, BR24, BC01, BC06, BLM11, BL24a, BS22, Bla05, BR01, BG86b, BMS05, BCF15, Buc91b, Buc92, BN94, BP99, BQR00, CC99, CN88, CGM08, Cer05, CF09, Cha24, CT86, CWZ95, CMG95, CJ94, CIK99, CV98, CK12, Cos92, CH88a, DR02, DR09, DKN09, DR14a, DP98, DT10, Der03, DGT12, Dok04, DT12, DNS90, Dyn91b, DK04, EF93, Fag90b, FZ05, FIKP13, Fer89, Fil00, FG95, FR08, FL19, FL21, FS00, GG19, GGLS20, GM16, GM05c, GMM17, GP93, GK96b, HMS11, Hai12, HW13a, HW13b, HKR18, HLP24, Hol96, HP95, HMY02, IKS13, IL01, Ker89, KKM99, Kim04, KS88a, Kot92, Kot95, KK10, Kry00, KR05, Kry14, LS20, LS97b, LM01]. equations [Lee90, Léo95, Li94b, Li94a, LDS02, Li16, LQ19, MPY94, MY99, MZ02, MN93, MM17, Mél01, MR12, Mou14, MP11, OP89, Oel89, OB04, Osa12, OT20, Osa22, Oss05, Pap90, PP94, PT99, Pes94, PZ00, Pic89, Rei89b, RZ09, RSZ23, RS23, Roz03, Rut90, Sai87, Sch23, Sch90, Tak87, Tan99, Tay89, TTV03, Tsa16, Wan18b, Wat88, Yon97, Yon08, Zha10, Zha20, dA00]. Equiaffine [BR97]. Equilibrium [Cha94, CLO01, And86, BMRS20, CP12, DZ95, DGZ05, GR10]. Equivalence [BFQ18, Nie89, DR06, DR07, Fit90, Geo94, GN98, MR13, Sri00, Sun98, Wat08, Fer94]. equivalent [PPS96]. equivalents [BR92, Fer94]. erased [BLV16, Sch09]. Erdos [Bas23, Ein89a, Gri87a, HK21, LZ24, San94a, San94b, Sha92b]. Ergodic [Bug91, Rol90, AK87b, ACS21, BCL20, BCF19a, Ber90a, CFG96, DR06, DR07, DNS18, Dow88, Eri86, FS86, Fuk08, Häg05, KV08, LOY98, Neu90, Owh03, Ros88, Str16, Tho14]. Ergodicity [Dar88, DDL90, JR95, LO96, Ros89, Zha16, CK00, CK02b, EdHM15, Klo13, KM12b, Mou92b, RZ09, Suz24]. Erinnerung [Pen90]. erosion [GLPP17]. Erratum [Ano98a, BC11c, BR10, Ded13, DR16, HW13a, Kif13a, Li14a, Li14b, LN15a, Pic10, Tao13b]. error [DJ94, HM87, Hal90b, Lou97]. errors [KM93, MR03, MR13]. Escape [Bow18, Law92, SS87a]. espace [JMP89, Jou87, LY87, NW91]. Esseen [BCG14a, Bob18, CS01, DT23, DH23, Kla09, Neu24, Rio96, Tan86]. Essential [Wu04]. estimate [Bañ87, CK02a, CLW11, Dev88, Ein87, GS90, Hin13, LV96, LV86, Maj06]. Estimates [BKK08, Dok04, GT05a, Kum93, Wan95, Wan18a, Zai88, ABXY24, AK16, Ass08, BCL20, BK08, Blu05, BCH81, BCH88, CP07a, CFS01, CK08a, CKS10, CT11b, CM24, Com89, CV18, Con24, Cro08, CM08, Dar92, DHZ19, DMS05, EL94, FJK23, For09, FTY21, GRT20, HM88a, HM92, HK03, Hsi93, IR89, Kop19, LBE18, Li08, Li14a, Lie89, MMF92, MO15, Neu24, Osa90, Pic02, PPZ99, QRZ03, Stu95, Vel09, Wan97b, vdBT91]. Estimating [BB09, Gaj89, Hal88, HLO23, Ji89, MSW09, Bir86, CRZ13, DD05]. Estimation [Kar87, LR90, Mil89, Win91, Bar21, BLV12, Bir14, BCZ^{+}10, Bru16, CMW15, Chi12, Dah90, DH22, Div22, DM16b, Efr99, EM21, EP92, EGL14, Fab90, GN09, GP00, GL14b, GT98b, HM87, Hal90a, HH90, HM91, HJM20, HI86, Joh86, KLP01, KKM99, KT04, Kle06, KV19, Kon90, LM19a, LNS99, LV12, LMW89, MNS15, Par90, Par91, RB03, Rio94, Str16, Vie97, Wan97b, WZ99, WT04, vH11]. Estimator [Bar11, CR86b, GN09, HM88b, LS86, LMW89, Zho88]. estimators [Bab89, Bab91, BB09, CG92, CH88b, El 18, GKS04, GN08, HP94, Hös96, HR95, KNP12, KM93, LM19b, Mas92, Nic07, Nic08, NM89, Yos92, Yos93a]. Étude [Wei86b, Ber86b, Bié98, LY86, LY87]. Euclidean [AS92, BH00, HN97, KMM98, Pen23, SY01, Yuk95]. Euler [BT96, FL19, HLP24]. evaluated [Alt89]. Eve [CCL^{+}24]. even [CHL18, OV19]. event [KN03]. every [Müc94]. everywhere [FKN24]. everywhere-blowup [FKN24]. evolution [AV24, CSS23, CM11b, EF93, HL21, Pes94, TTV03, Zha04]. Evolutionary [DMR11, CM11b, Fin02]. evolutions [MQ20]. Evolving [MP05, BBH24, DK17, EM13, PWW11]. Exact [Ber03, DS87, DD17, Ein92, FM90, GT98b, HM88b, Haw98, Iof95, KT04, LU03, MRT23, NN04, Pin07, Shi12, Str16, BR92, BJ14, LT00, Liu96, MR23, MS09, PR08, CS23]. example [Eig86, GD90, MZZ24, Spe90]. Examples [DP05, Bér05, GY05, Key87]. exceedance [HHL88]. Exceptional [AB22, BS98a, KLMH06, DHHL23, PR20]. Excessive [BB00, FM86, Sal90]. exchangeability [Kal88, Pan15, Sun98]. Exchangeable [BP86, Cra17, Pit95, CFV23, FHP18, Kal89]. Excited [ABK08, RS12, Zer05, vdHH10]. Exclusion [BGS13b, And86, AK87b, ABL88, AM08, AJV18, Bah98, BGS16, CGRS16, CLO01, DFP15, Dit90, FJL09, Fag22, Fer92, FF94, GJ18, Gre90, LY97, QRV99, Yau97]. Excursion [Dub06, VA92, AD22, BMM20, CS04, HRS17, Le 93b, MS87, McG86, Rog87, Sal86b]. Excursions [Ber90b, DL03a, Fit87, Hsi93, Lap90, PPY92, Sal86a]. Existence [ASZ09, CC99, CR93, CMG95, DHR21, DR14a, DDG12, Eva89a, GK96b, LL17, LW00, TW93, CWZ95, CK12, Ebe97, Feh17, FS17, Hau07, HLP24, Li94b, Mue91b, Mue98, NS23a, Pic96, Pic10, RZ09, Shi03, Tak87, Zam00, ZD90, vV89]. Exit [ADH14, Dar92, DeB87a, DeB88, DO05, DM09, KMM98, AD02b, BGR15, BZ07, BCF15, Dok04, DK04, Kry16, SV99, Von92, Von95]. expanding [ANV15, BGR00, CC94, CMS02]. expansion [BKP21, CF24, Cas93, Chi01, DG21a, DvZ04, FvdH17, Fuk08, GM16, GM05b, Jen89, KY00, LP11a, Sug94, Yos93a, Yos97, Yos04, vdHdHS98, vdH01, vdHS02]. expansions [ASS00, AC87, BM08, Bar86, BKN14, BC04, Bob18, CG13, DH23, Fil89, GH89, Gu17, Jen91, Jir16, Kee88, KM09, Myk95, RR03, Roz86, Slu91, Wen96, Wen99, Wen04, Yos92]. expectation [Art01, JP10]. expectations [Bor03, CHMP02, KN98]. Expected [Tal05, DS13, GT16, HS11a, LM22]. experiment [Con19]. experiments [DV90, LS89, Lus92b]. Explicit [AJV18, Ino23]. explicitly [MPY94]. exploration [AMP04, AMP05, CN07]. Explosion [ADGO17, KR16a]. exponent [CGN14, CMS05, CS23, DM23, DZ19, DGS21, EdHM15, GHS19, HK92, JRS20, Key87, Shi12, Ter16]. Exponential [BB07b, CV16, CMS02, Din13, DCRT19, Fuk90, Hai02, LV96, LT21, Oda08, Sch89b, Shi97, TW20, vdBT91, BBS21, BLS87, BK12, BBPS21, BKdS18, BL97, BP18, CX96, CM08, DZ18, DÜ23, GN09, Kor17, Mam87, Swa18, AL91a, AL91b]. Exponentially [Kem05, BGGM13]. exponentials [ABM05, Gho15]. exponentielle [AL91a, AL91b, FL93]. exponents [Bax89, BGS16, Ber21b, BL90, GH23, KMZ11, Lia00, Qia21, Wan01, Wer97]. extendable [ACLP20]. Extended [JS86, CdH21, Che90]. Extension [KS94, AMP04, Bal89, CM91, Kal00a, LS00, LZ24, MRT19, MRT20, Mur24, NZ90, Val91]. extensions [AD02a, CPY94, CF08, Eri90]. Extent [HM87]. exterior [CMS00]. external [CDP17, Cro13, DK17, DSS23, Vel04]. Extinction [SV17, BBF12]. Extrema [Kon90, Dav99, KMP22]. Extremal [Ahn23, BB24, Dab90, GKdV89, GdH14, LS16, SB03, And86, ABK13, BGS13a, CT18, DNS22, OTV90, RS94, SZ17]. Extreme [Ber87c, KL87a, KL87b, AK10, BGGM12, FFT10, GSW86, HK21, JR88, Mar91, Naj18]. extremes [Fal90, Gri88a, JM90, Mij88b, Wei88]. Eyring [LS22].

F [LWR92]. F-valued [LWR92]. FA [HMT23]. FA-2f [HMT23]. Faces [HS11b]. factor [DSS07]. factorisation [McG89]. factorization [Ces01, Kal11]. factorizations [FLT22]. facts [Bir86]. faible [DLP87]. faiblement [Rio96]. fails [Bra89]. failure [BV95]. families [Arn97, Au22, BLS87, CM08, IR89, LS87b, Mat92, Sch15]. family [Ben10, CK12, HH23, Mam87, Osa01]. FARIMA [BM12]. farthest [CK12]. Fast [CF19, LS21, Vir02, Yan24, BP89, CFM20, EKZ22, HSWZ18, IN16, Mör01]. Fast-oscillating [Yan24]. fast-slow [CFM20]. Faster [FR07, Kem05]. fastest [OTZ24]. fat [AB19]. Fatou [BY05, Zie88]. fault [HR02]. FBSDEs [CD13a]. features [KS91]. feedback [Iha90, PT17]. Fefferman [CSK99]. Feller [Blu05, Cra17, LPW13, MS00, Sch98]. Fermion [Lin86]. Fermionic [Ben21]. Ferrari [IVW18]. ferromagnetic [Hig91]. ferromagnets [DG24]. few [Agg19]. Feynman [AGM03, CS03, CK04, GKS16, Léo22]. fiber [Léa87a]. Fibonacci [JRdlR08]. fibrations [BW17a]. fibre [Léa87a]. field [Aru15, ALS20, Aug21, BCL20, BM17a, BPR20, BGP14, BDS93, BEGK01, BV22, CX96, CDP17, CN20, CGH24, CD01, Cro13, DHP23, Din13, Duf92, DCL19, Eld20, Erd98, Erd01, Fen94, FM90, FR95, GS24, Har90, HKR18, HN14, HKK23, IK13, JRS20, KLZ23, Lac16, LLP10, LW18, Mac21, MWW15, MS02, NT86, Pan15, Pra93, Rod17, Sch86, SS13, Swi90, Tal00b, Tal06, Tal09, Tal10, Tal16, Vel04, Wan01]. Fields [Ano98a, DeB05, Fon09, AM96, AS89, AS91, Art01, BSS18, Bau13, BMM20, BPR10, Bre08, Buc21, CSS23, Ces01, CCKR07, Chi01, CF03, Ded98, DZ18, DSS23, DPR22, DF20, DJ12, EGL14, Gan94, GZ93, GH89, GMM17, GT05b, Guy87, HP94, HK03, Hig86, KW15, Kol87, LP19, LP11b, LWR92, LQ97, MN87, Mor92, Mor87, NW90, NUZ90, Oll88, OT20, PT04, Rod19, Sep93, SW95, She07, Tre15, Vel09, Xia97, Yos99, You89]. filament [FG02]. filling [MS17a]. filtered [DV90, Höp93, Jac90, KLS86]. filtering [CMM90, CC05, CL97c, DJP01, Flo90, Kry15, vH09]. filters [KV08]. Filtration [CHMP02, Cei12, CL17, Imk96, Rog87]. Filtration-consistent [CHMP02]. Filtrations [CL14, Las94, Str86b]. Finding [Yon97]. Fine [LST20]. Finetti [AP14, CFV23, Pet90]. Finetti-type [Pet90]. finies [GS89, Mic99]. Finite [Ald88, CGS95, MSS22, Ros88, AK87a, Ass01, BZ20, Bau13, BL12b, Bor03, BV22, BNPS18, CIvV03, CFV23, CK20, CP93, CV12, CG90, DV04, DO05, Ebe97, EG95, FS17, GS89, GT98a, GM05a, GLL20, Häg95, Imk94, Kaj13, Kif95, Kry00, Léo22, Lig87, Luc91, Maj92, Mic99, Ohl89, PTZ20, PS93, Ruk86, Sta01, Tei11]. finite-difference [Kry00]. finite-dimensional [Ass01]. Finite-time [Ald88]. finite-volume [CV12]. finitely [GL99]. finiteness [GS24]. finitization [FFN12]. fires [vdBN21]. First [ALS20, Boi90, AD13, AG23, BP94, CQ23, CT11a, CD13b, DHS15, DHHL23, DV15, DG24, DZ18, DZ19, DGS21, Dok04, Don89, Don12, DR13, DR16, DFS20, GS89, GM05b, HN97, KKR88, KZ97, LNP96, Pes02, Wan95, Yao19]. first-order [DV15]. first-passage [AD13, CD13b, DHS15, DHHL23, HN97, KZ97, LNP96, Pes02, Yao19]. Fisher [BCG14b, Dun24, GKW01, JB04, LS89, Tri95]. Fixed [AM13, Bar00a, CGS23, CGH24, FLT22, HRS17, JPRR18, LBE18, LT00, Men16, Mir13, vdBJ12]. FK [CM11a, FPW24, Li19]. FK-percolation [CM11a]. FK-representation [Li19]. FKG [LO24, LS88]. Flat [Shi92, Kre97]. flats [Aff88]. Fleming [BBF12, Han02, Kwa24, Sta02]. flips [CS20]. Floodings [BP20]. Flots [Aro89]. Flow [LQ98, Aru15, Bax89, Ben21, CC99, CT11a, CF19, DHR21, HRvR17, LOY98, LST20, Mac21, Tha96, War05, vdBS90]. flows [Bax86, BS88, BFF18, BL03, Cas93, Cra00, IL01, IN16, KM02, KO01, Kop19, Kor04, LL04, LR04, Li94b, Lia93, Lia00, LW00, RS23, Tre15, ÜZ92, Zha18b]. fluctuating [Ber01a, DF91]. Fluctuation [LY97, Sil17, BG86b, Cla87, Oel87]. Fluctuation-dissipation [LY97]. Fluctuations [Ahn22, BCD21, BS24b, CP16, Cha09, CDP17, DH96, DGRZ20, GT98a, HK21, IS19, Uch88, Zam04, AG14, Bel13, Bou10, CCC91, Cha94, CLO01, DMB20, DŚ19, DGO20, Fer92, FF94, FN15, FR95, GL99, GJ18, Hab18, JLS09, Joh00, KK14, Kar13, LLS24, Mai16, NST23, Pop09, Wan94b, Wat88, YY21, Zes88]. fluids [CF19]. flux [FG88]. Fock [Jou87, LP11a, Lie98, PS87]. focusing [dBD02]. Fokker [ASZ09, BR24, RXZ20, Zha07]. Föllmer [EM20]. following [KPyWmW88]. fonctions [Fer91]. force [Kif97]. forced [Shi06]. forces [BFHM19]. forcing [DV15]. Forest [Mor03, HP19, vdBN21]. forests [HN17, Hut20b]. forever [Kwa24]. forget [LS19]. form [HMS11, Kim04, Lej01, LQ19, Mij88a, Roz03, Tal09]. formal [Jen89]. Formation [GLPP17, DG93b, Tha96]. forme [Wer94a]. forms [AR89, AR91, AI98, Avr88, Bau13, BG97a, CQ24, CG07b, FT87, GS90, KN98, KS88d, KY92, LP11b, Mat94, MT87, NO15, Oli16, SB03, Tan97, Tan86, WZ87, Yos96, dJ87]. formula [ABF23, BC11a, CC05, Dub06, ERV02, FSX19a, FSX19b, FP00, Gal22, GKS16, GHM20, HN14, IK13, KM02, Kry10, Kry11, LPS16, LY88, LS22, Léo22, Ndu91, Nic16, RT02, RV96, Ust88, ZD90]. Formulae [FI07, Nor93, Zam02]. formulas [AM96, BM19, GT20, Ino23, PS93, Thi91]. formules [Thi91]. fortes [Rio00]. Forward [BS14b, PT99, RV93, HP95, MPY94, MZ11, Yon97]. Forward-backward [PT99, HP95, MPY94, MZ11, Yon97]. four [ASS19, CS23, MPY94, Sch09, Shi12, Yil11]. four-dimensional [Sch09]. Fourier [CH93, DM91a, Hou90b, Ngu88]. fourth [OT17]. Fractal [DvdW01, Arb91, BR23, DCMT21, DL05, Ham92, HK03, Hu94, KZ98, MPV01, TZ01, Tel90, XZ13, Zäh98, vdB94]. fractal-like [KZ98]. fractals [Gra87, Hin13, Kaj13, Kum93, KK96, KY92, Tel89]. Fractional [GM06, Wat93, AL02, BR24, BC11b, BC11c, BH07, BG24, BNS14, CT11b, CHKN18, CQ02, DK99a, DK99b, Dri13, DvZ04, GHM23, Gri91b, HS13, HN09, MN02, OB04, Ort89, PS16, PT00, RR00, Tal98a, TTV03, Wan03, XZ02]. Fragmentation [AD08, AD22, Ber00, Ber01b, JN08]. fragmentations [BBCK18, Mie03, Mie05]. Fréchet [Cla86, HLW24]. Free [AK10, BLQ23, HPU04, Le 95b, MMS14, Mło99, Tal06, Ans99, Aru15, ALS20, BNT05, BCD21, Bel08, BG07a, BV95, BPR20, BS98b, BS24b, Cam87, CJN21, CDP17, CN20, CG13, Col05, DFP15, Din13, DPR22, DHK22, Fag91, FW24, FT19, Fre19, GOZ21, GW86, HH23, HHW18, HN14, IK13, KS91, Kar07, Kar08, LS88, LW18, MSS22, Neu24, Ngu88, RAS14, Rod17, Ros88, SS13, She07, Sub23, Zes88]. freedom [KSS86]. freeness [HNY99]. Freezing [MS99]. French [AL94, Aro89, AL91a, AL91b, Asp89, AFZ87, Bar99, Ber86b, Ber89, Ber90c, BW93, Bia92, Bié98, Bou88, Bre05, Car86, CRZ96, CSC93, DD88, DM91a, DLP87, DS88a, Eis90, Fer91, Fer94, FS88, FL93, Fou98, GS89, Gol96, Guy87, HT86, HS90a, Has88, Jac87, JMP89, Jou87, KPyWmW88, Lap90, Lar95, LY86, LY87, Le 87, Le 88, LM92, Léa87a, Léa87b, Léa88, LR90, LT88, Leu10, Mai89, Mal95, Mat95, MP87, Mic99, NW91, Pet92, Pra93, Rau88, Rio96, Rio98, Rio00, Rio01, Thi91, Val91, Wei86b, WY93, Wer94a, Wer94b, dlR95]. frequency [NRST16, PRT13]. friendly [TY03]. frog [KZ17]. front [FL96c]. Frozen [Kis15, vdBN21]. full [CP16, DLTV24, HKK23]. Fully [Der22, BK04, MPS19]. Fully-connected [Der22]. function [AB15, BLV16, BMS08, CH02, CK02a, CR86b, EP92, GN09, GH10, Hal90c, Ham00, HI86, Hös91b, Hua24, JP10, Kot92, Kry16, Lah99, Lam22, LS01a, LNS99, LM00, Lie89, Maj88a, MO15, MR07a, Naj18, Son04, Sta17, Tan99, WHZY07, Wat96, dT07a, vdBB94]. Functional [DJPP13, Gra05, KM86, KP86b, Wat87, BCL20, Bal86, BT03, Boy19, BMS08, Cap95, CMS07, CMS23, CMN92, CGZ00, CGZ03, DIT14, DKP22, Fra05, HLØ^{+}93, KL93, LM01, LQ98, MM90b, Mij88a, Mil89, Mil12, MR95, MN02, MO87, Nie89, NN20, San94a, San94b, Ver90]. Functionals [Ito88, ADN22, BZ12, CG94, CS89, Che00, Cla86, DPR22, EL94, Eri86, Gin94, GS07, GRS94, GP00, HI86, HS94, Ji89, Kal95, Kle06, KY00, Mit86, Num90, RS91, Ros87, SY01, Shi94b, Slu91, ÜZ94a, WZ93, vW97]. functions [AW15a, Aeb93, AA88, BLS87, BLBE92, Bar21, Bar86, BY05, BW18, Bla88, BKK08, BGS13b, CP12, Cer24, CKS11, CG92, CR95, CLL20, CM97, Dab90, DG10, Düm93, EM97a, ERV02, Fab90, Fer91, GRAS17b, GW86, GH89, Gro89, GT98b, HS95b, HM24, Ino08, KN98, KMZ11, Mas92, Men18, MW14, RZ18, RV96, Sal90, Shc13, Shi94a, Sug94, Tro12, Urb90, Wee88, Wee90, Yuk87, Zäh98]. functions-skeleton [Sug94]. Fundamental [LM19a]. Furstenberg [BE17b, Car87]. Future [Tho88, GM05a, HS95c, KLL94a].

G [ABZ23, Coo23, Han12]. G-inequality [Han12]. Galton [ADN22, Kor17, Aid08, Aïd14, AdR17, Bow18, CFK13, DGPZ02, FW08, JRS20, LRW23, LS00, LPP96, PZ08]. game [AB11, Bal98, vH18]. games [Fin02, HMM19, Kry14, Lac16, LM17, Mor86, Nag87, RSV01]. Gamma [AB15, Bon90, BCD21, Bon88, CN16]. gap [ABXY24, CM01, DKW23, EKZ22, GJS21, KM12b, Mic99, Tao13a]. gaps [AF12, Fuk11, NTV17]. gas [BBT23, CM01, Cha19b, Con19, Der03, FM03, GS20, Qua99, Zes88]. gases [Leb17, SS15]. gasket [BP88, HHH02, HHK90, HH91, HK92, HHW94, HS90b, LQ19, PP91]. gaskets [Ham00, HHW94]. Gauge [Maz99, Stu91, CS03, Gan99, GRS94]. Gauss [FNS97, Nic16, WZ89]. Gaussian [AKTW18, AG22, AHR01, Aru15, ALS20, ABF23, Avr88, AFZ87, AW01, BG86a, BBG12, BJ20, Bau13, BLM24, BMM20, BSC02, BWW18, BPR20, Ber87c, BW93, BDN17, BM07, Blu05, BG24, BG86b, BFO09, Bor08, BCF18, BNPS18, BF18, CY15, CCC91, CX96, CH02, CF24, Cas12, CS89, CHKN18, CCK15, CN20, CL91, CM96, DS18, DR14a, DM17, DV00, DMB20, Din13, DZ18, DŚ19, DPR22, Ein88, EM97a, Eis03, Epl86, EGL14, Fel03, Fer91, Fle12, FM12b, Fuk90, GN22, GS24, Gin94, GG96, GK91, Gor92, Guy87, HJM20, Har04, HN14, HT24, Iha90, Kol87, Kon90, KL87a, KLL94b, LP19, LW18, Mam87, MR95, NN04, OT17, RT86, Rod17, Ros87, SS13, Sch05, Sha92b, SW95, She07, Sow92, SB03, Tan86, Web89, Wei86a, Won20, Xia97]. Gaussian [Zai87, dlR95]. Gaussian-embedded [AKTW18]. gaussien [dlR95]. gaussiennes [BW93, Fer91]. gaussiens [AFZ87, Guy87]. Gelfand [Met13, Pet14]. Geman [BY86]. genealogies [FRS23]. genealogy [BL00, Lam02]. General [FW89, Kal89, Pop09, PS93, ADN22, AH20, AEK17, Au22, BL21, BCD21, CMS03, DHS15, Fou98, FT92, GH23, GT98a, HMS11, HH05, Hoo91, Hös96, HL19, Jon09, KM96, Kho08, Kif97, Kry94, Lac16, LLS24, Led00, Liu22a, LM19b, Men18, Müc94, MS90, RR95, SV17, Tal09, Yin14, dA00]. générale [Fou98]. généralisant [Ber86b]. generalisation [FI19]. generalised [FM12b, vdHS02]. generalization [FHS90, HS87]. generalizations [Blu05, FR05, MR23]. Generalized [Bla88, BRS96, Bon90, CM08, FvdH17, Ito88, KS18, Mit87, NZ86, PZ98, Qia21, Agg19, AA88, BL24b, Ber95b, Bon88, CGRS16, Chi01, CLDDMG99, El 18, EYY12, Fal90, GN98, HJS88, HM24, JN91, JZ10, KT07, KN98, Lov95, Mir13, Sha92a, Tho14, WHZY07, Wei86a, Yau97, dJ87, vdBG13]. generalizing [Ber86b]. generate [Kol11]. generated [CL91, KMTT19, Yas02]. generating [DSC96, RS91, vdBB94]. generation [CSS23, FLT22]. generations [FW08]. generator [BL91, GP20]. Generators [Ans13, BRS14, BH08, CMG02, Ros88]. generic [HLR23]. Genetic [Bér05, Sta04a]. genetics [EP19, FRS23]. genus [Cha10a, FLT22, Lou22]. geodesic [Dar92, DZ19, GH23, Le 95b, Liu22b, Ndu91]. Geodesics [GRAS17a, GPS22, ACLP20, MQ20]. Geography [DW93a, Jon09]. Geometric [Bur89, Hol11, KM12b, NU91, OTZ24, Cra00, Gor92, Joh19, KY00, LB08, LQ98, PSSS13, YSA17]. geometrically [Häg05]. Geometry [ABS21, BEES18, BS24a, DFSX20, AJLS15, BCH20, BBN20, CFV23, DDG12, El 18, GH23, GdH14, KMM98, LR21, MS16, MS17a]. geometry-dependent [DDG12]. German [Pen90]. giant [JPRR18]. Gibbs [AD02a, AKRT97, BS05, BK07, BGP94, COP93, Ces01, Com89, CDCIV14, CK15a, DRZ02, DZ96, Deu87, FG02, Hig91, Hor00, Ji89, LS22, LMV98, MS10, Oll88, Pap87, Rob10, Tzv10]. Gibbsian [CRZ96, DDG12, Geo94, Kül01, LS88, Mas92, Mer03, You89]. Gibbsienne [CRZ96]. Ginibre [BD20, OS16, OS22]. Ginzburg [Fri89, Fun89, Fun91, FN01, LN05, Nis03]. Girsanov [Buc91a, BF93, GRT20, Pri96a]. given [BPR07, HLO23, Mór91a, NRS23, Sch89a, Shi94b]. glass [AG95a, AG95b, Cat96, Gru96, Gui97, PT04, Pan15, SZ17, Tal16, Tou98]. glasses [AC13, ADG01, Cha10b, FW24, GKN92, Gan99, JT17, Mac21, Maz99, Tal09, Tal10]. Glauber [Gia95, BKMP05, DS20, Gru96, LT15, LLP10, LS21, Mor11, Nac03, Sta17]. Glivenko [HJS88, Yu93, vH13]. Glivenko-Cantelli [HJS88]. Global [ADN22, ABL12, CT11b, DLTV24, GS15b, HLP24, Mar91, RS23, BCF93, CC99, CFR93, DMB20, GdH92, NS23a, RSX13]. Gordon [LRV23]. governed [BFQ18]. Grad [Uch89]. Gradient [Cha94, Pic02, BK07, Buc21, CV18, CK15a, Kop19, Vel09, Wan97b]. grafting [EPW06]. Graham [EV21]. grain [LS01a]. grands [Rio00]. granular [CGM08]. graph [ABS21, AKLP23, BGJ96, Cra17, DNS18, DMX21, GKN92, GHS20, HvdH11, HvdH19, Hor00, HP19, JPRR18, LW18, LW20, Ruc88, vdHL10]. graph-valued [Cra17]. Graphical [Eic12]. graphon [KV19]. graphs [ABBG12, ADS16, AL00, BR21, Bas23, Bel98, BvE24, BKMP05, BBW14, BSW17, BvdHS18, BLS13, BC15, BBN20, BDW21, BP20, CF19, DHR21, Dal88, DV04, Dom17, DJPP13, EL21, EM21, Ers09, EGL14, FS00, Gam14, GIM22, GIM23, GRST14, GM14, GKS16, HP99, HK21, HRvR17, Hut24, JN91, Jia12, Jon09, Lam22, LW20, MRT23, MR07b, MR09, MMS12, NRS23, O’C98, PSSS13, Sam22, SV17, Tel89, Tel90, TY23, Vir02]. gravity [BHSY23, GHS19, GPS22, MQ20, MS21, MSW21]. Gray [BD88]. greater [Zyg09]. greatest [PR12]. Green [BLV16, CK02a, KMZ11, MO15, RZ18, Son04]. grid [RST^{+}13]. Griffeath [BD88]. Grothendieck [GV16]. Ground [GS20, Mer03]. Group [AD02a, AS10, BS19, Bia92, CR95, Hou16]. groupe [Bia92]. groupes [Rau88]. Groups [BB07b, Dun07, AUY08, Ale02, AG23, BCS11, BL91, BE17b, BSC02, BP94, BLV12, DSC96, DO05, DM09, DG10, DG21b, Ers19, Fel03, Fel05, FNS97, Fre19, GW86, GM13, Hou19, Jaw95, Mat04, PTZ20, Rai97b, Rau88, Ros88, SCZ21]. grow [JRdlR08]. growing [AS88, BB21]. Growth [AD22, BB07b, DW07, GS03, Sch98, Ale87b, Ale02, BE17b, BSV10, BBCK18, BCF18, CM13, Cro08, DS97, DHB11, EL21, EPW06, GRAS17a, GRAS17b, Gol08, GdH92, GdH94, Hut24, KS95, LPW13, Sil17, dHNOS03, Fou98]. Growth-fragmentation [AD22]. growth-fragmentations [BBCK18]. GRR [FTY21]. GUE [ACJvM15, CP16, NST24]. GUE-minor [ACJvM15]. GUEs [Bar01]. Gumbel [Bel13, Joh07]. Gundy [CSK99].

H [Ove94]. H-transforms [Ove94]. Haar [Jia05, Jia09a, Par22, PR05]. Hadamard [HN14, IK13]. Hahn [Kor22]. hairpin [ST88]. Half [BGP24, AR18, BGN91, Fuk04, IR10, KMTT19]. half-line [Fuk04]. Half-space [BGP24, IR10]. half-spaces [BGN91]. half-spheres [KMTT19]. halfspace [Bru19]. Hamilton [BRM24, GG19, GGLS20]. Hamiltonian [DL20, FFL94, Fun91, GNW23, Kor04, LO96, Yan24]. Hamiltonians [GGLS20]. Hammersley [AD95, BGGS18]. Hamming [EI20, vdHL10]. hard [BG02b, HMM19, Osa98, Pap87, RR17, Yur11]. hard-core [HMM19]. hard-wall [BG02b]. hardly [FR92]. hardness [MWW09]. Hardy [Wei90]. Harmonic [DT10, KX09, BY05, BH13, Bla88, BKK08, GW86, Grü86, Shi94b, Tha96]. Harmonicity [FM19b]. Harmonizability [Hou90a, MM90a]. harmonizable [MS95, RS94]. Harnack [ADS16, BvE24, Wan18b]. Harris [Che00, HMS11, Lev88]. Harry [Ano21b, DCG21, Gri21]. Hastings [Sil17]. Hausdorff [KSX08, KSX09, LS00, Liu96, Mat04, She93, Urb90, Xia97]. Having [BB07b, FT87, Gyö86, Mit87]. Hawkes [LS00]. hazard [LV86, LMW89, MW90]. head [GSW86]. Heat [CK08a, CZ16, Mur24, vdBS90, vdB94, vdB00, AHR01, ALN99, AK16, AL91a, AL91b, BCK21, BG99a, CKS10, CT11b, CHKN18, CMS03, CK12, CJKS13, CS12, Cro08, DKN09, DHZ19, DGT12, Dey16, FL93, FJK23, Hai12, HK03, HN09, Kaj13, Luc14, MP05, Mue91b, MP92, MS93, Mue98, MP11, Ndu91, Nor93, Osa90, PZ00, Tay89, Tha96, Tri95, Wan97b]. Heavy [HM07, Mir11, BS24a, BG13a, Men17, MW16, Rob10, Sch00, vdBN21]. heavy-tailed [BS24a, BG13a, Men17, MW16, Sch00, vdBN21]. Heckman [Sch07]. hedging [DS14, Gei05, Kra96]. Height [Lam22, DÜ23, Hua24, Las19]. Heisenberg [DG10, GM13]. Hellinger [Bar11, Bir86, DV90]. Hellinger-type [Bar11]. Help [Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano90n, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano91k, Ano91l, Ano91m, Ano91n, Ano92a, Ano92b, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j]. Help [Ano92k, Ano92l, Ano92m, Ano92c, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano94k, Ano94l, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95l, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k]. Help [Ano99l, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e]. Help [Ano08f, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f]. Hermitian [Ben10, BWW18, CES21, CES24, LT20, Péc06a, Péc06b, Shc13]. heteroclinic [Bak11]. heterogeneous [Bah98, Cha16, DS23]. Hidden [Kor22]. Hierarchical [DG93b, GLT10, Lac10, Pan15, AS10, AP14, Cha19b, GS20, HS95b]. hierarchically [FG94]. Hierarchies [Ber93a, FHP18]. High [DM16b, FL21, Gu17, Kot88, NRST16, Rai97b, Tou98, ABXY24, BCF19b, BM11a, CJN21, Dah90, DP96, DZ19, El 18, EK23, EKZ22, GW04, Har90, HvdH11, HvdH19, HP20, Hsi93, Hut20b, KM22, KS00, LBE18, Lou22, MNS21, PV20, DM16a, SCC19, Szn11, Tal01a, Ter16, Zha18a, vdHH10]. high-density [DP96]. high-dimensional [BCF19b, BM11a, El 18, HvdH11, HvdH19, Hut20b, KM22, PV20, SCC19, Zha18a]. High-frequency [NRST16]. high-temperature [EKZ22]. Higher [DeB04, DeB05, DCvEH18, AW15a, Ber93a, DGRZ20, Mar08, Roh11, Yil11, IMS20]. highly [DF91]. Higuchi [CV12]. Hilbert [Asp89, CMG95, DR02, DR09, DDP86, Ein89b, Ein92, GP00, Hol96, MP91]. hilbertienne [Asp89]. Hipster [ABCM20]. histogram [BB09, CDR86, Hal90a]. histograms [YS92]. historical [BL01, BM05b, Dyn91a]. Hit [Kem05]. Hitting [DKN09, FFT10, Fuk04, Ram88, Ram90, Gol08, HP18, PSS15, QR13]. hives [NST24]. Hodge [Li11, Li14b]. hold [Aru15, HHT21]. Hölder [BR92, BK01, CJ94, Ebe97, FJK23, FT96, HLN13, MW89, MP11, RR00, Sch98, Xia97, Zha19]. Hölder-continuity [MW89, RR00]. Hole [BNPS18]. holes [Rév92]. holomorphic [DG10, Sug94]. homogène [Has88]. homogeneity [BDG^{+}09, Kol13]. Homogeneous [Ber01b, BF11, BJ99, CM02, CV98, EHJT13, FV00, FS08, GT05b, Häg96, Ham92, Has88, HL00, KNP12, Léo95, Lig96, Pel12, Sta01]. Homogenization [Cas01, Gra95, KK96, Lej01, AL02, Coo23, DGO20, Fag22, Gu17, KO01, Mou14]. homogenized [GM12]. homology [Le 95b]. homomorphisms [Mat04]. homotopic [Gru98]. Hoover [AP14]. Hopf [BW17a, Bax94, McG89, Uek89]. Hopfield [BGP94, BGP95, BG97b, EL04, GL99, Lou97, Tal98b, Tal01b]. Hörmander [BH07, CHM22]. hot [BPP04, Con19, DF01]. hot-spots [BPP04]. Huber [Buj86]. hubs [BB21]. hull [BF88, BH98, CQ23, CL16, Gol96, HLO23, Kho92]. hulls [BR97, CG94, Dwy90, Gro88, KMTT19, KM22]. Hunt [BL91]. Hwang [BY86]. Hydrodynamic [FM03, FJL09, Fag22, Fun03, Nis03, SU93, Fri89, Fun91]. Hydrodynamical [Bah98, Fun89, Zes88]. Hyperbolic [LS97a, AS22, BKP21, BKMP05, BJ99, Bud18, Cur16, DM22b, DG21b, EFL01, HHT21, MRT19, MRT20, MSS94]. Hypercontraction [KS88d]. Hypercontractivité [Lar95]. Hypercontractivity [CS08a, BBG12, HP18, Lar95]. hypercube [BB06]. Hyperfinite [Rei89a, Str86b]. hypermixing [CK88]. hyperplane [Nes19]. hyperplanes [HHT21]. hypersurfaces [Has89, Wie86]. hyperuniformity [GS20]. Hypocoercivity [Mon18]. Hypoelliptic [CM02]. Hypoellipticity [Kry15]. hypothesis [BB03, GM94, LT00].

I.I.D. [Ein89a, KM86, KM88, BD14, BMP04, Ein88, EM97b, GH89, Jan10, Muk91]. ice [vdBB04]. ideal [RR92]. Identical [Bra87, MiS09]. identically [DD88, Pru98]. identically-distributed [DD88]. Identification [MS90, DJ00, Guy87]. Identifying [MU90]. identiquement [DD88]. identities [Kal89]. identity [AMP04, ST16, ÜZ07, Wüt98]. if [JPRR18]. II [AL91b, HvdH19, Bor86b, BYY14b, BD87, BM05b, CS09, CJKS13, CKM93, DE92, DR99b, Eba86, GTT98, GdH94, HvdH11, HRS17, Hut24, KS88c, Le 15, Leh03, Mie05, Muk93, Por96, Szn96b, Wen99, Wit87a, Zäh86]. iid [Jon09, Par22, Tao13b, Tao13c]. III [MS21, Wen04, YY21, Yin14]. ilinear [Sha92a]. ill [CIK99, Gol04]. ill-posed [CIK99, Gol04]. image [BLV12, PT87]. Images [Rai03]. Imaginary [MS16, MS17a, AJJ22]. imbedding [Fug88, Gar95]. immigration [CR90, DL03a, Lam02, Win91]. Immortal [GK12]. impact [El 18]. imperfectly [You89]. impinged [Kni01]. implemented [BCZ^{+}10]. implications [Ald88]. implies [Reg23]. Importance [HSWZ18, ELY99]. Improved [RST^{+}13, Kop19]. improvement [BCH88]. improves [FL21]. impurities [vdBN21]. incipient [DS11, Kes86]. inclusion [Kim21, Kim23]. inclusions [ABR08]. incomplete [Kra96]. incompressible [CF19, Oss05, Zha10]. increase [Ber93b, DM97, Léa87b]. increasing [ADH14, AD95, Ass05, DHZ19, Eri86, Joh00, Sep98, Wat96, vEH88]. Increment [CCFS00, EM97b, MM13b]. increments [Bes91, DS87, Gri91b, Gri95, JS86, Kön94, Li92, Ort89, iS91, Sch90, Wee88]. indepéndantes [DD88, WY93]. Independence [CG07b, NU91, Spe90, Sun98]. independent [Ber99, Bes91, Bol86, Bol87, Bra89, Bry08, CP88, CD90, DD88, GK01, HS88, JS86, KL00, MT87, MWW09, Pan90, Pru98, PR08, RST^{+}13, iS91, Sch05, Sch90, Sep93, Swa07, Szn89, Tho88, Wee88, WY93]. index [Kal00b, KLP01, MW14]. indexed [Ale87b, BP94, Ble93, EM97a, EGL14, FW08, MP86, RRO24, Yuk87]. indicators [DP15]. indifference [FMS11]. indirect [GP00]. Indistinguishability [HN17]. induced [Dar88, LQ97, LQ98, ÜZ96]. Induction [vdHdHS07]. inductive [vdHdHS98, vdHS02]. Inegalités [Mat95, Rio01]. Inequalities [BGV07, Bor08, Vie97, AW15a, ADS16, AL00, AS22, AM95, BCL20, BGM18, BMN12, BC12, BL97, BG99b, BG07b, BRS14, CGG07, Ces01, Cha07, Cha09, CCKR07, CSK99, CMS02, DNS22, ELP08, EI20, GGM05, GL09, GL07, GRST14, GLWY09, HP18, HPU04, JB04, JZ15, KNP12, LP11a, LPS16, LB08, Liu17, MMS14, Mat95, Mil12, Pin07, RB03, Rio01, Tei16, TY23, Wan97a, Yur11, vdGL13, vdBG13]. inequalitiy [CV18]. inequality [BvE24, CGW10, CLW11, CP93, Con24, DSS23, DS09, EM20, Fle12, FHS90, GN09, GV16, Han12, Har04, Hil87b, HS87, JP10, Kar12, Kla09, LN05, LS02, MPR11, Ose08, Wan18b, Wu00, Yau97, Yos99, Zai87, vdBJ12]. inert [BBZQC10, Kni01]. inf [Gri91a, Gri95, Li92]. Infante [BEN23]. infections [DS23]. inference [BM19, Coc18, DIT14, Höp88, KSvdVvZ16, RS17, RD13, You89]. inferior [CS94]. infima [HS95c, KLL94a]. infimum [Föl89, Hil87b]. infini [CRZ96]. infini-dimensionnelles [CRZ96]. infinie [Lar95]. Infinite [Ass01, Bal89, BM01, BO09, CGH24, Deu87, JR08, KT07, KM12a, Osa12, OT20, Qua99, Tsa16, AR91, ACCH23, ADGO17, AA88, AC87, BBM00, BB24, BM08, BJMW11, Ber87a, BR01, BG86b, BLS13, CRZ96, CC21, CGS95, DT00, DRZ02, Dal88, DS11, DWXY23, DW90, Don97, DG10, Eis03, FG94, FM97, FT19, FS86, GKN92, GI20, GM13, GdH94, GKZ93, HP99, Häg03, HH90, HMT20, Has89, Hut24, Iwa87, Kes86, KS89, Lar95, LW00, LO96, MM90b, MR07b, MP88, MNS89, MS92, NN93, PS98, Ram98, RR00, DM16a, Sch99, Tho14, Xu91, Yos96, Zam00, Osa22]. Infinite-dimensional [Ass01, Bal89, BO09, Deu87, Osa12, OT20, CRZ96, DT00, DRZ02, DG10, GM13, MNS89, Osa22]. infinite-order [DWXY23, HH90]. Infinitely [BG07a, FQ88, AT86, BJ14, BOR16, BG13b, GHV90, KM12a, MiS09, Mło99, RR89, Sam87, Tan96, Tan97]. infinitesimal [GP20]. infinity [DSS14, KM96]. influence [Gri88a]. Information [AJLS15, BCG14b, Dun24, Eba86, GLWY09, Hal90a, Jac90, JB04, Liu17, LS89, WZ99]. inhomogeneity [Roh11]. Inhomogeneous [AP00, BP18, Mön24, AMP04, AMP05, AKLP23, BSW17, BvdHS18, BR23, BDW21, DH23, RB03]. Initial [CL07, AK08, BL21, HLR23, KR16b, KV08, LS19, LS21, Mél01, Oua22, Uch94]. initial-terminal [AK08]. inner [BR97, FM19a]. inner-product [FM19a]. innovation [CF03, MSW09]. Innovations [CHW88]. inset [Eba86]. Instantaneous [FKN24, CR90, CR93]. instruments [Bar89]. integers [Zer05]. integrability [Fuk90, Kor22]. integrable [Asp89, DG10, MT20, Asp89]. Integral [Res88, CT86, CC05, FV00, GRT20, GM05b, Imk92, Imk94, Jur88, Kal90, KS88b, KS88c, Nor93, PP87, Pic89, Slu91, Yon08]. integrals [AGM03, BG24, DK99a, DK99b, GMW24, IW94, JMP89, KM88, Kry02, KKT10, Kuw09, Lie98, Lio05, Maj88a, Maj06, MO86, MO87, NZ86, NZ90, Str86a, Tan99, ZD90]. integrands [Imk94, Lie98, NP88]. integrated [CS04, FT04, HM87, NN04, PS15]. Integration [AM96, BC11a, Fin02, FI07, Léa87a, PT00, Zäh98, Zam02, NW91, Asp89, Ble89, DM03, Pra93, RT02, RV93, Asp89, Pra93, Léa87a, NW91]. integrators [Ble93]. intensity [BB09, RB03, Rol90, Tei11, vdBBdH05]. inter [Pfe87]. inter-record [Pfe87]. Interacting [GKW01, O’C21, RS93, AD95, BLR02, BCG86, BS95b, BW17b, CP10, CGS95, CFG96, CK00, CK02b, DG93b, DG93a, DG94, DH22, DG21a, FG94, FGM09, GM89, GM94, GMM17, HD21, Kol03, LRW23, MS16, Oel87, Oel89, Osa98, Rai09, Rez96, Suz24, Swa00, Tro95, Uch88, Uch94, Yos96, Gia95]. interaction [BDS93, Com89, Fen94, Geo94, KLZ23, LWX04, NT86, Rai97a, Tal00b, vdHdHK03]. interactions [DDG12, DW90, Fun91, Kol03, NT87, NP12, Per92, Uel02]. interdependency [Hos91a]. interesting [Hal88]. interface [BG02b, BKM24, DD05, DGI00, DR24, FN01, Fun03, GLPP17, GRAS17a, Li19, Nis03, Tri95, Vel04, Zam04]. Interfaces [COP93, FPW24, Gia95, GT05a, Izy17, MSW21]. interior [MS17a]. interlacement [Szn11]. interlacements [LS15b, MS24b, Rod19, Szn09, Szn17, Tei11]. Interlacing [GZ18]. intermediate [BHR01, Dab90, IN16]. Intermittency [CHKN18, CM95, GdH99]. intermittent [Ded10, Ded13, Gou04, Ter16]. Internal [BB07b, AG14, KSS86, Luc14, MS99, Nán11, LS19]. internally [Mou92c]. Interpolation [BC12, BM19, DDMP19, TKPK86]. interpretation [BG07a, Fon09, MPS19]. Interpretations [FR05]. intersecting [Aff88, Joh02, KW04]. Intersection [Ber99, AC07, Ass08, BK12, BPR10, Ber89, Ber90c, BL90, CL04, Hut20a, Mat95, WY93, Wer97]. Intersections [MW17, Wie86, Ald86]. interval [AM88, Bug85, CMS02, Sch15, Ano87a]. intervals [BD87, CPW95, Hal89, Naj18]. intransitive [SS24]. intransitivity [HMZ20]. intricacy [BZ12]. intrinsic [HLW24, RW91]. Invariance [ABDH13, BM12, MP86, BBT16, BPS18, BK92, BT03, Bau16, Ber90a, BB07a, BD14, BD89, BCW21, DNS18, DLP87, Dup10, Ein87, Ers19, FM19b, GVV22, GZ12b, Han02, KLMH06, Kif13a, Kif13b, LPR97, Li19, OS95, Pow19, PRS16, RAS05, Rio94, Sam87, Sch15, SS04, Str64]. Invariant [BFPP19, DV15, Fil00, Jon09, Kal07a, Kal07b, Kal11, Off14, ARV13, Ano87a, AL09, BC22, BB24, BS19, BdT10, BdTR19, Bug85, CMG95, DDG02, DF20, EK87, FW24, Feh17, FW20, GS89, Gar95, Jaw95, Kal00b, Kal14, Kol13, LS22, NP12, OT17, PZ93, Sow92, VA92, Wan18a, vV89]. invariante [GS89]. invasion [DS11]. inverse [CT02, KSvdVvZ16]. inverses [Eva00a, Ino23]. inversion [GY05]. invertibility [BR21, ÜZ07]. invertible [Hof86, ÜZ94b]. Inverting [ST16]. investigating [CL97b]. involving [SU86, SV86b]. irregular [BRR11, GD90]. Isaacs [Kry14]. Ising [Ale01, Bod05, BGW13, BdT10, BdTR19, CGN14, CJN21, CIvV03, CM01, CM11a, CGH24, CV12, DMR11, DG24, DSS23, Dom17, EKZ22, FW24, FPW24, For21, Gal22, Gan99, Hig91, Hig93, Hry98, Iof95, Izy17, LLP10, Li19, MS99, Maz99, MMS12, OV19, PV97, Pis96, DM16a]. isometries [Dau23]. isometry [PRT13]. isomorphic [Lem88]. isomorphism [DPR22, Eis94, GM13, Soo17]. Isomorphisms [Mar87]. Isoperimetric [Osa90, SCZ21, BJ20, BG99b, Mil12, Tei16]. isopérimétrie [Mic99]. isoperimetry [Hut24, Mic99]. isoradial [GM14, dT07a]. isotonic [Dur02]. Isotropic [HKR18, WZ89, DCH24, Feh17, GLZ11, Swa04]. itô [Slu91, AP91, BNT05, ERV02, FP00, GKS16, Gyö86, GK96b, Kry10, Kry11, LB95, Lyo10, NZ90, Sal86b, Ust88]. Iterated [EP92, AF12, AGOZ88, Bal86, CG92, DDP86, Ein89b, Ein92, Ein93, Gri88a, Gri88b, Gri87b, Gri91a, KL87b, LS87b, Lio05, Mir11, MR95, MO86, MO87, Mue91a, Sch88a, Sch00, Sch15, Sha92b, Str64, Ver90, Wee92, Wei86a, Wit87b]. Iterates [BGR00]. Ito [KM02, RV96]. IV [MS17a].

Jack [DŚ19, GS15a]. Jack-deformed [DŚ19]. Jackknife [CK20]. Jacobi [Ahn22, ABZ23, BRM24, Col05, FR04, GG19, GGLS20, GZ18, Jia09a, LQ97, Tan99]. Jensen [JP10]. Jeulin [AMP04]. Johnson [BR10, BR08]. join [Gra05]. Joint [BNS14, LWXZ12, DG93c, Kül01, Pfe87, Rog93, Wei88]. Jonction [HS90a]. Jong [DKP22]. Jordan [Che92b]. jump [Bas88, BP18, CK08a, CKK13, CZ16, DV02, DV04, Imk91, KR16b, MRT19, MRT20, MM13b, MMW15, Pic96, Pic10]. jumping [Son04]. jumps [BC11a, Ber93b, BCW21, BRS14, Bor03, CD20, Fen94, GJ18, LDS02, CD21]. junction [HS90a].

Kac [GKS16, BGP95, COP93, CS03, CK04, FI19, GR10, Le 88, Léo22]. Kallianpur [Mör00]. Kanter [MR07a]. Kantorovich [Mou14]. Kantorovitch [FÜ04]. Kaplan [LMW89, Zho88]. Kawasaki [Gia95, dHNOS03]. KCM [HMT20]. Kernel [GKS04, AK16, AL91a, AL91b, BCK21, Blu05, CG92, CK08a, CKS10, CT11b, CMS03, CS12, Cro08, CL91, DHZ19, Dev88, FM19a, FL93, GN08, GN09, HM88a, HM92, HK03, MMF92, MP05, Mur24, Ndu91, SS24, WT04]. kernels [BBG12, Ber93a, BB00, BFQ18, CZ16, DLM03, Fin02, Hen07, Kaj13, Kal07a, Kal07b, Kal14, Nor93, Osa90]. Kesten [Ano21b, DCG21, ESZ09, Gri21]. Khintchine [Kol11]. Kiefer [CK08b]. Killed [GQZ88, ABB11, SV03, Son04]. killing [BPP04, Mad89]. Kinetic [Rez96, BLM11, BKM24, Mon18]. Kinetically [CMRT08, CMRT15, HMT23]. kinetics [BC11b, BC11c]. Kingman [DK17]. Kirkpatrick [Cat96, Tal98c, Tou98]. Kiyoshi [Lyo10]. KMT [DW21]. Knight [BFF18, DR24, Eis90, Eis94, LPW13, ST16, Val91]. Koch [vdB00]. Kolmogorov [Bra86, Cha24, FNS87, Pen90, RXZ20]. Komlós [Bal89, Tra90]. Komlós-Révész-theorem [Tra90]. KPP [CR88]. KPZ [Aru15, BGS16, BKWW23, CH16, DGRZ20, GK24, LLT23, Yan22]. Kraichnan [GM05c]. Kramers [LS22]. Kullback [Hal90a]. Kuramochi [Kan00]. Kushner [CL99].

l [Nis97]. labeling [DL23]. labelled [Cha10a]. labyrinth [BCF19b]. lace [FvdH17, vdHdHS98, vdH01, vdHS02]. lack [BG24]. lacunary [Fuk91]. LAD [Bab89, KM93]. LAD-estimators [KM93]. ladder [DG93c, Don95]. Lagrangian [BJT11, Tre15]. Laguerre [FR04]. Lah [KM22]. Lamperti [JY03]. lamplighter [DDMP19]. LAN [Mas92]. Landau [FGM09, Fri89, Fun89, Fun91, FN01, LN05, Nis03]. landscape [BDN17, CHL18, Dau23]. Langevin [AG95a, AG95b, Ass01, BG86b, CV18, LL24]. Laplace [Bol86, Bol87, KL00, Kuw09, Lio05, RS91]. Laplacian [AB11, BR24, GM06, Jia12, LM17]. Laplacians [CT11b, Hor00]. Large [AK08, AN92, AG95a, AG95b, AG97, AC03, Ass08, Bal86, BS88, BGGM12, BC15, Bor03, BDE00, BGP95, CX96, CL04, Com89, CD90, DGPZ02, DGI00, DMRYZ04, DE92, DS88b, Fen94, FW08, FL96b, FN01, GN22, GZ93, Geo94, HV15, JLLV93, KKT10, LS15b, Lip96, Mar10, MS02, Müc94, Num90, Oll87, Oll88, Pes94, PR05, PV97, QSY24, QRV99, Rio00, Sep93, Sep98, SY01, Sow92, Tri95, Yat13, Zha14, AKLP23, ASS19, Bai86, BL12a, BR90, BDN17, BGdH10, BGdH23, BCG88, Bre08, CT11a, CR95, CH19, CK88, CLDDMG99, CGZ00, CGZ03, CGM10, CM88, DHR21, DG94, DS87, DP96, Dob88, DH96, Don97, DK87, EK96, EK23, EZ22, FLT19, Fil89, FW89, Fin02, FL96c, FS00, FKW19, Gan94, GL07, Gra05, HK01, HK03, HLLS23, HM24]. large [IR06, Iof95, Jia12, Kif95, KV14, Kon90, LR86, LP19, LM89, LS88, LP11c, Léo95, Mal12, MN04, MN87, MW13, MW16, Mou92c, NRS23, O’C98, Pis96, PPZ99, RST15, Roz86, San94a, San94b, Sch96, SV86a, Shi06, Swi90, Ver03, Wat87, Yil11, YBK88, Yos93b, Zhe95, dA00]. large-deviation [Fin02, HK01]. large-deviations [AKLP23]. large-spin [MN04]. large-time [DHR21]. larger [DZ19]. largest [GN22, LLR10, Péc06a, Péc06b, Péc09, YBK88, Yur99]. Last [Dau23, BBS21, CP12, FN15, FW20, GH23, HM07, Joh19, Liu22b]. last-passage [BBS21, HM07, Joh19]. late [MS17b]. latent [BBN20]. lattice [AKRT97, Ale98, CM01, CK19, CLM94, DMR11, DL23, FM03, FW89, Fri89, GT05b, HP20, KW04, Kül01, LW20, Maj92, Qua99, Sep93, Sta17, SU93, Yos99, Zha18b]. lattice-gas [CM01]. lattices [CS04, Häg03, IMŞK93, Shi03]. Law [EZ22, Rid04, Rio00, Sch15, Shi06, AL23, AF12, AGOZ88, AG97, ASS19, AT86, Bal86, BT96, BL12a, Bas88, Ber87b, BD89, BCC12, BYY14a, BYY14b, BHR01, Bru13, BM11b, CMRT15, CPY94, Cha92, CG92, CC94, CLDDMG99, CP03, CM88, DDP86, Dob88, DM05, DS88a, Ein88, Ein89b, Ein92, Ein93, Fil89, FW89, GT03, GT16, Gri88a, Gri88b, Gri87b, Gri91a, HS13, HNY99, Hut21, JMP89, JM90, Kar12, KZ17, Kry03, KL87b, LR86, LP19, LS87b, LS18, LLP10, Maj92, MM90b, MSW21, MO86, MO87, Rod17, RS93, Rog93, San94a, San94b, Sch88a, Sch00, Sha92b, Str64, Wee92, Wei86a, Wit87b, Yin14, dHHKR91]. Lawler [CM91]. Laws [LM89, Ber95b, BD87, BDG10, CFV23, CD13a, Che00, Chr86, DV15, Deh86, DS87, Dub02, FKL04, FNS97, FW04, FW07, GR19, Gou04, Has88, HS94, Hor90, Jan88, Jan10, KR16b, Kho92, KY17, Lac90, LPR97, Mar10, MN87, Mir11, MP91, MR95, Mör00, Mue91a, Ogu01, Oua22, RST15, Röm96, Sch89a, Szu92, Ver90]. layers [CQ23]. LDP [AJV18]. Learning [Men18, CLL20, MST14]. least [HM87, Oli16]. least-squares [HM87]. leaves [Alt89]. Lebesgue [Bel08, Yan09]. Legendre [Cat96]. Leibler [Hal90a]. lemma [AP91, JZ10, Yu93, Zie88]. length [DK17, DZ19, Gri87a, Gru98, Har90, PWW11]. Lengths [BP15, BD87, Lah01, TW01]. lent [BD11]. letter [BGdH10, BGdH23]. Level [Hig91, Alt89, Aru15, Bor91, Bru19, CN20, DW93a, Hsi93, Slu91, Xia97]. level-crossing-count [Slu91]. level-regular [Alt89]. level-sets [CN20]. levels [GZ18, Tal00a]. Levitov [Sil17]. Lévy [AD08, ADH14, AE14, And08, BR24, BNT05, BFF18, Ber86b, Ber93b, BGR15, BDW21, BH09, BMS08, BM11b, CKS11, Che18, CMS05, DM05, DR13, DR16, DSS14, DL05, DW07, Dur09, DJ12, Eva87, Eva89b, Eva00a, Fag91, FHS90, Fou98, Jaf99, Jur88, KSX08, KX09, KSX09, Kol11, Luc91, Mal95, NRST16, RRO24, Sav09, Sim99, Sto86, Wee90, Wee92, Yan09]. lié [Eis90, Ale02, BL91, BLV12, Def16, Hou19, Rai97b]. life [CKM93]. lifetime [DeB87b, KP89, Xu91]. Lifschitz [Erd98, Erd01, PP91]. liftings [Tal88a]. like [Bar24, CKS10, Dri13, HW13a, HW13b, KZ98, MMS12]. Likelihood [HR02, CM08, Höp93, HR95, Mas92, Nic07, Nic08, RS17, SCC19, Yos92]. Likelihood-based [HR02, RS17]. LIL [EM93, Sav09]. Lim [Li92, Gri91a, Gri95]. Limit [ABM05, Bar02, BCG14b, BD87, CG94, CQ23, Che95, CP03, DH91, DG14, Fuk94, GS07, GT97, GTT98, GLZ11, Gro88, HHW18, HR91, HTV89, Jan06, Jan10, KTT89, Lac90, LT88, LRW94, Mat94, NOT22, Pes02, Pet92, Szu92, WY93, Zha18b, ABBG12, AdR17, ANV15, AB10, AS88, Ale87a, AGOZ88, ACS21, Ans99, AL09, AGL17, ABBN04, Bah98, Bar86, BLM11, BCF19b, BLV16, BV95, BLQ23, BL24b, Ber95a, BHS10, BS05, BSW17, BG07b, BCG14a, Bob18, BKM24, BMP97, Bon20, Boy19, Bra89, Bre05, BDW21, BMS08, CG09, Car05, CS89, CD13b, CG07a, Che00, CS08b, CS09, CP88, CES24, CHR18, CC94, CFR93, CH88b, CS94, DS18, Ded98, Ded10, Ded13, DG98, DL01, DL03b, DG22, DKW23, DJPP13, Dun24, FM03, Fag22]. limit [FT87, Fra05, Fri89, Fuk91, Fuk11, Fun91, Fun95, Fun03, GR10, GN08, GS90, Gol07, Gou04, Gra05, GLP17, GLL20, GKR07, Häg05, HW86, HHT21, Hil87a, HLLS23, Hor00, Hor90, HLW24, Hou19, HL19, JR08, JW24, JB04, Jur88, Kar08, KM86, KZ97, KP86b, Kif10, Kol10, KM00, Kot88, Kou89, KL87a, Lac16, LLT23, Lee90, LS18, Lev88, LS86, LST20, MW89, MN98a, Man91, MMŚ21, Mar08, MN04, Mik04, MM13b, MW14, MW16, MMS12, Mor87, NP22, Nie89, Nis03, Nis97, Oel89, PPZ14, Pel12, Pen99, Pen96, PZ08, PV97, Por86, RR92, Rai98, Rei05, Ros87, San94a, San94b, Sch86, Sch88b, Shi06, SU93, Swa00, Swi90, Tho14, Uch89, Wac12, Wat87, Wen96, Wen99, Wen04, Wit87a, Yas02, YBK88, Yuk87, Zes88, dJ87]. limit [dHHKR91, vW97, Bou88]. limite [Bou88, Bre05, LT88, Pet92, WY93]. Limited [BB07b, BPP97, BW18, CRS18, Luc14]. Limiting [BY86, Bai86, Le96, Le99, Mou90, And08, Jia09a, LV86, Luc14, Pop09, Szn89]. Limits [BV13, BDW21, Li16, Ric18, ADP16, BJMW11, Ben10, BG86b, BFO09, Bor20, BO09, BCF18, BRM24, CY15, CFM20, DGZ05, DR24, FLT22, GW86, GVV89, HS21, HS20, HTV12, Kol03, KK10, Las19, LM19a, MiS09, NST23, Off14, PS18, RS12, Rez96, Sam87, Tho01, Uch94, Wei93, vdHdHK03]. line [Ban16, BJP22, BG24, CH16, DR86, FW20, Fuk04, HR91, Hry98, KKT10, Kuw09, Léa01, LMU03, Naj18, RSX13, SS13]. linéaires [Bou88]. Linear [BGGM13, Buc91b, BN94, Gam04, Gol08, Hou90b, OP89, ADGO17, Bab89, Bai86, BM12, BMN12, BSS14, Ber20, Ber21a, Ble89, Bou88, Boy19, CMS07, CMS23, CN88, CMM90, CT86, CG07b, CIK99, CK08b, DKN09, DGT12, EL94, FZ99, FK10, FO14, GS90, GP00, GN98, GM05c, GP93, HM97, HMY02, KKM99, KP86b, Kol11, Kou89, KM93, LDS02, MY99, Mil89, Swa04, WHZY07, WZ99, Yat13, vW97]. linearization [Ans99]. linearly [MR07b, Mor92, Mor87, Swa00]. lines [Aru15, BGGS18, EK23, HR02]. link [WHZY07]. Liouville [AK16, BCF19a, BHSY23, DZ19, DGS21, DF20, DFS20, GHS21, GHS19, GPS22, KT07, MQ20, MS21, MSW21]. Lipschitz [AE14, AW15a, BBB90, BY05, BS14a, BSS18, Cer03, Cha24, CLL20, DeB87b, FT96, Hau07, Kol11, Mij88a, Mir11, Tru03]. Lipschitzian [FZ05]. live [Kwa24]. Local [AL23, BBS21, BB90, BK92, Bor20, BYY14a, CL19, DGPS07, Don12, Duf92, FK16, GW86, HJL90, Kho92, KM00, Las19, Leb17, LS18, Lus92b, Nol89, QR13, Rio94, VW09, Wac12, XZ02, AMP04, AZ20, AGOZ88, ACS21, AHJ21, AC07, Ass08, AFZ87, BMRS20, BBB90, BK12, BNP11, Ber89, BW93, Bor86a, Bor86b, BYY14b, Bou90, Bre05, CG09, Car05, CS03, CL04, CK04, CFR93, CCFS00, DS95, DG22, Don97, EM97a, Eis90, ELY99, Eva00a, Föl89, FR95, GR19, GLL20, GdH92, GdH94, HM97, Höp93, Hor90, Hou19, Hry98, HN09, IW94, Kal03, Kal09, Kar12, KY17, Lac90, LZ10, Le 15, LS87a, Leu10, MW89, Mam87, Mat95, MU12, NP22, NRW87, NW91, Pal12, Pit96, RRO24, RXZ20, RW91, Roh11, Sug94, SV86b, Tei16, Val91, VW24, War05]. local [Wee92, Xia97, Yin14, Ber90c]. localisation [CM17]. localisées [Wei86b]. Localization [BSS14, BG13a, CN16, Day88, DL20, TY03, AGKW09, CV10, Imk86, Var07]. localized [Klo13, Wei86b]. Locally [Dun07, MW90, Chi12, Dob88, Fel05, FM97, Jaw95, LB08, MMS12, PZ93, Rho09]. Location [Bra86, CFR87, JR88, Lus92a, Mar91, Mat92, NM89, Par90]. location-adaptive [NM89]. location-scale [Par90]. locaux [Ber89, Ber90c, Eis90, Leu10, Mat95, Val91]. locus [CKM93]. Loewner [HL21, MQ20, Zha04]. log [ASZ09, BJ14, BCD21, BGP24, CN16, DZ18, Höp93, LP19, SS15, TY23, Yos99]. log-concave [ASZ09, BGP24, Har04]. log-correlated [DZ18, LP19]. log-gamma [BCD21, CN16]. log-infinitely [BJ14]. log-likelihood [Höp93]. log-Sobolev [TY23, Yos99]. logarithm [AF12, AGOZ88, Bal86, CG92, DDP86, Ein89b, Ein92, Ein93, Gri88a, Gri88b, Gri87b, Gri91a, KL87b, LS87b, MR95, MO86, MO87, Mue91a, Sch88a, Sch00, Sch15, Sha92b, Str64, Ver90, Wee92, Wei86a, Wit87b]. Logarithmic [Wan97a, Yau97, AL00, CLM^{+}12, CGG07, CGW10, Ces01, CV18, Con24, DS09, GGM05, LN05, Wan97b, Wu00]. logistic [LPW13, SCC19]. Lognormal [ARV13]. loi [DS88a, JMP89]. Lois [Rio00, Has88]. Long [CV98, EK23, GM05c, Mue91b, Mue98, Sta02, AM08, BGS16, BG99a, BGP14, BCW21, CS08b, CS09, CG90, CS12, CK08b, Dit90, FT87, GJ18, GSW86, Hut21, KM93, LS19, Mön24, Mou92a, MT95, Rob94, Swa00, TT91]. long-range [AM08, BGS16, BCW21, CS08b, CS09, CK08b, FT87, Hut21, Mön24, TT91]. Long-time [CV98, Mue98, Sta02, Dit90, Swa00]. longest [AD95, Gri87a, TW01]. loop [BLV16, CS16, CLW11, Ebe97, HK11, KW16, LW04a, MWW15, MSW21, Sch09, Sri00, Zha21, vdBCL16]. loop-erased [BLV16, Sch09]. loops [FL96a, Hab18, SS18]. looptrees [CK15b]. Lorentz [BBT23, Qua99]. loss [Bar21, Bir14, CLL20, Don94, Hal90a, LS89, Men18, TW20]. lot [Wer94b].

Lotka [CP07b]. Low [Jia12, CMT15, GT98a, GG96, LM19a, DM16a, Tal00b, Zyg09]. low-rank [LM19a]. low-temperature [Tal00b]. Lower [AG14, CT11a, HM91, KH03, Lou97, UA03, Wee88, Wee90, Bre08, Lej01, Mij88a, Oli16, Por95, Szn11, CSC93]. lower-order [Lej01]. lozenge [AG22, Hua24, Las19, Pet14]. LQG [DFS20]. Lyapounov [Zyg09]. Lyapunov [Bax89, Cha24, Cra00, CMS05, EdHM15, Key87, KMZ11, Lia00, Sta17, Wan01].

M. [Le 88]. MA [Ino08]. Macdonald [BC14]. Machlup [Cap95, CMN92, MN02, WZ93]. Macroscopic [KK10, OS13, Var07]. magnetic [Erd98, Erd01, GKS16, Pra93]. magnétique [Pra93]. magnetization [CGN14, CX96]. magnitude [KN98]. Major [NN20]. Majoration [Léa87b]. majority [DF16]. Majorization [Gor92]. makes [CHM86]. Malliavin [AJJ22, DNS90, Flo90, Imk96, KY00, NZ86, Sch23, Ust88, Yos92, Yos97]. Mallows [BP15]. Mandelbrot [Bar99, BPW09, CCD88, CCGS91]. Manifold [Zha07, BG17, Box89, Li03]. manifolds [AKTW18, Arn97, Bla05, BM15, CKP14, Cra93, Div22, Fil00, HLW24, Hsu90, KNP12, KZ98, LB95, Le96, LB08, LM01, Li94b, Li94a, Li08, Li11, Li16, LL24, Ogu01, Wan97a, Li14a, Li14b]. many [AK08, BL12a, BG13b, GL99, HL18, KM12a, LW90, Mło99, Szn89, Tal98b, Tan96, Tan97]. map [Dup10, Le 18, MS24a, MS21, Tro21]. Mapping [PS22, CL97b]. mappings [AMP05, Ers09, FQ88]. maps [AD02a, Ano87a, BGR00, Bar02, BBCK18, Bug85, Bug91, Cha10a, Cha19a, CC94, CMS02, Ded10, Ded13, GW04, GLT10, Gou04, GH20, GHM20, GHS20, Lov95, Mir11, Rai03, Ric18, Sch15, Ter16, Tha96]. marche [KPyWmW88]. marges [HT86]. Marginal [LRW23, BCH20, COT19, Dud94, Gyö86, RR95]. marginally [Lac10]. marginals [vW97]. margins [HT86]. marked [GZ93, Kal90]. markets [Kra96]. Markov [Ano87a, CD21, Uma06, AD02a, Alh04, AS89, ABBP02, BL12a, BP89, Bas88, BKU10, Ber87a, BC04, Ber97, BK93, Bla88, BMP97, BCC12, BCS19, Bor88b, BG13b, Bou88, BV22, BKWW23, Bug85, Bug91, CD20, CR90, CR93, Che00, CFS01, CF08, CFKZ08, Che95, CSC93, Dar88, DLM03, DL01, DL03b, DH23, Duf92, DE92, Dyn93a, DK95, DK04, EM13, EP97, FR03, Fen94, FM86, Fit87, Fit90, FR08, Ful09, GY05, GOZ21, GL14a, GT98a, GLL20, GVA86, Gre87a, Gre87b, GK01, GLWY09, GT05b, HS21, Häg05, HS90a, HS95a, HS95b, HS94, Höp88, HJL90, Höp93, Hor90, IR06, JY03, Jen91, Jou87, KR16b, Kel87, Kes88, Kol11, Kol87, KM00, KM09, KM12b, Le 93a, LR94, LM00, Lev88, Lov95, Mic99, Mic08, MW14, MR16]. Markov [Nag87, NOT22, NW90, OTZ24, OS95, Pel12, RRO24, Ros89, SS87b, Sch07, Shi90, Shi94a, Shi94b, Stu91, Tho88, Von95, VA92, Wit86, WZ89, Wu04, vH11]. Markovian [BK17a, BB03, CG04, KMP22, Oll87, Uch88]. markovien [HS90a]. markoviens [Bou88, Jou87]. Martin [CK04, Cra93, DG21b, IR10, Kan00]. Martingale [BHW19, DS14, EM90, FG95, GMW24, Kyp00, Li08, Str86a, Wei93, Wei90, Asp89, BJP22, BL15, BJL17, BK93, BHR01, DHR21, Fag90a, FS86, GPW13, Hin13, JZ15, McG86, MZ87, MU12, Mou14, RST15, Rog93, Tho01, WZ87, Zam00, Li14a]. Martingale-driven [GMW24]. Martingales [BBCK18, Arn97, AB07, BGM18, Bar99, BT03, BC12, Bla88, CSK99, Cla87, Dar92, DDP86, DS95, ELY99, Hit90, Imk91, Jia09b, KS91, Kaz88, LR86, Le99, Lin86, MU12, MP91, Myk95, Ose08, PX88, Rog87, Tei88, Yos97]. Mass [BKdS18, FM97, Hil87b, Kot95, Zha20, dBM13]. mass- [Zha20]. massive [Rod17]. matching [AST19, CM24, DMX21, DWXY23, HMO23, MRT23]. Matchings [BLS13, Hol11, HJW22]. mathematicians [She07, Tal98c]. mating [GHS20]. mating-of-trees [GHS20]. matrices [Agg19, AEK17, AEK19, Au22, ABZ23, Avr88, BY86, BE17a, BZ20, BR21, BG09, BGGM12, BWW18, BDN17, BV13, BKYY16, BCC12, BG13a, BJ02, BYY14a, BD20, Bry08, CMS15, CRZ13, CMW15, CT18, CL19, CES21, CES23, CES24, Coo17, CP00, DMB20, EYY12, FM19a, FS88, FI19, Fug88, GG96, GLP17, HKR18, HK20, Ino23, Jia05, Jia09a, Jia12, Joh02, Kar12, KY13, KY17, LP19, LLS24, LL04, LS16, LS18, LV12, LT20, MMS14, Mal12, MP02, MM13a, Muk91, Muk93, NST24, NS23b, NO15, NTV17, O’C21, Osa12, Osa22, Par22, Péc06a, Péc06b, PR05, PRT13, Rai98, Sch05, Shc13, Tao13b, Tao13c, Ver11, Wat87, Wie02, YY21]. matricial [BG07a]. Matrix [ABO23, Ahn23, AEK19, AZ06, BGM18, Ben10, BLZ20, BCGZ22, Cha92, DG21a, FS88, FR04, FR05, GR19, GM05b, HPU04, KLT17, LP11c, LM19a, LLR10, LLT^{+}19, Osa98, Péc09, PS13, Pop09, RR17, Tao13a, YBK88]. matrix-valued [Cha92]. matter [GMM17]. Max [GHV90]. Max-infinitely [GHV90]. max-stable [GHV90]. Maxima [Jia05, CCK15, KLZ23, Leu10, Pan90, dHHKR91]. Maximal [BH09, BK17a, CT11a, EM97b, HP18, Key87, LS15a, LV86, LS89, NZ17, NZ21]. maximale [HS90a]. maximization [Nut12]. maximizers [BZ12]. Maximizing [LM22]. Maximum [DGPS07, Sep95, AW01, Ber87c, BHR01, CFR87, CM08, Din13, GZ93, HR95, LP19, Mas92, MM13b, MS90, Nic07, Nic08, Rog93, Wac12, Yos92, Zha18b, HS90a]. McConnell [Bañ87]. McKean [BR24, BJT11, CM18, DH22, DR14b, HLLS23, Kot95, KK10]. McMillan [Bre08]. MD [KM93]. MD- [KM93]. Mean [Har90, JRS20, Wan01, AA88, Aug21, BCL20, BM17a, BC04, BGP14, BDS93, BEGK01, CD01, DHR21, Don97, Duf92, Eld20, Fen94, FM90, Fle12, FR95, Fuk91, HK92, HKR18, HRvR17, HD21, JLS09, KLZ23, Lac16, LLP10, LM19b, Mac21, MS02, Mül89, NT86, PT04, Pan15, PSS15, Sch86, Swi90, Tal00b, Tal06, Tal09, Tal10, Tal16, dlR95]. Mean-field [Har90, JRS20, Aug21, BM17a, BGP14, BEGK01, CD01, Duf92, Eld20, HKR18, KLZ23, LLP10, Mac21, MS02, Sch86, Tal16]. mean-fields [PT04]. mean-zero [JLS09]. meander [PR12]. meanders [KT07]. means [CM88, GT98a, HLW24]. Measurability [FGM09]. measurable [Dud94, Eba86, LW90]. Measure [Div22, FÜ04, Kol03, Mar98, Per92, AL94, AD02b, AS11, ASZ09, And86, BB09, BT03, BFPP19, Ble89, BKWW23, CK08a, CKK13, CL97c, Dar88, DM22a, DV15, DG14, DHK22, DD17, DKS94, Eig86, EK23, Eld20, Feh17, FW20, FG15, FS86, FT92, GS89, GM12, GRW92, GPW09, GD90, HH23, Har04, Hau07, IMS20, Kal00a, KS89, KS88a, KLL94b, LS22, Mas88, Mél01, MS09, PT87, Rei89b, RT86, She93, Sow92, Tzv10, ÜZ92, ÜZ94b, Wan97c, Xia97, Yan09]. Measure-valued [Kol03, Per92, CL97c, Dar88, DKS94, KS88a, Wan97c]. measured [Isc86]. measured-valued [Isc86]. measurement [DLP87, Hos91a]. measurements [Wei86b]. Measures [BGV07, Eba86, FI07, ÜZ96, ARV13, Ano87a, Arb91, AL09, AC15, BC22, BL24a, Bel86, BS05, BB00, Bob10, BSV24, BR01, BDS93, Bor08, BGP24, BFQ18, Bug85, BDG^{+}09, CP12, CMG95, CMS02, Com89, CK15a, DDG02, Deu87, Dow88, DK04, Dzi91, EM13, ES03, EM90, Eva90, FM12a, Fer94, FM86, GG11, Grü86, Hil87b, Hin05, Jia09b, KLS86, Kal07a, Kal07b, Kar87, Kee88, Kes88, Kol13, Kül01, Kül03, LS88, LMV98, Luc91, MM13a, NRST16, OT17, OS16, OS22, PZ93, Rob10, SV99, ST88, Shi94b, Sri00, Tha91, Tho01, Thu86, Wan18a, Wat00, Zäh86, Zäh88, Zha21, vV89]. mechanical [BS95a, BS97b]. media [BRR11, Ber01a, CGM08, FKW19, KL93, Lej01, Owh03, Sei00]. median [DS20, Swa07]. medium [BCFN89, DF91, EF98, GdH14, GdH94, GKW01, Mat94, Wan97c]. mediums [KPyWmW88]. Mehl [BR10, BR08]. Mehler [BRS96, LPS16]. Meier [LMW89, Zho88]. Meixner [BW05]. mélangeante [DLP87]. Mellin [FT04]. memory [BGP94, NN93, TW20]. memoryless [Mar87]. mesh [LST20]. Mesoscopic [Bou10, CES24, HK20, HL19]. message [DM16b, Duf92]. message-switching [Duf92]. Mesure [CR86a, AL94, DLP87, GS89]. mesures [Fer94, Wei86b]. Metastability [AGL17, BL12b, BGW13, BEGK01, Dom17, BL15, CLM^{+}12, DCH24, Hol03, Kim21, Kim23, LLP10, Mic22]. metastable [LS22, Mon18]. Meteor [Bur15]. method [BM19, Bar90, Bau13, Bon20, BD11, Cat98, Cha07, Cha10b, CS01, DJP01, FSX19a, FSX19b, Gaj89, GM05b, Lah99, Lio05, MR23, NP09, RR03, Sch88b, Yon97]. methods [EM13, GT24, Kot92, Lah01, Rei89a, Wan94a]. Metric [Dud94, ACLP20, BP20, CK08a, CKK13, DM22a, DF20, GPW09, GPS22, LW18, RR92]. metrics [DFS20, GR10]. Metropolis [Cat98, CW17, Gay19]. Meyers [Pen99]. Microscopic [FG88, Leb17]. migration [Neu90]. Mimicking [Gyö86]. min [dHP86]. min-stable [dHP86]. Minimal [BM07, GRS24, HJW22, ABS21, AS92, BS24a, Cla87, DH91, DG24, Dow88, Tay89, vH11]. Minimax [Chi12, DJ94, DJ00, Don94, GP05, GL14b, GT98b, Par91]. minimaxity [WZ13]. minimises [HM87]. minimization [LM09]. minimize [LS89]. minimizers [BJ20]. minimum [Gaj89]. Minkowski [Bor08, Zha19]. minor [ACJvM15]. minorant [PR12]. minorants [AE14]. Minoration [Liu22a]. Minorations [CSC93]. Mises [Cas12, DG14]. mismatched [Iha90]. mixed [ABXY24, CK08a, CP13, CDP17, CHL18, HJL90, JT17, Liu22a, Lus92b, PT02, RSX13, SRMC93]. mixes [CP00]. Mixing [CMRT15, EV21, FR07, LS24, PS13, PSS20, Ter16, Ale98, BBPS21, BM03, Ber87b, BCT12, Bra87, CS20, DLP87, EKZ22, Gri87a, GVV22, Hai02, HP18, HvdH11, HvdH19, Hig86, KY00, MP05, Oda08, OTZ24, PSS15, Ram98, Reg23, RST^{+}13, Sam87, Sch89b, SSSY17, Yos04]. mixtures [Kes88]. Mobile [PSSS13]. Mod [BKN14]. Mod-discrete [BKN14]. mode [FL21]. Model [Bar00a, BM07, Bir14, Lou97, TA96, WZ13, Ale01, AZ06, ABDH13, ADS16, ACM06, BBT16, BBM99, BM17a, BDC12, BS22, BGL19, BL24b, BBF12, BBC^{+}04, BKdS18, Bod05, BGW13, BMP97, BGJ96, BP18, BdT10, BdTR19, BGP94, BGP95, BG97b, BCG88, BS95b, CJN21, CMRT15, CMT15, CSZ16, CM95, Cat96, CM11a, CM13, CW17, CHL18, CHS22, CV12, CDCIV14, CMS05, CGM10, CM17, DDD19, DMR11, DS97, DD05, DNS18, DL23, Dit88, Dom17, DCMT21, DM91b, EL04, EdHM15, FLT19, FM90, For21, Fri89, FR95, Fun89, FN01, Fun03, FM21, Gal22, GM98, GdH99, GKM00, Gei05, GL99, GRAS17a, GRAS17b, GJMM24, GKZ93, Häg96, Ham15, HMT23, Hig91, HS20, HMM19, HKK23, Hry98, HNY08, Iof95, Jac90, JR08, JLLV93, KK14, KW04, Kif97]. model [KSvdVvZ16, Kol10, Kor22, Kül98, Lac10, LRV23, LO24, LLP10, Li19, MS99, MM15, MT20, MRT19, MRT20, MNS15, MPS18, Myt98, Nis03, NRW87, OV19, PV97, Qua99, RD13, Rom14, DM16a, SU86, SZ17, Tal98b, Tal98c, Tal00b, Tal01b, Tal06, Tal16, Tou98, Tsa16, WHZY07, WZ99, Zam04, Zha14, vdBB04, IMS20]. modell [Uch89]. modelled [Dey16]. modelling [Eic12]. models [ABXY24, AK87a, Ale98, Ale01, Bab89, BCW21, BR08, BJT11, BEGK01, BGS13a, BS24b, CIvV03, CMRT08, CQ20, CMS07, CMS23, CP12, CP13, CDP17, CD01, CP07b, DIT14, DG93b, DH22, DJ00, DGI00, DGZ05, DR24, DM23, DSS23, DCM16, EKZ22, FN15, FG19, GLT10, GN98, GMM17, Häg03, HK11, HJM20, HP20, HJL90, Höp93, HN97, IPS13, JR88, KS95, Kou89, KM93, Kül01, LV96, LS01a, LR21, LT21, Lus92a, Man91, Mar91, MU90, MPV01, MMS12, Neu90, PT04, Pan15, Par90, Par91, Pis96, Pop09, PRS16, DM16a, Sha92a, Sub23, Tal09, Tal10, dT07a, BR10]. Moderate [BW17b, EL04, AS21, CL04, CP03, ES03, MPR11]. moderately [Oel87, Oel89]. Modified [GGM05, DF16, Mou92c, Wu00]. moduli [Ale87b]. modulus [GNSS95]. moment [BC12, Ben21, FT19, JP91, Li94a, Tei88, Wit87b]. Moments [Bar99, Agg19, BLM24, BK12, Bry88, BW05, GW04, GS24, GTS84, GTS87, GS04, GHK22, GHK23, Hit88, KMM98, Lin87, Mat92, Bar99]. Monge [FÜ04, Mik04]. Monotone [HK11, Ald88, BP15, HMO23, KR16b, Nac03, Soo17, TW01, ÜZ96]. Monotonic [Pen99]. Monotonicity [HP99, vdHH10, Bon90, HP91]. Monte [DJP01, EM13, Mél01]. mosaics [Gol96, HS11b]. mosaïques [Gol96]. Most [AB19, BES00, Eis97, Erh00, Hal88, HS00]. Motion [HD21, AL94, AD02b, AE14, ABBL24, AH20, AL23, Ahn23, ABBS13, ABK13, AC03, BK04, BR92, BS97a, BPP04, BP88, BB90, BB95, BH00, BH07, BBS14, Ber89, Ber00, BS05, BS98b, BM05a, Bor91, BL01, BN02, BM05b, BS14b, BNS14, CPY94, CD99, CR88, CGS23, CMS00, CDK18, CKM93, Dav99, DL03a, DeB87a, DeB87b, DT93, DeB12, Del99, DF01, DL97, DM03, Dri13, DHK22, DP99, DvZ04, EF98, Eva00b, FL95, FM97, FR92, FP00, GHM23, Gan94, GVV93, Gro89, Ham92, HS13, HS95c, HL19, KP89, KMP22, Kre95, Kry03, Kum93, Kwo92, Kwo95, LS97a, Le96, LY86, LY87, Le 87, LM92, Le 15, LQ13, Leu10, Mai16, Mal95, McG89, MW95, MN02, Mör00, Mör01, Mou90, MP19, NRW87, Ove93, Ove94, PPS96, PS16]. motion [PT87, PT00, PR12, Pov99, RW91, SV99, Sal90, SS18, Sch96, Ser95, Shi92, SV03, Sto86, Szn93b, Tal98a, Tha96, TTV03, TW98, Tri91, Von92, Wan03, Wer94a, Wer94b, Wil87, Wüt98, Xu91, dlR95, vdBS90, vdB94, DeB88]. motions [BCR04, Ber90c, CP10, Che92b, CQ02, Def16, DD17, FW20, GS15a, HH23, Has89, Mat95, PW97, RS12, RW88, Swa07, TW93, RW89]. Motoo [Sch88b]. Mouvement [Le 87, dlR95, AL94, Ber89, LY87, LM92, Leu10, Mal95, Wer94b]. mouvements [Ber90c, LY86, Mat95]. move [BP89]. movement [Cro13]. moving [CH93, DS23, HH90, PT02, SRMC93]. moyen [dlR95]. moyennables [Rau88]. Moyennes [KPyWmW88]. Mueller [Kry97]. Multi [Ble89, HT10, JY03, Lia22, RSZ23, Zer05, BCH20, Ble93, COT19, HS21, HLLS23, KLP01, Man86, Sai87]. multi-armed [Man86]. Multi-dimensional [HT10, Ble93, Sai87]. Multi-excited [Zer05]. multi-index [KLP01]. Multi-linear [Ble89]. multi-marginal [BCH20, COT19]. multi-parameter [Man86]. Multi-point [Lia22]. multi-scale [HLLS23]. Multi-self-similar [JY03]. multi-type [HS21]. Multidimensional [Kif95, BG97a, BDG^{+}09, BDG10, CW93, DR07, Ein87, FP00, Ker89, Kol13, Mir13, Pal12, Sim11, Str16, Stu93, SU93, SV86b, Tou98, Zai87, Zai88]. Multifractal [BM02, DJ12, Bar99, GH10, Jaf99, Sch20]. multifractale [Bar99]. Multilevel [DG94, GS15a]. multilinear [KS88d, MT87, NO15]. multimodality [MMF92]. Multinomial [LS89]. Multiparameter [WZ87, FS86, HS95a, Tal98a]. Multiple [DK99b, DG93b, DG93a, Eva87, NZ90, Slu91, Tal98a, Tal00a, ABR08, Ble89, DS11, DK99a, FPW24, KS88b, KS88c, Maj88a, MO86, MO87, Roh11]. multiple-armed [DS11]. multiples [AB22]. Multiplicative [PW20, AJJ22, BLM24, BWW18, BBW14, BvdHS18, BM16, BDW21, Cer03, Chi12, DKN09, DLTV24, DHK22, Fuk91, GGLS20, HW13a, HW13b, HH23, Jeg21, LQ98, Sal90, Shi94b, Won20]. multiplicities [HM24]. multiplier [CK20]. multiply [Izy17, Thu86]. multiply-connected [Izy17]. Multiscale [BZ07, QSY24]. multitype [FV00, GRW92]. Multivariate [FSX19a, FSX19b, Bon20, CP93, DKP22, Eic12, Kal95, LZ24, Men16, MW14, MS90, Pou88, QD91, RR92, TA96, WT04]. mutation [SU86]. mutations [CMMM11]. mutual [Sun98]. mutually [KM12a, Myt98, Tan96]. myopic [HTV12].

N [Zai87, WT04]. NAE [SSZ22]. NAE-SAT [SSZ22]. nary [Kol03].

Natural [Wit86, BLS87, Mit86]. naturally [Rol03]. nature [Jaf99]. Navier [Fon09, BBPS21, BFHM19, BCF15, CC99, DR14a, FG95, FR08, FL21, LS97b, Mél01, MR12, Oss05, RZ09, Zha10]. Near [LM19b, vdBN21, AB11, CM11a, Hsi93, NZ21, Yao19, Zha18a]. Near-critical [vdBN21, NZ21, Yao19]. Near-optimal [LM19b]. Nearby [MP91, Röm96]. nearest [CR86b, LM24, Lig87, Mou92a]. nearly [SY01]. Necessary [LLR10]. needlet [KNP12]. negative [AUY08, Bor91, CKS11, MZZ24]. neighbor [LM24]. neighborhood [BC15]. neighborhoods [RR03]. neighbour [CR86b]. neighbour-estimator [CR86b]. Nelson [Lim99]. NESS [GJMM24]. nested [CH13, GI20, KW16, Kum93, KK96, LS20, MiS09]. nesting [DCL19, MWW15]. nets [CM88]. network [Gra05]. networks [BBM00, Bak11, BB21, BBH24, Cha16, Duf92, GN22, GR19, GV16, LMU03, MST14, MS02, Tel89]. Neumann [AD04, CWZ95, FW12, PZ98, Wan94a]. neural [MS02]. neutral [Szn99]. Neyman [JZ10]. nice [Maj06]. Nikolaevitsch [Pen90]. nilpotent [AG23, Hou19]. no [Ber93b, HMO23, Pap87]. nodal [BW18]. nodes [Mie05]. Noise [BLZ20, Mat94, PR20, AHR01, BC22, BR24, BG02a, BH13, Bié98, BEN23, BDE00, BCF15, CT02, Cer03, Cer05, CHM22, CV10, CHKN18, CV98, CMS05, DKN09, Day88, DR14a, DLTV24, DMP93, FGS17, FL19, FL21, GG19, GP00, HW13b, HJM20, HLP24, HLØ^{+}93, HKK23, HN09, KSvdVvZ16, LLT23, LW00, LO96, Mik04, Mue91b, MP92, MS93, Mue98, Myt02, MP11, NP92, Oua22, Spe90, Tri95, TW20, VW24, dBD02, HW13a]. noises [Coq00, Fag91, LR04]. Noisy [Bak11, GH10, GS04]. nombres [Rio00]. Non [BCP19, BS22, BHS10, BBF12, BGV07, BL90, DV00, DGT12, DZ18, GTT98, GVV89, Gri88b, Hut20a, Joh02, Lac14, LS22, MSW21, Mor86, Nag87, Sch05, Shi03, AW15a, Aff88, And86, Ans13, AG97, AP91, Bau16, Ben10, BD14, BM11b, CN88, CM02, CGM08, Cer03, Cha24, CMM90, CT86, CDR86, CS01, CS03, CK04, CZ16, CES21, CES24, DDG02, DKN09, DM17, DF16, Dud94, DGL87, Eig86, EMS13, Fab90, FZ05, FvdH17, FL93, FS17, GL14a, GM05c, GMM17, Hau07, Hil87b, HLP24, Jan95, KP89, KW04, KM12b, Kwo92, LR21, LS88, Lou22, LT20, MMS14, MR13, MT20, Mou92a, Mou92c, Oda08, OS95, PV20, Par22, Pel12, RXZ20, Sow92, Thi91, Tru03, ÜZ94a, ÜZ94b, Wei93, Wer97, WZ93, Yan22, dA00]. non-additive [Oda08]. non-anticipatory [AP91]. non-atomic [Hil87b, Wei93]. non-attractive [Mou92a]. non-backtracking [FvdH17]. non-ballistic [Bau16]. non-Brownian [EMS13]. Non-central [BHS10, GTT98]. Non-classical [Gri88b]. Non-coincidence [Lac14]. Non-commutative [GVV89, Sch05, Ans13, AG97, Par22]. Non-compact [BGV07, LR21]. non-constant [FS17, Kwo92]. non-convex [MMS14, dA00]. non-elliptic [BD14]. Non-existence [Shi03]. Non-extinction [BBF12]. Non-Gaussian [DV00, Sow92]. non-Gibbsian [LS88]. non-Hermitian [Ben10, CES21, CES24, LT20]. non-homogeneous [CM02, Pel12]. non-integrable [MT20]. non-internally [Mou92c]. Non-intersecting [Joh02, Aff88, KW04]. Non-intersection [BL90, Hut20a, Wer97]. non-invertible [ÜZ94b]. Non-linear [DGT12, CN88, CMM90, CT86, DKN09, GM05c]. non-Lipschitz [AW15a, Cer03, Cha24]. non-Lipschitzian [FZ05]. non-local [CS03, CK04, RXZ20]. non-measurable [Dud94]. non-null [PV20]. non-parametric [Fab90, Jan95]. non-reflection [Tru03]. non-regular [MR13]. non-relativistic [GMM17]. Non-reversible [LS22, KM12b]. non-separable [Eig86]. non-separating [Lou22]. Non-simple [MSW21, Yan22]. non-smooth [KP89]. non-summable [DM17]. non-symmetric [CZ16, OS95]. non-uniform [CDR86, CS01, DGL87]. Non-uniformly [BS22, CGM08]. non-uniqueness [DF16, HLP24]. Non-universality [DZ18]. non-vanishing [BM11b]. Non-zero-sum [Mor86]. nonadapted [Lie98]. nonamenable [Hut24]. nonclassical [Erd98]. Noncoalescence [CL90]. noncolliding [KT07]. Noncommutative [JZ15, Sch90, BC12, Leh03, Ose08]. noncompact [Arb91, Dow88, Li94b, Li94a, Wan97a]. Nonconventional [Kif10, KV14, Kif13a, Kif13b, KR14]. noncritical [CDCIV14]. nondegenerate [Kry14, Wei86a]. nondeterminism [Nol89]. nondifferentiable [Düm93]. Nonequilibrium [JLS09]. Nonexponential [FL93]. nonhomogeneous [KM09]. nonintersecting [BG13b]. nonisotropic [GW86]. Nonlinear [AK87a, BR24, CL97c, FK10, KLP01, PZ00, BMS05, BHW19, Car87, CS89, CC05, CHMP02, DLTV24, Dun24, Dyn91b, Flo90, GM89, Gra95, HLN13, Kol11, Kot88, LWXZ12, MPS19, MU90, MRT19, MRT20, OT17, PZ98, Pen99, RSZ23, Ros87, Zha20, dBD02, vH09]. nonmonotonous [Ust88]. nonnegative [Muk93]. Nonparametric [Dah90, DH22, GH10, CST24, Coc18, DJ00, GN98, HM87, HP94, HI86, LT00, Lie89, LS89, MR13, MN98b, MSW09, Nic07, Nic08]. nonplanar [Häg03]. Nonsmooth [Cla86]. nonstandard [Sto86]. nonstationary [BM12]. norm [BY86, BR92, DMS05, Don94, FM19a, FT04, FJK23, GL09, Gri95, Kar07, Kon90, Kry10, LNS99, LT00, Li92, Li08, Li14a, Mal12, Par22, RT86, Ver11, vdGL13]. Normal [AI98, LPS16, Nol14, Rai98, AB07, Bon20, Cek98, CHM86, DL01, Hal90b, LS02, MS90]. normal-Poisson [Cek98]. normality [Bry08, GW04, GS90, HJL90, Lus92b, Mam92, Ohl89, QD91]. normalized [BC04]. normally [HS88, Ruc88]. normed [BMS08]. Norms [CH88b, Bob10, Cap95, GKS04, HJM20, Hit90, Jan10, LU03, LM19b, Zyg09]. note [Ano87a, Bug85, CGW10, CT86, Ein89b, EM91, Fal90, Imk86, IV00, Röm96, YK88, Zie88]. Notes [Ber19]. notion [Lov95]. Novikov [Stu93]. noyau [AL91a, AL91b, FL93]. nucleation [BJL17, CM13, DS97]. nucleation-and-growth [DS97]. nuisance [Ruk86, Sha92a]. null [LR94, PV20]. Number [Nán11, ABF23, BCP19, BMM20, CLM94, CP88, GS24, HL18, HMT20, HS11a, KN03, Maj92, PS98, SSZ22, Suz24, Wei88]. numbers [ASS19, BL12a, BW17a, CLDDMG99, CM88, Dob88, EZ22, Fil89, FW89, KM22, LR86, LP19, LM89, MP02, MN87, RST15, San94a, San94b, Shi06, Rio00]. object [Kni01]. oblique [Cos92, HT10, Kwo92]. Observability [vH09]. observable [Kry15]. observations [DJP01, GP00]. observed [Coc18, MR03, You89]. obstacle [LM17]. obstacles [DFSX20, Pov99]. obtuse [ADP16]. occupancy [GI20]. Occupation [BGS16, BCG88, Imk92, LR94, CR86a, DM91a, EP97, IW94, Isc86, LS15b, MW95, Mou90]. occurs [Mór91a]. ODEs [Tea08]. odometer [CHR18]. off [BV22, DDMP19, Fre19, LLP10, Por96]. On-diagonal [Kaj13]. On-line [LMU03]. Once [DKL02, Maj88b]. One [CF08, Don97, Kry03, Liu22b, Rei89b, ABBL24, AW01, BLM11, BvE24, Bol90, COP93, CGZ00, CGZ03, DL03a, DM23, DGL87, Dyn91b, FR95, Fun89, GKdV89, Gol07, Gyö86, HHW94, HS94, HLN13, JLS09, Ji89, KR16a, KTT89, Kni01, Kön93, Kön94, KZ17, Lal09, LS87a, LST20, MR09, Pap87, PPZ99, PR20, PS93, Sch15, Xia04, Xia07, vdH01, vdHdHK03]. one-arm [DM23]. one-dependent [GKdV89]. one-dimensional [BLM11, COP93, CGZ00, CGZ03, DL03a, Fun89, Gol07, Gyö86, HHW94, JLS09, Ji89, KR16a, KTT89, Kön93, Kön94, MR09, Pap87, PPZ99, vdH01, vdHdHK03]. one-endedness [BvE24]. one-parameter [AW01, Sch15]. One-point [CF08, Liu22b]. One-sided [Don97]. online [DFGGR19]. only [Coq00]. Ono [Tzv10]. Onsager [Cap95, CMN92, MN02, WZ93]. Onset [Gia95]. Opdam [Sch07]. open [BKWW23, CT11b, Eba86, Yan22]. Operator [Luc91, RW19, Ans99, GH91, LU03, Mit87, Par22]. Operator-semistable [Luc91]. operators [Bar89, BL07, DT00, DL01, FS88, FV08, GKS16, Kar08, Klo13, Kol11, Lej01, Mer03, RXZ20]. Optical [Mam87]. Optimal [CRZ13, CMW15, GH23, GM05a, GT24, GT16, Jir16, KV19, LW90, Nes19, Roh11, Zha19, Zha20, AB11, Bal98, BEES18, COT19, Div22, DS14, Don94, Efr99, FGM09, HT10, KN03, LM17, LM19b, NW22, Rob94, RD13, SSSY17, Tal05, Zha18a]. optimality [Mas92, PV20]. optimization [GT24]. Optional [FK97, Kra96]. oracle [BMN12, GL09, Yur11]. orbits [GRS24]. Order [DeB05, KN98, AW15a, BK17b, Ber93a, Bre08, Cha09, CW93, Che00, Con19, DeB04, DV15, Deh86, DD88, DP96, DWXY23, DCvEH18, EHM88, Gu17, GM05b, HH90, Hos91a, Jan88, LPS16, Lej01, Mar08, Mar91, OT17, Pap90, Pis96, STZ12, Wat93, vEH88, vH11]. order-preservation [CW93]. Ordered [CD01, IVW18]. ordering [BC22]. Orderings [HL18, LU03]. ordinary [Oli16]. organized [BBC^{+}04, For21]. oriented [CS08b, CS09, DCTT18, Shi03, vdHdHS07]. Orlicz [vdGL13]. Ornstein [BCL23, AC87, CIvV03, CS08a, CMG02, DMRYZ04, ELY99, FZ99, GRT20, MS92, Mit87, Shi90, Sim11]. orthant [TW93]. orthogonal [FR04, Fre19, PX88]. orthogonally [FW24]. oscillating [Wat88, Yan24]. oscillation [Kaj13]. oscillators [OS13]. oscillatory [Tan99]. other [Bir86, Ein93, IR89, RS23, SY01]. out-off [Por96]. Outlets [DS11]. Outliers [BZ20, BGR16, Tao13c, Tao13b]. output [AL94, Mai89]. oval [CLM94]. overcomplete [ASS00]. overfitting [CST24]. overlap [AC13, BG97b, EL04, GJS21, GL99]. overlapping [Alh04, CP88]. overlaps [BD20, CD01]. overparametrization [MZZ24]. p [Yos93b]. p.d.e [MT95]. Packing [CPW95, CLSW01, Ber03, DD17, FT92, MS09, Tal05]. Pairs [Aff88, LM22]. pairwise [Bra89, Sun98, Uch88]. Palm [BFQ18, Kal00a, Kal03, Kal07a, Kal07b, Kal11, Kar87, Nie89, OS16, OS22]. Parabolic [GM98, LQ19, BCF19a, BS22, BKdS18, BRS14, CM95, CV98, CMS05, CM07, CGM10, DDD19, Dok04, EdHM15, FKN24, FK10, GdH99, GKM00, Gro89, GM16, HR95, Mou14, MPS18, PT99, Wat88]. parallel [BR97]. parameter [AW01, Ble93, CF94a, CL91, EL04, FS10, FR04, FR05, FL96b, FK10, FW12, HM97, HS95b, Imk86, Imk91, IW94, IR89, KS95, Man86, MZ87, Mor86, Ruk86, Sch15, Szn11]. parameterization [MP88]. parameters [BCD21, Car86, KKM99, LR90, MS90, Sha92a, Thi91]. paramètres [Car86, LR90, Thi91]. Parametric [EGL14, You89, Fab90, Jan95]. parametrization [Zha19]. Parametrized [Sep93]. Pareto [BRM24, Fal90]. Parisi [AC15, LW04b]. Parking [ACCH23]. Part [GT97, CKM93, HRS17, Tru03, GTT98, GdH94]. Partial [LV12, Yos04, BM01, BP99, BQR00, CPY94, CJ94, CIK99, CHM86, DP98, DS87, DT12, Ein88, EM93, EG95, EM97b, GP93, HMY02, HT24, Ino08, Jac90, Kim04, KS88a, Kot92, Kot95, KK10, LWX04, MY99, MP86, MPS18, Myt02, Rei89b, Sch23, Wan18b, dBM13]. partially [Kry15, Pit95]. particle [AD95, AS05, AS09, BB24, BBF12, BG86b, BCFN89, BD11, BW17b, Cha94, CG90, CK00, CK02b, CL99, DH22, FG88, FV00, Fon09, FGM09, GS89, Geo94, Gia95, GRW92, GdH94, GK12, GKR07, HD21, JLS09, Kni01, KLZ23, Kol03, KV18, Léo95, Lig87, Mai16, Mou92a, Rez96, Suz24, Uch89]. particles [CL97a, DGL87, EMS13, GM89, Mör01, NT86, NT87, Osa98, PR08, RS93, Uch88]. particule [GS89]. parties [NW91, Rio01]. Partition [dT07a, CH02, MNS15]. partitioning [Hil87b, LW90]. partitioning-inequality [Hil87b]. partitions [BO09, Pit95]. partly [Jen89]. Parts [FI07, AM96, BC11a, LB95, NW91, Rio01, RT02, Zam02]. Passage [DM05, ALS20, AD13, AG23, BBS21, BP94, Boi90, CP12, CT11a, CD13b, DHS15, DHHL23, Dau23, DG24, DZ18, DZ19, DGS21, Don89, Don12, DR13, DR16, DFS20, FN15, FW20, GH23, HM07, HN97, Joh19, KKR88, KZ97, LNP96, Liu22b, Pes02, Yao19]. passing [DM16b]. past [GM05a]. Path [Blu05, CL91, Dyn91a, FI07, MZ02, Nor93, AGM03, Ber95b, BK01, BDS93, CN07, CQ02, Ebe97, FO14, GT05a, HSWZ18, IR06, KP89, KW04, Kni01, Le 93a, Li03, LN15a, LN15b, LQ97, MR03, Mik04, RT02, Sch96, ÜZ95, Wan18a]. path-dependent [Wan18a]. path-valued [Le 93a]. Paths [BJ02, MW17, AFZ87, BG24, BM01, Bur89, BL90, BL01, BM05b, Che18, DD16, DH96, Eva87, HHK90, Joh02, LQ13, LS87b, Qua99, SS18, Sch98, Ver97, Zam02]. Pathwise [BH00, BG02a, BCF93, CD99, MP11, Ove94, DGO20, Vel09]. Pattern [HRS17, CP88, DG93b, KW15]. Pattern-avoiding [HRS17]. patterns [GL99, Met13, Mór91a, Tal98b]. PDE [AST19, FL96b, FK10, FW12, Fun95, HR95, MT94, Shi06]. PDEs [DeB05, Cas01, DeB04, Hai02, MPS19, Oda08, PT99]. peaks [CK12].

Pearson [GTS87, GTS84, JZ10]. Pearson-type [GTS87, GTS84]. peeling [BRM24, CQ23]. penalization [BBM99]. penalized [Gol04]. Penalties [BM07]. penalty [vH11]. pentads [DSS07]. perceptron [MZZ24]. Percolation [BGN91, BR08, CK15b, HMM19, Qui10, Roy91, Tei16, TW12, ATT18, AD13, AG23, BBS21, BBS00, BBS16, BB06, BP94, BS98a, BM03, BNP11, BGP03, BB07a, Bod05, Boi90, CN07, CCC91, CG09, CM01, CP12, CM11a, CT11a, CS16, CD13b, CCD88, CCGS91, CS08b, CS09, CS12, DS11, DHS15, DHHL23, DvdW01, DG24, Der22, DP96, DM23, DZ18, DZ19, DGS21, DFS20, DCTT18, DCvEH18, DCRT19, DCH24, DS88b, FN15, FW20, GKN92, GH23, GHS21, GHM12, GH19b, GKZ93, GM14, HP99, Häg03, HM07, Har90, Hig86, Hig93, HST15, Hol03, HS11a, HN97, Hut21, Hut24, Joh19, Kes86, KZ97, KS00, Kho08, Kis15, Lac14, LNP96, Liu22b, Maz99, MPV01, MSW21, Mön24, Mou92c, Ngu88, PS98, PSSS13, PSS15, PSS20, Pis96, PRS16]. percolation [Sch99, Shi03, SS04, Szn11, UA03, Yao19, vdBN21, vdHS05, vdHdHS07, BR10]. Perfect [AS95, DM16a, BGP94, Dal88]. Performance [Hal90c, El 18]. perimeter [BH98]. Periodic [BL21, BBT23, Sch86, BD23, BdT10, Fra05, GK24, Lia22, MR09, Rol90, Tzv10, dT07a]. periodization [Owh03]. Perkins [Kry97]. permanents [RW05]. permutation [BS19, BP15, FM21]. permutations [BU11, Bor20, Fic91, HRS17]. permuted [Au22]. permuton [BHSY23]. Persistence [BB21, DM17, EF98, GRW92, PS15, BMN12]. Persistent [Tót86]. Personal [Pen90]. Persönliche [Pen90]. Perturbation [CFKZ08, BH13, EK87]. perturbations [BZ20, CK04, DG98, FW04, FW07, FK10, Hai12, Kor04, Las19, Péc06a, Péc06b, Sow92, Tao13c, Tea08, ÜZ07, Yan24, Tao13b]. Perturbed [PW97, AHR01, Ber90c, CD99, Dav99, EF93, Eva89b, FIKP13, HLP24, KMP22, Pic89, Son93]. perturbés [Ber90c]. petit [Léa87b]. petites [LT88]. Pfaffian [FM12b]. Phase [BBC^{+}04, BK07, BBN20, Buc21, CdH21, CS16, Cha16, CCGS91, MW01, RS17, ATT18, BD02, BJL17, BG97b, BS95b, CLM^{+}12, CP10, DG93b, DDMP19, DWXY23, DCM16, DCMT21, Gam04, Gre91, GdH92, Ham15, Hry98, JT17, Klo13, MT94, MT11, Pap87, Shi03, Tan99, YY21, vdHL10]. phases [Fun95]. phenomena [BGL19, DG93b, DG21b]. phenomenon [BL97, BM11a, BV22, Fre19, Mir11, Por96]. phylogeny [RS17]. physical [NS23a]. Pick [LM00]. Piecewise [Hof86, BP87, SCZ21]. pinned [BFO09, DV00, Rod19]. Pinning [BT03, CSZ16, CdH21, GLT10, IV00, Lac10, Szn96a, Szn96b]. Pirogov [HPR20]. pitchfork [BG02a, BEN23]. Pitman [CC21]. pivoting [HT24]. plan [Gol96, LM92, Wer94b]. Planar [Cur16, DCMT21, FIKP13, AD22, BB95, BMM20, BBCK18, Cha10a, CS04, DCM16, EL21, For21, Gol96, GH20, GHS20, Han12, HS95c, Lac14, Lam22, LM92, LW20, Mas92, Mat95, NST23, Ric18, Sil17, Ver97, Wer94a, Wer94b, Xu91]. Planarity [Lou22]. Plancherel [FM12a]. Planck [RXZ20, ASZ09, BR24, Zha07]. Plane [dT07b, AR18, BGP14, BMS02, BCG86, BR97, Bud18, CL16, Dal88, Joh00, Kee88, MS17a, Mör00, Rév92, Roy91, Sep98, Wer94a, Zha15]. planes [BS98a, NW90]. plans [Mat95]. planted [DWXY23, MNS15]. plates [JR95]. pockets [FKW19, Szn99]. Poincaré [BL97, BG99b, Cha09, CLW11, ELP08, Fle12, JZ15, KY92, LN05, LPS16, TY23, Uek89]. Poincaré-Hopf [Uek89]. Poincaré-type [BG99b]. Point [FW20, Xia04, Xia07, AK10, BDC12, BKP21, Ber87c, BCT12, BFQ18, Cer24, CF08, CGH24, Dab90, DDG12, DL01, DL03b, DGL87, FL95, FM12b, GZ93, GT24, HJW22, HHL88, Jac87, KMTT19, Kal09, Kar87, Kel87, Las94, LT00, Leu10, Lia22, Liu22b, Mas92, MW17, Müc94, Nie89, OS16, OS22, PPY92, RAS14, SK86, Wu00, ZD90]. Point-to-line [FW20]. point-to-point [RAS14]. Points [DM97, Fou98, LM92, AB22, AS92, AM13, BF88, Bar00b, BB96, BM05a, BR97, CGS23, CLM94, Dal88, EK96, Eva87, FR92, GS24, Hil94, HRS17, KMTT19, LM22, Maj92, Men16, MS17b, Mir13, MS09, Shi92, SZ17, Tal98a, Wer94b, vdHS05]. Pointwise [AK87b, GM16, KT04, Str16]. Poisson [Bai08, BE17b, BD11, Cek98, CR95, Coc18, Den00, Fag90a, FS10, FM89, Hau07, HV09, HHT21, Hol11, HMO23, HS11b, Ito88, KMTT19, KR14, LP11a, LPS16, Maj92, MMŚ21, PPY92, Pri96b, Pri96a, RB03, Rol90, Roy91, TW12, Wu00]. Poisson-driven [BD11, Den00]. Poissonian [Bié98, Gol96, MW01, Pov99, Szn93a, Szn93b, Szn96a, Szn96b, Wüt98, vdBBdH05]. poissoniennes [Gol96]. poissonnien [Bié98]. Poland [BL24b]. polar [She93]. polaron [BDS93, Man91]. policies [SS87b]. Polluted [GH19b]. Pólya [HS13, PS18]. polygonal [AS89, AS91, vdBS90]. polygons [Hua24]. polyhedral [Wil87]. Polymer [CLM^{+}12, BCD21, CH02, Kor22]. polymers [BC22, CN16, DD16, DR92, HTV12, Kif97, Var07, vdHdHK03]. Polymorphic [CM11b]. Polynomial [CS20, EI20, GVV22, Ale02, ABZ23, BLBE92, BCGZ22, Bor88a, EL21, FI19, Mit87]. polynomials [AT86, BCP19, Bry08, DS18, GT97, GTT98, GS15a, Mal12, MSS22, Par22, Sch05, Shc13, TT91, vEH88]. polynuclear [BSV10]. polytopal [Bru16]. polytopes [CY14, CY15, Fle12, Mül89, Rei05]. ponctuel [Leu10]. ponctuels [Jac87]. pont [Gol96]. population [BL12a, DS23, EP19, FRS23, Ohl89]. porous [BRR11, GdH14]. portée [Lar95]. portfolio [IPS13]. portion [CHM86]. posed [CIK99, Gol04]. posedness [DLTV24, Yon08]. position [CK12]. Positive [JP91, Ber93b, Car05, Cha10a, CW93, CWZ95, DGS21, FQ88, HP91, O’C21, Ros89, VW09]. Positivity [Osa98, LW00, Wan18b]. possessing [NZ90]. possibilities [DS95]. possible [CM88, FNS87]. post [Kaj13]. post-critically [Kaj13]. posterior [Bar24, DFM24]. Potential [SV03, AH20, BH13, DV00, Fit90, FM90, GH89, Mas92, MW01, MPS18, Szn93b, Szn96a, Szn96b, Wüt98, Yat13]. potentials [COP93, Con24, GNW23, Häg95, HL19, KMZ11, MMS14, MS87, Mer03, NW22, RAS14]. Potts [Ale01, BM17a, BCT12, CDCIV14, DCM16, Pis96, DM16a]. pour [Bou88, Car86, CSC93, DLP87, Lar95, LR90, Mat95, MP87, Mic99, Pet92, Rau88, Rio96, Rio00, Rio01, Thi91, WY93]. Power [HS13, Hut21, BLS87, CMRT15, DNS22, DM05, DFGGR19, Jan10, LLP10, Nut12, Rai03, Yuk95].

Power-law [Hut21]. powers [LM22, Rai97b]. pre [HHK90, Osa90]. pre-Sierpinski [HHK90, Osa90]. Precise [MW13, PPZ99]. Predictability [IMŞK93, SK87]. Predictable [Las94, Fer89, Mai89]. Prediction [MW05, Ber92, GM05a, Lah99, MP88, RS94, TKPK86, Car86, Car86]. predictor [El 18, Lah99]. predictors [CP93]. preferential [CCL^{+}24]. première [GS89]. prescribed [Fun91]. presence [Cro13]. preservation [Cha24, CW93, SK87]. preserves [Mit86]. preserving [DG14]. presque [DD88, MP87, Rio00]. pressureless [Der03]. Previsible [Str86b]. prévisibles [Mai89]. prime [HL18]. Principal [Jan95, BKYY16, CCFS00, EK87, GJS21, MW01]. Principes [DLP87]. Principle [Lig87, AKLP23, ABDH13, BPS18, Bau16, Ber90a, BB07a, BD14, BP86, BGdH10, BGdH23, BCW21, CF09, CK88, DRZ02, DZ96, Dup10, Ein87, EF93, FL96b, FS00, Gan94, GL14a, GZ93, Gol04, GVV22, GZ12b, HLLS23, Kif13a, Kif13b, KS88d, Kry14, MT20, OS95, Pes94, Pow19, PRS16, QSY24, RAS05, RXZ20, Sam87, Sch15, Str64, Tro12, Yos93b]. principles [BM12, BBT16, BK92, BD89, BGP95, BW17b, CK08b, DNS18, DLP87, Fit90, Joh86, LPR97, Mar10, MP86, Rio94, Sep95, SY01, SS04, Wat87]. priority [Fon09, Xia07]. priors [Cas12]. Private [BSV24]. Prob [Fon09]. Probab. [Ano98a]. Probabilistic [BRR11, CGM08, DL05, MS00, MPS19, Son93, AMS94, AS22, Cra93, Dyn91b, ESZ09, Fon09, HMM19, LRV23, Oss05, Pap90, Wan18b]. probabilities [BR90, Bor03, CH19, CP03, DKN09, Don12, DCRT19, EP19, FPW24, Fil89, JZ10, KS18, Law92, QR13, Roz86, SW95, SB03, VW09, Wit86]. Probability [Bar89, DeB05, AM96, Ans99, BNT05, BGN91, Bel08, BNP11, BBS14, BPR07, BM15, BGR15, BGP24, Bry88, BNPS18, CMT15, CSS23, CHS22, CG13, Col05, Com89, Dar92, EM13, FJK23, For09, Gri21, HM88a, HM92, HHW18, HLO23, HMZ20, Hil87b, HNY99, JS86, KLS86, Kes88, KS00, Leh03, Mar87, MM90b, Mło99, Mou92c, Neu24, Pic89, QRZ03, RS91, Shi97, Thu86, Urb90, Wan18a, vdHdHS07]. Problem [DR99a, DR99b, Ald92, AST19, Ass05, BCL20, BH05, BPP04, BD02, BJ20, BL07, BJL17, CM24, Cos92, COT19, CM07, DWXY23, EK87, FM12b, FNS87, FL96b, FW12, Fug88, GS04, Gre90, GdH94, HT86, Hig86, HS00, JP91, Kry94, Kwo92, LM17, LW04b, Liu22a, MW89, Mik04, NS23a, Ndu91, PW20, Pap90, Roz03, Son93, Tal01a, Wan94a]. problème [HT86]. Problèmes [Car86]. problems [AW15b, BY86, Bal98, BEES18, BDN17, Car86, CT02, CIK99, Col05, DO05, DM09, Dud94, EF93, EP92, Gam04, GM98, Gol04, GPW13, Hal89, Hal90b, Hal90c, HLØ^{+}93, KMM98, KSvdVvZ16, Kry15, Mor91b, NN04, PZ98, RS94, RR95, Ruk86, Zam00, ZD90, dBM13]. procedures [Ber86a, KSvdVvZ16]. process [ADH14, AD22, AR89, AD95, AMP04, AMP05, And86, AK87b, ABL88, ABK13, AGL17, ABF23, Ass05, AJV18, BMRS20, BG86a, Ber87c, Ber86b, Ber00, BK93, BCL23, BL19, Bur15, CGRS16, Car86, CF24, Cas12, Cei12, CV16, Cha16, CCD88, CCGS91, CK20, CM97, CDS91, CFR87, CM96, CL16, DS20, DFP15, Dit90, DS88b, Dyn93a, EV21, Eis97, EM20, ERV02, Eva87, Eva89b, FS10, FF94, FV00, FW08, FFN12, Föl89, FM12b, FL96c, Gre91, GK12, Gri91b, Gri95, Han02, HH91, HHL88, HNY08, IW94, Jac90, Kat15, Las94, LP11a, Leu10, Li92, LN15a, LN15b, Mit87, MT94, NV21, OS16, OS22, Pen96, Pri96b, QRV99, QR13, Rév92, Sal86b, SV17, SV86a, Shi94a, Sta01, SZ17, Swi90, SV86b, Uch89, Wee92, Win91, Yuk87, dHHKR91, vdHS05]. Processes [DeB05, GM06, AD08, AE14, ACJvM15, AN92, Ahn22, Ald88, Ale87a, Ale87b, AGOZ88, AM08, And08, Ans13, AC87, ABO23, AL09, AK10, AS10, AFZ87, AW01, Bah98, Bal86, BK08, BM12, BL12a, Bar00b, Bas88, BES00, BC01, BC06, BL12b, BJL17, BKP21, BFF18, Ber89, Ber93b, BL00, Ber01b, BL03, Ber20, Ber21a, Bes91, BK01, BBC03, BGR15, BG86b, Bor88b, BG13b, BC14, BCF18, BGS13a, BM16, BW05, BKWW23, BH09, BMS08, BM11b, BF18, BFQ18, CH93, CF94a, CD13a, CL17, CMN92, CS89, CKS11, CLO01, CCM16, CR90, CR93, CW93, Che93, CFS01, CK02a, CK08a, CF08, CFKZ08, CKS10, CKK13, CP93, Che18, CK88, Coc18, CMMM11, Cra17, CL97c, CH88a, CL91, CS94, CK08b, Dab90, DV02, DV04, DeB04, Deh88, DG88, DS95]. processes [DM17, DDG12, Deu87, DDL90, DMRYZ04, DM05, DR13, DSS14, Dri13, Duf92, DE92, Dur09, DP99, Dyn91a, Dyn93a, DKS94, ER96, EM93, EG95, EM97a, Eis03, ELY99, Epl86, Erh00, Eva00a, EPW06, FJL09, Fag22, Fag90a, FF91, FZ99, FKL04, Fen94, FS10, Fer89, FM86, Fit87, Fit90, FS08, Fou98, FL96b, FS00, FW04, FW07, FW12, FM89, GY05, GS89, GKdV89, GOZ21, Gho15, GM05a, GHV90, Gin94, GS07, GJ18, GHK22, GHK23, GK91, GRT20, Gor92, GZ18, GRW92, GVA86, GK01, GLWY09, Gyö86, HH90, HK03, HSWZ18, Haw98, HS95a, HS95b, HS94, HV15, HJW22, Hoo91, Höp93, Hor90, Hos91a, Hou90a, Hu94, HM24, Imk86, Imk92, Imk94, Isc86, JY03, Jac87, Jaf99, JS86, JR88, JLS09, KMTT19, Kal89, Kal00b]. processes [Kal03, Kal09, KS94, Kar87, KM88, KRST11, Kel87, KLL94a, KSX08, KX09, KSX09, Kim21, Kim23, KM02, Kol08, Kol11, Kon90, KZ17, Kry10, Kry11, Kry15, Kry21, KU97, Lam02, LY97, LN05, Le 87, Le 93a, Lee90, LR94, Lig96, Lus92a, MM90a, Man86, Mar98, Mas92, MRT19, MRT20, MP88, MS95, Mic08, Mik04, MS92, MMW15, MR95, MO86, MO87, MP86, Mou92b, Müc94, MT95, Nag87, Nag89, NN04, Nie89, Nol89, Num90, OTV90, Oel87, Oel89, Ogu01, Oll87, OB04, Ort89, OS95, Osa01, PPY92, Pic96, Pit96, Pou88, PZ11, PS15, QRZ03, RR89, RS94, RW05, Res88, RB03, RRO24, Rio98, Rio01, RT02, Rol90, Sal86a, iS91, Sav09, Sch07, Sch98, SK86, SK87, Sch90, Sei00, SY01, Sha92b, Shi90, Shi94b]. processes [Sim99, Sim11, Son93, Sta02, Stu93, Stu91, Swa18, Tal88a, Tan86, Thi91, Tro95, TKPK86, Ust88, Von95, VA92, Wan95, Wat96, Web91, Web89, Wee88, Wee90, Wen96, Wen99, Wen04, Wu00, Xia04, Xia07, Yan09, Yau97, Yos96, Yu93, Zhe95, dHP86, DR16, Pic10]. processes.- [Wen04]. Processus [Rio98, AFZ87, Ber86b, Ber89, Car86, Fou98, GS89, Jac87, Le 87, Leu10, Rio01, Thi91]. produce [EHJT13]. Product [Col05, Ahn22, Cli86, Dwy90, Eri90, EZ22, FM19a, Fer94, GG96, KS89, KN03, LS86, PTZ20]. product-limit [LS86]. Products [HS88, LL04, MP02, Ahn23, BY86, BM02, BN94, FS88, GLP17, Kar07, Kar08, Muk91, Ver11, Wat87]. produit [Fer94]. Produits [FS88]. profile [BN22, LS19, Won20]. profiles [DMX21, LS15b, NOT22, SCZ21]. progeny [ABB11]. programming [Kry14, SS87b]. Prohorov [Mas88]. projections [Col05, DG88, Las94, PPZ14]. Projective [Tho01, LM22]. projectors [JW24, NSU19]. Proof [DR14b, BD88, CN07, DMR11, Gay19, LW04b, MR07a, Sch88b, Swa18, Tra90]. propagating [CHM22]. Propagation [Der03, KLZ23, NT86, AN92, FL96c, GGLS20, Gra95, Gui97, KZ98, LL23, MMW15, NT87, Szn86]. Properties [HP94, Mil12, Sug94, AD02b, Alt89, AC13, AC15, Bai86, BF11, BLS87, BT03, Bry88, Bug91, Bur89, Cer99, CCD88, CF03, Cra00, CM18, CL91, DDG02, DLM03, Del99, DGI00, DCMT21, Hai02, Hal89, HJS88, HS95a, HS95b, Hol11, HR95, JT17, Kal09, KP86a, Ker89, Kes88, MPV01, Met13, PZ11, Ram98, Rol90, Sta02, Zha19]. property [AB07, Bon88, BQR00, CPY94, DCTT18, Eba86, GY05, GJS21, Kel87, Kol87, Kwo95, MS00, MU12, MP92, Mur24, PRT13, RW88, RW89, vH13]. prove [BM19]. proximity [SZ10]. pruning [BFF18]. Pseudo [Che92b, LS88, LW18]. Pseudo-free [LS88]. pseudo-metric [LW18]. Publications [Ano21b]. pulses [BM02]. punctured [Gru98]. pure [Bas88, Imk91, KR16b, Pan15, SZ17, Sub23, Tal10]. pure-jump [KR16b]. pursuits [LS01b].

quadrangulations [CS20, Lou22]. quadrant [Ros89]. Quadratic [BH08, Bau13, BG97a, Bou90, CM24, FT87, GLT10, Gin94, GS07, GS90, Kle06, LP11b, Liu17, Mik04, Oli16, SB03, Tan86, Tan99, dJ87]. quadraticity [Kou89]. Quadrature [Web91]. Quadri [dT07b]. Quadri-tilings [dT07b]. quadric [Has89]. quantile [CK08b, Dab90, FR89, HM88b, Hal90c]. Quantitative [BGV07, CDM24, Coo23, DNS22, EM13, JW24, CM24, MMW15, Mos12, Mou14]. quantization [DT00, HKK23]. quantized [GMM17]. Quantum [Bia91, CC21, FM89, Kem05, AKRT97, BCS11, Bar89, BL91, BFPP19, BHSY23, CS08a, Coq00, DZ95, Erd01, Eva89a, Fag90a, Fag90b, FR03, FNS97, Fre19, GHS19, GPS22, HKK23, JP91, Li19, Lie98, LW00, MN04, MQ20, MS21, MSW21, NP12, Pri96b, Sch90, Uma06]. Quasi [Ces01, CMMM11, GNW23, Han02, Hen07, MN93, OT17, CV16, CCM16, DeB12, FK10, GP93, Müc94, QSY24]. Quasi-compactness [Hen07]. quasi-cone [DeB12]. quasi-every [Müc94]. Quasi-factorization [Ces01]. Quasi-invariance [Han02]. Quasi-invariant [OT17]. quasi-linear [FK10, GP93]. Quasi-stationary [CMMM11, GNW23, CV16, CCM16, QSY24]. quasilinear [DMS05, Kot95, NP92, PP94, PT99]. Quasipotential [BCF15]. quasisymmetric [Tro21]. quaternionic [Por96]. Quenched [ACS21, BMRS20, BCK21, BB07a, BGdH10, BCW21, CES23, CGZ00, CGZ03, CM07, CM17, DNS18, DG22, GZ12b, Mer03, PRS16, RAS14, SS04, ANV15, BBT16, BD14, BG02b, BMP04, EdHM15, GLT10, Gui97, Kül01, Lac14, Yil11, BGdH23]. question [GK96a]. questions [MS24b, PT00]. queueing [BBM00]. queues [Bar01, Gra05]. quickly [LT15].

R [Iwa87, Tan97, Yos93b, KM93]. R- [KM93]. r.v [Ein88]. races [HL18]. radial [CKM93, ELY99, HL21, HM24, LB95, NS23a]. radiation [GMM17]. radii [BNPS18]. radius [Bas23, BCGZ22, Wu04]. Radon [GS04]. Raikov [Fuk94, Pet92, Rio00]. Random [AS22, Arb91, AC13, AC07, BC17, BWW18, BS17, BMP04, BCS18, CN88, CS04, Che18, CP00, DHP23, DGPS07, DHZ19, DKW23, Dun07, Dwy90, Ers09, FLT22, Fic91, FT19, FW04, FW07, GR19, GKZ93, HK97, HV16, HvdH11, Hil94, Jur88, Kal90, Kal95, KM88, Kor04, KSS86, Mad89, NP12, NTV17, Nor86, Pen23, PSS15, Sab11, Szn09, Tel89, ÜZ95, Wil97, YY21, YSA17, vdBCL16, vdHL10, vdHHN20, AB22, ABB11, ABBG12, ABCM20, ABS21, Aff88, AG22, Ahn23, Aid08, Aïd14, AdR17, ANV15, AGKW09, AEK19, Ald86, AS92, Ald92, AP00, AMP04, AMP05, Ale87a, Ale01, ARV13, ALN99, ABK08, AKLP23, ABDH13, ADS16, ACS21, AL00, AHJ21, AL02, ABM05, AC03, Ass08, AS17, ADP16, ABZ23, AT86, Avr88, Bab91, BY86, Bai86]. random [BCP19, BB21, BE17a, BF88, BB09, BC11b, BC11c, BBT16, BJMW11, BZ20, BR21, BS08, BSS14, BS14a, BSS18, Bau16, BK12, BDC12, BW18, BS22, Bel98, BG09, BGGM12, Ben10, BLV16, BP94, BM03, Bér05, BV95, BS19, BLQ23, BvE24, Ber90a, BGP03, BB07a, BD14, Ber19, BP86, BD89, BHS10, Ber03, Ber01a, BGP14, BBCK18, Ber21b, BU11, BBW14, BSW17, BDN17, BvdHS18, Bia91, BCW21, BR23, BV13, BSV24, BMP97, BGJ96, BR08, BR10, Bol86, Bol87, Bol90, BZ07, BFO09, BCC12, BG13a, BC15, BCGZ22, BO09, BG13b, Bor03, BJ02, BYY14a, Bou90, BGS13a, Bow18, Boy19, BH98, BBN20, Bre08, BDW21, BCF93, BCH81, BCH88, CY14, CQ23, Cam87, CR14, Car05, CX96, CH02, Cas01, CT02, CW17]. random [Ces01, Cha19a, Cha16, CCKR07, CL19, CV10, CH19, CCK15, Chi01, CG04, CG07b, Chr86, CES21, CES24, CM24, CPW95, CLSW01, CGH24, Col05, Com89, CGZ00, CP03, CGZ03, Coo17, CD90, CF94b, CS12, Cro08, CFK13, Cro13, CS23, CM88, CK15b, Dar88, Dav90, Dav99, DF91, Ded98, DR86, DG88, DH91, DD05, DGPZ02, DG21a, DD88, DG98, DNS18, DG22, DW21, DFSX20, DMX21, DH96, DT23, DŚ19, Dom17, Don89, DG93c, Don95, DM05, DLP87, DR11, DM22b, Dub02, DCM16, DCL19, DCMT21, DJPP13, DE92, DKL02, EV21, Ein89a, EK96, EM97b, EF93, EdHM15, Ers19, Eva90, FJL09, Fag22, FM19a, FHS12, FKL04, Feh17, Fer91, FM86, FKT88, FM90, FM06, For09, FHP18, FT87, Fre19, FGP10, FK16, FFL94, Fuk04]. random [Fun91, FM21, Gam04, Gam14, GKN92, Gan94, GW04, GIM22, GIM23, GZ93, GW86, Ger86, Gho15, GI20, GT24, GG96, Gol07, Gol08, GH89, GRS24, GLZ11, GLP17, GLL20, GS03, Gre90, GdH92, GdH94, GPW09, GM16, GVV22, GD90, GZ12b, GT05b, GH20, GHS20, Häg96, HS88, HP94, HK01, Ham92, Ham00, Ham15, HL18, HK92, Hau07, HLO23, HKR18, HPU04, Hig86, Hil87a, Hit88, HTV12, Hou16, HV04, HS00, HS07, HTV89, HL00, Hut20a, IR10, Imk96, IL01, JN91, JN08, JRdlR08, Jaw95, Jia12, Joh02, Jon09, JPRR18, KK14, KMTT19, KM22, KPyWmW88, Kar07, Kar12, KM86, KTT89, KM96, KLMH06, Kif95, Kif97, KN98, Klo13, KY17, KH03, Kol87, Kom01, KO01, Kön93, Kön94, KMZ11, KMP22, KS88d, Kül98, Kül03, KL00]. random [Kyp00, LZ10, LS15a, LP19, LOY98, LN05, LP11b, Le 93b, LR21, LY88, LS18, Lej01, LS24, LWR92, LS15b, Li16, LM24, Lig96, LRW94, LW04b, LLT^{+}19, LMU03, LW18, LPP96, LW20, LT20, MMS14, Maj88b, Maj92, Mal12, MRT23, MS87, Mar98, Mat94, MM15, MR03, MWZ05, MWZ08, MT87, MM13a, Mer03, MR07b, MR09, MN87, MM13b, MS17b, Mor92, Mou14, MS02, MS90, Muk91, Muk93, Mül89, Naj18, Nán11, Nes19, NS23b, NZ17, NZ21, NO15, Nol14, NRW87, O’C98, Off14, Oli16, Oll88, OS95, Osa12, Osa22, PS18, Pan90, PZ93, Péc06a, Péc06b, Pem92, PZ08, PSS20, Pet14, PRT13, Pin07, PT17, PPZ99, Pit95, Pop09, Por95, Por96, PR20, PRS16, RS12, Rai97b, Rai98, RR17, RAS05, RAS14, Rei05, RW05, Rho09, Ric18, Rod19]. random [Rol03, RS23, Ruc88, SCZ21, DM16a, Sas86, SS87a, Sch00, Sch89a, Sch05, Sch09, Sei00, Shi97, Shi12, SS04, SSZ22, Sta04b, Ste97, Szn99, Szn02, Szn17, Tak98, Tal01a, Tea08, Tei11, Tha91, Tót86, TW01, Tro12, VW09, Ver11, VW24, Vir02, Wac12, Wan94b, Wat88, Wat08, Wat87, Wie86, Wie02, Xia97, XZ13, Yan24, Yil11, Zäh86, Zäh88, Zer05, Zha14, Zyg09, vdBBdH05, vdBG13, vdHdHK03, vdHH10, vW97, FS88, HvdH19]. Random-cluster [BS17, BDC12, BGJ96, DCM16, DCMT21, Häg96, vdBG13]. Random-field [DHP23]. Random-walk [BC17]. Randomized [BSV10, BG07b, Dal88, Gre87a, Gre87b, KP86a, RSV01]. randomly [AS88, Alt89, BR97, CLM94, Shi06, vdBJ12]. randomly-growing [AS88]. randomness [BG02b]. range [AM08, AL09, AGL17, Ass05, AS17, AS21, BMRS20, Bau13, BL12b, BJL17, BGS16, BCW21, CIvV03, CP07a, CS08b, CS09, CS12, Cro13, CM96, CK08b, Del99, DFSX20, DD17, Eva89b, FT87, Häg95, HK01, HP20, Hut21, JLS09, KM93, Lar95, Mön24, Mou92a, MT95, Rob94, Swi90, TT91, Yan09]. ranges [XZ13]. Rank [ES03, BB24, CMW15, CP16, FIKP13, IPS13, IKS13, LM19a, LS87a, Péc06a, Péc06b, Tao13b, Tao13c]. rank-based [FIKP13, IPS13, IKS13]. Rank-dependent [ES03]. ranks [CP10, KP86b]. rapidly [BM08, CP00, MS10, Wat88]. rapport [Asp89]. rare [GT98a]. Rate [CH88a, GT03, LR86, ABBN04, Ass01, DR14b, Efr99, FW89, FGM09, FG15, GdH94, Gri87b, HM88b, JM90, KM12a, Kry00, LV86, LMW89, Mij88b, SS87a, Shi97, Zha18a]. Rates [Ale87b, Bel98, Dun24, EM93, GK91, Mou92c, Rob94, dBM13, ACLP20, BG93, BG97a, Ber87b, Bra87, BCG86, CRZ13, DS87, DFM24, Ebe16, Ein92, GdH92, HH90, Hal90b, Hal91, IPS13, MM13a, Ter16]. ratio [Eri86, Höp93, SCC19]. rational [HK11]. Ray [RR00, BFF18, DR24, Eis90, Eis94, LPW13, ST16, Val91]. Rayleigh [EPW06]. rays [MS17a]. re [EPW06]. re-grafting [EPW06]. reach [AKTW18]. reaction [Cer03, Cer05, CF09, DDL90, Dit88, FLT19, FV00, FG19, GJMM24, JLLV93, LV96, Mou92b, MT11, Oel89, Sow92]. reaction-diffusion [Cer03, Cer05, CF09, FV00, FG19, GJMM24, Mou92b, Oel89, Sow92]. reactions [Kot88]. Reactive [LN15a, LN15b]. Real [Bor88a, BCP19, BKP21, EPW06, FHP18, FM12b, Sch05]. realizations [AS89]. rearrangements [DSS14]. Reciprocal [Kre97, RT02]. recollections [Pen90]. Reconstructing [MR03]. Reconstruction [MNS15, GH10]. record [BPS18, Deh88, Lin87, Pfe87]. records [Deh88, KM22]. recovering [GS04]. recovery [Don94, GJS21]. rectangle [HS11a, Tal05]. rectangles [CLSW01]. Rectangular [BG09, BG07a]. Recurrence [CRS18, CK00, CK02b, Cha19a, FR03, GL14a, GS03, Joh86, KZ17, Kül98, PT17, Shi90]. recurrences [KN03]. Recurrent [Muk93, AdR17, AR18, BM01, Che00, HV04, HS07, Law92, LR94, Lev88, Mor03, MS02]. recursion [Mir13]. recursions [BDG^{+}09, BDG10, Kol13]. recursive [Arb91, Ber03, EP92, Ham00]. reducing [DFGGR19]. Reduction [CK08b]. reference [ASZ09]. refined [FTY21]. Refinements [DZ96, Ste97]. réfléchie [Lap90]. Reflected [GM06, Kwo95, Wil87, BN02, CD99, Che92a, IR10, Lap90, Mur24, Tru03]. reflecting [BH00, BCR04, BL01, BM05b, Che92b, Che93, DT93, FW20, FT96, GM89, HH05, Kan00, OS95, RW88, RW89, Sai87, Tan96, TW93]. Reflection [Ebe16, BPP04, Cer24, Cos92, DMP93, DO05, DM09, Gra95, HT10, Kwo92, NP92, Tru03, Zam02]. reflections [Por96]. Regeneration [NN93]. Regenerative [Ber97, Ber99, Kal03]. regime [ADN22, BW22, BS24a, BGP94, CCC91, Erd98, HKK23, Mön24, Tou98, YSA17]. regimes [AS21, Bow18]. region [LR88, vdBS90]. Regional [GM06]. regions [Ber92, vdB94]. registration [BLV12]. regression [AM88, BLS87, Bar00a, BMN12, CMS07, CMS23, CG92, CST24, Chi12, Dur02, El 18, Fab90, GP05, GS15b, Hal89, Hal90b, Kou03, KM93, LM89, LBE18, LNS99, Lie89, MU90, MR13, SCC19]. regular [ABS21, Alt89, Bel98, Bra89, Coo17, DS20, Dom17, DJPP13, Dyn93a, GIM22, GIM23, Hor00, HS07, KRST11, LLT^{+}19, MR13, Rio98, SSZ22, TY23, Ver90, Vie97, Yat13]. régularisation [AFZ87]. Régularité [Fer91]. Regularity [Cha24, Dun07, FTY21, Imk94, Jac90, AW01, Blu05, BH09, Dyn93b, MZ02, MM90a, Sta02, Yon08, Fer91]. Regularization [BF11, GG19, AFZ87, CHM22, Tal88a]. regularized [BMN12, El 18]. regularly [BPS18, MM13b, MW13, MW14]. réguliers [Rio98]. Reinforced [Dav90, BSS14, Ber21b, DV02, DV04, DKL02, MR07b, MR09, MRT19, MRT20, Pem92, Rol03, Zha14]. reinforcement [Ber20, Ber21a]. Rel [Fon09]. Relat [Ano98a]. Related [DeB05, Aeb93, AD13, BT03, BG09, BR08, BR10, BKWW23, CHMP02, Cra00, DP98, Eis90, ERV02, FV00, GRST14, Gre90, Kal07a, Kal07b, Kal11, Klü89, Kry11, Lee90, LT21, MT94, Osa12, Osa22, PT00, Yu23, Yur11]. relation [Aru15, BS97b, FI19, Häg95]. Relations [Mic99, DS95, Ful09, NUZ90, Mic99]. relationship [Bel86, EHM88, Mil89]. Relative [Ram98, Bax89, BS05, Cer99]. relatively [Dev88, DG21b]. relativistic [Bai08, GMM17]. relaxation [Ald88]. relaxed [Art01]. relevant [Lac10]. REM [Gay19]. Remark [DeB88, Gan99, Sta04b, Thr87]. remarkable [Bon88]. Remarks [Bry88, Ros87, GK96a]. Renewal [Lal86, AA88, Don97, ESZ09, Erh00, Grü86, Kee88, KM96]. renormalised [AZ20]. Renormalization [AL02, CGH24, LRV23]. renormalized [SS15]. Rényi [Bas23, HK21, San94a, San94b]. Repeated [Hec86, NP12]. repellent [Bol90, HTV12, Kön94]. repelling [BG17, HHH02, TW98]. replacement [PTZ20]. replica [BM19, FW24]. replica-symmetric [FW24]. replicas [Cat96]. Representation [Höp93, Mor92, AGM03, AB07, BBB90, BRR11, BL91, Ber87a, BFF18, BLQ23, BPR07, BPR11, CFV23, CC05, CF03, DT93, ESZ09, FHP18, Hig91, Ino08, Kal90, LP11a, LPW13, Li19, LMW89, MM90a, Oss05, Zha10]. representations [AP14, Bab89, GJV88, JY03, Jur88, Kal88, Kül01, LR21, LS86, Lo89, Pou88, RR89, Res88, vdBG13]. represented [MO86, MO87]. repulsion [BG02b, CL97a, CN20, Vel04, Zam04]. repulsive [KV18]. resampled [Szu92]. resampling [GPW13, Hal90c, Lah01]. Rescaled [DP99, SCC19]. resistance [CS23, KY92, Shi12]. resolution [Dah90, Pra93]. resolvents [BB03, Stu91]. resonance [HK11, KW04]. resources [GM94]. respect [BG86a, Bar11, BS98b, DM03, LM19b, MM90b, Zäh98]. response [HR02]. restitution [Lou97]. restricted [PRT13]. restriction [Qia18]. result [CF94a, CM91, Gri91a, Gri95, HH05, HK01, Kry97, Szn86, Szn89, Zhe95, dA00]. results [AM08, AD13, BBM00, BSC02, Bou90, Che90, Day88, DS94, Gru96, Hau07, Haw98, LP11b, LBE18, Li92, LM24, Mar08, O’C98, PV20, Péc09, PT87, RSX13, Ser95, Sil17, Tal98b, Tal00b]. Retaux [CHS22]. retrieval [BG97b]. return [HV09]. reversal [COT19, MNS89]. Reversibility [Zha15, Han02, MS17a]. reversible [BL12b, BvE24, Cha94, DDL90, GL14a, Kim21, Kim23, KM12b, LS22, Lig87, Mou92b, NOT22]. Révész [Tra90, Gri87a, Sha92b]. Revuz [BB00]. reward [Erh00]. rewards [Erh00]. Ricci [Kop19]. ridge [El 18]. ridge-regularized [El 18]. Riemann [DDD19, Kar13, KT07, KW16, Naj18]. Riemannian [Li14a, Li14b, BG17, Hab18, HLW24, Hsu90, Li03, Li08, Li11, LL24, Ogu01, Wan97a, Zha07]. Riesz [Li14a, Fuk94, KT04, Li08, Pet92, Rio00]. Right [Eva00a, Gam14, CDS91, FF91, Fit90, Müc94, Sal86a]. Right-convergence [Gam14]. rights [LY87]. rigid [Häg03]. Rigidity [BK17a, Cha19b, HL19, AL23, Suz24]. Rigorous [Tal98b, Tal00b, BG97b]. Ring [BGR16, GZ12a]. rising [EP92]. Risk [BBM99, Nut12, Don94, DJ94, Gol04, GS15b, LM09]. risks [Bar11]. Robbins [Mör00]. Robinson [MMŚ21]. Robust [CLL20, KLT17, TKPK86, Bar24, CC05, DS14, DM16b, El 18]. Robustness [DGO20]. role [HM97]. Root [COT19, WT04, BCZ^{+}10, BMS08, EPW06, Kry03, GOZ21]. root-unroot [BCZ^{+}10]. roots [BCP19]. rotation [VA92]. rotational [Kal95]. rotations [DG98, ÜZ95]. rotators [BGP14]. Rough [DD16, FO14, HW13b, BG24, CHKN18, Che18, CQ02, DGT12, Dey16, GGLS20, HW13a]. roughness [UA03]. routes [MMŚ21]. routing [LMU03]. rules [BBW14, DF16, DCH24]. run [Gri87a, Mór91a]. runs [GSW86]. RW [AJV18].

S. [Zai87]. sûres [Rio00]. Saddlepoint [Jen91]. Sample [Ber95b, Sch96, Wat96, Ale87b, BH05, BC04, BKYY16, CDR86, Che90, CK20, CHM86, Eva87, FR89, GHV90, HLO23, Hsi93, IR06, LT15, LQ13, LP11c, LLR10, Mij88b, NSU19, Par90, Péc09, RS17, Sch98, Shc13, Wei88, YBK88]. samples [JR88, Jan10, LPR97, MS10]. sampling [FHP18, HSWZ18, MWW09, Ohl89, PPY92, RST^{+}13]. sandpile [JR08]. sandpiles [CHR18, Hut20b, JRS20]. Sandwiching [GIM22, GIM23]. Sanov [Gru96]. sat [Tal01a, SSZ22]. satisfaction [Gam04]. satisfy [Sch90]. saucisses [Wei86b, WY93]. sausage [ASS19, EL87, HK01, vdBT91, vdBB94]. sausages [MS24b, Wei86b, WY93]. Scalar [KR16b, Oua22, CD13a, DR06, DV15, Mar10]. scale [ARV13, CKS11, DG93b, DG93a, DF20, GM16, Gu17, HLLS23, Kim21, Kim23, Par90, PZ93, Sha92a, Ust88]. scaled [Ers09, Lip96, MW01, Mer03, Swa07]. scales [BS24b, GP00, IN16]. Scaling [AdR17, BCF19b, BLV16, BL24b, Ber21b, CHR18, DGZ05, HS21, HS20, NST23, PS18, Uch94, Wüt98, BJ14, Ben10, BLQ23, BFO09, BRM24, CY15, CM01, CMRT15, CPY94, DR24, DF20, FLT22, Fun95, Hor00, Off14]. scatterers [BBT23]. scenery [AC03, AC07, MR03, Wan03]. Schelling [HS20]. scheme [BT96, BM19, GP05, GI20, MPY94]. schemes [Hec86, Jan06, Pet14, Sei00]. Schensted [MMŚ21]. Scheraga [BL24b]. Schlögl [Neu90]. Scholes [Gei05]. Schramm [HL21, MQ20]. Schramm-Loewner [MQ20]. Schrödinger [AN92, BCL20, BHW19, Con19, Con24, DLTV24, GKS16, MS87, Mer03, Nag89, OT17, PW20, Pra93, RSZ23, Sam22, Zha20, dBD02]. score [CR86b]. SDE [DD16]. SDEs [BR24, BD11, BDG21, Cer99, CM18, DHB11, Den00, GHM23, HLLS23, MZ11, Roz03, STZ12, Ver03]. search [LR88]. Second [Kim21, Kim23, Cha09, Che00, Con19, Hos91a, Imk92, LS01a, LPS16, Pap90, STZ12]. second-order [Hos91a]. security [Kra96]. seen [ABBS13, CDS91, GS89]. segment [BN22]. Selection [BM07, Bar00a, Bar11, BBM99, Bir14, HM91, Kou03, Mai16, SU86, WZ99, WZ13]. selections [BHS10, FR08]. selective [GT05a]. selectors [WT04]. Self [Ald86, BK12, BLR02, BG17, DT00, DGL87, Gan94, HHH02, HH19, HHK90, HH91, Mie03, Mie05, NRW87, PZ93, Rai97a, Rai09, iS91, Zäh88, Arb91, AC07, Ass08, BDC12, BBCK18, Bol90, BS95b, Chi01, For21, Ger86, GK91, Gra87, GVA86, HK92, HKR18, Hin05, Hin13, HTV12, JY03, Kaj13, Kön93, Kön94, MO86, OTV90, Osa98, PT02, Thu86, TW98, Uel02, Urb90, Wat00, YK88, vdHdHS98, vdH01, vdHS05]. Self-adjointness [DT00]. self-affine [Urb90]. Self-attracting [HH19, Rai97a]. Self-avoiding [HHK90, HH91, NRW87, HH19, HK92, HTV12, Kön93, Uel02, YK88, vdHdHS98, vdH01, vdHS05]. self-consistent [HKR18]. self-decomposable [Ger86, Thu86]. Self-diffusion [DGL87, Osa98]. self-dual [BDC12]. Self-interacting [BLR02, Rai09, BS95b]. Self-intersection [BK12, AC07, Ass08]. Self-intersections [Ald86]. self-organized [For21]. self-repellent [Bol90, HTV12, Kön94]. Self-repelling [BG17, HHH02, TW98]. Self-similar [Mie03, Mie05, PZ93, iS91, Zäh88, Arb91, BBCK18, Chi01, GK91, Gra87, Hin05, Hin13, Kaj13, MO86, OTV90, PT02, Wat00]. Self-similarity [Gan94]. selfaveraging [Kül03]. Selfdecomposable [Pan90, MN98a, Wat00]. Semi [MN98a, Cha92, FJK23, FO14, GT03, GT16, HMY02, Wat00]. semi- [FO14]. semi-circle [Cha92]. semi-circular [GT03, GT16]. semi-linear [HMY02]. semi-norm [FJK23]. Semi-selfdecomposable [MN98a, Wat00]. semicircle [HNY99]. Semicircular [Au22]. semiconvexity [Con24]. semigroup [CS08a]. semigroups [Bar89, BL91, BSC02, Blu05, BRS96, CSS23, Cer99, CZ16, CMG02, Dzi91, FR03, FT19, FNS97, FM89, IR06, Kol11, Pic02, Sta02, Wan97b, Wu04, Uma06]. semilinear [AHR01, Cas01, CWZ95, CMG95, Hai12, Lee90, PZ11, Roz03]. Semimartingale [Kwo95, CM97, DT93, Lus92b, MU90, RW88, RW89, TW93]. Semimartingales [LB08, BOR16, Bes91, ER96, LB95, LM89, Nis97, PS93, RV96]. Semiparametric [BLV12, Cas12, Par90, Par91]. semistable [Luc91]. sens [DD88]. sense [DD88, FNS97, Sim99]. sensitivity [BLZ20, BGS13b, PR20]. separable [Eig86]. separating [Lou22]. separation [Fun95, GK01, Hry98]. sequence [AKTW18, BGdH10, BGdH23, Bra89, CM11b, CP88, CM96, DLP87, Gri87a, HHL88, JPRR18, NM89]. sequences [BPS18, BS14a, Ber87b, Ber90a, Bou90, CHW88, CK08b, Hit88, JMP89, JRdlR08, Kal89, Kal00b, Lal86, MPR11, MW13, MW14, Rio96, Ros87, Sep98, Vie97, Yas02, Yu93]. Sequential [RST15, EM13, KP86b]. series [AF12, Aro89, BNPS18, CT86, DNS22, DvZ04, Eic12, Fuk91, Fuk11, MW16, MW05, PX88, Aro89]. set [BL12b, BH98, Cam87, Dow88, Dud94, Hal88, Hig91, Liu96, LW18, Maj88b, MPV01, MS24b, TW12]. set-valued [Dud94]. sets [Ale87b, ALS20, BMM20, Ber97, Ber99, Bor91, Bru19, BN13, CT11b, CN20, DW93a, DG88, DSC96, Dur09, Fic91, GT24, Gre90, HW86, Hin05, Hu94, IMŞK93, Kal00b, Kry02, Law92, LW90, MP05, MP86, MWW09, NSU19, RST^{+}13, She93, Str86b, Tal88b, VW24, Xia97, Yat13, Zäh86, Zam02]. setting [AV24]. settings [Mon18]. setups [KR14]. several [Gra05, MiS09, Tro95, Wu00]. severely [Gol04]. sewing [BDG21]. Shannon [Bre08, EM20]. Shannon-McMillan [Bre08]. shape [BF88, CG04, GS04, LLT23, Luc14, Wer94a]. shape-from-moments [GS04]. shapes [AG23, DH96, vdBBdH05]. shared [GM94]. Sharp [BR21, CT02, CCM16, DCH24, Dur02, For09, Ham15, HMT23, Hol03, Kle06, LL23, Son04, Stu95, TY23, DWXY23, DCM16, Efr99, Hig93, Hil87b, WZ13]. sharper [GHM12]. Sharpness [ATT18]. sheet [DW93a, DM97, Lac90, RW91]. sheets [FT04, XZ02]. Sherrington [Cat96, Tal98c, Tou98]. Shift [Tho94]. Shift-coupling [Tho94]. shifted [CLM94, KLL94b, LLT^{+}19]. Shifts [ÜZ94a, BLV12, Pri96a, ÜZ94b, ÜZ96].

Shock [Fer92, FF94, FN15]. Shocks [ABL88]. short [AB19, CM96, Sch88b]. short-range [CM96]. shorter [MR07a]. shortest [Gra05]. Shotgun [DL23]. shrinkage [HM97]. Shy [BBC07]. side [Kni01]. sided [AM13, CKS10, Don97, HL21, MR16, PP87, Sav09, Zho88]. Sierpinski [BP88, BB90, CQ24, Ham00, HHK90, HH91, HK92, HHW94, LQ19, Osa90, Osa01, PP91, Shi03, HHH02, HS90b]. sigma [MRT19, MRT20]. signal [PV20]. signature [CF24]. signatures [LQ13]. signed [BNS14, Cek98, Mue98]. significance [Hec86]. similar [Arb91, BBCK18, Chi01, GK91, Gra87, GVA86, Hin05, Hin13, JY03, Jur88, Kaj13, Mie03, Mie05, MO86, OTV90, PZ93, PT02, iS91, Wat00, Zäh88]. similarity [Gan94]. Simple [CS12, DR86, Gol07, Ass08, Bah98, BM19, Bau13, BCF19b, BM03, BB07a, BD88, CLO01, DFP15, Fag22, Fer92, FF94, Kal09, KW16, LY97, MSW21, PR20, PRS16, QRV99, SS18, Swa18, Tal05, Yan22, Yau97]. Simulated [Cat98, Mon18, Rai09]. simulation [Epl86, DM16a]. Simultaneous [Ber92]. simultaneously [HP99]. Sinai [CP03, HPR20]. sine [CPY94, LRV23]. Singer [Léa88]. Single [LLS24, FL95, GZ12a, Tao13a, BGR16]. Singular [CD13a, DR02, DR09, GH19a, Hai12, Aeb93, Ahn23, EK87, GRT20, GMW24, GNW23, KR05, LWX04, LLT^{+}19, LW20, NT86, QSY24, Sch23, Stu93, Tak87, Tru03, Wan18a]. singularities [Tha96]. Singularity [Dur09, Coo17, Hin05, Kry16, OS16, OS22, PP91]. site [BMRS20, HS00, Lac10, PT17]. site-based [PT17]. sites [BES00, Eis97, Nán11]. Six [IMS20]. Size [PPY92, AB15, BBW14, GG11, JN08, JRS20, Kül98, Tei11, vdBJ12]. Size-biased [PPY92, GG11]. sizes [CCFS00, MM13b]. SK [BGS13a, BS24b]. skeleton [HHT21, Sug94]. skew [Eri90, Kal11, Wil87]. Skorohod [BPR07, BPR11, Buc91b, Buc92, Cos92, CL90, HS95b, Imk94, Thi91]. Skorokhod [BCH20, Che93, DR99a, DR99b, Hoo91, JMP89, MR16, Thi91, Tru03]. Slab [Bod05]. SLE [AS11, CN07, Dub06, FS17, FTY21, GHM20, MW17, RZ18, RZ16, SZ10, Zha15, Zha16, Zha19, Zha21]. SLEs [IK13, MS16]. slices [FM19b]. Slightly [Hut24]. slit [BMS02]. slope [AM88, Mil89]. slopes [BCD21]. Slow [Cro13, Rob10, BK04, BB96, CFM20, HSWZ18]. Slowdown [Szn99, FGP10]. slowed [AJV18]. slowly [CM17, Law92]. Small [BH13, FJK23, FR07, Hab18, KM09, MR95, Sav09, SW95, BBT23, BU11, BDE00, BM11b, Day88, FZ99, FT04, FL96b, FK10, FW12, KS95, LLT23, Léa87b, LT88, NN04, Neu90, Péc06a, Péc06b, Ruc88, Tea08, Tei11, TW20, Yos92, Yos93a]. small-noise [Day88]. Small-time [Hab18, BM11b]. smallest [EM97b, LLT^{+}19]. Smoluchovski [Kol10]. smooth [Bar86, CY14, Cap95, FR89, HLW24, KP89, Kry94, LS01a, Li94b, Pic96, Pic10, Web91]. Smoothed [HMP92]. Smoothing [Cer99, CM18, AM13, HM97, MM17, Men16, MW90]. Smoothness [CN88, CKS11, Dri13, BG93]. snake [AD04, AD08, DL97, Le 95a, Ver97]. Snakes [Eva00b, DK95]. snowflake [vdB00]. Sobolev [AL00, CGG07, CGW10, Ces01, CV18, Con24, DS09, GGM05, LN05, Li03, RR00, Shi94a, TY23, Wan97b, Wan97a, Wat93, Wu00, Yau97, Yos99]. social [MST14]. soft [Mas92]. soft-core [Mas92]. Sojourns [HS95c, Ber87c]. solid [LO24]. solid-on-solid [LO24]. solidification [vdBB04]. solitary [RSZ23]. Solution [Don94, HP95, COT19, CL99, DT12, FL19, GM05c, MR23, MSS94, Pap90]. Solutions [AD04, MM17, AR91, AM13, AM95, BRR11, BFHM19, BHW19, CN88, CWZ95, CV98, DHR21, Den00, DMS05, DR14b, DGRZ20, Fil00, FG95, FMS11, GR10, GK96b, HRvR17, HLN13, IKS13, Ker89, KR05, Le 95a, LWXZ12, MZ02, MZ11, MPS19, MP92, Mue98, NS23a, Oss05, RS23, Rut90, SSZ22, Tak87, Tay89, Tri95, Yon97, ZD90, dBD02]. solvable [RSX13]. Solving [MPY94, Pra93]. Some [AD02b, AM08, BR92, BSC02, Bou90, CPY94, DDG02, Del99, FS10, GY05, GK96a, HS94, Hu94, JT17, Kal88, KS91, Kal09, LP11b, Lee90, LS86, MMF92, MS87, Nis97, NUZ90, O’C98, QD91, Ser95, Von95, Yuk87, ADS16, Ans13, Bar21, BM12, Bar00b, BPR10, Bir86, BDG^{+}09, BCH81, BCH88, Cap95, CP13, CIK99, CG90, DT00, FR05, GL14a, GK91, GL07, HJS88, Hsi93, KS95, KLL94a, Kon90, KS88a, LP11c, Lo89, Mór91a, Mou92a, Ndu91, Off14, Péc09, Pic02, RS94, SCZ21, Wat00, vdB94]. sommes [Pet92, Rio00]. sortie [AL94, Mai89]. sorting [GR19]. SOS [CMT15]. soup [LW04a]. soups [vdBCL16]. sources [Mar87]. Souriau [Léa01]. Space [FÜ04, FI07, AK87a, AHR01, BLM24, BKP21, Ber90a, BPR07, BS98b, BMP04, Bor91, BGP24, BCF15, BHW19, Cap95, CDM24, Che90, CLW11, DRZ02, DLTV24, Def16, DDP86, Ein89b, Ein92, EdHM15, FSS10, HSWZ18, HHT21, HNY99, Hol96, HKK23, IR10, Iwa87, JMP89, Jen91, Jou87, KLS86, Kaz88, KMM98, KH03, Kol08, Kre97, LOY98, LP11a, LY87, LW90, LL24, Lie98, LM22, MWZ05, MWZ08, MN87, MS17a, MP91, NUZ90, NU91, NW91, PS87, Pri96a, RAS05, RT02, Sri00, Tal05, ÜZ93, ÜZ96, ÜZ97, ÜZ07, Wat93, Yos93b]. space-filling [MS17a]. space-time [Def16, HKK23]. Spaces [BGV07, ACLP20, AM96, AR89, AS22, BGN91, BR90, BG93, BG97a, BR01, BJ99, CSS23, Cer99, Che92a, CK08a, CKK13, CMG95, CJ94, DR02, DR09, DM22a, Dob88, Ebe97, FS86, GT20, GPW09, GNSS95, Har04, KN03, LPS16, Li03, Luc91, LQ97, Mar87, Müc94, RR00, Shi94a, Thu86, Tre15, Wat93, Wei90]. spaces-the [AR89]. spanned [Mou92c]. Spanning [Mor03, ABS21, AS92, BCK21, BS24a, HN17, HP19, Hut20b, KW15, MNS21, NP22, Sch09]. Sparse [DP15, AKLP23, BR21, Bas23, BC15, BCS18, BCS19, CMW15, CQ20, Gam14, GV16, HK21, LS18, O’C98]. Spatial [BU11, CS09, Lal09, LZ10, SSSY17, ABL12, BK01, CR88, CP07b, DL03a, DLTV24, FM21, GHK22, GHK23, Lah99, Mai16, PPS96, Per92, Reg23]. spatially [Bah98, CdH21, FV00, Léo95]. SPDE [DMS05, HLN13, MSS94, WZ93]. SPDEs [AZ20, CF19, DMP93, FKN24, FO14, GH19a, GMW24, Hau07, Kry94, Kry03, LWXZ12, NP92, PP94, PZ11, Wan18a, Zam02]. special [Por95]. Spectra [Tel90]. Spectral [ABXY24, CF03, GM12, RR89, Ver11, AGKW09, AW15b, AK16, Bas23, BK01, BCGZ22, CM01, CL19, CDK18, Dah90, DKW23, EKZ22, FM19a, GT16, HI86, Hor00, JW24, KM12b, MM90a, MM13a, MS95, Mic99, NSU19, NN04, Rob94, Wu04]. spectrally [CKS11]. spectre [Gol96]. Spectrum [Yur99, CP16, CV10, Cra00, Gol96, IR06, Sch20, Tao13b, Tao13c]. Speed [Aïd14, GGLS20, Mai16, BS08, BGP03, CR88, GR10, PT17]. Speeding [CD20, CD21]. speedup [FGP10]. sphere [BRZ20b, Gru98, KW16, Ndu91, BRZ20a]. spheres [KMTT19, PV20]. spherical [ADG01, LS01a, SZ17, Sub23, Tal06]. spiders [Eva00b]. spiked [BV13, CMW15]. Spiking [RR17]. Spin [Cha10b, IMS20, ABXY24, AC13, AG95a, AG95b, ADG01, BS24b, CMRT08, CMRT15, Cat96, CP13, CDP17, CHL18, DZ95, FW24, FIN99, GKN92, Gan99, Gru96, Gui97, JT17, Kül01, Lar95, LT21, Mac21, Maz99, MN04, PT04, Pan15, SZ17, Sub23, SU93, Tal09, Tal10, Tal16, Tou98]. spin-glass [Gru96]. spinal [FRS23]. Spine [LR21]. spins [Lar95, MN04, Sep95, Tal00b]. Spitzer [Don95]. spline [Fab90, Kou03]. split [Cei12]. split-words [Cei12]. Splitting [BD87, Mie05]. Spohn [IVW18]. spots [BPP04, DF01]. Spread [DS23, vdHS05, vdHdHS07]. spread-out [vdHS05, vdHdHS07]. Spreading [Kal00b]. Spreading-invariant [Kal00b]. Square [DG10, Asp89, BMS08, HM87, HK92, Kry03, LT15, LLT^{+}19, PPS96, SCC19]. square-integrable [Asp89]. squared [Eis03, PSS15]. squares [DMRYZ04, HM87, Oli16]. Stability [AEK19, CK04, DG21b, EM20, Sch99, ASZ09, BD88, CDM24, EHM88, FO14, Li94a, LDS02, SU86, Yu23]. Stabilization [Cer05, LPS16]. Stable [BJMW11, CH93, DeB05, EFL01, GM06, SRMC93, BK08, BM12, BES00, BC06, BD89, BBC03, BDG10, CHW88, CK02a, CKS10, Chr86, CHM86, CH13, CK15b, Day88, DeB04, Dev88, Dzi91, Eis97, GHV90, Gou04, HMT20, Haw98, Hu94, Jan88, JM90, KM88, Kor17, KS88b, KS88c, Las19, LPR97, MiS09, MM90a, Mie03, Mie05, Mij88a, Mir11, Myt02, Nol89, PT02, PZ11, PS15, RR92, Son93, Szu92, Wac12, dHP86]. stable-like [CKS10]. stables [Le 87]. stacks [KMP22]. stage [Ohl89]. Stam [EM20]. standard [CL17]. standardness [CL14]. standing [BCFN89]. star [DF20]. star-scale [DF20]. started [DL01, DL03b]. starting [Müc94]. State [Wan97c, BL00, BCFN89, Hor00, Iwa87, Jen91, Lam02, Mer03, Müc94]. states [AKRT97, BK07, BGP94, CMS15, CDCIV14, FG88, FFL94, GS20, Hig91, Ji89, Mło99, Nán11, NP12, Pan15, Pet90, PP91, SU86, Tal10]. Static [FLT19]. stationarity [BBS21]. Stationary [BBZQC10, BL24a, BFHM19, BP99, CQ20, CM95, CK19, FFL94, GRAS17b, IMS20, Kal14, OTV90, SS87b, SU86, VW24, dHP86, ABF23, AG23, AFZ87, BPS18, BCF19a, BKP21, Ber90a, Boi90, BCFN89, Bra89, BKWW23, BCF93, BF18, BDG^{+}09, CL17, CS89, CV16, CCM16, Chi01, CMMM11, CM96, Ded98, DD05, Erh00, EP97, Eva00a, FKL04, Fer91, FG88, FG95, GKdV89, Gin94, GS07, GS03, Gri87a, GNW23, Guy87, HP94, Hos91a, HHL88, HMO23, IS19, KS94, Kar87, Lal86, LW20, MP88, Nes19, Nie89, Num90, Pit96, Pou88, QSY24, Rho09, Sam87, Sch90, Sha92b, Tan86, TKPK86, Wee88]. stationnaires [AFZ87, Fer91, Guy87]. statistic [Deh86]. Statistical [CIK99, DIT14, RD13, CLL20, DR06, DR07, ER96, EM21, HJL90, Jac90, Kan90, Ruk86]. Statistically [Gra87]. statistics [Ano89a, AJLS15, BC04, BDE00, CL19, DP05, DDP86, DG14, Dun24, DE92, EHM88, FFT10, GJV88, GR19, HJS88, Höp93, Hua24, IR89, JN91, Jan88, KP86b, KV18, Maj06, Mar91, Szu92, Tal16, Yat13]. stay [Car05, VW09]. staying [CMT15]. Steel [DMR11]. Stefan [BD02, NS23a]. Stein [Bar90, Bon20, Cha07, Cha10b, CS01, FSX19a, FSX19b, LU03, NP09, RR03, Sch88b]. Stein-type [LU03]. step [Ber21b, Höp93, MPY94]. step-reinforced [Ber21b]. stepping [SU86]. steps [MP87]. sticks [Roy91]. Sticky [LR04, GM89, Gra95, LL04]. Stieltjes [CMS15]. Stirling [KM22]. stirring [Ham15]. Stochastic [ASS00, AUY08, AR91, ALN99, BCS11, BC01, BW17a, BL03, BS98b, Ble93, BMS05, BCF15, Cer03, Che90, CJ94, CQ02, DM03, Gar95, GMM17, HLØ^{+}93, HKK23, HN09, KS88a, LR88, LS97b, Léa01, LM01, Li94a, Lia93, LQ97, MT95, Myt02, NP88, RW19, RZ09, Sai87, Tan99, TTV03, Uek89, Zha04, AV24, Ale87a, Ans13, AS95, AI98, Aro89, Asp89, AM95, ABR08, ABBP02, BT96, BD02, Bax86, BS88, Bax89, Bax94, BBPS21, BL24a, Ber86a, BG99a, Bla05, Ble89, BJT11, BEGK01, Box89, BFHM19, BH09, BHW19, Buc91b, Buc92, BN94, BP99, BQR00, BDG^{+}09, BDG10, BDG21, CC99, CGRS16, Car87, Cas93, Cer05, CF09, CF19, CHKN18, CP93, CMG95, CV98, CK12, CJKS13, Coo23, Cos92, CP07b, Cra00, CH88a, CS94, Cur16, DT00]. stochastic [DR02, DDG02, DR09, DHR21, DKN09, DV15, DP98, DT10, DG21a, Dit88, DT12, DNS90, DGO20, DK87, EM90, Fag90b, FZ05, FG88, Fer89, Fil00, FG95, FR08, FL19, Fon09, FJK23, FS00, FK10, FV08, Fug88, Fun95, FG19, GG19, GM05a, GKR07, Gu17, GP20, GP93, GK96b, Hai02, HMS11, Hai12, HP91, HRvR17, Hol96, Hou90a, Hou90b, HP95, HMY02, HR95, HS87, IKS13, IMS20, Imk92, IL01, Iwa87, JR95, JMP89, KS94, Ker89, KKM99, Kim04, KM02, Kol13, Kot92, Kot95, KK10, Kry02, KR05, Kry10, Kry14, KKT10, Kuw09, Lac16, LM89, LU03, Li94b, LDS02, LWX04, Lia00, Lie98, LW00, Lus92a, MPY94, MY99, MZ02, Maj88a, Maj06, MN93, Mar10, MPS19, MP87, MR12, MO87, Mor86, MT94, MT11, MP11, Nic16, NZ86]. stochastic [OP89, Oda08, Oel89, Ogu01, Osa12, OT20, Osa22, PP87, PP94, PT99, PS87, Pes94, PZ00, Pic89, Pou88, Pra93, Pri96b, RS94, Rei89b, RW05, Res88, RSZ23, Rom14, RV93, Rut90, Sch86, Sch23, Sch90, ST88, Str86a, Tak87, Tal88a, TZ97, Tsa16, Wan18b, War05, Web91, Wei93, Wen96, Wen99, Wen04, Yon97, Yon08, Zäh98, ZD90, Zha10, Zha20, dA00, dHP86, Thi91]. stochastically [Ald88, FHS12]. stochastique [Asp89, Pra93, Thi91]. stochastiques [Aro89, JMP89, MP87]. Stokes [Fon09, BBPS21, BFHM19, BCF15, CC99, DR14a, FG95, FR08, FL21, LS97b, Mél01, MR12, Oss05, RZ09, Zha10]. stone [SU86]. Stop [PS87]. Stopped [EP97, Eva00a, Kyp00]. Stopping [FF91, MZ87, RSV01, BEES18, COT19, Dal88, Dyn93a, Gre87a, Gre87b, IMŞK93, Kal90, LM17, Mor86, Mor91b, Nag87]. strange [Bir86]. Strassen [Buj86, EM93, GK96a, Gri87b, Gri91a, Mue91a]. strategies [Bar21, RSV01]. Stratified [AMS94]. Stratonovich [LQ13, MN93, ZD90]. Stratonovitch [CV98, CL99, Thi91]. strict [Wan18b]. strictly [DZ19, ELY99]. strings [FKT88]. strip [Gol08, Hua24, Maj92]. Strong [ASS19, Bab89, Bes91, BCH81, BCH88, CHM22, Deh86, Deh88, DFM24, DT12, ER96, Ein88, Hor90, IKS13, KR05, LPR97, Li94b, Lie89, Man91, MN87, Var07, dHHKR91, AA88, BMRS20, Ber87b, Cha12, CLDDMG99, CM88, DS87, ES03, Fil89, FW89, GJV88, GVV22, GK96b, HJS88, Kif13a, Kif13b, LR86, LMW89, MS00, MP91, Nes19, Pfe87, PX88, Reg23, GLZ11, Rio00]. strongly [GS90, Hig86, WZ99]. Stroock [CMM90]. Structural [GL09, KP86a, PZ11]. Structure [Jou87, AMS94, BKK08, Cam87, Cha10a, CHM22, DW93b, DGO20, FM21, GdH99, Gia95, MS21, RRO24, Szn89, Wil88]. structured [CMMM11, PRT13]. structures [AC13, RR03]. Stuck [ETW12]. studies [Bax86]. study [FZ05, LY86, LY87, War05, Wei86b, Ber86b, Bié98]. style [CFV23]. Sub [BB03, Kor17, BSS14, Fuk90, Gol08, Hab18, Zha20]. Sub-exponential [Kor17]. sub-Gaussian [Fuk90]. sub-linear [BSS14, Gol08]. Sub-Markovian [BB03]. sub-Riemannian [Hab18]. subclasses [MiS09]. Subcritical [BLM24, BK12, CCC91, CG09, CR88, DCRT19, NZ17, NST23, Swa18]. subdiffusive [FJL09, HS07]. subelliptic [FV08, GM13]. Subexponential [Klü89, BM08, Wil88]. subgraphs [Ruc88, vdHL10]. subleading [BK17b]. Sublinear [DHS15]. submanifolds [Fra05, Ram88, Ram90]. submap [GW04]. submartingal [Kwo92]. submartingale [BL07]. submatrix [BDN17, GJS21]. subordinate [SV03, Son04]. subordinated [Ose08]. Subordination [Le 18, Sim99, CM96]. subordinator [FT92, Mij88a]. subsampling [Lah99]. subsequence [BP86]. subsequences [AD95, BP15, BNS14, Joh00, LT86, TW01]. subsets [EK23, PT87, vdBJ12]. subshift [Dow88]. subshifts [Häg95, Kif95]. subspace [Tro12]. substituting [IR89]. substitutions [LS87a]. successive [Zai88]. Sudakov [vW97]. sufficiency [Mar91]. Sufficient [CL17, ÜZ07, Wit87b, AJLS15, LLR10]. suitably [BC04]. suite [DLP87]. suites [JMP89, Rio96]. suivant [KPyWmW88]. sum [Cam87, CHM86, Eba86, EM93, EG95, FNS87, GH91, Kar12, Mij88b, Mor86, MP86, Nag87]. sum-free [Cam87]. summable [DM17]. Sums [NST24, ABM05, BJMW11, BPS18, BG07b, Bol86, Bol87, BM01, Chr86, DP15, DS87, EHM88, Ein88, Ein89a, EM97b, FW08, Fuk94, Gri88b, HTV89, Jan88, JM90, Jan10, Jen91, KM86, KM88, Kif13b, KL00, Neu24, Pet92, Pin07, Rio00, Sch88a, TT91, Thr87, Wan94b, Wat08, dHHKR91, Kif13a]. sup [Don94, GKS04, LT00]. sup-norm [Don94, LT00]. sup-norms [GKS04]. Super [BL01, BM05b, DD17, Kop19, AL94, AD02b, ABBP02, BM03, BM05a, Del99, DF01, DL97, DP99, EF98, FL95, FM97, Le 15, Mör01, MP19, Ove93, Ove94, SV99, Sch96, Ser95, Tri91]. Super-Brownian [BL01, BM05b, BM05a, Del99, DF01, DL97, DP99, EF98, FL95, FM97, Le 15, Mör01, MP19, Ove93, Ove94, Sch96, Ser95]. super-Markov [ABBP02]. Super-Ricci [Kop19]. Superbranching [DG88]. superBrownian [CS04]. Superconcentration [DG24]. Superconvergence [BV95]. supercritical [AL09, CR88, CH19, GS89, Hut24, Lac14, PSS20, Sch99, dBD02, vdHL10, vdHHN20]. superdiffusions [She93]. Superdiffusive [CFM20]. Superdiffusivity [LNP96]. supermartingale [Kry02]. supermartingales [Kra96]. Superposition [RXZ20]. superpositions [FR04]. superprocesses [DL03a, Dyn91a, Dyn93b, DK95, EM91, HLN13, Le 93a, Led00, LWX04]. superquadratic [DHB11]. superstable [Geo94]. supersymmetric [MRT19, MRT20]. Support [GZ12a, Bru16, CMM90, GNSS95, MSS94, MP92, ST88, Tri91]. supported [Hil94]. Supporting [Bar00b]. Supporting-points [Bar00b]. suprema [Hal91]. Supremum [Ger86, CK20, Eis90, GH91, Maj06, Web89]. sur-critique [GS89]. sure [BBPS21, Ber90a, BD89, Ber95a, BMP97, Boy19, BMS08, CG07a, CFR93, CS94, DD88, Dup10, EHM88, FHS12, GKM00, LWR92, LDS02, MN93, MP87, Mör00, RAS05, Rio00, Thr87]. surely [Sch89a]. Surface [DP96, Pis96, BW93, Bre08, DV00, NS23a]. surface-order [Bre08]. surfaces [ABR08, BBB90, BW93, BM17b, DDD19, DHP23, DG24, EFL01, HR02]. Survival [AS10, CP07b, BBS14, BGR15, DFSX20, SV86a, Shi97, vdBBdH05, vdHdHS07]. survive [CFK13]. sustainability [CHS22]. Swendsen [BCT12]. switch [TY23]. switching [Duf92, HT10]. symbols [BZ20]. Symmetric [ABBL24, AT86, BP89, CMG02, DeB05, DFP15, DSS14, Gre90, GM06, HS95b, AK87b, AJV18, BES00, BKU10, CFS01, CFKZ08, CKK13, CZ16, DDG02, DeB04, DD05, DCH24, Eis97, Fag22, FW24, FM06, Hou16, Kan00, KSX08, KSX09, KU97, LB08, LM19a, Lie98, MP02, MT87, Müc94, Nag87, OS95, PR20, QRV99, Rai98, Res88, RV93, Stu95, Szu92, Wil87, Zai88, vEH88]. Symmetries [TZ97, Gan99, Kal95, Maz99]. symmetrised [AK08]. Symmetry [BJ20, NS23a, RZ16, Tal00a]. symplectic [Sch05]. Synchronization [BGP14, FGS17]. synchronous [BC02]. synthetic [BSV24]. System [GM89, Bai86, BS95a, BS97b, BFHM19, DGL87, FW20, FG94, FV00, FIN99, Fun91, Gia95, GM94, GdH94, MT11, Qui10, SU93, Tan96, Uch88, dlR95]. système [dlR95]. Systèmes [Mai89, Bou88, Lar95]. Systems [BC06, ANV15, AKRT97, AS05, AS09, BB24, BCF19a, BG86b, Bou88, BCG86, BCF93, BW17b, CRS18, Car87, Cer03, Cha94, CG07a, Che95, CFM20, CG90, CGS95, CK00, CK02b, CF94b, DZ95, DDG02, DKN09, DH22, DK04, EP92, EMS13, FG88, FGM09, FL96b, FW04, FW07, FK10, FFL94, Fuk91, Geo94, GKR07, HV09, Hof86, HD21, KT07, KLZ23, Kol03, KV18, Lar95, LR94, LT21, Léo95, Lig87, LO96, Mai89, Mou92a, Oel89, Pap87, PP94, Rez96, Sep93, Suz24, Tan97, Yan24]. Sznitman [BL00].

tacnode [BL19, ACJvM15]. tagged [JLS09, Uch89]. Tail [BDG^{+}09, Kol13, AHJ21, Bas23, BG24, Erd98, Erd01, GH23, KN03, Kor17, LLT23, Maj88a, Mir11, NTV17, Oli16, OT20, PR08, Won20]. Tail-homogeneity [BDG^{+}09]. tailed [BS24a, BG13a, Men17, MW16, Sch00, vdBN21]. tails [BCK21, BKdS18, Cli86, Din13, DF20, HM07, Rob10]. take [LS19]. Talagrand [BL97, CGW10]. tale [ARS23]. tamed [RZ09]. Taming [Eld20]. Tanaka [FIKP13]. tandem [BBM00]. tangent [Hit88]. TAP [Sub23]. TASEP [BL21, FN15, IS19, Lia22, MR23, PS22]. Taylor [Aro89, BNPS18, DG21a, GM13, LZ24, Sug94]. techniques [HP91]. telegraph [OB04]. Temperature [CL07, ABXY24, CMT15, DS20, Dom17, EKZ22, GT98a, GL99, Mor11, Sam22, SS15, DM16a, Tal00b, Tal01b, Tal01a, Tou98]. temperatures [DZ19]. Temps [Ber90c, AFZ87, Ber89, BW93, DM91a, Eis90, Léa87b, Mat95, NW91, Val91]. tending [KM96, KMZ11]. tension [NS23a]. tensors [DT23]. term [Ber90b, Cer03, CG90, Mue91b, MS93, Mue98]. terminal [AK08, BH08, Rog93]. terms [LB95, Lej01]. tessellations [BR08, BR10, Cha19a, TZ01]. test [DL97, FV00, SCC19]. testing [GP05, LT00, Roh11, TA96]. Tests [Bar21, Bar24, Hec86, Jan95, MMF92, QD91, Ruk86]. tetrads [DSS07]. Th [Fon09, Deh86, Deh88]. their [ADP16, Bar00b, BBCK18, CP10, CMG02, CFG96, DS95, DK04, GQZ88, JY03, Kif95, LR04, Nie89, Nis97, Nor86, Nor93, Pal12, Rio94, SU86, Swa00, Wat87, Xia04, Xia07]. Theorem [Eig86, AB10, AGOZ88, AA88, ACS21, ABBN04, Bal89, Bar86, BL91, BY05, BLM11, BH07, BCF19a, BV95, Ber95a, BPR11, Ber20, Ber21a, BS05, BCG14a, Bob18, BMP97, Bon20, BCC12, Boy19, Bra89, Bre05, BMS08, Buj86, CG09, Car05, Car87, Cas12, CMM90, CC21, CD13b, CS03, CP88, CES24, CV12, DS18, Ded98, Ded10, Ded13, DG98, DL01, DL03b, DG22, DR24, DH23, Dun24, DM91b, Ein89a, Eis90, Eis94, ESZ09, Eri86, Fag91, Fel05, Fra05, Fuk91, Fuk11, GS90, GRS94, Gol07, Gou04, GZ12a, Häg05, HMS11, HW86, HHT21, HLLS23, HLW24, Hou19, HL19, Imk91, JB04, Kar08, KZ97, Kol10, Kou89, KL87a, Lal86, LS00, Léa88, LZ24, Mar08, MP91, Mor87, Mos12, Neu90, NN20, Oel87, PPZ14, Pel12]. theorem [Pen99, PZ08, Pet90, Pfe87, Por86, Pri96a, QRZ03, RR95, Rau88, Rio96, San94a, San94b, Sch88b, Shi06, Soo17, Tho14, Uek89, ÜZ93, Wac12, Yas02, dJ87, dHHKR91, vW97, BGR16, BCG14b, Fuk94, GLZ11, LT88, Pet92, WY93, Tra90]. théoreme [Léa88, Bre05, Eis90, LT88, Pet92, Rau88, Rio96, WY93]. Théorèmes [Bou88, Val91, MP87]. theorems [ANV15, Ale87a, AK87b, ABM05, Bar02, BHS10, BG07b, Bou88, Bre08, BDW21, CG94, CS89, CG07a, Che95, CFR93, CH88b, DG14, DKW23, Don97, DPR22, DJPP13, Eis94, FT87, FS86, GR10, GL14a, GN08, GS07, GT97, GTT98, Gra05, GLP17, GLL20, Gro88, GNSS95, HHW18, HR91, Hor90, Jan06, JW24, KM86, KTT89, KM96, KP86b, Kif10, KV14, KM00, Kot88, Lee90, Lev88, LRW94, MN98a, MP87, MN04, Nic07, Nic08, NRST16, Nie89, Nis97, RR92, Rai98, Rei05, RW05, Ros87, Stu91, Tan86, Val91, Wat87, Wen96, Wen99, Wen04, Wit87a, Yuk87, Zha18b]. theoretic [Yat13]. théorie [Fou98, Lap90]. Theory [Ano98a, DeB05, Ans99, Bab91, Ble89, Box89, CQ23, CIvV03, Cer24, CMM90, CFG96, DK87, Epl86, Fou98, FFT10, GS86, GSW86, Gri21, HHW18, HPR20, Hoo91, IPS13, Kim04, Kry94, Lap90, Leh03, Li03, MS87, MW14, MW16, MS02, Neu24, Sha92a, SV03, Tho14, ÜZ97, VA92, Yos92]. There [HMO23, Coq00]. Thermodynamic [AL09, AGL17, YSA17]. thermostatted [BKM24]. Thick [BM05a]. thin [CD13b]. thinning [DFGGR19]. third [Pap90]. Thomas [CKP14]. Three [Cer24, Cha19b, CCGS91, Coq00, DGRZ20, Fit90, Fon09, HST15, Qui10, Zyg09, dHNOS03]. three-dimensional [Cha19b, dHNOS03]. Three-point [Cer24]. Threshold [FS08, CM01, CL14, Cha16, DWXY23, DL23, DCH24, GHM12, GH19b, HMT23, Hol03, MWW09, Pen96]. thresholding [BCZ^{+}10, Yur11]. thrice [Gru98]. Tight [BCT12]. Tightness [BRZ20b, Muk91, Che90, BRZ20a]. tiling [LT15]. tilings [AG22, Hua24, Joh02, Las19, Pet14, dT07b]. till [Mór91a]. tilts [IVW18]. Time [DGPS07, Dun07, FR95, MNS89, OB04, AB22, AUY08, Ald88, AMP04, ACS21, AL00, AHJ21, ADP16, AFZ87, BGM18, BBS16, BBB90, BRZ20a, BRZ20b, BM03, BGGM13, BG99a, BGP14, Ber89, Ber93b, BW93, BGR15, BMP04, BCS19, Bor91, Bor86a, Bor86b, BCG88, BCF15, BM11b, CMRT15, CQ20, CS20, CT86, CD13b, CHKN18, CV98, COT19, CCFS00, CL91, DHR21, DRZ02, DV02, DG93b, DG93a, DS13, Def16, DJP01, DD16, DHZ19, DG22, Dit90, DS14, DM91a, DR13, DR16, EV21, Eic12, EHJT13, EdHM15, FZ99, Flo90, Föl89, FIN99, FFT10, Fuk04, GY05, GGLS20, Gei05, GM05a, GS07, GM05c, Hab18, HvdH11, Höp88, Hor90, HKK23, HN09, Isc86, IN16, Joh19, Kal90, KR16a, Kim21, Kim23, KMM98, KV08, KO01, KM09, KR05, Kry16, LL23]. time [LOY98, Léa87b, LS24, LS01b, LS15b, Lia22, MW16, Mit86, Mór91a, Mor91b, MW05, Mue91b, Mue98, NRW87, NW91, Off14, PSS20, PS22, PRT13, Pit96, PR08, RAS05, RRO24, RW91, Sav09, SV17, Sta02, Swa00, SV86b, Tho94, Tri95, Ust88, Wee92, Wu04, Yor92, HvdH19]. time-dependent [ACS21, DG22, KO01]. Time-fractional [OB04]. time-frequency [PRT13]. time-homogeneous [EHJT13]. time-inversion [GY05]. time-reversal [COT19]. time-varying [Off14]. times [ADH14, Ald88, AZ20, ADGO17, AC07, Ass08, BB90, BPS18, BK92, BK12, Bel13, BK17b, BGS16, Ber90c, CL04, CKM93, DHHL23, DeB87a, DeB88, Deh88, Dok04, Don89, DM05, Eis90, EP97, Eva00a, Gol08, Gre87a, Gre87b, HV09, HP18, IW94, Kal90, KP86a, KKR88, KLMH06, Lac90, LZ10, Le 15, LR94, Mat95, MW95, Mor86, Nes19, Nol89, Pal12, PS87, PSS15, Pfe87, PR20, Tho88, Val91, War05, Xia97, XZ02]. tip [ABBS13, Zha16]. tips [Has88]. Tits [AS22]. Toda [Cer24]. Toeplitz [Avr88, BZ20, CRZ13, Gin94, GS07, Ino23]. top [BLZ20]. topics [Kry11]. topological [AR89, Li94a]. topologies [BVBE20, ES03]. topology [BJP22, BM15, HS95b]. tori [DDMP19, HvdH11, HvdH19, KSW16]. torus [Bel13, Boy19, HKK23, Sch09]. tossing [HK97]. Total [ABB11, DLP10, DK17, HNY08]. tourne [Wer94b]. Trace [BK08, BCR04, CS23, FS17, PPS96, Shi12, WZ93]. traces [Bry08].

Tractability [MZZ24, EM21]. Tracy [Joh07, LS18]. trajectoires [AFZ87]. trajectories [AB11, BFPP19, CC94, GK01, LN15a, LN15b, Tal98a]. Transfer [KW15, FS88]. transfert [FS88]. transform [AM13, CMS15, Cat96, KT04, Men16, Ngu88]. Transformation [DM91a, Mal95, ÜZ92, ÜZ94b, BF93, Cat98, DG14, GHM20, Imk91, Yas02]. Transformations [LMV98, Nag89, Ver90, Buc91a, Dar88, GM05a, LY88, Lem88]. transforming [CM97]. transforms [CH93, CS03, FT04, Li08, Li11, Li14a, Li14b, Ove94, Wei93]. Transience [FR03, PT17, KZ17, Kül98, PR20]. Transient [Aid08, AS17, AS21, Don89, FGP10, Gol07, HR91, Hut20a, KLL94a, Sab11]. transit [ADGO17]. Transition [EP19, ATT18, BW22, BR21, BBC^{+}04, BCT12, Cer99, CS16, CP10, Cha16, CP03, CM07, DG93b, DDMP19, DL01, DWXY23, DCH24, Gam04, Gre91, Ham15, Hig93, KM00, KS18, Kum93, LN15a, LN15b, MS99, MW01, MT94, Pap87, QRZ03, RS17, Shi03, Sta02, TY03, Wit86]. Transitions [DHHL23, BBN20, Buc21, CdH21, CCGS91, DCM16, GT98a, GdH92, HP19, LRW94]. translated [Mat92]. transmutation [FL96b, FL96c]. Transport [DM22a, Bob18, CH88a, DP98, Div22, DS14, FL21, GT24, HLP24, Kop19, LS20, MMS14, MS24a, NW22, PW20, Mil12]. transport-entropy [MMS14]. Transportation [FÜ04, GLWY09, Aug21, CGW10, Con19, FGM09, GGM05, GL07, HPU04, Liu17]. Transportation-information [GLWY09, Liu17]. transpositions [FLT22]. Transversal [Joh00]. transverse [LT21]. trap [CM17]. trapping [FK16]. traps [CM17, vdBBdH05]. Travelling [Dit90, Bra86]. Tree [BP94, GPW13, ADH14, ABS21, Aid08, Aïd14, AdR17, ACCH23, BPP97, BCK21, BS24a, CCL^{+}24, Cro08, DS20, DKL02, Gan94, Häg96, Has88, HS07, Kre95, LS00, Le 93b, LM24, Mie03, Mie05, MMS12, MWW09, PWW11, SS87a, Sch09]. Tree-indexed [BP94]. tree-like [MMS12]. Tree-valued [GPW13, ADH14]. Trees [DÜ23, LPW13, ADN22, AB19, AS88, AS92, AP00, AMP04, AMP05, Alt89, ADGO17, BGGS18, BFF18, BKMP05, BvdHS18, BR23, BJ99, Bow18, BDW21, CMRT15, Cha10a, CR88, CFK13, CH13, DMR11, DV04, DGPZ02, DL05, DW07, Eva00b, EPW06, FHS12, FKL04, FS08, FHP18, GPW09, GHS20, HS21, Ham15, HL00, HN17, JN08, JRS20, KW15, Kho08, Kor17, LRW23, Le 18, Lig96, LPP96, MNS21, Mic08, MS17a, NP22, PS18, PZ08, RRO24, Sta01, Tak98, Tro21, Zha14]. trends [Kon90]. trials [CP88]. triangle [NRS23]. triangular [CP00, Jan06, NS23b, Pel12, PS13, Sam87]. triangulations [CK15b, Cur16, Sta17]. trichordal [Qia18]. tridiagonal [Pop09]. trigonometric [AF12, BCP19, Fuk91, Fuk11]. trimmed [GJV88, Gri88b]. triple [CG09]. trou [Mic99]. true [HTV12, Luc14, TW98]. truncated [AM88, FI19, GTS84, GTS87, PR05, Wit87b]. Tsetlin [Met13, Pet14]. Tsirel’son [Yor92]. TSP [Yuk95]. tug [DS13, LM17]. tug-of-war [DS13, LM17]. Tukey [Bru19]. tuplewise [Pru98]. turbulence [FP96]. turbulent [DP98]. turns [Wer94b]. Two [Bax86, CKS10, KM96, Zho88, AB22, AHR01, AM13, ARS23, ACM06, AS21, BY86, BBS00, BBS16, BD02, BCK21, BDC12, Bel08, BK17b, BRZ20a, BRZ20b, Bér05, BS24b, BCG86, CF94a, Car86, Cat96, CLM94, CC94, CL91, Din13, DCTT18, DCH24, DS88b, DP99, EZ22, EF98, FS10, Fuk04, FI07, GKdV89, GT24, GHM12, GH19b, GM16, Gu17, HH05, Hig93, HRvR17, Hol03, Imk86, Imk91, Iof95, Joh19, Kar12, Kes86, KZ97, Kis15, LR90, Lip96, Mat95, MZ87, MR16, NT86, Ohl89, PP87, Par90, Ros89, Sav09, Shi92, Thi91, WY93, ZD90]. two-correlations [GKdV89]. two-dimensional [AB22, ACM06, BCK21, BDC12, Din13, DCTT18, DCH24, EZ22, EF98, Fuk04, Hig93, HRvR17, Hol03, Kes86, Ros89, Shi92]. two-parameter [Imk86, Imk91]. two-point [ZD90]. two-sample [Par90]. two-scale [GM16, Gu17]. Two-sided [CKS10, Zho88, AM13, MR16, PP87, Sav09]. two-space [AHR01]. two-stage [Ohl89]. two-time [CL91, Joh19]. two-valued [GKdV89]. type [AEK17, BB09, Bar11, BRR11, BDS93, Bor08, Buc92, DG10, DS88a, DV90, ELP08, FGM09, Fun91, Gin94, GTS84, GTS87, Gri87a, HS21, Häg95, HLLS23, JR88, JZ15, Kal89, Kat15, Kes88, Kif95, Kla09, Kol11, KM09, Kot95, KK10, LLS24, LU03, Léo95, LQ19, Mar91, MPR11, Neu24, Nic07, Nic08, Ose08, Pen99, Pet90, Sha92b, Shi90, Thi91, vdBG13, BG99b]. types [BL12a, CK08a, NT86]. Typical [NRS23, COP93, vW97].

U [ES03]. U-empirical [ES03]. ubiquity [AW15b]. Uhlenbeck [AC87, BCL23, CS08a, CMG02, DMRYZ04, ELY99, FZ99, GRT20, MS92, Mit87, Shi90, Sim11]. UIPT [AR18]. ultraviolet [LRV23]. unbalanced [DM22a]. unbiased [MR12]. unbounded [BC11b, BC11c, BH08, CR95, CT11b, CH19, Dok04, Fag90b, FZ99, Fen94, Wei88]. uncertainty [DGPZ02]. unconditional [Kla09]. unconstrained [CQ24]. underlying [Jen89]. uni [CR93, CSC93]. uni-dimensional [CSC93]. uni-instantaneous [CR93]. unicellular [Cha10a]. unidimensionnelles [CSC93]. Uniform [BG97a, GN08, Jan88, Lev88, Mas92, NM89, PT87, RT86, Bañ87, BCK21, BK92, BBPS21, BPR11, Bor20, CDR86, CS01, DMS05, DGL87, EL21, HN17, HP19, Jir16, LL23, LT15, Le 93b, Lie89, LT86, Lou22, Mas88, MNS21, Mur24, NP22, RST15, Ros88, Sch09, KPyWmW88]. uniform-in-diffusivity [BBPS21]. uniform-in-time [LL23]. uniformes [KPyWmW88]. Uniformity [MS17b, FKT88]. Uniformly [Dub02, BS22, CGM08, DH23, FV08, KH03, Kry14, Tro12]. unimodal [BGR15]. unimodular [LRW23]. Uniqueness [AKRT97, BP87, Bas88, Ber87a, Bru13, CK15a, DF16, GKN92, Häg03, Lim99, Myt98, Rom14, Swa04, Tan97, Ass01, BH00, CD99, CR93, CMG95, FNS97, HP99, HLP24, Mat92, MS24b, MP11, RZ09, TW93, Zam00]. unit [CQ23, Mas88, PPS96]. unitary [AL23, AG22, CDK18, FI19, Kar08, Par22, PR05, Wie02]. univariate [GS15b]. Universal [BCL23, Kül03, Won20, AB15, CFV23, Mat04, Rio98, vH13]. Universality [AEK17, BM17a, HMT20, HV04, Hut20b, Met13, Péc09, RW19, Agg19, BCP19, BvdHS18, BS24a, CES21, CES23, Col05, DG21a, DZ18, EYY12, FG19, GM14, NO15, Swa00]. Universe [AMS94]. universelle [Rio98]. unknown [WHZY07]. unlikely [BGGM13]. unroot [BCZ^{+}10]. unsuitable [Thi91]. Unusual [Hal89]. upcrossings [Le 15]. Upper [Bas23, DM23, Gri91b, Hin13, Hit90, Ort89, CP00, Fal90, Hut21, LLT23, NS23b, PS13, vdHS05]. upper-critical [Hut21]. urn [HS13, Jan06]. useful [Ein87, Kou89]. using [AD04, AD08, Bir86, Dar92, Sch88b, Wan95]. UST [BvE24]. utility [Nut12].

vacant [MS24b, Tei11, TW12]. vacuity [Yur99]. validation [HM87, HMP92]. Validity [Jen89, Lah01]. Valleys [DGPS07]. valuation [FMS11]. value [CS23, CCFS00, Fle12, FFT10, GSW86, HLØ^{+}93, JR88, KMM98, LO24, LLT^{+}19, Mou92a, NN04, Pap90, PZ98, Rog93, Shi12, ZD90]. Valued [BG86b, ADH14, Ber90a, Cha92, Cra17, CL97c, CS94, CM88, Dar88, DDP86, Dud94, DKS94, Ein88, GKdV89, GPW13, Hig86, Isc86, Jia09b, KS91, KM02, Kol03, KS88a, Kot92, Kry10, Kry11, Le 93a, LWR92, MWZ05, MWZ08, MT87, MN87, MP91, Nis97, Per92, Shi94a, Wan97c, vV89]. values [Ahn23, AK10, Cer99, DW90, KL87a, KL87b, Lin87, MR95, Naj18]. vanishing [BM11b, SS15]. Varadhan [CMM90, LR90, Mue91a]. Variable [HM88a, HM92, HS95b, Imk96, Kol11, Kry00, MS90, PS16, PS93]. variables [AGKW09, AT86, Avr88, BJMW11, BV95, Ber90a, BP86, BD89, Chr86, CM88, DD88, DG93c, Don95, DLP87, Dub02, Dud94, Ein89a, EM97b, FT87, GS90, HS88, HL18, HSWZ18, Hil87a, Hit88, HTV89, Jen91, Kar07, KM86, KM88, KH03, MWZ05, MWZ08, MT87, MN87, MP91, Pin07, Röm96, Sas86, Sch89a, Thi91, Wit87a, ZD90]. Variance [CY14, CY15, ABF23, BCP19, BLS87, BLBE92, BJMW11, Ber87c, DHS15, DM16b, EG95, FHS90, HM88b, Hös91b, Hös96, MP88, Web89]. variances [Win91]. variation [Bou90, BNS14, DLP10, KLS86, Lov95, PX88, Ver90]. Variational [Zha07, AV24, DRZ02, GdH94, HN14, MT20, Sta04a]. Variations [Guy87, TZ97, Yan22, Guy87]. varying [BM08, BPS18, CM17, Hös96, LQ97, MM13b, MW13, MW14, Off14]. vector [AM96, AR89, CSS23, HLO23, Jia09b, KS91, Luc91, LQ97, Tha91, TKPK86]. vector-valued [Jia09b, KS91]. vectors [Bol86, Bol87, CCK15, EK96, Fuk90, Ger86, KL00, Pan90, Sch00, Wan94b]. version [AP14, BH07, Box89, CC05, CV12, Rai09]. versions [CD99, CG07a, Lie98]. versus [CH93, vdBB04]. Vertex [DV04, IMS20, Pem92, BSS14, DV02, Kor22, MRT19, MRT20]. Vertex-reinforced [DV04, Pem92, DV02, MRT19, MRT20]. via [AR91, AGM03, AJJ22, BBM99, BK01, Blu05, BdT10, BdTR19, BCZ^{+}10, CCKR07, CS01, DNS22, DG21a, DMX21, Dok04, DFM24, Don94, DM16b, DK87, EM20, Fag22, FT04, Gaj89, GH23, GG11, GL09, GS15a, GR19, GV16, GK96b, HP18, HPU04, HM24, HS87, IR06, Imk96, Kal11, LT15, LM09, LM17, Lin87, Liu22a, LMW89, LW18, Mil12, Pet14, RB03, RT02, Ust88, Yos92].

Viability [BQR00]. Viot [BBF12, Han02, Kwa24, Sta02]. virtually [AG23]. viscous [CF19, Rom14]. visibility [Wie86]. visited [BES00, Eis97, FR92, HS00, Maj88b, Nán11]. Vlasov [CM18, HLLS23, BR24, DH22, Kot95, KK10]. Voiculescu [AG97]. voisinage [CR86a]. Volterra [CP07b, Yon08]. Volume [BW18, Cro08, Ale02, BLM24, CV12, HvdH11, HvdH19, JR08, vdBB94]. volumes [Liu22a]. Voronoi [DCRT19, HS11b]. vortex [FG02]. vorticity [FL21]. voter [BCG88, MT95]. VRJP [LST20]. vs [DF16, Kül98]. vu [GS89].

w.r.t [Müc94]. waiting [Mór91a]. wald [Kal89]. wald-type [Kal89]. Walk [DGPS07, ABB11, Aïd14, AdR17, ABK08, AHJ21, AC07, Ass08, AS17, AS21, BC17, BS08, BSS14, Bel98, BLV16, BM03, BLQ23, BGP03, BB07a, BD14, Bia91, BMP97, BCS18, Boy19, CKP14, CGZ00, CP03, CGZ03, CP00, CS12, CFK13, Cro13, CS23, Dav90, Dav99, DR86, DW21, DFSX20, Don89, DG93c, DR11, DKL02, Fuk04, Gol08, GdH92, GKZ93, GH20, HK01, HHH02, HH19, HK92, Hou16, HS07, IR10, KPyWmW88, Kön93, Kön94, Kyp00, LS15a, LS24, Maj88b, MR03, MR09, MM13b, MS17b, Nán11, NS23b, NRW87, Pem92, PS13, PSS20, PT17, PPZ99, PR20, PRS16, Rod19, Rol03, Sch09, Shi12, Sta04b, Ste97, Szn99, Uel02, Zha14, vdBCL16, vdHS05, vdHH10, vdHHN20]. walker [Vir02]. walkers [TY03]. Walks [BMS02, DSC96, Dun07, Kem05, AB22, ABCM20, Aid08, Ald86, ACS21, AL00, AS22, ADP16, BC11b, BC11c, Bau16, BK12, BP94, BS19, Ber19, Ber01a, Ber21b, BSV24, BMP04, Bol90, BZ07, BFO09, BO09, BG13b, Bor03, Bow18, BS95b, CR14, Car05, CK19, CH19, Che18, CD90, DH91, DGPZ02, DHZ19, DG22, DKW23, DH96, Don95, DM05, DM22b, DE92, ETW12, Ers19, FHS12, FM06, For09, FT19, Fre19, FGP10, FK16, GW86, Gol07, GLZ11, GLL20, GS03, Gre90, GVV22, GZ12b, HV16, Hil94, HTV12, HV04, HL00, Hut20a, IVW18, Jaw95, KM22, KTT89, KM96, KLMH06, KMP22, KSS86, LZ10, Lig96, LRW94, LPP96, Mad89, MR07b, Muk93, Nes19, NZ17, NZ21, PZ08, PSS15, Por95, RS12, RAS05, RAS14, Sab11, SCZ21]. walks [SS87a, SS04, Szn02, Szn09, Szn17, Tak98, Tel89, Tót86, VW09, Wac12, Wil97, XZ13, YK88, Yil11, Zer05, Zyg09, vdHdHS98, vdH01]. wall [ABK08, BG02b, CMT15, Fun03, Zam04]. Wang [BCT12]. war [DS13, LM17]. Wasserstein [Bon20, CDM24, CSS23, DFM24, DS09, FSS10, FSX19a, FSX19b, FG15, LL17, LL24]. wasted [Tal05]. Watanabe [Yos92]. Watson [ADN22, Aid08, Aïd14, AdR17, Bow18, CFK13, DGPZ02, FW08, JRS20, Kor17, LRW23, LS00, LPP96, PZ08]. Watts [Dub06]. Wave [FL96c, Aeb93, Bra86, CN88, CK12, PZ00]. wavelet [ASS00, BCZ^{+}10, Chi01]. Waves [KZ98, Dit90, RSZ23, Aro89]. Weak [AMP05, BK93, CGG07, Con24, DFS20, FG19, GJV88, HRvR17, Kol08, KU97, Ogu01, Ose08, SK86, SK87, Swa07, Wan03, vdHdHK03, Ale98, BJP22, BD89, BB03, CHM22, CLW11, DV00, Fil00, GR10, Li11, Li14b, MZ11, MZ87, Mön24, MMS12, Yu93, DLP87]. Weakly [Kül01, Lem88, BFO09, BW17b, Dit90, Dub02, Fuk91, HL18, KM09, MPR11, Rio96, Wan94b, YK88, Yos99, vdH01]. wedge [DS13]. wedges [Mou90]. Weibull [JR88, Mar91]. weight [BSS14, DÜ23, HM24]. Weighted [Chr86, Gei05, KM88, Neu24, Ale87b, ADS16, BM08, EHM88, GKS04, Isc86, KM86, Thr87, Yuk95]. weights [ACS21, BU11, DNS18]. welding [RZ16]. Well [Yon08, DLTV24]. Well-posedness [Yon08, DLTV24]. Wellposedness [STZ12]. Wentzell [Kry11]. Wetting [Vel09, DGZ05]. Weyl [BJ02, Def16, DF20, Yas02]. where [Erh00]. which [Bra89, GY05, Gra05, HM87, Hal88, LS89, Mit86, Wer94b]. White [DMP93, FL19, Mat94, NP92, AHR01, BCF15, CV98, DLTV24, GP00, HJM20, HLØ^{+}93, HKK23, KSvdVvZ16, MP11, Oua22, Spe90, Tri95]. Whittaker [ABO23, BCF18, IMS20]. Whittle [GS90]. whole [MS17a, Zha15]. whole-plane [MS17a, Zha15]. Wichura [BM11b]. Wick [BN94]. wide [Sim99]. Widom [Joh07, LS18]. width [HM88a, HM92, Mül89]. Wiener [ASS19, BT03, BLM24, Cap95, Car86, CF24, CM97, CFR87, Deh88, DL97, EL87, FSS10, FÜ04, Föl89, FI07, Gri91b, Gri95, HK01, Har04, Imk92, IW94, Imk94, Imk96, KH03, Kol08, Li92, MWZ05, MWZ08, MM90b, McG89, MO86, MO87, MS24b, NP09, NZ90, NUZ90, NU91, NW91, Ort89, Rév92, Slu91, Sug94, Tre15, ÜZ92, ÜZ93, ÜZ94a, ÜZ94b, ÜZ95, ÜZ96, ÜZ97, ÜZ07, VW24, Wat93, Wei86b, WY93, Yos93b, vdBT91, vdBB94]. Wiener-itô [Slu91]. Wigner [Agg19, AEK17, AG97, Au22, BS98b, BLZ20, CMS15, Cha92, CES23, EYY12, HK20, KY13, LLS24, LS16, RS93, Tao13a]. Wigner-type [AEK17, LLS24]. wild [Mam92]. Wilks [BM11a]. winding [BW17a, BF18]. windings [EFL01]. window [HM88a, HM92]. Wired [Mor03]. Wishart [Bry08]. within [Gei05, Pen23]. without [AP91, BF11, GS86, Iha90, MS99, Men18]. Wold [CHW88, MP88]. word [Gru98]. words [BGdH10, BGdH23, Cei12, CG04, TW01]. work [Gri21]. wreath [EZ22]. Wright [GKW01, Tri95]. wrong [KV08].

Yor [Mat95]. Young [DŚ19].

Z [BP89]. Zariski [GG96]. Zeno [Ste97]. Zernike [CIvV03]. Zero [CL07, DS20, Mat94, Mor11, And08, AL09, AGL17, Ass05, BMRS20, BLS87, BL12b, BJL17, BMS08, CP07a, Dom17, Erh00, GRS94, HS94, JLS09, KMZ11, KZ17, LW18, Mik04, Mor86, Nag87, Pes02, Sam22, Sav09]. zero-noise [Mik04]. zeroone [HS94]. zero-range [AGL17, Ass05, BMRS20, CP07a, JLS09]. zeroset [LW18]. zero-sum [Nag87]. Zero-temperature [DS20, Mor11]. zeroes [ABF23, BNPS18]. zeros [Bou10, FI19, KK14, Kar13, Mal95, Reg23, Mal95]. zeta [Bou10, Kar13, Naj18].

References

Anderson:1988:SRT

[AA88] Kevin K. Anderson and Krishna B. Athreya. A strong renewal theorem for generalized renewal functions in the infinite mean case. Probability Theory and Related Fields, 77(4):471–479, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00959611.

Attal:2007:CRP

[AB07] Stéphane Attal and Alexander C. R. Belton. The chaoticrepresentation property for a class of normal martingales. Probability Theory and Related Fields, 139(3–4):543–562, November 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-006-0052-z.

Aistleitner:2010:CLT

[AB10] Christoph Aistleitner and István Berkes. On the central limit theorem for f(n_{k}x). Probability Theory and Related Fields, 146(1–2):267–289, January 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0190-6.

Atar:2011:NOT

[AB11] Rami Atar and Amarjit Budhiraja. On near optimal trajectories for a game associated with the ∞-Laplacian. Probability Theory and Related Fields, 151(3–4):509–528, December 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-010-0306-7.

Arratia:2015:BSB

[AB15] Richard Arratia and Peter Baxendale. Bounded size bias coupling: a gamma function bound, and universal Dickmanfunction behavior. Probability Theory and Related Fields, 162(3– 4):411–429, August 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0572-x.

Addario-Berry:2019:MTS

[AB19] Louigi Addario-Berry. Most trees are short and fat. Probability Theory and Related Fields, 173(1–2):1–26, February 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0829-x.

Abe:2022:EPT

[AB22] Yoshihiro Abe and Marek Biskup. Exceptional points of twodimensional random walks at multiples of the cover time. Probability Theory and Related Fields, 183(1–2):1–55, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01113-4.

Addario-Berry:2011:TPK

[ABB11] L. Addario-Berry and N. Broutin. Total progeny in killed branching random walk. Probability Theory and Related Fields, 151(1–2):265–295, October 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0299-2.

Addario-Berry:2012:CLC

[ABBG12] L. Addario-Berry, N. Broutin, and C. Goldschmidt. The continuum limit of critical random graphs. Probability Theory and Related Fields, 152(3–4):367–406, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-010-0325-4.

Addario-Berry:2024:SCM

[ABBL24] Louigi Addario-Berry, Erin Beckman, and Jessica Lin. Symmetric cooperative motion in one dimension. Probability Theory and Related Fields, 188(1–2):625–666, February 2024. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-023-01244-2.

Artstein:2004:RCE

[ABBN04] Shiri Artstein, Keith M. Ball, Franck Barthe, and Assaf Naor. On the rate of convergence in the entropic central limit theorem. Probability Theory and Related Fields, 129(3):381–390, July 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0329-4.

Athreya:2002:DSD

[ABBP02] S. R. Athreya, M. T. Barlow, R. F. Bass, and E. A. Perkins. Degenerate stochastic differential equations and super-Markov chains. Probability Theory and Related Fields, 123(4):484–520, August 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400100191.

Aidekon:2013:BBM

[ABBS13] E. Aïdékon, J. Berestycki, É. Brunet, and Z. Shi. Branching Brownian motion seen from its tip. Probability Theory and Related Fields, 157(1–2):405–451, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0461-0.

Addario-Berry:2020:HRW

[ABCM20] L. Addario-Berry, H. Cairns, and R. Mitchell. Hipster random walks. Probability Theory and Related Fields, 178(1–2): 437–473, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-020-00980-z.

Andres:2013:IPR

[ABDH13] S. Andres, M. T. Barlow, J.-D. Deuschel, and B. M. Hambly. Invariance principle for the random conductance model. Probability Theory and Related Fields, 156(3–4):535–580, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-012-0435-2.

Assaf:2023:AFV

[ABF23] Eran Assaf, Jeremiah Buckley, and Naomi Feldheim. An asymptotic formula for the variance of the number of zeroes of a stationary Gaussian process. Probability Theory and Related Fields, 187(3–4):999–1036, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01218-4.

Amir:2008:ERW

[ABK08] Gideon Amir, Itai Benjamini, and Gady Kozma. Excited random walk against a wall. Probability Theory and Related Fields, 140(1–2):83–102, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0058-1.

Arguin:2013:EPB

[ABK13] Louis-Pierre Arguin, Anton Bovier, and Nicola Kistler. The extremal process of branching Brownian motion. Probability Theory and Related Fields, 157(3–4):535–574, December 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0464-x.

Andjel:1988:SAE

[ABL88] E. D. Andjel, M. D. Bramson, and T. M. Liggett. Shocks in the asymmetric exclusion process. Probability Theory and Related Fields, 78(2):231–247, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00322020.

Angel:2012:GDS

[ABL12] Omer Angel, Nathanaël Berestycki, and Vlada Limic. Global divergence of spatial coalescents. Probability Theory and Related Fields, 152(3–4):625–679, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0332-5.

Arous:2005:LTS

[ABM05] Gérard Ben Arous, Leonid V. Bogachev, and Stanislav A. Molchanov. Limit theorems for sums of random exponentials. Probability Theory and Related Fields, 132(4):579–612, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-004-0406-3.

Arista:2023:MWP

[ABO23] Jonas Arista, Elia Bisi, and Neil O’Connell. Matrix Whittaker processes. Probability Theory and Related Fields, 187(1– 2):203–257, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01210-y.

Atar:2008:DSD

[ABR08] Rami Atar, Amarjit Budhiraja, and Kavita Ramanan. Deterministic and stochastic differential inclusions with multiple surfaces of discontinuity. Probability Theory and Related Fields, 142(1–2):249–283, September 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0104-z.

Addario-Berry:2021:GMS

[ABS21] Louigi Addario-Berry and Sanchayan Sen. Geometry of the minimal spanning tree of a random 3-regular graph. Probability Theory and Related Fields, 180(3–4):553–620, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01071-3.

Adhikari:2024:SGE

[ABXY24] Arka Adhikari, Christian Brennecke, Changji Xu, and Horng-Tzer Yau. Spectral gap estimates for mixed p-spin models at high temperature. Probability Theory and Related Fields, 189(3– 4):879–907, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-024-01261-9.

Augeri:2023:CCP

[ABZ23] Fanny Augeri, Raphael Butez, and Ofer Zeitouni. A CLT for the characteristic polynomial of random Jacobi matrices, and the GβE. Probability Theory and Related Fields, 186(1–2):1– 89, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01194-9.

Antoniadis:1987:EEI

[AC87] Anestis Antoniadis and Rene Carmona. Eigenfunction expansions for infinite dimensional Ornstein–Uhlenbeck processes. Probability Theory and Related Fields, 74(1):31–54, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01845638.

Asselah:2003:LDB

[AC03] Amine Asselah and Fabienne Castell. Large deviations for Brownian motion in a random scenery. Probability Theory and Related Fields, 126(4):497–527, August 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0265-3.

Asselah:2007:RWR

[AC07] Amine Asselah and Fabienne Castell. Random walk in random scenery and self-intersection local times in dimensions d ≥ 5. Probability Theory and Related Fields, 138(1–2):1–32, May 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-006-0014-5.

Arguin:2013:ROS

[AC13] Louis-Pierre Arguin and Sourav Chatterjee. Random overlap structures: properties and applications to spin glasses. Probability Theory and Related Fields, 156(1–2):375–413, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-012-0431-6.

Auffinger:2015:PPM

[AC15] Antonio Auffinger and Wei-Kuo Chen. On properties of Parisi measures. Probability Theory and Related Fields, 161(3–4): 817–850, April 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0563-y.

Aldous:2023:PIB

[ACCH23] David Aldous, Alice Contat, Nicolas Curien, and Olivier Hénard. Parking on the infinite binary tree. Probability Theory and Related Fields, 187(1–2):481–504, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01189-6.

Adler:2015:TGM

[ACJvM15] Mark Adler, Sunil Chhita, Kurt Johansson, and Pierre van Moerbeke. Tacnode GUE-minor processes and double Aztec diamonds. Probability Theory and Related Fields, 162(1–2): 275–325, June 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0573-9.

Ahidar-Coutrix:2020:CRE

[ACLP20] A. Ahidar-Coutrix, T. Le Gouic, and Q. Paris. Convergence rates for empirical barycenters in metric spaces: curvature, convexity and extendable geodesics. Probability Theory and Related Fields, 177(1–2):323–368, June 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-019-00950-0.

Arous:2006:ATD

[ACM06] Gérard Ben Arous, Jirí Cerný, and Thomas Mountford. Aging in two-dimensional Bouchaud’s model. Probability Theory and Related Fields, 134(1):1–43, January 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0408-1.

Andres:2021:QLL

[ACS21] Sebastian Andres, Alberto Chiarini, and Martin Slowik. Quenched local limit theorem for random walks among timedependent ergodic degenerate weights. Probability Theory and Related Fields, 179(3–4):1145–1181, April 2021. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-021-01028-6.

Aldous:1995:HIP

[AD95] D. Aldous and P. Diaconis. Hammersley’s interacting particle process and longest increasing subsequences. Probability Theory and Related Fields, 103(2):199–213, ???? 1995. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01204214.

Aaronson:2002:GEG

[AD02a] Jon Aaronson and Manfred Denker. Group extensions of Gibbs– Markov maps. Probability Theory and Related Fields, 123 (1):28–40, May 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400100173.

Abraham:2002:SPE

[AD02b] Romain Abraham and Jean-François Delmas. Some properties of the exit measure for super Brownian motion. Probability Theory and Related Fields, 122(1):71–107, January 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100181.

Abraham:2004:SNC

[AD04] Romain Abraham and Jean-François Delmas. Solutions of Δu = 4u^{2} with Neumann’s conditions using the Brownian snake. Probability Theory and Related Fields, 128(4):475–516, April 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0302-2.

Abraham:2008:FAL

[AD08] Romain Abraham and Jean-François Delmas. Fragmentation associated with Lévy processes using snake. Probability Theory and Related Fields, 141(1–2):113–154, May 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0081-2.

Auffinger:2013:DEP

[AD13] Antonio Auffinger and Michael Damron. Differentiability at the edge of the percolation cone and related results in firstpassage percolation. Probability Theory and Related Fields, 156 (1–2):193–227, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0425-4.

Aidekon:2022:GFP

[AD22] Elie Aïdékon and William Da Silva. Growth-fragmentation process embedded in a planar Brownian excursion. Probability Theory and Related Fields, 183(1–2):125–166, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01119-y.

Arous:2001:ASS

[ADG01] G. Ben Arous, A. Dembo, and A. Guionnet. Aging of spherical spin glasses. Probability Theory and Related Fields, 120(1):1–67, May 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/PL00008774.

Amini:2017:ELT

[ADGO17] Omid Amini, Luc Devroye, Simon Griffiths, and Neil Olver. Explosion and linear transit times in infinite trees. Probability Theory and Related Fields, 167(1–2):325–347, February 2017. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0683-z;   http://link.springer.com/
article/10.1007/s00440-015-0683-z.

Abraham:2014:ETI

[ADH14] Romain Abraham, Jean-François Delmas, and Patrick Hoscheit. Exit times for an increasing Lévy tree-valued process. Probability Theory and Related Fields, 159(1–2):357–403, June 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-013-0509-9.

Abraham:2022:GRG

[ADN22] Romain Abraham, Jean-François Delmas, and Michel Nassif. Global regime for general additive functionals of conditioned Bienaymé–Galton–Watson trees. Probability Theory and Related Fields, 182(1–2):277–351, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01095-
9.

Attal:2016:COR

[ADP16] S. Attal, J. Deschamps, and C. Pellegrini. Complex obtuse random walks and their continuous-time limits. Probability Theory and Related Fields, 165(1–2):65–116, June 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
015-0627-7.

Aidekon:2017:SLR

[AdR17] Elie Aïdékon and Loïc de Raphélis. Scaling limit of the recurrent biased random walk on a Galton–Watson tree. Probability Theory and Related Fields, 169(3–4):643–666, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-016-0739-8.

Andres:2016:HIW

[ADS16] Sebastian Andres, Jean-Dominique Deuschel, and Martin Slowik. Harnack inequalities on weighted graphs and some applications to the random conductance model. Probability Theory and Related Fields, 164(3–4):931–977, April 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-015-
0623-y.

Abramson:2014:LMB

[AE14] Joshua Abramson and Steven N. Evans. Lipschitz minorants of Brownian motion and Lévy processes. Probability Theory and Related Fields, 158(3–4):809–857, April 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0497-9.

Aebi:1993:DSD

[Aeb93] Robert Aebi. Diffusions with singular drift related to wave functions. Probability Theory and Related Fields, 96(1):107–121, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01195885.

Ajanki:2017:UGW

[AEK17] Oskari H. Ajanki, László Erdős, and Torben Krüger. Universality for general Wigner-type matrices. Probability Theory and Related Fields, 169(3–4):667–727, December 2017. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s00440-
016-0740-2.pdf;   http://link.springer.com/article/
10.1007/s00440-016-0740-2.

Ajanki:2019:SMD

[AEK19] Oskari H. Ajanki, László Erdős, and Torben Krüger. Stability of the matrix Dyson equation and random matrices with correlations. Probability Theory and Related Fields, 173 (1–2):293–373, February 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0835-z;
 http://link.springer.com/content/pdf/10.1007/s00440-
018-0835-z.pdf.

Aistleitner:2012:LIL

[AF12] Christoph Aistleitner and Katusi Fukuyama. On the law of the iterated logarithm for trigonometric series with bounded gaps. Probability Theory and Related Fields, 154(3–4):607–620, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-011-0378-z.

Affentranger:1988:PNI

[Aff88] F. Affentranger. Pairs of non-intersecting random flats. Probability Theory and Related Fields, 79(1):47–49, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00319102.

Azais:1987:ADT

[AFZ87] J. M. Azaïs and D. Florens-Zmirou. Approximation du temps local des processus gaussiens stationnaires par régularisation des trajectoires. (French) [Approximation of the local time of stationary Gaussian processes by regularization of paths]. Probability Theory and Related Fields, 76(1):121–132, September 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00390279.

Arous:1995:LDLa

[AG95a] G. B. Arous and A. Guionnet. Large deviations for Langevin spin glass dynamics. Probability Theory and Related Fields, 102 (4):455–509, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01198846.

Arous:1995:LDLb

[AG95b] G. Ben Arous and A. Guionnet. Large deviations for Langevin spin glass dynamics. Probability Theory and Related Fields, 103(3):431, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01195482.

Arous:1997:LDW

[AG97] G. Ben Arous and A. Guionnet. Large deviations for Wigner’s law and Voiculescu’s non-commutative entropy. Probability Theory and Related Fields, 108(4):517–542, August 1997. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050119.

Asselah:2014:LBF

[AG14] Amine Asselah and Alexandre Gaudillière. Lower bounds on fluctuations for internal DLA. Probability Theory and Related Fields, 158(1–2):39–53, February 2014. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0476-6.

Aggarwal:2022:GUE

[AG22] Amol Aggarwal and Vadim Gorin. Gaussian unitary ensemble in random lozenge tilings. Probability Theory and Related Fields, 184(3–4):1139–1166, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-022-01168-3.

Auffinger:2023:ASS

[AG23] Antonio Auffinger and Christian Gorski. Asymptotic shapes for stationary first passage percolation on virtually nilpotent groups. Probability Theory and Related Fields, 186(1–2):285– 326, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01196-7.

Aggarwal:2019:BUG

[Agg19] Amol Aggarwal. Bulk universality for generalized Wigner matrices with few moments. Probability Theory and Related Fields, 173(1–2):375–432, February 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-018-0836-y.

Aizenman:2009:BDR

[AGKW09] Michael Aizenman, François Germinet, Abel Klein, and Simone Warzel. On Bernoulli decompositions for random variables, concentration bounds, and spectral localization. Probability Theory and Related Fields, 143(1–2):219–238, January 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0125-7.

Armendariz:2017:MCZ

[AGL17] Inés Armendáriz, Stefan Grosskinsky, and Michail Loulakis. Metastability in a condensing zero-range process in the thermodynamic limit. Probability Theory and Related Fields, 169(1–2): 105–175, October 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-016-0728-y.

Albeverio:2003:RBE

[AGM03] S. Albeverio, G. Guatteri, and S. Mazzucchi. A representation of the Belavkin equation via Feynman path integrals. Probability Theory and Related Fields, 125(3):365–380, March 2003. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
002-0231-5.

Andersen:1988:CLT

[AGOZ88] Niels T. Andersen, Evarist Giné, Mina Ossiander, and Joel Zinn. The central limit theorem and the law of iterated logarithm for empirical processes under local conditions. Probability Theory and Related Fields, 77(2):271–305, February 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00334041.

Adhikari:2020:DBM

[AH20] Arka Adhikari and Jiaoyang Huang. Dyson Brownian motion for general β and potential at the edge. Probability Theory and Related Fields, 178(3–4):893–950, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00992-9.

Angel:2021:TBR

[AHJ21] Omer Angel, Tom Hutchcroft, and Antal Járai. On the tail of the branching random walk local time. Probability Theory and Related Fields, 180(1–2):467–494, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01014-4.

Ahn:2022:FBJ

[Ahn22] Andrew Ahn. Fluctuations of β-Jacobi product processes. Probability Theory and Related Fields, 183(1–2):57–123, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01109-0.

Ahn:2023:ESV

[Ahn23] Andrew Ahn. Extremal singular values of random matrix products and Brownian motion on GL(N, C). Probability Theory and Related Fields, 187(3–4):949–997, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01217-5.

Albeverio:2001:TSD

[AHR01] Sergio Albeverio, Zbignew Haba, and Francesco Russo. A twospace dimensional semilinear heat equation perturbed by (Gaussian) white noise. Probability Theory and Related Fields, 121(3): 319–366, November 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400100153.

Arnold:1998:NFS

[AI98] Ludwig Arnold and Peter Imkeller. Normal forms for stochastic differential equations. Probability Theory and Related Fields, 110(4):559–588, May 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050159.

Aidekon:2008:TRW

[Aid08] Elie Aidékon. Transient random walks in random environment on a Galton–Watson tree. Probability Theory and Related Fields, 142(3–4):525–559, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0114-x.

Aidekon:2014:SBR

[Aïd14] Elie Aïdékon. Speed of the biased random walk on a Galton– Watson tree. Probability Theory and Related Fields, 159(3–4): 597–617, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0515-y.

Aru:2022:DIM

[AJJ22] Juhan Aru, Antoine Jego, and Janne Junnila. Density of imaginary multiplicative chaos via Malliavin calculus. Probability Theory and Related Fields, 184(3–4):749–803, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01135-y.

Ay:2015:IGS

[AJLS15] Nihat Ay, Jürgen Jost, Hông Vân Lê, and Lorenz Schwachhöfer. Information geometry and sufficient statistics. Probability Theory and Related Fields, 162(1–2):327–364, June 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0574-8.

Avena:2018:ELS

[AJV18] L. Avena, M. Jara, and F. Völlering. Explicit LDP for a slowed RW driven by a symmetric exclusion process. Probability Theory and Related Fields, 171(3–4):865–915, August 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0797-6.

Akcoglu:1987:NMD

[AK87a] M. A. Akcoglu and U. Krengel. Nonlinear models of diffusion on a finite space. Probability Theory and Related Fields, 76 (4):411–420, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00960065.

Andjel:1987:PET

[AK87b] Enrique D. Andjel and Claude P. Kipnis. Pointwise ergodic theorems for the symmetric exclusion process. Probability Theory and Related Fields, 75(4):545–550, August 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00320333.

Adams:2008:LDM

[AK08] Stefan Adams and Wolfgang König. Large deviations for many Brownian bridges with symmetrised initial-terminal condition. Probability Theory and Related Fields, 142(1–2):79– 124, September 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00440-007-0099-5.pdf.

Arous:2010:FPP

[AK10] G. Ben Arous and V. Kargin. Free point processes and free extreme values. Probability Theory and Related Fields, 147(1– 2):161–183, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0204-z.

Andres:2016:CEL

[AK16] Sebastian Andres and Naotaka Kajino. Continuity and estimates of the Liouville heat kernel with applications to spectral dimensions. Probability Theory and Related Fields, 166(3–4): 713–752, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0670-4;
http://link.springer.com/article/10.1007/s00440-015-
0670-4.

Andreis:2023:LDP

[AKLP23] Luisa Andreis, Wolfgang König, Heide Langhammer, and Robert I. A. Patterson. A large-deviations principle for all the components in a sparse inhomogeneous random graph. Probability Theory and Related Fields, 186(1–2):521–620, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01180-7.

Albeverio:1997:UGS

[AKRT97] S. Albeverio, Y. G. Kondratiev, M. Röckner, and T. V. Tsikalenko. Uniqueness of Gibbs states for quantum lattice systems. Probability Theory and Related Fields, 108(2):193–218, June 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050107.

Adler:2018:CRS

[AKTW18] Robert J. Adler, Sunder Ram Krishnan, Jonathan E. Taylor, and Shmuel Weinberger. Convergence of the reach for a sequence of Gaussian-embedded manifolds. Probability Theory and Related Fields, 171(3–4):1045–1091, August 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0801-1.

Arous:1991:DEDa

[AL91a] G. Ben Arous and R. Léandre. Décroissance exponentielle du noyau de la chaleur sur la diagonale (i). (French) [Exponential decay of kernel heat on the diagonal (I)]. Probability Theory and Related Fields, 90(2):175–202, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01192161.

Arous:1991:DEDb

[AL91b] G. Ben Arous and R. Léandre. Décroissance exponentielle du noyau de la chaleur sur la diagonale (II). (French) [Exponential decay of kernel heat on the diagonal. (II)]. Probability Theory and Related Fields, 90(3):377–402, ???? 1991. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01193751.

Abraham:1994:MSD

[AL94] Romain Abraham and Jean-François Le Gall. Sur la mesure de sortie du super mouvement brownien. (French) [On the output measure of super Brownian motion]. Probability Theory and Related Fields, 99(2):251–275, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01199025.

Ane:2000:LSI

[AL00] Cécile Ané and Michel Ledoux. On logarithmic Sobolev inequalities for continuous time random walks on graphs. Probability Theory and Related Fields, 116(4):573–602, April 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050263.

Anh:2002:RHF

[AL02] V. V. Anh and N. N. Leonenko. Renormalization and homogenization of fractional diffusion equations with random data. Probability Theory and Related Fields, 124(3):381–408, November 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400200217.

Armendariz:2009:TLI

[AL09] Inés Armendáriz and Michail Loulakis. Thermodynamic limit for the invariant measures in supercritical zero range processes. Probability Theory and Related Fields, 145(1–2):175– 188, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0165-7.

Adhikari:2023:LLR

[AL23] Arka Adhikari and Benjamin Landon. Local law and rigidity for unitary Brownian motion. Probability Theory and Related Fields, 187(3–4):753–815, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-023-01230-
8.

Aldous:1986:SID

[Ald86] David J. Aldous. Self-intersections of 1-dimensional random walks. Probability Theory and Related Fields, 72(4):559–587, July 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00344721.

Aldous:1988:FTI

[Ald88] David J. Aldous. Finite-time implications of relaxation times for stochastically monotone processes. Probability Theory and Related Fields, 77(1):137–145, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01848136.

Aldous:1992:ARA

[Ald92] David Aldous. Asymptotics in the random assignment problem. Probability Theory and Related Fields, 93(4):507–534, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01192719.

Alexander:1987:CLT

[Ale87a] Kenneth S. Alexander. Central limit theorems for stochastic processes under random entropy conditions. Probability Theory and Related Fields, 75(3):351–378, July 1987. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00318707.

Alexander:1987:RGS

[Ale87b] Kenneth S. Alexander. Rates of growth and sample moduli for weighted empirical processes indexed by sets. Probability Theory and Related Fields, 75(3):379–423, July 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00318708.

Alexander:1998:WML

[Ale98] Kenneth S. Alexander. On weak mixing in lattice models. Probability Theory and Related Fields, 110(4):441–471, May 1998.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050155.

Alexander:2001:ARC

[Ale01] Kenneth S. Alexander. The asymmetric random cluster model and comparison of Ising and Potts models. Probability Theory and Related Fields, 120(3):395–444, July 2001. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008788.

Alexopoulos:2002:CDL

[Ale02] Georgios K. Alexopoulos. Centered densities on Lie groups of polynomial volume growth. Probability Theory and Related Fields, 124(1):112–150, September 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400200212.

Alhakim:2004:EEO

[Alh04] Abbas Alhakim. On the eigenvalues and eigenvectors of an overlapping Markov chain. Probability Theory and Related Fields, 128(4):589–605, April 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0321-z.

Alos:1999:SHE

[ALN99] Elisa Alòs, Jorge A. León, and David Nualart. Stochastic heat equation with random coefficients. Probability Theory and Related Fields, 115(1):41–94, August 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050236.

Aru:2020:FPS

[ALS20] Juhan Aru, Titus Lupu, and Avelio Sepúlveda. First passage sets of the 2D continuum Gaussian free field. Probability Theory and Related Fields, 176(3–4):1303–1355, April 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00941-1.

Althofer:1989:APL

[Alt89] Ingo Althöfer. Asymptotic properties of level-regular decision trees with randomly evaluated leaves. Probability Theory and Related Fields, 80(3):381–394, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01794430.

Antille:1988:CIS

[AM88] André Antille and Philip Milasevic. A confidence interval for the slope of a truncated regression. Probability Theory and Related Fields, 78(1):63–72, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00718035.

Assing:1995:BSS

[AM95] S. Assing and R. Manthey. The behavior of solutions of stochastic differential inequalities. Probability Theory and Related Fields, 103(4):493–514, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01246336.

Airault:1996:IPF

[AM96] Helene Airault and Paul Malliavin. Integration by parts formulas and dilatation vector fields on elliptic probability spaces. Probability Theory and Related Fields, 106(4):447–494, December 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050072.

Andjel:2008:SRL

[AM08] Enrique Andjel and T. S. Mountford. Some results for longrange exclusion processes. Probability Theory and Related Fields, 142(1–2):189–217, September 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0102-1.

Alsmeyer:2013:FPS

[AM13] Gerold Alsmeyer and Matthias Meiners. Fixed points of the smoothing transform: two-sided solutions. Probability Theory and Related Fields, 155(1–2):165–199, February 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0395-y.

Aldous:2004:EPI

[AMP04] David Aldous, Grégory Miermont, and Jim Pitman. The exploration process of inhomogeneous continuum random trees, and an extension of Jeulin’s local time identity. Probability Theory and Related Fields, 129(2):182–218, June 2004. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0334-7.

Aldous:2005:WCR

[AMP05] David Aldous, Grégory Miermont, and Jim Pitman. Weak convergence of random p-mappings and the exploration process of inhomogeneous continuum random trees. Probability Theory and Related Fields, 133(1):1–17, September 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0407-2.

Albeverio:1994:SSU

[AMS94] Sergio Albeverio, Stanislav A. Molchanov, and Donatas Surgailis. Stratified structure of the Universe and Burgers’ equation — a probabilistic approach. Probability Theory and Related Fields, 100(4):457–484, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01268990.

Aebi:1992:LDP

[AN92] Robert Aebi and Masao Nagasawa. Large deviations and the propagation of chaos for Schrödinger processes. Probability Theory and Related Fields, 94(1):53–68, ???? 1992. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01222509.

Andjel:1986:CNE

[And86] Enrique D. Andjel. Convergence to a non extremal equilibrium measure in the exclusion process. Probability Theory and Related Fields, 73(1):127–134, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01845996.

Andrew:2008:LBL

[And08] P. Andrew. On the limiting behaviour of Lévy processes at zero. Probability Theory and Related Fields, 140(1–2):103–127, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00440-007-0059-0.pdf.

Anonymous:1986:A

[Ano86a] Anonymous. Announcement. Probability Theory and Related Fields, 71(1):A3, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00366268.

Anonymous:1986:HCa

[Ano86b] Anonymous. Help & contacts. Probability Theory and Related Fields, 71(1):??, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCb

[Ano86c] Anonymous. Help & contacts. Probability Theory and Related Fields, 71(2):??, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCc

[Ano86d] Anonymous. Help & contacts. Probability Theory and Related Fields, 71(3):??, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCd

[Ano86e] Anonymous. Help & contacts. Probability Theory and Related Fields, 71(4):??, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCe

[Ano86f] Anonymous. Help & contacts. Probability Theory and Related Fields, 72(1):??, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCf

[Ano86g] Anonymous. Help & contacts. Probability Theory and Related Fields, 72(2):??, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCg

[Ano86h] Anonymous. Help & contacts. Probability Theory and Related Fields, 72(3):??, June 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCh

[Ano86i] Anonymous. Help & contacts. Probability Theory and Related Fields, 72(4):??, July 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCi

[Ano86j] Anonymous. Help & contacts. Probability Theory and Related Fields, 73(1):??, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCj

[Ano86k] Anonymous. Help & contacts. Probability Theory and Related Fields, 73(2):??, September 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCk

[Ano86l] Anonymous. Help & contacts. Probability Theory and Related Fields, 73(3):??, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1986:HCl

[Ano86m] Anonymous. Help & contacts. Probability Theory and Related Fields, 73(4):??, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:CAN

[Ano87a] Anonymous. Correction and addendum to: “A note on invariant measures for Markov maps of an interval. Probability Theory and Related Fields, 76(2):255–256, October 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00319987. See [Bug85].

Anonymous:1987:HCa

[Ano87b] Anonymous. Help & contacts. Probability Theory and Related Fields, 74(1):??, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCb

[Ano87c] Anonymous. Help & contacts. Probability Theory and Related Fields, 74(2):??, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCc

[Ano87d] Anonymous. Help & contacts. Probability Theory and Related Fields, 74(3):??, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCd

[Ano87e] Anonymous. Help & contacts. Probability Theory and Related Fields, 74(4):??, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCe

[Ano87f] Anonymous. Help & contacts. Probability Theory and Related Fields, 75(1):??, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCf

[Ano87g] Anonymous. Help & contacts. Probability Theory and Related Fields, 75(2):??, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCg

[Ano87h] Anonymous. Help & contacts. Probability Theory and Related Fields, 75(3):??, July 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCh

[Ano87i] Anonymous. Help & contacts. Probability Theory and Related Fields, 75(4):??, August 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCi

[Ano87j] Anonymous. Help & contacts. Probability Theory and Related Fields, 76(1):??, September 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCj

[Ano87k] Anonymous. Help & contacts. Probability Theory and Related Fields, 76(2):??, October 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCk

[Ano87l] Anonymous. Help & contacts. Probability Theory and Related Fields, 76(3):??, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1987:HCl

[Ano87m] Anonymous. Help & contacts. Probability Theory and Related Fields, 76(4):??, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCa

[Ano88a] Anonymous. Help & contacts. Probability Theory and Related Fields, 77(1):??, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCb

[Ano88b] Anonymous. Help & contacts. Probability Theory and Related Fields, 77(2):??, February 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCc

[Ano88c] Anonymous. Help & contacts. Probability Theory and Related Fields, 77(3):??, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCd

[Ano88d] Anonymous. Help & contacts. Probability Theory and Related Fields, 77(4):??, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCe

[Ano88e] Anonymous. Help & contacts. Probability Theory and Related Fields, 78(1):??, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCf

[Ano88f] Anonymous. Help & contacts. Probability Theory and Related Fields, 78(2):??, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCg

[Ano88g] Anonymous. Help & contacts. Probability Theory and Related Fields, 78(3):??, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCh

[Ano88h] Anonymous. Help & contacts. Probability Theory and Related Fields, 78(4):??, August 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCi

[Ano88i] Anonymous. Help & contacts. Probability Theory and Related Fields, 79(1):??, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCj

[Ano88j] Anonymous. Help & contacts. Probability Theory and Related Fields, 79(2):??, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCk

[Ano88k] Anonymous. Help & contacts. Probability Theory and Related Fields, 79(3):??, October 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCl

[Ano88l] Anonymous. Help & contacts. Probability Theory and Related Fields, 79(4):??, November 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCm

[Ano88m] Anonymous. Help & contacts. Probability Theory and Related Fields, 80(1):??, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1988:HCn

[Ano88n] Anonymous. Help & contacts. Probability Theory and Related Fields, 80(2):??, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:AS

[Ano89a] Anonymous. The annals of statistics. Probability Theory and Related Fields, 81(3):A3–A5, April 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00340062.

Anonymous:1989:HCa

[Ano89b] Anonymous. Help & contacts. Probability Theory and Related Fields, 80(3):??, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCb

[Ano89c] Anonymous. Help & contacts. Probability Theory and Related Fields, 80(4):??, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCc

[Ano89d] Anonymous. Help & contacts. Probability Theory and Related Fields, 81(1):??, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCd

[Ano89e] Anonymous. Help & contacts. Probability Theory and Related Fields, 81(2):??, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCe

[Ano89f] Anonymous. Help & contacts. Probability Theory and Related Fields, 81(3):??, April 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCf

[Ano89g] Anonymous. Help & contacts. Probability Theory and Related Fields, 81(4):??, May 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCg

[Ano89h] Anonymous. Help & contacts. Probability Theory and Related Fields, 82(1):??, June 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCh

[Ano89i] Anonymous. Help & contacts. Probability Theory and Related Fields, 82(2):??, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCi

[Ano89j] Anonymous. Help & contacts. Probability Theory and Related Fields, 82(3):??, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCj

[Ano89k] Anonymous. Help & contacts. Probability Theory and Related Fields, 82(4):??, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCk

[Ano89l] Anonymous. Help & contacts. Probability Theory and Related Fields, 83(1–2):??, September 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCl

[Ano89m] Anonymous. Help & contacts. Probability Theory and Related Fields, 83(3):??, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1989:HCm

[Ano89n] Anonymous. Help & contacts. Probability Theory and Related Fields, 83(4):??, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCa

[Ano90a] Anonymous. Help & contacts. Probability Theory and Related Fields, 84(1):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCb

[Ano90b] Anonymous. Help & contacts. Probability Theory and Related Fields, 84(2):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCc

[Ano90c] Anonymous. Help & contacts. Probability Theory and Related Fields, 84(3):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCd

[Ano90d] Anonymous. Help & contacts. Probability Theory and Related Fields, 84(4):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCe

[Ano90e] Anonymous. Help & contacts. Probability Theory and Related Fields, 85(1):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCf

[Ano90f] Anonymous. Help & contacts. Probability Theory and Related Fields, 85(2):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCg

[Ano90g] Anonymous. Help & contacts. Probability Theory and Related Fields, 85(3):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCh

[Ano90h] Anonymous. Help & contacts. Probability Theory and Related Fields, 85(4):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCi

[Ano90i] Anonymous. Help & contacts. Probability Theory and Related Fields, 86(1):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCj

[Ano90j] Anonymous. Help & contacts. Probability Theory and Related Fields, 86(2):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCk

[Ano90k] Anonymous. Help & contacts. Probability Theory and Related Fields, 86(3):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCl

[Ano90l] Anonymous. Help & contacts. Probability Theory and Related Fields, 86(4):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCm

[Ano90m] Anonymous. Help & contacts. Probability Theory and Related Fields, 87(1):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1990:HCn

[Ano90n] Anonymous. Help & contacts. Probability Theory and Related Fields, 87(2):??, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCa

[Ano91a] Anonymous. Help & contacts. Probability Theory and Related Fields, 87(3):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCb

[Ano91b] Anonymous. Help & contacts. Probability Theory and Related Fields, 87(4):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCc

[Ano91c] Anonymous. Help & contacts. Probability Theory and Related Fields, 88(1):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCd

[Ano91d] Anonymous. Help & contacts. Probability Theory and Related Fields, 88(2):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCe

[Ano91e] Anonymous. Help & contacts. Probability Theory and Related Fields, 88(3):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCf

[Ano91f] Anonymous. Help & contacts. Probability Theory and Related Fields, 88(4):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCg

[Ano91g] Anonymous. Help & contacts. Probability Theory and Related Fields, 89(1):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCh

[Ano91h] Anonymous. Help & contacts. Probability Theory and Related Fields, 89(2):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCi

[Ano91i] Anonymous. Help & contacts. Probability Theory and Related Fields, 89(3):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCj

[Ano91j] Anonymous. Help & contacts. Probability Theory and Related Fields, 89(4):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCk

[Ano91k] Anonymous. Help & contacts. Probability Theory and Related Fields, 90(1):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCl

[Ano91l] Anonymous. Help & contacts. Probability Theory and Related Fields, 90(2):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCm

[Ano91m] Anonymous. Help & contacts. Probability Theory and Related Fields, 90(3):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1991:HCn

[Ano91n] Anonymous. Help & contacts. Probability Theory and Related Fields, 90(4):??, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCa

[Ano92a] Anonymous. Help & contacts. Probability Theory and Related Fields, 91(1):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCb

[Ano92b] Anonymous. Help & contacts. Probability Theory and Related Fields, 91(2):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCm

[Ano92c] Anonymous. Help & contacts. Probability Theory and Related Fields, 91(3–4):??, September 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCc

[Ano92d] Anonymous. Help & contacts. Probability Theory and Related Fields, 92(1):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCd

[Ano92e] Anonymous. Help & contacts. Probability Theory and Related Fields, 92(2):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCe

[Ano92f] Anonymous. Help & contacts. Probability Theory and Related Fields, 92(3):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCf

[Ano92g] Anonymous. Help & contacts. Probability Theory and Related Fields, 92(4):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCg

[Ano92h] Anonymous. Help & contacts. Probability Theory and Related Fields, 93(1):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCh

[Ano92i] Anonymous. Help & contacts. Probability Theory and Related Fields, 93(2):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCi

[Ano92j] Anonymous. Help & contacts. Probability Theory and Related Fields, 93(3):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCj

[Ano92k] Anonymous. Help & contacts. Probability Theory and Related Fields, 93(4):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCk

[Ano92l] Anonymous. Help & contacts. Probability Theory and Related Fields, 94(1):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1992:HCl

[Ano92m] Anonymous. Help & contacts. Probability Theory and Related Fields, 94(2):??, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCa

[Ano93a] Anonymous. Help & contacts. Probability Theory and Related Fields, 94(3):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCb

[Ano93b] Anonymous. Help & contacts. Probability Theory and Related Fields, 94(4):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCc

[Ano93c] Anonymous. Help & contacts. Probability Theory and Related Fields, 95(1):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCd

[Ano93d] Anonymous. Help & contacts. Probability Theory and Related Fields, 95(2):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCe

[Ano93e] Anonymous. Help & contacts. Probability Theory and Related Fields, 95(3):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCf

[Ano93f] Anonymous. Help & contacts. Probability Theory and Related Fields, 95(4):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCg

[Ano93g] Anonymous. Help & contacts. Probability Theory and Related Fields, 96(1):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCh

[Ano93h] Anonymous. Help & contacts. Probability Theory and Related Fields, 96(2):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCi

[Ano93i] Anonymous. Help & contacts. Probability Theory and Related Fields, 96(3):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCj

[Ano93j] Anonymous. Help & contacts. Probability Theory and Related Fields, 96(4):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCk

[Ano93k] Anonymous. Help & contacts. Probability Theory and Related Fields, 97(1–2):??, March 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCl

[Ano93l] Anonymous. Help & contacts. Probability Theory and Related Fields, 97(3):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1993:HCm

[Ano93m] Anonymous. Help & contacts. Probability Theory and Related Fields, 97(4):??, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCa

[Ano94a] Anonymous. Help & contacts. Probability Theory and Related Fields, 98(1):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCb

[Ano94b] Anonymous. Help & contacts. Probability Theory and Related Fields, 98(2):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCc

[Ano94c] Anonymous. Help & contacts. Probability Theory and Related Fields, 98(3):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCd

[Ano94d] Anonymous. Help & contacts. Probability Theory and Related Fields, 98(4):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCe

[Ano94e] Anonymous. Help & contacts. Probability Theory and Related Fields, 99(1):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCf

[Ano94f] Anonymous. Help & contacts. Probability Theory and Related Fields, 99(2):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCg

[Ano94g] Anonymous. Help & contacts. Probability Theory and Related Fields, 99(3):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCh

[Ano94h] Anonymous. Help & contacts. Probability Theory and Related Fields, 99(4):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCi

[Ano94i] Anonymous. Help & contacts. Probability Theory and Related Fields, 100(1):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCj

[Ano94j] Anonymous. Help & contacts. Probability Theory and Related Fields, 100(2):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCk

[Ano94k] Anonymous. Help & contacts. Probability Theory and Related Fields, 100(3):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1994:HCl

[Ano94l] Anonymous. Help & contacts. Probability Theory and Related Fields, 100(4):??, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCa

[Ano95a] Anonymous. Help & contacts. Probability Theory and Related Fields, 101(1):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCb

[Ano95b] Anonymous. Help & contacts. Probability Theory and Related Fields, 101(2):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCc

[Ano95c] Anonymous. Help & contacts. Probability Theory and Related Fields, 101(3):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCd

[Ano95d] Anonymous. Help & contacts. Probability Theory and Related Fields, 101(4):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCe

[Ano95e] Anonymous. Help & contacts. Probability Theory and Related Fields, 102(1):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCf

[Ano95f] Anonymous. Help & contacts. Probability Theory and Related Fields, 102(2):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCg

[Ano95g] Anonymous. Help & contacts. Probability Theory and Related Fields, 102(3):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCh

[Ano95h] Anonymous. Help & contacts. Probability Theory and Related Fields, 102(4):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCi

[Ano95i] Anonymous. Help & contacts. Probability Theory and Related Fields, 103(1):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCj

[Ano95j] Anonymous. Help & contacts. Probability Theory and Related Fields, 103(2):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCk

[Ano95k] Anonymous. Help & contacts. Probability Theory and Related Fields, 103(3):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1995:HCl

[Ano95l] Anonymous. Help & contacts. Probability Theory and Related Fields, 103(4):??, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCa

[Ano96a] Anonymous. Help & contacts. Probability Theory and Related Fields, 104(1):??, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCb

[Ano96b] Anonymous. Help & contacts. Probability Theory and Related Fields, 104(2):??, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCc

[Ano96c] Anonymous. Help & contacts. Probability Theory and Related Fields, 104(3):??, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCd

[Ano96d] Anonymous. Help & contacts. Probability Theory and Related Fields, 104(4):??, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCe

[Ano96e] Anonymous. Help & contacts. Probability Theory and Related Fields, 105(1):??, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCf

[Ano96f] Anonymous. Help & contacts. Probability Theory and Related Fields, 105(2):??, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCg

[Ano96g] Anonymous. Help & contacts. Probability Theory and Related Fields, 105(3):??, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCh

[Ano96h] Anonymous. Help & contacts. Probability Theory and Related Fields, 105(4):??, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCi

[Ano96i] Anonymous. Help & contacts. Probability Theory and Related Fields, 106(1):??, September 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCj

[Ano96j] Anonymous. Help & contacts. Probability Theory and Related Fields, 106(2):??, October 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCk

[Ano96k] Anonymous. Help & contacts. Probability Theory and Related Fields, 106(3):??, November 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1996:HCl

[Ano96l] Anonymous. Help & contacts. Probability Theory and Related Fields, 106(4):??, December 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCa

[Ano97a] Anonymous. Help & contacts. Probability Theory and Related Fields, 107(1):??, January 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCb

[Ano97b] Anonymous. Help & contacts. Probability Theory and Related Fields, 107(2):??, February 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCc

[Ano97c] Anonymous. Help & contacts. Probability Theory and Related Fields, 107(3):??, March 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCd

[Ano97d] Anonymous. Help & contacts. Probability Theory and Related Fields, 107(4):??, April 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCe

[Ano97e] Anonymous. Help & contacts. Probability Theory and Related Fields, 108(1):??, May 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCf

[Ano97f] Anonymous. Help & contacts. Probability Theory and Related Fields, 108(2):??, June 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCg

[Ano97g] Anonymous. Help & contacts. Probability Theory and Related Fields, 108(3):??, July 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCh

[Ano97h] Anonymous. Help & contacts. Probability Theory and Related Fields, 108(4):??, August 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCi

[Ano97i] Anonymous. Help & contacts. Probability Theory and Related Fields, 109(1):??, September 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCj

[Ano97j] Anonymous. Help & contacts. Probability Theory and Related Fields, 109(2):??, October 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCk

[Ano97k] Anonymous. Help & contacts. Probability Theory and Related Fields, 109(3):??, November 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1997:HCl

[Ano97l] Anonymous. Help & contacts. Probability Theory and Related Fields, 109(4):??, November 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:EPT

[Ano98a] Anonymous. Erratum: Probab. Theory Relat. Fields (1991) 89: 447–455. Probability Theory and Related Fields, 112(2): 297, October 1998. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s004400050191. See [Hal91].

Anonymous:1998:HCa

[Ano98b] Anonymous. Help & contacts. Probability Theory and Related Fields, 110(1):??, January 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCb

[Ano98c] Anonymous. Help & contacts. Probability Theory and Related Fields, 110(2):??, February 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCc

[Ano98d] Anonymous. Help & contacts. Probability Theory and Related Fields, 110(3):??, March 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCd

[Ano98e] Anonymous. Help & contacts. Probability Theory and Related Fields, 110(4):??, May 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCe

[Ano98f] Anonymous. Help & contacts. Probability Theory and Related Fields, 111(1):??, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCf

[Ano98g] Anonymous. Help & contacts. Probability Theory and Related Fields, 111(2):??, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCg

[Ano98h] Anonymous. Help & contacts. Probability Theory and Related Fields, 111(3):??, July 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCh

[Ano98i] Anonymous. Help & contacts. Probability Theory and Related Fields, 111(4):??, August 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCi

[Ano98j] Anonymous. Help & contacts. Probability Theory and Related Fields, 112(1):??, September 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCj

[Ano98k] Anonymous. Help & contacts. Probability Theory and Related Fields, 112(2):??, October 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCk

[Ano98l] Anonymous. Help & contacts. Probability Theory and Related Fields, 112(3):??, November 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1998:HCl

[Ano98m] Anonymous. Help & contacts. Probability Theory and Related Fields, 112(4):??, December 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCa

[Ano99a] Anonymous. Help & contacts. Probability Theory and Related Fields, 113(1):??, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCb

[Ano99b] Anonymous. Help & contacts. Probability Theory and Related Fields, 113(2):??, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCc

[Ano99c] Anonymous. Help & contacts. Probability Theory and Related Fields, 113(3):??, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCd

[Ano99d] Anonymous. Help & contacts. Probability Theory and Related Fields, 113(4):??, March 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCe

[Ano99e] Anonymous. Help & contacts. Probability Theory and Related Fields, 114(1):??, April 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCf

[Ano99f] Anonymous. Help & contacts. Probability Theory and Related Fields, 114(2):??, May 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCg

[Ano99g] Anonymous. Help & contacts. Probability Theory and Related Fields, 114(3):??, June 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCh

[Ano99h] Anonymous. Help & contacts. Probability Theory and Related Fields, 114(4):??, July 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCi

[Ano99i] Anonymous. Help & contacts. Probability Theory and Related Fields, 115(1):??, August 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCj

[Ano99j] Anonymous. Help & contacts. Probability Theory and Related Fields, 115(2):??, September 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCk

[Ano99k] Anonymous. Help & contacts. Probability Theory and Related Fields, 115(3):??, October 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:1999:HCl

[Ano99l] Anonymous. Help & contacts. Probability Theory and Related Fields, 115(4):??, November 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCa

[Ano00a] Anonymous. Help & contacts. Probability Theory and Related Fields, 116(1):??, January 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCb

[Ano00b] Anonymous. Help & contacts. Probability Theory and Related Fields, 116(2):??, February 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCc

[Ano00c] Anonymous. Help & contacts. Probability Theory and Related Fields, 116(3):??, March 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCd

[Ano00d] Anonymous. Help & contacts. Probability Theory and Related Fields, 116(4):??, April 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCe

[Ano00e] Anonymous. Help & contacts. Probability Theory and Related Fields, 117(1):??, May 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCf

[Ano00f] Anonymous. Help & contacts. Probability Theory and Related Fields, 117(2):??, June 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCg

[Ano00g] Anonymous. Help & contacts. Probability Theory and Related Fields, 117(3):??, July 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCh

[Ano00h] Anonymous. Help & contacts. Probability Theory and Related Fields, 117(4):??, August 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCi

[Ano00i] Anonymous. Help & contacts. Probability Theory and Related Fields, 118(1):??, September 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCj

[Ano00j] Anonymous. Help & contacts. Probability Theory and Related Fields, 118(2):??, October 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCk

[Ano00k] Anonymous. Help & contacts. Probability Theory and Related Fields, 118(3):??, November 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2000:HCl

[Ano00l] Anonymous. Help & contacts. Probability Theory and Related Fields, 118(4):??, December 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCa

[Ano01a] Anonymous. Help & contacts. Probability Theory and Related Fields, 119(1):??, January 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCb

[Ano01b] Anonymous. Help & contacts. Probability Theory and Related Fields, 119(2):??, February 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCc

[Ano01c] Anonymous. Help & contacts. Probability Theory and Related Fields, 119(3):??, March 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCd

[Ano01d] Anonymous. Help & contacts. Probability Theory and Related Fields, 119(4):??, April 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCe

[Ano01e] Anonymous. Help & contacts. Probability Theory and Related Fields, 120(1):??, May 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCf

[Ano01f] Anonymous. Help & contacts. Probability Theory and Related Fields, 120(2):??, June 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCg

[Ano01g] Anonymous. Help & contacts. Probability Theory and Related Fields, 120(3):??, July 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCh

[Ano01h] Anonymous. Help & contacts. Probability Theory and Related Fields, 120(4):??, August 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCi

[Ano01i] Anonymous. Help & contacts. Probability Theory and Related Fields, 121(1):??, September 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCj

[Ano01j] Anonymous. Help & contacts. Probability Theory and Related Fields, 121(2):??, October 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCk

[Ano01k] Anonymous. Help & contacts. Probability Theory and Related Fields, 121(3):??, November 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2001:HCl

[Ano01l] Anonymous. Help & contacts. Probability Theory and Related Fields, 121(4):??, December 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCa

[Ano02a] Anonymous. Help & contacts. Probability Theory and Related Fields, 122(1):??, January 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCb

[Ano02b] Anonymous. Help & contacts. Probability Theory and Related Fields, 122(2):??, February 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCc

[Ano02c] Anonymous. Help & contacts. Probability Theory and Related Fields, 122(3):??, March 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCd

[Ano02d] Anonymous. Help & contacts. Probability Theory and Related Fields, 122(4):??, April 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCe

[Ano02e] Anonymous. Help & contacts. Probability Theory and Related Fields, 123(1):??, May 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCf

[Ano02f] Anonymous. Help & contacts. Probability Theory and Related Fields, 123(2):??, June 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCg

[Ano02g] Anonymous. Help & contacts. Probability Theory and Related Fields, 123(3):??, July 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCh

[Ano02h] Anonymous. Help & contacts. Probability Theory and Related Fields, 123(4):??, August 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCi

[Ano02i] Anonymous. Help & contacts. Probability Theory and Related Fields, 124(1):??, September 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCj

[Ano02j] Anonymous. Help & contacts. Probability Theory and Related Fields, 124(2):??, October 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCk

[Ano02k] Anonymous. Help & contacts. Probability Theory and Related Fields, 124(3):??, November 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2002:HCl

[Ano02l] Anonymous. Help & contacts. Probability Theory and Related Fields, 124(4):??, December 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCa

[Ano03a] Anonymous. Help & contacts. Probability Theory and Related Fields, 125(1):??, January 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCb

[Ano03b] Anonymous. Help & contacts. Probability Theory and Related Fields, 125(2):??, February 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCc

[Ano03c] Anonymous. Help & contacts. Probability Theory and Related Fields, 125(3):??, March 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCd

[Ano03d] Anonymous. Help & contacts. Probability Theory and Related Fields, 125(4):??, April 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCe

[Ano03e] Anonymous. Help & contacts. Probability Theory and Related Fields, 126(1):??, May 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCf

[Ano03f] Anonymous. Help & contacts. Probability Theory and Related Fields, 126(2):??, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCg

[Ano03g] Anonymous. Help & contacts. Probability Theory and Related Fields, 126(3):??, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCh

[Ano03h] Anonymous. Help & contacts. Probability Theory and Related Fields, 126(4):??, August 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCi

[Ano03i] Anonymous. Help & contacts. Probability Theory and Related Fields, 127(1):??, September 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCj

[Ano03j] Anonymous. Help & contacts. Probability Theory and Related Fields, 127(2):??, October 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCk

[Ano03k] Anonymous. Help & contacts. Probability Theory and Related Fields, 127(3):??, November 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2003:HCl

[Ano03l] Anonymous. Help & contacts. Probability Theory and Related Fields, 127(4):??, December 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCa

[Ano04a] Anonymous. Help & contacts. Probability Theory and Related Fields, 128(1):??, January 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCb

[Ano04b] Anonymous. Help & contacts. Probability Theory and Related Fields, 128(2):??, February 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCc

[Ano04c] Anonymous. Help & contacts. Probability Theory and Related Fields, 128(3):??, March 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCd

[Ano04d] Anonymous. Help & contacts. Probability Theory and Related Fields, 128(4):??, April 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCe

[Ano04e] Anonymous. Help & contacts. Probability Theory and Related Fields, 129(1):??, May 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCf

[Ano04f] Anonymous. Help & contacts. Probability Theory and Related Fields, 129(2):??, June 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCg

[Ano04g] Anonymous. Help & contacts. Probability Theory and Related Fields, 129(3):??, July 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCh

[Ano04h] Anonymous. Help & contacts. Probability Theory and Related Fields, 129(4):??, August 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCi

[Ano04i] Anonymous. Help & contacts. Probability Theory and Related Fields, 130(1):??, September 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCj

[Ano04j] Anonymous. Help & contacts. Probability Theory and Related Fields, 130(2):??, October 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCk

[Ano04k] Anonymous. Help & contacts. Probability Theory and Related Fields, 130(3):??, November 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2004:HCl

[Ano04l] Anonymous. Help & contacts. Probability Theory and Related Fields, 130(4):??, December 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCa

[Ano05a] Anonymous. Help & contacts. Probability Theory and Related Fields, 131(1):??, January 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCb

[Ano05b] Anonymous. Help & contacts. Probability Theory and Related Fields, 131(2):??, February 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCc

[Ano05c] Anonymous. Help & contacts. Probability Theory and Related Fields, 131(3):??, March 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCd

[Ano05d] Anonymous. Help & contacts. Probability Theory and Related Fields, 131(4):??, April 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCe

[Ano05e] Anonymous. Help & contacts. Probability Theory and Related Fields, 132(1):??, May 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCf

[Ano05f] Anonymous. Help & contacts. Probability Theory and Related Fields, 132(2):??, June 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCg

[Ano05g] Anonymous. Help & contacts. Probability Theory and Related Fields, 132(3):??, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCh

[Ano05h] Anonymous. Help & contacts. Probability Theory and Related Fields, 132(4):??, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCi

[Ano05i] Anonymous. Help & contacts. Probability Theory and Related Fields, 133(1):??, September 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCj

[Ano05j] Anonymous. Help & contacts. Probability Theory and Related Fields, 133(2):??, October 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCk

[Ano05k] Anonymous. Help & contacts. Probability Theory and Related Fields, 133(3):??, November 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2005:HCl

[Ano05l] Anonymous. Help & contacts. Probability Theory and Related Fields, 133(4):??, December 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2006:HCa

[Ano06a] Anonymous. Help & contacts. Probability Theory and Related Fields, 134(1):??, January 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2006:HCb

[Ano06b] Anonymous. Help & contacts. Probability Theory and Related Fields, 134(2):??, February 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2006:HCc

[Ano06c] Anonymous. Help & contacts. Probability Theory and Related Fields, 134(3):??, March 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2006:HCd

[Ano06d] Anonymous. Help & contacts. Probability Theory and Related Fields, 134(4):??, April 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2007:HCa

[Ano07a] Anonymous. Help & contacts. Probability Theory and Related Fields, 137(1–2):??, January 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2007:HCb

[Ano07b] Anonymous. Help & contacts. Probability Theory and Related Fields, 137(3–4):??, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2007:HCc

[Ano07c] Anonymous. Help & contacts. Probability Theory and Related Fields, 138(1–2):??, May 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2007:HCd

[Ano07d] Anonymous. Help & contacts. Probability Theory and Related Fields, 138(3–4):??, July 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2007:HCe

[Ano07e] Anonymous. Help & contacts. Probability Theory and Related Fields, 139(1–2):??, September 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2007:HCf

[Ano07f] Anonymous. Help & contacts. Probability Theory and Related Fields, 139(3–4):??, November 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2008:HCa

[Ano08a] Anonymous. Help & contacts. Probability Theory and Related Fields, 140(1–2):??, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2008:HCb

[Ano08b] Anonymous. Help & contacts. Probability Theory and Related Fields, 140(3–4):??, March 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2008:HCc

[Ano08c] Anonymous. Help & contacts. Probability Theory and Related Fields, 141(1–2):??, May 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2008:HCd

[Ano08d] Anonymous. Help & contacts. Probability Theory and Related Fields, 141(3–4):??, July 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2008:HCe

[Ano08e] Anonymous. Help & contacts. Probability Theory and Related Fields, 142(1–2):??, September 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2008:HCf

[Ano08f] Anonymous. Help & contacts. Probability Theory and Related Fields, 142(3–4):??, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2009:HCa

[Ano09a] Anonymous. Help & contacts. Probability Theory and Related Fields, 143(1–2):??, January 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2009:HCb

[Ano09b] Anonymous. Help & contacts. Probability Theory and Related Fields, 143(3–4):??, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2009:HCc

[Ano09c] Anonymous. Help & contacts. Probability Theory and Related Fields, 144(1–2):??, May 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2009:HCd

[Ano09d] Anonymous. Help & contacts. Probability Theory and Related Fields, 144(3–4):??, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2009:HCe

[Ano09e] Anonymous. Help & contacts. Probability Theory and Related Fields, 145(1–2):??, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2009:HCf

[Ano09f] Anonymous. Help & contacts. Probability Theory and Related Fields, 145(3–4):??, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2010:HCa

[Ano10a] Anonymous. Help & contacts. Probability Theory and Related Fields, 146(1–2):??, January 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2010:HCb

[Ano10b] Anonymous. Help & contacts. Probability Theory and Related Fields, 146(3–4):??, March 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2010:HCc

[Ano10c] Anonymous. Help & contacts. Probability Theory and Related Fields, 147(1–2):??, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2010:HCd

[Ano10d] Anonymous. Help & contacts. Probability Theory and Related Fields, 147(3–4):??, July 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2010:HCe

[Ano10e] Anonymous. Help & contacts. Probability Theory and Related Fields, 148(1–2):??, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2010:HCf

[Ano10f] Anonymous. Help & contacts. Probability Theory and Related Fields, 148(3–4):??, November 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2011:HCa

[Ano11a] Anonymous. Help & contacts. Probability Theory and Related Fields, 149(1–2):??, February 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2011:HCb

[Ano11b] Anonymous. Help & contacts. Probability Theory and Related Fields, 149(3–4):??, April 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2011:HCc

[Ano11c] Anonymous. Help & contacts. Probability Theory and Related Fields, 150(1–2):??, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2011:HCd

[Ano11d] Anonymous. Help & contacts. Probability Theory and Related Fields, 150(3–4):??, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2011:HCe

[Ano11e] Anonymous. Help & contacts. Probability Theory and Related Fields, 151(1–2):??, October 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2011:HCf

[Ano11f] Anonymous. Help & contacts. Probability Theory and Related Fields, 151(3–4):??, December 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2012:HCa

[Ano12a] Anonymous. Help & contacts. Probability Theory and Related Fields, 152(1–2):??, February 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2012:HCb

[Ano12b] Anonymous. Help & contacts. Probability Theory and Related Fields, 152(3–4):??, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2012:HCc

[Ano12c] Anonymous. Help & contacts. Probability Theory and Related Fields, 153(1–2):??, June 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2012:HCd

[Ano12d] Anonymous. Help & contacts. Probability Theory and Related Fields, 153(3–4):??, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2012:HCe

[Ano12e] Anonymous. Help & contacts. Probability Theory and Related Fields, 154(1–2):??, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2012:HCf

[Ano12f] Anonymous. Help & contacts. Probability Theory and Related Fields, 154(3–4):??, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2013:HCa

[Ano13a] Anonymous. Help & contacts. Probability Theory and Related Fields, 155(1–2):??, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2013:HCb

[Ano13b] Anonymous. Help & contacts. Probability Theory and Related Fields, 155(3–4):??, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2013:HCc

[Ano13c] Anonymous. Help & contacts. Probability Theory and Related Fields, 156(1–2):??, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2013:HCd

[Ano13d] Anonymous. Help & contacts. Probability Theory and Related Fields, 156(3–4):??, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2013:HCe

[Ano13e] Anonymous. Help & contacts. Probability Theory and Related Fields, 157(1–2):??, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2013:HCf

[Ano13f] Anonymous. Help & contacts. Probability Theory and Related Fields, 157(3–4):??, December 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2014:HCa

[Ano14a] Anonymous. Help & contacts. Probability Theory and Related Fields, 158(1–2):??, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2014:HCb

[Ano14b] Anonymous. Help & contacts. Probability Theory and Related Fields, 158(3–4):??, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2014:HCc

[Ano14c] Anonymous. Help & contacts. Probability Theory and Related Fields, 159(1–2):??, June 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2014:HCd

[Ano14d] Anonymous. Help & contacts. Probability Theory and Related Fields, 159(3–4):??, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2014:HCe

[Ano14e] Anonymous. Help & contacts. Probability Theory and Related Fields, 160(1–2):??, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2014:HCf

[Ano14f] Anonymous. Help & contacts. Probability Theory and Related Fields, 160(3–4):??, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Anonymous:2017:Ea

[Ano17a] Anonymous. Editorial. Probability Theory and Related Fields, 167(3–4):553–554, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s00440-
017-0764-2.

Anonymous:2017:Eb

[Ano17b] Anonymous. Editorial. Probability Theory and Related Fields, 169(1–2):1–2, October 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0795-8.

Anonymous:2021:E

[Ano21a] Anonymous. Editorial. Probability Theory and Related Fields, 179(3–4):541–542, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-020-01023-3.

Anonymous:2021:PHK

[Ano21b] Anonymous. Publications of Harry Kesten. Probability Theory and Related Fields, 181(1-3):3–16, November 2021. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-021-01047-3.

Anshelevich:1999:LCL

[Ans99] Michael Anshelevich. The linearization of the central limit operator in free probability theory. Probability Theory and Related Fields, 115(3):401–416, October 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050243.

Anshelevich:2013:GSN

[Ans13] Michael Anshelevich. Generators of some non-commutative stochastic processes. Probability Theory and Related Fields, 157(3–4):777–815, December 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0470-z.

Aimino:2015:AQL

[ANV15] Romain Aimino, Matthew Nicol, and Sandro Vaienti. Annealed and quenched limit theorems for random expanding dynamical systems. Probability Theory and Related Fields, 162(1– 2):233–274, June 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0571-y.

Asch:1991:ILN

[AP91] Joachim Asch and Jürgen Potthoff. Itô’s lemma without nonanticipatory conditions. Probability Theory and Related Fields, 88(1):17–46, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193581.

Aldous:2000:ICR

[AP00] David Aldous and Jim Pitman. Inhomogeneous continuum random trees and the entrance boundary of the additive coalescent. Probability Theory and Related Fields, 118(4):455–482, December 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/PL00008751.

Austin:2014:HVF

[AP14] Tim Austin and Dmitry Panchenko. A hierarchical version of the de Finetti and Aldous–Hoover representations. Probability Theory and Related Fields, 159(3–4):809–823, August 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0521-0.

Albeverio:1989:CDF

[AR89] Sergio Albeverio and Michael Röckner. Classical Dirichlet forms on topological vector spaces-the construction of the associated diffusion process. Probability Theory and Related Fields, 83 (3):405–434, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00964372.

Albeverio:1991:SDE

[AR91] S. Albeverio and M. Röckner. Stochastic differential equations in infinite dimensions: solutions via Dirichlet forms. Probability Theory and Related Fields, 89(3):347–386, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198791.

Angel:2018:HPU

[AR18] Omer Angel and Gourab Ray. The half plane UIPT is recurrent. Probability Theory and Related Fields, 170(3–4):657–683, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-017-0767-z;   http://link.springer.com/
content/pdf/10.1007/s00440-017-0767-z.pdf.

Arbeiter:1991:RRC

[Arb91] Matthias Arbeiter. Random recursive construction of selfsimilar fractal measures. the noncompact case. Probability Theory and Related Fields, 88(4):497–520, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192554.

Arnaudon:1997:DAF

[Arn97] Marc Arnaudon. Differentiable and analytic families of continuous martingales in manifolds with connection. Probability Theory and Related Fields, 108(2):219–257, June 1997. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050108.

Arous:1989:FST

[Aro89] Gérard Ben Arous. Flots et séries de Taylor stochastiques. (French) [Waves and stochastic Taylor series]. Probability Theory and Related Fields, 81(1):29–77, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00343737.

Angel:2023:TTB

[ARS23] Omer Angel, Gourab Ray, and Yinon Spinka. A tale of two balloons. Probability Theory and Related Fields, 185(3–4):815– 837, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-022-01165-6.

Artstein:2001:CCF

[Art01] Zvi Artstein. Compact convergence of σ-fields and relaxed conditional expectation. Probability Theory and Related Fields, 120 (3):369–394, July 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008787.

Aru:2015:KRD

[Aru15] Juhan Aru. KPZ relation does not hold for the level lines and SLE_{κ} flow lines of the Gaussian free field. Probability Theory and Related Fields, 163(3–4):465–526, December 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0597-1.

Allez:2013:LSS

[ARV13] Romain Allez, Rémi Rhodes, and Vincent Vargas. Lognormal ⋆-scale invariant random measures. Probability Theory and Related Fields, 155(3–4):751–788, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0412-9.

Aldous:1988:DLC

[AS88] David Aldous and Paul Shields. A diffusion limit for a class of randomly-growing binary trees. Probability Theory and Related Fields, 79(4):509–542, November 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00318784.

Arak:1989:MFP

[AS89] T. Arak and D. Surgailis. Markov fields with polygonal realizations. Probability Theory and Related Fields, 80(4):543– 579, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00318906.

Arak:1991:CPF

[AS91] T. Arak and D. Surgailis. Consistent polygonal fields. Probability Theory and Related Fields, 89(3):319–346, ???? 1991.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01198790.

Aldous:1992:AEM

[AS92] David Aldous and J. Michael Steele. Asymptotics for Euclidean minimal spanning trees on random points. Probability Theory and Related Fields, 92(2):247–258, ???? 1992. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01194923.

Arnold:1995:PCT

[AS95] Ludwig Arnold and Michael Scheutzow. Perfect cocycles through stochastic differential equations. Probability Theory and Related Fields, 101(1):65–88, ???? 1995. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192196.

Arthreya:2005:BCP

[AS05] Siva R. Arthreya and Jan M. Swart. Branching-coalescing particle systems. Probability Theory and Related Fields, 131(3): 376–414, March 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0377-4.

Athreya:2009:BCP

[AS09] Siva R. Athreya and Jan M. Swart. Branching-coalescing particle systems. Probability Theory and Related Fields, 145(3–4): 639–640, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00440-009-0232-8.pdf.

Athreya:2010:SCP

[AS10] Siva R. Athreya and Jan M. Swart. Survival of contact processes on the hierarchical group. Probability Theory and Related Fields, 147(3–4):529–563, July 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-009-0214-x.

Alberts:2011:CMS

[AS11] Tom Alberts and Scott Sheffield. The covariant measure of SLE on the boundary. Probability Theory and Related Fields, 149 (3–4):331–371, April 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0252-4.

Asselah:2017:BRT

[AS17] Amine Asselah and Bruno Schapira. Boundary of the range of transient random walk. Probability Theory and Related Fields, 168(3–4):691–719, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-016-0722-4.

Asselah:2021:TRM

[AS21] Amine Asselah and Bruno Schapira. The two regimes of moderate deviations for the range of a transient walk. Probability Theory and Related Fields, 180(1–2):439–465, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01063-3.

Aoun:2022:RWH

[AS22] Richard Aoun and Cagri Sert. Random walks on hyperbolic spaces: Concentration inequalities and probabilistic Tits alternative. Probability Theory and Related Fields, 184(1–2):323– 365, October 2022. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-022-01116-1.

Asperti:1989:ISP

[Asp89] Carlo Asperti. Sur l’intégration stochastique par rapport à une martingale hilbertienne de carré intégrable. (French) [On the stochastic integration compared to a square-integrable Hilbert martingale]. Probability Theory and Related Fields, 81(1):139– 158, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00343740.

Abramovich:2000:SEO

[ASS00] F. Abramovich, T. Sapatinas, and B. W. Silverman. Stochastic expansions in an overcomplete wavelet dictionary. Probability Theory and Related Fields, 117(1):133–144, May 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050268.

Assing:2001:IDL

[Ass01] Sigurd Assing. Infinite-dimensional Langevin equations: uniqueness and rate of convergence for finite-dimensional approximations. Probability Theory and Related Fields, 120(2): 143–167, June 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008778.

Asselah:2005:DPA

[Ass05] Amine Asselah. On the Dirichlet problem for asymmetric zerorange process on increasing domains. Probability Theory and Related Fields, 131(1):62–96, January 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0367-6.

Asselah:2008:LDE

[Ass08] Amine Asselah. Large deviations estimates for self-intersection local times for simple random walk in Z^{3}. Probability Theory and Related Fields, 141(1–2):19–45, May 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0078-x.

Asselah:2019:SLL

[ASS19] Amine Asselah, Bruno Schapira, and Perla Sousi. Strong law of large numbers for the capacity of the Wiener sausage in dimension four. Probability Theory and Related Fields, 173(3– 4):813–858, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0842-0.

Ambrosio:2019:PAD

[AST19] Luigi Ambrosio, Federico Stra, and Dario Trevisan. A PDE approach to a 2-dimensional matching problem. Probability Theory and Related Fields, 173(1–2):433–477, February 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0837-x.

Ambrosio:2009:ESF

[ASZ09] Luigi Ambrosio, Giuseppe Savaré, and Lorenzo Zambotti. Existence and stability for Fokker–Planck equations with logconcave reference measure. Probability Theory and Related Fields, 145(3–4):517–564, November 2009. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-008-0177-3.

Avram:1986:SPR

[AT86] Florin Avram and Murad S. Taqqu. Symmetric polynomials of random variables attracted to an infinitely divisible law. Probability Theory and Related Fields, 71(4):491–500, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00699038.

Ahlberg:2018:SPT

[ATT18] Daniel Ahlberg, Vincent Tassion, and Augusto Teixeira. Sharpness of the phase transition for continuum percolation in R^{2}. Probability Theory and Related Fields, 172(1–2):525–581, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-017-0815-8.

Au:2022:SFG

[Au22] Benson Au. Semicircular families of general covariance from Wigner matrices with permuted entries. Probability Theory and Related Fields, 184(3–4):1167–1196, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01143-y.

Augeri:2021:TAM

[Aug21] Fanny Augeri. A transportation approach to the mean-field approximation. Probability Theory and Related Fields, 180(1–2): 1–32, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01056-2.

Akahori:2008:SEC

[AUY08] Jirô Akahori, Chihiro Uenishi, and Kouji Yano. Stochastic equations on compact groups in discrete negative time. Probability Theory and Related Fields, 140(3–4):569–593, March 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0076-z.

Agresti:2024:CVS

[AV24] Antonio Agresti and Mark Veraar. The critical variational setting for stochastic evolution equations. Probability Theory and Related Fields, 188(3–4):957–1015, April 2024. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-023-01249-x.

Avram:1988:BFG

[Avr88] F. Avram. On bilinear forms in Gaussian random variables and Toeplitz matrices. Probability Theory and Related Fields, 79(1): 37–45, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319101.

Azais:2001:RDM

[AW01] Jean-Marc Azaïs and Mario Wschebor. On the regularity of the distribution of the maximum of one-parameter Gaussian processes. Probability Theory and Related Fields, 119(1): 70–98, January 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00012739.

Adamczak:2015:CIN

[AW15a] Radoslaw Adamczak and Pawel Wolff. Concentration inequalities for non-Lipschitz functions with bounded derivatives of higher order. Probability Theory and Related Fields, 162(3–4):531–586, August 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-014-0579-3;
 http://link.springer.com/content/pdf/10.1007/s00440-
014-0579-3.pdf.

Aizenman:2015:UCD

[AW15b] Michael Aizenman and Simone Warzel. On the ubiquity of the Cauchy distribution in spectral problems. Probability Theory and Related Fields, 163(1–2):61–87, October 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0587-3.

Anderson:2006:CBM

[AZ06] Greg W. Anderson and Ofer Zeitouni. A CLT for a band matrix model. Probability Theory and Related Fields, 134(2): 283–338, February 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0422-3.

Altman:2020:BSR

[AZ20] Henri Elad Altman and Lorenzo Zambotti. Bessel SPDEs and renormalised local times. Probability Theory and Related Fields, 176(3–4):757–807, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-019-00926-0.

Babu:1989:SRL

[Bab89] Gutti Jogesh Babu. Strong representations for LAD estimators in linear models. Probability Theory and Related Fields, 83 (4):547–558, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01845702.

Babu:1991:ATE

[Bab91] Gutti Jogesh Babu. Asymptotic theory for estimators under random censorship. Probability Theory and Related Fields, 90 (2):275–290, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192165.

Bahadoran:1998:HLS

[Bah98] C. Bahadoran. Hydrodynamical limit for spatially heterogeneous simple exclusion processes. Probability Theory and Related Fields, 110(3):287–331, March 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050150.

Bai:1986:LPL

[Bai86] Z. D. Bai. Limiting properties of large system of random linear equations. Probability Theory and Related Fields, 73(4): 539–553, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00324851.

Bailleul:2008:PBR

[Bai08] Ismael Bailleul. Poisson boundary of a relativistic diffusion. Probability Theory and Related Fields, 141(1–2):283–329, May 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-007-0086-x.

Bakhtin:2011:NHN

[Bak11] Yuri Bakhtin. Noisy heteroclinic networks. Probability Theory and Related Fields, 150(1–2):1–42, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0264-0.

Baldi:1986:LDF

[Bal86] Paolo Baldi. Large deviations and functional iterated logarithm law for diffusion processes. Probability Theory and Related Fields, 71(3):435–453, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01000215.

Balder:1989:IDE

[Bal89] Erik J. Balder. Infinite-dimensional extension of a theorem of Komlós. Probability Theory and Related Fields, 81(2):185–188, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00319549.

Bally:1998:CAO

[Bal98] Vlad Bally. Construction of asymptotically optimal controls for control and game problems. Probability Theory and Related Fields, 111(3):453–467, July 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050173.

Banuelos:1987:ECM

[Bañ87] Rodrigro Bañuelos. On an estimate of cranston and McConnell for elliptic diffusions in uniform domains. Probability Theory and Related Fields, 76(3):311–323, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01297488.

Banerjee:2016:BCL

[Ban16] Sayan Banerjee. The Brownian Conga line. Probability Theory and Related Fields, 165(3–4):901–961, August 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0649-1;   http://link.springer.com/
article/10.1007/s00440-015-0649-1.

Barbour:1986:AEB

[Bar86] A. D. Barbour. Asymptotic expansions based on smooth functions in the central limit theorem. Probability Theory and Related Fields, 72(2):289–303, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00699108.

Barchielli:1989:POC

[Bar89] Alberto Barchielli. Probability operators and convolution semigroups of instruments in quantum probability. Probability Theory and Related Fields, 82(1):1–8, June 1989. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00340008.

Barbour:1990:SMD

[Bar90] A. D. Barbour. Stein’s method for diffusion approximations. Probability Theory and Related Fields, 84(3):297–322, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01197887.

Barral:1999:MCA

[Bar99] Julien Barral. Moments, continuité, et analyse multifractale des martingales de Mandelbrot. (French) [Moments, continuity, and multifractal analysis of Mandelbrot martingales]. Probability Theory and Related Fields, 113(4):535–569, March 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050217.

Baraud:2000:MSR

[Bar00a] Yannick Baraud. Model selection for regression on a fixed design. Probability Theory and Related Fields, 117(4):467–493, August 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/PL00008731.

Baryshnikov:2000:SPP

[Bar00b] Y. Baryshnikov. Supporting-points processes and some of their applications. Probability Theory and Related Fields, 117(2): 163–182, June 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050002.

Baryshnikov:2001:GQ

[Bar01] Yu. Baryshnikov. GUEs and queues. Probability Theory and Related Fields, 119(2):256–274, February 2001. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008760.

Bardet:2002:LTC

[Bar02] Jean-Baptiste Bardet. Limit theorems for coupled analytic maps. Probability Theory and Related Fields, 124(2):151–177, October 2002. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/s004400200206.

Baraud:2011:ESR

[Bar11] Yannick Baraud. Estimator selection with respect to Hellingertype risks. Probability Theory and Related Fields, 151(1–2): 353–401, October 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0302-y.

Baraud:2021:TES

[Bar21] Yannick Baraud. Tests and estimation strategies associated to some loss functions. Probability Theory and Related Fields, 180(3–4):799–846, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-021-01065-1.

Baraud:2024:RTB

[Bar24] Yannick Baraud. From robust tests to Bayes-like posterior distributions. Probability Theory and Related Fields, 188(1–2): 159–234, February 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01222-8.

Bass:1988:ULP

[Bas88] R. F. Bass. Uniqueness in law for pure jump Markov processes. Probability Theory and Related Fields, 79(2):271–287, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00320922.

Basak:2023:UTS

[Bas23] Anirban Basak. Upper tail of the spectral radius of sparse Erdős–Rényi graphs. Probability Theory and Related Fields, 187(3–4):885–947, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01232-6.

Bauerschmidt:2013:SMF

[Bau13] Roland Bauerschmidt. A simple method for finite range decomposition of quadratic forms and Gaussian fields. Probability Theory and Related Fields, 157(3–4):817–845, December 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-012-0471-y.

Baur:2016:IPC

[Bau16] Erich Baur. An invariance principle for a class of nonballistic random walks in random environment. Probability Theory and Related Fields, 166(1–2):463–514, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-015-0664-2.

Baxendale:1986:ABS

[Bax86] Peter H. Baxendale. Asymptotic behaviour of stochastic flows of diffeomorphisms: Two case studies. Probability Theory and Related Fields, 73(1):51–85, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01845993.

Baxendale:1989:LER

[Bax89] Peter H. Baxendale. Lyapunov exponents and relative entropy for a stochastic flow of diffeomorphisms. Probability Theory and Related Fields, 81(4):521–554, May 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00367301.

Baxendale:1994:SHB

[Bax94] Peter H. Baxendale. A stochastic Hopf bifurcation. Probability Theory and Related Fields, 99(4):581–616, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01206233.

Barlow:1990:LTB

[BB90] Martin T. Barlow and Richard F. Bass. Local times for Brownian motion on the Sierpinski carpet. Probability Theory and Related Fields, 85(1):91–104, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01377631.

Bose:1991:ABA

[BB91] Arup Bose and G. J. Babu. Accuracy of the bootstrap approximation. Probability Theory and Related Fields, 90(3):301–316, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01193748.

Bass:1995:CBM

[BB95] Richard F. Bass and Krzysztof Burdzy. Conditioned Brownian motion in planar domains. Probability Theory and Related Fields, 101(4):479–493, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01202781.

Bass:1996:CCB

[BB96] Richard F. Bass and Krzysztof Burdzy. A critical case for Brownian slow points. Probability Theory and Related Fields, 105 (1):85–108, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192072.

Beznea:2000:EKR

[BB00] Lucian Beznea and Nicu Boboc. Excessive kernels and Revuz measures. Probability Theory and Related Fields, 117(2): 267–288, June 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050007.

Beznea:2003:SMR

[BB03] Lucian Beznea and Nicu Boboc. Sub-Markovian resolvents under weak duality hypothesis. Probability Theory and Related Fields, 126(3):339–363, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0255-5.

Balogh:2006:BPH

[BB06] József Balogh and Béla Bollobás. Bootstrap percolation on the hypercube. Probability Theory and Related Fields, 134 (4):624–648, April 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-005-0451-6.

Berger:2007:QIP

[BB07a] Noam Berger and Marek Biskup. Quenched invariance principle for simple random walk on percolation clusters. Probability Theory and Related Fields, 137(1–2):83–120, January 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0498-z.

Blachere:2007:IDL

[BB07b] Sébastien Blachère and Sara Brofferio. Internal diffusion limited aggregation on discrete groups having exponential growth. Probability Theory and Related Fields, 137(3–4):323–343, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-006-0009-2.

Baraud:2009:EIR

[BB09] Yannick Baraud and Lucien Birgé. Estimating the intensity of a random measure by histogram type estimators. Probability Theory and Related Fields, 143(1–2):239–284, January 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0126-6.

Banerjee:2021:PHG

[BB21] Sayan Banerjee and Shankar Bhamidi. Persistence of hubs in growing random networks. Probability Theory and Related Fields, 180(3–4):891–953, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01066-
0.

Banerjee:2024:EID

[BB24] Sayan Banerjee and Amarjit Budhiraja. Extremal invariant distributions of infinite Brownian particle systems with rank dependent drifts. Probability Theory and Related Fields, 190(1– 2):79–117, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-024-01305-0.

Banuelos:1990:RLT

[BBB90] Rodrigo Bañuelos, Richard F. Bass, and Krzysztof Burdzy. A representation of local time for Lipschitz surfaces. Probability Theory and Related Fields, 84(4):521–547, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198318.

Bogdan:2003:CSP

[BBC03] Krzysztof Bogdan, Krzysztof Burdzy, and Zhen-Qing Chen. Censored stable processes. Probability Theory and Related Fields, 127(1):89–152, September 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0275-1.

Biskup:2004:PTC

[BBC^{+}04] M. Biskup, Ph. Blanchard, L. Chayes, D. Gandolfo, and T. Krüger. Phase transition and critical behavior in a model of organized criticality. Probability Theory and Related Fields, 128(1):1–41, January 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0269-z.

Benjamini:2007:SC

[BBC07] Itai Benjamini, Krzysztof Burdzy, and Zhen-Qing Chen. Shy couplings. Probability Theory and Related Fields, 137(3–4): 345–377, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0008-3.

Bertoin:2018:MSS

[BBCK18] Jean Bertoin, Timothy Budd, Nicolas Curien, and Igor Kortchemski. Martingales in self-similar growth-fragmentations and their connections with random planar maps. Probability Theory and Related Fields, 172(3–4):663–724, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-017-0818-5.

Bieniek:2012:NEF

[BBF12] Mariusz Bieniek, Krzysztof Burdzy, and Sam Finch. Nonextinction of a Fleming–Viot particle model. Probability Theory and Related Fields, 153(1–2):293–332, June 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0372-5.

Bakry:2012:DDH

[BBG12] Dominique Bakry, François Bolley, and Ivan Gentil. Dimension dependent hypercontractivity for Gaussian kernels. Probability Theory and Related Fields, 154(3–4):845–874, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-011-0387-y.

Banerjee:2024:CED

[BBH24] Sayan Banerjee, Shankar Bhamidi, and Xiangying Huang. Coevolving dynamic networks. Probability Theory and Related Fields, 189(1–2):369–445, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-024-01274-4.

Barron:1999:RBM

[BBM99] Andrew Barron, Lucien Birgé, and Pascal Massart. Risk bounds for model selection via penalization. Probability Theory and Related Fields, 113(3):301–413, February 1999. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050210.

Baccelli:2000:ARI

[BBM00] F. Baccelli, A. Borovkov, and J. Mairesse. Asymptotic results on infinite tandem queueing networks. Probability Theory and Related Fields, 118(3):365–405, November 2000. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008747.

Brennan:2020:PTD

[BBN20] Matthew Brennan, Guy Bresler, and Dheeraj Nagaraj. Phase transitions for detecting latent geometry in random graphs. Probability Theory and Related Fields, 178(3–4):1215–1289, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-020-00998-3.

Bedrossian:2021:ASE

[BBPS21] Jacob Bedrossian, Alex Blumenthal, and Sam Punshon-Smith. Almost-sure enhanced dissipation and uniform-in-diffusivity exponential mixing for advection–diffusion by stochastic Navier– Stokes. Probability Theory and Related Fields, 179(3–4):777– 834, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-020-01010-8.

Balister:2000:DPT

[BBS00] P. N. Balister, B. Bollobás, and A. M. Stacey. Dependent percolation in two dimensions. Probability Theory and Related Fields, 117(4):495–513, August 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/PL00008732.

Berestycki:2014:CBB

[BBS14] Julien Berestycki, Nathanaël Berestycki, and Jason Schweinsberg. Critical branching Brownian motion with absorption: survival probability. Probability Theory and Related Fields, 160(3–4):489–520, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-013-0533-9.

Balister:2016:TBP

[BBS16] Paul Balister, Béla Bollobás, and Paul Smith. The time of bootstrap percolation in two dimensions. Probability Theory and Related Fields, 166(1–2):321–364, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0657-1.

Balazs:2021:LSE

[BBS21] Márton Balázs, Ofer Busani, and Timo Seppäläinen. Local stationarity in exponential last-passage percolation. Probability Theory and Related Fields, 180(1–2):113–162, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01035-7.

Barlow:2016:CQA

[BBT16] Martin Barlow, Krzysztof Burdzy, and Ádám Timár. Comparison of quenched and annealed invariance principles for random conductance model. Probability Theory and Related Fields, 164 (3–4):741–770, April 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0618-8.

Balint:2023:PLG

[BBT23] Péter Bálint, Henk Bruin, and Dalia Terhesiu. Periodic Lorentz gas with small scatterers. Probability Theory and Related Fields, 186(1–2):159–219, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01197-6.

Bhamidi:2014:AMC

[BBW14] Shankar Bhamidi, Amarjit Budhiraja, and Xuan Wang. The augmented multiplicative coalescent, bounded size rules and critical dynamics of random graphs. Probability Theory and Related Fields, 160(3–4):733–796, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0540-x.

Bass:2010:SDD

[BBZQC10] Richard F. Bass, Krzysztof Burdzy, Zhen-Qing, and Martin Hairer Chen. Stationary distributions for diffusions with inert drift. Probability Theory and Related Fields, 146(1–2): 1–47, January 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0182-6.

Bass:2001:SDE

[BC01] Richard F. Bass and Zhen-Qing Chens. Stochastic differential equations for Dirichlet processes. Probability Theory and Related Fields, 121(3):422–446, November 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400100151.

Burdzy:2002:CSC

[BC02] Krzysztof Burdzy and Zhen-Qing Chen. Coalescence of synchronous couplings. Probability Theory and Related Fields, 123 (4):553–578, August 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400200202.

Bertail:2004:EES

[BC04] Patrice Bertail and Stéphan Clémençon. Edgeworth expansions of suitably normalized sample mean statistics for atomic Markov chains. Probability Theory and Related Fields, 130(3): 388–414, November 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0360-0.

Bass:2006:SED

[BC06] Richard F. Bass and Zhen-Qing Chen. Systems of equations driven by stable processes. Probability Theory and Related Fields, 134(2):175–214, February 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0426-z.

Bally:2011:IPF

[BC11a] Vlad Bally and Emmanuelle Clément. Integration by parts formula and applications to equations with jumps. Probability Theory and Related Fields, 151(3–4):613–657, December 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0310-y.

Barlow:2011:CFK

[BC11b] Martin T. Barlow and Jirí Cerný. Convergence to fractional kinetics for random walks associated with unbounded conductances. Probability Theory and Related Fields, 149(3–4):639– 673, April 2011. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/s00440-009-0257-z. See erratum [BC11c].

Barlow:2011:ECF

[BC11c] Martin T. Barlow and Jirí Cerný. Erratum to: Convergence to fractional kinetics for random walks associated with unbounded conductances. Probability Theory and Related Fields, 149(3–4):675–677, April 2011. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s00440-
011-0344-9.pdf. See [BC11b].

Bekjan:2012:IMI

[BC12] Turdebek N. Bekjan and Zeqian Chen. Interpolation and Φmoment inequalities of noncommutative martingales. Probability Theory and Related Fields, 152(1–2):179–206, February 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-010-0319-2.

Borodin:2014:MP

[BC14] Alexei Borodin and Ivan Corwin. Macdonald processes. Probability Theory and Related Fields, 158(1–2):225–400, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-013-0482-3.

Bordenave:2015:LDE

[BC15] Charles Bordenave and Pietro Caputo. Large deviations of empirical neighborhood distribution in sparse random graphs. Probability Theory and Related Fields, 163(1–2):149–222, October 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-014-0590-8.

Barraquand:2017:RWB

[BC17] Guillaume Barraquand and Ivan Corwin. Random-walk in beta-distributed random environment. Probability Theory and Related Fields, 167(3–4):1057–1116, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-016-0699-z.

Bakhtin:2022:DPI

[BC22] Yuri Bakhtin and Hong-Bin Chen. Dynamic polymers: invariant measures and ordering by noise. Probability Theory and Related Fields, 183(1–2):167–227, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-021-01099-5.

Bordenave:2012:CLT

[BCC12] Charles Bordenave, Pietro Caputo, and Djalil Chafaï. Circular law theorem for random Markov matrices. Probability Theory and Related Fields, 152(3–4):751–779, April 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0336-1.

Barraquand:2021:FLG

[BCD21] Guillaume Barraquand, Ivan Corwin, and Evgeni Dimitrov. Fluctuations of the log-gamma polymer free energy with general parameters and slopes. Probability Theory and Related Fields, 181(1-3):113–195, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-021-01073-1.

Brzezniak:1993:PGA

[BCF93] Zdzislaw Brzeźniak, Marek Capiński, and Franco Flandoli. Pathwise global attractors for stationary random dynamical systems. Probability Theory and Related Fields, 95(1):87–102, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01197339.

Brzezniak:2015:QET

[BCF15] Z. Brzeźniak, S. Cerrai, and M. Freidlin. Quasipotential and exit time for 2D stochastic Navier–Stokes equations driven by space time white noise. Probability Theory and Related Fields, 162(3– 4):739–793, August 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0584-6.

Borodin:2018:ADG

[BCF18] Alexei Borodin, Ivan Corwin, and Patrik L. Ferrari. Anisotropic (2 + 1)d growth and Gaussian limits of q-Whittaker processes. Probability Theory and Related Fields, 172(1–2):245–321, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-017-0809-6.

Bella:2019:LTS

[BCF19a] Peter Bella, Alberto Chiarini, and Benjamin Fehrman. A Liouville theorem for stationary and ergodic ensembles of parabolic systems. Probability Theory and Related Fields, 173(3–4):759–812, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0843-z;
 http://link.springer.com/content/pdf/10.1007/s00440-
018-0843-z.pdf.

BenArous:2019:SLA

[BCF19b] Gérard Ben Arous, Manuel Cabezas, and Alexander Fribergh. Scaling limit for the ant in a simple high-dimensional labyrinth. Probability Theory and Related Fields, 174(1–2):553–646, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-018-0876-3.

Boldrighini:1989:CSS

[BCFN89] C. Boldrighini, G. C. Cosimi, S. Frigio, and A. Nogueira. Convergence to a stationary state and diffusion for a charged particle in a standing medium. Probability Theory and Related Fields, 80(4):481–500, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00318904.

Bramson:1986:CRT

[BCG86] Maury Bramson, J. Theodore Cox, and David Griffeath. Consolidation rates for two interacting systems in the plane. Probability Theory and Related Fields, 73(4):613–625, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00324856.

Bramson:1988:OTL

[BCG88] Maury Bramson, J. Theodore Cox, and David Griffeath. Occupation time large deviations of the voter model. Probability Theory and Related Fields, 77(3):401–413, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00319297.

Bobkov:2014:BEB

[BCG14a] Sergey G. Bobkov, Gennadiy P. Chistyakov, and Friedrich Götze. Berry–Esseen bounds in the entropic central limit theorem. Probability Theory and Related Fields, 159(3–4):435–478, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-013-0510-3.

Bobkov:2014:FIC

[BCG14b] Sergey G. Bobkov, Gennadiy P. Chistyakov, and Friedrich Götze. Fisher information and the Central Limit Theorem. Probability Theory and Related Fields, 159(1–2):1–59, June 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-013-0500-5.

Bordenave:2022:CSR

[BCGZ22] Charles Bordenave, Djalil Chafaï, and David García-Zelada. Convergence of the spectral radius of a random matrix through its characteristic polynomial. Probability Theory and Related Fields, 182(3–4):1163–1181, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01079-
9.

Burke:1981:SAS

[BCH81] Murray D. Burke, Sándor Csörgő, and Lajos Horváth. Strong approximations of some biometric estimates under random censorship. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 56(1):87–112, ???? 1981. CODEN ZWV-

GAA. ISSN 0044-3719. URL http://link.springer.com/
article/10.1007/BF00531976. See correction and improve-

ment [BCH88].

Burke:1988:CIS

[BCH88] Murray D. Burke, Sándor Csörgő, and Lajos Horváth. A correction to and improvement of “Strong approximations of some biometric estimates under random censorship”. Probability Theory and Related Fields, 79(1):51–57, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF00319103. See [BCH81].

Beiglbock:2020:GMM

[BCH20] Mathias Beiglböck, Alexander M. G. Cox, and Martin Huesmann. The geometry of multi-marginal Skorokhod embedding. Probability Theory and Related Fields, 176(3–4):1045–1096, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-019-00935-z;   https://link.springer.
com/content/pdf/10.1007/s00440-019-00935-z.pdf.

Barlow:2021:QAT

[BCK21] M. T. Barlow, D. A. Croydon, and T. Kumagai. Quenched and averaged tails of the heat kernel of the two-dimensional uniform spanning tree. Probability Theory and Related Fields, 181(1-3): 57–111, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01078-w.

Backhoff:2020:MFS

[BCL20] Julio Backhoff, Giovanni Conforti, and Christian Léonard. The mean field Schrödinger problem: ergodic behavior, entropy estimates and functional inequalities. Probability Theory and Related Fields, 178(1–2):475–530, October 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-020-00977-8.

Boursier:2023:UCD

[BCL23] Jeanne Boursier, Djalil Chafaï, and Cyril Labbé. Universal cutoff for Dyson Ornstein Uhlenbeck process. Probability Theory and Related Fields, 185(1–2):449–512, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01158-5.

Bally:2019:NUV

[BCP19] Vlad Bally, Lucia Caramellino, and Guillaume Poly. Non universality for the variance of the number of real roots of random trigonometric polynomials. Probability Theory and Related Fields, 174(3–4):887–927, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0869-2.

Benjamini:2004:BTR

[BCR04] Itai Benjamini, Zhen-Qing Chen, and Steffen Rohde. Boundary trace of reflecting Brownian motions. Probability Theory and Related Fields, 129(1):1–17, May 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0318-7.

Banica:2011:SAE

[BCS11] Teodor Banica, Stephen Curran, and Roland Speicher. Stochastic aspects of easy quantum groups. Probability Theory and Related Fields, 149(3–4):435–462, April 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-010-0260-4.pdf.

Bordenave:2018:RWS

[BCS18] Charles Bordenave, Pietro Caputo, and Justin Salez. Random walk on sparse random digraphs. Probability Theory and Related Fields, 170(3–4):933–960, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0796-7.

Bordenave:2019:CET

[BCS19] Charles Bordenave, Pietro Caputo, and Justin Salez. Cutoff at the “entropic time” for sparse Markov chains. Probability Theory and Related Fields, 173(1–2):261–292, February 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0834-0.

Borgs:2012:TBM

[BCT12] Christian Borgs, Jennifer T. Chayes, and Prasad Tetali. Tight bounds for mixing of the Swendsen–Wang algorithm at the Potts transition point. Probability Theory and Related Fields, 152(3– 4):509–557, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0329-0.

Biskup:2021:QIP

[BCW21] Marek Biskup, Xin Chen, and Jian Wang. Quenched invariance principle for a class of random conductance models with longrange jumps. Probability Theory and Related Fields, 180(3–4):

847–889, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01059-z.

Brown:2010:RUA

[BCZ^{+}10] Lawrence Brown, Tony Cai, Ren Zhang, Linda Zhao, and Harrison Zhou. The root-unroot algorithm for density estimation as implemented via wavelet block thresholding. Probability Theory and Related Fields, 146(3–4):401–433, March 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0194-2.

Brennan:1987:SII

[BD87] Michael D. Brennan and Richard Durrett. Splitting intervals II: Limit laws for lengths. Probability Theory and Related Fields, 75(1):109–127, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00320085.

Bramson:1988:SPS

[BD88] Maury Bramson and Rick Durrett. A simple proof of the stability criterion of Gray and Griffeath. Probability Theory and Related Fields, 80(2):293–298, December 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00356107.

Berkes:1989:ASW

[BD89] István Berkes and H. Dehling. Almost sure and weak invariance principles for random variables attracted by a stable law. Probability Theory and Related Fields, 83(3):331–353, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00964369.

Barbu:2002:TPS

[BD02] Viorel Barbu and Giuseppe Da Prato. The two phase stochastic Stefan problem. Probability Theory and Related Fields, 124(4): 544–560, December 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-002-0232-4.

Bouleau:2011:ALP

[BD11] Nicolas Bouleau and Laurent Denis. Application of the lent particle method to Poisson-driven SDEs. Probability Theory and Related Fields, 151(3–4):403–433, December 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0303-x.

Berger:2014:QIP

[BD14] Noam Berger and Jean-Dominique Deuschel. A quenched invariance principle for non-elliptic random walk in i.i.d. balanced random environment. Probability Theory and Related Fields, 158(1– 2):91–126, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0478-4.

Bourgade:2020:DOB

[BD20] P. Bourgade and G. Dubach. The distribution of overlaps between eigenvectors of Ginibre matrices. Probability Theory and Related Fields, 177(1–2):397–464, June 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00953-x.

Borodin:2023:BTP

[BD23] Alexei Borodin and Maurice Duits. Biased 2 × 2 periodic Aztec diamond and an elliptic curve. Probability Theory and Related Fields, 187(1–2):259–315, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-023-01195-
8.

Beffara:2012:SDP

[BDC12] Vincent Beffara and Hugo Duminil-Copin. The self-dual point of the two-dimensional random-cluster model is critical for q ≥ 1. Probability Theory and Related Fields, 153(3–4):511–542, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-011-0353-8.

Boue:2000:LDS

[BDE00] Michelle Boué, Paul Dupuis, and Richard S. Ellis. Large deviations for small noise diffusions with discontinuous statistics.

Probability Theory and Related Fields, 116(1):125–149, January 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008720.

Buraczewski:2009:THS

[BDG^{+}09] Dariusz Buraczewski, Ewa Damek, Yves Guivarc’h, Andrzej Hulanicki, and Roman Urban. Tail-homogeneity of stationary measures for some multidimensional stochastic recursions. Probability Theory and Related Fields, 145(3–4):385–420, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-008-0172-8.

Buraczewski:2010:CSL

[BDG10] Dariusz Buraczewski, Ewa Damek, and Yves Guivarc’h. Convergence to stable laws for a class of multidimensional stochastic recursions. Probability Theory and Related Fields, 148(3–4): 333–402, November 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0233-7.

Butkovsky:2021:ASS

[BDG21] Oleg Butkovsky, Konstantinos Dareiotis, and Máté Gerencsér. Approximation of SDEs: a stochastic sewing approach. Probability Theory and Related Fields, 181(4):975–1034, December 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01080-2.

Bhamidi:2017:ELL

[BDN17] Shankar Bhamidi, Partha S. Dey, and Andrew B. Nobel. Energy landscape for large average submatrix detection problems in Gaussian random matrices. Probability Theory and Related Fields, 168(3–4):919–983, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0766-0.

Bolthausen:1993:CPM

[BDS93] Erwin Bolthausen, Jean-Dominique Deuschel, and Uwe Schmock. Convergence of path measures arising from a mean field or polaron type interaction. Probability Theory and Related Fields, 95(3):283–310, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192166.

Boutillier:2010:CII

[BdT10] Cédric Boutillier and Béatrice de Tilière. The critical Zinvariant Ising model via dimers: the periodic case. Probability Theory and Related Fields, 147(3–4):379–413, July 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
009-0210-1.

Boutillier:2019:IIM

[BdTR19] Cédric Boutillier, Béatrice de Tilière, and Kilian Raschel. The Z-invariant Ising model via dimers. Probability Theory and Related Fields, 174(1–2):235–305, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0861-x.

Broutin:2021:LMI

[BDW21] Nicolas Broutin, Thomas Duquesne, and Minmin Wang. Limits of multiplicative inhomogeneous random graphs and Lévy trees: limit theorems. Probability Theory and Related Fields, 181(4): 865–973, December 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01075-z.

Bao:2017:DCR

[BE17a] Zhigang Bao and László Erdős. Delocalization for a class of random block band matrices. Probability Theory and Related Fields, 167(3–4):673–776, April 2017. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s00440-
015-0692-y.pdf.

Bartholdi:2017:PFB

[BE17b] Laurent Bartholdi and Anna Erschler. Poisson–Furstenberg boundary and growth of groups. Probability Theory and Related Fields, 168(1–2):347–372, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Beiglbock:2018:GDC

[BEES18] Mathias Beiglböck, Manu Eder, Christiane Elgert, and Uwe Schmock. Geometry of distribution-constrained optimal stopping problems. Probability Theory and Related Fields, 172 (1–2):71–101, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0805-x;
 http://link.springer.com/content/pdf/10.1007/s00440-
017-0805-x.pdf.

Bovier:2001:MSD

[BEGK01] Anton Bovier, Michael Eckhoff, Véronique Gayrard, and Markus Klein. Metastability in stochastic dynamics of disordered mean-field models. Probability Theory and Related Fields, 119(1):99–161, January 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/PL00012740.

Bell:1986:RBD

[Bel86] Denis Bell. On the relationship between differentiability and absolute continuity of measures on R^{n}. Probability Theory and Related Fields, 72(3):417–424, June 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00334194.

Belsley:1998:RCR

[Bel98] Eric David Belsley. Rates of convergence of random walk on distance regular graphs. Probability Theory and Related Fields, 112(4):493–533, December 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050198.

Belinschi:2008:LDF

[Bel08] Serban Teodor Belinschi. The Lebesgue decomposition of the free additive convolution of two probability distributions. Probability Theory and Related Fields, 142(1–2):125–150, September 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-007-0100-3.

Belius:2013:GFC

[Bel13] David Belius. Gumbel fluctuations for cover times in the discrete torus. Probability Theory and Related Fields, 157(3–4): 635–689, December 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0467-7.

Bender:2010:ESL

[Ben10] Martin Bender. Edge scaling limits for a family of non-Hermitian random matrix ensembles. Probability Theory and Related Fields, 147(1–2):241–271, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
009-0207-9.

Benigni:2021:FEM

[Ben21] Lucas Benigni. Fermionic eigenvector moment flow. Probability Theory and Related Fields, 179(3–4):733–775, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01018-0.

Blumenthal:2023:PBC

[BEN23] Alex Blumenthal, Maximilian Engel, and Alexandra Neamtu. On the pitchfork bifurcation for the Chafee–Infante equation with additive noise. Probability Theory and Related Fields, 187(3–4):603–627, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01235-3.

Berger:1986:ABC

[Ber86a] Erich Berger. Asymptotic behaviour of a class of stochastic approximation procedures. Probability Theory and Related Fields, 71(4):517–552, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00699040.

Berthuet:1986:EPG

[Ber86b] R. Berthuet. Étude de processus généralisant l’Aire de Lévy. (French) [Study of a process generalizing the Lévy area]. Probability Theory and Related Fields, 73(3):463–480, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00776243.

Berbee:1987:CIC

[Ber87a] Henry Berbee. Chains with infinite connections: Uniqueness and Markov representation. Probability Theory and Related Fields, 76(2):243–253, October 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00319986.

Berbee:1987:CRS

[Ber87b] Henry Berbee. Convergence rates in the strong law for bounded mixing sequences. Probability Theory and Related Fields, 74 (2):255–270, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00569992.

Berman:1987:ESG

[Ber87c] Simeon M. Berman. Extreme sojourns of a Gaussian process with a point of maximum variance. Probability Theory and Related Fields, 74(1):113–124, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01845642.

Bertoin:1989:APD

[Ber89] Jean Bertoin. Applications des processus de Dirichlet aux temps locaux et temps locaux d’intersection d’un mouvement Brownien. (French) [Applications of Dirichlet processes in local time and local intersection time of Brownian motion]. Probability Theory and Related Fields, 80(3):433–460, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01794433.

Berger:1990:ASI

[Ber90a] Erich Berger. An almost sure invariance principle for stationary ergodic sequences of Banach space valued random variables. Probability Theory and Related Fields, 84(2):161–201, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01197844.

Bertoin:1990:EDT

[Ber90b] Jean Bertoin. Excursions of a BES_{0}(d) and its drift term (0 < d < 1). Probability Theory and Related Fields, 84(2):231–250, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01197846.

Bertoin:1990:TLI

[Ber90c] Jean Bertoin. Temps locaux d’intersection de mouvements browniens perturbés et applications. (French) [Local intersection times of perturbed Brownian motions and applications]. Probability Theory and Related Fields, 85(1):33–42, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01377626.

Beran:1992:SPR

[Ber92] Rudolf Beran. Simultaneous prediction regions. Probability Theory and Related Fields, 92(2):231–246, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01194922.

Berlinet:1993:HHO

[Ber93a] Alain Berlinet. Hierarchies of higher order kernels. Probability Theory and Related Fields, 94(4):489–504, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192560.

Bertoin:1993:LPN

[Ber93b] Jean Bertoin. Lévy processes with no positive jumps at an increase time. Probability Theory and Related Fields, 96(1): 123–135, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01195886.

Berkes:1995:ASC

[Ber95a] István Berkes. On the almost sure central limit theorem and domains of attraction. Probability Theory and Related Fields, 102(1):1–17, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01295218.

Bertoin:1995:SPB

[Ber95b] Jean Bertoin. Sample path behaviour in connection with generalized arcsine laws. Probability Theory and Related Fields, 103 (3):317–327, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01195477.

Bertoin:1997:REM

[Ber97] Jean Bertoin. Regenerative embedding of Markov sets. Probability Theory and Related Fields, 108(4):559–571, August 1997.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050121.

Bertoin:1999:IIR

[Ber99] Jean Bertoin. Intersection of independent regenerative sets. Probability Theory and Related Fields, 114(1):97–121, April 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050223.

Bertoin:2000:FPC

[Ber00] Jean Bertoin. A fragmentation process connected to Brownian motion. Probability Theory and Related Fields, 117(2):289–301, June 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050008.

Bernabei:2001:ABR

[Ber01a] M. S. Bernabei. Anomalous behaviour for the random corrections to the cumulants of random walks in fluctuating random media. Probability Theory and Related Fields, 119(3):410–432, March 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/PL00008765.

Bertoin:2001:HFP

[Ber01b] Jean Bertoin. Homogeneous fragmentation processes. Probability Theory and Related Fields, 121(3):301–318, November 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100152.

Berlinkov:2003:EPD

[Ber03] Artemi Berlinkov. Exact packing dimension in random recursive constructions. Probability Theory and Related Fields, 126(4): 477–496, August 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0281-3.

Berard:2005:GAR

[Bér05] Jean Bérard. Genetic algorithms in random environments: two examples. Probability Theory and Related Fields, 133(1):123– 140, September 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0419-y.

Berger:2019:NRW

[Ber19] Quentin Berger. Notes on random walks in the Cauchy domain of attraction. Probability Theory and Related Fields, 175(1– 2):1–44, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0887-0.

Bertoin:2020:HLR

[Ber20] Jean Bertoin. How linear reinforcement affects Donsker’s theorem for empirical processes. Probability Theory and Related Fields, 178(3–4):1173–1192, December 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-020-01001-9.

Bertoin:2021:CHL

[Ber21a] Jean Bertoin. Correction to: How linear reinforcement affects Donsker’s theorem for empirical processes. Probability Theory and Related Fields, 181(1-3):291–310, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01048-2.

Bertoin:2021:SES

[Ber21b] Jean Bertoin. Scaling exponents of step-reinforced random walks. Probability Theory and Related Fields, 179(1–2):295–315, February 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-020-01008-2.

Besdziek:1991:SAS

[Bes91] N. Besdziek. Strong approximations of semimartingales by processes with independent increments. Probability Theory and Related Fields, 87(4):489–520, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01304277.

Bass:2000:MVS

[BES00] Richard F. Bass, Nathalie Eisenbaum, and Zhan Shi. The most visited sites of symmetric stable processes. Probability Theory and Related Fields, 116(3):391–404, March 2000.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050255.

Barany:1988:SCH

[BF88] Imre Bárány and Zoltán Füredi. On the shape of the convex hull of random points. Probability Theory and Related Fields, 77(2): 231–240, February 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00334039.

Buckdahn:1993:CAA

[BF93] Rainer Buckdahn and Hans Föllmer. A conditional approach to the anticipating Girsanov transformation. Probability Theory and Related Fields, 95(3):311–330, ???? 1993. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192167.

Bally:2011:RPH

[BF11] Vlad Bally and Nicolas Fournier. Regularization properties of the 2D homogeneous Boltzmann equation without cutoff. Probability Theory and Related Fields, 151(3–4):659–704, December 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-010-0311-x.

Buckley:2018:WSG

[BF18] Jeremiah Buckley and Naomi Feldheim. The winding of stationary Gaussian processes. Probability Theory and Related Fields, 172(1–2):583–614, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0816-7;   http://link.springer.com/content/pdf/
10.1007/s00440-017-0816-7.pdf.

Berestycki:2018:RKR

[BFF18] J. Berestycki, M. C. Fittipaldi, and J. Fontbona. Ray–Knight representation of flows of branching processes with competition by pruning of Lévy trees. Probability Theory and Related Fields, 172(3–4):725–788, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0819-4.

Breit:2019:SSC

[BFHM19] Dominic Breit, Eduard Feireisl, Martina Hofmanová, and Bohdan Maslowski. Stationary solutions to the compressible Navier–Stokes system driven by stochastic forces. Probability Theory and Related Fields, 174(3–4):981–1032, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-018-0875-4.

Bolthausen:2009:CUS

[BFO09] Erwin Bolthausen, Tadahisa Funaki, and Tatsushi Otobe. Concentration under scaling limits for weakly pinned Gaussian random walks. Probability Theory and Related Fields, 143(3–4): 441–480, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0132-8.

Benoist:2019:IMQ

[BFPP19] T. Benoist, M. Fraas, Y. Pautrat, and C. Pellegrini. Invariant measure for quantum trajectories. Probability Theory and Related Fields, 174(1–2):307–334, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0862-9.

Bufetov:2018:EPM

[BFQ18] Alexander I. Bufetov, Shilei Fan, and Yanqi Qiu. Equivalence of Palm measures for determinantal point processes governed by Bergman kernels. Probability Theory and Related Fields, 172 (1–2):31–69, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0803-z.

Baker:1986:DRG

[BG86a] C. R. Baker and A. F. Gualtierotti. Discrimination with respect to a Gaussian process. Probability Theory and Related Fields, 71 (2):159–182, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00332309.

Bojdecki:1986:LEV

[BG86b] Tomasz Bojdecki and Luis G. Gorostiza. Langevin equations for L′-valued Gaussian processes and fluctuation limits of infinite particle systems. Probability Theory and Related Fields, 73(2):

227–244, September 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00339938.

Bentkus:1993:SCC

[BG93] Vidmantas Bentkus and Friedrich Götze. On smoothness conditions and convergence rates in the CLT in Banach spaces. Probability Theory and Related Fields, 96(2):137–151, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192130.

Bentkus:1997:URC

[BG97a] V. Bentkus and F. Götze. Uniform rates of convergence in the CLT for quadratic forms in multidimensional spaces. Probability Theory and Related Fields, 109(3):367–416, November 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050136.

Bovier:1997:RPH

[BG97b] Anton Bovier and Véronique Gayrard. The retrieval phase of the Hopfield model: A rigorous analysis of the overlap distribution^{∗}. Probability Theory and Related Fields, 107(1):61–98, January 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050077.

Bertini:1999:LTB

[BG99a] Lorenzo Bertini and Giambattista Giacomin. On the long time behavior of the stochastic heat equation. Probability Theory and Related Fields, 114(3):279–289, June 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050226.

Bobkov:1999:DIP

[BG99b] S. G. Bobkov and F. Götze. Discrete isoperimetric and Poincarétype inequalities. Probability Theory and Related Fields, 114 (2):245–277, May 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050225.

Berglund:2002:PDD

[BG02a] Nils Berglund and Barbara Gentz. Pathwise description of dynamic pitchfork bifurcations with additive noise. Probability Theory and Related Fields, 122(3):341–388, March 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100174.

Bertacchi:2002:EIR

[BG02b] Daniela Bertacchi and Giambattista Giacomin. Enhanced interface repulsion from quenched hard-wall randomness. Probability Theory and Related Fields, 124(4):487–516, December 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200223.

Benaych-Georges:2007:IDD

[BG07a] Florent Benaych-Georges. Infinitely divisible distributions for rectangular free convolution: classification and matricial interpretation. Probability Theory and Related Fields, 139(1–2): 143–189, September 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0042-1.

Bobkov:2007:CIL

[BG07b] Sergey G. Bobkov and Friedrich Götze. Concentration inequalities and limit theorems for randomized sums. Probability Theory and Related Fields, 137(1–2):49–81, January 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0500-9.

Benaych-Georges:2009:RRM

[BG09] Florent Benaych-Georges. Rectangular random matrices, related convolution. Probability Theory and Related Fields, 144 (3–4):471–515, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0152-z.

Bordenave:2013:LDE

[BG13a] Charles Bordenave and Alice Guionnet. Localization and delocalization of eigenvectors for heavy-tailed random matrices.

Probability Theory and Related Fields, 157(3–4):885–953, December 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-012-0473-9.

Borodin:2013:MPI

[BG13b] Alexei Borodin and Vadim Gorin. Markov processes of infinitely many nonintersecting random walks. Probability Theory and Related Fields, 155(3–4):935–997, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0417-4.

Benaim:2017:SRD

[BG17] Michel Benaïm and Carl-Erik Gauthier. Self-repelling diffusions on a Riemannian manifold. Probability Theory and Related Fields, 169(1–2):63–104, October 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-016-0717-1.

Boedihardjo:2024:LGT

[BG24] H. Boedihardjo and X. Geng. On the lack of Gaussian tail for rough line integrals along fractional Brownian paths. Probability Theory and Related Fields, 188(3–4):1287–1313, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01242-4.

Birkner:2010:QLD

[BGdH10] Matthias Birkner, Andreas Greven, and Frank den Hollander. Quenched large deviation principle for words in a letter sequence. Probability Theory and Related Fields, 148(3–4):403– 456, November 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0235-5. See correction

[BGdH23].

Birkner:2023:CQL

[BGdH23] Matthias Birkner, Andreas Greven, and Frank den Hollander. Correction to: Quenched large deviation principle for words in a letter sequence. Probability Theory and Related Fields, 187(1– 2):523–569, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01212-w. See [BGdH10].

Benaych-Georges:2012:LDE

[BGGM12] F. Benaych-Georges, A. Guionnet, and M. Maida. Large deviations of the extreme eigenvalues of random deformations of matrices. Probability Theory and Related Fields, 154(3–4): 703–751, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0382-3.

Benjamini:2013:LCT

[BGGM13] Itai Benjamini, Ori Gurel-Gurevich, and Ben Morris. Linear cover time is exponentially unlikely. Probability Theory and Related Fields, 155(1–2):451–461, February 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0403-2.

Basdevant:2018:HLH

[BGGS18] A.-L. Basdevant, L. Gerin, J.-B. Gouéré, and A. Singh. From Hammersley’s lines to Hammersley’s trees. Probability Theory and Related Fields, 171(1–2):1–51, June 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0772-2.

Bollobas:1996:RCM

[BGJ96] B. Bollobás, G. Grimmett, and S. Janson. The random-cluster model on the complete graph. Probability Theory and Related Fields, 104(3):283–317, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01213683.

Berger:2019:DCP

[BGL19] Quentin Berger, Giambattista Giacomin, and Hubert Lacoin. Disorder and critical phenomena: the α = 0 copolymer model. Probability Theory and Related Fields, 174(3–4):787–819, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-018-0870-9.

Bacry:2018:CIM

[BGM18] Emmanuel Bacry, Stéphane Gaïffas, and Jean-François Muzy. Concentration inequalities for matrix martingales in continuous time. Probability Theory and Related Fields, 170(1–2):

525–553, February 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0786-9.

Barsky:1991:PHS

[BGN91] David J. Barsky, Geoffrey R. Grimmett, and Charles M. Newman. Percolation in half-spaces: equality of critical densities and continuity of the percolation probability. Probability Theory and Related Fields, 90(1):111–148, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01321136.

Bovier:1994:GSH

[BGP94] Anton Bovier, Véronique Gayrard, and Pierre Picco. Gibbs states of the Hopfield model in the regime of perfect memory. Probability Theory and Related Fields, 100(3):329–363, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01193704.

Bovier:1995:LDP

[BGP95] Anton Bovier, Véronique Gayrard, and Pierre Picco. Large deviation principles for the Hopfield model and the Kac–Hopfield model. Probability Theory and Related Fields, 101(4):511–546, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01202783.

Berger:2003:SBR

[BGP03] Noam Berger, Nina Gantert, and Yuval Peres. The speed of biased random walk on percolation clusters. Probability Theory and Related Fields, 126(2):221–242, June 2003. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0258-2.

Bertini:2014:SRL

[BGP14] Lorenzo Bertini, Giambattista Giacomin, and Christophe Poquet. Synchronization and random long time dynamics for mean-field plane rotators. Probability Theory and Related Fields, 160(3–4):593–653, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-013-0536-6.

Brazitikos:2024:HSD

[BGP24] Silouanos Brazitikos, Apostolos Giannopoulos, and Minas Pafis. Half-space depth of log-concave probability measures. Probability Theory and Related Fields, 188(1–2):309–336, February 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01236-2.

Barbour:2000:IEM

[BGR00] A. D. Barbour, R. M. Gerrard, and G. Reinert. Iterates of expanding maps. Probability Theory and Related Fields, 116(2): 151–180, February 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008725.

Bogdan:2015:BET

[BGR15] Krzysztof Bogdan, Tomasz Grzywny, and Michal Ryznar. Barriers, exit time and survival probability for unimodal Lévy processes. Probability Theory and Related Fields, 162(1–2):155–198, June 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00440-014-0568-6.pdf.

Benaych-Georges:2016:OSR

[BGR16] Florent Benaych-Georges and Jean Rochet. Outliers in the Single Ring Theorem. Probability Theory and Related Fields, 165 (1–2):313–363, June 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0632-x.

Bovier:2013:CEP

[BGS13a] Anton Bovier, Véronique Gayrard, and Adéla Svejda. Convergence to extremal processes in random environments and extremal ageing in SK models. Probability Theory and Related Fields, 157(1–2):251–283, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0456-x.

Broman:2013:ESB

[BGS13b] Erik I. Broman, Christophe Garban, and Jeffrey E. Steif. Exclusion sensitivity of Boolean functions. Probability Theory and Related Fields, 155(3–4):621–663, April 2013. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0409-9.

Bernardin:2016:OTL

[BGS16] Cédric Bernardin, Patrícia Gonçalves, and Sunder Sethuraman. Occupation times of long-range exclusion and connections to KPZ class exponents. Probability Theory and Related Fields, 166(1–2):365–428, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0661-5.

Bolley:2007:QCI

[BGV07] François Bolley, Arnaud Guillin, and Cédric Villani. Quantitative concentration inequalities for empirical measures on noncompact spaces. Probability Theory and Related Fields, 137(3– 4):541–593, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0004-7.

Bodineau:2013:MDI

[BGW13] Thierry Bodineau, Benjamin Graham, and Marc Wouts. Metastability in the dilute Ising model. Probability Theory and Related Fields, 157(3–4):955–1009, December 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0474-8.

Braker:1998:APR

[BH98] H. Bräker and T. Hsing. On the area and perimeter of a random convex hull in a bounded convex set. Probability Theory and Related Fields, 111(4):517–550, August 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050176.

Bass:2000:PUR

[BH00] Richard F. Bass and Elton P. Hsu. Pathwise uniqueness for reflecting Brownian motion in Euclidean domains. Probability Theory and Related Fields, 117(2):183–200, June 2000. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050003.

Bai:2005:BSP

[BH05] Zhidong Bai and Tailen Hsing. The broken sample problem. Probability Theory and Related Fields, 131(4):528–552, April 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-004-0384-5.

Baudoin:2007:VHT

[BH07] Fabrice Baudoin and Martin Hairer. A version of Hörmander’s theorem for the fractional Brownian motion. Probability Theory and Related Fields, 139(3–4):373–395, November 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0035-0.

Briand:2008:QBC

[BH08] Philippe Briand and Ying Hu. Quadratic BSDEs with convex generators and unbounded terminal conditions. Probability Theory and Related Fields, 141(3–4):543–567, July 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0093-y.

Brzezniak:2009:MRS

[BH09] Zdzislaw Brzeźniak and Erika Hausenblas. Maximal regularity for stochastic convolutions driven by Lévy processes. Probability Theory and Related Fields, 145(3–4):615–637, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-008-0181-7.

Bernardin:2013:SPD

[BH13] Cédric Bernardin and François Huveneers. Small perturbation of a disordered harmonic chain by a noise and an anharmonic potential. Probability Theory and Related Fields, 157(1– 2):301–331, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0458-8.

Brown:2001:MMM

[BHR01] Haydyn Brown, David Hobson, and L. C. G. Rogers. The maximum maximum of a martingale constrained by an intermediate law. Probability Theory and Related Fields, 119(4):558–578, April 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/PL00008771.

Berkes:2010:NCL

[BHS10] István Berkes, Lajos Horváth, and Johannes Schauer. Noncentral limit theorems for random selections. Probability Theory and Related Fields, 147(3–4):449–479, July 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
009-0212-z.

Borga:2023:BPL

[BHSY23] Jacopo Borga, Nina Holden, Xin Sun, and Pu Yu. Baxter permuton and Liouville quantum gravity. Probability Theory and Related Fields, 186(3–4):1225–1273, August 2023. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-023-01193-w.

Brzezniak:2019:MSS

[BHW19] Zdzisław Brzeźniak, Fabian Hornung, and Lutz Weis. Martingale solutions for the stochastic nonlinear Schrödinger equation in the energy space. Probability Theory and Related Fields, 174(3–4):1273–1338, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0882-5;   http://link.springer.com/content/pdf/
10.1007/s00440-018-0882-5.pdf.

Biane:1991:QRW

[Bia91] Philippe Biane. Quantum random walk on the dual of SU ( n ). Probability Theory and Related Fields, 89(1):117–129, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01225828.

Biane:1992:ECD

[Bia92] Ph. Biane. Équation de Choquet–Deny sur le dual d’un groupe compact. (French) [The Choquet–Deny equation on the dual of a compact group]. Probability Theory and Related Fields, 94(1):39–51, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01222508.

Bie:1998:EEC

[Bié98] Erwan Saint Loubert Bié. Étude d’une EDPS conduite par un bruit poissonnien. (French) [Study of an EDPS conducted by Poissonian noise]. Probability Theory and Related Fields, 111 (2):287–321, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050169.

Birge:1986:EDU

[Bir86] Lucien Birgé. On estimating a density using Hellinger distance and some other strange facts. Probability Theory and Related Fields, 71(2):271–291, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00332312.

Birge:2014:MSD

[Bir14] Lucien Birgé. Model selection for density estimation with L_{2}- loss. Probability Theory and Related Fields, 158(3–4):533–574, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-013-0488-x.

Bougerol:1999:BBH

[BJ99] Philippe Bougerol and Thierry Jeulin. Brownian bridge on hyperbolic spaces and on homogeneous trees. Probability Theory and Related Fields, 115(1):95–120, August 1999. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050237.

Bougerol:2002:PWC

[BJ02] Philippe Bougerol and Thierry Jeulin. Paths in Weyl chambers and random matrices. Probability Theory and Related Fields, 124(4):517–543, December 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400200221.

Barral:2014:ESL

[BJ14] Julien Barral and Xiong Jin. On exact scaling log-infinitely divisible cascades. Probability Theory and Related Fields, 160 (3–4):521–565, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-013-0534-8.

Barchiesi:2020:SMG

[BJ20] Marco Barchiesi and Vesa Julin. Symmetry of minimizers of a Gaussian isoperimetric problem. Probability Theory and Related Fields, 177(1–2):217–256, June 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-019-00947-9.

Beltran:2017:MPA

[BJL17] J. Beltrán, M. Jara, and C. Landim. A martingale problem for an absorbed diffusion: the nucleation phase of condensing zero range processes. Probability Theory and Related Fields, 169(3–4):1169–1220, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-016-0749-6.

Bartkiewicz:2011:SLS

[BJMW11] Katarzyna Bartkiewicz, Adam Jakubowski, Thomas Mikosch, and Olivier Wintenberger. Stable limits for sums of dependent infinite variance random variables. Probability Theory and Related Fields, 150(3–4):337–372, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0276-9.

Beiglbock:2022:AMC

[BJP22] M. Beiglböck, B. Jourdain, and G. Pammer. Approximation of martingale couplings on the line in the adapted weak topology. Probability Theory and Related Fields, 183(1–2):359–413, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01103-y.

Bossy:2011:CML

[BJT11] Mireille Bossy, Jean-François Jabir, and Denis Talay. On conditional McKean Lagrangian stochastic models. Probability Theory and Related Fields, 151(1–2):319–351, October 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0301-z.

Bass:1992:LTC

[BK92] Richard F. Bass and Davar Khoshnevisan. Local times on curves and uniform invariance principles. Probability Theory and Related Fields, 92(4):465–492, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01274264.

Bhatt:1993:WCM

[BK93] Abhay G. Bhatt and Rajeeva L. Karandikar. Weak convergence to a Markov process: The martingale approach. Probability Theory and Related Fields, 96(3):335–351, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01292676.

Blount:2001:HCS

[BK01] Douglas Blount and Michael A. Kouritzin. Hölder continuity for spatial and path processes via spectral analysis. Probability Theory and Related Fields, 119(4):589–603, April 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008773.

Bakhtin:2004:DAS

[BK04] Victor Bakhtin and Yuri Kifer. Diffusion approximation for slow motion in fully coupled averaging. Probability Theory and Related Fields, 129(2):157–181, June 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0326-7.

Biskup:2007:PCG

[BK07] Marek Biskup and Roman Kotecký. Phase coexistence of gradient Gibbs states. Probability Theory and Related Fields, 139(1– 2):1–39, September 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0013-6.

Banuelos:2008:TES

[BK08] Rodrigo Bañuelos and Tadeusz Kulczycki. Trace estimates for stable processes. Probability Theory and Related Fields, 142 (3–4):313–338, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0106-x.

Becker:2012:SIL

[BK12] Mathias Becker and Wolfgang König. Self-intersection local times of random walks: exponential moments in subcritical dimensions. Probability Theory and Related Fields, 154(3–4):

585–605, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0377-0.

Banerjee:2017:RMM

[BK17a] Sayan Banerjee and Wilfrid S. Kendall. Rigidity for Markovian maximal couplings of elliptic diffusions. Probability Theory and Related Fields, 168(1–2):55–112, June 2017. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/content/pdf/10.
1007/s00440-016-0706-4.pdf.

Belius:2017:SOT

[BK17b] David Belius and Nicola Kistler. The subleading order of two dimensional cover times. Probability Theory and Related Fields, 167(1–2):461–552, February 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0689-6;   http://link.springer.com/
article/10.1007/s00440-015-0689-6.

Biskup:2018:MCA

[BKdS18] Marek Biskup, Wolfgang König, and Renato S. dos Santos. Mass concentration and aging in the parabolic Anderson model with doubly-exponential tails. Probability Theory and Related Fields, 171(1–2):251–331, June 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0777-x.

Bogdan:2008:ESH

[BKK08] Krzysztof Bogdan, Tadeusz Kulczycki, and Mateusz Kwaśnicki. Estimates and structure of α-harmonic functions. Probability Theory and Related Fields, 140(3–4):345–381, March 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0067-0.

Bogdan:2024:ADL

[BKM24] Krzysztof Bogdan, Tomasz Komorowski, and Lorenzo Marino. Anomalous diffusion limit for a kinetic equation with a thermostatted interface. Probability Theory and Related Fields, 189 (1–2):721–769, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01251-3.

Berger:2005:GDT

[BKMP05] Noam Berger, Claire Kenyon, Elchanan Mossel, and Yuval Peres. Glauber dynamics on trees and hyperbolic graphs. Probability Theory and Related Fields, 131(3):311–340, March 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0369-4.

Barbour:2014:MDE

[BKN14] A. D. Barbour, E. Kowalski, and A. Nikeghbali. Mod-discrete expansions. Probability Theory and Related Fields, 158(3–4): 859–893, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0498-8.

Benjamini:2021:AED

[BKP21] Itai Benjamini, Yoav Krauz, and Elliot Paquette. Anchored expansion of Delaunay complexes in real hyperbolic space and stationary point processes. Probability Theory and Related Fields, 181(1-3):197–209, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-021-01076-y.

Bass:2010:CSM

[BKU10] Richard F. Bass, Takashi Kumagai, and Toshihiro Uemura. Convergence of symmetric Markov chains on Z^{d}. Probability Theory and Related Fields, 148(1–2):107–140, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-009-0224-8.

Bryc:2023:MPR

[BKWW23] Włodek Bryc, Alexey Kuznetsov, Yizao Wang, and Jacek Wesołowski. Markov processes related to the stationary measure for the open KPZ equation. Probability Theory and Related Fields, 185(1–2):353–389, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-022-01110-
7.

Bloemendal:2016:PCS

[BKYY16] Alex Bloemendal, Antti Knowles, Horng-Tzer Yau, and Jun Yin. On the principal components of sample covariance matrices. Probability Theory and Related Fields, 164(1–2):459– 552, February 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0616-x.

Burdzy:1990:NIE

[BL90] Krzysztof Burdzy and Gregory F. Lawler. Non-intersection exponents for Brownian paths. Probability Theory and Related Fields, 84(3):393–410, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01197892.

Barchielli:1991:QAH

[BL91] A. Barchielli and G. Lupieri. A quantum analogue of Hunt’s representation theorem for the generator of convolution semigroups on Lie groups. Probability Theory and Related Fields, 88 (2):167–194, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01212558.

Bobkov:1997:PIT

[BL97] S. Bobkov and M. Ledoux. Poincaré’s inequalities and Talagrand’s concentration phenomenon for the exponential distribution. Probability Theory and Related Fields, 107(3):383–400, March 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050090.

Bertoin:2000:BSC

[BL00] Jean Bertoin and Jean-François Le Gall. The Bolthausen– Sznitman coalescent and the genealogy of continuous-state branching processes. Probability Theory and Related Fields, 117 (2):249–266, June 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050006.

Burdzy:2001:SBM

[BL01] Krzysztof Burdzy and Jean-François Le Gall. Super-Brownian motion with reflecting historical paths. Probability Theory and Related Fields, 121(4):447–491, December 2001. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100157.

Bertoin:2003:SFA

[BL03] Jean Bertoin and Jean-François Le Gall. Stochastic flows associated to coalescent processes. Probability Theory and Related Fields, 126(2):261–288, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0264-4.

Bass:2007:SPC

[BL07] Richard F. Bass and Alexander Lavrentiev. The submartingale problem for a class of degenerate elliptic operators. Probability Theory and Related Fields, 139(3–4):415–449, November 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-006-0047-9.

Barbour:2012:LLN

[BL12a] A. D. Barbour and M. J. Luczak. A law of large numbers approximation for Markov population processes with countably many types. Probability Theory and Related Fields, 153(3–4): 727–757, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0359-2.

Beltran:2012:MRC

[BL12b] J. Beltrán and C. Landim. Metastability of reversible condensed zero range processes on a finite set. Probability Theory and Related Fields, 152(3–4):781–807, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-010-0337-0.

Beltran:2015:MAM

[BL15] J. Beltrán and C. Landim. A martingale approach to metastability. Probability Theory and Related Fields, 161(1–2):267– 307, February 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0549-9.

Buckingham:2019:TP

[BL19] Robert Buckingham and Karl Liechty. The k-tacnode process. Probability Theory and Related Fields, 175(1–2):341–395, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-018-0885-2.

Baik:2021:PTG

[BL21] Jinho Baik and Zhipeng Liu. Periodic TASEP with general initial conditions. Probability Theory and Related Fields, 179(3– 4):1047–1144, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-020-01004-6.

Bedrossian:2024:SMS

[BL24a] Jacob Bedrossian and Kyle Liss. Stationary measures for stochastic differential equations with degenerate damping. Probability Theory and Related Fields, 189(1–2):101–178, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01265-5.

Berger:2024:SLD

[BL24b] Quentin Berger and Alexandre Legrand. Scaling limit of the disordered generalized Poland–Scheraga model for DNA denaturation. Probability Theory and Related Fields, 190(1–2): 179–258, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-024-01304-1.

Blake:1988:GMG

[Bla88] Louis H. Blake. Generalized martingales, generalized Markov chains and generalized harmonic functions. Probability Theory and Related Fields, 79(3):405–430, October 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00342233.

Blache:2005:BSD

[Bla05] Fabrice Blache. Backward stochastic differential equations on manifolds. Probability Theory and Related Fields, 132(3): 391–437, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0400-9.

Bar-Lev:1992:PVF

[BLBE92] Shaul K. Bar-Lev, Daoud Bshouty, and Peter Enis. On polynomial variance functions. Probability Theory and Related Fields, 94(1):69–82, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01222510.

Blei:1989:MLM

[Ble89] Ron C. Blei. Multi-linear measure theory and multiple stochastic integration. Probability Theory and Related Fields, 81(4): 569–584, May 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00367304.

Blei:1993:SII

[Ble93] Ron C. Blei. Stochastic integrators indexed by a multidimensional parameter. Probability Theory and Related Fields, 95(2):141–153, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192267.

Bassetti:2011:CLT

[BLM11] Federico Bassetti, Lucia Ladelli, and Daniel Matthes. Central limit theorem for a class of one-dimensional kinetic equations. Probability Theory and Related Fields, 150(1–2):77–109, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-010-0269-8.

Bazaes:2024:SGM

[BLM24] Rodrigo Bazaes, Isabel Lammers, and Chiranjib Mukherjee. Subcritical Gaussian multiplicative chaos in the Wiener space: construction, moments and volume decay. Probability Theory and Related Fields, 190(3-4):753–801, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01271-7.

Berestycki:2023:FBD

[BLQ23] Nathanaël Berestycki, Marcin Lis, and Wei Qian. Free boundary dimers: random walk representation and scaling limit. Probability Theory and Related Fields, 186(3–4):735–812, August 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01203-x.

Benaim:2002:SID

[BLR02] Michel Benaïm, Michel Ledoux, and Olivier Raimond. Selfinteracting diffusions. Probability Theory and Related Fields, 122(1):1–41, January 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400100161.

Bar-Lev:1987:CNE

[BLS87] Shaul K. Bar-Lev and Osnat Stramer. Characterizations of natural exponential families with power variance functions by zero regression properties. Probability Theory and Related Fields, 76 (4):509–522, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00960071.

Bordenave:2013:MIG

[BLS13] Charles Bordenave, Marc Lelarge, and Justin Salez. Matchings on infinite graphs. Probability Theory and Related Fields, 157(1– 2):183–208, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0453-0.

Blunck:2005:PRF

[Blu05] Sönke Blunck. Path regularity for Feller semigroups via Gaussian kernel estimates and generalizations to arbitrary semigroups on C_{0}. Probability Theory and Related Fields, 133(1): 71–97, September 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0415-2.

Bigot:2012:SES

[BLV12] Jérémie Bigot, Jean-Michel Loubes, and Myriam Vimond. Semiparametric estimation of shifts on compact Lie groups for image registration. Probability Theory and Related Fields, 152(3– 4):425–473, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0327-2.

Benes:2016:SLL

[BLV16] Christian Benes, Gregory F. Lawler, and Fredrik Viklund. Scaling limit of the loop-erased random walk Green’s function. Probability Theory and Related Fields, 166(1–2):271–319, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-015-0655-3.

Bordenave:2020:NST

[BLZ20] Charles Bordenave, Gábor Lugosi, and Nikita Zhivotovskiy. Noise sensitivity of the top eigenvector of a Wigner matrix. Probability Theory and Related Fields, 177(3–4):1103–1135, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-020-00970-1.

Booth:2001:IPB

[BM01] Lorna Booth and Ronald Meester. Infinite paths with bounded or recurrent partial sums. Probability Theory and Related Fields, 120(1):118–142, May 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008775.

Barral:2002:MPC

[BM02] Julien Barral and Benoît B. Mandelbrot. Multifractal products of cylindrical pulses. Probability Theory and Related Fields, 124(3):409–430, November 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400200220.

Benjamini:2003:MTS

[BM03] Itai Benjamini and Elchanan Mossel. On the mixing time of a simple random walk on the super critical percolation cluster. Probability Theory and Related Fields, 125(3):408–420, March 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-002-0246-y.

Blath:2005:TPS

[BM05a] Jochen Blath and Peter Mörters. Thick points of super-Brownian motion. Probability Theory and Related Fields, 131 (4):604–630, April 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0387-2.

Burdzy:2005:SBM

[BM05b] Krzysztof Burdzy and Leonid Mytnik. Super-Brownian motion with reflecting historical paths. II. Convergence of approximations. Probability Theory and Related Fields, 133(2): 145–174, October 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0413-4.

Birge:2007:MPG

[BM07] Lucien Birgé and Pascal Massart. Minimal penalties for Gaussian model selection. Probability Theory and Related Fields, 138 (1–2):33–73, May 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0011-8.

Barbe:2008:AEI

[BM08] Ph. Barbe and W. P. McCormick. Asymptotic expansions for infinite weighted convolutions of rapidly varying subexponential distributions. Probability Theory and Related Fields, 141 (1–2):155–180, May 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0082-1.

Boucheron:2011:HDW

[BM11a] Stéphane Boucheron and Pascal Massart. A high-dimensional Wilks phenomenon. Probability Theory and Related Fields, 150(3–4):405–433, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0278-7.

Buchmann:2011:STC

[BM11b] Boris Buchmann and Ross Maller. The small-time Chung– Wichura law for Lévy processes with non-vanishing Brownian component. Probability Theory and Related Fields, 149(1–2): 303–330, February 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0255-1.

Barbe:2012:IPS

[BM12] Ph. Barbe and W. P. McCormick. Invariance principles for some FARIMA and nonstationary linear processes in the domain of a stable distribution. Probability Theory and Related Fields, 154(3–4):429–476, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-011-0374-3.

Bobrowski:2015:TPD

[BM15] Omer Bobrowski and Sayan Mukherjee. The topology of probability distributions on manifolds. Probability Theory and Related Fields, 161(3–4):651–686, April 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0556-x.

Broutin:2016:NEC

[BM16] Nicolas Broutin and Jean-François Marckert. A new encoding of coalescent processes: applications to the additive and multiplicative cases. Probability Theory and Related Fields, 166(1–2): 515–552, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0665-1.

Basak:2017:UMF

[BM17a] Anirban Basak and Sumit Mukherjee. Universality of the meanfield for the Potts model. Probability Theory and Related Fields, 168(3–4):557–600, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-016-0718-0.

Bettinelli:2017:CBS

[BM17b] Jérémie Bettinelli and Grégory Miermont. Compact Brownian surfaces I: Brownian disks. Probability Theory and Related Fields, 167(3–4):555–614, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s00440-
016-0752-y.

Barbier:2019:AIM

[BM19] Jean Barbier and Nicolas Macris. The adaptive interpolation method: a simple scheme to prove replica formulas in Bayesian inference. Probability Theory and Related Fields, 174(3–4): 1133–1185, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0879-0.

Beliaev:2020:NES

[BMM20] Dmitry Beliaev, Michael McAuley, and Stephen Muirhead. On the number of excursion sets of planar Gaussian fields. Probability Theory and Related Fields, 178(3–4):655–698, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00984-9.

Bartlett:2012:RLR

[BMN12] Peter L. Bartlett, Shahar Mendelson, and Joseph Neeman. l_{1}- regularized linear regression: persistence and oracle inequalities. Probability Theory and Related Fields, 154(1–2):193–224, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-011-0367-2.

Boldrighini:1997:ASC

[BMP97] C. Boldrighini, R. A. Minlos, and A. Pellegrinotti. Almost-sure central limit theorem for a Markov model of random walk in dynamical random environment. Probability Theory and Related Fields, 109(2):245–273, October 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050132.

Boldrighini:2004:RWQ

[BMP04] C. Boldrighini, R. A. Minlos, and A. Pellegrinotti. Random walks in quenched i.i.d. space–time random environment are always a.s. diffusive. Probability Theory and Related Fields, 129 (1):133–156, May 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0331-x.

Bahadoran:2020:QCS

[BMRS20] C. Bahadoran, T. Mountford, K. Ravishankar, and E. Saada. Quenched convergence and strong local equilibrium for asymmetric zero-range process with site disorder. Probability Theory and Related Fields, 176(1–2):149–202, February 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00916-2.

Bousquet-Melou:2002:WSP

[BMS02] Mireille Bousquet-Mélou and Gilles Schaeffer. Walks on the slit plane. Probability Theory and Related Fields, 124(3):305– 344, November 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400200205.

Brzezniak:2005:SNB

[BMS05] Zdzislaw Brzeźniak, Bohdan Maslowski, and Jan Seidler. Stochastic nonlinear beam equations. Probability Theory and Related Fields, 132(1):119–149, May 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0392-5.

Buchmann:2008:ASF

[BMS08] Boris Buchmann, Ross Maller, and Alex Szimayer. An almost sure functional limit theorem at zero for a class of Lévy processes normed by the square root function, and applications. Probability Theory and Related Fields, 142(1–2):219–247, September 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-007-0103-0.

Buckdahn:1994:LSD

[BN94] R. Buckdahn and D. Nualart. Linear stochastic differential equations and Wick products. Probability Theory and Related Fields, 99(4):501–526, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01206230.

Burdzy:2002:BMR

[BN02] Krzysztof Burdzy and David Nualart. Brownian motion reflected on Brownian motion. Probability Theory and Related Fields, 122(4):471–493, April 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400100165.

Bull:2013:ACS

[BN13] Adam D. Bull and Richard Nickl. Adaptive confidence sets in L^{2}. Probability Theory and Related Fields, 156(3–4):889–919, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-012-0446-z.

Bufetov:2022:CPA

[BN22] Alexey Bufetov and Peter Nejjar. Cutoff profile of ASEP on a segment. Probability Theory and Related Fields, 183(1–2):229– 253, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01104-x.

Benjamini:2011:CPP

[BNP11] Itai Benjamini, Asaf Nachmias, and Yuval Peres. Is the critical percolation probability local? Probability Theory and Related Fields, 149(1–2):261–269, February 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-009-0251-5.

Buckley:2018:HPZ

[BNPS18] Jeremiah Buckley, Alon Nishry, Ron Peled, and Mikhail Sodin. Hole probability for zeroes of Gaussian Taylor series with finite radii of convergence. Probability Theory and Related Fields, 171 (1–2):377–430, June 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0782-0.

Burdzy:2014:JCA

[BNS14] Krzysztof Burdzy, David Nualart, and Jason Swanson. Joint convergence along different subsequences of the signed cubic variation of fractional Brownian motion. Probability Theory and Related Fields, 159(1–2):237–272, June 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0511-2.

Barndorff-Nielsen:2005:LID

[BNT05] O. E. Barndorff-Nielsen and S. Thorbjørnsen. The Lévy–Itô decomposition in free probability. Probability Theory and Related Fields, 131(2):197–228, February 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0322-y.

Borodin:2009:IDD

[BO09] Alexei Borodin and Grigori Olshanski. Infinite-dimensional diffusions as limits of random walks on partitions. Probability Theory and Related Fields, 144(1–2):281–318, May 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0148-8.

Bobkov:2010:CBN

[Bob10] Sergey G. Bobkov. Convex bodies and norms associated to convex measures. Probability Theory and Related Fields, 147(1– 2):303–332, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0209-7.

Bobkov:2018:BEB

[Bob18] Sergey G. Bobkov. Berry–Esseen bounds and Edgeworth expansions in the central limit theorem for transport distances. Probability Theory and Related Fields, 170(1–2):229–262, February 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-017-0756-2.

Bodineau:2005:SPI

[Bod05] T. Bodineau. Slab percolation for the Ising model. Probability Theory and Related Fields, 132(1):83–118, May 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0391-6.

Boivin:1990:FPP

[Boi90] Daniel Boivin. First passage percolation: The stationary case. Probability Theory and Related Fields, 86(4):491–499, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01198171.

Bolthausen:1986:LAS

[Bol86] E. Bolthausen. Laplace approximations for sums of independent random vectors. Probability Theory and Related Fields, 72 (2):305–318, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00699109.

Bolthausen:1987:LAS

[Bol87] E. Bolthausen. Laplace approximations for sums of independent random vectors. Probability Theory and Related Fields, 76(2): 167–206, October 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319983.

Bolthausen:1990:SRO

[Bol90] Erwin Bolthausen. On self-repellent one dimensional random walks. Probability Theory and Related Fields, 86(4):423–441, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01198167.

Bondesson:1988:RPG

[Bon88] L. Bondesson. A remarkable property of generalized gamma convolutions. Probability Theory and Related Fields, 78(3):321–333, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00334198.

Bondesson:1990:GGC

[Bon90] L. Bondesson. Generalized gamma convolutions and complete monotonicity. Probability Theory and Related Fields, 85(2): 181–194, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01277981.

Bonis:2020:SMN

[Bon20] Thomas Bonis. Stein’s method for normal approximation in Wasserstein distances with application to the multivariate central limit theorem. Probability Theory and Related Fields, 178 (3–4):827–860, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-020-00989-4.

Borodin:1986:CCBa

[Bor86a] A. N. Borodin. On the character of convergence to Brownian local time. I. Probability Theory and Related Fields, 72 (2):231–250, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00699105.

Borodin:1986:CCBb

[Bor86b] A. N. Borodin. On the character of convergence to Brownian local time. II. Probability Theory and Related Fields, 72 (2):251–277, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00699106.

Borell:1988:RPC

[Bor88a] Christer Borell. Real polynomial chaos and absolute continuity. Probability Theory and Related Fields, 77(3):397–400, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00319296.

Borkar:1988:CAA

[Bor88b] Vivek S. Borkar. A convex analytic approach to Markov decision processes. Probability Theory and Related Fields, 78(4):583–602, August 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00353877.

Borell:1991:BMN

[Bor91] Christer Borell. Brownian motion with negative drift and convex level sets in space–time. Probability Theory and Related Fields, 87(3):403–409, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01312218.

Borovkov:2003:LDP

[Bor03] A. A. Borovkov. Large deviations probabilities for random walks in the absence of finite expectations of jumps. Probability Theory and Related Fields, 125(3):421–446, March 2003. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
002-0243-1.

Borell:2008:IBM

[Bor08] Christer Borell. Inequalities of the Brunn–Minkowski type for Gaussian measures. Probability Theory and Related Fields, 140(1–2):195–205, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0062-5.

Basse-OConnor:2016:IDS

[BOR16] Andreas Basse-O’Connor and Jan Rosiński. On infinitely divisible semimartingales. Probability Theory and Related Fields, 164(1–2):133–163, February 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0609-1.

Borga:2020:LCP

[Bor20] Jacopo Borga. Local convergence for permutations and local limits for uniform ρ-avoiding permutations with |ρ| = 3. Probability Theory and Related Fields, 176(1–2):449–531, February 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-019-00922-4.

Bougerol:1988:TLP

[Bou88] Philippe Bougerol. Théorèmes limite pour les systèmes linéaires à coefficients markoviens. (French) [Limit theorems for linear systems with Markov coefficients]. Probability Theory and Related Fields, 78(2):193–221, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00322018.

Bouzar:1990:SLR

[Bou90] Nadjib Bouzar. Some local results on the convergence and the quadratic variation of random sequences. Probability Theory and Related Fields, 86(2):265–275, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01474646.

Bourgade:2010:MFZ

[Bou10] P. Bourgade. Mesoscopic fluctuations of the zeta zeros. Probability Theory and Related Fields, 148(3–4):479–500, November 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-009-0237-3.

Bowditch:2018:ERB

[Bow18] Adam Bowditch. Escape regimes of biased random walks on Galton–Watson trees. Probability Theory and Related Fields, 170(3–4):685–768, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0768-y;
 http://link.springer.com/content/pdf/10.1007/s00440-
017-0768-y.pdf.

Boxler:1989:SVC

[Box89] Petra Boxler. A stochastic version of center manifold theory. Probability Theory and Related Fields, 83(4):509–545, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01845701.

Boyer:2019:ASF

[Boy19] Jean-Baptiste Boyer. Almost sure functional central limit theorem for the linear random walk on the torus. Probability Theory and Related Fields, 173(1–2):651–696, February 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0871-8.

Berkes:1986:ERV

[BP86] István Berkes and Erika Péter. Exchangeable random variables and the subsequence principle. Probability Theory and Related Fields, 73(3):395–413, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00776240.

Bass:1987:UDP

[BP87] R. F. Bass and E. Pardoux. Uniqueness for diffusions with piecewise constant coefficients. Probability Theory and Related Fields, 76(4):557–572, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00960074.

Barlow:1988:BMS

[BP88] Martin T. Barlow and Edwin A. Perkins. Brownian motion on the Sierpinski gasket. Probability Theory and Related Fields, 79(4):543–623, November 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00318785.

Barlow:1989:SMC

[BP89] Martin T. Barlow and Edwin A. Perkins. Symmetric Markov chains in z^{d}: How fast can they move? Probability Theory and Related Fields, 82(1):95–108, June 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00340013.

Benjamini:1994:TIR

[BP94] Itai Benjamini and Yuval Peres. Tree-indexed random walks on groups and first passage percolation. Probability Theory and Related Fields, 98(1):91–112, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01311350.

Buckdahn:1999:SBS

[BP99] Rainer Buckdahn and Shige Peng. Stationary backward stochastic differential equations and associated partial differential equations. Probability Theory and Related Fields, 115(3):383–399, October 1999. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/s004400050242.

Bhatnagar:2015:LMS

[BP15] Nayantara Bhatnagar and Ron Peled. Lengths of monotone subsequences in a Mallows permutation. Probability Theory and Related Fields, 161(3–4):719–780, April 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-014-0559-7.

Borodin:2018:IEJ

[BP18] Alexei Borodin and Leonid Petrov. Inhomogeneous exponential jump model. Probability Theory and Related Fields, 172(1–2): 323–385, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0810-0.

Burdzy:2020:FMG

[BP20] Krzysztof Burdzy and Soumik Pal. Floodings of metric graphs. Probability Theory and Related Fields, 177(1–2):577–620, June 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-020-00974-x.

Barlow:1997:DLA

[BPP97] Martin T. Barlow, Robin Pemantle, and Edwin A. Perkins. Diffusion-limited aggregation on a tree. Probability Theory and Related Fields, 107(1):1–60, January 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050076.

Banuelos:2004:BMK

[BPP04] Rodrigo Banuelos, Michael Pang, and Mihai Pascu. Brownian motion with killing and reflection and the “hot-spots” problem. Probability Theory and Related Fields, 130(1):56–68, September 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-003-0323-x.

Berti:2007:SRG

[BPR07] Patrizia Berti, Luca Pratelli, and Pietro Rigo. Skorohod representation on a given probability space. Probability Theory and Related Fields, 137(3–4):277–288, March 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0018-1.

Berti:2010:AIF

[BPR10] Patrizia Berti, Luca Pratelli, and Pietro Rigo. Atomic intersection of σ-fields and some of its consequences. Probability Theory and Related Fields, 148(1–2):269–283, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-009-0230-x.

Berti:2011:SRT

[BPR11] Patrizia Berti, Luca Pratelli, and Pietro Rigo. A Skorohod representation theorem for uniform distance. Probability Theory and Related Fields, 150(1–2):321–335, June 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0279-6.

Berestycki:2020:CGF

[BPR20] Nathanaël Berestycki, Ellen Powell, and Gourab Ray. A characterisation of the Gaussian free field. Probability Theory and Related Fields, 176(3–4):1259–1301, April 2020. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-019-
00939-9;   https://link.springer.com/content/pdf/10.
1007/s00440-019-00939-9.pdf.

Basrak:2018:IPS

[BPS18] Bojan Basrak, Hrvoje Planinić, and Philippe Soulier. An invariance principle for sums and record times of regularly varying stationary sequences. Probability Theory and Related Fields, 172(3–4):869–914, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0822-9.

Barral:2009:DMC

[BPW09] Julien Barral, Jacques Peyrière, and Zhi-Ying Wen. Dynamics of Mandelbrot cascades. Probability Theory and Related Fields, 144(3–4):615–631, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0156-8.

Buckdahn:2000:VPB

[BQR00] Rainer Buckdahn, Marc Quincampoix, and Aurel Rascanu. Viability property for a backward stochastic differential equation and applications to partial differential equations. Probability Theory and Related Fields, 116(4):485–504, April 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050260.

Bentkus:1990:PLD

[BR90] V. Bentkus and A. Rackauskas. On probabilities of large deviations in Banach spaces. Probability Theory and Related Fields, 86(2):131–154, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01474639.

Baldi:1992:SEE

[BR92] P. Baldi and B. Roynette. Some exact equivalents for the Brownian motion in Hölder norm. Probability Theory and Related Fields, 93(4):457–484, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192717.

Buchta:1997:EIP

[BR97] Christian Buchta and Matthias Reitzner. Equiaffine inner parallel curves of a plane convex body and the convex hulls of randomly chosen points. Probability Theory and Related Fields, 108(3):385–415, July 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050114.

Bogachev:2001:EEM

[BR01] Vladimir I. Bogachev and Michael Röckner. Elliptic equations for measures on infinite dimensional spaces and applications. Probability Theory and Related Fields, 120(4):445–496, August 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008789.

Bollobas:2008:PRJ

[BR08] Béla Bollobás and Oliver Riordan. Percolation on random Johnson–Mehl tessellations and related models. Probability Theory and Related Fields, 140(3–4):319–343, March 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-

007-0066-1. See erratum [BR10].

Bollobas:2010:EPR

[BR10] Béla Bollobás and Oliver Riordan. Erratum to: Percolation on random Johnson–Mehl tessellations and related models. Probability Theory and Related Fields, 146(3–4):567–570, March 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-009-0247-1.pdf. See [BR08].

Basak:2021:STI

[BR21] Anirban Basak and Mark Rudelson. Sharp transition of the invertibility of the adjacency matrices of sparse random graphs. Probability Theory and Related Fields, 180(1–2):233–308, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01038-4.

Blanc-Renaudie:2023:CFD

[BR23] Arthur Blanc-Renaudie. Compactness and fractal dimensions of inhomogeneous continuum random trees. Probability Theory and Related Fields, 185(3–4):961–991, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01138-9.

Barbu:2024:NFP

[BR24] Viorel Barbu and Michael Röckner. Nonlinear Fokker–Planck equations with fractional Laplacian and McKean–Vlasov SDEs with Lévy noise. Probability Theory and Related Fields, 189(3– 4):849–878, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-024-01277-1.

Bramson:1986:LTW

[Bra86] Maury Bramson. Location of the travelling wave for the Kolmogorov equation. Probability Theory and Related Fields, 73(4): 481–515, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00324848.

Bradley:1987:IMR

[Bra87] Richard C. Bradley. Identical mixing rates. Probability Theory and Related Fields, 74(4):497–503, April 1987. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00363512.

Bradley:1989:SPI

[Bra89] Richard C. Bradley. A stationary, pairwise independent, absolutely regular sequence for which the central limit theorem fails. Probability Theory and Related Fields, 81(1):1–10, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00343735.

Breuillard:2005:DDT

[Bre05] E. Breuillard. Distributions diophantiennes et théorème limite local sur R^{d}. (French) [Diophantine distributions and local limit theorem on R^{d}]. Probability Theory and Related Fields, 132(1):13–38, May 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0388-1.

Brettschneider:2008:SMT

[Bre08] Julia Brettschneider. Shannon-McMillan theorems for discrete random fields along curves and lower bounds for surface-order large deviations. Probability Theory and Related Fields, 142 (3–4):443–473, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0112-z.

Bou-Rabee:2024:HJS

[BRM24] Ahmed Bou-Rabee and Peter S. Morfe. Hamilton–Jacobi scaling limits of Pareto peeling in 2D. Probability Theory and Related Fields, 188(1–2):235–307, February 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-023-01234-
4.

Barbu:2011:PRS

[BRR11] Viorel Barbu, Michael Röckner, and Francesco Russo. Probabilistic representation for solutions of an irregular porous media type equation: the degenerate case. Probability Theory and Related Fields, 151(1–2):1–43, October 2011. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-010-0291-x.

Bogachev:1996:GMS

[BRS96] Vladimir I. Bogachev, Michael Röckner, and Byron Schmuland. Generalized Mehler semigroups and applications. Probability Theory and Related Fields, 105(2):193–225, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01203835.

Bogachev:2014:PIG

[BRS14] Vladimir I. Bogachev, Michael Röckner, and Stanislav V. Shaposhnikov. On parabolic inequalities for generators of diffusions with jumps. Probability Theory and Related Fields, 158(1–2): 465–476, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0485-0.

Brunick:2013:ULC

[Bru13] Gerard Brunick. Uniqueness in law for a class of degenerate diffusions with continuous covariance. Probability Theory and Related Fields, 155(1–2):265–302, February 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0398-8.

Brunel:2016:AEC

[Bru16] Victor-Emmanuel Brunel. Adaptive estimation of convex and polytopal density support. Probability Theory and Related Fields, 164(1–2):1–16, February 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0605-5.

Brunel:2019:CEL

[Bru19] Victor-Emmanuel Brunel. Concentration of the empirical level sets of Tukey’s halfspace depth. Probability Theory and Related Fields, 173(3–4):1165–1196, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0850-0.

Bryc:1988:RPP

[Bry88] Włodzimierz Bryc. Remarks on properties of probability distributions determined by conditional moments. Probability Theory and Related Fields, 78(1):51–62, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00718034.

Bryc:2008:ANT

[Bry08] Włodzimierz Bryc. Asymptotic normality for traces of polynomials in independent complex Wishart matrices. Probability Theory and Related Fields, 140(3–4):383–405, March 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0068-z.

Belius:2020:CTC

[BRZ20a] David Belius, Jay Rosen, and Ofer Zeitouni. Correction to: Tightness for the cover time of the two dimensional sphere. Probability Theory and Related Fields, 176(3–4):1439–1444, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-020-00965-y;   https://link.springer.
com/content/pdf/10.1007/s00440-020-00965-y.pdf. See

[BRZ20b].

Belius:2020:TCT

[BRZ20b] David Belius, Jay Rosen, and Ofer Zeitouni. Tightness for the cover time of the two dimensional sphere. Probability Theory and Related Fields, 176(3–4):1357–1437, April 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-

019-00940-2. See correction [BRZ20a].

Baxendale:1988:LDS

[BS88] P. H. Baxendale and D. W. Stroock. Large deviations and stochastic flows of diffeomorphisms. Probability Theory and Related Fields, 80(2):169–215, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00356102.

Boldrighini:1995:DDM

[BS95a] C. Boldrighini and M. Soloveitchik. Drift and diffusion for a mechanical system. Probability Theory and Related Fields, 103 (3):349–379, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01195479.

Brydges:1995:DPM

[BS95b] D. C. Brydges and G. Slade. The diffusive phase of a model of self-interacting walks. Probability Theory and Related Fields, 103(3):285–315, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01195476.

Banuelos:1997:BMC

[BS97a] Rodrigo Bañuelos and Robert G. Smits. Brownian motion in cones. Probability Theory and Related Fields, 108(3):299–319, July 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050111.

Boldrighini:1997:ERM

[BS97b] C. Boldrighini and M. Soloveitchik. On the Einstein relation for a mechanical system. Probability Theory and Related Fields, 107 (4):493–515, April 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050095.

Benjamini:1998:EPP

[BS98a] Itai Benjamini and Oded Schramm. Exceptional planes of percolation. Probability Theory and Related Fields, 111(4):551–564, August 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050177.

Biane:1998:SCR

[BS98b] Philippe Biane and Roland Speicher. Stochastic calculus with respect to free Brownian motion and analysis on Wigner space. Probability Theory and Related Fields, 112(3):373–409, November 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050194.

Betz:2005:CLT

[BS05] Volker Betz and Herbert Spohn. A central limit theorem for Gibbs measures relative to Brownian motion. Probability Theory and Related Fields, 131(3):459–478, March 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0381-8.

Basdevant:2008:SCR

[BS08] Anne-Laure Basdevant and Arvind Singh. On the speed of a cookie random walk. Probability Theory and Related Fields, 141 (3–4):625–645, July 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0096-8.

Basu:2014:LER

[BS14a] Riddhipratim Basu and Allan Sly. Lipschitz embeddings of random sequences. Probability Theory and Related Fields, 159(3– 4):721–775, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0519-7.

Burdzy:2014:FBM

[BS14b] Krzysztof Burdzy and Michael Scheutzow. Forward Brownian motion. Probability Theory and Related Fields, 160(1–2): 95–126, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0524-x.

Blanca:2017:RCD

[BS17] Antonio Blanca and Alistair Sinclair. Random-cluster dynamics in Z^{2}. Probability Theory and Related Fields, 168(3–4):821–847, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-016-0725-1.

Berestycki:2019:CCI

[BS19] Nathanaël Berestycki and Bati Sengül. Cutoff for conjugacyinvariant random walks on the permutation group. Probability Theory and Related Fields, 173(3–4):1197–1241, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-018-0844-y;   http://link.springer.com/
content/pdf/10.1007/s00440-018-0844-y.pdf.

Bella:2022:NUP

[BS22] Peter Bella and Mathias Schäffner. Non-uniformly parabolic equations and applications to the random conductance model. Probability Theory and Related Fields, 182(1–2):353–397, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01081-1.

Bhamidi:2024:GMS

[BS24a] Shankar Bhamidi and Sanchayan Sen. Geometry of the minimal spanning tree in the heavy-tailed regime: new universality classes. Probability Theory and Related Fields, 188(3– 4):729–804, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-024-01259-3.

Bovier:2024:FFE

[BS24b] Anton Bovier and Adrien Schertzer. Fluctuations of the free energy in p-spin SK models on two scales. Probability Theory and Related Fields, 189(3–4):771–810, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01296-y.

Bendikov:2002:SDR

[BSC02] A. Bendikov and L. Saloff-Coste. Some dichotomy results for central Gaussian convolution semigroups on compact groups. Probability Theory and Related Fields, 124(4):561–573, December 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-002-0228-0.

Basdevant:2014:LVR

[BSS14] Anne-Laure Basdevant, Bruno Schapira, and Arvind Singh. Localization of a vertex reinforced random walk on Z with sublinear weight. Probability Theory and Related Fields, 159(1– 2):75–115, June 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0502-3.

Basu:2018:LER

[BSS18] Riddhipratim Basu, Vladas Sidoravicius, and Allan Sly. Lipschitz embeddings of random fields. Probability Theory and Related Fields, 172(3–4):1121–1179, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0826-5.

Beffara:2010:RPG

[BSV10] Vincent Beffara, Vladas Sidoravicius, and Maria Eulalia Vares. Randomized polynuclear growth with a columnar defect. Probability Theory and Related Fields, 147(3–4):565–581, July 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-009-0216-8.

Boedihardjo:2024:PMR

[BSV24] March Boedihardjo, Thomas Strohmer, and Roman Vershynin. Private measures, random walks, and synthetic data. Probability Theory and Related Fields, 189(1–2):569–611, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01279-z.

Bhamidi:2017:CLC

[BSW17] Shankar Bhamidi, Sanchayan Sen, and Xuan Wang. Continuum limit of critical inhomogeneous random graphs. Probability Theory and Related Fields, 169(1–2):565–641, October 2017. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0737-x.

Bally:1996:LES

[BT96] V. Bally and D. Talay. The law of the Euler scheme for stochastic differential equations. Probability Theory and Related Fields, 104(1):43–60, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01303802.

Baudoin:2003:PCW

[BT03] Fabrice Baudoin and Michèle Thieullen. Pinning class of the Wiener measure by a functional: related martingales and invariance properties. Probability Theory and Related Fields, 127 (1):1–36, September 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0280-4.

Betz:2011:SRP

[BU11] Volker Betz and Daniel Ueltschi. Spatial random permutations with small cycle weights. Probability Theory and Related Fields, 149(1–2):191–222, February 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-009-0248-0.

Buckdahn:1991:AGT

[Buc91a] Rainer Buckdahn. Anticipative Girsanov transformations. Probability Theory and Related Fields, 89(2):211–238, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01366907.

Buckdahn:1991:LSS

[Buc91b] Rainer Buckdahn. Linear Skorohod stochastic differential equations. Probability Theory and Related Fields, 90(2):223–240, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01192163.

Buckdahn:1992:SSD

[Buc92] Rainer Buckdahn. Skorohod stochastic differential equations of diffusion type. Probability Theory and Related Fields, 93 (3):297–323, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193054.

Buchholz:2021:PTC

[Buc21] Simon Buchholz. Phase transitions for a class of gradient fields. Probability Theory and Related Fields, 179(3–4):969–1022, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01021-5.

Budzinski:2018:HBP

[Bud18] Thomas Budzinski. The hyperbolic Brownian plane. Probability Theory and Related Fields, 171(1–2):503–541, June 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0785-x.

Bugiel:1985:NIM

[Bug85] Piotr Bugiel. A note on invariant measures for Markov maps of an interval. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 70(3):345–349, ???? 1985. CODEN ZWVGAA.

ISSN 0044-3719. URL http://link.springer.com/article/

10.1007/BF00534867. See correction and addendum [Ano87a].

Bugiel:1991:EPM

[Bug91] Piotr Bugiel. Ergodic properties of Markov maps in R^{d}. Probability Theory and Related Fields, 88(4):483–496, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192553.

Buja:1986:HST

[Buj86] Andreas Buja. On the Huber–Strassen theorem. Probability Theory and Related Fields, 73(1):149–152, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01845998.

Burdzy:1989:GPD

[Bur89] Krzysztof Burdzy. Geometric properties of 2-dimensional Brownian paths. Probability Theory and Related Fields, 81(4): 485–505, May 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00367299.

Burdzy:2015:MPM

[Bur15] Krzysztof Burdzy. Meteor process on Z^{d}. Probability Theory and Related Fields, 163(3–4):667–711, December 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0602-8.

Bercovici:1995:SCL

[BV95] H. Bercovici and D. Voiculescu. Superconvergence to the central limit and failure of the Cramér theorem for free random variables. Probability Theory and Related Fields, 103(2):215–222, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01204215.

Bloemendal:2013:LSR

[BV13] Alex Bloemendal and Bálint Virág. Limits of spiked random matrices I. Probability Theory and Related Fields, 156(3–4): 795–825, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0443-2.

Breuillard:2022:CPA

[BV22] Emmanuel Breuillard and Péter P. Varjú. Cut-off phenomenon for the ax+b Markov chain over a finite field. Probability Theory and Related Fields, 184(1–2):85–113, October 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01161-w.

Backhoff-Veraguas:2020:AAT

[BVBE20] Julio Backhoff-Veraguas, Daniel Bartl, and Manu Eder. All adapted topologies are equal. Probability Theory and Related Fields, 178(3–4):1125–1172, December 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-020-00993-8.

Bhamidi:2018:MCI

[BvdHS18] Shankar Bhamidi, Remco van der Hofstad, and Sanchayan Sen. The multiplicative coalescent, inhomogeneous continuum random trees, and new universality classes for critical random graphs. Probability Theory and Related Fields, 170 (1–2):387–474, February 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0760-6;
 http://link.springer.com/content/pdf/10.1007/s00440-
017-0760-6.pdf.

Berestycki:2024:HIO

[BvE24] Nathanaël Berestycki and Diederik van Engelenburg. Harnack inequality and one-endedness of UST on reversible random graphs. Probability Theory and Related Fields, 188(1–2): 487–548, February 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01239-z.

Berzin:1993:ADT

[BW93] Corinne Berzin and Mario Wschebor. Approximation du temps local des surfaces gaussiennes. (French) [Approximation of Gaussian surface local time]. Probability Theory and Related Fields, 96(1):1–32, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01195880.

Bryc:2005:CMM

[BW05] Włodzimierz Bryc and Jacek Wesolowski. Conditional moments of q-Meixner processes. Probability Theory and Related Fields, 131(3):415–441, March 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0379-2.

Baudoin:2017:SAW

[BW17a] Fabrice Baudoin and Jing Wang. Stochastic areas, winding numbers and Hopf fibrations. Probability Theory and Related Fields, 169(3–4):977–1005, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-016-0745-x.

Budhiraja:2017:MDP

[BW17b] Amarjit Budhiraja and Ruoyu Wu. Moderate deviation principles for weakly interacting particle systems. Probability Theory and Related Fields, 168(3–4):721–771, August 2017. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0723-3.

Beliaev:2018:VDN

[BW18] Dmitry Beliaev and Igor Wigman. Volume distribution of nodal domains of random band-limited functions. Probability Theory and Related Fields, 172(1–2):453–492, October 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0813-x;   http://link.springer.com/content/pdf/
10.1007/s00440-017-0813-x.pdf.

Bao:2022:EDC

[BW22] Zhigang Bao and Dong Wang. Eigenvector distribution in the critical regime of BBP transition. Probability Theory and Related Fields, 182(1–2):399–479, February 2022. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-021-01062-4.

Berestycki:2018:RHM

[BWW18] Nathanaël Berestycki, Christian Webb, and Mo Dick Wong. Random Hermitian matrices and Gaussian multiplicative chaos.

Probability Theory and Related Fields, 172(1–2):103–189, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-017-0806-9;   http://link.springer.com/
content/pdf/10.1007/s00440-017-0806-9.pdf.

Bai:1986:LBN

[BY86] Z. D. Bai and Y. Q. Yin. Limiting behavior of the norm of products of random matrices and two problems of Geman–Hwang. Probability Theory and Related Fields, 73(4):555–569, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00324852.

Bass:2005:FTH

[BY05] Richard F. Bass and Dahae You. A Fatou theorem for αharmonic functions in Lipschitz domains. Probability Theory and Related Fields, 133(3):391–408, November 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
005-0431-x.

Bourgade:2014:LCLa

[BYY14a] Paul Bourgade, Horng-Tzer Yau, and Jun Yin. Local circular law for random matrices. Probability Theory and Related Fields, 159(3–4):545–595, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-013-0514-z.

Bourgade:2014:LCLb

[BYY14b] Paul Bourgade, Horng-Tzer Yau, and Jun Yin. The local circular law II: the edge case. Probability Theory and Related Fields, 159(3–4):619–660, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-013-0516-x.

Bolthausen:2007:MAE

[BZ07] Erwin Bolthausen and Ofer Zeitouni. Multiscale analysis of exit distributions for random walks in random environments. Probability Theory and Related Fields, 138(3–4):581–645, July 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-006-0032-3.

Buzzi:2012:AMI

[BZ12] J. Buzzi and L. Zambotti. Approximate maximizers of intricacy functionals. Probability Theory and Related Fields, 153(3–4): 421–440, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0350-y.

Basak:2020:ORP

[BZ20] Anirban Basak and Ofer Zeitouni. Outliers of random perturbations of Toeplitz matrices with finite symbols. Probability Theory and Related Fields, 178(3–4):771–826, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00990-x.

Cameron:1987:SRS

[Cam87] Peter J. Cameron. On the structure of a random sum-free set. Probability Theory and Related Fields, 76(4):523–531, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00960072.

Capitaine:1995:OMF

[Cap95] Mireille Capitaine. Onsager–Machlup functional for some smooth norms on Wiener space. Probability Theory and Related Fields, 102(2):189–201, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01213388.

Carraro:1986:PPP

[Car86] L. Carraro. Problèmes de prédiction pour le processus de Wiener à deux paramètres. (French) [Prediction problems for the Wiener process with two parameters]. Probability Theory and Related Fields, 72(4):619–635, July 1986. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00344724.

Carverhill:1987:FTN

[Car87] Andrew Carverhill. Furstenberg’s theorem for nonlinear stochastic systems. Probability Theory and Related Fields, 74 (4):529–534, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00363514.

Caravenna:2005:LLT

[Car05] Francesco Caravenna. A local limit theorem for random walks conditioned to stay positive. Probability Theory and Related Fields, 133(4):508–530, December 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-005-0444-5.

Castell:1993:AES

[Cas93] Fabienne Castell. Asymptotic expansion of stochastic flows. Probability Theory and Related Fields, 96(2):225–239, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01192134.

Castell:2001:HRS

[Cas01] Fabienne Castell. Homogenization of random semilinear PDEs. Probability Theory and Related Fields, 121(4):492–524, December 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400100164.

Castillo:2012:SBM

[Cas12] Ismaël Castillo. A semiparametric Bernstein–von Mises theorem for Gaussian process priors. Probability Theory and Related Fields, 152(1–2):53–99, February 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-010-0316-5.

Catoni:1996:LTT

[Cat96] Olivier Catoni. The Legendre transform of two replicas of the Sherrington–Kirkpatrick spin glass model. Probability Theory and Related Fields, 105(3):369–392, ???? 1996. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192213.

Catoni:1998:ETM

[Cat98] Olivier Catoni. The energy transformation method for the Metropolis algorithm compared with simulated annealing. Probability Theory and Related Fields, 110(1):69–89, January 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050145.

Coelho:1994:ALL

[CC94] Zaqueu Coelho and Pierre Collet. Asymptotic limit law for the close approach of two trajectories in expanding maps of the circle. Probability Theory and Related Fields, 99(2):237–250, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01199024.

Capinski:1999:EGS

[CC99] Marek Capiński and Nigel J. Cutland. Existence of global stochastic flow and attractors for Navier–Stokes equations. Probability Theory and Related Fields, 115(1):121–151, August 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050238.

Clark:2005:RVI

[CC05] J. M. C. Clark and D. Crisan. On a robust version of the integral representation formula of nonlinear filtering. Probability Theory and Related Fields, 133(1):43–56, September 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0412-5.

Chapon:2021:QSS

[CC21] François Chapon and Reda Chhaibi. Quantum SL_{2}, infinite curvature and Pitman’s 2M − X theorem. Probability Theory and Related Fields, 179(3–4):835–888, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01002-8.

Campanino:1991:GFC

[CCC91] Massimo Campanino, J. T. Chayes, and L. Chayes. Gaussian fluctuations of connectivities in the subcritical regime of percolation. Probability Theory and Related Fields, 88(3):269–341, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01418864.

Chayes:1988:CPM

[CCD88] J. T. Chayes, L. Chayes, and R. Durrett. Connectivity properties of Mandelbrot’s percolation process. Probability Theory and Related Fields, 77(3):307–324, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00319291.

Csaki:2000:ISP

[CCFS00] Endre Csáki, Miklós Csörgő, Antónia Földes, and Zhan Shi. Increment sizes of the principal value of Brownian local time. Probability Theory and Related Fields, 117(4):515–531, August 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008733.

Chayes:1991:PTM

[CCGS91] J. T. Chayes, L. Chayes, E. Grannan, and G. Swindle. Phase transitions in Mandelbrot’s percolation process in three dimensions. Probability Theory and Related Fields, 90(3):291–300, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01193747.

Chernozhukov:2015:CAC

[CCK15] Victor Chernozhukov, Denis Chetverikov, and Kengo Kato. Comparison and anti-concentration bounds for maxima of Gaussian random vectors. Probability Theory and Related Fields, 162 (1–2):47–70, June 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0565-9.

Chazottes:2007:CIR

[CCKR07] J.-R. Chazottes, P. Collet, C. Külske, and F. Redig. Concentration inequalities for random fields via coupling. Probability Theory and Related Fields, 137(1–2):201–225, January 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0026-1.

Contat:2024:EAP

[CCL^{+}24] Alice Contat, Nicolas Curien, Perrine Lacroix, Etienne Lasalle, and Vincent Rivoirard. Eve, Adam and the preferential attachment tree. Probability Theory and Related Fields, 190(1–2): 321–336, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01253-1.

Chazottes:2016:SAQ

[CCM16] J.-R. Chazottes, P. Collet, and S. Méléard. Sharp asymptotics for the quasi-stationary distribution of birth-and-death processes. Probability Theory and Related Fields, 164(1–2): 285–332, February 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0612-6.

Cox:1990:LDI

[CD90] J. Theodore Cox and Richard Durrett. Large deviations for independent random walks. Probability Theory and Related Fields, 84(1):67–82, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01288559.

Chaumont:1999:PUP

[CD99] L. Chaumont and R. A. Doney. Pathwise uniqueness for perturbed versions of Brownian motion and reflected Brownian motion. Probability Theory and Related Fields, 113(4):519–534, March 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050216.

Comets:2001:OOD

[CD01] Francis Comets and Amir Dembo. Ordered overlaps in disordered mean-field models. Probability Theory and Related Fields, 121(1):1–29, September 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/PL00008794.

Carmona:2013:SFS

[CD13a] René Carmona and François Delarue. Singular FBSDEs and scalar conservation laws driven by diffusion processes. Probability Theory and Related Fields, 157(1–2):333–388, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-012-0459-7.

Chatterjee:2013:CLT

[CD13b] Sourav Chatterjee and Partha S. Dey. Central limit theorem for first-passage percolation time across thin cylinders. Probability Theory and Related Fields, 156(3–4):613–663, August 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0438-z.

Chatterjee:2020:SMC

[CD20] Sourav Chatterjee and Persi Diaconis. Speeding up Markov chains with deterministic jumps. Probability Theory and Related Fields, 178(3–4):1193–1214, December 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-020-01006-4. See correction [CD21].

Chatterjee:2021:CSM

[CD21] Sourav Chatterjee and Persi Diaconis. Correction to: Speeding up Markov chains with deterministic jumps. Probability Theory and Related Fields, 181(1-3):377–400, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01049-1. See [CD20].

Coquille:2014:GSN

[CDCIV14] Loren Coquille, Hugo Duminil-Copin, Dmitry Ioffe, and Yvan Velenik. On the Gibbs states of the noncritical Potts model on Z^{2}. Probability Theory and Related Fields, 158(1–2):477– 512, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0486-z.

Caravenna:2021:PTS

[CdH21] Francesco Caravenna and Frank den Hollander. Phase transitions for spatially extended pinning. Probability Theory and Related Fields, 181(1-3):329–375, November 2021. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-021-01068-y.

Collins:2018:SEU

[CDK18] Benoît Collins, Antoine Dahlqvist, and Todd Kemp. The spectral edge of unitary Brownian motion. Probability Theory and Related Fields, 170(1–2):49–93, February 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0753-x.

Carlier:2024:QSB

[CDM24] Guillaume Carlier, Alex Delalande, and Quentin Mérigot. Quantitative stability of barycenters in the Wasserstein space. Probability Theory and Related Fields, 188(3–4):1257–1286, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01241-5.

Chen:2017:FFE

[CDP17] Wei-Kuo Chen, Partha Dey, and Dmitry Panchenko. Fluctuations of the free energy in the mixed p-spin models with external field. Probability Theory and Related Fields, 168(1–2): 41–53, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Chen:1986:HNU

[CDR86] Jeesen Chen, Burgess Davis, and Herman Rubin. How nonuniform can a uniform sample be: a histogram approach. Probability Theory and Related Fields, 73(2):245–254, September 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00339939.

Cox:1991:CCP

[CDS91] J. T. Cox, R. Durrett, and R. Schinazi. The critical contact process seen from the right edge. Probability Theory and Related Fields, 87(3):325–332, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01312213.

Ceillier:2012:FSW

[Cei12] Gaël Ceillier. The filtration of the split-words process. Probability Theory and Related Fields, 153(1–2):269–292, June 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-011-0346-7.

Cekanavicius:1998:SNP

[Cek98] Vydas Cekanavicius. On signed normal-Poisson approximations. Probability Theory and Related Fields, 111(4):565–583, August 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050178.

Cerrai:1999:SPT

[Cer99] Sandra Cerrai. Smoothing properties of transition semigroups relative to SDEs with values in Banach spaces. Probability Theory and Related Fields, 113(1):85–114, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050203.

Cerrai:2003:SRD

[Cer03] Sandra Cerrai. Stochastic reaction-diffusion systems with multiplicative noise and non-Lipschitz reaction term. Probability Theory and Related Fields, 125(2):271–304, February 2003. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
002-0230-6.

Cerrai:2005:SNC

[Cer05] Sandra Cerrai. Stabilization by noise for a class of stochastic reaction-diffusion equations. Probability Theory and Related Fields, 133(2):190–214, October 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0421-4.

Cercle:2024:TPC

[Cer24] Baptiste Cerclé. Three-point correlation functions in the ∫↕_{∋} Toda theory I: reflection coefficients. Probability Theory and Related Fields, 188(1–2):89–158, February 2024. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-023-01219-3.

Cesi:2001:QFE

[Ces01] Filippo Cesi. Quasi-factorization of the entropy and logarithmic Sobolev inequalities for Gibbs random fields. Probability Theory and Related Fields, 120(4):569–584, August 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008792.

Cipolloni:2021:EUN

[CES21] Giorgio Cipolloni, László Erdős, and Dominik Schröder. Edge universality for non-Hermitian random matrices. Probability Theory and Related Fields, 179(1–2):1–28, February 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01003-7.

Cipolloni:2023:QUD

[CES23] Giorgio Cipolloni, László Erdős, and Dominik Schröder. Quenched universality for deformed Wigner matrices. Probability Theory and Related Fields, 185(3–4):1183–1218, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01156-7.

Cipolloni:2024:MCL

[CES24] Giorgio Cipolloni, László Erdős, and Dominik Schröder. Mesoscopic central limit theorem for non-Hermitian random matrices. Probability Theory and Related Fields, 188(3–4):1131– 1182, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01229-1.

Carmona:1994:DAR

[CF94a] René A. Carmona and Jean Pierre Fouque. A diffusion approximation result for two parameter processes. Probability Theory and Related Fields, 98(3):277–298, ???? 1994. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192255.

Crauel:1994:ARD

[CF94b] Hans Crauel and Franco Flandoli. Attractors for random dynamical systems. Probability Theory and Related Fields, 100 (3):365–393, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193705.

Cohen:2003:SRA

[CF03] Guy Cohen and Joseph M. Francos. Spectral representation and asymptotic properties of certain deterministic fields with innovation components. Probability Theory and Related Fields, 127 (2):277–304, October 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0287-x.

Chen:2008:OPE

[CF08] Zhen-Qing Chen and Masatoshi Fukushima. One-point extensions of Markov processes by darning. Probability Theory and Related Fields, 141(1–2):61–112, May 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0080-3.

Cerrai:2009:APC

[CF09] Sandra Cerrai and Mark Freidlin. Averaging principle for a class of stochastic reaction-diffusion equations. Probability Theory and Related Fields, 144(1–2):137–177, May 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0144-z.

Cerrai:2019:FFA

[CF19] Sandra Cerrai and Mark Freidlin. Fast flow asymptotics for stochastic incompressible viscous fluids in R^{2} and SPDEs on graphs. Probability Theory and Related Fields, 173(1–2):491– 535, February 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0839-8.

Cass:2024:WCE

[CF24] Thomas Cass and Emilio Ferrucci. On the Wiener chaos expansion of the signature of a Gaussian process. Probability Theory and Related Fields, 189(3–4):909–947, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01255-z.

Cox:1996:CID

[CFG96] J. Theodore Cox, Klaus Fleischmann, and Andreas Greven. Comparison of interacting diffusions and an application to their ergodic theory. Probability Theory and Related Fields, 105(4): 513–528, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01191911.

Croydon:2013:BRW

[CFK13] D. A. Croydon, A. Fribergh, and T. Kumagai. Biased random walk on critical Galton–Watson trees conditioned to survive. Probability Theory and Related Fields, 157(1–2):453–507, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00440-012-0462-z.pdf.

Chen:2008:PSM

[CFKZ08] Z.-Q. Chen, P. J. Fitzsimmons, K. Kuwae, and T.-S. Zhang. Perturbation of symmetric Markov processes. Probability Theory and Related Fields, 140(1–2):239–275, January 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0065-2.

Chevyrev:2020:SLD

[CFM20] Ilya Chevyrev, Peter K. Friz, and Ian Melbourne. Superdiffusive limits for deterministic fast-slow dynamical systems. Probability Theory and Related Fields, 178(3–4):735–770, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00988-5.

Csaki:1987:MWP

[CFR87] E. Csáki, A. Földes, and P. Révész. On the maximum of a Wiener process and its location. Probability Theory and Related Fields, 76(4):477–497, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00960069.

Csaki:1993:ASL

[CFR93] E. Csáki, A. Földes, and P. Révész. On almost sure local and global central limit theorems. Probability Theory and Related Fields, 97(3):321–337, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01195069.

Chen:2001:CES

[CFS01] Z.-Q. Chen, P. J. Fitzsimmons, and R. Song. Crossing estimates for symmetric Markov processes. Probability Theory and Related Fields, 120(1):68–84, May 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/PL00008776.

Carlier:2023:CGF

[CFV23] Guillaume Carlier, Gero Friesecke, and Daniela Vögler. Convex geometry of finite exchangeable laws and de Finetti style representation with universal correlated corrections. Probability Theory and Related Fields, 185(1–2):311–351, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01115-2.

Cox:1990:LTB

[CG90] J. T. Cox and A. Greven. On the long term behavior of some finite particle systems. Probability Theory and Related Fields, 85(2):195–237, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01277982.

Chen:1992:LIL

[CG92] Zhao-Guo Chen and Theo Gasser. A law of the iterated logarithm for kernel estimators of regression functions. Probability Theory and Related Fields, 93(3):285–296, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01193053.

Cabo:1994:LTF

[CG94] A. J. Cabo and P. Groeneboom. Limit theorems for functionals of convex hulls. Probability Theory and Related Fields, 100 (1):31–55, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01204952.

Chistyakov:2004:DSM

[CG04] G. P. Chistyakov and F. Götze. Distribution of the shape of Markovian random words. Probability Theory and Related Fields, 129(1):18–36, May 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0327-6.

Chazottes:2007:ASV

[CG07a] J.-R. Chazottes and S. Gouëzel. On almost-sure versions of classical limit theorems for dynamical systems. Probability Theory and Related Fields, 138(1–2):195–234, May 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0021-6.

Chistyakov:2007:ILF

[CG07b] G. P. Chistyakov and F. Götze. Independence of linear forms with random coefficients. Probability Theory and Related Fields, 137(1–2):1–24, January 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-006-0503-6.

Campanino:2009:LLT

[CG09] M. Campanino and M. Gianfelice. A local limit theorem for triple connections in subcritical Bernoulli percolation. Probability Theory and Related Fields, 143(3–4):353–378, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-007-0129-3.

Chistyakov:2013:AEC

[CG13] G. P. Chistyakov and F. Götze. Asymptotic expansions in the CLT in free probability. Probability Theory and Related Fields, 157(1–2):107–156, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0451-2.

Cattiaux:2007:WLS

[CGG07] P. Cattiaux, I. Gentil, and A. Guillin. Weak logarithmic Sobolev inequalities and entropic convergence. Probability Theory and Related Fields, 139(3–4):563–603, November 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0054-5.

Collin:2024:IDR

[CGH24] Orphée Collin, Giambattista Giacomin, and Yueyun Hu. Infinite disorder renormalization fixed point for the continuum random field Ising chain. Probability Theory and Related Fields, 190(3-4):881–939, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-024-01284-2.

Cattiaux:2008:PAG

[CGM08] P. Cattiaux, A. Guillin, and F. Malrieu. Probabilistic approach for granular media equations in the non-uniformly convex case. Probability Theory and Related Fields, 140(1–2):19–40, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-007-0056-3.

Cranston:2010:LDP

[CGM10] M. Cranston, D. Gauthier, and T. S. Mountford. On large deviations for the parabolic Anderson model. Probability Theory and Related Fields, 147(1–2):349–378, May 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-009-0249-z.pdf.

Camia:2014:IME

[CGN14] Federico Camia, Christophe Garban, and Charles M. Newman. The Ising magnetization exponent on Z^{2} is 1/15. Probability Theory and Related Fields, 160(1–2):175–187, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-013-0526-8.

Carinci:2016:GAE

[CGRS16] Gioia Carinci, Cristian Giardinà, Frank Redig, and Tomohiro Sasamoto. A generalized asymmetric exclusion process with U_{q}(∫↕_{∈}) stochastic duality. Probability Theory and Related Fields, 166(3–4):887–933, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
015-0674-0;   http://link.springer.com/content/pdf/
10.1007/s00440-015-0674-0.pdf.

Cox:1995:FIS

[CGS95] J. T. Cox, Andreas Greven, and Tokuzo Shiga. Finite and infinite systems of interacting diffusions. Probability Theory and Related Fields, 103(2):165–197, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01204213.

Chen:2023:FPB

[CGS23] Xinxin Chen, Christophe Garban, and Atul Shekhar. The fixed points of branching Brownian motion. Probability Theory and Related Fields, 185(3–4):839–884, April 2023. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-022-01183-4.

Cattiaux:2010:NTT

[CGW10] Patrick Cattiaux, Arnaud Guillin, and Li-Ming Wu. A note on Talagrand’s transportation inequality and logarithmic Sobolev inequality. Probability Theory and Related Fields, 148(1–2): 285–304, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0231-9.

Comets:2000:QAF

[CGZ00] Francis Comets, Nina Gantert, and Ofer Zeitouni. Quenched, annealed and functional large deviations for one-dimensional random walk in random environment. Probability Theory and Related Fields, 118(1):65–114, September 2000. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400000074.

Comets:2003:QAF

[CGZ03] Francis Comets, Nina Gantert, and Ofer Zeitouni. Quenched, annealed and functional large deviations for one-dimensional random walk in random environment. Probability Theory and Related Fields, 125(1):42–44, January 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-002-0234-2.

Csorgo:1988:RCT

[CH88a] M. Csörgő and L. Horváth. Rate of convergence of transport processes with an application to stochastic differential equations. Probability Theory and Related Fields, 78(3):379–387, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00334201.

Csorgo:1988:CLT

[CH88b] Miklós Csörgő and Lajos Horváth. Central limit theorems for L_{p}-norms of density estimators. Probability Theory and Related Fields, 80(2):269–291, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00356106.

Cambanis:1993:SPM

[CH93] Stamatis Cambanis and Christian Houdré. Stable processes: moving averages versus Fourier transforms. Probability Theory and Related Fields, 95(1):75–85, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01197338.

Carmona:2002:PFD

[CH02] Philippe Carmona and Yueyun Hu. On the partition function of a directed polymer in a Gaussian random environment. Probability Theory and Related Fields, 124(3):431–457, November 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400200213.

Curien:2013:STN

[CH13] Nicolas Curien and Bénédicte Haas. The stable trees are nested. Probability Theory and Related Fields, 157(3–4):847–883, December 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-012-0472-x.

Corwin:2016:KLE

[CH16] Ivan Corwin and Alan Hammond. KPZ line ensemble. Probability Theory and Related Fields, 166(1–2):67–185, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-015-0651-7.

Chen:2019:LDP

[CH19] Xinxin Chen and Hui He. On large deviation probabilities for empirical distribution of supercritical branching random walks with unbounded displacements. Probability Theory and Related Fields, 175(1–2):255–307, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0891-4.

Chan:1992:WSC

[Cha92] Terence Chan. The Wigner semi-circle law and eigenvalues of matrix-valued diffusions. Probability Theory and Related Fields, 93(2):249–272, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01195231.

Chang:1994:EFG

[Cha94] Chih-Chung Chang. Equilibrium fluctuations of gradient reversible particle systems. Probability Theory and Related Fields, 100(3):269–283, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193701.

Chatterjee:2007:SMC

[Cha07] Sourav Chatterjee. Stein’s method for concentration inequalities. Probability Theory and Related Fields, 138(1–2):305–321, May 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-006-0029-y.

Chatterjee:2009:FES

[Cha09] Sourav Chatterjee. Fluctuations of eigenvalues and second order Poincaré inequalities. Probability Theory and Related Fields, 143 (1–2):1–40, January 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0118-6.

Chapuy:2010:SUM

[Cha10a] Guillaume Chapuy. The structure of unicellular maps, and a connection between maps of positive genus and planar labelled trees. Probability Theory and Related Fields, 147(3–4):415–447, July 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-009-0211-0.

Chatterjee:2010:SGS

[Cha10b] Sourav Chatterjee. Spin glasses and Stein’s method. Probability Theory and Related Fields, 148(3–4):567–600, November 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00440-009-0240-8.pdf.

Chatterjee:2012:NAS

[Cha12] Sourav Chatterjee. A new approach to strong embeddings. Probability Theory and Related Fields, 152(1–2):231–264, February 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00440-010-0321-8.pdf.

Chatterjee:2016:PTT

[Cha16] Shirshendu Chatterjee. Phase transition for the threshold contact process, an approximation of heterogeneous random Boolean networks. Probability Theory and Related Fields, 165(3–4):985–1023, August 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s00440-015-
0656-2;   http://link.springer.com/article/10.1007/
s00440-015-0656-2.

Chapuy:2019:TRM

[Cha19a] Guillaume Chapuy. On tessellations of random maps and the t_{g}-recurrence. Probability Theory and Related Fields, 174(1– 2):477–500, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0865-6.

Chatterjee:2019:RTD

[Cha19b] Sourav Chatterjee. Rigidity of the three-dimensional hierarchical Coulomb gas. Probability Theory and Related Fields, 175 (3–4):1123–1176, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-019-00912-6.

Chak:2024:RPK

[Cha24] Martin Chak. Regularity preservation in Kolmogorov equations for non-Lipschitz coefficients under Lyapunov conditions. Probability Theory and Related Fields, 190(1–2):259–319, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01313-0.

Chen:1990:SAE

[Che90] Zhen Qing Chen. Stochastic analysis on extended sample space and a tightness results. Probability Theory and Related Fields, 86(4):517–549, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01198173.

Chen:1992:RDS

[Che92a] Zhen-Qing Chen. On reflected Dirichlet spaces. Probability Theory and Related Fields, 94(2):135–162, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192442.

Chen:1992:PJD

[Che92b] Zhen-Qing Chen. Pseudo Jordan domains and reflecting Brownian motions. Probability Theory and Related Fields, 94(2): 271–280, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192446.

Chen:1993:RDP

[Che93] Zhen-Qing Chen. On reflecting diffusion processes and Skorokhod decompositions. Probability Theory and Related Fields, 94(3):281–315, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199246.

Chernov:1995:LTM

[Che95] N. I. Chernov. Limit theorems and Markov approximations for chaotic dynamical systems. Probability Theory and Related Fields, 101(3):321–362, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01200500.

Chen:2000:LLS

[Che00] Xia Chen. On the limit laws of the second order for additive functionals of Harris recurrent Markov chains. Probability Theory and Related Fields, 116(1):89–123, January 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008724.

Chevyrev:2018:RWL

[Che18] Ilya Chevyrev. Random walks and Lévy processes as rough paths. Probability Theory and Related Fields, 170 (3–4):891–932, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0781-1;
 http://link.springer.com/content/pdf/10.1007/s00440-
017-0781-1.pdf.

Chi:2001:CSS

[Chi01] Zhiyi Chi. Construction of stationary self-similar generalized fields by random wavelet expansion. Probability Theory and Related Fields, 121(2):269–300, October 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/PL00008805.

Chichignoud:2012:MMA

[Chi12] M. Chichignoud. Minimax and minimax adaptive estimation in multiplicative regression: locally Bayesian approach. Probability Theory and Related Fields, 153(3–4):543–586, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-011-0354-7.

Chen:2018:ISH

[CHKN18] Le Chen, Yaozhong Hu, Kamran Kalbasi, and David Nualart. Intermittency for the stochastic heat equation driven by a rough time fractional Gaussian noise. Probability Theory and Related Fields, 171(1–2):431–457, June 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0783-z.

Chen:2018:ELM

[CHL18] Wei-Kuo Chen, Madeline Handschy, and Gilad Lerman. On the energy landscape of the mixed even p-spin model. Probability Theory and Related Fields, 171(1–2):53–95, June 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-017-0773-1.

Csorgo:1986:WPS

[CHM86] Sándor Csörgő, Lajos Horváth, and David M. Mason. What portion of the sample makes a partial sum asymptotically stable or normal? Probability Theory and Related Fields, 72(1):1–16, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00343893.

ChaudrudeRaynal:2022:SRB

[CHM22] Paul-Éric Chaudru de Raynal, Igor Honoré, and Stéphane Menozzi. Strong regularization by Brownian noise propagating through a weak Hörmander structure. Probability Theory and Related Fields, 184(1–2):1–83, October 2022. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-022-01150-z.

Coquet:2002:FCN

[CHMP02] François Coquet, Ying Hu, Jean Mémin, and Shige Peng. Filtration-consistent nonlinear expectations and related gexpectations. Probability Theory and Related Fields, 123(1): 1–27, May 2002. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/s004400100172.

Christoph:1986:WSR

[Chr86] Gerd Christoph. Weighted sums of random variables attracted to stable laws. Probability Theory and Related Fields, 71(3): 335–340, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01000209.

Cipriani:2018:SLO

[CHR18] Alessandra Cipriani, Rajat Subhra Hazra, and Wioletta M. Ruszel. Scaling limit of the odometer in divisible sandpiles. Probability Theory and Related Fields, 172(3–4): 829–868, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0821-x;
 http://link.springer.com/content/pdf/10.1007/s00440-
017-0821-x.pdf.

Chen:2022:SPC

[CHS22] Xinxing Chen, Yueyun Hu, and Zhan Shi. The sustainability probability for the critical Derrida–Retaux model. Probability Theory and Related Fields, 182(1–2):641–684, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01091-z.

Cambanis:1988:IWD

[CHW88] Stamatis Cambanis, Clyde D. Hardin Jr., and Aleksander Weron. Innovations and Wold decompositions of stable sequences. Probability Theory and Related Fields, 79(1):1–27, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/BF00319099.

Chow:1999:SAS

[CIK99] Pao-Liu Chow, Ildar A. Ibragimov, and Rafail Z. Khasminskii. Statistical approach to some ill-posed problems for linear partial differential equations. Probability Theory and Related Fields, 113(3):421–441, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400050212.

Campanino:2003:OZT

[CIvV03] Massimo Campanino, Dmitry Ioffe, and Y. van Velenik. Ornstein–Zernike theory for finite range Ising models above T_{c}. Probability Theory and Related Fields, 125(3):305–349, March 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-002-0229-z.

Chow:1994:SPD

[CJ94] Pao-Liu Chow and Jing-Lin Jiang. Stochastic partial differential equations in Hölder spaces. Probability Theory and Related Fields, 99(1):1–27, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199588.

Conus:2013:CCS

[CJKS13] Daniel Conus, Mathew Joseph, Davar Khoshnevisan, and Shang-Yuan Shiu. On the chaotic character of the stochastic heat equation, II. Probability Theory and Related Fields, 156(3– 4):483–533, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0434-3.

Camia:2021:EFB

[CJN21] Federico Camia, Jianping Jiang, and Charles M. Newman. The effect of free boundary conditions on the Ising model in high dimensions. Probability Theory and Related Fields, 181(1-3): 311–328, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01041-9.

Chiyonobu:1988:LDP

[CK88] T. Chiyonobu and S. Kusuoka. The large deviation principle for hypermixing processes. Probability Theory and Related Fields, 78(4):627–649, August 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00353880.

Cox:2000:REI

[CK00] J. Theodore Cox and Achim Klenke. Recurrence and ergodicity of interacting particle systems. Probability Theory and Related Fields, 116(2):239–255, February 2000. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008728.

Chen:2002:GFE

[CK02a] Zhen-Qing Chen and Panki Kim. Green function estimate for censored stable processes. Probability Theory and Related Fields, 124(4):595–610, December 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-002-0226-2.

Cox:2002:REI

[CK02b] J. Theodore Cox and Achim Klenke. Recurrence and ergodicity of interacting particle systems. Probability Theory and Related Fields, 122(4):613–615, April 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400200197.

Chen:2004:SMB

[CK04] Zhen-Qing Chen and Panki Kim. Stability of Martin boundary under non-local Feynman–Kac perturbations. Probability Theory and Related Fields, 128(4):525–564, April 2004. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0317-8.

Chen:2008:HKE

[CK08a] Zhen-Qing Chen and Takashi Kumagai. Heat kernel estimates for jump processes of mixed types on metric measure spaces. Probability Theory and Related Fields, 140(1–2):277–317, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-007-0070-5.

Csorgo:2008:RPQ

[CK08b] Miklós Csörgő and Rafal Kulik. Reduction principles for quantile and Bahadur–Kiefer processes of long-range dependent linear sequences. Probability Theory and Related Fields, 142(3–4): 339–366, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0107-9.

Conus:2012:EPF

[CK12] Daniel Conus and Davar Khoshnevisan. On the existence and position of the farthest peaks of a family of stochastic heat and wave equations. Probability Theory and Related Fields, 152(3– 4):681–701, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0333-4.

Cotar:2015:UGG

[CK15a] Codina Cotar and Christof Külske. Uniqueness of gradient Gibbs measures with disorder. Probability Theory and Related Fields, 162(3–4):587–635, August 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0580-x;   http://link.springer.com/content/pdf/
10.1007/s00440-014-0580-x.pdf.

Curien:2015:PRT

[CK15b] Nicolas Curien and Igor Kortchemski. Percolation on random triangulations and stable looptrees. Probability Theory and Related Fields, 163(1–2):303–337, October 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0593-5.

Chaika:2019:SCW

[CK19] Jon Chaika and Arjun Krishnan. Stationary coalescing walks on the lattice. Probability Theory and Related Fields, 175(3–4): 655–675, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0893-2.

Chen:2020:JMB

[CK20] Xiaohui Chen and Kengo Kato. Jackknife multiplier bootstrap: finite sample approximations to the U-process supremum with applications. Probability Theory and Related Fields, 176(3–4): 1097–1163, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00936-y.

Chen:2013:DAS

[CKK13] Zhen-Qing Chen, Panki Kim, and Takashi Kumagai. Discrete approximation of symmetric jump processes on metric measure spaces. Probability Theory and Related Fields, 155(3–4): 703–749, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0411-x.

Cranston:1993:RPB

[CKM93] Michael Cranston, Wilfrid S. Kendall, and Peter March. The radial part of Brownian motion II. Its life and times on the cut locus. Probability Theory and Related Fields, 96(3):353–368, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01292677.

Castillo:2014:TBW

[CKP14] Ismaël Castillo, Gérard Kerkyacharian, and Dominique Picard. Thomas Bayes’ walk on manifolds. Probability Theory and Related Fields, 158(3–4):665–710, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-013-0493-0.

Chen:2010:TSH

[CKS10] Zhen-Qing Chen, Panki Kim, and Renming Song. Two-sided heat kernel estimates for censored stable-like processes. Probability Theory and Related Fields, 146(3–4):361–399, March 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-008-0193-3.

Chan:2011:SSF

[CKS11] T. Chan, A. E. Kyprianou, and M. Savov. Smoothness of scale functions for spectrally negative Lévy processes. Probability Theory and Related Fields, 150(3–4):691–708, August 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0289-4.

Cranston:1990:NSE

[CL90] M. Cranston and Y. Le Jan. Noncoalescence for the Skorohod equation in a convex domain of R^{2}. Probability Theory and Related Fields, 87(2):241–252, ???? 1990. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01198431.

Csorgo:1991:PPK

[CL91] Miklós Csörgő and Zhengyan Lin. Path properties of kernel generated two-time parameter Gaussian processes. Probability Theory and Related Fields, 89(4):423–445, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01199787.

Cepa:1997:DPE

[CL97a] Emmanuel Cépa and Dominique Lépingle. Diffusing particles with electrostatic repulsion. Probability Theory and Related Fields, 107(4):429–449, April 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050092.

Chen:1997:MAI

[CL97b] Kani Chen and Shaw-Hwa Lo. On a mapping approach to investigating the bootstrap accuracy. Probability Theory and Related Fields, 107(2):197–217, February 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400050083.

Crisan:1997:NFM

[CL97c] Dan Crisan and Terry Lyons. Nonlinear filtering and measurevalued processes. Probability Theory and Related Fields, 109(2): 217–244, October 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050131.

Crisan:1999:PAS

[CL99] D. Crisan and T. Lyons. A particle approximation of the solution of the Kushner–Stratonovitch equation. Probability Theory and Related Fields, 115(4):549–578, November 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050249.

Chen:2004:LMD

[CL04] Xia Chen and Wenbo V. Li. Large and moderate deviations for intersection local times. Probability Theory and Related Fields, 128(2):213–254, February 2004. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0298-7.

Cerf:2007:IDD

[CL07] Raphaël Cerf and Sana Louhichi. The initial drift of a 2D droplet at zero temperature. Probability Theory and Related Fields, 137(3–4):379–428, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-006-0007-4.

Ceillier:2014:FTS

[CL14] Gaël Ceillier and Christophe Leuridan. Filtrations at the threshold of standardness. Probability Theory and Related Fields, 158 (3–4):785–808, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0496-x.

Curien:2016:HPB

[CL16] Nicolas Curien and Jean-François Le Gall. The hull process of the Brownian plane. Probability Theory and Related Fields, 166(1–2):187–231, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s00440-
015-0652-6.

Ceillier:2017:SCF

[CL17] Gaël Ceillier and Christophe Leuridan. Sufficient conditions for the filtration of a stationary processes to be standard. Probability Theory and Related Fields, 167(3–4):979–999, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-016-0696-2.

Che:2019:LSS

[CL19] Ziliang Che and Benjamin Landon. Local spectral statistics of the addition of random matrices. Probability Theory and Related Fields, 175(1–2):579–654, October 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00932-2.

Clarke:1986:NAF

[Cla86] Brenton R. Clarke. Nonsmooth analysis and Fréchet differentiability of M-functionals. Probability Theory and Related Fields, 73(2):197–209, September 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00339936.

Clark:1987:CMA

[Cla87] J. M. C. Clark. Convergent martingales of asymptotically minimal fluctuation. Probability Theory and Related Fields, 75(4): 531–543, August 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00320332.

Colubi:1999:GSL

[CLDDMG99] Ana Colubi, Miguel López-Díiaz, J. Santos Domínguez-Menchero, and M. Angeles Gil. A generalized strong law of large numbers. Probability Theory and Related Fields, 114 (3):401–417, June 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050229.

Cline:1986:CTP

[Cli86] Daren B. H. Cline. Convolution tails, product tails and domains of attraction. Probability Theory and Related Fields, 72 (4):529–557, July 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00344720.

Chinot:2020:RSL

[CLL20] Geoffrey Chinot, Guillaume Lecué, and Matthieu Lerasle. Robust statistical learning with Lipschitz and convex loss functions. Probability Theory and Related Fields, 176(3–4):897–940, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-019-00931-3.

Cheng:1994:NLP

[CLM94] Zheming Cheng, Joel L. Lebowitz, and Péter Major. On the number of lattice points between two enlarged and randomly shifted, copies of an oval. Probability Theory and Related Fields, 100(2):253–268, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199268.

Caputo:2012:PDD

[CLM^{+}12] Pietro Caputo, Hubert Lacoin, Fabio Martinelli, François Simenhaus, and Fabio Lucio Toninelli. Polymer dynamics in the depinned phase: metastability with logarithmic barriers. Probability Theory and Related Fields, 153(3–4):587–641, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-011-0355-6.

Chang:2001:EFA

[CLO01] Chih-Chung Chang, Claudio Landim, and Stefano Olla. Equilibrium fluctuations of asymmetric simple exclusion processes in dimension d ≥ 3. Probability Theory and Related Fields, 119 (3):381–409, March 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008764.

Coffman:2001:PRR

[CLSW01] E. G. Coffman, Jr., George S. Lueker, Joel Spencer, and Peter M. Winkler. Packing random rectangles. Probability Theory and Related Fields, 120(4):585–599, August 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008793.

Chen:2011:CEW

[CLW11] Xin Chen, Xue-Mei Li, and Bo Wu. A concrete estimate for the weak Poincaré inequality on loop space. Probability Theory and Related Fields, 151(3–4):559–590, December 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0308-5.

Cuesta:1988:SLL

[CM88] J. A. Cuesta and C. Matran. The strong law of large numbers for k-means and best possible nets of Banach valued random variables. Probability Theory and Related Fields, 78(4):523–534, August 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00353875.

Cranston:1991:ERB

[CM91] M. C. Cranston and T. S. Mountford. An extension of a result of Burdzy and Lawler. Probability Theory and Related Fields, 89(4):487–502, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199790.

Carmona:1995:SPA

[CM95] R. A. Carmona and S. A. Molchanov. Stationary parabolic Anderson model and intermittency. Probability Theory and Related Fields, 102(4):433–453, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01198845.

Csorgo:1996:EPS

[CM96] Sándor Csörgó and Jan Mielniczuk. The empirical process of a short-range dependent stationary sequence under Gaussian subordination. Probability Theory and Related Fields, 104(1):15–25, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01303800.

Chitashvili:1997:FTW

[CM97] R. Chitashvili and M. Mania. On functions transforming a Wiener process into a semimartingale. Probability Theory and Related Fields, 109(1):57–76, September 1997. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050125.

Cancrini:2001:DSS

[CM01] N. Cancrini and F. Martinelli. Diffusive scaling of the spectral gap for the dilute Ising lattice-gas dynamics below the percolation threshold. Probability Theory and Related Fields, 120(4): 497–534, August 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008790.

Cattiaux:2002:HNH

[CM02] Patrick Cattiaux and Laurent Mesnager. Hypoelliptic nonhomogeneous diffusions. Probability Theory and Related Fields, 123(4):453–483, August 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400100194.

Cranston:2007:QAT

[CM07] M. Cranston and S. Molchanov. Quenched to annealed transition in the parabolic Anderson problem. Probability Theory and Related Fields, 138(1–2):177–193, May 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0020-7.

Csiszar:2008:GML

[CM08] Imre Csiszár and Frantisek Matús. Generalized maximum likelihood estimates for exponential families. Probability Theory and Related Fields, 141(1–2):213–246, May 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0084-z.

Cerf:2011:IMN

[CM11a] R. Cerf and R. J. Messikh. The 2D-Ising model near criticality: a FK-percolation analysis. Probability Theory and Related Fields, 150(1–2):193–217, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0272-0.

Champagnat:2011:PES

[CM11b] Nicolas Champagnat and Sylvie Méléard. Polymorphic evolution sequence and evolutionary branching. Probability Theory and Related Fields, 151(1–2):45–94, October 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0292-9.

Cerf:2013:DNG

[CM13] Raphaël Cerf and Francesco Manzo. A d-dimensional nucleation and growth model. Probability Theory and Related Fields, 155(1–2):427–449, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-011-0402-3.

Croydon:2017:QLB

[CM17] David A. Croydon and Stephen Muirhead. Quenched localisation in the bouchaud trap model with slowly varying traps. Probability Theory and Related Fields, 168(1–2):269–315, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Crisan:2018:SPM

[CM18] Dan Crisan and Eamon McMurray. Smoothing properties of McKean–Vlasov SDEs. Probability Theory and Related Fields, 171(1–2):97–148, June 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0774-0;
 http://link.springer.com/content/pdf/10.1007/s00440-
017-0774-0.pdf.

Clozeau:2024:AQE

[CM24] Nicolas Clozeau and Francesco Mattesini. Annealed quantitative estimates for the quadratic 2D-discrete random matching problem. Probability Theory and Related Fields, 190(1–2): 485–541, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01254-0.

Chojnowska-Michalik:1995:EUI

[CMG95] A. Chojnowska-Michalik and B. Goldys. Existence, uniqueness and invariant measures for stochastic semilinear equations on Hilbert spaces. Probability Theory and Related Fields, 102 (3):331–356, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192465.

Chojnowska-Michalik:2002:SOU

[CMG02] Anna Chojnowska-Michalik and Beniamin Goldys. Symmetric Ornstein–Uhlenbeck semigroups and their generators. Probability Theory and Related Fields, 124(4):459–486, December 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200222.

Chaleyat-Maurel:1990:SVS

[CMM90] M. Chaleyat-Maurel and D. Michel. A Stroock–Varadhan support theorem in non-linear filtering theory. Probability Theory and Related Fields, 84(1):119–139, ???? 1990. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01288562.

Collet:2011:QSD

[CMMM11] Pierre Collet, Servet Martínez, Sylvie Méléard, and Jaime San Martín. Quasi-stationary distributions for structured birth and death processes with mutations. Probability Theory and Related Fields, 151(1–2):191–231, October 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-010-0297-4.

Chaleyat-Maurel:1992:OMF

[CMN92] Mireille Chaleyat-Maurel and David Nualart. The Onsager– Machlup functional for a class of anticipating processes. Probability Theory and Related Fields, 94(2):247–270, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192445.

Cancrini:2008:KCS

[CMRT08] N. Cancrini, F. Martinelli, C. Roberto, and C. Toninelli. Kinetically constrained spin models. Probability Theory and Related Fields, 140(3–4):459–504, March 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0072-3.

Cancrini:2015:MTK

[CMRT15] N. Cancrini, F. Martinelli, C. Roberto, and C. Toninelli. Mixing time of a kinetically constrained spin model on trees: power law scaling at criticality. Probability Theory and Related Fields, 161(1–2):247–266, February 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0548-x.

Collet:2000:ABB

[CMS00] Pierre Collet, Servet Martínez, and Jaime San Martín. Asymptotic behaviour of a Brownian motion on exterior domains. Probability Theory and Related Fields, 116(3):303–316, March 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050251.

Collet:2002:EID

[CMS02] P. Collet, S. Martinez, and B. Schmitt. Exponential inequalities for dynamical measures of expanding maps of the interval. Probability Theory and Related Fields, 123(3):301–322, July 2002.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200204.

Collet:2003:AHK

[CMS03] Pierre Collet, Servet Martínez, and Jaime San Martín. Asymptotic of the heat kernel in general Benedicks domains. Probability Theory and Related Fields, 125(3):350–364, March 2003. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
002-0241-3.

Cranston:2005:LEP

[CMS05] M. Cranston, T. S. Mountford, and T. Shiga. Lyapunov exponent for the parabolic Anderson model with Lévy noise. Probability Theory and Related Fields, 132(3):321–355, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0346-y.

Cardot:2007:CFL

[CMS07] Hervé Cardot, André Mas, and Pascal Sarda. CLT in functional linear regression models. Probability Theory and Related Fields, 138(3–4):325–361, July 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-006-0025-2. See

correction [CMS23].

Cacciapuoti:2015:BST

[CMS15] Claudio Cacciapuoti, Anna Maltsev, and Benjamin Schlein. Bounds for the Stieltjes transform and the density of states of Wigner matrices. Probability Theory and Related Fields, 163 (1–2):1–59, October 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0586-4.

Cardot:2023:CCF

[CMS23] Hervé Cardot, André Mas, and Pascal Sarda. Correction: CLT in functional linear regression models. Probability Theory and Related Fields, 187(1–2):519–522, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01215-7. See [CMS07].

Caputo:2015:PSA

[CMT15] Pietro Caputo, Fabio Martinelli, and Fabio Lucio Toninelli. On the probability of staying above a wall for the (2+1)-dimensional SOS model at low temperature. Probability Theory and Related Fields, 163(3–4):803–831, December 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-015-0658-0.

Cai:2015:OER

[CMW15] Tony Cai, Zongming Ma, and Yihong Wu. Optimal estimation and rank detection for sparse spiked covariance matrices. Probability Theory and Related Fields, 161(3–4):781–815, April 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-014-0562-z.

Carmona:1988:RNL

[CN88] Rene Carmona and David Nualart. Random non-linear wave equations: Smoothness of the solutions. Probability Theory and Related Fields, 79(4):469–508, November 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00318783.

Camia:2007:CPE

[CN07] Federico Camia and Charles M. Newman. Critical percolation exploration path and SLE_{6}: a proof of convergence. Probability Theory and Related Fields, 139(3–4):473–519, November 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00440-006-0049-7.pdf.

Comets:2016:LLG

[CN16] Francis Comets and Vu-Lan Nguyen. Localization in loggamma polymers with boundaries. Probability Theory and Related Fields, 166(1–2):429–461, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0662-4.

Chiarini:2020:ERG

[CN20] Alberto Chiarini and Maximilian Nitzschner. Entropic repulsion for the Gaussian free field conditioned on disconnection by levelsets. Probability Theory and Related Fields, 177(1–2):525–575, June 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-019-00957-7.

Coca:2018:ENI

[Coc18] Alberto J. Coca. Efficient nonparametric inference for discretely observed compound Poisson processes. Probability Theory and Related Fields, 170(1–2):475–523, February 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0761-5.

Collins:2005:PRP

[Col05] Benoît Collins. Product of random projections, Jacobi ensembles and universality problems arising from free probability. Probability Theory and Related Fields, 133(3):315–344, November 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-005-0428-5.

Comets:1989:LDE

[Com89] F. Comets. Large deviation estimates for a conditional probability distribution. applications to random interaction Gibbs measures. Probability Theory and Related Fields, 80(3):407–432, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01794432.

Conforti:2019:SOE

[Con19] Giovanni Conforti. A second order equation for Schrödinger bridges with applications to the hot gas experiment and entropic transportation cost. Probability Theory and Related Fields, 174 (1–2):1–47, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0856-7.

Conforti:2024:WSE

[Con24] Giovanni Conforti. Weak semiconvexity estimates for Schrödinger potentials and logarithmic Sobolev inequality for Schrödinger bridges. Probability Theory and Related Fields, 189(3–4):1045– 1071, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-024-01264-6.

Cook:2017:SAM

[Coo17] Nicholas A. Cook. On the singularity of adjacency matrices for random regular digraphs. Probability Theory and Related Fields, 167(1–2):143–200, February 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0679-8;   http://link.springer.com/
article/10.1007/s00440-015-0679-8.

Cooperman:2023:QSH

[Coo23] William Cooperman. Quantitative stochastic homogenization of the G equation. Probability Theory and Related Fields, 186 (1–2):493–520, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01175-4.

Cassandro:1993:ITG

[COP93] Marzio Cassandro, Enza Orlandi, and Errico Presutti. Interfaces and typical Gibbs configurations for one-dimensional Kac potentials. Probability Theory and Related Fields, 96(1):57–96, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01195883.

Coquio:2000:WTO

[Coq00] Agnès Coquio. Why are there only three quantum noises? Probability Theory and Related Fields, 118(3):349–364, November 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008746.

Costantini:1992:SOR

[Cos92] C. Costantini. The Skorohod oblique reflection problem in domains with corners and application to stochastic differential equations. Probability Theory and Related Fields, 91(1):43–70, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01194489.

Cox:2019:RSM

[COT19] Alexander M. G. Cox, Jan Oblój, and Nizar Touzi. The root solution to the multi-marginal embedding problem: an optimal stopping and time-reversal approach. Probability Theory and Related Fields, 173(1–2):211–259, February 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0833-1;   http://link.springer.com/content/pdf/
10.1007/s00440-018-0833-1.pdf.

Chrysaphinou:1988:LTN

[CP88] O. Chrysaphinou and S. Papastavridis. A limit theorem on the number of overlapping appearances of a pattern in a sequence of independent trials. Probability Theory and Related Fields, 79(1): 129–143, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319109.

Cheng:1993:BIC

[CP93] R. Cheng and M. Pourahmadi. Baxter’s inequality and convergence of finite predictors of multivariate stochastic processes. Probability Theory and Related Fields, 95(1):115–124, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01197341.

Coppersmith:2000:RWU

[CP00] Don Coppersmith and Igor Pak. Random walk on upper triangular matrices mixes rapidly. Probability Theory and Related Fields, 117(3):407–417, July 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050012.

Comets:2003:LLT

[CP03] Francis Comets and Serguei Popov. Limit law for transition probabilities and moderate deviations for Sinai’s random walk in random environment. Probability Theory and Related Fields, 126(4):571–609, August 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0273-3.

Caputo:2007:EDE

[CP07a] Pietro Caputo and Gustavo Posta. Entropy dissipation estimates in a zero-range dynamics. Probability Theory and Related Fields, 139(1–2):65–87, September 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-006-0039-9.

Cox:2007:SCS

[CP07b] J. Theodore Cox and Edwin A. Perkins. Survival and coexistence in stochastic spatial Lotka–Volterra models. Probability Theory and Related Fields, 139(1–2):89–142, September 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-006-0040-3.

Chatterjee:2010:PTB

[CP10] Sourav Chatterjee and Soumik Pal. A phase transition behavior for Brownian motions interacting through their ranks. Probability Theory and Related Fields, 147(1–2):123–159, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-009-0203-0.

Cator:2012:BFE

[CP12] Eric Cator and Leandro P. R. Pimentel. Busemann functions and equilibrium measures in last passage percolation models. Probability Theory and Related Fields, 154(1–2):89–125, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00440-011-0363-6.pdf.

Chen:2013:ACS

[CP13] Wei-Kuo Chen and Dmitry Panchenko. An approach to chaos in some mixed p-spin models. Probability Theory and Related Fields, 157(1–2):389–404, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0460-1.

Capitaine:2016:FES

[CP16] Mireille Capitaine and Sandrine Péché. Fluctuations at the edges of the spectrum of the full rank deformed GUE. Probability Theory and Related Fields, 165(1–2):117–161, June 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-015-0628-6.

Coffman:1995:PRI

[CPW95] E. G. Coffman, Jr., Bjorn Poonen, and Peter Winkler. Packing random intervals. Probability Theory and Related Fields, 102 (1):105–121, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01295224.

Carmona:1994:SEA

[CPY94] Ph. Carmona, F. Petit, and M. Yor. Some extensions of the arc sine law as partial consequences of the scaling property of Brownian motion. Probability Theory and Related Fields, 100(1):1–29, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01204951.

Coutin:2002:SAR

[CQ02] Laure Coutin and Zhongmin Qian. Stochastic analysis, rough path analysis and fractional Brownian motions. Probability Theory and Related Fields, 122(1):108–140, January 2002. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100158.

Caputo:2020:SDC

[CQ20] Pietro Caputo and Matteo Quattropani. Stationary distribution and cover time of sparse directed configuration models. Probability Theory and Related Fields, 178(3–4):1011–1066, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00995-6.

Calka:2023:LTF

[CQ23] Pierre Calka and Gauthier Quilan. Limit theory for the first layers of the random convex hull peeling in the unit ball. Probability Theory and Related Fields, 187(3–4):1037–1091, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01224-6.

Cao:2024:DFU

[CQ24] Shiping Cao and Hua Qiu. Dirichlet forms on unconstrained Sierpinski carpets. Probability Theory and Related Fields, 189 (1–2):613–657, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-024-01280-6.

Csorgo:1986:MDV

[CR86a] M. Csörgő and P. Révész. Mesure du voisinage and occupation density. Probability Theory and Related Fields, 73(2):211– 226, September 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00339937.

Csorgo:1986:NNE

[CR86b] M. Csörgo and P. Révész. A nearest neighbour-estimator for the score function. Probability Theory and Related Fields, 71 (2):293–305, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00332313.

Chauvin:1988:KES

[CR88] Brigitte Chauvin and Alain Rouault. KPP equation and supercritical branching Brownian motion in the subcritical speed area. Application to spatial trees. Probability Theory and Related Fields, 80(2):299–314, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00356108.

Chen:1990:MBP

[CR90] A. Y. Chen and E. Renshaw. Markov branching processes with instantaneous immigration. Probability Theory and Related Fields, 87(2):209–240, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01198430.

Chen:1993:EUC

[CR93] Anyue Chen and Eric Renshaw. Existence and uniqueness criteria for conservative uni-instantaneous denumerable Markov processes. Probability Theory and Related Fields, 94(4):427–456, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01192557.

Chen:1995:CPA

[CR95] L. H. Y. Chen and M. Roos. Compound Poisson approximation for unbounded functions on a group, with application to large deviations. Probability Theory and Related Fields, 103 (4):515–528, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01246337.

Campos:2014:ECB

[CR14] David Campos and Alejandro F. Ramírez. Ellipticity criteria for ballistic behavior of random walks in random environment. Probability Theory and Related Fields, 160(1–2):189–251, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-013-0527-7.

Cranston:1993:PAM

[Cra93] M. Cranston. A probabilistic approach to Martin boundaries for manifolds with ends. Probability Theory and Related Fields, 96(3):319–334, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01292675.

Cranston:2000:GPS

[Cra00] M. Cranston. On geometric properties of stochastic flows related to the Lyapunov spectrum. Probability Theory and Related Fields, 118(1):1–16, September 2000. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic).  URL http://link.springer.com/article/10.
1007/PL00008737;   http://link.springer.com/article/
10.1007/s004400000067.

Crane:2017:EGV

[Cra17] Harry Crane. Exchangeable graph-valued Feller processes. Probability Theory and Related Fields, 168(3–4):849–899, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-016-0726-0.

Croydon:2008:VGH

[Cro08] David A. Croydon. Volume growth and heat kernel estimates for the continuum random tree. Probability Theory and Related Fields, 140(1–2):207–238, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0063-4.

Croydon:2013:SMR

[Cro13] David A. Croydon. Slow movement of a random walk on the range of a random walk in the presence of an external field. Probability Theory and Related Fields, 157(3–4):515–534, December 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00440-012-0463-y.pdf.

Cabezas:2018:RDD

[CRS18] M. Cabezas, L. T. Rolla, and V. Sidoravicius. Recurrence and density decay for diffusion-limited annihilating systems. Probability Theory and Related Fields, 170(3–4):587–615, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-017-0763-3.

Cattiaux:1996:AGD

[CRZ96] P. Cattiaux, S. Roelly, and H. Zessin. Une approche gibbsienne des diffusions browniennes infini-dimensionnelles. (French) [Gibbsian approach of Brownian infinite-dimensional diffusian]. Probability Theory and Related Fields, 104(2):147–179, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01247836.

Cai:2013:ORC

[CRZ13] T. Tony Cai, Zhao Ren, and Harrison H. Zhou. Optimal rates of convergence for estimating Toeplitz covariance matrices. Probability Theory and Related Fields, 156(1–2):101–143, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-012-0422-7.

Chambers:1989:CLT

[CS89] Daniel Chambers and Eric Slud. Central limit theorems for nonlinear functionals of stationary Gaussian processes. Probability Theory and Related Fields, 80(3):323–346, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01794427.

Csorgo:1994:ASL

[CS94] Miklós Csörgő and Qi-Man Shao. On almost sure limit inferior for b-valued stochastic processes and applications. Probability Theory and Related Fields, 99(1):29–54, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01199589.

Chen:2001:NUB

[CS01] Louis H. Y. Chen and Qi-Man Shao. A non-uniform Berry– Esseen bound via Stein’s method. Probability Theory and Related Fields, 120(2):236–254, June 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/PL00008782.

Chen:2003:CGT

[CS03] Zhen-Qing Chen and Renming Song. Conditional gauge theorem for non-local Feynman–Kac transforms. Probability Theory and Related Fields, 125(1):45–72, January 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400200219.

Chassaing:2004:RPL

[CS04] Philippe Chassaing and Gilles Schaeffer. Random planar lattices and integrated superBrownian excursion. Probability Theory and Related Fields, 128(2):161–212, February 2004. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0297-8.

Carbone:2008:HQO

[CS08a] Raffaella Carbone and Emanuela Sasso. Hypercontractivity for a quantum Ornstein–Uhlenbeck semigroup. Probability Theory and Related Fields, 140(3–4):505–522, March 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0073-2.

Chen:2008:CBL

[CS08b] Lung-Chi Chen and Akira Sakai. Critical behavior and the limit distribution for long-range oriented percolation. I. Probability Theory and Related Fields, 142(1–2):151–188, September 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-007-0101-2.

Chen:2009:CBL

[CS09] Lung-Chi Chen and Akira Sakai. Critical behavior and the limit distribution for long-range oriented percolation. II: Spatial correlation. Probability Theory and Related Fields, 145(3–4): 435–458, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0174-6.

Crawford:2012:SRW

[CS12] Nicholas Crawford and Allan Sly. Simple random walk on long range percolation clusters i: heat kernel bounds. Probability Theory and Related Fields, 154(3–4):753–786, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-011-0383-2.

Chang:2016:PTL

[CS16] Yinshan Chang and Artëm Sapozhnikov. Phase transition in loop percolation. Probability Theory and Related Fields, 164(3– 4):979–1025, April 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0624-x.

Caraceni:2020:PMT

[CS20] Alessandra Caraceni and Alexandre Stauffer. Polynomial mixing time of edge flips on quadrangulations. Probability Theory and Related Fields, 176(1–2):35–76, February 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-019-
00913-5;   https://link.springer.com/content/pdf/10.
1007/s00440-019-00913-5.pdf.

Croydon:2023:CEV

[CS23] D. A. Croydon and D. Shiraishi. Correction to: Exact value of the resistance exponent for four dimensional random walk trace. Probability Theory and Related Fields, 185(1–2):699–704, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-022-01160-x. See [Shi12].

Coulhon:1993:MPC

[CSC93] Th. Coulhon and L. Saloff-Coste. Minorations pour les chaînes de Markov unidimensionnelles. (French) [Lower bounds for unidimensional Markov chains]. Probability Theory and Related Fields, 97(3):423–431, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01195074.

Choulli:1999:FBD

[CSK99] Tahir Choulli, Christoph Stricker, and Leszek Krawczyk. On Fefferman and Burkholder–Davis–Gundy inequalities for Emartingales. Probability Theory and Related Fields, 113(4): 571–597, March 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050218.

Cavagnari:2023:DPV

[CSS23] Giulia Cavagnari, Giuseppe Savaré, and Giacomo Enrico Sodini. Dissipative probability vector fields and generation of evolution semigroups in Wasserstein spaces. Probability Theory and Related Fields, 185(3–4):1087–1182, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01148-7.

Chhor:2024:BOA

[CST24] Julien Chhor, Suzanne Sigalla, and Alexandre B. Tsybakov. Benign overfitting and adaptive nonparametric regression. Probability Theory and Related Fields, 189(3–4):949–980, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01278-0.

Caravenna:2016:CDP

[CSZ16] Francesco Caravenna, Rongfeng Sun, and Nikos Zygouras. The continuum disordered pinning model. Probability Theory and Related Fields, 164(1–2):17–59, February 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0606-4;   http://link.springer.com/content/pdf/
10.1007/s00440-014-0606-4.pdf.

Chan:1986:NCI

[CT86] K. S. Chan and H. Tong. A note on certain integral equations associated with non-linear time series analysis. Probability Theory and Related Fields, 73(1):153–158, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01845999.

Cavalier:2002:SAI

[CT02] Laurent Cavalier and Alexandre Tsybakov. Sharp adaptation for inverse problems with random noise. Probability Theory and Related Fields, 123(3):323–354, July 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400100169.

Cerf:2011:LLD

[CT11a] Raphaël Cerf and Marie Théret. Lower large deviations for the maximal flow through a domain of R^{d} in first passage percolation. Probability Theory and Related Fields, 150(3–4):635–661, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-010-0287-6.

Chen:2011:GHK

[CT11b] Zhen-Qing Chen and Joshua Tokle. Global heat kernel estimates for fractional Laplacians in unbounded open sets. Probability Theory and Related Fields, 149(3–4):373–395, April 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
009-0256-0.

Chafai:2018:CEE

[CT18] Djalil Chafaï and Konstantin Tikhomirov. On the convergence of the extremal eigenvalues of empirical covariance matrices with dependence. Probability Theory and Related Fields, 170 (3–4):847–889, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0778-9.

Curien:2016:PSH

[Cur16] Nicolas Curien. Planar stochastic hyperbolic triangulations. Probability Theory and Related Fields, 165(3–4):509– 540, August 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0638-4;
http://link.springer.com/article/10.1007/s00440-015-
0638-4.

Chueshov:1998:LTB

[CV98] Igor D. Chueshov and Pierre A. Vuillermot. Long-time behavior of solutions to a class of stochastic parabolic equations with homogeneous white noise: Stratonovitch’s case. Probability Theory and Related Fields, 112(2):149–202, October 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050186.

Cheliotis:2010:SRE

[CV10] Dimitris Cheliotis and Bálint Virág. The spectrum of the random environment and localization of noise. Probability Theory and Related Fields, 148(1–2):141–158, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-009-0225-7.pdf.

Coquille:2012:FVV

[CV12] Loren Coquille and Yvan Velenik. A finite-volume version of Aizenman–Higuchi theorem for the 2D Ising model. Probability Theory and Related Fields, 153(1–2):25–44, June 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0339-6.

Champagnat:2016:ECQ

[CV16] Nicolas Champagnat and Denis Villemonais. Exponential convergence to quasi-stationary distribution and Q-process. Probability Theory and Related Fields, 164(1–2):243–283, February 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-014-0611-7.

Conforti:2018:CGE

[CV18] Giovanni Conforti and Max Von Renesse. Couplings, gradient estimates and logarithmic Sobolev inequalitiy for Langevin bridges. Probability Theory and Related Fields, 172(1–2):493– 524, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0814-9.

Chen:1993:OPP

[CW93] Mu-Fa Chen and Feng-Yu Wang. On order-preservation and positive correlations for multidimensional diffusion processes. Probability Theory and Related Fields, 95(3):421–428, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01192172.

Cerny:2017:AMD

[CW17] Jirí Cerný and Tobias Wassmer. Aging of the Metropolis dynamics on the random energy model. Probability Theory and Related Fields, 167(1–2):253–303, February 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0681-1;   http://link.springer.com/
article/10.1007/s00440-015-0681-1.

Chen:1995:EPS

[CWZ95] Z. Q. Chen, R. J. Williams, and Z. Zhao. On the existence of positive solutions for semilinear elliptic equations with Neumann boundary conditions. Probability Theory and Related Fields, 101(2):251–276, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01375828.

Carmona:1996:LDE

[CX96] René A. Carmona and Lin Xu. Large deviations and exponential decay for the magnetization in a Gaussian random field. Probability Theory and Related Fields, 106(2):233–247, October 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050063.

Calka:2014:VAR

[CY14] Pierre Calka and J. E. Yukich. Variance asymptotics for random polytopes in smooth convex bodies. Probability Theory and Related Fields, 158(1–2):435–463, February 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0484-1.

Calka:2015:VAS

[CY15] Pierre Calka and J. E. Yukich. Variance asymptotics and scaling limits for Gaussian polytopes. Probability Theory and Related Fields, 163(1–2):259–301, October 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-014-0592-6.

Chen:2016:HKA

[CZ16] Zhen-Qing Chen and Xicheng Zhang. Heat kernels and analyticity of non-symmetric jump diffusion semigroups. Probability Theory and Related Fields, 165(1–2):267–312, June 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
015-0631-y.

deAcosta:2000:GNC

[dA00] A. de Acosta. A general non-convex large deviation result with applications to stochastic equations. Probability Theory and Related Fields, 118(4):483–521, December 2000. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008752.

Dabrowski:1990:EPP

[Dab90] André Robert Dabrowski. Extremal point processes and intermediate quantile functions. Probability Theory and Related Fields, 85(3):365–386, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01193943.

Dahlhaus:1990:NHR

[Dah90] R. Dahlhaus. Nonparametric high resolution spectral estimation. Probability Theory and Related Fields, 85(2):147–180, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01277980.

Dalang:1988:IPG

[Dal88] Robert C. Dalang. On infinite perfect graphs and randomized stopping points on the plane. Probability Theory and Related Fields, 78(3):357–378, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00334200.

Darling:1988:EMV

[Dar88] R. W. R. Darling. Ergodicity of a measure-valued Markov chain induced by random transformations. Probability Theory and Related Fields, 77(2):211–229, February 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00334038.

Darling:1992:EPE

[Dar92] R. W. R. Darling. Exit probability estimates for martingales in geodesic balls, using curvature. Probability Theory and Related Fields, 93(2):137–152, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01195225.

Dauvergne:2023:LPI

[Dau23] Duncan Dauvergne. Last passage isometries for the directed landscape. Probability Theory and Related Fields, 186(1–2):391– 437, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-022-01173-6.

Davis:1990:RRW

[Dav90] Burgess Davis. Reinforced random walk. Probability Theory and Related Fields, 84(2):203–229, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01197845.

Davis:1999:BMR

[Dav99] Burgess Davis. Brownian motion and random walk perturbed at extrema. Probability Theory and Related Fields, 113(4):501–518, March 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050215.

Day:1988:LRD

[Day88] Martin V. Day. Localization results for densities associated with stable small-noise diffusions. Probability Theory and Related Fields, 77(3):457–470, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00319299.

deBouard:2002:ENS

[dBD02] A. de Bouard and A. Debussche. On the effect of a noise on the solutions of the focusing supercritical nonlinear Schrödinger equation. Probability Theory and Related Fields, 123(1):76–96, May 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400100183.

delBarrio:2013:RCP

[dBM13] Eustasio del Barrio and Carlos Matrán. Rates of convergence for partial mass problems. Probability Theory and Related Fields, 155(3–4):521–542, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-011-0406-z.

Duminil-Copin:2021:HK

[DCG21] Hugo Duminil-Copin and Geoffrey R. Grimmett. Harry Kesten (1931–2019). Probability Theory and Related Fields, 181(1-3):1, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01077-x.

Duminil-Copin:2024:SMT

[DCH24] Hugo Duminil-Copin and Ivailo Hartarsky. Sharp metastability transition for two-dimensional bootstrap percolation with symmetric isotropic threshold rules. Probability Theory and Related Fields, 190(1–2):445–483, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-024-01310-
3.

Duminil-Copin:2019:DRC

[DCL19] Hugo Duminil-Copin and Marcin Lis. On the double random current nesting field. Probability Theory and Related Fields, 175(3–4):937–955, December 2019. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-019-
00899-0;   https://link.springer.com/content/pdf/10.
1007/s00440-019-00899-0.pdf.

Duminil-Copin:2016:PTP

[DCM16] Hugo Duminil-Copin and Ioan Manolescu. The phase transitions of the planar random-cluster and Potts models with q ≥ 1 are sharp. Probability Theory and Related Fields, 164 (3–4):865–892, April 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0621-0.

Duminil-Copin:2021:PRC

[DCMT21] Hugo Duminil-Copin, Ioan Manolescu, and Vincent Tassion. Planar random-cluster model: fractal properties of the critical phase. Probability Theory and Related Fields, 181(1-3): 401–449, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01060-6.

Duminil-Copin:2019:EDC

[DCRT19] Hugo Duminil-Copin, Aran Raoufi, and Vincent Tassion. Exponential decay of connection probabilities for subcritical Voronoi percolation in R^{d}. Probability Theory and Related Fields, 173(1– 2):479–490, February 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0838-9.

Duminil-Copin:2018:BCP

[DCTT18] H. Duminil-Copin, V. Tassion, and A. Teixeira. The boxcrossing property for critical two-dimensional oriented percolation. Probability Theory and Related Fields, 171(3–4):685–708, August 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-017-0790-0.

Duminil-Copin:2018:HOC

[DCvEH18] Hugo Duminil-Copin, Aernout C. D. van Enter, and Tim Hulshof. Higher order corrections for anisotropic bootstrap percolation. Probability Theory and Related Fields, 172 (1–2):191–243, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0808-7;
 http://link.springer.com/content/pdf/10.1007/s00440-
017-0808-7.pdf.

Deniel:1988:CPS

[DD88] Y. Déniel and Y. Derriennic. Sur la convergence presque sure, au sens de Cesaro d’ordre α, 0 < α < 1, de variables aléatoires indepéndantes et identiquement distribuées. (French) [On the almost-sure convergence, in the sense of Cesaro, of order α, 0 < α < 1, of independent and identically-distributed random variables]. Probability Theory and Related Fields, 79(4): 629–636, November 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00318787.

Delmotte:2005:EDS

[DD05] T. Delmotte and J.-D. Deuschel. On estimating the derivatives of symmetric diffusions in stationary random environment, with applications to Δφ interface model. Probability Theory and Related Fields, 133(3):358–390, November 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
005-0430-y.

Delarue:2016:RPS

[DD16] François Delarue and Roland Diel. Rough paths and 1D SDE with a time dependent distributional drift: application to polymers. Probability Theory and Related Fields, 165(1–2):1–63, June 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-015-0626-8.

Duhalde:2017:EPM

[DD17] Xan Duhalde and Thomas Duquesne. Exact packing measure of the range of ψ-super Brownian motions. Probability Theory and Related Fields, 167(1–2):201–252, February 2017. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0680-2;   http://link.springer.com/
article/10.1007/s00440-015-0680-2.

Dahlqvist:2019:PAM

[DDD19] Antoine Dahlqvist, Joscha Diehl, and Bruce K. Driver. The parabolic Anderson model on Riemann surfaces. Probability Theory and Related Fields, 174(1–2):369–444, June 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0857-6;   http://link.springer.com/content/pdf/
10.1007/s00440-018-0857-6.pdf.

DaPrato:2002:SPI

[DDG02] Giuseppe Da Prato, Arnaud Debussche, and Beniamin Goldys. Some properties of invariant measures of non symmetric dissipative stochastic systems. Probability Theory and Related Fields, 123(3):355–380, July 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400100188.

Dereudre:2012:EGP

[DDG12] David Dereudre, Remy Drouilhet, and Hans-Otto Georgii. Existence of Gibbsian point processes with geometry-dependent interactions. Probability Theory and Related Fields, 153(3–4): 643–670, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0356-5.

Ding:1990:ERR

[DDL90] Wan-Ding Ding, Richard Durrett, and Thomas M. Liggett. Ergodicity of reversible reaction diffusion processes. Probability Theory and Related Fields, 85(1):13–26, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01377624.

Dembo:2019:CLC

[DDMP19] Amir Dembo, Jian Ding, Jason Miller, and Yuval Peres. Cutoff for lamplighter chains on tori: dimension interpolation and phase transition. Probability Theory and Related Fields, 173(1– 2):605–650, February 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0883-4.

Dehling:1986:BLI

[DDP86] Herold Dehling, Manfred Denker, and Walter Philipp. A bounded law of the iterated logarithm for Hilbert space valued martingales and its application to U-statistics. Probability Theory and Related Fields, 72(1):111–131, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00343899.

Dupuis:1992:LDM

[DE92] Paul Dupuis and Richard S. Ellis. Large deviations for Markov processes with discontinuous statistics, II: random walks. Probability Theory and Related Fields, 91(2):153–194, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01291423.

DeBlassie:1987:ETC

[DeB87a] R. Dante DeBlassie. Exit times from cones in R^{n} of Brownian motion. Probability Theory and Related Fields, 74(1):1–29, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.

1007/BF01845637. See remark [DeB88].

DeBlassie:1987:LCB

[DeB87b] R. Dante DeBlassie. The lifetime of conditioned Brownian motion in certain Lipschitz domains. Probability Theory and Related Fields, 75(1):55–65, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00320080.

DeBlassie:1988:RET

[DeB88] R. Dante DeBlassie. Remark on “Exit times from cones in R^{n} of Brownian motion. Probability Theory and Related Fields, 79(1): 95–97, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319106. See [DeB87a].

DeBlassie:2004:HOP

[DeB04] R. Dante DeBlassie. Higher order PDEs and symmetric stable processes. Probability Theory and Related Fields, 129(4): 495–536, August 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0347-x. See correction

[DeB05].

DeBlassie:2005:CHO

[DeB05] R. Dante DeBlassie. Correction to “Higher Order PDEs and Symmetric Stable Processes,” Probability Theory and Related Fields 129, 495–536 (2004). Probability Theory and Related Fields, 133(1):141–143, September 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-

005-0443-6. See [DeB04].

DeBlassie:2012:BMQ

[DeB12] Dante DeBlassie. Brownian motion in a quasi-cone. Probability Theory and Related Fields, 154(1–2):127–148, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-011-0364-5.

Dedecker:1998:CLT

[Ded98] Jérôme Dedecker. A central limit theorem for stationary random fields. Probability Theory and Related Fields, 110(3):397–426, March 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050153.

Dedecker:2010:ECL

[Ded10] J. Dedecker. An empirical central limit theorem for intermittent maps. Probability Theory and Related Fields, 148 (1–2):177–195, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-009-0227-5. See er-

ratum [Ded13].

Dedecker:2013:EEC

[Ded13] J. Dedecker. Erratum to: “An empirical central limit theorem for intermittent maps”. Probability Theory and Related Fields, 155(1–2):487–491, February 2013. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s00440-
011-0393-0.pdf. See [Ded10].

Defosseux:2016:ALA

[Def16] Manon Defosseux. Affine Lie algebras and conditioned spacetime Brownian motions in affine Weyl chambers. Probability Theory and Related Fields, 165(3–4):649–665, August 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0642-8;   http://link.springer.com/
article/10.1007/s00440-015-0642-8.

Deheuvels:1986:SLO

[Deh86] Paul Deheuvels. Strong laws for the k-th order statistic when k ≤ c log_{2} n. Probability Theory and Related Fields, 72(1): 133–154, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00343900.

Deheuvels:1988:SAR

[Deh88] Paul Deheuvels. Strong approximations of k-th records and k-th record times by Wiener processes. Probability Theory and Related Fields, 77(2):195–209, February 1988. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00334037.

Delmas:1999:SPR

[Del99] Jean-François Delmas. Some properties of the range of super-Brownian motion. Probability Theory and Related Fields, 114 (4):505–547, July 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050233.

Denis:2000:CDS

[Den00] Laurent Denis. A criterion of density for solutions of Poissondriven SDEs. Probability Theory and Related Fields, 118(3): 406–426, November 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008748.

Dermoune:2003:PCP

[Der03] Azzouz Dermoune. Propagation and conditional propagation of chaos for pressureless gas equations. Probability Theory and Related Fields, 126(4):459–476, August 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0274-2.

Dereudre:2022:FCB

[Der22] David Dereudre. Fully-connected bond percolation on Z^{d}. Probability Theory and Related Fields, 183(1–2):547–579, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01088-8.

Deuschel:1987:IDD

[Deu87] J. D. Deuschel. Infinite-dimensional diffusion processes as Gibbs measures on C[0, 1]^{Z^{d}}. Probability Theory and Related Fields, 76(3):325–340, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01297489.

Devroye:1988:KER

[Dev88] Luc Devroye. The kernel estimate is relatively stable. Probability Theory and Related Fields, 77(4):521–536, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00959615.

Deya:2016:MRH

[Dey16] Aurélien Deya. On a modelled rough heat equation. Probability Theory and Related Fields, 166(1–2):1–65, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-015-0650-8.

Dawson:1991:CBH

[DF91] Donald A. Dawson and Klaus Fleischmann. Critical branching in a highly fluctuating random medium. Probability Theory and Related Fields, 90(2):241–274, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01192164.

Delmas:2001:HSC

[DF01] Jean-François Delmas and Klaus Fleischmann. On the hot spots of a catalytic super-Brownian motion. Probability Theory and Related Fields, 121(3):389–421, November 2001. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100156.

Dias:2016:UVN

[DF16] J. C. A. Dias and S. Friedli. Uniqueness vs. non-uniqueness for complete connections with modified majority rules. Probability Theory and Related Fields, 164(3–4):893–929, April 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
015-0622-z.

Dubedat:2020:LMS

[DF20] Julien Dubédat and Hugo Falconet. Liouville metric of starscale invariant fields: tails and Weyl scaling. Probability Theory and Related Fields, 176(1–2):293–352, February 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00919-z.

Dwivedi:2019:POT

[DFGGR19] Raaz Dwivedi, Ohad N. Feldheim, Ori Gurel-Gurevich, and Aaditya Ramdas. The power of online thinning in reducing discrepancy. Probability Theory and Related Fields, 174(1–2):103–131, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-018-0860-y.

Dolera:2024:SPC

[DFM24] Emanuele Dolera, Stefano Favaro, and Edoardo Mainini. Strong posterior contraction rates via Wasserstein dynamics. Probability Theory and Related Fields, 189(1–2):659–720, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01260-w.

DeMasi:2015:SSE

[DFP15] Anna De Masi, Pablo A. Ferrari, and Errico Presutti. Symmetric simple exclusion process with free boundaries. Probability Theory and Related Fields, 161(1–2):155–193, February 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-014-0546-z.

Dubedat:2020:WLM

[DFS20] Julien Dubédat, Hugo Falconet, and Xin Sun. Weak LQG metrics and Liouville first passage percolation. Probability Theory and Related Fields, 178(1–2):369–436, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00979-6.

Ding:2020:GRW

[DFSX20] Jian Ding, Ryoki Fukushima, Rongfeng Sun, and Changji Xu. Geometry of the random walk range conditioned on survival among Bernoulli obstacles. Probability Theory and Related Fields, 177(1–2):91–145, June 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-019-00943-z.

Dekking:1988:SPP

[DG88] F. M. Dekking and G. R. Grimmett. Superbranching processes and projections of random Cantor sets. Probability Theory and Related Fields, 78(3):335–355, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00334199.

Dawson:1993:MTS

[DG93a] Donald Dawson and Andreas Greven. Multiple time scale analysis of interacting diffusions. Probability Theory and Related Fields, 95(4):467–508, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01196730.

Dawson:1993:HMI

[DG93b] Donald A. Dawson and Andreas Greven. Hierarchical models of interacting diffusions: Multiple time scale phenomena, phase transition and pattern of cluster-formation. Probability Theory and Related Fields, 96(4):435–473, ???? 1993. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01200205.

Doney:1993:JDL

[DG93c] R. A. Doney and P. E. Greenwood. On the joint distribution of ladder variables of random walk. Probability Theory and Related Fields, 94(4):457–472, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192558.

Dawson:1994:MLD

[DG94] D. A. Dawson and J. Gärtner. Multilevel large deviations and interacting diffusions. Probability Theory and Related Fields, 98 (4):423–487, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192835.

Denker:1998:CLT

[DG98] Manfred Denker and Mikhail Gordin. The central limit theorem for random perturbations of rotations. Probability Theory and Related Fields, 111(1):1–16, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050160.

Driver:2010:SIH

[DG10] Bruce K. Driver and Maria Gordina. Square integrable holomorphic functions on infinite-dimensional Heisenberg type groups. Probability Theory and Related Fields, 147(3–4):481–528, July 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00440-009-0213-y.pdf.

Denker:2014:LTM

[DG14] Manfred Denker and Mikhail Gordin. Limit theorems for von Mises statistics of a measure preserving transformation. Probability Theory and Related Fields, 160(1–2):1–45, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-013-0522-z.

Dembo:2021:DIT

[DG21a] Amir Dembo and Reza Gheissari. Diffusions interacting through a random matrix: universality via stochastic Taylor expansion. Probability Theory and Related Fields, 180(3–4):1057– 1097, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01027-7.

Dussaule:2021:SPM

[DG21b] Matthieu Dussaule and Ilya Gekhtman. Stability phenomena for Martin boundaries of relatively hyperbolic groups. Probability Theory and Related Fields, 179(1–2):201–259, February 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01000-w.

Deuschel:2022:QLC

[DG22] Jean-Dominique Deuschel and Xiaoqin Guo. Quenched local central limit theorem for random walks in a time-dependent balanced random environment. Probability Theory and Related Fields, 182(1–2):111–156, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01097-
7.

Dembin:2024:SMS

[DG24] Barbara Dembin and Christophe Garban. Superconcentration for minimal surfaces in first passage percolation and disordered Ising ferromagnets. Probability Theory and Related Fields, 190(3-4):675–702, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01252-2.

Deuschel:2000:LDC

[DGI00] Jean-Dominique Deuschel, Giambattista Giacomin, and Dmitry Ioffe. Large deviations and concentration properties for Δφ interface models. Probability Theory and Related Fields, 117 (1):49–111, May 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050266.

Durr:1987:SDN

[DGL87] Detlef Dürr, Sheldon Goldstein, and Joel L. Lebowitz. Selfdiffusion in a non-uniform one dimensional system of point particles with collisions. Probability Theory and Related Fields, 75 (2):279–290, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00354038.

Duerinckx:2020:RPS

[DGO20] Mitia Duerinckx, Antoine Gloria, and Felix Otto. Robustness of the pathwise structure of fluctuations in stochastic homogenization. Probability Theory and Related Fields, 178(1–2): 531–566, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-020-00983-w.

Dembo:2007:VML

[DGPS07] Amir Dembo, Nina Gantert, Yuval Peres, and Zhan Shi. Valleys and the maximum local time for random walk in random environment. Probability Theory and Related Fields, 137(3–4): 443–473, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0005-6.

Dembo:2002:LDR

[DGPZ02] Amir Dembo, Nina Gantert, Yuval Peres, and Ofer Zeitouni. Large deviations for random walks on Galton–Watson trees: averaging and uncertainty. Probability Theory and Related Fields, 122(2):241–288, February 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400100162.

Dunlap:2020:FSK

[DGRZ20] Alexander Dunlap, Yu Gu, Lenya Ryzhik, and Ofer Zeitouni. Fluctuations of the solutions to the KPZ equation in dimensions three and higher. Probability Theory and Related Fields, 176(3– 4):1217–1258, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00938-w.

Ding:2021:DEL

[DGS21] Jian Ding, Ewain Gwynne, and Avelio Sepúlveda. The distance exponent for Liouville first passage percolation is positive. Probability Theory and Related Fields, 181(4):1035–1051, December 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01093-x.

Deya:2012:NLR

[DGT12] A. Deya, M. Gubinelli, and S. Tindel. Non-linear rough heat equations. Probability Theory and Related Fields, 153(1–2): 97–147, June 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0341-z.

Deuschel:2005:SLE

[DGZ05] Jean-Dominique Deuschel, Giambattista Giacomin, and Lorenzo Zambotti. Scaling limits of equilibrium wetting models in (1 + 1)-dimension. Probability Theory and Related Fields, 132 (4):471–500, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0401-8.

Dekking:1991:LDM

[DH91] F. M. Dekking and B. Host. Limit distributions for minimal displacement of branching random walks. Probability Theory and Related Fields, 90(3):403–426, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01193752.

Dobrushin:1996:FSL

[DH96] R. Dobrushin and O. Hryniv. Fluctuations of shapes of large areas under paths of random walks. Probability Theory and Related Fields, 105(4):423–458, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01191908.

DellaMaestra:2022:NEI

[DH22] Laetitia Della Maestra and Marc Hoffmann. Nonparametric estimation for interacting particle systems: McKean–Vlasov models. Probability Theory and Related Fields, 182(1–2):551–613, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01044-6.

Dolgopyat:2023:BET

[DH23] Dmitry Dolgopyat and Yeor Hafouta. A Berry–Esseen theorem and Edgeworth expansions for uniformly elliptic inhomogeneous Markov chains. Probability Theory and Related Fields, 186(1– 2):439–476, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01177-2.

Delbaen:2011:BSS

[DHB11] Freddy Delbaen, Ying Hu, and Xiaobo Bao. Backward SDEs with superquadratic growth. Probability Theory and Related Fields, 150(1–2):145–192, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0271-1.

denHollander:1991:SLC

[dHHKR91] F. den Hollander, G. Hooghiemstra, M. Keane, and J. Resing. Strong law and central limit theorem for a process between maxima and sums. Probability Theory and Related Fields, 90 (1):37–55, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01321133.

Damron:2023:TET

[DHHL23] Michael Damron, Jack Hanson, David Harper, and Wai-Kit Lam. Transitions for exceptional times in dynamical firstpassage percolation. Probability Theory and Related Fields, 185(3–4):1039–1085, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-022-01178-1.

Driver:2022:BMF

[DHK22] Bruce K. Driver, Brian Hall, and Todd Kemp. The Brown measure of the free multiplicative Brownian motion. Probability Theory and Related Fields, 184(1–2):209–273, October 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01142-z.

denHollander:2003:DGT

[dHNOS03] F. den Hollander, F. R. Nardi, E. Olivieri, and E. Scoppola. Droplet growth for three-dimensional Kawasaki dynamics. Probability Theory and Related Fields, 125(2):153–194, February 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-002-0233-3.

deHaan:1986:SMS

[dHP86] L. de Haan and J. Pickands III. Stationary min-stable stochastic processes. Probability Theory and Related Fields, 72(4):477–492, July 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00344716.

Dario:2023:RFR

[DHP23] Paul Dario, Matan Harel, and Ron Peled. Random-field random surfaces. Probability Theory and Related Fields, 186(1– 2):91–158, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01179-0.

Dabrock:2021:EMS

[DHR21] Nils Dabrock, Martina Hofmanová, and Matthias Röger. Existence of martingale solutions and large-time behavior for a stochastic mean curvature flow of graphs. Probability Theory and Related Fields, 179(1–2):407–449, February 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01012-6.

Damron:2015:SVF

[DHS15] Michael Damron, Jack Hanson, and Philippe Sosoe. Sublinear variance in first-passage percolation for general distributions. Probability Theory and Related Fields, 163(1–2):223–258, October 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-014-0591-7.

Dembo:2019:RWA

[DHZ19] Amir Dembo, Ruojun Huang, and Tianyi Zheng. Random walks among time increasing conductances: heat kernel estimates. Probability Theory and Related Fields, 175(1–2):397–445, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-018-0894-1.

Ding:2013:EDE

[Din13] Jian Ding. Exponential and double exponential tails for maximum of two-dimensional discrete Gaussian free field. Probability Theory and Related Fields, 157(1–2):285–299, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-012-0457-9.

Dittrich:1988:SMC

[Dit88] P. Dittrich. A stochastic model of a chemical reaction with diffusion. Probability Theory and Related Fields, 79(1):115– 128, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319108.

Dittrich:1990:TWL

[Dit90] Peter Dittrich. Travelling waves and long-time behaviour of the weakly asymmetric exclusion process. Probability Theory and Related Fields, 86(4):443–455, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01198168.

Dalalyan:2014:SIC

[DIT14] Arnak Dalalyan, Yuri Ingster, and Alexandre B. Tsybakov. Statistical inference in compound functional models. Probability Theory and Related Fields, 158(3–4):513–532, April 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0487-y.

Divol:2022:MEM

[Div22] Vincent Divol. Measure estimation on manifolds: an optimal transport approach. Probability Theory and Related Fields, 183 (1–2):581–647, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01118-z.

deJong:1987:CLT

[dJ87] Peter de Jong. A central limit theorem for generalized quadratic forms. Probability Theory and Related Fields, 75(2):261–277, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00354037.

Donoho:1994:MRB

[DJ94] David L. Donoho and Iain M. Johnstone. Minimax risk over l_{p}-balls for l_{p}-error. Probability Theory and Related Fields, 99 (2):277–303, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199026.

Delyon:2000:MIN

[DJ00] Bernard Delyon and Anatoli Juditsky. On minimax identification of nonparametric autoregressive models. Probability Theory and Related Fields, 116(1):21–39, January 2000. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008721.

Durand:2012:MAL

[DJ12] Arnaud Durand and Stéphane Jaffard. Multifractal analysis of Lévy fields. Probability Theory and Related Fields, 153(1– 2):45–96, June 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0340-0.

DelMoral:2001:MCM

[DJP01] Pierre Del Moral, Jean Jacod, and Philip Protter. The Monte– Carlo method for filtering with discrete-time observations. Probability Theory and Related Fields, 120(3):346–368, July 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
PL00008786.

Dumitriu:2013:FLT

[DJPP13] Ioana Dumitriu, Tobias Johnson, Soumik Pal, and Elliot Paquette. Functional limit theorems for random regular graphs.

Probability Theory and Related Fields, 156(3–4):921–975, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-012-0447-y.

Dupuis:1987:ABC

[DK87] Paul Dupuis and Harold J. Kushner. Asymptotic behavior of constrained stochastic approximations via the theory of large deviations. Probability Theory and Related Fields, 75(2):223–244, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00354035.

Dynkin:1995:MSS

[DK95] E. B. Dynkin and S. E. Kuznetsov. Markov snakes and superprocesses. Probability Theory and Related Fields, 103(4):433–473, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01246334.

Dasgupta:1999:CDM

[DK99a] A. Dasgupta and G. Kallianpur. Chaos decomposition of multiple fractional integrals and applications. Probability Theory and Related Fields, 115(4):527–548, November 1999. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050248.

Dasgupta:1999:MFI

[DK99b] A. Dasgupta and G. Kallianpur. Multiple fractional integrals. Probability Theory and Related Fields, 115(4):505–525, November 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050247.

Dynkin:2004:MBE

[DK04] E. B. Dynkin and S. E. Kuznetsov. -measures for branching exit Markov systems and their applications to differential equations. Probability Theory and Related Fields, 130(1):135–150, September 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0333-8.

Dahmer:2017:TEL

[DK17] Iulia Dahmer and Götz Kersting. The total external length of the evolving Kingman coalescent. Probability Theory and Related Fields, 167(3–4):1165–1214, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-016-0703-7.

Durrett:2002:OER

[DKL02] Rick Durrett, Harry Kesten, and Vlada Limic. Once edgereinforced random walk on a tree. Probability Theory and Related Fields, 122(4):567–592, April 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400100179.

Dalang:2009:HPS

[DKN09] Robert C. Dalang, Davar Khoshnevisan, and Eulalia Nualart. Hitting probabilities for systems of non-linear stochastic heat equations with multiplicative noise. Probability Theory and Related Fields, 144(3–4):371–427, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-008-0150-1.

Dobler:2022:MFJ

[DKP22] Christian Döbler, Mikołaj Kasprzak, and Giovanni Peccati. The multivariate functional de Jong CLT. Probability Theory and Related Fields, 184(1–2):367–399, October 2022. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-022-01114-3.

Dynkin:1994:BMV

[DKS94] E. B. Dynkin, S. E. Kuznetsov, and A. V. Skorokhod. Branching measure-valued processes. Probability Theory and Related Fields, 99(1):55–96, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199590.

Dinh:2023:RWS

[DKW23] Tien-Cuong Dinh, Lucas Kaufmann, and Hao Wu. Random walks on SL_{2}(C): spectral gap and limit theorems. Probability Theory and Related Fields, 186(3–4):877–955, August 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01191-y.

Dhersin:1997:WTS

[DL97] Jean-Stéphane Dhersin and Jean-François Le Gall. Wiener’s test for super-Brownian motion and the Brownian snake. Probability Theory and Related Fields, 108(1):103–129, May 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050103.

Derriennic:2001:CLT

[DL01] Yves Derriennic and Michael Lin. The central limit theorem for Markov chains with normal transition operators, started at a point. Probability Theory and Related Fields, 119(4):508–528, April 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/PL00008769.

Dawson:2003:CIS

[DL03a] Donald A. Dawson and Zenghu Li. Construction of immigration superprocesses with dependent spatial motion from onedimensional excursions. Probability Theory and Related Fields, 127(1):37–61, September 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0278-y.

Derriennic:2003:CLT

[DL03b] Yves Derriennic and Michael Lin. The central limit theorem for Markov chains started at a point. Probability Theory and Related Fields, 125(1):73–76, January 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400200215.

Duquesne:2005:PFA

[DL05] Thomas Duquesne and Jean-François Le Gall. Probabilistic and fractal aspects of Lévy trees. Probability Theory and Related Fields, 131(4):553–603, April 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0385-4.

Dumaz:2020:LCA

[DL20] Laure Dumaz and Cyril Labbé. Localization of the continuous Anderson Hamiltonian in 1-D. Probability Theory and Related Fields, 176(1–2):353–419, February 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00920-6.

Ding:2023:SAT

[DL23] Jian Ding and Haoyu Liu. Shotgun assembly threshold for lattice labeling model. Probability Theory and Related Fields, 187(1–2):423–442, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01199-4.

DelMoral:2003:CPM

[DLM03] P. Del Moral, M. Ledoux, and L. Miclo. On contraction properties of Markov kernels. Probability Theory and Related Fields, 126(3):395–420, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0270-6.

Doukhan:1987:PIF

[DLP87] P. Doukhan, J. R. Leon, and F. Portal. Principes d’invariance faible pour la mesure empirique d’une suite de variables aléatoires mélangeante. (French) [Weak invariance principles for empirical measurement of a sequence of random mixing variables]. Probability Theory and Related Fields, 76(1):51– 70, September 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00390275.

Ding:2010:TVC

[DLP10] Jian Ding, Eyal Lubetzky, and Yuval Peres. Total variation cutoff in birth-and-death chains. Probability Theory and Related Fields, 146(1–2):61–85, January 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0185-3.

delaRue:1995:MMS

[dlR95] Thierry de la Rue. Mouvement moyen et système dynamique gaussien. (French) [Mean motion and Gaussian dynamic system]. Probability Theory and Related Fields, 102(1):45–56, ????

1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01295220.

Debussche:2024:GWP

[DLTV24] Arnaud Debussche, Ruoyuan Liu, Nikolay Tzvetkov, and Nicola Visciglia. Global well-posedness of the 2D nonlinear Schrödinger equation with multiplicative spatial white noise on the full space. Probability Theory and Related Fields, 189(3–4):1161– 1218, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-024-01288-y.

Donati-Martin:1991:TFT

[DM91a] C. Donati-Martin. Transformation de Fourier et temps d’occupation browniens. (French) [Fourier transformation and Brownian occupation time]. Probability Theory and Related Fields, 88(2):137–166, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01212557.

Durrett:1991:CCT

[DM91b] R. Durrett and A. M. Møller. Complete convergence theorem for a competition model. Probability Theory and Related Fields, 88(1):121–136, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193585.

Dalang:1997:PIB

[DM97] Robert C. Dalang and T. Mountford. Points of increase of the Brownian sheet. Probability Theory and Related Fields, 108(1):1–27, May 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050099.

Donati-Martin:2003:SIR

[DM03] C. Donati-Martin. Stochastic integration with respect to q Brownian motion. Probability Theory and Related Fields, 125 (1):77–95, January 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-002-0224-4.

Doney:2005:PTR

[DM05] R. A. Doney and R. A. Maller. Passage times of random walks and Lévy processes across power law boundaries. Probability Theory and Related Fields, 133(1):57–70, September 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0414-3.

Doumerc:2009:EPA

[DM09] Yan Doumerc and John Moriarty. Exit problems associated with affine reflection groups. Probability Theory and Related Fields, 145(3–4):351–383, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0171-9.

Santis:2016:PSI

[DM16a] Emilio De Santis and Andrea Maffei. Perfect simulation for the infinite random cluster model, Ising and Potts models at low or high temperature. Probability Theory and Related Fields, 164(1– 2):109–131, February 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0608-2.

Donoho:2016:HDR

[DM16b] David Donoho and Andrea Montanari. High dimensional robust M-estimation: asymptotic variance via approximate message passing. Probability Theory and Related Fields, 166 (3–4):935–969, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-015-0675-z;
 http://link.springer.com/content/pdf/10.1007/s00440-
015-0675-z.pdf.

Dembo:2017:PGP

[DM17] Amir Dembo and Sumit Mukherjee. Persistence of Gaussian processes: non-summable correlations. Probability Theory and Related Fields, 169(3–4):1007–1039, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0746-9.

DePonti:2022:ETD

[DM22a] Nicolò De Ponti and Andrea Mondino. Entropy-transport distances between unbalanced metric measure spaces. Probability Theory and Related Fields, 184(1–2):159–208, October 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01159-4.

Dubail:2022:AAR

[DM22b] Bastien Dubail and Laurent Massoulié. Accelerating abelian random walks with hyperbolic dynamics. Probability Theory and Related Fields, 184(3–4):939–968, December 2022. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-022-01128-x.

Dewan:2023:UBO

[DM23] Vivek Dewan and Stephen Muirhead. Upper bounds on the one-arm exponent for dependent percolation models. Probability Theory and Related Fields, 185(1–2):41–88, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01176-3.

Diaz:2020:GFB

[DMB20] Mario Diaz, James A. Mingo, and Serban T. Belinschi. On the global fluctuations of block Gaussian matrices. Probability Theory and Related Fields, 176(1–2):599–648, February 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00925-1.

Donati-Martin:1993:WND

[DMP93] C. Donati-Martin and E. Pardoux. White noise driven SPDEs with reflection. Probability Theory and Related Fields, 95 (1):1–24, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01197335.

Daskalakis:2011:ETI

[DMR11] Constantinos Daskalakis, Elchanan Mossel, and Sébastien Roch. Evolutionary trees and the Ising model on the Bethe lattice: a proof of Steel’s conjecture. Probability Theory and Related Fields, 149(1–2):149–189, February 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-009-0246-2.pdf.

Donati-Martin:2004:LDS

[DMRYZ04] C. Donati-Martin, A. Rouault, M. Yor, and M. Zani. Large deviations for squares of Bessel and Ornstein–Uhlenbeck processes. Probability Theory and Related Fields, 129(2):261–289, June 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-004-0338-y.

Denis:2005:EUN

[DMS05] Laurent Denis, Anis Matoussi, and Lucretiu Stoica. L^{p} estimates for the uniform norm of solutions of quasilinear SPDE’s. Probability Theory and Related Fields, 133(4):437–463, December 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-005-0436-5.

Ding:2021:ERG

[DMX21] Jian Ding, Zongming Ma, and Jiaming Xu. Efficient random graph matching via degree profiles. Probability Theory and Related Fields, 179(1–2):29–115, February 2021. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-020-00997-4.

Duc:1990:AMC

[DNS90] Nguyen Minh Duc, D. Nualart, and M. Sanz. Application of Malliavin calculus to a class of stochastic differential equations. Probability Theory and Related Fields, 84(4):549–571, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01198319.

Deuschel:2018:QIP

[DNS18] Jean-Dominique Deuschel, Tuan Anh Nguyen, and Martin Slowik. Quenched invariance principles for the random conductance model on a random graph with degenerate ergodic weights. Probability Theory and Related Fields, 170(1–2):363– 386, February 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0759-z.

De:2022:QCI

[DNS22] Anindya De, Shivam Nadimpalli, and Rocco A. Servedio. Quantitative correlation inequalities via extremal power series. Probability Theory and Related Fields, 183(1–2):649–675, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01120-5.

Doumerc:2005:EPA

[DO05] Yan Doumerc and Neil O’Connell. Exit problems associated with finite reflection groups. Probability Theory and Related Fields, 132(4):501–538, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0402-7.

Dobric:1988:LLN

[Dob88] V. Dobrić. The law of large numbers in locally convex spaces. Probability Theory and Related Fields, 78(3):403–417, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00334203.

Dokuchaev:2004:EDB

[Dok04] Nikolai Dokuchaev. Estimates for distances between first exit times via parabolic equations in unbounded cylinders. Probability Theory and Related Fields, 129(2):290–314, June 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0341-3.

Dommers:2017:MIM

[Dom17] Sander Dommers. Metastability of the Ising model on random regular graphs at zero temperature. Probability Theory and Related Fields, 167(1–2):305–324, February 2017. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0682-0;   http://link.springer.com/
article/10.1007/s00440-015-0682-0.

Doney:1989:ABF

[Don89] R. A. Doney. On the asymptotic behaviour of first passage times for transient random walk. Probability Theory and Related Fields, 81(2):239–246, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00319553.

Donoho:1994:AMR

[Don94] David L. Donoho. Asymptotic minimax risk for sup-norm loss: Solution via optimal recovery. Probability Theory and Related Fields, 99(2):145–170, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199020.

Doney:1995:SCL

[Don95] R. A. Doney. Spitzer’s condition and ladder variables in random walks. Probability Theory and Related Fields, 101(4):577–580, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01202785.

Doney:1997:OSL

[Don97] R. A. Doney. One-sided local large deviation and renewal theorems in the case of infinite mean. Probability Theory and Related Fields, 107(4):451–465, April 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050093.

Doney:2012:LBF

[Don12] R. A. Doney. Local behaviour of first passage probabilities. Probability Theory and Related Fields, 152(3–4):559–588, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-010-0330-7.

Downarowicz:1988:MSN

[Dow88] Tomasz Downarowicz. A minimal 0–1 subshift with noncompact set of ergodic measures. Probability Theory and Related Fields, 79(1):29–35, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00319100.

Deuschel:1996:SOL

[DP96] Jean-Dominique Deuschel and Agoston Pisztora. Surface order large deviations for high-density percolation. Probability Theory and Related Fields, 104(4):467–482, ???? 1996. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198162.

Deck:1998:CSP

[DP98] T. Deck and J. Potthoff. On a class of stochastic partial differential equations related to turbulent transport. Probability Theory and Related Fields, 111(1):101–122, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050163.

Durrett:1999:RCP

[DP99] Richard Durrett and Edwin A. Perkins. Rescaled contact processes converge to super-Brownian motion in two or more dimensions. Probability Theory and Related Fields, 114(3):309–399, June 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050228.

Dedecker:2005:NDC

[DP05] Jérôme Dedecker and Clémentine Prieur. New dependence coefficients. examples and applications to statistics. Probability Theory and Related Fields, 132(2):203–236, June 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0394-3.

Daskalakis:2015:SCS

[DP15] Constantinos Daskalakis and Christos Papadimitriou. Sparse covers for sums of indicators. Probability Theory and Related Fields, 162(3–4):679–705, August 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-014-0582-8.

Drewitz:2022:CCF

[DPR22] Alexander Drewitz, Alexis Prévost, and Pierre-François Rodriguez. Cluster capacity functionals and isomorphism theorems for Gaussian free fields. Probability Theory and Related Fields, 183(1–2):255–313, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-021-01090-0.

Deheuvels:1986:SRW

[DR86] Paul Deheuvels and Pál Révész. Simple random walk on the line in random environment. Probability Theory and Related Fields, 72(2):215–230, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00699104.

Durrett:1992:ABB

[DR92] R. T. Durrett and L. C. G. Rogers. Asymptotic behavior of Brownian polymers. Probability Theory and Related Fields, 92 (3):337–349, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01300560.

Dupuis:1999:CDSa

[DR99a] Paul Dupuis and Kavita Ramanan. Convex duality and the Skorokhod problem. I. Probability Theory and Related Fields, 115(2):153–195, September 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050269.

Dupuis:1999:CDSb

[DR99b] Paul Dupuis and Kavita Ramanan. Convex duality and the Skorokhod problem. II. Probability Theory and Related Fields, 115(2):197–236, September 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050270.

DaPrato:2002:SDS

[DR02] Giuseppe Da Prato and Michael Röckner. Singular dissipative stochastic equations in Hilbert spaces. Probability Theory and Related Fields, 124(2):261–303, October 2002. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200214.

Dalalyan:2006:ASE

[DR06] Arnak Dalalyan and Markus Reiß. Asymptotic statistical equivalence for scalar ergodic diffusions. Probability Theory and Related Fields, 134(2):248–282, February 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0416-1.

Dalalyan:2007:ASE

[DR07] Arnak Dalalyan and Markus Reiß. Asymptotic statistical equivalence for ergodic diffusions: the multidimensional case. Probability Theory and Related Fields, 137(1–2):25–47, January 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-006-0502-7.

DaPrato:2009:SDS

[DR09] Giuseppe Da Prato and Michael Röckner. Singular dissipative stochastic equations in Hilbert spaces. Probability Theory and Related Fields, 143(3–4):659–664, March 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-008-0179-1.pdf.

Drewitz:2011:BCR

[DR11] A. Drewitz and A. F. Ramírez. Ballisticity conditions for random walk in random environment. Probability Theory and Related Fields, 150(1–2):61–75, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-010-0268-9.

Doney:2013:ABF

[DR13] R. A. Doney and V. Rivero. Asymptotic behaviour of first passage time distributions for Lévy processes. Probability Theory and Related Fields, 157(1–2):1–45, October 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-

012-0448-x. See erratum [DR16].

Debussche:2014:EDN

[DR14a] Arnaud Debussche and Marco Romito. Existence of densities for the 3D Navier–Stokes equations driven by Gaussian noise. Probability Theory and Related Fields, 158(3–4):575–596, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-013-0490-3.

Dolera:2014:PMC

[DR14b] Emanuele Dolera and Eugenio Regazzini. Proof of a McKean conjecture on the rate of convergence of Boltzmann-equation solutions. Probability Theory and Related Fields, 160(1–2): 315–389, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0530-z.

Doney:2016:EAB

[DR16] R. A. Doney and V. Rivero. Erratum to: Asymptotic behaviour of first passage time distributions for Lévy processes. Probability Theory and Related Fields, 164(3–4):1079–1083, April 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-015-0660-6;   http://link.springer.com/
content/pdf/10.1007/s00440-015-0660-6.pdf. See [DR13].

Deuschel:2024:RKT

[DR24] Jean-Dominique Deuschel and Pierre-François Rodriguez. A Ray–Knight theorem for ∇ϕ interface models and scaling limits. Probability Theory and Related Fields, 189(1–2):447–499, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01275-3.

Driscoll:2013:SDA

[Dri13] Patrick Driscoll. Smoothness of densities for area-like processes of fractional Brownian motion. Probability Theory and Related Fields, 155(1–2):1–34, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-011-0389-9.

DaiPra:2002:GVP

[DRZ02] Paolo Dai Pra, Sylvie Roelly, and Hans Zessin. A Gibbs variational principle in space–time for infinite-dimensional diffusions. Probability Theory and Related Fields, 122(2):289–315, February 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400100170.

Deheuvels:1987:ECR

[DS87] Paul Deheuvels and Josef Steinebach. Exact convergence rates in strong approximation laws for large increments of partial sums. Probability Theory and Related Fields, 76(3):369–393, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01297492.

Dunau:1988:CDT

[DS88a] Jean-Louis Dunau and Henri Sénateur. Une caracterisation du type de la loi de Cauchy-conforme sur R^{n}. (French) [A characterization of the type of Cauchy-compliant law on R^{n}]. Probability Theory and Related Fields, 77(1):129–135, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01848135.

Durrett:1988:LDC

[DS88b] R. Durrett and R. H. Schonmann. Large deviations for the contact process and two dimensional percolation. Probability Theory and Related Fields, 77(4):583–603, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00959619.

Durrett:1994:CRC

[DS94] Rick Durrett and Glen Swindle. Coexistence results for catalysts. Probability Theory and Related Fields, 98(4):489–515, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01192836.

Delbaen:1995:APB

[DS95] F. Delbaen and W. Schachermayer. Arbitrage possibilities in Bessel processes and their relations to local martingales. Probability Theory and Related Fields, 102(3):357–366, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192466.

Dehghanpour:1997:NGM

[DS97] Pouria Dehghanpour and Roberto H. Schonmann. A nucleationand-growth model. Probability Theory and Related Fields, 107 (1):123–135, January 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050079.

Doring:2009:LSI

[DS09] Maik Döring and Wilhelm Stannat. The logarithmic Sobolev inequality for the Wasserstein diffusion. Probability Theory and Related Fields, 145(1–2):189–209, September 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0166-6.

Damron:2011:OIP

[DS11] Michael Damron and Artëm Sapozhnikov. Outlets of 2D invasion percolation and multiple-armed incipient infinite clusters. Probability Theory and Related Fields, 150(1–2):257–294, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00440-010-0274-y.pdf.

DeBlassie:2013:ETE

[DS13] Dante DeBlassie and Robert G. Smits. The expected time to end the tug-of-war in a wedge. Probability Theory and Related Fields, 155(1–2):347–378, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0400-5.

Dolinsky:2014:MOT

[DS14] Yan Dolinsky and H. Mete Soner. Martingale optimal transport and robust hedging in continuous time. Probability Theory and Related Fields, 160(1–2):391–427, October 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0531-y.

De:2018:NCL

[DS18] Anindya De and Rocco A. Servedio. A new central limit theorem and decomposition for Gaussian polynomials, with an application to deterministic approximate counting. Probability Theory and Related Fields, 171(3–4):981–1044, August 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0804-y.

Dolega:2019:GFJ

[DŚ19] Maciej Dołega and Piotr Śniady. Gaussian fluctuations of Jack-deformed random Young diagrams. Probability Theory and Related Fields, 174(1–2):133–176, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0854-9;   http://link.springer.com/content/pdf/
10.1007/s00440-018-0854-9.pdf.

Damron:2020:ZTG

[DS20] Michael Damron and Arnab Sen. Zero-temperature Glauber dynamics on the 3-regular tree and the median process. Probability Theory and Related Fields, 178(1–2):25–68, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00968-9.

Dauvergne:2023:SIH

[DS23] Duncan Dauvergne and Allan Sly. Spread of infections in a heterogeneous moving population. Probability Theory and Related Fields, 187(1–2):73–131, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-023-01216-
6.

Diaconis:1996:WGS

[DSC96] P. Diaconis and L. Saloff-Coste. Walks on generating sets of Abelian groups. Probability Theory and Related Fields, 105 (3):393–421, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192214.

Drton:2007:AFA

[DSS07] Mathias Drton, Bernd Sturmfels, and Seth Sullivant. Algebraic factor analysis: tetrads, pentads and beyond. Probability Theory and Related Fields, 138(3–4):463–493, July 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0033-2.

Drewitz:2014:SRA

[DSS14] Alexander Drewitz, Perla Sousi, and Rongfeng Sun. Symmetric rearrangements around infinity with applications to Lévy processes. Probability Theory and Related Fields, 158(3–4):637–664, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-013-0492-1.

Ding:2023:NCI

[DSS23] Jian Ding, Jian Song, and Rongfeng Sun. A new correlation inequality for Ising models with external fields. Probability Theory and Related Fields, 186(1–2):477–492, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01132-1.

DeBlassie:1993:SRR

[DT93] R. Dante DeBlassie and Ellen H. Toby. On the semimartingale representation of reflecting Brownian motion in a cusp. Probability Theory and Related Fields, 94(4):505–524, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192561.

DaPrato:2000:SAS

[DT00] Giuseppe Da Prato and Luciano Tubaro. Self-adjointness of some infinite-dimensional elliptic operators and application to stochastic quantization. Probability Theory and Related Fields, 118(1):131–145, September 2000. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008739;
 http://link.springer.com/article/10.1007/s004400000073.

deTiliere:2007:PFP

[dT07a] Béatrice de Tilière. Partition function of periodic isoradial dimer models. Probability Theory and Related Fields, 138(3– 4):451–462, July 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0041-2.

deTiliere:2007:QTP

[dT07b] Béatrice de Tilière. Quadri-tilings of the plane. Probability Theory and Related Fields, 137(3–4):487–518, March 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0002-9.

Delbaen:2010:HAS

[DT10] Freddy Delbaen and Shanjian Tang. Harmonic analysis of stochastic equations and backward stochastic differential equations. Probability Theory and Related Fields, 146(1–2):291– 336, January 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0191-5.

Du:2012:SSB

[DT12] Kai Du and Shanjian Tang. Strong solution of backward stochastic partial differential equations in C^{2} domains. Probability Theory and Related Fields, 154(1–2):255–285, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-011-0369-0.

Dodos:2023:ABE

[DT23] Pandelis Dodos and Konstantinos Tyros. Anticoncentration and Berry–Esseen bounds for random tensors. Probability Theory and Related Fields, 187(1–2):317–384, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01211-x.

Durhuus:2023:TEH

[DÜ23] Bergfinnur Durhuus and Meltem Ünel. Trees with exponential height dependent weight. Probability Theory and Related Fields, 186(3–4):999–1043, August 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01188-7.

Dubischar:2002:UCR

[Dub02] Daniel Dubischar. Uniformly converging random variables for weakly converging laws. Probability Theory and Related Fields, 123(4):601–605, August 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400200199.

Dubedat:2006:EDS

[Dub06] Julien Dubédat. Excursion decompositions for SLE and Watts’ crossing formula. Probability Theory and Related Fields, 134 (3):453–488, March 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-005-0446-3.

Dudley:1994:MMP

[Dud94] R. M. Dudley. Metric marginal problems for set-valued or nonmeasurable variables. Probability Theory and Related Fields, 100 (2):175–189, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199264.

Duffield:1992:LMF

[Duf92] N. G. Duffield. Local mean-field Markov processes: An application to message-switching networks. Probability Theory and Related Fields, 93(4):485–505, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01192718.

Dumbgen:1993:NFB

[Düm93] Lutz Dümbgen. On nondifferentiable functions and the bootstrap. Probability Theory and Related Fields, 95(1):125–140, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01197342.

Dungey:2007:TRR

[Dun07] Nick Dungey. Time regularity for random walks on locally compact groups. Probability Theory and Related Fields, 137(3–4): 429–442, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0006-5.

Dung:2024:RFI

[Dun24] Nguyen Tien Dung. Rates of Fisher information convergence in the central limit theorem for nonlinear statistics. Probability Theory and Related Fields, 190(3-4):625–673, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01331-y.

Dupont:2010:BCM

[Dup10] Christophe Dupont. Bernoulli coding map and almost sure invariance principle for endomorphisms of P^{k}. Probability Theory and Related Fields, 146(3–4):337–359, March 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0192-4.

Durot:2002:SAI

[Dur02] Cécile Durot. Sharp asymptotics for isotonic regression. Probability Theory and Related Fields, 122(2):222–240, February 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400100171.

Durand:2009:SSL

[Dur09] Arnaud Durand. Singularity sets of Lévy processes. Probability Theory and Related Fields, 143(3–4):517–544, March 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0134-6.

Dzhaparidze:1990:HTD

[DV90] K. Dzhaparidze and E. Valkeila. On the Hellinger type distances for filtered experiments. Probability Theory and Related Fields, 85(1):105–117, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01377632.

Deuschel:2000:NGS

[DV00] J.-D. Deuschel and Y. Velenik. Non-Gaussian surface pinned by a weak potential. Probability Theory and Related Fields, 116 (3):359–377, March 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400070004.

Davis:2002:CTV

[DV02] Burgess Davis and Stanislav Volkov. Continuous time vertexreinforced jump processes. Probability Theory and Related Fields, 123(2):281–300, June 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400100189.

Davis:2004:VRJ

[DV04] Burgess Davis and Stanislav Volkov. Vertex-reinforced jump processes on trees and finite graphs. Probability Theory and Related Fields, 128(1):42–62, January 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0286-y.

Debussche:2015:IMS

[DV15] A. Debussche and J. Vovelle. Invariant measure of scalar firstorder conservation laws with stochastic forcing. Probability Theory and Related Fields, 163(3–4):575–611, December 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0599-z.

Dekking:2001:FPB

[DvdW01] F. M. Dekking and P. v. d. Wal. Fractal percolation and branching cellular automata. Probability Theory and Related Fields, 120(2):277–308, June 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008784.

Dzhaparidze:2004:SEF

[DvZ04] Kacha Dzhaparidze and Harry van Zanten. A series expansion of fractional Brownian motion. Probability Theory and Related Fields, 130(1):39–55, September 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0310-2.

Dobrushin:1990:CAI

[DW90] R. L. Dobrushin and V. Warstat. Completely analytic interactions with infinite values. Probability Theory and Related Fields, 84(3):335–359, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01197889.

Dalang:1993:GLS

[DW93a] Robert C. Dalang and John B. Walsh. Geography of the level sets of the Brownian sheet. Probability Theory and Related Fields, 96(2):153–176, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192131.

Dalang:1993:SBB

[DW93b] Robert C. Dalang and John B. Walsh. The structure of a Brownian bubble. Probability Theory and Related Fields, 96 (4):475–501, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01200206.

Duquesne:2007:GLT

[DW07] Thomas Duquesne and Matthias Winkel. Growth of Lévy trees. Probability Theory and Related Fields, 139(3–4):313– 371, November 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0064-3.

Dimitrov:2021:KCR

[DW21] Evgeni Dimitrov and Xuan Wu. KMT coupling for random walk bridges. Probability Theory and Related Fields, 179(3– 4):649–732, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01030-y.

Ding:2023:PMP

[DWXY23] Jian Ding, Yihong Wu, Jiaming Xu, and Dana Yang. The planted matching problem: sharp threshold and infinite-order phase transition. Probability Theory and Related Fields, 187 (1–2):1–71, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01208-6.

Dwyer:1990:RCH

[Dwy90] Rex A. Dwyer. Random convex hulls in a product of balls. Probability Theory and Related Fields, 86(4):457–467, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01198169.

Dynkin:1991:PPH

[Dyn91a] E. B. Dynkin. Path processes and historical superprocesses. Probability Theory and Related Fields, 90(1):1–36, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01321132.

Dynkin:1991:PAO

[Dyn91b] E. B. Dynkin. A probabilistic approach to one class of nonlinear differential equations. Probability Theory and Related Fields, 89 (1):89–115, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01225827.

Dynkin:1993:BDM

[Dyn93a] E. B. Dynkin. Birth delay of a Markov process and the stopping distributions for regular processes. Probability Theory and Related Fields, 94(3):399–411, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01199250.

Dynkin:1993:RS

[Dyn93b] E. B. Dynkin. On regularity of superprocesses. Probability Theory and Related Fields, 95(2):263–281, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192273.

DaPrato:1995:CEC

[DZ95] G. Da Prato and J. Zabczyk. Convergence to equilibrium for classical and quantum spin systems. Probability Theory and Related Fields, 103(4):529–552, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01246338.

Dembo:1996:RGC

[DZ96] A. Dembo and O. Zeitouni. Refinements of the Gibbs conditioning principle. Probability Theory and Related Fields, 104 (1):1–14, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01303799.

Ding:2018:NUF

[DZ18] Jian Ding and Fuxi Zhang. Non-universality for first passage percolation on the exponential of log-correlated Gaussian fields. Probability Theory and Related Fields, 171(3–4):1157–1188, August 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-017-0811-z.

Ding:2019:LFP

[DZ19] Jian Ding and Fuxi Zhang. Liouville first passage percolation: geodesic length exponent is strictly larger than 1 at high temperatures. Probability Theory and Related Fields, 174(1– 2):335–367, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00905-5.

Dziubanski:1991:ABD

[Dzi91] Jacek Dziubański. Asymptotic behaviour of densities of stable semigroups of measures. Probability Theory and Related Fields, 87(4):459–467, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01304275.

Ebanks:1986:MII

[Eba86] Bruce R. Ebanks. Measures of inset information on open domains-II: Additive inset entropies with measurable sum property. Probability Theory and Related Fields, 73(4):517–528, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/BF00324849.

Eberle:1997:DPL

[Ebe97] Andreas Eberle. Diffusions on path and loop spaces: existence, finite dimensional approximation and Hölder continuity. Probability Theory and Related Fields, 109(1):77–99, September 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050126.

Eberle:2016:RCC

[Ebe16] Andreas Eberle. Reflection couplings and contraction rates for diffusions. Probability Theory and Related Fields, 166(3–4): 851–886, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0673-1;
http://link.springer.com/article/10.1007/s00440-015-
0673-1.

Erhard:2015:PAM

[EdHM15] D. Erhard, F. den Hollander, and G. Maillard. The parabolic Anderson model in a dynamic random environment: space– time ergodicity for the quenched Lyapunov exponent. Probability Theory and Related Fields, 162(1–2):1–46, June 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0564-x.

Eizenberg:1993:APP

[EF93] Alexander Eizenberg and Mark Freidlin. Averaging principle for perturbed random evolution equations and corresponding Dirichlet problems. Probability Theory and Related Fields, 94 (3):335–374, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199248.

Etheridge:1998:PTD

[EF98] Alison M. Etheridge and Klaus Fleischmann. Persistence of a two-dimensional super-Brownian motion in a catalytic medium. Probability Theory and Related Fields, 110(1):1–12, January 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050142.

Enriquez:2001:SWH

[EFL01] Nathanaël Enriquez, Jacques Franchi, and Yves Le Jan. Stable windings on hyperbolic surfaces. Probability Theory and Related Fields, 119(2):213–255, February 2001. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008759.

Efromovich:1999:RSO

[Efr99] Sam Efromovich. On rate and sharp optimal estimation. Probability Theory and Related Fields, 113(3):415–419, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050211.

Einmahl:1995:CBF

[EG95] U. Einmahl and V. Goodman. Clustering behavior of finite variance partial sum processes. Probability Theory and Related Fields, 102(4):547–565, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01198849.

Espinasse:2014:PEG

[EGL14] T. Espinasse, F. Gamboa, and J.-M. Loubes. Parametric estimation for Gaussian fields indexed by graphs. Probability Theory and Related Fields, 159(1–2):117–155, June 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0503-2.

Ekstrom:2013:CTH

[EHJT13] Erik Ekström, David Hobson, Svante Janson, and Johan Tysk. Can time-homogeneous diffusions produce any distribution? Probability Theory and Related Fields, 155(3–4):493–520, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-011-0405-0.

Einmahl:1988:RBA

[EHM88] J. H. J. Einmahl, E. Haeusler, and D. M. Mason. On the relationship between the almost sure stability of weighted empirical distributions and sums of order statistics. Probability Theory and Related Fields, 79(1):59–74, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00319104.

Eskenazis:2020:PIH

[EI20] Alexandros Eskenazis and Paata Ivanisvili. Polynomial inequalities on the Hamming cube. Probability Theory and Related Fields, 178(1–2):235–287, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-020-00973-
y.

Eichler:2012:GMM

[Eic12] Michael Eichler. Graphical modelling of multivariate time series. Probability Theory and Related Fields, 153(1–2):233–268, June 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00440-011-0345-8.pdf.

Eigen:1986:CED

[Eig86] S. J. Eigen. A counter-example to Dye’s theorem for all nonseparable measure algebras. Probability Theory and Related Fields, 72(3):471–475, June 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00334197.

Einmahl:1987:UEM

[Ein87] Uwe Einmahl. A useful estimate in the multidimensional invariance principle. Probability Theory and Related Fields, 76(1): 81–101, September 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00390277.

Einmahl:1988:SAP

[Ein88] Uwe Einmahl. Strong approximations for partial sums of i.i.d. B-valued r.v.’s in the domain of attraction of a Gaussian law.

Probability Theory and Related Fields, 77(1):65–85, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01848131.

Einmahl:1989:DET

[Ein89a] Uwe Einmahl. The Darling–Erdős theorem for sums of I.I.D. random variables. Probability Theory and Related Fields, 82 (2):241–257, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00354762.

Einmahl:1989:NLI

[Ein89b] Uwe Einmahl. A note on the law of the iterated logarithm in Hilbert space. Probability Theory and Related Fields, 82 (2):213–223, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00354760.

Einmahl:1992:ECR

[Ein92] Uwe Einmahl. Exact convergence rates for the bounded law of the iterated logarithm in Hilbert space. Probability Theory and Related Fields, 92(2):177–194, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01194920.

Einmahl:1993:OLI

[Ein93] Uwe Einmahl. On the other law of the iterated logarithm. Probability Theory and Related Fields, 96(1):97–106, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01195884.

Eisenbaum:1990:TRK

[Eis90] N. Eisenbaum. Un théorème de Ray–Knight lié au supremum des temps locaux browniens. (French) [A Ray–Knight theorem related supremum of Brownian local times]. Probability Theory and Related Fields, 87(1):79–95, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01217747.

Eisenbaum:1994:DIT

[Eis94] Nathalie Eisenbaum. Dynkin’s isomorphism theorem and the Ray–Knight theorems. Probability Theory and Related Fields, 99(2):321–335, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199028.

Eisenbaum:1997:MVS

[Eis97] Nathalie Eisenbaum. On the most visited sites by a symmetric stable process. Probability Theory and Related Fields, 107 (4):527–535, April 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050097.

Eisenbaum:2003:IDS

[Eis03] Nathalie Eisenbaum. On the infinite divisibility of squared Gaussian processes. Probability Theory and Related Fields, 125 (3):381–392, March 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-002-0245-z.

Eizenberg:1987:ABP

[EK87] Alexander Eizenberg and Yuri Kifer. The asymptotic behavior of the principal eigenvalue in a singular perturbation problem with invariant boundaries. Probability Theory and Related Fields, 76(4):439–476, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00960068.

Einmahl:1996:DPL

[EK96] U. Einmahl and J. Kuelbs. Dominating points and large deviations for random vectors. Probability Theory and Related Fields, 105(4):529–543, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01191912.

Elboim:2023:LLS

[EK23] Dor Elboim and Bo’az Klartag. Long lines in subsets of large measure in high dimension. Probability Theory and Related Fields, 187(3–4):657–695, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-023-01231-
7.

Eldan:2022:SCS

[EKZ22] Ronen Eldan, Frederic Koehler, and Ofer Zeitouni. A spectral condition for spectral gap: fast mixing in high-temperature Ising models. Probability Theory and Related Fields, 182(3– 4):1035–1051, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01085-x.

ElKaroui:2018:IPG

[El 18] Noureddine El Karoui. On the impact of predictor geometry on the performance on high-dimensional ridge-regularized generalized robust regression estimators. Probability Theory and Related Fields, 170(1–2):95–175, February 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0754-9;   http://link.springer.com/content/pdf/
10.1007/s00440-016-0754-9.pdf.

Eisele:1987:AWS

[EL87] T. Eisele and R. Lang. Asymptotics for the Wiener sausage with drift. Probability Theory and Related Fields, 74(1):125–140, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01845643.

Efromovich:1994:AEL

[EL94] Sam Efromovich and Mark G. Low. Adaptive estimates of linear functionals. Probability Theory and Related Fields, 98 (2):261–275, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192517.

Eichelsbacher:2004:MDO

[EL04] Peter Eichelsbacher and Matthias Löwe. Moderate deviations for the overlap parameter in the Hopfield model. Probability Theory and Related Fields, 130(4):441–472, December 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0349-8.

Ebrahimnejad:2021:PGU

[EL21] Farzam Ebrahimnejad and James R. Lee. On planar graphs of uniform polynomial growth. Probability Theory and Related Fields, 180(3–4):955–984, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01045-

Eldan:2020:TCT

[Eld20] Ronen Eldan. Taming correlations through entropy-efficient measure decompositions with applications to mean-field approximation. Probability Theory and Related Fields, 176(3– 4):737–755, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00924-2.

Efraim:2008:PTI

[ELP08] L. Ben Efraim and F. Lust-Piquard. Poincaré type inequalities on the discrete cube and in the CAR algebra. Probability Theory and Related Fields, 141(3–4):569–602, July 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0094-x.

Elworthy:1999:ISL

[ELY99] K. D. Elworthy, Xue-Mei Li, and M. Yor. The importance of strictly local martingales; applications to radial Ornstein– Uhlenbeck processes. Probability Theory and Related Fields, 115 (3):325–355, October 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050240.

ElKaroui:1990:MMS

[EM90] N. El Karoui and S. Méléard. Martingale measures and stochastic calculus. Probability Theory and Related Fields, 84(1): 83–101, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01288560.

Etheridge:1991:NS

[EM91] Alison Etheridge and Peter March. A note on superprocesses. Probability Theory and Related Fields, 89(2):141–147, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01366902.

Einmahl:1993:RCS

[EM93] Uwe Einmahl and David M. Mason. Rates of clustering in Strassen’s LIL for partial sum processes. Probability Theory and Related Fields, 97(4):479–487, ???? 1993. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192960.

Einmahl:1997:GAL

[EM97a] Uwe Einmahl and David M. Mason. Gaussian approximation of local empirical processes indexed by functions. Probability Theory and Related Fields, 107(3):283–311, March 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050086.

Einmahl:1997:SMI

[EM97b] Uwe Einmahl and David M. Mason. On the smallest maximal increment of partial sums of i.i.d. random variables. Probability Theory and Related Fields, 108(1):67–86, May 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050101.

Eberle:2013:QAE

[EM13] Andreas Eberle and Carlo Marinelli. Quantitative approximations of evolving probability measures and sequential Markov chain Monte Carlo methods. Probability Theory and Related Fields, 155(3–4):665–701, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0410-y.

Eldan:2020:SSS

[EM20] Ronen Eldan and Dan Mikulincer. Stability of the Shannon– Stam inequality via the Föllmer process. Probability Theory and Related Fields, 177(3–4):891–922, August 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-020-00967-w.

ElAlaoui:2021:CTS

[EM21] Ahmed El Alaoui and Andrea Montanari. On the computational tractability of statistical estimation on amenable graphs. Probability Theory and Related Fields, 181(4):815–864, December 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01092-y.

Evans:2013:CSN

[EMS13] Steven N. Evans, Ben Morris, and Arnab Sen. Coalescing systems of non-Brownian particles. Probability Theory and Related Fields, 156(1–2):307–342, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0429-0.

Elton:1992:IFS

[EP92] J. H. Elton and M. Piccioni. Iterated function systems a rising from recursive estimation problems. Probability Theory and Related Fields, 91(1):103–114, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01194492.

Evans:1997:SMC

[EP97] Steven N. Evans and Jim Pitman. Stopped Markov chains with stationary occupation times. Probability Theory and Related Fields, 109(3):425–433, November 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050138.

Epstein:2019:TPD

[EP19] Charles L. Epstein and Camelia A. Pop. Transition probabilities for degenerate diffusions arising in population genetics. Probability Theory and Related Fields, 173(1–2):537–603, February 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-018-0840-2.

Eplett:1986:ATS

[Epl86] W. J. R. Eplett. Approximation theory for the simulation of continuous Gaussian processes. Probability Theory and Related Fields, 73(2):159–181, September 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00339934.

Evans:2006:RPR

[EPW06] Steven N. Evans, Jim Pitman, and Anita Winter. Rayleigh processes, real trees, and root growth with re-grafting. Probability Theory and Related Fields, 134(1):81–126, January 2006. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0411-6.

Eberlein:1996:SAS

[ER96] E. Eberlein and M. Römersperger. Strong approximation of semimartingales and statistical processes. Probability Theory and Related Fields, 104(4):539–567, ???? 1996. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198166.

Erdos:1998:LTM

[Erd98] László Erdős. Lifschitz tail in a magnetic field: the nonclassical regime. Probability Theory and Related Fields, 112(3): 321–371, November 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050193.

Erdos:2001:LTM

[Erd01] László Erdős. Lifschitz tail in a magnetic field: coexistence of classical and quantum behavior in the borderline case. Probability Theory and Related Fields, 121(2):219–236, October 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
PL00008803.

Erhardsson:2000:SRR

[Erh00] Torkel Erhardsson. On stationary renewal reward processes where most rewards are zero. Probability Theory and Related Fields, 117(2):145–161, June 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050001.

Erickson:1986:RET

[Eri86] K. Bruce Erickson. A ratio ergodic theorem for increasing additive functionals. Probability Theory and Related Fields, 72 (4):493–504, July 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00344717.

Erickson:1990:CES

[Eri90] K. Bruce Erickson. Continuous extensions of skew product diffusions. Probability Theory and Related Fields, 85(1):73–89, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01377630.

Erschler:2009:RMS

[Ers09] Anna Erschler. Random mappings of scaled graphs. Probability Theory and Related Fields, 144(3–4):543–579, July 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0154-x.

Erschler:2019:AID

[Ers19] Anna Erschler. Almost invariance of distributions for random walks on groups. Probability Theory and Related Fields, 174 (1–2):445–476, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00915-3.

Errami:2002:IFF

[ERV02] Mohammed Errami, Francesco Russo, and Pierre Vallois. Itô’s formula for C^{1,λ}-functions of a Càdlàg process and related calculus. Probability Theory and Related Fields, 122(2):191– 221, February 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400100168.

Eichelsbacher:2003:RDM

[ES03] Peter Eichelsbacher and Uwe Schmock. Rank-dependent moderate deviations of u-empirical measures in strong topologies. Probability Theory and Related Fields, 126(1):61–90, May 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-003-0254-6.

Enriquez:2009:PRC

[ESZ09] Nathanaël Enriquez, Christophe Sabot, and Olivier Zindy. A probabilistic representation of constants in Kesten’s renewal theorem. Probability Theory and Related Fields, 144(3–4): 581–613, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0155-9.

Erschler:2012:SW

[ETW12] Anna Erschler, Bálint Tóth, and Wendelin Werner. Stuck walks. Probability Theory and Related Fields, 154(1–2):149–163, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-011-0365-4.

Eberhard:2021:MTC

[EV21] Sean Eberhard and Péter P. Varjú. Mixing time of the Chung–Diaconis–Graham random process. Probability Theory and Related Fields, 179(1–2):317–344, February 2021. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-020-01009-1.

Evans:1987:MPS

[Eva87] Steven N. Evans. Multiple points in the sample paths of a Lévy process. Probability Theory and Related Fields, 76(3):359–367, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01297491.

Evans:1989:EQD

[Eva89a] M. P. Evans. Existence of quantum diffusions. Probability Theory and Related Fields, 81(4):473–483, May 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00367298.

Evans:1989:RPL

[Eva89b] S. N. Evans. The range of a perturbed Lévy process. Probability Theory and Related Fields, 81(4):555–557, May 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00367302.

Evans:1990:ARM

[Eva90] Steven N. Evans. Association and random measures. Probability Theory and Related Fields, 86(1):1–19, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01207510.

Evans:2000:RIL

[Eva00a] Steven N. Evans. Right inverses of Lévy processes and stationary stopped local times. Probability Theory and Related Fields, 118(1):37–48, September 2000. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008741;
 http://link.springer.com/article/10.1007/s004400000072.

Evans:2000:SSB

[Eva00b] Steven N. Evans. Snakes and spiders: Brownian motion on Rtrees. Probability Theory and Related Fields, 117(3):361–386, July 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050010.

Erdos:2012:BUG

[EYY12] László Erdős, Horng-Tzer Yau, and Jun Yin. Bulk universality for generalized Wigner matrices. Probability Theory and Related Fields, 154(1–2):341–407, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0390-3.

Erschler:2022:LLN

[EZ22] Anna Erschler and Tianyi Zheng. Law of large numbers for the drift of the two-dimensional wreath product. Probability Theory and Related Fields, 182(3–4):999–1033, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01098-6.

Fabian:1990:CCS

[Fab90] Václav Fabian. Complete cubic spline estimation of nonparametric regression functions. Probability Theory and Related Fields, 85(1):57–64, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01377628.

Fagnola:1990:MCQ

[Fag90a] Franco Fagnola. A martingale characterization of quantum Poisson processes. Probability Theory and Related Fields, 84 (3):323–333, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01197888.

Fagnola:1990:QSD

[Fag90b] Franco Fagnola. On quantum stochastic differential equations with unbounded coefficients. Probability Theory and Related Fields, 86(4):501–516, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01198172.

Fagnola:1991:LTF

[Fag91] Franco Fagnola. A Lévy theorem for free noises. Probability Theory and Related Fields, 90(4):491–504, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192140.

Faggionato:2022:HLS

[Fag22] Alessandra Faggionato. Hydrodynamic limit of simple exclusion processes in symmetric random environments via duality and homogenization. Probability Theory and Related Fields, 184(3–4):1093–1137, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-022-01163-8.

Falk:1990:NGP

[Fal90] Michael Falk. A note on generalized Pareto distributions and the k upper extremes. Probability Theory and Related Fields, 85(4):499–503, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01203167.

Fehrman:2017:EIM

[Feh17] Benjamin J. Fehrman. On the existence of an invariant measure for isotropic diffusions in random environment. Probability Theory and Related Fields, 168(1–2):409–453, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/content/pdf/10.
1007/s00440-016-0714-4.pdf.

Feldman:2003:CGD

[Fel03] G. M. Feldman. A characterization of the Gaussian distribution on Abelian groups. Probability Theory and Related Fields, 126 (1):91–102, May 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0256-4.

Feldman:2005:CTL

[Fel05] G. M. Feldman. On a characterization theorem for locally compact abelian groups. Probability Theory and Related Fields, 133(3):345–357, November 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-005-0429-4.

Feng:1994:LDM

[Fen94] Shui Feng. Large deviations for Markov processes with mean field interaction and unbounded jumps. Probability Theory and Related Fields, 100(2):227–252, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01199267.

Ferreyra:1989:ASE

[Fer89] Guillermo Ferreyra. Approximation of stochastic equations driven by predictable processes. Probability Theory and Related Fields, 83(3):391–403, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00964371.

Fernique:1991:RFA

[Fer91] X. Fernique. Régularité de fonctions aléatoires gaussiennes stationnaires. (French) [Regularity of stationary random Gaussian functions]. Probability Theory and Related Fields, 88(4): 521–536, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192555.

Ferrari:1992:SFA

[Fer92] Pablo A. Ferrari. Shock fluctuations in asymmetric simple exclusion. Probability Theory and Related Fields, 91(1):81–101, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01194491.

Fernique:1994:ECM

[Fer94] X. Fernique. Sur l’équivalence de certaines mesures produit. (French) [On the equivalents of certain product measures]. Probability Theory and Related Fields, 98(1):77–90, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01311349.

Falkner:1991:SDR

[FF91] Neil Falkner and P. J. Fitzsimmons. Stopping distributions for right processes. Probability Theory and Related Fields, 89 (3):301–318, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01198789.

Ferrari:1994:SFA

[FF94] P. A. Ferrari and L. R. G. Fontes. Shock fluctuations in the asymmetric simple exclusion process. Probability Theory and Related Fields, 99(2):305–319, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01199027.

Fritz:1994:SSR

[FFL94] J. Fritz, T. Funaki, and J. L. Lebowitz. Stationary states of random Hamiltonian systems. Probability Theory and Related Fields, 99(2):211–236, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199023.

Fleming:2012:FBP

[FFN12] Benjamin J. Fleming, Peter J. Forrester, and Eric Nordenstam. A finitization of the bead process. Probability Theory and Related Fields, 152(1–2):321–356, February 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0324-5.

Freitas:2010:HTS

[FFT10] Ana Cristina Moreira Freitas, Jorge Milhazes Freitas, and Mike Todd. Hitting time statistics and extreme value theory. Probability Theory and Related Fields, 147(3–4):675–710, July 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-009-0221-y.

Ferrari:1988:MSS

[FG88] P. A. Ferrari and S. Goldstein. Microscopic stationary states for stochastic systems with particle flux. Probability Theory and Related Fields, 78(3):455–471, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00334207.

Fleischmann:1994:DCI

[FG94] Klaus Fleischmann and Andreas Greven. Diffusive clustering in an infinite system of hierarchically interacting diffusions. Probability Theory and Related Fields, 98(4):517–566, ???? 1994.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192837.

Flandoli:1995:MSS

[FG95] Franco Flandoli and Dariusz Gatarek. Martingale and stationary solutions for stochastic Navier–Stokes equations. Probability Theory and Related Fields, 102(3):367–391, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192467.

Flandoli:2002:GEV

[FG02] Franco Flandoli and Massimiliano Gubinelli. The Gibbs ensemble of a vortex filament. Probability Theory and Related Fields, 122(3):317–340, March 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400100163.

Fournier:2015:RCW

[FG15] Nicolas Fournier and Arnaud Guillin. On the rate of convergence in Wasserstein distance of the empirical measure. Probability Theory and Related Fields, 162(3–4):707–738, August 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0583-7.

Furlan:2019:WUC

[FG19] M. Furlan and M. Gubinelli. Weak universality for a class of 3 d stochastic reaction-diffusion models. Probability Theory and Related Fields, 173(3–4):1099–1164, April 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0849-6.

Fontbona:2009:MOT

[FGM09] Joaquin Fontbona, Hélène Guérin, and Sylvie Méléard. Measurability of optimal transportation and convergence rate for Landau type interacting particle systems. Probability Theory and Related Fields, 143(3–4):329–351, March 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0128-4.

Fribergh:2010:SST

[FGP10] Alexander Fribergh, Nina Gantert, and Serguei Popov. On slowdown and speedup of transient random walks in random environment. Probability Theory and Related Fields, 147(1– 2):43–88, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0201-2.

Flandoli:2017:SN

[FGS17] Franco Flandoli, Benjamin Gess, and Michael Scheutzow. Synchronization by noise. Probability Theory and Related Fields, 168(3–4):511–556, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-016-0716-2.

Forman:2018:REH

[FHP18] Noah Forman, Chris Haulk, and Jim Pitman. A representation of exchangeable hierarchies by sampling from random real trees. Probability Theory and Related Fields, 172(1–2):1–29, October 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-017-0799-4.

Foley:1990:GLC

[FHS90] R. D. Foley, T. P. Hill, and M. C. Spruill. A generalization of Lévy’s concentration-variance inequality. Probability Theory and Related Fields, 86(1):53–62, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01207513.

Faraud:2012:ASC

[FHS12] Gabriel Faraud, Yueyun Hu, and Zhan Shi. Almost sure convergence for stochastically biased random walks on trees. Probability Theory and Related Fields, 154(3–4):621–660, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-011-0379-y.

Funaki:2007:IPF

[FI07] Tadahisa Funaki and Kensuke Ishitani. Integration by parts formulae for Wiener measures on a path space between two curves. Probability Theory and Related Fields, 137(3–4):289–321, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-006-0010-9.

Forrester:2019:GRB

[FI19] Peter J. Forrester and Jesper R. Ipsen. A generalisation of the relation between zeros of the complex Kac polynomial and eigenvalues of truncated unitary matrices. Probability Theory and Related Fields, 175(3–4):833–847, December 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00903-7.

Fichtner:1991:RPC

[Fic91] Karl-Heinz Fichtner. Random permutations of countable sets. Probability Theory and Related Fields, 89(1):35–60, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01225824.

Fernholz:2013:PDR

[FIKP13] E. Robert Fernholz, Tomoyuki Ichiba, Ioannis Karatzas, and Vilmos Prokaj. Planar diffusions with rank-based characteristics and perturbed Tanaka equations. Probability Theory and Related Fields, 156(1–2):343–374, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0430-7.

Fill:1989:AEL

[Fil89] James Allen Fill. Asymptotic expansions for large deviation probabilities in the strong law of large numbers. Probability Theory and Related Fields, 81(2):213–233, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00319551.

Filipovic:2000:IMW

[Fil00] Damir Filipović. Invariant manifolds for weak solutions to stochastic equations. Probability Theory and Related Fields, 118(3):323–341, November 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/PL00008744.

Fontes:1999:CTD

[FIN99] L. R. G. Fontes, M. Isopi, and C. M. Newman. Chaotic time dependence in a disordered spin system. Probability Theory and Related Fields, 115(3):417–443, October 1999. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050244.

Finnoff:2002:ILD

[Fin02] William Finnoff. Integration of large-deviation kernels and applications to large deviations for evolutionary games. Probability Theory and Related Fields, 122(1):141–162, January 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100167.

Fitzsimmons:1987:EMP

[Fit87] P. J. Fitzsimmons. On the excursions of Markov processes in classical duality. Probability Theory and Related Fields, 75 (2):159–178, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00354031.

Fitzsimmons:1990:ETP

[Fit90] P. J. Fitzsimmons. On the equivalence of three potential principles for right Markov processes. Probability Theory and Related Fields, 84(2):251–265, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01197847.

Foondun:2023:SBP

[FJK23] Mohammud Foondun, Mathew Joseph, and Kunwoo Kim. Small ball probability estimates for the Hölder semi-norm of the stochastic heat equation. Probability Theory and Related Fields, 185(1–2):553–613, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-022-01153-w.

Faggionato:2009:HBS

[FJL09] A. Faggionato, M. Jara, and C. Landim. Hydrodynamic behavior of 1D subdiffusive exclusion processes with random conductances. Probability Theory and Related Fields, 144(3–4): 633–667, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0157-7.

Follmer:1997:ODU

[FK97] H. Föllmer and D. Kramkov. Optional decompositions under constraints. Probability Theory and Related Fields, 109(1): 1–25, September 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050122.

Freidlin:2010:NSP

[FK10] M. Freidlin and L. Koralov. Nonlinear stochastic perturbations of dynamical systems and quasi-linear parabolic PDE’s with a small parameter. Probability Theory and Related Fields, 147 (1–2):273–301, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0208-8.

Fribergh:2016:LTE

[FK16] Alexander Fribergh and Daniel Kious. Local trapping for elliptic random walks in random environments in Z^{d}. Probability Theory and Related Fields, 165(3–4):795–834, August 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0646-4;   http://link.springer.com/
article/10.1007/s00440-015-0646-4.

Fayolle:2004:BDP

[FKL04] Guy Fayolle, Maxim Krikun, and Jean-Marc Lasgouttes. Birth and death processes on certain random trees: classification and stationary laws. Probability Theory and Related Fields, 128(3): 386–418, March 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0311-1.

Foondun:2024:IEB

[FKN24] Mohammud Foondun, Davar Khoshnevisan, and Eulalia Nualart. Instantaneous everywhere-blowup of parabolic SPDEs. Probability Theory and Related Fields, 190(1–2):601–624, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-024-01263-7.

Flajolet:1988:DUR

[FKT88] P. Flajolet, P. Kirschenhofer, and R. F. Tichy. Deviations from uniformity in random strings. Probability Theory and Related Fields, 80(1):139–150, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00348756.

Freidlin:2019:DMP

[FKW19] Mark Freidlin, Leonid Koralov, and Alexander Wentzell. On diffusions in media with pockets of large diffusivity. Probability Theory and Related Fields, 175(1–2):559–578, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-018-00897-8.

Florchinger:1993:DNE

[FL93] Patrick Florchinger and Rémi Léandre. Décroissance non exponentielle du noyau de la chaleur. (French) [Nonexponential decay of the heat kernel]. Probability Theory and Related Fields, 95(2):237–262, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192272.

Fleischmann:1995:NAS

[FL95] Klaus Fleischmann and Jean-François Le Gall. A new approach to the single point catalytic super-Brownian motion. Probability Theory and Related Fields, 102(1):63–82, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01295222.

Franchi:1996:BCA

[FL96a] J. Franchi and Y. Le Jan. Brownian charges around loops. Probability Theory and Related Fields, 104(4):501–514, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01198164.

Freidlin:1996:LDP

[FL96b] M. I. Freidlin and T.-Y. Lee. Large deviation principle for the diffusion-transmutation processes and Dirichlet problem for PDE systems with small parameter. Probability Theory and Related Fields, 105(2):227–254, ???? 1996. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01203836.

Freidlin:1996:WFP

[FL96c] M. I. Freidlin and T.-Y. Lee. Wave front propagation and large deviations for diffusion-transmutation process. Probability Theory and Related Fields, 106(1):39–70, September 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050057.

Flandoli:2019:WNS

[FL19] Franco Flandoli and Dejun Luo. ρ-white noise solution to 2D stochastic Euler equations. Probability Theory and Related Fields, 175(3–4):783–832, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00902-8.

Flandoli:2021:HMT

[FL21] Franco Flandoli and Dejun Luo. High mode transport noise improves vorticity blow-up control in 3D Navier–Stokes equations. Probability Theory and Related Fields, 180(1–2):309–363, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01037-5.

Fleury:2012:PIM

[Fle12] B. Fleury. Poincaré inequality in mean value for Gaussian polytopes. Probability Theory and Related Fields, 152(1–2): 141–178, February 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0318-3.

Florchinger:1990:MCT

[Flo90] Patrick Florchinger. Malliavin calculus with time dependent coefficients and application to nonlinear filtering. Probability Theory and Related Fields, 86(2):203–223, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01474642.

Farfan:2019:SLD

[FLT19] J. Farfán, C. Landim, and K. Tsunoda. Static large deviations for a reaction–diffusion model. Probability Theory and Related Fields, 174(1–2):49–101, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-018-0858-5.

Feray:2022:RGS

[FLT22] Valentin Féray, Baptiste Louf, and Paul Thévenin. Random generation and scaling limits of fixed genus factorizations into transpositions. Probability Theory and Related Fields, 184(3–4): 681–748, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01170-9.

Fitzsimmons:1986:EMM

[FM86] P. J. Fitzsimmons and B. Maisonneuve. Excessive measures and Markov processes with random birth and death. Probability Theory and Related Fields, 72(3):319–336, June 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00334189.

Frigerio:1989:QPP

[FM89] Alberto Frigerio and Hans Maassen. Quantum Poisson processes and dilations of dynamical semigroups. Probability Theory and Related Fields, 83(4):489–508, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01845700.

Fleischmann:1990:EAM

[FM90] Klaus Fleischmann and Stanislav Alekseevich Molchanov. Exact asymptotics in a mean field model with random potential. Probability Theory and Related Fields, 86(2):239–251, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01474644.

Fleischmann:1997:SBM

[FM97] Klaus Fleischmann and Carl Mueller. A super-Brownian motion with a locally infinite catalytic mass. Probability Theory and Related Fields, 107(3):325–357, March 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050088.

Faggionato:2003:HLD

[FM03] Alessandra Faggionato and Fabio Martinelli. Hydrodynamic limit of a disordered lattice gas. Probability Theory and Related Fields, 127(4):535–608, December 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0305-z.

Fontes:2006:SRW

[FM06] L. R. G. Fontes and P. Mathieu. On symmetric random walks with random conductances on Z^{d}. Probability Theory and Related Fields, 134(4):565–602, April 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-005-0448-1.

Feray:2012:APM

[FM12a] Valentin Féray and Pierre-Loïc Méliot. Asymptotics of q-Plancherel measures. Probability Theory and Related Fields, 152 (3–4):589–624, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0331-6.

Forrester:2012:PPP

[FM12b] Peter J. Forrester and Anthony Mays. Pfaffian point process for the Gaussian real generalised eigenvalue problem. Probability Theory and Related Fields, 154(1–2):1–47, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-011-0361-8.

Fan:2019:SNR

[FM19a] Zhou Fan and Andrea Montanari. The spectral norm of random inner-product kernel matrices. Probability Theory and Related Fields, 173(1–2):27–85, February 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0830-4.

Filmus:2019:HIS

[FM19b] Yuval Filmus and Elchanan Mossel. Harmonicity and invariance on slices of the Boolean cube. Probability Theory and Related Fields, 175(3–4):721–782, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00900-w.

Fyodorov:2021:BSM

[FM21] Yan V. Fyodorov and Stephen Muirhead. The band structure of a model of spatial random permutation. Probability Theory and Related Fields, 179(3–4):543–587, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01019-z.

Frei:2011:CBB

[FMS11] Christoph Frei, Semyon Malamud, and Martin Schweizer. Convexity bounds for BSDE solutions, with applications to indifference valuation. Probability Theory and Related Fields, 150 (1–2):219–255, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0273-z.

Funaki:2001:LDG

[FN01] T. Funaki and T. Nishikawa. Large deviations for the Ginzburg– Landau Δϕ interface model. Probability Theory and Related Fields, 120(4):535–568, August 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/PL00008791.

Ferrari:2015:ASF

[FN15] Patrik L. Ferrari and Peter Nejjar. Anomalous shock fluctuations in TASEP and last passage percolation models. Probability Theory and Related Fields, 161(1–2):61–109, February 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-013-0544-6.

Frank:1987:BPB

[FNS87] M. J. Frank, R. B. Nelsen, and B. Schweizer. Best-possible bounds for the distribution of a sum — a problem of Kolmogorov. Probability Theory and Related Fields, 74(2):199–211, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00569989.

Franz:1997:GLS

[FNS97] U. Franz, D. Neuenschwander, and R. Schott. Gauss laws in the sense of Bernstein and uniqueness of embedding into convolution semigroups on quantum groups and braided groups. Probability Theory and Related Fields, 109(1):101–127, September 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050127.

Friz:2014:RPS

[FO14] Peter Friz and Harald Oberhauser. Rough path stability of (semi-)linear SPDEs. Probability Theory and Related Fields, 158(1–2):401–434, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-013-0483-2.

Foldes:1989:ILT

[Föl89] Antónia Földes. On the infimum of the local time of a Wiener process. Probability Theory and Related Fields, 82(4):545–563, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00341283.

Fontbona:2009:APC

[Fon09] J. Fontbona. Acknowledgement of priority concerning the article “A probabilistic interpretation and stochastic particle approximations of the three dimensional Navier–Stokes equation” Prob. Th. Rel. Fields. 136:102–156 (2006). Probability Theory and Related Fields, 143(3–4):657, March 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-008-0176-4.pdf.

Ford:2009:SPE

[For09] Kevin Ford. Sharp probability estimates for random walks with barriers. Probability Theory and Related Fields, 145(1–2):269– 283, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0168-4.

Forien:2021:PIM

[For21] Nicolas Forien. A planar Ising model of self-organized criticality. Probability Theory and Related Fields, 180(1–2):163–198, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01025-9.

Fourati:1998:PCP

[Fou98] S. Fourati. Points de croissance des processus de Lévy et théorie générale des processus. (French) [Growth points of Lévy processes and general theory of processes]. Probability Theory and Related Fields, 110(1):13–49, January 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050143.

Fannjiang:1996:DT

[FP96] Albert Fannjiang and George Papanicolaou. Diffusion in turbulence. Probability Theory and Related Fields, 105(3):279–334, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01192211.

Follmer:2000:IFM

[FP00] Hans Föllmer and Philip Protter. On Itô’s formula for multidimensional Brownian motion. Probability Theory and Related Fields, 116(1):1–20, January 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/PL00008719.

Feng:2024:CPM

[FPW24] Yu Feng, Eveliina Peltola, and Hao Wu. Connection probabilities of multiple FK–Ising interfaces. Probability Theory and Related Fields, 189(1–2):281–367, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-024-01269-
1.

Fannes:1988:IDC

[FQ88] M. Fannes and J. Quaegebeur. Infinitely divisible completely positive mappings. Probability Theory and Related Fields, 79 (3):369–403, October 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00342232.

Falk:1989:BDB

[FR89] M. Falk and R.-D. Reiss. Bootstrapping the distance between smooth bootstrap and sample quantile distribution. Probability Theory and Related Fields, 82(2):177–186, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00354758.

Foldes:1992:HVP

[FR92] Antónia Földes and Pál Révész. On hardly visited points of the Brownian motion. Probability Theory and Related Fields, 91(1):71–80, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01194490.

Fritz:1995:TDC

[FR95] J. Fritz and B. Rüdiger. Time dependent critical fluctuations of a one dimensional local mean field model. Probability Theory and Related Fields, 103(3):381–407, ???? 1995. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01195480.

Fagnola:2003:TRQ

[FR03] Franco Fagnola and Rolando Rebolledo. Transience and recurrence of quantum Markov semigroups. Probability Theory and Related Fields, 126(2):289–306, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0268-0.

Forrester:2004:CSD

[FR04] Peter J. Forrester and Eric M. Rains. Correlations for superpositions and decimations of Laguerre and Jacobi orthogonal matrix ensembles with a parameter. Probability Theory and Related Fields, 130(4):518–576, December 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0374-7.

Forrester:2005:ISP

[FR05] Peter J. Forrester and Eric M. Rains. Interpretations of some parameter dependent generalizations of classical matrix ensembles. Probability Theory and Related Fields, 131(1):1–61, January 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-004-0375-6.

Fountoulakis:2007:FMS

[FR07] N. Fountoulakis and B. A. Reed. Faster mixing and small bottlenecks. Probability Theory and Related Fields, 137(3–4):475–486, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-006-0003-8.

Flandoli:2008:MSS

[FR08] Franco Flandoli and Marco Romito. Markov selections for the 3D stochastic Navier–Stokes equations. Probability Theory and Related Fields, 140(3–4):407–458, March 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0069-y.

Franke:2005:FCL

[Fra05] Brice Franke. A functional central limit theorem for diffusions on periodic submanifolds of R^{N}. Probability Theory and Related Fields, 133(2):236–244, October 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0425-0.

Freslon:2019:CPR

[Fre19] Amaury Freslon. Cut-off phenomenon for random walks on free orthogonal quantum groups. Probability Theory and Related Fields, 174(3–4):731–760, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0863-8.

Fritz:1989:HLG

[Fri89] J. Fritz. On the hydrodynamic limit of a Ginzburg–Landau lattice model. Probability Theory and Related Fields, 81(2): 291–318, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319556.

Foutel-Rodier:2023:CGT

[FRS23] Félix Foutel-Rodier and Emmanuel Schertzer. Convergence of genealogies through spinal decomposition with an application to population genetics. Probability Theory and Related Fields, 187(3–4):697–751, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01223-7.

Frangos:1986:MEM

[FS86] Nikos E. Frangos and Louis Sucheston. On multiparameter ergodic and martingale theorems in infinite measure spaces. Probability Theory and Related Fields, 71(4):477–490, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00699037.

Ferrero:1988:PAO

[FS88] P. Ferrero and B. Schmitt. Produits aléatoires d’opérateurs matrices de transfert. (French) [Random products of transfer matrix operators]. Probability Theory and Related Fields, 79(2): 227–248, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00320920.

Freidlin:2000:DPG

[FS00] Mark Freidlin and Shuenn-Jyi Sheu. Diffusion processes on graphs: stochastic differential equations, large deviation principle. Probability Theory and Related Fields, 116(2):181–220, February 2000. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/PL00008726.

Fontes:2008:TCP

[FS08] Luiz Renato Fontes and Roberto H. Schonmann. Threshold θ ≥ 2 contact processes on homogeneous trees. Probability Theory and Related Fields, 141(3–4):513–541, July 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0092-z.

Feng:2010:SDP

[FS10] Shui Feng and Wei Sun. Some diffusion processes associated with two parameter Poisson–Dirichlet distribution and Dirichlet process. Probability Theory and Related Fields, 148(3–4): 501–525, November 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0238-2.

Friz:2017:EST

[FS17] Peter K. Friz and Atul Shekhar. On the existence of SLE trace: finite energy drivers and non-constant κ. Probability Theory and Related Fields, 169(1–2):353–376, October 2017. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0731-3.

Fang:2010:WSW

[FSS10] Shizan Fang, Jinghai Shao, and Karl-Theodor Sturm. Wasserstein space over the Wiener space. Probability Theory and Related Fields, 146(3–4):535–565, March 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-009-0199-5.

Fang:2019:CMA

[FSX19a] Xiao Fang, Qi-Man Shao, and Lihu Xu. Correction to: Multivariate approximations in Wasserstein distance by Stein’s method and Bismut’s formula. Probability Theory and Related Fields, 175(3–4):1177–1181, December 2019. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-019-
00930-4;   https://link.springer.com/content/pdf/10.
1007/s00440-019-00930-4.pdf.

Fang:2019:MAW

[FSX19b] Xiao Fang, Qi-Man Shao, and Lihu Xu. Multivariate approximations in Wasserstein distance by Stein’s method and Bismut’s formula. Probability Theory and Related Fields, 174(3–4): 945–979, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0874-5.

Fox:1987:CLT

[FT87] Robert Fox and Murad S. Taqqu. Central limit theorems for quadratic forms in random variables having long-range dependence. Probability Theory and Related Fields, 74(2):213–240, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00569990.

Fristedt:1992:PMG

[FT92] Bert E. Fristedt and S. James Taylor. The packing measure of a general subordinator. Probability Theory and Related Fields, 92(4):493–510, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01274265.

Fukushima:1996:CDR

[FT96] Masatoshi Fukushima and Matsuyo Tomisaki. Construction and decomposition of reflecting diffusions on Lipschitz domains with Hölder cusps. Probability Theory and Related Fields, 106(4): 521–557, December 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050074.

Fill:2004:AAM

[FT04] James Allen Fill and Fred Torcaso. Asymptotic analysis via Mellin transforms for small deviations in L^{2}-norm of integrated Brownian sheets. Probability Theory and Related Fields, 130(2): 259–288, October 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0363-x.

Forghani:2019:RWI

[FT19] Behrang Forghani and Giulio Tiozzo. Random walks of infinite moment on free semigroups. Probability Theory and Related Fields, 175(3–4):1099–1122, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00911-7.

Friz:2021:RSR

[FTY21] Peter K. Friz, Huy Tran, and Yizheng Yuan. Regularity of SLE in (t, κ) and refined GRR estimates. Probability Theory and Related Fields, 180(1–2):71–112, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01058-0.

Feyel:2004:MKM

[FÜ04] D. Feyel and A. S. Üstünel. Monge–Kantorovitch measure transportation and Monge–Ampère equation on Wiener space. Probability Theory and Related Fields, 128(3):347–385, March 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-003-0307-x.

Fuglede:1988:IPS

[Fug88] B. Fuglede. On the imbedding problem for stochastic and doubly stochastic matrices. Probability Theory and Related Fields, 80(2):241–260, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00356104.

Fukuda:1990:EIS

[Fuk90] Ryoji Fukuda. Exponential integrability of sub-Gaussian vectors. Probability Theory and Related Fields, 85(4):505–521, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01203168.

Fukuyama:1991:MCL

[Fuk91] Katusi Fukuyama. A mean central limit theorem for weakly multiplicative systems and its application to lacunary trigonometric series. Probability Theory and Related Fields, 89(2):159–179, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01366904.

Fukuyama:1994:CLT

[Fuk94] Katusi Fukuyama. The Central Limit Theorem for Riesz– Raikov sums. Probability Theory and Related Fields, 100(1): 57–75, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01204953.

Fukai:2004:HTH

[Fuk04] Yasunari Fukai. Hitting time of a half-line by two-dimensional random walk. Probability Theory and Related Fields, 128(3): 323–346, March 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0306-y.

Fukasawa:2008:EEE

[Fuk08] Masaaki Fukasawa. Edgeworth expansion for ergodic diffusions. Probability Theory and Related Fields, 142(1–2):1–20, September 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-007-0097-7.

Fukuyama:2011:CLT

[Fuk11] Katusi Fukuyama. A central limit theorem for trigonometric series with bounded gaps. Probability Theory and Related Fields, 149(1–2):139–148, February 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-009-0245-3.

Fulman:2009:CRM

[Ful09] Jason Fulman. Commutation relations and Markov chains. Probability Theory and Related Fields, 144(1–2):99–136, May 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-008-0143-0.

Funaki:1989:DHE

[Fun89] Tadahisa Funaki. Derivation of the hydrodynamical equation for one-dimensional Ginzburg–Landau model. Probability Theory and Related Fields, 82(1):39–93, June 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00340012.

Funaki:1991:HLS

[Fun91] Tadahisa Funaki. The hydrodynamic limit for a system with interactions prescribed by Ginzburg–Landau type random Hamiltonian. Probability Theory and Related Fields, 90(4):519–562, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01192142.

Funaki:1995:SLS

[Fun95] T. Funaki. The scaling limit for a stochastic PDE and the separation of phases. Probability Theory and Related Fields, 102 (2):221–288, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01213390.

Funaki:2003:HLI

[Fun03] Tadahisa Funaki. Hydrodynamic limit for Δϕ interface model on a wall. Probability Theory and Related Fields, 126(2):155–183, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-002-0238-y.

Fleischmann:2000:ITC

[FV00] Klaus Fleischmann and Vladimir A. Vatutin. An integral test for a critical multitype spatially homogeneous branching particle process and a related reaction-diffusion system. Probability Theory and Related Fields, 116(4):545–572, April 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050262.

Friz:2008:USO

[FV08] Peter Friz and Nicolas Victoir. On uniformly subelliptic operators and stochastic area. Probability Theory and Related Fields, 142(3–4):475–523, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0113-y.

Fitzner:2017:GAN

[FvdH17] Robert Fitzner and Remco van der Hofstad. Generalized approach to the non-backtracking lace expansion. Probability Theory and Related Fields, 169(3–4):1041–1119, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s00440-016-0747-8.pdf;   http://
link.springer.com/article/10.1007/s00440-016-0747-8.

Fill:1989:CRS

[FW89] James Allen Fill and Michael J. Wichura. The convergence rate for the strong law of large numbers: General lattice distributions. Probability Theory and Related Fields, 81(2):189–212, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00319550.

Freidlin:2004:RPD

[FW04] Mark Freidlin and Matthias Weber. Random perturbations of dynamical systems and diffusion processes with conservation laws. Probability Theory and Related Fields, 128(3):441–466, March 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0312-0. See correction [FW07].

Freidlin:2007:CRP

[FW07] Mark Freidlin and Matthias Weber. Correction to “Random perturbations of dynamical systems and diffusion processes with conservation laws”. Probability Theory and Related Fields, 137(3–4):595–596, March 2007. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s00440-006-0027-0. See [FW04].

Fleischmann:2008:LDS

[FW08] Klaus Fleischmann and Vitali Wachtel. Large deviations for sums indexed by the generations of a Galton–Watson process. Probability Theory and Related Fields, 141(3–4):445–470, July 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-007-0090-1.

Freidlin:2012:NPP

[FW12] M. I. Freidlin and A. D. Wentzell. On the Neumann problem for PDE’s with a small parameter and the corresponding diffusion processes. Probability Theory and Related Fields, 152(1–2): 101–140, February 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0317-4.

FitzGerald:2020:PLL

[FW20] Will FitzGerald and Jon Warren. Point-to-line last passage percolation and the invariant measure of a system of reflecting Brownian motions. Probability Theory and Related Fields, 178(1–2):121–171, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-020-00972-z.

Fan:2024:RSF

[FW24] Zhou Fan and Yihong Wu. The replica-symmetric free energy for Ising spin glasses with orthogonally invariant couplings. Probability Theory and Related Fields, 190(1–2):1–77, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01309-w.

Fang:1999:STB

[FZ99] S. Fang and T. S. Zhang. On the small time behavior of Ornstein–Uhlenbeck processes with unbounded linear drifts. Probability Theory and Related Fields, 114(4):487–504, July 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050232.

Fang:2005:SCS

[FZ05] Shizan Fang and Tusheng Zhang. A study of a class of stochastic differential equations with non-Lipschitzian coefficients. Probability Theory and Related Fields, 132(3):356–390, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0398-z.

Gajek:1989:EDD

[Gaj89] Leslaw Gajek. Estimating a density and its derivatives via the minimum distance method. Probability Theory and Related Fields, 80(4):601–617, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00318908.

Galashin:2022:FBC

[Gal22] Pavel Galashin. A formula for boundary correlations of the critical Ising model. Probability Theory and Related Fields, 182(1–2): 615–640, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01086-w.

Gamarnik:2004:LPT

[Gam04] David Gamarnik. Linear phase transition in random linear constraint satisfaction problems. Probability Theory and Related Fields, 129(3):410–440, July 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0345-z.

Gamarnik:2014:RCS

[Gam14] David Gamarnik. Right-convergence of sparse random graphs. Probability Theory and Related Fields, 160(1–2):253–278, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-013-0528-6.

Gantert:1994:SSB

[Gan94] Nina Gantert. Self-similarity of Brownian motion and a large deviation principle for random fields on a binary tree. Probability Theory and Related Fields, 98(1):7–20, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01311346.

Gandolfi:1999:RGS

[Gan99] Alberto Gandolfi. A remark on gauge symmetries in Ising spin glasses. Probability Theory and Related Fields, 114(4):419–430, July 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050230.

Garnier:1995:SII

[Gar95] J. Garnier. Stochastic invariant imbedding. Probability Theory and Related Fields, 103(2):249–271, ???? 1995. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01204217.

Gayrard:2019:AMD

[Gay19] Véronique Gayrard. Aging in Metropolis dynamics of the REM: a proof. Probability Theory and Related Fields, 174(1– 2):501–551, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0873-6.

Guerre-Delabriere:1990:EIR

[GD90] Sylvie Guerre-Delabriere. An example of an irregular random measure. Probability Theory and Related Fields, 86(2):155–165, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01474640.

Greven:1992:BRW

[GdH92] Andreas Greven and Frank den Hollander. Branching random walk in random environment: phase transitions for local and global growth rates. Probability Theory and Related Fields, 91 (2):195–249, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01291424.

Greven:1994:VPI

[GdH94] A. Greven and F. den Hollander. On a variational problem for an infinite particle system in a random medium. Part II: The local growth rate. Probability Theory and Related Fields, 100 (3):301–328, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193703.

Gartner:1999:CSI

[GdH99] J. Gärtner and F. den Hollander. Correlation structure of intermittency in the parabolic Anderson model. Probability Theory and Related Fields, 114(1):1–54, April 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050220.

Goodman:2014:EGB

[GdH14] Jesse Goodman and Frank den Hollander. Extremal geometry of a Brownian porous medium. Probability Theory and Related Fields, 160(1–2):127–174, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-013-0525-9.

Geiss:2005:WBD

[Gei05] Stefan Geiss. Weighted BMO and discrete time hedging within the Black–Scholes model. Probability Theory and Related Fields, 132(1):39–73, May 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0389-0.

Georgii:1994:LDE

[Geo94] Hans-Otto Georgii. Large deviations and the equivalence of ensembles for Gibbsian particle systems with superstable interaction. Probability Theory and Related Fields, 99(2):171–195, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01199021.

Gerritse:1986:SSD

[Ger86] Gerard Gerritse. Supremum self-decomposable random vectors. Probability Theory and Related Fields, 72(1):17–33, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00343894.

Goldsheid:1996:ZCD

[GG96] I. Ya. Goldsheid and Y. Guivarc’h. Zariski closure and the dimension of the Gaussian low of the product of random matrices. I. Probability Theory and Related Fields, 105(1):109–142, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01192073.

Ghosh:2011:CMS

[GG11] Subhankar Ghosh and Larry Goldstein. Concentration of measures via size-biased couplings. Probability Theory and Related Fields, 149(1–2):271–278, February 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-009-0253-3.

Gassiat:2019:RNS

[GG19] Paul Gassiat and Benjamin Gess. Regularization by noise for stochastic Hamilton–Jacobi equations. Probability Theory and Related Fields, 173(3–4):1063–1098, April 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0848-7.

Gassiat:2020:SPH

[GGLS20] Paul Gassiat, Benjamin Gess, Pierre-Louis Lions, and Panagiotis E. Souganidis. Speed of propagation for Hamilton–Jacobi equations with multiplicative rough time dependence and convex Hamiltonians. Probability Theory and Related Fields, 176(1– 2):421–448, February 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00921-5.

Gentil:2005:MLS

[GGM05] Ivan Gentil, Arnaud Guillin, and Laurent Miclo. Modified logarithmic Sobolev inequalities and transportation inequalities. Probability Theory and Related Fields, 133(3):409–436, November 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-005-0432-9.

Gotze:1989:AEP

[GH89] F. Götze and C. Hipp. Asymptotic expansions for potential functions of i.i.d. random fields. Probability Theory and Related Fields, 82(3):349–370, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00339992.

Greenwood:1991:DAO

[GH91] Priscilla E. Greenwood and Gerard Hooghiemstra. On the domain of attraction of an operator between supremum and sum. Probability Theory and Related Fields, 89(2):201–210, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01366906.

Gloter:2010:NRM

[GH10] Arnaud Gloter and Marc Hoffmann. Nonparametric reconstruction of a multifractal function from noisy data. Probability Theory and Related Fields, 146(1–2):155–187, January 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0187-1.

Gerencser:2019:SSD

[GH19a] Máté Gerencsér and Martin Hairer. Singular SPDEs in domains with boundaries. Probability Theory and Related Fields, 173(3–4):697–758, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0841-1;
 http://link.springer.com/content/pdf/10.1007/s00440-
018-0841-1.pdf.

Gravner:2019:PBP

[GH19b] Janko Gravner and Alexander E. Holroyd. Polluted bootstrap percolation with threshold two in all dimensions. Probability Theory and Related Fields, 175(1–2):467–486, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-018-0892-3.

Gwynne:2020:ADR

[GH20] Ewain Gwynne and Tom Hutchcroft. Anomalous diffusion of random walk on random planar maps. Probability Theory and Related Fields, 178(1–2):567–611, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00986-7.

Ganguly:2023:OTE

[GH23] Shirshendu Ganguly and Milind Hegde. Optimal tail exponents in general last passage percolation via bootstrapping & geodesic geometry. Probability Theory and Related Fields, 186 (1–2):221–284, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01204-w.

Gonzalez:2022:AMS

[GHK22] Isaac Gonzalez, Emma Horton, and Andreas E. Kyprianou. Asymptotic moments of spatial branching processes. Probability Theory and Related Fields, 184(3–4):805–858, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01131-2. See correction [GHK23].

Gonzalez:2023:CAM

[GHK23] Isaac Gonzalez, Emma Horton, and Andreas E. Kyprianou. Correction to: Asymptotic moments of spatial branching processes. Probability Theory and Related Fields, 187(1–2):505–515, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01220-w. See [GHK22].

Gravner:2012:STB

[GHM12] Janko Gravner, Alexander E. Holroyd, and Robert Morris. A sharper threshold for bootstrap percolation in two dimensions. Probability Theory and Related Fields, 153(1–2):1–23, June 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00440-010-0338-z.pdf.

Gwynne:2020:DTF

[GHM20] Ewain Gwynne, Nina Holden, and Jason Miller. Dimension transformation formula for conformal maps into the complement of an SLE curve. Probability Theory and Related Fields, 176(1–2):649–667, February 2020. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-019-
00952-y;   https://link.springer.com/content/pdf/10.
1007/s00440-019-00952-y.pdf.

Galeati:2023:DDS

[GHM23] Lucio Galeati, Fabian A. Harang, and Avi Mayorcas. Distribution dependent SDEs driven by additive fractional Brownian motion. Probability Theory and Related Fields, 185(1–2): 251–309, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01145-w.

Ghosh:2015:DPC

[Gho15] Subhroshekhar Ghosh. Determinantal processes and completeness of random exponentials: the critical case. Probability Theory and Related Fields, 163(3–4):643–665, December 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0601-9.

Gwynne:2019:DEL

[GHS19] Ewain Gwynne, Nina Holden, and Xin Sun. A distance exponent for Liouville quantum gravity. Probability Theory and Related Fields, 173(3–4):931–997, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-018-0846-9.

Gwynne:2020:MTA

[GHS20] Ewain Gwynne, Nina Holden, and Xin Sun. A mating-of-trees approach for graph distances in random planar maps. Probability Theory and Related Fields, 177(3–4):1043–1102, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00969-8.

Garban:2021:LDP

[GHS21] Christophe Garban, Nina Holden, and Xin Sun. Liouville dynamical percolation. Probability Theory and Related Fields, 180(3–4):621–678, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-021-01057-1.

Gine:1990:MID

[GHV90] Evarist Giné, Marjorie G. Hahn, and Pirooz Vatan. Maxinfinitely divisible and max-stable sample continuous processes. Probability Theory and Related Fields, 87(2):139–165, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01198427.

Gnedin:2020:NIO

[GI20] Alexander Gnedin and Alexander Iksanov. On nested infinite occupancy scheme in random environment. Probability Theory and Related Fields, 177(3–4):855–890, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00963-0.

Giacomin:1995:OSI

[Gia95] G. Giacomin. Onset and structure of interfaces in a Kawasaki + Glauber interacting particle system. Probability Theory and Related Fields, 103(1):1–24, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01199029.

Gao:2022:SDR

[GIM22] Pu Gao, Mikhail Isaev, and Brendan D. McKay. Sandwiching dense random regular graphs between binomial random graphs. Probability Theory and Related Fields, 184(1–2):115– 158, October 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01157-6. See correction

[GIM23].

Gao:2023:CSD

[GIM23] Pu Gao, Mikhail Isaev, and Brendan D. McKay. Correction: Sandwiching dense random regular graphs between binomial random graphs. Probability Theory and Related Fields, 187 (1–2):517, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01221-9. See [GIM22].

Ginovian:1994:TTQ

[Gin94] M. S. Ginovian. On Toeplitz type quadratic functionals of stationary Gaussian processes. Probability Theory and Related Fields, 100(3):395–406, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01193706.

Goncalves:2018:DFE

[GJ18] Patrícia Gonçalves and Milton Jara. Density fluctuations for exclusion processes with long jumps. Probability Theory and Related Fields, 170(1–2):311–362, February 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0758-0.

Goncalves:2024:CNR

[GJMM24] P. Gonçalves, M. Jara, R. Marinho, and O. Menezes. CLT for NESS of a reaction-diffusion model. Probability Theory and Related Fields, 190(1–2):337–377, October 2024. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-024-01293-1.

Gamarnik:2021:OGP

[GJS21] David Gamarnik, Aukosh Jagannath, and Subhabrata Sen. The overlap gap property in principal submatrix recovery. Probability Theory and Related Fields, 181(4):757–814, December 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01089-7.

Gijbels:1988:WSR

[GJV88] Irène Gijbels, Paul Janssen, and Noël Veraverbeke. Weak and strong representations for trimmed U-statistics. Probability Theory and Related Fields, 77(2):179–194, February 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00334036.

Goodman:1991:RCS

[GK91] Victor Goodman and James Kuelbs. Rates of clustering for some Gaussian self-similar processes. Probability Theory and Related Fields, 88(1):47–75, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01193582.

Griffin:1996:SRQ

[GK96a] P. Griffin and J. Kuelbs. Some remarks on a question of Strassen. Probability Theory and Related Fields, 104(2):211–229, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01247838. See [Str64].

Gyongy:1996:ESS

[GK96b] István Gyöngy and Nicolai Krylov. Existence of strong solutions for Itô’s stochastic equations via approximations. Probability Theory and Related Fields, 105(2):143–158, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01203833.

Grigoryan:2001:ASI

[GK01] Alexander Grigor’yan and Mark Kelbert. Asymptotic separation for independent trajectories of Markov processes. Probability Theory and Related Fields, 119(1):31–69, January 2001.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
PL00012738.

Grigorescu:2012:IPC

[GK12] Ilie Grigorescu and Min Kang. Immortal particle for a catalytic branching process. Probability Theory and Related Fields, 153 (1–2):333–361, June 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0347-6.

Gu:2024:EDP

[GK24] Yu Gu and Tomasz Komorowski. Effective diffusivities in periodic KPZ. Probability Theory and Related Fields, 190(3-4): 1007–1061, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-024-01297-x.

Gandolfi:1989:ETC

[GKdV89] A. Gandolfi, M. Keane, and V. de Valk. Extremal twocorrelations of two-valued stationary one-dependent processes. Probability Theory and Related Fields, 80(3):475–480, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01794435.

Gartner:2000:ASA

[GKM00] J. Gärtner, W. König, and S. A. Molchanov. Almost sure asymptotics for the continuous parabolic Anderson model. Probability Theory and Related Fields, 118(4):547–573, December 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008754.

Gandolfi:1992:UIC

[GKN92] A. Gandolfi, M. S. Keane, and C. M. Newman. Uniqueness of the infinite component in a random graph with applications to percolation and spin glasses. Probability Theory and Related Fields, 92(4):511–527, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01274266.

Grothaus:2007:LSD

[GKR07] Martin Grothaus, Yuri G. Kondratiev, and Michael Röckner. N/V -limit for stochastic dynamics in continuous particle systems. Probability Theory and Related Fields, 137(1–2):121– 160, January 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0499-y.

Gine:2004:KDE

[GKS04] Evarist Giné, Vladimir Koltchinskii, and Lyudmila Sakhanenko. Kernel density estimators: convergence in distribution for weighted sup-norms. Probability Theory and Related Fields, 130(2):167–198, October 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0339-x.

Guneysu:2016:FKI

[GKS16] Batu Güneysu, Matthias Keller, and Marcel Schmidt. A Feynman–Kac–Itô formula for magnetic Schrödinger operators on graphs. Probability Theory and Related Fields, 165(1–2): 365–399, June 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0633-9.

Greven:2001:IFW

[GKW01] Andreas Greven, Achim Klenke, and Anton Wakolbinger. Interacting Fisher–Wright diffusions in a catalytic medium. Probability Theory and Related Fields, 120(1):85–117, May 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
PL00008777.

Grimmett:1993:RWI

[GKZ93] G. R. Grimmett, H. Kesten, and Y. Zhang. Random walk on the infinite cluster of the percolation model. Probability Theory and Related Fields, 96(1):33–44, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01195881.

Gentz:1999:FOH

[GL99] Barbara Gentz and Matthias Löwe. The fluctuations of the overlap in the Hopfield model with finitely many patterns at the critical temperature. Probability Theory and Related Fields, 115 (3):357–381, October 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050241.

Gozlan:2007:LDA

[GL07] Nathael Gozlan and Christian Léonard. A large deviation approach to some transportation cost inequalities. Probability Theory and Related Fields, 139(1–2):235–283, September 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0045-y.

Goldenshluger:2009:SAM

[GL09] Alexander Goldenshluger and Oleg Lepski. Structural adaptation via L_{p}-norm oracle inequalities. Probability Theory and Related Fields, 143(1–2):41–71, January 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0119-5.

Gaudilliere:2014:DPN

[GL14a] A. Gaudillière and C. Landim. A Dirichlet principle for non reversible Markov chains and some recurrence theorems. Probability Theory and Related Fields, 158(1–2):55–89, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-012-0477-5.

Goldenshluger:2014:AMD

[GL14b] A. Goldenshluger and O. Lepski. On adaptive minimax density estimation on R^{d}. Probability Theory and Related Fields, 159(3– 4):479–543, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0512-1.

Grama:2020:CLL

[GLL20] Ion Grama, Ronan Lauvergnat, and Émile Le Page. Conditioned local limit theorems for random walks defined on finite Markov chains. Probability Theory and Related Fields, 176(1–2): 669–735, February 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00948-8.

Grama:2017:CLT

[GLP17] Ion Grama, Émile Le Page, and Marc Peigné. Conditioned limit theorems for products of random matrices. Probability Theory and Related Fields, 168(3–4):601–639, August 2017. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0719-z.

Ganguly:2017:FIC

[GLPP17] Shirshendu Ganguly, Lionel Levine, Yuval Peres, and James Propp. Formation of an interface by competitive erosion. Probability Theory and Related Fields, 168(1–2):455–509, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Giacomin:2010:HPM

[GLT10] Giambattista Giacomin, Hubert Lacoin, and Fabio Lucio Toninelli. Hierarchical pinning models, quadratic maps and quenched disorder. Probability Theory and Related Fields, 147 (1–2):185–216, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0205-y.

Guillin:2009:TII

[GLWY09] Arnaud Guillin, Christian Léonard, Liming Wu, and Nian Yao. Transportation-information inequalities for Markov processes. Probability Theory and Related Fields, 144(3–4):669–695, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-008-0159-5.

Graczyk:2011:SCL

[GLZ11] Piotr Graczyk, Jean-Jacques Loeb, and Tomasz Zak. Strong Central Limit Theorem for isotropic random walks in R^{d}. Probability Theory and Related Fields, 151(1–2):153–172, October 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-010-0295-6.

Graham:1989:SIP

[GM89] Carl Graham and Michel Métivier. System of interacting particles and nonlinear diffusion reflecting in a domain with sticky boundary. Probability Theory and Related Fields, 82(2):225–240, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00354761.

Graham:1994:CHS

[GM94] Carl Graham and Sylvie Méléard. Chaos hypothesis for a system interacting through shared resources. Probability Theory and Related Fields, 100(2):157–174, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01199263.

Gartner:1998:PPA

[GM98] J. Gärtner and S. A. Molchanov. Parabolic problems for the Anderson model. Probability Theory and Related Fields, 111 (1):17–55, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050161.

Gidas:2005:OTP

[GM05a] Basilis Gidas and Alejandro Murua. Optimal transformations for prediction in continuous-time stochastic processes: finite past and future. Probability Theory and Related Fields, 131 (4):479–492, April 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0371-x.

Guionnet:2005:CEM

[GM05b] Alice Guionnet and Mylène Maïda. Character expansion method for the first order asymptotics of a matrix integral. Probability Theory and Related Fields, 132(4):539–578, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0403-6.

Guionnet:2005:LTB

[GM05c] Alice Guionnet and Christian Mazza. Long time behaviour of the solution to non-linear Kraichnan equations. Probability Theory and Related Fields, 131(4):493–518, April 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0382-7.

Guan:2006:RSS

[GM06] Qing-Yang Guan and Zhi-Ming Ma. Reflected symmetric αstable processes and regional fractional Laplacian. Probability Theory and Related Fields, 134(4):649–694, April 2006. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
005-0438-3.

Gloria:2012:SMA

[GM12] Antoine Gloria and Jean-Christophe Mourrat. Spectral measure and approximation of homogenized coefficients. Probability Theory and Related Fields, 154(1–2):287–326, October 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0370-7.

Gordina:2013:STI

[GM13] Maria Gordina and Tai Melcher. A subelliptic Taylor isomorphism on infinite-dimensional Heisenberg groups. Probability Theory and Related Fields, 155(1–2):379–426, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-011-0401-4.

Grimmett:2014:BPI

[GM14] Geoffrey R. Grimmett and Ioan Manolescu. Bond percolation on isoradial graphs: criticality and universality. Probability Theory and Related Fields, 159(1–2):273–327, June 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0507-y.

Gu:2016:PTS

[GM16] Yu Gu and Jean-Christophe Mourrat. Pointwise two-scale expansion for parabolic equations with random coefficients. Probability Theory and Related Fields, 166(1–2):585–618, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-015-0667-z.

Guneysu:2017:SDE

[GMM17] B. Güneysu, O. Matte, and J. S. Møller. Stochastic differential equations for models of non-relativistic matter interacting with quantized radiation fields. Probability Theory and Related Fields, 167(3–4):817–915, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-016-0694-4.

Grazieschi:2024:MDI

[GMW24] P. Grazieschi, K. Matetski, and H. Weber. Martingale-driven integrals and singular SPDEs. Probability Theory and Related Fields, 190(3-4):1063–1120, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-024-01311-
2.

Grama:1998:AEN

[GN98] Ion Grama and Michael Nussbaum. Asymptotic equivalence for nonparametric generalized linear models. Probability Theory and Related Fields, 111(2):167–214, June 1998. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050166.

Gine:2008:UCL

[GN08] Evarist Giné and Richard Nickl. Uniform central limit theorems for kernel density estimators. Probability Theory and Related Fields, 141(3–4):333–387, July 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0087-9.

Gine:2009:EID

[GN09] Evarist Giné and Richard Nickl. An exponential inequality for the distribution function of the kernel density estimator, with applications to adaptive estimation. Probability Theory and Related Fields, 143(3–4):569–596, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-008-0137-y.

Ganguly:2022:LDL

[GN22] Shirshendu Ganguly and Kyeongsik Nam. Large deviations for the largest eigenvalue of Gaussian networks with constant average degree. Probability Theory and Related Fields, 184(3–4): 613–679, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01164-7.

Gyongy:1995:AST

[GNSS95] István Gyöngy, David Nualart, and Marta Sanz-Sole. Approximation and support theorems in modulus spaces. Probability Theory and Related Fields, 101(4):495–509, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01202782.

Guillin:2023:QSD

[GNW23] Arnaud Guillin, Boris Nectoux, and Liming Wu. Quasistationary distribution for Hamiltonian dynamics with singular potentials. Probability Theory and Related Fields, 185(3– 4):921–959, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01154-9.

Goldman:1996:SCM

[Gol96] André Goldman. Le spectre de certaines mosaïques poissoniennes du plan et l’enveloppe convexe du pont brownien. (French) [The spectrum of certain planar Poissonian mosaics and the convex hull of the Brownian bridge]. Probability Theory and Related Fields, 105(1):57–83, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192071.

Golubev:2004:PPE

[Gol04] Yu Golubev. The principle of penalized empirical risk in severely ill-posed problems. Probability Theory and Related Fields, 130 (1):18–38, September 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0362-y.

Goldsheid:2007:STR

[Gol07] Ilya Ya. Goldsheid. Simple transient random walks in onedimensional random environment: the central limit theorem. Probability Theory and Related Fields, 139(1–2):41–64, September 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-006-0038-x.

Goldsheid:2008:LSL

[Gol08] Ilya Ya. Goldsheid. Linear and sub-linear growth and the CLT for hitting times of a random walk in random environment on a strip. Probability Theory and Related Fields, 141(3–4):471–511, July 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-007-0091-0.

Gordon:1987:ECD

[Gor87] Yehoram Gordon. Elliptically contoured distributions. Probability Theory and Related Fields, 76(4):429–438, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00960067.

Gordon:1992:MGP

[Gor92] Yehoram Gordon. Majorization of Gaussian processes and geometric applications. Probability Theory and Related Fields, 91 (2):251–267, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01291425.

Gouezel:2004:CLT

[Gou04] Sébastien Gouëzel. Central limit theorem and stable laws for intermittent maps. Probability Theory and Related Fields, 128 (1):82–122, January 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0300-4.

Gassiat:2021:FBC

[GOZ21] Paul Gassiat, Harald Oberhauser, and Christina Z. Zou. A free boundary characterisation of the Root barrier for Markov processes. Probability Theory and Related Fields, 180(1–2):33– 69, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01052-6.

Gyongy:1993:QLS

[GP93] I. Gyöngy and E. Pardoux. On quasi-linear stochastic partial differential equations. Probability Theory and Related Fields, 94 (4):413–425, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192556.

Goldenshluger:2000:AEL

[GP00] Alexander Goldenshluger and Sergei V. Pereverzev. Adaptive estimation of linear functionals in Hilbert scales from indirect white noise observations. Probability Theory and Related Fields, 118(2):169–186, October 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s440-000-8013-3.

Gayraud:2005:AMT

[GP05] Ghislaine Gayraud and Christophe Pouet. Adaptive minimax testing in the discrete regression scheme. Probability Theory and Related Fields, 133(4):531–558, December 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
005-0445-4.

Gubinelli:2020:IGS

[GP20] Massimiliano Gubinelli and Nicolas Perkowski. The infinitesimal generator of the stochastic Burgers equation. Probability Theory and Related Fields, 178(3–4):1067–1124, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00996-5.

Gwynne:2022:GMB

[GPS22] Ewain Gwynne, Joshua Pfeffer, and Scott Sheffield. Geodesics and metric ball boundaries in Liouville quantum gravity. Probability Theory and Related Fields, 182(3–4):905–954, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01112-5.

Greven:2009:CDR

[GPW09] Andreas Greven, Peter Pfaffelhuber, and Anita Winter. Convergence in distribution of random metric measure spaces (Λcoalescent measure trees). Probability Theory and Related Fields, 145(1–2):285–322, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-008-0169-3.

Greven:2013:TVR

[GPW13] Andreas Greven, Peter Pfaffelhuber, and Anita Winter. Treevalued resampling dynamics martingale problems and applications. Probability Theory and Related Fields, 155(3–4):789–838, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-012-0413-8.

Gong:1988:KDT

[GQZ88] Guanglu Gong, Minping Qian, and Zhongxin Zhao. Killed diffusions and their conditioning. Probability Theory and Related Fields, 80(1):151–167, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00348757.

Gabetta:2010:CLT

[GR10] Ester Gabetta and Eugenio Regazzini. Central limit theorems for solutions of the Kac equation: speed of approach to equilibrium in weak metrics. Probability Theory and Related Fields, 146(3–4):451–480, March 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0196-0.

Gorin:2019:RSN

[GR19] Vadim Gorin and Mustazee Rahman. Random sorting networks: local statistics via random matrix laws. Probability Theory and Related Fields, 175(1–2):45–96, October 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0886-1.

Graf:1987:SSS

[Gra87] Siegfried Graf. Statistically self-similar fractals. Probability Theory and Related Fields, 74(3):357–392, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00699096.

Graham:1995:HPC

[Gra95] Carl Graham. Homogenization and propagation of chaos to a nonlinear diffusion with sticky reflection. Probability Theory and Related Fields, 101(3):291–302, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01200497.

Graham:2005:FCL

[Gra05] Carl Graham. Functional central limit theorems for a large network in which customers join the shortest of several queues. Probability Theory and Related Fields, 131(1):97–120, January 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-004-0372-9.

Georgiou:2017:GCI

[GRAS17a] Nicos Georgiou, Firas Rassoul-Agha, and Timo Seppäläinen. Geodesics and the competition interface for the corner growth model. Probability Theory and Related Fields, 169(1–2):223– 255, October 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-016-0734-0.

Georgiou:2017:SCB

[GRAS17b] Nicos Georgiou, Firas Rassoul-Agha, and Timo Seppäläinen. Stationary cocycles and Busemann functions for the corner growth model. Probability Theory and Related Fields, 169(1– 2):177–222, October 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-016-0729-x.

Greven:1987:CMCa

[Gre87a] Andreas Greven. Couplings of Markov chains by randomized stopping times. Probability Theory and Related Fields, 75(2): 195–212, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00354033.

Greven:1987:CMCb

[Gre87b] Andreas Greven. Couplings of Markov chains by randomized stopping times. Probability Theory and Related Fields, 75 (3):431–458, July 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00318710.

Greven:1990:SER

[Gre90] Andreas Greven. Symmetric exclusion on random sets and a related problem for random walks in random environment. Probability Theory and Related Fields, 85(3):307–364, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01193942.

Greven:1991:PTC

[Gre91] Andreas Greven. A phase transition for the coupled branching process. Probability Theory and Related Fields, 87(4):417–458, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01304274.

Grill:1987:ERT

[Gri87a] Karl Grill. Erdős–Révész type bounds for the length of the longest run from a stationary mixing sequence. Probability Theory and Related Fields, 75(1):77–85, May 1987. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00320082.

Grill:1987:RCS

[Gri87b] Karl Grill. On the rate of convergence in Strassen’s law of the iterated logarithm. Probability Theory and Related Fields, 74 (4):583–589, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00363517.

Griffin:1988:IEL

[Gri88a] Philip S. Griffin. The influence of extremes on the law of the iterated logarithm. Probability Theory and Related Fields, 77(2): 241–270, February 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00334040.

Griffin:1988:NCL

[Gri88b] Phillip S. Griffin. Non-classical law of the iterated logarithm behaviour for trimmed sums. Probability Theory and Related Fields, 78(2):293–319, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00322025.

Grill:1991:LIR

[Gri91a] Karl Grill. A lim inf result in Strassen’s law of the iterated logarithm. Probability Theory and Related Fields, 89(2):149–157, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01366903.

Grill:1991:UCI

[Gri91b] Karl Grill. Upper classes for the increments of the fractional Wiener process. Probability Theory and Related Fields, 87(4): 411–416, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01304273.

Grill:1995:LIR

[Gri95] K. Grill. A lim inf result for the increments of the Wiener process under the L_{2}-norm. Probability Theory and Related Fields, 102 (4):567–573, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01198850.

Grimmett:2021:HKW

[Gri21] Geoffrey R. Grimmett. Harry Kesten’s work in probability theory. Probability Theory and Related Fields, 181(1-3):17–55, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01046-4.

Groeneboom:1988:LTC

[Gro88] Piet Groeneboom. Limit theorems for convex hulls. Probability Theory and Related Fields, 79(3):327–368, October 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00342231.

Groeneboom:1989:BMP

[Gro89] Piet Groeneboom. Brownian motion with a parabolic drift and Airy functions. Probability Theory and Related Fields, 81(1): 79–109, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00343738.

Glover:1994:GTC

[GRS94] Joseph Glover, Murali Rao, and Renming Song. The gauge theorem for a class of additive functionals of zero energy. Probability Theory and Related Fields, 98(4):567, ???? 1994. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01192838.

Gouezel:2024:MDB

[GRS24] Sébastien Gouëzel, Jérôme Rousseau, and Manuel Stadlbauer. Minimal distance between random orbits. Probability Theory and Related Fields, 189(3–4):811–847, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01283-3.

Gozlan:2014:DCE

[GRST14] Nathael Gozlan, Cyril Roberto, Paul-Marie Samson, and Prasad Tetali. Displacement convexity of entropy and related inequalities on graphs. Probability Theory and Related Fields, 160(1– 2):47–94, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0523-y.

Gordina:2020:OUP

[GRT20] Maria Gordina, Michael Röckner, and Alexander Teplyaev. Ornstein–Uhlenbeck processes with singular drifts: integral estimates and Girsanov densities. Probability Theory and Related Fields, 178(3–4):861–891, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-020-00991-
w.

Grubel:1986:HRM

[Grü86] Rudolf Grübel. On harmonic renewal measures. Probability Theory and Related Fields, 71(3):393–404, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01000213.

Grunwald:1996:SRG

[Gru96] M. Grunwald. Sanov results for Glauber spin-glass dynamics. Probability Theory and Related Fields, 106(2):187–232, October 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050062.

Gruet:1998:LHB

[Gru98] Jean-Claude Gruet. On the length of the homotopic Brownian word in the thrice punctured sphere. Probability Theory and Related Fields, 111(4):489–516, August 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050175.

Gorostiza:1992:PCM

[GRW92] L. G. Gorostiza, S. Roelly, and A. Wakolbinger. Persistence of critical multitype particle and measure branching processes.

Probability Theory and Related Fields, 92(3):313–335, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01300559.

Getoor:1986:CTD

[GS86] R. K. Getoor and J. Steffens. Capacity theory without duality. Probability Theory and Related Fields, 73(3):415–445, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00776241.

Galves:1989:AFM

[GS89] Antonio Galves and Rinaldo Schinazi. Approximations finies de la mesure invariante de processus de contact sur-critique vu par la première particule. (French) [Approximations of the finite invariant measure of supercritical contact processes seen by the first particle]. Probability Theory and Related Fields, 83 (4):435–445, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01845698.

Giraitis:1990:CLT

[GS90] L. Giraitis and D. Surgailis. A central limit theorem for quadratic forms in strongly dependent linear variables and its application to asymptotical normality of Whittle’s estimate. Probability Theory and Related Fields, 86(1):87–104, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01207515.

Greschonig:2003:GRS

[GS03] Gernot Greschonig and Klaus Schmidt. Growth and recurrence of stationary random walks. Probability Theory and Related Fields, 125(2):266–270, February 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400200237.

Goldenshluger:2004:SMP

[GS04] A. Goldenshluger and V. Spokoiny. On the shape-from-moments problem and recovering edges from noisy Radon data. Probability Theory and Related Fields, 128(1):123–140, January 2004.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-003-0303-1.

Ginovyan:2007:LTT

[GS07] M. S. Ginovyan and A. A. Sahakyan. Limit theorems for Toeplitz quadratic functionals of continuous-time stationary processes. Probability Theory and Related Fields, 138(3–4): 551–579, July 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0037-y.

Gorin:2015:MDB

[GS15a] Vadim Gorin and Mykhaylo Shkolnikov. Multilevel Dyson Brownian motions via Jack polynomials. Probability Theory and Related Fields, 163(3–4):413–463, December 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0596-2.

Guntuboyina:2015:GRB

[GS15b] Adityanand Guntuboyina and Bodhisattva Sen. Global risk bounds and adaptation in univariate convex regression. Probability Theory and Related Fields, 163(1–2):379–411, October 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-014-0595-3.

Ganguly:2020:GSH

[GS20] Shirshendu Ganguly and Sourav Sarkar. Ground states and hyperuniformity of the hierarchical Coulomb gas in all dimensions. Probability Theory and Related Fields, 177(3–4):621–675, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-019-00955-9.

Gass:2024:NCP

[GS24] Louis Gass and Michele Stecconi. The number of critical points of a Gaussian field: finiteness of moments. Probability Theory and Related Fields, 190(3-4):1167–1197, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01273-5.

Gordon:1986:EVT

[GSW86] Louis Gordon, Mark F. Schilling, and Michael S. Waterman. An extreme value theory for long head runs. Probability Theory and Related Fields, 72(2):279–287, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00699107.

Giraitis:1997:LTB

[GT97] Liudas Giraitis and Murad S. Taqqu. Limit theorems for bivariate Appell polynomials. part i: Central limit theorems. Probability Theory and Related Fields, 107(3):359–381, March 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050089.

Gaudron:1998:FEM

[GT98a] Isabelle Gaudron and Alain Trouvé. Fluctuations of empirical means at low temperature for finite Markov chains with rare transitions in the general case. Probability Theory and Related Fields, 111(2):215–251, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050167.

Guerre:1998:EAM

[GT98b] Emmanuel Guerre and Alexander B. Tsybakov. Exact asymptotic minimax constants for the estimation of analytical functions in L_{p}. Probability Theory and Related Fields, 112(1): 33–51, September 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050182.

Gotze:2003:RCS

[GT03] F. Götze and A. Tikhomirov. Rate of convergence to the semicircular law. Probability Theory and Related Fields, 127(2): 228–276, October 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0285-z.

Giacomin:2005:EPD

[GT05a] Giambattista Giacomin and Fabio Lucio Toninelli. Estimates on path delocalization for copolymers at selective interfaces. Probability Theory and Related Fields, 133(4):464–482, December 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-005-0439-2.

Gurevich:2005:MAH

[GT05b] B. M. Gurevich and A. A. Tempelman. Markov approximation of homogeneous lattice random fields. Probability Theory and Related Fields, 131(4):519–527, April 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0383-6.

Gotze:2016:OBC

[GT16] F. Götze and A. Tikhomirov. Optimal bounds for convergence of expected spectral distributions to the semi-circular law. Probability Theory and Related Fields, 165(1–2):163–233, June 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-015-0629-5.

Gigli:2020:BBD

[GT20] Nicola Gigli and Luca Tamanini. Benamou–Brenier and duality formulas for the entropic cost on RCD^{∗}(K, N) spaces. Probability Theory and Related Fields, 176(1–2):1–34, February 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-019-00909-1.

Goldman:2024:OTM

[GT24] Michael Goldman and Dario Trevisan. Optimal transport methods for combinatorial optimization over two random point sets. Probability Theory and Related Fields, 188(3–4):1315– 1384, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01245-1.

Glanzel:1984:CTM

[GTS84] W. Glänzel, A. Teles, and A. Schubert. Characterization by truncated moments and its application to Pearson-type distributions. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 66(2):173–183, ???? 1984. CODEN ZWVGAA.

ISSN 0044-3719. URL http://link.springer.com/article/

10.1007/BF00531527. See correction [GTS87].

Glanzel:1987:CCT

[GTS87] W. Glänzel, A. Telcs, and A. Schubert. Correction to “Characterization by truncated moments and its application to Pearsontype distributions”. Probability Theory and Related Fields, 74 (2):317, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00569996. See [GTS84].

Giraitis:1998:LTB

[GTT98] Liudas Giraitis, Murad S. Taqqu, and Norma Terrin. Limit theorems for bivariate Appell polynomials. Part II: Non-central limit theorems. Probability Theory and Related Fields, 110(3): 333–367, March 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050151.

Gu:2017:HOC

[Gu17] Yu Gu. High order correctors and two-scale expansions in stochastic homogenization. Probability Theory and Related Fields, 169(3–4):1221–1259, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0750-0.

Guionnet:1997:AQP

[Gui97] A. Guionnet. Averaged and quenched propagation of chaos for spin glass dynamics. Probability Theory and Related Fields, 109 (2):183–215, October 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050130.

Guyon:1987:VCG

[Guy87] Xavier Guyon. Variations de champs gaussiens stationnaires: application à l’identification. (French) [Variations of Gaussian stationary fields: application to identification]. Probability Theory and Related Fields, 75(2):179–193, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00354032.

Guedon:2016:CDS

[GV16] Olivier Guédon and Roman Vershynin. Community detection in sparse networks via Grothendieck’s inequality. Probability Theory and Related Fields, 165(3–4):1025–1049, August 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0659-z;   http://link.springer.com/
article/10.1007/s00440-015-0659-z.

Graversen:1986:SSM

[GVA86] S. E. Graversen and J. Vuolle-Apiala. α-self-similar Markov processes. Probability Theory and Related Fields, 71(1):149– 158, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00366277.

Goderis:1989:NCC

[GVV89] D. Goderis, A. Verbeure, and P. Vets. Non-commutative central limits. Probability Theory and Related Fields, 82(4):527–544, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00341282.

Griffin:1993:DAC

[GVV93] P. S. Griffin, G. C. Verchota, and A. L. Vogel. Distoriton of area and conditioned Brownian motion. Probability Theory and Related Fields, 96(3):385–413, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01292679.

Guerra:2022:PBC

[GVV22] Enrique Guerra, Glauco Valle, and Maria Eulália Vares. Polynomial ballisticity conditions and invariance principle for random walks in strong mixing environments. Probability Theory and Related Fields, 182(3–4):685–750, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01106-9.

Gerl:1986:LLH

[GW86] Peter Gerl and Wolfgang Woess. Local limits and harmonic functions for nonisotropic random walks on free groups. Probability Theory and Related Fields, 71(3):341–355, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01000210.

Gao:2004:AND

[GW04] Zhicheng Gao and Nicholas C. Wormald. Asymptotic normality determined by high moments, and submap counts of random maps. Probability Theory and Related Fields, 130(3):368– 376, November 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0356-9.

Gallardo:2005:SNE

[GY05] Léonard Gallardo and Marc Yor. Some new examples of Markov processes which enjoy the time-inversion property. Probability Theory and Related Fields, 132(1):150–162, May 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0399-y.

Gyongy:1986:MOD

[Gyö86] I. Gyöngy. Mimicking the one-dimensional marginal distributions of processes having an Itô differential. Probability Theory and Related Fields, 71(4):501–516, ???? 1986. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00699039.

Georgii:1993:LDM

[GZ93] Hans-Otto Georgii and Hans Zessin. Large deviations and the maximum entropy principle for marked point random fields. Probability Theory and Related Fields, 96(2):177–204, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01192132.

Guionnet:2012:SCS

[GZ12a] Alice Guionnet and Ofer Zeitouni. Support convergence in the single ring theorem. Probability Theory and Related Fields, 154(3–4):661–675, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-011-0380-5.

Guo:2012:QIP

[GZ12b] Xiaoqin Guo and Ofer Zeitouni. Quenched invariance principle for random walks in balanced random environment. Probability Theory and Related Fields, 152(1–2):207–230, February 2012.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-010-0320-9.

Gorin:2018:IAL

[GZ18] Vadim Gorin and Lingfu Zhang. Interlacing adjacent levels of β-Jacobi corners processes. Probability Theory and Related Fields, 172(3–4):915–981, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0823-8.

Habermann:2018:STF

[Hab18] Karen Habermann. Small-time fluctuations for sub-Riemannian diffusion loops. Probability Theory and Related Fields, 171 (3–4):617–652, August 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0788-7;
 http://link.springer.com/content/pdf/10.1007/s00440-
017-0788-7.pdf.

Haggstrom:1995:RBF

[Häg95] Olle Häggström. On the relation between finite range potentials and subshifts of finite type. Probability Theory and Related Fields, 101(4):469–478, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01202780.

Haggstrom:1996:RCM

[Häg96] Olle Häggström. The random-cluster model on a homogeneous tree. Probability Theory and Related Fields, 104(2):231–253, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01247839.

Haggstrom:2003:UIR

[Häg03] Olle Häggström. Uniqueness of infinite rigid components in percolation models: the case of nonplanar lattices. Probability Theory and Related Fields, 127(4):513–534, December 2003. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0290-2.

Haggstrom:2005:CLT

[Häg05] Olle Häggström. On the central limit theorem for geometrically ergodic Markov chains. Probability Theory and Related Fields, 132(1):74–82, May 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0390-7.

Hairer:2002:EMP

[Hai02] M. Hairer. Exponential mixing properties of stochastic PDEs through asymptotic coupling. Probability Theory and Related Fields, 124(3):345–380, November 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400200216.

Hairer:2012:SPS

[Hai12] Martin Hairer. Singular perturbations to semilinear stochastic heat equations. Probability Theory and Related Fields, 152(1–2): 265–297, February 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0322-7.

Hall:1988:EDW

[Hal88] Peter Hall. Estimating the direction in which a data set is most interesting. Probability Theory and Related Fields, 80(1): 51–77, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00348752.

Hall:1989:UPB

[Hal89] Peter Hall. Unusual properties of bootstrap confidence intervals in regression problems. Probability Theory and Related Fields, 81(2):247–273, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319554.

Hall:1990:AIC

[Hal90a] Peter Hall. Akaike’s information criterion and Kullback–Leibler loss for histogram density estimation. Probability Theory and Related Fields, 85(4):449–467, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01203164.

Hall:1990:EDE

[Hal90b] Peter Hall. Effects of design and error on normal convergence rates in regression problems. Probability Theory and Related Fields, 85(3):283–305, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01193941.

Hall:1990:PBB

[Hal90c] Peter Hall. Performance of balanced bootstrap resampling in distribution function and quantile problems. Probability Theory and Related Fields, 85(2):239–260, ???? 1990. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01277983.

Hall:1991:CRS

[Hal91] Peter Hall. On convergence rates of suprema. Probability Theory and Related Fields, 89(4):447–455, ???? 1991. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01199788. See erratum [Ano98a].

Hambly:1992:BMH

[Ham92] B. M. Hambly. Brownian motion on a homogeneous random fractal. Probability Theory and Related Fields, 94(1):1–38, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01222507.

Hambly:2000:AEC

[Ham00] B. M. Hambly. On the asymptotics of the eigenvalue counting function for random recursive Sierpinski gaskets. Probability Theory and Related Fields, 117(2):221–247, June 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050005.

Hammond:2015:SPT

[Ham15] Alan Hammond. Sharp phase transition in the random stirring model on trees. Probability Theory and Related Fields, 161(3– 4):429–448, April 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0543-7.

Handa:2002:QIR

[Han02] Kenji Handa. Quasi-invariance and reversibility in the Fleming– Viot process. Probability Theory and Related Fields, 122(4): 545–566, April 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400100178.

Hansen:2012:IPD

[Han12] Wolfhard Hansen. 3G-inequality for planar domains. Probability Theory and Related Fields, 152(1–2):357–366, February 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-010-0335-2.

Hara:1990:MFC

[Har90] Takashi Hara. Mean-field critical behaviour for correlation length for percolation in high dimensions. Probability Theory and Related Fields, 86(3):337–385, ???? 1990. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01208256.

Harge:2004:CLC

[Har04] Gilles Hargé. A convex/log-concave correlation inequality for Gaussian measure and an application to abstract Wiener spaces. Probability Theory and Related Fields, 130(3):415–440, November 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-004-0365-8.

Hassenforder:1988:LCB

[Has88] Claudie Hassenforder. Les lois de Cauchy sur les bouts de l’arbre homogène. (French) [The Cauchy laws on the tips of the homogeneous tree]. Probability Theory and Related Fields, 78(4): 603–616, August 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00353878.

Hasegawa:1989:BMI

[Has89] Yoshihei Hasegawa. Brownian motions on infinite dimensional quadric hypersurfaces. Probability Theory and Related Fields, 80(3):347–364, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01794428.

Hausenblas:2007:SDP

[Hau07] Erika Hausenblas. SPDEs driven by Poisson random measure with non Lipschitz coefficients: existence results. Probability Theory and Related Fields, 137(1–2):161–200, January 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-006-0501-8.

Hawkes:1998:ECR

[Haw98] John Hawkes. Exact capacity results for stable processes. Probability Theory and Related Fields, 112(1):1–11, September 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050180.

Huang:2021:MMC

[HD21] Xiangying Huang and Rick Durrett. Motion by mean curvature in interacting particle systems. Probability Theory and Related Fields, 181(1-3):489–532, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01082-
0.

Heckman:1986:RST

[Hec86] Nancy E. Heckman. Repeated significance tests with biased coin allocation schemes. Probability Theory and Related Fields, 73(4):627–635, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00324857.

Hennion:2007:QCA

[Hen07] Hubert Hennion. Quasi-compactness and absolutely continuous kernels. Probability Theory and Related Fields, 139(3–4): 451–471, November 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0048-8.

Hall:1990:CRD

[HH90] Peter Hall and Jeffrey D. Hart. Convergence rates in density estimation for data from infinite-order moving average processes. Probability Theory and Related Fields, 87(2):253–274, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01198432.

Hattori:1991:SAP

[HH91] Kumiko Hattori and Tetsuya Hattori. Self-avoiding process on the Sierpinski gasket. Probability Theory and Related Fields, 88 (4):405–428, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192550.

Hamadene:2005:BTR

[HH05] S. Hamadène and M. Hassani. BSDEs with two reflecting barriers: the general result. Probability Theory and Related Fields, 132(2):237–264, June 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0395-2.

Hammond:2019:SAS

[HH19] Alan Hammond and Tyler Helmuth. Self-attracting selfavoiding walk. Probability Theory and Related Fields, 175(3–4): 677–719, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-00898-7.

Hall:2023:BMF

[HH23] Brian C. Hall and Ching-Wei Ho. The Brown measure of a family of free multiplicative Brownian motions. Probability Theory and Related Fields, 186(3–4):1081–1166, August 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01166-5.

Hambly:2002:SRW

[HHH02] B. M. Hambly, Kumiko Hattori, and Tetsuya Hattori. Selfrepelling walk on the Sierpiński gasket. Probability Theory and Related Fields, 124(1):1–25, September 2002. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100192.

Hattori:1990:SAP

[HHK90] Kumiko Hattori, Tetsuya Hattori, and Shigeo Kusuoka. Selfavoiding paths on the pre-Sierpinski gasket. Probability Theory and Related Fields, 84(1):1–26, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01288555.

Hsing:1988:EPP

[HHL88] T. Hsing, J. Hüsler, and M. R. Leadbetter. On the exceedance point process for a stationary sequence. Probability Theory and Related Fields, 78(1):97–112, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00718038.

Herold:2021:DCL

[HHT21] Felix Herold, Daniel Hug, and Christoph Thäle. Does a central limit theorem hold for the k-skeleton of Poisson hyperplanes in hyperbolic space? Probability Theory and Related Fields, 179 (3–4):889–968, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01032-w.

Hattori:1994:AOD

[HHW94] Kumiko Hattori, Tetsuya Hattori, and Hiroshi Watanabe. Asymptotically one-dimensional diffusions on the Sierpinski gasket and the abc-gaskets. Probability Theory and Related Fields, 100(1):85–116, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01204955.

Hasebe:2018:LTB

[HHW18] Takahiro Hasebe, Hao-Wei Huang, and Jiun-Chau Wang. Limit theorems in bi-free probability theory. Probability Theory and Related Fields, 172(3–4):1081–1119, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0825-6.

Hasminskii:1986:AEN

[HI86] R. Z. Hasminskii and I. A. Ibragimov. Asymptotically efficient nonparametric estimation of functionals of a spectral density function. Probability Theory and Related Fields, 73(3):447–461, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00776242.

Higuchi:1986:PPV

[Hig86] Yasunari Higuchi. A percolation problem for ±1-valued strongly mixing random fields on Z_{∗}^{d}. Probability Theory and Related Fields, 73(4):597–611, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00324855.

Higuchi:1991:LSR

[Hig91] Yasunari Higuchi. Level set representation for the Gibbs states of the ferromagnetic Ising model. Probability Theory and Related Fields, 90(2):203–221, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192162.

Higuchi:1993:STT

[Hig93] Yasunari Higuchi. A sharp transition for the two-dimensional Ising percolation. Probability Theory and Related Fields, 97 (4):489–514, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192961.

Hildebrand:1987:LDD

[Hil87a] Adolf Hildebrand. On the limit distribution of discrete random variables. Probability Theory and Related Fields, 75(1):67–76, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00320081.

Hill:1987:SPI

[Hil87b] Theodore P. Hill. A sharp partitioning-inequality for non-atomic probability measures based on the mass of the infimum of the measures. Probability Theory and Related Fields, 75(1):143–147, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00320087.

Hildebrand:1994:RWS

[Hil94] Martin Hildebrand. Random walks supported on random points of Z/nZ. Probability Theory and Related Fields, 100(2):191–203, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01199265.

Hino:2005:SEM

[Hin05] Masanori Hino. On singularity of energy measures on self-similar sets. Probability Theory and Related Fields, 132(2):265–290, June 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-004-0396-1.

Hino:2013:UEM

[Hin13] Masanori Hino. Upper estimate of martingale dimension for selfsimilar fractals. Probability Theory and Related Fields, 156(3– 4):739–793, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0442-3.

Hitczenko:1988:CMT

[Hit88] Pawel Hitczenko. Comparison of moments for tangent sequences of random variables. Probability Theory and Related Fields, 78 (2):223–230, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00322019.

Hitczenko:1990:UBN

[Hit90] Pawel Hitczenko. Upper bounds for the L_{p}-norms of martingales. Probability Theory and Related Fields, 86(2):225–238, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01474643.

Hopfner:1990:LAN

[HJL90] Reinhard Höpfner, Jean Jacod, and Lucia Ladelli. Local asymptotic normality and mixed normality for Markov statistical models. Probability Theory and Related Fields, 86(1):105–129, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01207516.

Han:2020:ENG

[HJM20] Yanjun Han, Jiantao Jiao, and Rajarshi Mukherjee. On estimation of L_{r}-norms in Gaussian white noise models. Probability Theory and Related Fields, 177(3–4):1243–1294, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00982-x.

Helmers:1988:GCP

[HJS88] R. Helmers, P. Janssen, and R. Serfling. Glivenko-Cantelli properties of some generalized empirical DF’s and strong convergence of generalized L-statistics. Probability Theory and Related Fields, 79(1):75–93, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00319105.

Holroyd:2022:MMP

[HJW22] Alexander E. Holroyd, Svante Janson, and Johan Wästlund. Minimal matchings of point processes. Probability Theory and Related Fields, 184(1–2):571–611, October 2022. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-022-01151-y.

Hattori:1992:EMS

[HK92] Tetsuya Hattori and Shigeo Kusuoka. The exponent for the mean square displacement of self-avoiding random walk on the Sierpinski gasket. Probability Theory and Related Fields, 93 (3):273–284, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193052.

Harris:1997:RCT

[HK97] Matthew Harris and Michael Keane. Random coin tossing. Probability Theory and Related Fields, 109(1):27–37, September 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050123.

Hamana:2001:LDR

[HK01] Yuji Hamana and Harry Kesten. A large-deviation result for the range of random walk and for the Wiener sausage. Probability Theory and Related Fields, 120(2):183–208, June 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008780.

Hambly:2003:DPF

[HK03] B. M. Hambly and T. Kumagai. Diffusion processes on fractal fields: heat kernel estimates and large deviations. Probability Theory and Related Fields, 127(3):305–352, November 2003.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-003-0284-0.

Hammond:2011:MLM

[HK11] Alan Hammond and Richard Kenyon. Monotone loop models and rational resonance. Probability Theory and Related Fields, 150(3–4):613–633, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0285-8.

He:2020:MED

[HK20] Yukun He and Antti Knowles. Mesoscopic eigenvalue density correlations of Wigner matrices. Probability Theory and Related Fields, 177(1–2):147–216, June 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-019-00946-w.

He:2021:FEE

[HK21] Yukun He and Antti Knowles. Fluctuations of extreme eigenvalues of sparse Erdős–Rényi graphs. Probability Theory and Related Fields, 180(3–4):985–1056, August 2021. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-021-01054-4.

Hoshino:2023:SQA

[HKK23] Masato Hoshino, Hiroshi Kawabi, and Seiichiro Kusuoka. Stochastic quantization associated with the exp(Φ)_{2}-quantum field model driven by space-time white noise on the torus in the full L^{1} regime. Probability Theory and Related Fields, 185(1–2): 391–447, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01126-z.

He:2018:ISC

[HKR18] Yukun He, Antti Knowles, and Ron Rosenthal. Isotropic selfconsistent equations for mean-field random matrices. Probability Theory and Related Fields, 171(1–2):203–249, June 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0776-y.

Hueter:2000:ABR

[HL00] Irene Hueter and Steven P. Lalley. Anisotropic branching random walks on homogeneous trees. Probability Theory and Related Fields, 116(1):57–88, January 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/PL00008723.

Harper:2018:OWC

[HL18] Adam J. Harper and Youness Lamzouri. Orderings of weakly correlated random variables, and prime number races with many contestants. Probability Theory and Related Fields, 170(3–4): 961–1010, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0800-2.

Huang:2019:RMC

[HL19] Jiaoyang Huang and Benjamin Landon. Rigidity and a mesoscopic central limit theorem for Dyson Brownian motion for general β and potentials. Probability Theory and Related Fields, 175(1–2):209–253, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-018-0889-y.

Healey:2021:SRS

[HL21] Vivian Olsiewski Healey and Gregory F. Lawler. N-sided radial Schramm–Loewner evolution. Probability Theory and Related Fields, 181(1-3):451–488, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01033-
9.

Hong:2023:CLT

[HLLS23] Wei Hong, Shihu Li, Wei Liu, and Xiaobin Sun. Central limit type theorem and large deviation principle for multi-scale McKean–Vlasov SDEs. Probability Theory and Related Fields, 187(1–2):133–201, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01214-8.

Hu:2013:HCS

[HLN13] Yaozhong Hu, Fei Lu, and David Nualart. Hölder continuity of the solutions for a class of nonlinear SPDE’s arising from one dimensional superprocesses. Probability Theory and Related Fields, 156(1–2):27–49, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0419-2.

Holden:1993:SBV

[HLØ^{+}93] Helge Holden, Tom Lindstrøm, Bernt Øksendal, Jan Ubøe, and Tu-Sheng Zhang. Stochastic boundary value problems: a white noise functional approach. Probability Theory and Related Fields, 95(3):391–419, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192171.

Hayakawa:2023:EPG

[HLO23] Satoshi Hayakawa, Terry Lyons, and Harald Oberhauser. Estimating the probability that a given vector is in the convex hull of a random sample. Probability Theory and Related Fields, 185 (3–4):705–746, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01186-1.

Hofmanova:2024:GEN

[HLP24] Martina Hofmanová, Theresa Lange, and Umberto Pappalettera. Global existence and non-uniqueness of 3D Euler equations perturbed by transport noise. Probability Theory and Related Fields, 188(3–4):1183–1255, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-023-01233-
5.

Hairer:2023:ACE

[HLR23] Martin Hairer, Khoa Lê, and Tommaso Rosati. The Allen– Cahn equation with generic initial datum. Probability Theory and Related Fields, 186(3–4):957–998, August 2023. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-023-01198-5.

Hotz:2024:CLT

[HLW24] Thomas Hotz, Huiling Le, and Andrew T. A. Wood. Central limit theorem for intrinsic Fréchet means in smooth compact Riemannian manifolds. Probability Theory and Related Fields, 189(3–4):1219–1246, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-024-01291-3.

Hall:1987:EWL

[HM87] Peter Hall and James Stephen Marron. Extent to which leastsquares cross-validation minimises integrated square error in nonparametric density estimation. Probability Theory and Related Fields, 74(4):567–581, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00363516.

Hall:1988:VWW

[HM88a] Peter Hall and J. S. Marron. Variable window width kernel estimates of probability densities. Probability Theory and Related Fields, 80(1):37–49, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00348751.

Hall:1988:ECR

[HM88b] Peter Hall and Michael A. Martin. Exact convergence rate of bootstrap quantile variance estimator. Probability Theory and Related Fields, 80(2):261–268, December 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00356105.

Hall:1991:LBB

[HM91] Peter Hall and J. S. Marron. Lower bounds for bandwidth selection in density estimation. Probability Theory and Related Fields, 90(2):149–173, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192160.

Hall:1992:VWW

[HM92] Peter Hall and J. S. Marron. Variable window width kernel estimates of probability densities. Probability Theory and Related Fields, 91(1):133, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01194494.

Hall:1997:RSP

[HM97] Peter Hall and J. Stephen Marron. On the role of the shrinkage parameter in local linear smoothing. Probability Theory and Related Fields, 108(4):495–516, August 1997. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050118.

Hambly:2007:HTL

[HM07] Ben Hambly and James B. Martin. Heavy tails in last-passage percolation. Probability Theory and Related Fields, 137(1–2): 227–275, January 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0019-0.

Huang:2024:AGB

[HM24] Jiaoyang Huang and Colin McSwiggen. Asymptotics of generalized Bessel functions and weight multiplicities via large deviations of radial Dunkl processes. Probability Theory and Related Fields, 190(3-4):941–1006, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-024-01282-
4.

Holroyd:2019:PGP

[HMM19] Alexander E. Holroyd, Irène Marcovici, and James B. Martin. Percolation games, probabilistic cellular automata, and the hard-core model. Probability Theory and Related Fields, 174(3–4):1187–1217, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-018-0881-6.

Huesmann:2023:TNS

[HMO23] Martin Huesmann, Francesco Mattesini, and Felix Otto. There is no stationary cyclically monotone Poisson matching in 2d. Probability Theory and Related Fields, 187(3–4):629–656, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01225-5.

Hall:1992:SCV

[HMP92] Peter Hall, J. S. Marron, and Byeong U. Park. Smoothed crossvalidation. Probability Theory and Related Fields, 92(1):1–20, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01205233.

Hairer:2011:ACG

[HMS11] M. Hairer, J. C. Mattingly, and M. Scheutzow. Asymptotic coupling and a general form of Harris’ theorem with applications to stochastic delay equations. Probability Theory and Related Fields, 149(1–2):223–259, February 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-009-0250-6.

Hartarsky:2020:UCK

[HMT20] Ivailo Hartarsky, Laure Marêché, and Cristina Toninelli. Universality for critical KCM: infinite number of stable directions. Probability Theory and Related Fields, 178(1–2):289–326, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-020-00976-9.

Hartarsky:2023:STF

[HMT23] Ivailo Hartarsky, Fabio Martinelli, and Cristina Toninelli. Sharp threshold for the FA-2f kinetically constrained model. Probability Theory and Related Fields, 185(3–4):993–1037, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01169-2.

Hu:2002:SLD

[HMY02] Ying Hu, Jin Ma, and Jiongmin Yong. On semi-linear degenerate backward stochastic partial differential equations. Probability Theory and Related Fields, 123(3):381–411, July 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100193.

Hazla:2020:PID

[HMZ20] Jan Hazła, Elchanan Mossel, and Guangqu Zheng. The probability of intransitivity in dice and close elections. Probability Theory and Related Fields, 178(3–4):951–1009, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00994-7.

Howard:1997:EMF

[HN97] C. Douglas Howard and Charles M. Newman. Euclidean models of first-passage percolation. Probability Theory and Related Fields, 108(2):153–170, June 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050105.

Hu:2009:SHE

[HN09] Yaozhong Hu and David Nualart. Stochastic heat equation driven by fractional noise and local time. Probability Theory and Related Fields, 143(1–2):285–328, January 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0127-5.

Hedenmalm:2014:GFF

[HN14] Haakan Hedenmalm and Pekka J. Nieminen. The Gaussian free field and Hadamard’s variational formula. Probability Theory and Related Fields, 159(1–2):61–73, June 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0501-4.

Hutchcroft:2017:ITU

[HN17] Tom Hutchcroft and Asaf Nachmias. Indistinguishability of trees in uniform spanning forests. Probability Theory and Related Fields, 168(1–2):113–152, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Hiwatashi:1999:CSL

[HNY99] Osamu Hiwatashi, Masaru Nagisa, and Hiroaki Yoshida. The characterizations of a semicircle law by the certain freeness in a C *-probability space. Probability Theory and Related Fields, 113(1):115–133, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050204.

Hughes:2008:AMT

[HNY08] C. P. Hughes, A. Nikeghbali, and M. Yor. An arithmetic model for the total disorder process. Probability Theory and Related Fields, 141(1–2):47–59, May 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0079-9.

Hofbauer:1986:PID

[Hof86] Franz Hofbauer. Piecewise invertible dynamical systems. Probability Theory and Related Fields, 72(3):359–386, June 1986.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00334191.

Holevo:1996:DSE

[Hol96] A. S. Holevo. On dissipative stochastic equations in a Hilbert space. Probability Theory and Related Fields, 104(4):483–500, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01198163.

Holroyd:2003:SMT

[Hol03] Alexander E. Holroyd. Sharp metastability threshold for two-dimensional bootstrap percolation. Probability Theory and Related Fields, 125(2):195–224, February 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
002-0239-x.

Holroyd:2011:GPP

[Hol11] Alexander E. Holroyd. Geometric properties of Poisson matchings. Probability Theory and Related Fields, 150(3–4):511–527, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-010-0282-y.

Hoover:1991:CDS

[Hoo91] D. N. Hoover. Convergence in distribution and Skorokhod convergence for the general theory of processes. Probability Theory and Related Fields, 89(3):239–259, ???? 1991. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198786.

Hopfner:1988:AIC

[Höp88] R. Höpfner. Asymptotic inference for continuous-time Markov chains. Probability Theory and Related Fields, 77(4):537–550, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00959616.

Hopfner:1993:SMS

[Höp93] Reinhard Höpfner. On statistics of Markov step processes: Representation of log-likelihood ratio processes in filtered local models. Probability Theory and Related Fields, 94(3):375–398, ????

1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01199249.

Horvath:1990:SLL

[Hor90] Lajos Horváth. Strong laws and limit theorems for local time of Markov processes. Probability Theory and Related Fields, 85 (3):413–424, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193946.

Hora:2000:GSD

[Hor00] Akihito Hora. Gibbs state on a distance-regular graph and its application to a scaling limit of the spectral distributions of discrete Laplacians. Probability Theory and Related Fields, 118(1):115–130, September 2000. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008738;
 http://link.springer.com/article/10.1007/s004400000083.

Hosoya:1991:DMI

[Hos91a] Yuzo Hosoya. The decomposition and measurement of the interdependency between second-order stationary processes. Probability Theory and Related Fields, 88(4):429–444, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192551.

Hossjer:1991:CVF

[Hös91b] Ola Hössjer. The change-of-variance function for dependent data. Probability Theory and Related Fields, 90(4):447–467, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01192138.

Hossjer:1996:ABV

[Hös96] Ola Hössjer. Asymptotic bias and variance for a general class of varying bandwidth density estimators. Probability Theory and Related Fields, 105(2):159–192, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01203834.

Houdre:1990:HBB

[Hou90a] Christian Houdré. Harmonizability, V -boundedness, (2, p )- boundedness of stochastic processes. Probability Theory and Related Fields, 84(1):39–54, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01288557.

Houdre:1990:LFS

[Hou90b] Christian Houdré. Linear Fourier and stochastic analysis. Probability Theory and Related Fields, 87(2):167–188, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01198428.

Hough:2016:RCW

[Hou16] Bob Hough. The random k cycle walk on the symmetric group. Probability Theory and Related Fields, 165(1–2):447–482, June 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-015-0636-6.

Hough:2019:LLT

[Hou19] Robert Hough. The local limit theorem on nilpotent Lie groups. Probability Theory and Related Fields, 174(3–4):761–786, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-018-0864-7.

Herbst:1991:DET

[HP91] Ira Herbst and Loren Pitt. Diffusion equation techniques in stochastic monotonicity and positive correlations. Probability Theory and Related Fields, 87(3):275–312, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01312211.

Hall:1994:PNE

[HP94] Peter Hall and Prakash Patil. Properties of nonparametric estimators of autocovariance for stationary random fields. Probability Theory and Related Fields, 99(3):399–424, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01199899.

Hu:1995:SFB

[HP95] Y. Hu and S. Peng. Solution of forward-backward stochastic differential equations. Probability Theory and Related Fields, 103(2):273–283, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01204218.

Haggstrom:1999:MUP

[HP99] Olle Häggström and Yuval Peres. Monotonicity of uniqueness for percolation on Cayley graphs: all infinite clusters are born simultaneously. Probability Theory and Related Fields, 113(2): 273–285, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050208.

Hermon:2018:CLM

[HP18] Jonathan Hermon and Yuval Peres. A characterization of L_{2} mixing and hypercontractivity via hitting times and maximal inequalities. Probability Theory and Related Fields, 170(3–4): 769–800, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0769-x.

Hutchcroft:2019:CGU

[HP19] Tom Hutchcroft and Yuval Peres. The component graph of the uniform spanning forest: transitions in dimensions 9, 10, 11, …. Probability Theory and Related Fields, 175(1–2):141–208, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-018-0884-3;   https://link.springer.
com/content/pdf/10.1007/s00440-018-0884-3.pdf.

Holmes:2020:RLM

[HP20] Mark Holmes and Edwin Perkins. On the range of lattice models in high dimensions. Probability Theory and Related Fields, 176(3–4):941–1009, April 2020. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-019-
00933-1;   https://link.springer.com/content/pdf/10.
1007/s00440-019-00933-1.pdf.

Helmuth:2020:APS

[HPR20] Tyler Helmuth, Will Perkins, and Guus Regts. Algorithmic Pirogov–Sinai theory. Probability Theory and Related Fields, 176 (3–4):851–895, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00928-y.

Hiai:2004:FTC

[HPU04] Fumio Hiai, Dénes Petz, and Yoshimichi Ueda. Free transportation cost inequalities via random matrix approximation. Probability Theory and Related Fields, 130(2):199–221, October 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0351-1.

Hobson:1991:LTT

[HR91] D. G. Hobson and L. C. G. Rogers. Limit theorems for transient diffusions on the line. Probability Theory and Related Fields, 89 (1):61–74, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01225825.

Huebner:1995:APM

[HR95] M. Huebner and B. L. Rozovskii. On asymptotic properties of maximum likelihood estimators for parabolic stochastic PDE’s. Probability Theory and Related Fields, 103(2):143–163, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01204212.

Hall:2002:LBC

[HR02] Peter Hall and Christian Rau. Likelihood-based confidence bands for fault lines in response surfaces. Probability Theory and Related Fields, 124(1):26–49, September 2002. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100195.

Hoffman:2017:PAP

[HRS17] Christopher Hoffman, Douglas Rizzolo, and Erik Slivken. Pattern-avoiding permutations and Brownian excursion, part II: fixed points. Probability Theory and Related Fields, 169(1– 2):377–424, October 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-016-0732-2.

Hofmanova:2017:WSS

[HRvR17] Martina Hofmanová, Matthias Röger, and Max von Renesse. Weak solutions for a stochastic mean curvature flow of twodimensional graphs. Probability Theory and Related Fields, 168 (1–2):373–408, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Hryniv:1998:LBP

[Hry98] Ostap Hryniv. On local behaviour of the phase separation line in the 2D Ising model. Probability Theory and Related Fields, 110 (1):91–107, January 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050146.

Hwang:1987:GCI

[HS87] Chii-Ruey Hwang and Shuenn-Jyi Sheu. A generalization of Chernoff inequality via stochastic analysis. Probability Theory and Related Fields, 75(1):149–157, May 1987. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00320088.

Hall:1988:PIN

[HS88] Peter Hall and Eugene Seneta. Products of independent, normally attracted random variables. Probability Theory and Related Fields, 78(1):135–142, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00718041.

Harison:1990:JMD

[HS90a] Victor Harison and Serguey Nikolaievitch Smirnov. Jonction maximale en distribution dans le cas markovien. (French) [Maximum junction distribution in the Markov case]. Probability Theory and Related Fields, 84(4):491–503, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198316.

Hinz:1990:ADS

[HS90b] Andreas M. Hinz and Andreas Schief. The average distance on the Sierpiński gasket. Probability Theory and Related Fields, 87 (1):129–138, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01217750.

Hohnle:1994:SZO

[HS94] Rainer Höhnle and Karl-Theodor Sturm. Some zero-one laws for additive functionals of Markov processes. Probability Theory and Related Fields, 100(4):407–416, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01268986.

Hirsch:1995:MPM

[HS95a] F. Hirsch and S. Song. Markov properties of multiparameter processes and capacities. Probability Theory and Related Fields, 103(1):45–71, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199031.

Hirsch:1995:SST

[HS95b] F. Hirsch and S. Song. Symmetric Skorohod topology on nvariable functions and hierarchical Markov properties of nparameter processes. Probability Theory and Related Fields, 103(1):25–43, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199030.

Hu:1995:SFI

[HS95c] Y. Hu and Z. Shi. Sojourns and future infima of planar Brownian motion. Probability Theory and Related Fields, 103(3):329–348, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01195478.

Hu:2000:PMV

[HS00] Yueyun Hu and Zhan Shi. The problem of the most visited site in random environment. Probability Theory and Related Fields, 116(2):273–302, February 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/PL00008730.

Hu:2007:SBR

[HS07] Yueyun Hu and Zhan Shi. A subdiffusive behaviour of recurrent random walk in random environment on a regular tree. Probability Theory and Related Fields, 138(3–4):521–549, July 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-006-0036-z.

Hongler:2011:CPE

[HS11a] Clément Hongler and Stanislav Smirnov. Critical percolation: the expected number of clusters in a rectangle. Probability Theory and Related Fields, 151(3–4):735–756, December 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0313-8.

Hug:2011:FPV

[HS11b] Daniel Hug and Rolf Schneider. Faces of Poisson–Voronoi mosaics. Probability Theory and Related Fields, 151(1–2):125– 151, October 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0294-7.

Hammond:2013:PLP

[HS13] Alan Hammond and Scott Sheffield. Power law Pólya’s urn and fractional Brownian motion. Probability Theory and Related Fields, 157(3–4):691–719, December 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0468-6.

Holden:2020:SLS

[HS20] Nina Holden and Scott Sheffield. Scaling limits of the Schelling model. Probability Theory and Related Fields, 176(1–2):219– 292, February 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00918-0.

Haas:2021:SLM

[HS21] Bénédicte Haas and Robin Stephenson. Scaling limits of multitype Markov Branching trees. Probability Theory and Related Fields, 180(3–4):727–797, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01064-
2.

Hsing:1993:SEB

[Hsi93] Tailen Hsing. On some estimates based on sample behavior near high level excursions. Probability Theory and Related Fields, 95 (3):331–356, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192168.

Hilario:2015:CPT

[HST15] M. R. Hilário, V. Sidoravicius, and A. Teixeira. Cylinders’ percolation in three dimensions. Probability Theory and Related Fields, 163(3–4):613–642, December 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-014-0600-x.

Hsu:1990:BBR

[Hsu90] Pei Hsu. Brownian bridges on Riemannian manifolds. Probability Theory and Related Fields, 84(1):103–118, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01288561.

Hartmann:2018:ISP

[HSWZ18] Carsten Hartmann, Christof Schütte, Marcus Weber, and Wei Zhang. Importance sampling in path space for diffusion processes with slow–fast variables. Probability Theory and Related Fields, 170(1–2):177–228, February 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0755-3.

Hansel:1986:PMF

[HT86] Georges Hansel and Jean-Pierre Troallic. Sur le problème des marges. (French) [On the problem of margins]. Probability Theory and Related Fields, 71(3):357–366, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01000211.

Hu:2010:MDB

[HT10] Ying Hu and Shanjian Tang. Multi-dimensional BSDE with oblique reflection and optimal switching. Probability Theory and Related Fields, 147(1–2):89–121, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-009-0202-1.

Huang:2024:ACA

[HT24] Han Huang and Konstantin Tikhomirov. Average-case analysis of the Gaussian elimination with partial pivoting. Probability Theory and Related Fields, 189(1–2):501–567, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-024-01276-2.

Hudson:1989:LDS

[HTV89] William N. Hudson, Howard G. Tucker, and Jerry A. Veeh. Limit distributions of sums of m-dependent Bernoulli random variables. Probability Theory and Related Fields, 82(1):9–17, June 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00340009.

Horvath:2012:DLT

[HTV12] Illés Horváth, Bálint Tóth, and Bálint Vető. Diffusive limits for “true” (or myopic) self-avoiding random walks and self-repellent Brownian polymers in d ≥ 3. Probability Theory and Related Fields, 153(3–4):691–726, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0358-3.

Hu:1994:SFS

[Hu94] Xiaoyu Hu. Some fractal sets determined by stable processes. Probability Theory and Related Fields, 100(2):205–225, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01199266.

Huang:2024:ESL

[Hua24] Jiaoyang Huang. Edge statistics for lozenge tilings of polygons, I: concentration of height function on strip domains. Probability Theory and Related Fields, 188(1–2):337–485, February 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01238-0.

Hutchcroft:2020:NIT

[Hut20a] Tom Hutchcroft. Non-intersection of transient branching random walks. Probability Theory and Related Fields, 178(1–2):1– 23, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-020-00964-z.

Hutchcroft:2020:UHD

[Hut20b] Tom Hutchcroft. Universality of high-dimensional spanning forests and sandpiles. Probability Theory and Related Fields, 176(1–2):533–597, February 2020. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-019-
00923-3;   https://link.springer.com/content/pdf/10.
1007/s00440-019-00923-3.pdf.

Hutchcroft:2021:PLB

[Hut21] Tom Hutchcroft. Power-law bounds for critical long-range percolation below the upper-critical dimension. Probability Theory and Related Fields, 181(1-3):533–570, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01043-7.

Hutchcroft:2024:SSP

[Hut24] Tom Hutchcroft. Slightly supercritical percolation on nonamenable graphs II: growth and isoperimetry of infinite clusters. Probability Theory and Related Fields, 188(1–2):549–582, February 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01240-6.

Hryniv:2004:UCB

[HV04] O. Hryniv and Y. Velenik. Universality of critical behaviour in a class of recurrent random walks. Probability Theory and Related Fields, 130(2):222–258, October 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0353-z.

Haydn:2009:CPD

[HV09] Nicolai Haydn and Sandro Vaienti. The compound Poisson distribution and return times in dynamical systems. Probability Theory and Related Fields, 144(3–4):517–542, July 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0153-y.

Holcomb:2015:LDH

[HV15] Diane Holcomb and Benedek Valkó. Large deviations for the Sine_{β} and Sch_{τ} processes. Probability Theory and Related Fields, 163(1–2):339–378, October 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0594-4.

Herdade:2016:RWD

[HV16] Simão Herdade and Van Vu. Random walks with different directions. Probability Theory and Related Fields, 164(3–4): 1071–1078, April 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0635-7.

Heydenreich:2011:RGA

[HvdH11] Markus Heydenreich and Remco van der Hofstad. Random graph asymptotics on high-dimensional tori II: volume, diameter and mixing time. Probability Theory and Related Fields, 149 (3–4):397–415, April 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00440-009-0258-y.pdf. See cor-

rection [HvdH19].

Heydenreich:2019:CRG

[HvdH19] Markus Heydenreich and Remco van der Hofstad. Correction to: Random graph asymptotics on high-dimensional tori II: volume, diameter and mixing time. Probability Theory and Related Fields, 175(3–4):1183–1185, December 2019. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-019-
00929-x;   https://link.springer.com/content/pdf/10.
1007/s00440-019-00929-x.pdf. See [HvdH11].

Hall:1986:CDS

[HW86] Peter Hall and J. C. H. Wightwick. Convergence determining sets in the central limit theorem. Probability Theory and Related Fields, 71(1):1–17, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00366269.

Hairer:2013:ERB

[HW13a] Martin Hairer and Hendrik Weber. Erratum to: Rough Burgerslike equations with multiplicative noise. Probability Theory and Related Fields, 157(3–4):1011–1013, December 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-013-0538-4.pdf. See [HW13b].

Hairer:2013:RBL

[HW13b] Martin Hairer and Hendrik Weber. Rough Burgers-like equations with multiplicative noise. Probability Theory and Related Fields, 155(1–2):71–126, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-

011-0392-1. See erratum [HW13a].

Ihara:1990:CMG

[Iha90] S. Ihara. Capacity of mismatched Gaussian channels with and without feedback. Probability Theory and Related Fields, 84 (4):453–471, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01198314.

Izyurov:2013:HFC

[IK13] Konstantin Izyurov and Kalle Kytölä. Hadamard’s formula and couplings of SLEs with free field. Probability Theory and Related Fields, 155(1–2):35–69, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0391-2.

Ichiba:2013:SSS

[IKS13] Tomoyuki Ichiba, Ioannis Karatzas, and Mykhaylo Shkolnikov. Strong solutions of stochastic equations with rank-based coefficients. Probability Theory and Related Fields, 156(1–2):229–248, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-012-0426-3.

Imkeller:2001:CFS

[IL01] Peter Imkeller and Christian Lederer. On the cohomology of flows of stochastic and random differential equations. Probability Theory and Related Fields, 120(2):209–235, June 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008781.

Imkeller:1986:NLT

[Imk86] Peter Imkeller. A note on the localization of two-parameter processes. Probability Theory and Related Fields, 73(1):119–125, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01845995.

Imkeller:1991:TTT

[Imk91] Peter Imkeller. The transformation theorem for two-parameter pure jump martingales. Probability Theory and Related Fields, 89(3):261–283, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01198787.

Imkeller:1992:ODS

[Imk92] Peter Imkeller. Occupation densities for stochastic integral processes in the second Wiener chaos. Probability Theory and Related Fields, 91(1):1–24, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01194487.

Imkeller:1994:RSI

[Imk94] Peter Imkeller. Regularity of Skorohod integral processes based on integrands in a finite Wiener chaos. Probability Theory and Related Fields, 98(2):137–142, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01192510.

Imkeller:1996:EWF

[Imk96] Peter Imkeller. Enlargement of the Wiener filtration by an absolutely continuous random variable via Malliavin’s calculus. Probability Theory and Related Fields, 106(1):105–135, September 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050059.

Imamura:2020:SSH

[IMS20] Takashi Imamura, Matteo Mucciconi, and Tomohiro Sasamoto. Stationary stochastic Higher Spin Six Vertex Model and q-Whittaker measure. Probability Theory and Related Fields, 177(3–4):923–1042, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-020-00966-x.

Ivanoff:1993:PSL

[IMŞK93] B. Gail Ivanoff, Ely Merzbach, and Ioana Şchiopu-Kratina. Predictability and stopping on lattices of sets. Probability Theory and Related Fields, 97(4):433–446, ???? 1993. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192958.

Iyer:2016:ADF

[IN16] Gautam Iyer and Alexei Novikov. Anomalous diffusion in fast cellular flows at intermediate time scales. Probability Theory and Related Fields, 164(3–4):707–740, April 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
015-0617-9.

Inoue:2008:AMR

[Ino08] Akihiko Inoue. AR and MA representation of partial autocorrelation functions, with applications. Probability Theory and Related Fields, 140(3–4):523–551, March 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0074-1.

Inoue:2023:EFI

[Ino23] Akihiko Inoue. Explicit formulas for the inverses of Toeplitz matrices, with applications. Probability Theory and Related Fields, 185(1–2):513–552, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-022-01162-9.

Ioffe:1995:ELD

[Iof95] Dmitry Ioffe. Exact large deviation bounds up to T_{c} for the Ising model in two dimensions. Probability Theory and Related Fields, 102(3):313–330, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192464.

Ichiba:2013:CRR

[IPS13] Tomoyuki Ichiba, Soumik Pal, and Mykhaylo Shkolnikov. Convergence rates for rank-based models with applications to portfolio theory. Probability Theory and Related Fields, 156(1– 2):415–448, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0432-5.

Iverson:1989:ECS

[IR89] H. K. Iverson and R. H. Randles. The effects on convergence of substituting parameter estimates into U-statistics and other families of statistics. Probability Theory and Related Fields, 81 (3):453–471, April 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00340061.

Ignatiouk-Robert:2006:SMS

[IR06] Irina Ignatiouk-Robert. On the spectrum of Markov semigroups via sample path large deviations. Probability Theory and Related Fields, 134(1):44–80, January 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0410-7.

Ignatiouk-Robert:2010:MBR

[IR10] Irina Ignatiouk-Robert. Martin boundary of a reflected random walk on a half-space. Probability Theory and Related Fields, 148(1–2):197–245, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-009-0228-4.

Sato:1991:SSP

[iS91] Ken iti Sato. Self-similar processes with independent increments. Probability Theory and Related Fields, 89(3):285–300, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01198788.

Imamura:2019:FST

[IS19] Takashi Imamura and Tomohiro Sasamoto. Fluctuations for stationary q-TASEP. Probability Theory and Related Fields, 174 (1–2):647–730, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0868-3.

Iscoe:1986:WOT

[Isc86] I. Iscoe. A weighted occupation time for a class of measuredvalued branching processes. Probability Theory and Related Fields, 71(1):85–116, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00366274.

Ito:1988:GPF

[Ito88] Yoshifusa Ito. Generalized Poisson functionals. Probability Theory and Related Fields, 77(1):1–28, ???? 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01848128.

Ioffe:2000:NDC

[IV00] Dmitry Ioffe and Yvan Velenik. A note on the decay of correlations under δ-pinning. Probability Theory and Related Fields, 116(3):379–389, March 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050254.

Ioffe:2018:DFS

[IVW18] Dmitry Ioffe, Yvan Velenik, and Vitali Wachtel. Dyson Ferrari– Spohn diffusions and ordered walks under area tilts. Probability Theory and Related Fields, 170(1–2):11–47, February 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0751-z.

Imkeller:1994:ABL

[IW94] Peter Imkeller and Ferene Weisz. The asymptotic behaviour of local times and occupation integrals of the N-parameter Wiener process in R^{d}. Probability Theory and Related Fields, 98(1):47–75, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01311348.

Iwata:1987:IDS

[Iwa87] Koichiro Iwata. An infinite dimensional stochastic differential equation with state space C (r). Probability Theory and Related Fields, 74(1):141–159, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01845644.

Izyurov:2017:CII

[Izy17] Konstantin Izyurov. Critical Ising interfaces in multiplyconnected domains. Probability Theory and Related Fields, 167(1–2):379–415, February 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s00440-015-
0685-x;   http://link.springer.com/article/10.1007/
s00440-015-0685-x.

Jacod:1987:CPP

[Jac87] Jean Jacod. Sur la convergence des processus ponctuels. (French) [On the convergence of point processes]. Probability Theory and Related Fields, 76(4):573–586, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00960075.

Jacod:1990:RPR

[Jac90] Jean Jacod. Regularity, partial regularity, partial information process, for a filtered statistical model. Probability Theory and Related Fields, 86(3):305–335, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01208255.

Jaffard:1999:MNL

[Jaf99] Stéphane Jaffard. The multifractal nature of Lévy processes. Probability Theory and Related Fields, 114(2):207–227, May 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050224.

Janssen:1988:UCS

[Jan88] A. Janssen. Uniform convergence of sums of order statistics to stable laws. Probability Theory and Related Fields, 78(2): 261–271, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00322023.

Janssen:1995:PCD

[Jan95] Arnold Janssen. Principal component decomposition of nonparametric tests. Probability Theory and Related Fields, 101 (2):193–209, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01375824.

Janson:2006:LTT

[Jan06] Svante Janson. Limit theorems for triangular urn schemes. Probability Theory and Related Fields, 134(3):417–452, March 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-005-0442-7.

Janssen:2010:LLP

[Jan10] Anja Janßen. Limit laws for power sums and norms of i.i.d. samples. Probability Theory and Related Fields, 146(3–4):515–533, March 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00440-008-0198-y.pdf.

Jaworski:1995:AIA

[Jaw95] Wojciech Jaworski. On the asymptotic and invariant σ-algebras of random walks on locally compact groups. Probability Theory and Related Fields, 101(2):147–171, ???? 1995. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01375822.

Johnson:2004:FII

[JB04] Oliver Johnson and Andrew Barron. Fisher information inequalities and the central limit theorem. Probability Theory and Related Fields, 129(3):391–409, July 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0344-0.

Jego:2021:CBM

[Jeg21] Antoine Jego. Critical Brownian multiplicative chaos. Probability Theory and Related Fields, 180(1–2):495–552, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01051-7.

Jensen:1989:VFE

[Jen89] J. L. Jensen. Validity of the formal Edgeworth expansion when the underlying distribution is partly discrete. Probability Theory and Related Fields, 81(4):507–519, May 1989. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00367300.

Jensen:1991:SES

[Jen91] Jens Ledet Jensen. Saddlepoint expansions for sums of Markov dependent variables on a continuous state space. Probability Theory and Related Fields, 89(2):181–199, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01366905.

Ji:1989:EFO

[Ji89] Chuanshu Ji. Estimating functionals of one-dimensional Gibbs states. Probability Theory and Related Fields, 82(2):155–175, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00354757.

Jiang:2005:MEH

[Jia05] Tiefeng Jiang. Maxima of entries of Haar distributed matrices. Probability Theory and Related Fields, 131(1):121–144, January 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-004-0376-5.

Jiang:2009:AHD

[Jia09a] Tiefeng Jiang. Approximation of Haar distributed matrices and limiting distributions of eigenvalues of Jacobi ensembles. Probability Theory and Related Fields, 144(1–2):221–246, May 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-008-0146-x.

Jiao:2009:CMV

[Jia09b] Yong Jiao. Carleson measures and vector-valued BMO martingales. Probability Theory and Related Fields, 145(3–4):421– 434, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0173-7.

Jiang:2012:LEL

[Jia12] Tiefeng Jiang. Low eigenvalues of Laplacian matrices of large random graphs. Probability Theory and Related Fields, 153(3– 4):671–690, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0357-4.

Jirak:2016:OEE

[Jir16] Moritz Jirak. Optimal eigen expansions and uniform bounds. Probability Theory and Related Fields, 166(3–4):753–799, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0671-3;

http://link.springer.com/article/10.1007/s00440-015-
0671-3.

Jona-Lasinio:1993:LDR

[JLLV93] G. Jona-Lasinio, C. Landim, and M. E. Vares. Large deviations for a reaction diffusion model. Probability Theory and Related Fields, 97(3):339–361, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01195070.

Jara:2009:NFT

[JLS09] M. D. Jara, C. Landim, and S. Sethuraman. Nonequilibrium fluctuations for a tagged particle in mean-zero one-dimensional zero-range processes. Probability Theory and Related Fields, 145(3–4):565–590, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0178-2.

Janssen:1990:RCS

[JM90] Arnold Janssen and David M. Mason. On the rate of convergence of sums of extremes to a stable law. Probability Theory and Related Fields, 86(2):253–264, ???? 1990. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01474645.

Jakubowski:1989:CLS

[JMP89] A. Jakubowski, J. Mémin, and G. Pages. Convergence en loi des suites d’intégrales stochastiques sur l’espace D^{1} de Skorokhod. (French) [Convergence in law of sequences of stochastic integrals on the Skorokhod space D^{1}]. Probability Theory and Related Fields, 81(1):111–137, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00343739.

Janson:1991:ADG

[JN91] Svante Janson and Krzysztof Nowicki. The asymptotic distributions of generalized U-statistics with applications to random graphs. Probability Theory and Related Fields, 90(3):341–375, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01193750.

Janson:2008:SRF

[JN08] Svante Janson and Ralph Neininger. The size of random fragmentation trees. Probability Theory and Related Fields, 142 (3–4):399–442, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-007-0110-1.

Johnstone:1986:AED

[Joh86] Iain Johnstone. Admissible estimation, Dirichlet principles and recurrence of birth-death chains on Z_{+}^{p}. Probability Theory and Related Fields, 71(2):231–269, January 1986. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00332311.

Johansson:2000:TFI

[Joh00] Kurt Johansson. Transversal fluctuations for increasing subsequences on the plane. Probability Theory and Related Fields, 116 (4):445–456, April 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050258.

Johansson:2002:NIP

[Joh02] Kurt Johansson. Non-intersecting paths, random tilings and random matrices. Probability Theory and Related Fields, 123 (2):225–280, June 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400100187.

Johansson:2007:GTW

[Joh07] K. Johansson. From Gumbel to Tracy–Widom. Probability Theory and Related Fields, 138(1–2):75–112, May 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0012-7.

Johansson:2019:TTD

[Joh19] Kurt Johansson. The two-time distribution in geometric last-passage percolation. Probability Theory and Related Fields, 175(3–4):849–895, December 2019. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/s00440-019-
00901-9;   https://link.springer.com/content/pdf/10.
1007/s00440-019-00901-9.pdf.

Jonasson:2009:IRG

[Jon09] Johan Jonasson. Invariant random graphs with iid degrees in a general geography. Probability Theory and Related Fields, 143(3–4):643–656, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0160-z.

Journe:1987:SCM

[Jou87] J.-L. Journé. Structure des cocycles markoviens sur l’espace de Fock. (French) [Markov cocycle structure in Fock space]. Probability Theory and Related Fields, 75(2):291–316, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00354039.

Jorgensen:1991:PEA

[JP91] Palle E. T. Jorgensen and Robert T. Powers. Positive elements in the algebra of the quantum moment problem. Probability Theory and Related Fields, 89(2):131–139, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01366901.

Jia:2010:JIC

[JP10] Guangyan Jia and Shige Peng. Jensen’s inequality for g-convex function under g-expectation. Probability Theory and Related Fields, 147(1–2):217–239, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-009-0206-x.

Joos:2018:HDI

[JPRR18] Felix Joos, Guillem Perarnau, Dieter Rautenbach, and Bruce Reed. How to determine if a random graph with a fixed degree sequence has a giant component. Probability Theory and Related Fields, 170(1–2):263–310, February 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0757-1.

Janssen:1988:CLM

[JR88] A. Janssen and R.-D. Reiss. Comparison of location models of Weibull type samples and extreme value processes. Probability Theory and Related Fields, 78(2):273–292, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00322024.

Jacquot:1995:ESP

[JR95] S. Jacquot and G. Royer. Ergodicity of stochastic plates. Probability Theory and Related Fields, 102(1):19–44, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01295219.

Jarai:2008:IVL

[JR08] Antal A. Járai and Frank Redig. Infinite volume limit of the Abelian sandpile model in dimensions d ≥ 3. Probability Theory and Related Fields, 141(1–2):181–212, May 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0083-0.

Janvresse:2008:HDR

[JRdlR08] Élise Janvresse, Benoît Rittaud, and Thierry de la Rue. How do random Fibonacci sequences grow? Probability Theory and Related Fields, 142(3–4):619–648, November 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0117-7.

Jarai:2020:MFA

[JRS20] Antal A. Járai, Wioletta M. Ruszel, and Ellen Saada. Meanfield avalanche size exponent for sandpiles on Galton–Watson trees. Probability Theory and Related Fields, 177(1–2):369–396, June 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-019-00951-z;   https://link.springer.
com/content/pdf/10.1007/s00440-019-00951-z.pdf.

Jakubowski:1986:ECC

[JS86] Adam Jakubowski and Leszek Slomiński. Extended convergence to continuous in probability processes with independent increments. Probability Theory and Related Fields, 72(1):55–82, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00343896.

Jagannath:2017:SPP

[JT17] Aukosh Jagannath and Ian Tobasco. Some properties of the phase diagram for mixed p-spin glasses. Probability Theory and Related Fields, 167(3–4):615–672, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0691-z.

Jurek:1988:RIR

[Jur88] Zbigniew J. Jurek. Random integral representations for classes of limit distributions similar to Lévy class L_{0}. Probability Theory and Related Fields, 78(3):473–490, July 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00334208.

Jirak:2024:QLT

[JW24] Moritz Jirak and Martin Wahl. Quantitative limit theorems and bootstrap approximations for empirical spectral projectors. Probability Theory and Related Fields, 190(1–2):119–177, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-024-01290-4.

Jacobsen:2003:MSS

[JY03] Martin Jacobsen and Marc Yor. Multi-self-similar Markov processes on R_{+}^{n} and their Lamperti representations. Probability Theory and Related Fields, 126(1):1–28, May 2003. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0263-5.

Ji:2010:GNP

[JZ10] Shaolin Ji and Xun Yu Zhou. A generalized Neyman–Pearson lemma for g-probabilities. Probability Theory and Related Fields, 148(3–4):645–669, November 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-009-0244-4.

Junge:2015:NMD

[JZ15] Marius Junge and Qiang Zeng. Noncommutative martingale deviation and Poincaré type inequalities with applications. Probability Theory and Related Fields, 161(3–4):449–507, April 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-014-0552-1.

Kajino:2013:DOH

[Kaj13] Naotaka Kajino. On-diagonal oscillation of the heat kernels on post-critically finite self-similar fractals. Probability Theory and Related Fields, 156(1–2):51–74, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0420-9.

Kallenberg:1988:SNR

[Kal88] Olav Kallenberg. Some new representations in bivariate exchangeability. Probability Theory and Related Fields, 77(3): 415–455, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319298.

Kallenberg:1989:GWT

[Kal89] Olav Kallenberg. General wald-type identities for exchangeable sequences and processes. Probability Theory and Related Fields, 83(4):447–487, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01845699.

Kallenberg:1990:RTC

[Kal90] Olav Kallenberg. Random time change and an integral representation for marked stopping times. Probability Theory and Related Fields, 86(2):167–202, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01474641.

Kallenberg:1995:RAF

[Kal95] O. Kallenberg. Random arrays and functionals with multivariate rotational symmetries. Probability Theory and Related Fields, 103(1):91–141, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199033.

Kallenberg:2000:EBP

[Kal00a] Olav Kallenberg. An extension of the basic Palm measure correspondence. Probability Theory and Related Fields, 117 (1):113–131, May 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050267.

Kallenberg:2000:SIS

[Kal00b] Olav Kallenberg. Spreading-invariant sequences and processes on bounded index sets. Probability Theory and Related Fields, 118(2):211–250, October 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s440-000-8015-x.

Kallenberg:2003:PDL

[Kal03] Olav Kallenberg. Palm distributions and local approximation of regenerative processes. Probability Theory and Related Fields, 125(1):1–41, January 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400200218.

Kallenberg:2007:IMDa

[Kal07a] Olav Kallenberg. Invariant measures and disintegrations with applications to Palm and related kernels. Probability Theory and Related Fields, 139(1–2):285–310, September 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0053-y.

Kallenberg:2007:IMDb

[Kal07b] Olav Kallenberg. Invariant measures and disintegrations with applications to Palm and related kernels. Probability Theory and Related Fields, 139(1–2):311, September 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-007-0071-4.

Kallenberg:2009:SLA

[Kal09] Olav Kallenberg. Some local approximation properties of simple point processes. Probability Theory and Related Fields, 143(1– 2):73–96, January 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0120-z.

Kallenberg:2011:IPR

[Kal11] Olav Kallenberg. Invariant Palm and related disintegrations via skew factorization. Probability Theory and Related Fields, 149(1–2):279–301, February 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-009-0254-2.

Kallenberg:2014:SID

[Kal14] Olav Kallenberg. Stationary and invariant densities and disintegration kernels. Probability Theory and Related Fields, 160(3–4): 567–592, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0535-7.

Kanter:1990:DDB

[Kan90] Marek Kanter. Discrimination distance bounds and statistical applications. Probability Theory and Related Fields, 86(3): 403–422, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01208258.

Kaneko:2000:MKB

[Kan00] Hiroshi Kaneko. Martin–Kuramochi boundary and reflecting symmetric diffusion. Probability Theory and Related Fields, 117 (4):533–550, August 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008734.

Karr:1987:EPM

[Kar87] Alan F. Karr. Estimation of Palm measures of stationary point processes. Probability Theory and Related Fields, 74(1):55–69, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01845639.

Kargin:2007:NPF

[Kar07] Vladislav Kargin. The norm of products of free random variables. Probability Theory and Related Fields, 139(3–4):397– 413, November 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0046-x.

Kargin:2008:LTP

[Kar08] Vladislav Kargin. A limit theorem for products of free unitary operators. Probability Theory and Related Fields, 141(3– 4):603–623, July 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0095-9.

Kargin:2012:CIL

[Kar12] Vladislav Kargin. A concentration inequality and a local law for the sum of two random matrices. Probability Theory and Related Fields, 154(3–4):677–702, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0381-4.

Kargin:2013:FRZ

[Kar13] V. Kargin. On fluctuations of Riemann’s zeta zeros. Probability Theory and Related Fields, 157(3–4):575–604, December 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-012-0465-9.

Katori:2015:EDP

[Kat15] Makoto Katori. Elliptic determinantal process of type A. Probability Theory and Related Fields, 162(3–4):637–677, August 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-014-0581-9.

Kazamaki:1988:NAS

[Kaz88] Norihiko Kazamaki. A new aspect of L_{∞} in the space of BMO-martingales. Probability Theory and Related Fields, 78 (1):113–126, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00718039.

Keener:1988:AER

[Kee88] Robert Keener. Asymptotic expansions for renewal measures in the plane. Probability Theory and Related Fields, 80(1): 1–20, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00348749.

Kellerer:1987:MPP

[Kel87] Hans G. Kellerer. Markov property of point processes. Probability Theory and Related Fields, 76(1):71–80, September 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00390276.

Kempe:2005:DQW

[Kem05] Julia Kempe. Discrete quantum walks hit exponentially faster. Probability Theory and Related Fields, 133(2):215–235, October 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-004-0423-2.

Kersting:1989:APS

[Ker89] G. Kersting. Asymptotic properties of solutions of multidimensional stochastic differential equations. Probability Theory and Related Fields, 82(2):187–211, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00354759.

Kesten:1986:IIC

[Kes86] Harry Kesten. The incipient infinite cluster in two-dimensional percolation. Probability Theory and Related Fields, 73(3): 369–394, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00776239.

Kessler:1988:MTP

[Kes88] C. Kessler. Markov type properties for mixtures of probability measures. Probability Theory and Related Fields, 78(2):253–259, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00322022.

Key:1987:CEM

[Key87] Eric Key. Computable examples of the maximal Lyapunov exponent. Probability Theory and Related Fields, 75(1):97–107, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00320084.

Kohatsu-Higa:2003:LBD

[KH03] Arturo Kohatsu-Higa. Lower bounds for densities of uniformly elliptic random variables on Wiener space. Probability Theory and Related Fields, 126(3):421–457, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0272-4.

Khoshnevisan:1992:LAL

[Kho92] Davar Khoshnevisan. Local asymptotic laws for the Brownian convex hull. Probability Theory and Related Fields, 93(3): 377–392, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193057.

Khoshnevisan:2008:DPG

[Kho08] Davar Khoshnevisan. Dynamical percolation on general trees. Probability Theory and Related Fields, 140(1–2):169–193, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-007-0061-6.

Kifer:1995:MRS

[Kif95] Y. Kifer. Multidimensional random subshifts of finite type and their large deviations. Probability Theory and Related Fields, 103(2):223–248, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01204216.

Kifer:1997:BER

[Kif97] Yuri Kifer. The Burgers equation with a random force and a general model for directed polymers in random environments. Probability Theory and Related Fields, 108(1):29–65, May 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050100.

Kifer:2010:NLT

[Kif10] Yuri Kifer. Nonconventional limit theorems. Probability Theory and Related Fields, 148(1–2):71–106, September 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
009-0223-9.

Kifer:2013:ESI

[Kif13a] Yuri Kifer. Erratum to: A strong invariance principle for nonconventional sums. Probability Theory and Related Fields, 157 (1–2):509, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00440-013-0489-9.pdf.  See

[Kif13b].

Kifer:2013:SIP

[Kif13b] Yuri Kifer. A strong invariance principle for nonconventional sums. Probability Theory and Related Fields, 155(1–2):463–486, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/s00440-011-0404-1. See erratum [Kif13a].

Kim:2004:TSP

[Kim04] Kyeong-Hun Kim. On L_{p}-theory of stochastic partial differential equations of divergence form in C^{1} domains. Probability Theory and Related Fields, 130(4):473–492, December 2004. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0368-5.

Kim:2021:STS

[Kim21] Seonwoo Kim. Second time scale of the metastability of reversible inclusion processes. Probability Theory and Related Fields, 180(3–4):1135–1187, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01036-

6. See correction [Kim23].

Kim:2023:CST

[Kim23] Seonwoo Kim. Correction: Second time scale of the metastability of reversible inclusion processes. Probability Theory and Related Fields, 186(3–4):1275–1282, August 2023. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-023-01205-9. See [Kim21].

Kiss:2015:FPT

[Kis15] Demeter Kiss. Frozen percolation in two dimensions. Probability Theory and Related Fields, 163(3–4):713–768, December 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-014-0603-7.

Kumagai:1996:HNF

[KK96] T. Kumagai and S. Kusuoka. Homogenization on nested fractals. Probability Theory and Related Fields, 104(3):375–398, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01213686.

Kotelenez:2010:MLS

[KK10] Peter M. Kotelenez and Thomas G. Kurtz. Macroscopic limits for stochastic partial differential equations of McKean– Vlasov type. Probability Theory and Related Fields, 146(1–2): 189–222, January 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0188-0.

Kabluchko:2014:CRE

[KK14] Zakhar Kabluchko and Anton Klimovsky. Complex random energy model: zeros and fluctuations. Probability Theory and Related Fields, 158(1–2):159–196, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0480-5.

Khasminskii:1999:EPL

[KKM99] R. Khasminskii, N. Krylov, and N. Moshchuk. On the estimation of parameters for linear stochastic differential equations. Probability Theory and Related Fields, 113(3):443–472, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050213.

Keller:1988:ABF

[KKR88] G. Keller, G. Kersting, and U. Rösler. On the asymptotic behaviour of first passage times for discussions. Probability Theory and Related Fields, 77(3):379–395, March 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00319295.

Kusuoka:2010:LDS

[KKT10] Shigeo Kusuoka, Kazumasa Kuwada, and Yozo Tamura. Large deviation for stochastic line integrals as L^{p}-currents. Probability Theory and Related Fields, 147(3–4):649–674, July 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
009-0219-5.

Kuelbs:1987:EVG

[KL87a] J. Kuelbs and M. Ledoux. Extreme values and a Gaussian central limit theorem. Probability Theory and Related Fields, 74 (3):341–355, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00699095.

Kuelbs:1987:EVL

[KL87b] J. Kuelbs and M. Ledoux. Extreme values and the law of the iterated logarithm. Probability Theory and Related Fields, 74 (3):319–340, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00699094.

Korostelev:1993:AFD

[KL93] A. P. Korostelev and S. L. Leonov. Action functional for diffusions in discontinuous media. Probability Theory and Related Fields, 94(3):317–333, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199247.

Kusuoka:2000:LAS

[KL00] Shigeo Kusuoka and Song Liang. Laplace approximations for sums of independent random vectors. Probability Theory and Related Fields, 116(2):221–238, February 2000. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008727.

Klartag:2009:BET

[Kla09] Boáz Klartag. A Berry–Esseen type inequality for convex bodies with an unconditional basis. Probability Theory and Related Fields, 145(1–2):1–33, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-008-0158-6.

Klemela:2006:SAE

[Kle06] Jussi Klemelä. Sharp adaptive estimation of quadratic functionals. Probability Theory and Related Fields, 134(4):539–564, April 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-005-0447-2.

Khoshnevisan:1994:FIS

[KLL94a] Davar Khoshnevisan, Thomas M. Lewis, and Wenbo V. Li. On the future infima of some transient processes. Probability Theory and Related Fields, 99(3):337–360, ???? 1994. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01199896.

Kuelbs:1994:GMS

[KLL94b] James Kuelbs, Wenbo V. Li, and Werner Linde. The Gaussian measure of shifted balls. Probability Theory and Related Fields, 98(2):143–162, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192511.

Khoshnevisan:2006:ETI

[KLMH06] Davar Khoshnevisan, David A. Levin, and Pedro J. Méndez-Hernández. Exceptional times and invariance for dynamical random walks. Probability Theory and Related Fields, 134(3): 383–416, March 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-005-0435-6.

Klopp:2013:AEE

[Klo13] Frédéric Klopp. Asymptotic ergodicity of the eigenvalues of random operators in the localized phase. Probability Theory and Related Fields, 155(3–4):867–909, April 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0415-6.

Kerkyacharian:2001:NEA

[KLP01] Gérard Kerkyacharian, Oleg Lepski, and Dominique Picard. Nonlinear estimation in anisotropic multi-index denoising. Probability Theory and Related Fields, 121(2):137–170, October 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008800.

Kabanov:1986:VDP

[KLS86] Yu. M. Kabanov, R. Sh. Liptser, and A. N. Shiryaev. On the variation distance for probability measures defined on a filtered space. Probability Theory and Related Fields, 71(1):19–35, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00366270.

Klopp:2017:RMC

[KLT17] Olga Klopp, Karim Lounici, and Alexandre B. Tsybakov. Robust matrix completion. Probability Theory and Related Fields, 169(1–2):523–564, October 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-016-0736-y.

Kluppelberg:1989:SDC

[Klü89] Claudia Klüppelberg. Subexponential distributions and characterizations of related classes. Probability Theory and Related Fields, 82(2):259–269, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00354763.

Kolliopoulos:2023:PCM

[KLZ23] Nikolaos Kolliopoulos, Martin Larsson, and Zeyu Zhang. Propagation of chaos for maxima of particle systems with mean-field drift interaction. Probability Theory and Related Fields, 187 (3–4):1093–1127, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01213-9.

Kasahara:1986:FLT

[KM86] Yuji Kasahara and Makoto Maejima. Functional limit theorems for weighted sums of I.I.D. random variables. Probability Theory and Related Fields, 72(2):161–183, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00699101.

Kasahara:1988:WSD

[KM88] Yuji Kasahara and Makoto Maejima. Weighted sums of I.I.D. random variables attracted to integrals of stable processes. Probability Theory and Related Fields, 78(1):75–96, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00718037.

Koul:1993:ARM

[KM93] Hira L. Koul and Kanchan Mukherjee. Asymptotics of R-, MDand LAD-estimators in linear regression models with long range dependent errors. Probability Theory and Related Fields, 95 (4):535–553, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01196733.

Kesten:1996:TRT

[KM96] Harry Kesten and R. A. Maller. Two renewal theorems for general random walks tending to infinity. Probability Theory and Related Fields, 106(1):1–38, September 1996. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050056.

Konakov:2000:LLT

[KM00] Valentin Konakov and Enno Mammen. Local limit theorems for transition densities of Markov chains converging to diffusions. Probability Theory and Related Fields, 117(4):551–587, August 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008735.

Kinateder:2002:IFD

[KM02] Kimberly Kinateder and Patrick McDonald. An Ito formula for domain-valued processes driven by stochastic flows. Probability Theory and Related Fields, 124(1):73–99, September 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200201.

Konakov:2009:STE

[KM09] Valentin Konakov and Enno Mammen. Small time Edgeworthtype expansions for weakly convergent nonhomogeneous Markov chains. Probability Theory and Related Fields, 143(1–2):137– 176, January 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0123-9.

Klenke:2012:IRM

[KM12a] Achim Klenke and Leonid Mytnik. Infinite rate mutually catalytic branching in infinitely many colonies: construction, characterization and convergence. Probability Theory and Related Fields, 154(3–4):533–584, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0376-1.

Kontoyiannis:2012:GES

[KM12b] I. Kontoyiannis and S. P. Meyn. Geometric ergodicity and the spectral gap of non-reversible Markov chains. Probability Theory and Related Fields, 154(1–2):327–339, October 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0373-4.

Kabluchko:2022:LDS

[KM22] Zakhar Kabluchko and Alexander Marynych. Lah distribution: Stirling numbers, records on compositions, and convex hulls of high-dimensional random walks. Probability Theory and Related Fields, 184(3–4):969–1028, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-022-01146-
9.

Kinateder:1998:ETM

[KMM98] K. K. J. Kinateder, Patrick McDonald, and David Miller. Exit time moments, boundary value problems, and the geometry of domains in Euclidean space. Probability Theory and Related Fields, 111(4):469–487, August 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050174.

Kosygina:2022:CRW

[KMP22] Elena Kosygina, Thomas Mountford, and Jonathon Peterson. Convergence of random walks with Markovian cookie stacks to Brownian motion perturbed at extrema. Probability Theory and Related Fields, 182(1–2):189–275, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01055-3.

Kabluchko:2019:CGR

[KMTT19] Zakhar Kabluchko, Alexander Marynych, Daniel Temesvari, and Christoph Thäle. Cones generated by random points on half-spheres and convex hulls of Poisson point processes. Probability Theory and Related Fields, 175(3–4):1021–1061, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-019-00907-3.

Kosygina:2011:LEG

[KMZ11] Elena Kosygina, Thomas S. Mountford, and Martin P. W. Zerner. Lyapunov exponents of Green’s functions for random potentials tending to zero. Probability Theory and Related Fields, 150(1–2):43–59, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0266-y.

Klass:1998:OMB

[KN98] Michael J. Klass and Krzysztof Nowicki. Order of magnitude bounds for expectations of Δ_{2}-functions of generalized random bilinear forms. Probability Theory and Related Fields, 112(4): 457–492, December 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050197.

Klass:2003:OBT

[KN03] Michael J. Klass and Krzysztof Nowicki. An optimal bound on the tail distribution of the number of recurrences of an event in product spaces. Probability Theory and Related Fields, 126 (1):51–60, May 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-002-0252-0.

Knight:2001:PIO

[Kni01] Frank B. Knight. On the path of an inert object impinged on one side by a Brownian particle. Probability Theory and Related Fields, 121(4):577–598, December 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400100160.

Kerkyacharian:2012:CIC

[KNP12] Gerard Kerkyacharian, Richard Nickl, and Dominique Picard. Concentration inequalities and confidence bands for needlet density estimators on compact homogeneous manifolds. Probability Theory and Related Fields, 153(1–2):363–404, June 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0348-5.

Komorowski:2001:HTD

[KO01] Tomasz Komorowski and Stefano Olla. On homogenization of time-dependent random flows. Probability Theory and Related Fields, 121(1):98–116, September 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/PL00008799.

Kolsrud:1987:MPC

[Kol87] Torbjörn Kolsrud. On the Markov property for certain Gaussian random fields. Probability Theory and Related Fields, 74 (3):393–402, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00699097.

Kolokoltsov:2003:MVL

[Kol03] Vassili N. Kolokoltsov. Measure-valued limits of interacting particle systems with k-nary interactions. Probability Theory and Related Fields, 126(3):364–394, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0267-1.

Kolesnikov:2008:WCD

[Kol08] Alexander V. Kolesnikov. Weak convergence of diffusion processes on Wiener space. Probability Theory and Related Fields, 140(1–2):1–17, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-006-0023-4.

Kolokoltsov:2010:CLT

[Kol10] Vassili N. Kolokoltsov. The central limit theorem for the Smoluchovski coagulation model. Probability Theory and Related Fields, 146(1–2):87–153, January 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-008-0186-2.

Kolokoltsov:2011:LKT

[Kol11] Vassili N. Kolokoltsov. The Lévy–Khintchine type operators with variable Lipschitz continuous coefficients generate linear or nonlinear Markov processes and semigroups. Probability Theory and Related Fields, 151(1–2):95–123, October 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0293-8.

Kolesko:2013:THI

[Kol13] Konrad Kolesko. Tail homogeneity of invariant measures of multidimensional stochastic recursions in a critical case. Probability Theory and Related Fields, 156(3–4):593–612, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-012-0437-0.pdf.

Komorowski:2001:DAC

[Kom01]  Tomasz Komorowski.  Diffusion approximation for the

convection-diffusion equation with random drift. Probability Theory and Related Fields, 121(4):525–550, December 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100159.

Konakov:1990:ESG

[Kon90] V. D. Konakov. Extrema of some Gaussian processes with large trends and density estimation in L_{∞} norm. Probability Theory and Related Fields, 86(3):277–304, ???? 1990. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01208254.

Konig:1993:DOD

[Kön93] Wolfgang König. The drift of a one-dimensional self-avoiding random walk. Probability Theory and Related Fields, 96(4): 521–543, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01200208.

Konig:1994:DOD

[Kön94] Wolfgang König. The drift of a one-dimensional self-repellent random walk with bounded increments. Probability Theory and Related Fields, 100(4):513–544, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01268992.

Kopfer:2019:SRF

[Kop19] Eva Kopfer. Super-Ricci flows and improved gradient and transport estimates. Probability Theory and Related Fields, 175(3–4): 897–936, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00904-6.

Koralov:2004:RPD

[Kor04] Leonid Koralov. Random perturbations of 2-dimensional Hamiltonian flows. Probability Theory and Related Fields, 129(1):37– 62, May 2004. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/s00440-003-0320-0.

Kortchemski:2017:SET

[Kor17] Igor Kortchemski. Sub-exponential tail bounds for conditioned stable Bienaymé–Galton–Watson trees. Probability Theory and Related Fields, 168(1–2):1–40, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Korotkikh:2022:HDI

[Kor22] Sergei Korotkikh. Hidden diagonal integrability of q-Hahn vertex model and Beta polymer model. Probability Theory and Related Fields, 184(1–2):493–570, October 2022. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-022-01117-0.

Kotelenez:1988:HDL

[Kot88] Peter Kotelenez. High density limit theorems for nonlinear chemical reactions with diffusion. Probability Theory and Related Fields, 78(1):11–37, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00718032.

Kotelenez:1992:CMC

[Kot92] Peter Kotelenez. Comparison methods for a class of function valued stochastic partial differential equations. Probability Theory and Related Fields, 93(1):1–19, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01195385.

Kotelenez:1995:CQS

[Kot95] Peter Kotelenez. A class of quasilinear stochastic partial differential equations of McKean–Vlasov type with mass conservation. Probability Theory and Related Fields, 102(2):159–188, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01213387.

Koul:1989:QLT

[Kou89] Hira L. Koul. A quadraticity limit theorem useful in linear models. Probability Theory and Related Fields, 82(3):371–386, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00339993.

Kou:2003:ESC

[Kou03] S. C. Kou. On the efficiency of selection criteria in spline regression. Probability Theory and Related Fields, 127(2):153–176, October 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0277-z.

Karr:1986:SPR

[KP86a] A. F. Karr and A. O. Pittenger. Structural properties of randomized times. Probability Theory and Related Fields, 72(3): 395–415, June 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00334193.

Khmaladze:1986:FLT

[KP86b] E. V. Khmaladze and A. M. Parjanadze. Functional limit theorems for linear statistics from sequential ranks. Probability Theory and Related Fields, 73(4):585–595, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00324854.

Kenig:1989:PDL

[KP89] Carlos E. Kenig and Jill Pipher. The h-path distribution of the lifetime of conditioned Brownian motion for non-smooth domains. Probability Theory and Related Fields, 82(4):615–623, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00341286.

Kahane:1988:MUM

[KPyWmW88] Jean-Pierre Kahane, Jacques Peyrière, Zhi ying Wen, and Li ming Wu. Moyennes uniformes et moyennes suivant une marche aléatoire. (French) [Uniform mediums and mediums following a random walk]. Probability Theory and Related Fields, 79(4):625–628, November 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00318786.

Krylov:2005:SSS

[KR05] N. V. Krylov and M. Röckner. Strong solutions of stochastic equations with singular time dependent drift. Probability Theory and Related Fields, 131(2):154–196, February 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0361-z.

Kifer:2014:PCP

[KR14] Yuri Kifer and Ariel Rapaport. Poisson and compound Poisson approximations in conventional and nonconventional setups. Probability Theory and Related Fields, 160(3–4):797–831, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-013-0541-9.

Karatzas:2016:DTE

[KR16a] Ioannis Karatzas and Johannes Ruf. Distribution of the time to explosion for one-dimensional diffusions. Probability Theory and Related Fields, 164(3–4):1027–1069, April 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
015-0625-9.

Kaspar:2016:SCL

[KR16b] David C. Kaspar and Fraydoun Rezakhanlou. Scalar conservation laws with monotone pure-jump Markov initial conditions. Probability Theory and Related Fields, 165(3–4):867– 899, August 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0648-2;
http://link.springer.com/article/10.1007/s00440-015-
0648-2.

Kramkov:1996:ODS

[Kra96] D. O. Kramkov. Optional decomposition of supermartingales and hedging contingent claims in incomplete security markets. Probability Theory and Related Fields, 105(4):459–479, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01191909.

Krebs:1995:BMC

[Kre95] W. B. Krebs. Brownian motion on the continuum tree. Probability Theory and Related Fields, 101(3):421–433, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01200505.

Krener:1997:RDF

[Kre97] Arthur J. Krener. Reciprocal diffusions in flat space. Probability Theory and Related Fields, 107(2):243–281, February 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050085.

Keller-Ressel:2011:APR

[KRST11] Martin Keller-Ressel, Walter Schachermayer, and Josef Teichmann. Affine processes are regular. Probability Theory and Related Fields, 151(3–4):591–611, December 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0309-4.

Krylov:1994:TDP

[Kry94] N. V. Krylov. A W_{2}^{n}-theory of the Dirichlet problem for SPDEs in general smooth domains. Probability Theory and Related Fields, 98(3):389–421, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192260.

Krylov:1997:RCM

[Kry97] N. V. Krylov. On a result of C. Mueller and E. Perkins. Probability Theory and Related Fields, 108(4):543–557, August 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050120.

Krylov:2000:RCF

[Kry00] N. V. Krylov. On the rate of convergence of finite-difference approximations for Bellmans equations with variable coefficients. Probability Theory and Related Fields, 117(1):1–16, May 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050264.

Krylov:2002:SCS

[Kry02] N. V. Krylov. A supermartingale characterization of sets of stochastic integrals and applications. Probability Theory and Related Fields, 123(4):521–552, August 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400100190.

Krylov:2003:OMS

[Kry03] N. V. Krylov. One more square root law for Brownian motion and its application to SPDEs. Probability Theory and Related Fields, 127(4):496–512, December 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0301-3.

Krylov:2010:IFP

[Kry10] N. V. Krylov. Itô’s formula for the L_{p}-norm of stochastic W_{p}^{1}- valued processes. Probability Theory and Related Fields, 147 (3–4):583–605, July 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0217-7.

Krylov:2011:IWF

[Kry11] N. V. Krylov. On the Itô–Wentzell formula for distributionvalued processes and related topics. Probability Theory and Related Fields, 150(1–2):295–319, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-010-0275-x.

Krylov:2014:DPP

[Kry14] N. V. Krylov. On the dynamic programming principle for uniformly nondegenerate stochastic differential games in domains and the Isaacs equations. Probability Theory and Related Fields, 158(3–4):751–783, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-013-0495-y.

Krylov:2015:HFP

[Kry15] N. V. Krylov. Hypoellipticity for filtering problems of partially observable diffusion processes. Probability Theory and Related Fields, 161(3–4):687–718, April 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0557-9.

Krylov:2016:SFT

[Kry16] N. V. Krylov. On singularity as a function of time of a conditional distribution of an exit time. Probability Theory and Related Fields, 165(3–4):541–557, August 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0639-3;   http://link.springer.com/
article/10.1007/s00440-015-0639-3.

Krylov:2021:DPD

[Kry21] N. V. Krylov. On diffusion processes with drift in L_{d}. Probability Theory and Related Fields, 179(1–2):165–199, February 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01007-3.

Konno:1988:SPD

[KS88a] N. Konno and T. Shiga. Stochastic partial differential equations for some measure-valued diffusions. Probability Theory and Related Fields, 79(2):201–225, September 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00320919.

Krakowiak:1988:SMIa

[KS88b] W. Krakowiak and J. Szulga. On a p-stable multiple integral. I. Probability Theory and Related Fields, 78(3):437–448, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00334205.

Krakowiak:1988:SMIb

[KS88c] W. Krakowiak and J. Szulga. On a p-stable multiple integral. II. Probability Theory and Related Fields, 78(3):449–453, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00334206.

Krakowiak:1988:HPR

[KS88d] Wieslaw Krakowiak and Jerzy Szulga. Hypercontraction principle and random multilinear forms. Probability Theory and Related Fields, 77(3):325–342, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00319292.

Kitada:1989:ACI

[KS89] Kazuhiro Kitada and Hiroshi Sato. On the absolute continuity of infinite product measure and its convolution. Probability Theory and Related Fields, 81(4):609–627, May 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00367307.

Kallenberg:1991:SDF

[KS91] Olav Kallenberg and Rafal Sztencel. Some dimension-free features of vector-valued martingales. Probability Theory and Related Fields, 88(2):215–247, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01212560.

Kamm:1994:ESS

[KS94] Barbara Kamm and Andreas Schief. Extension of stationary stochastic processes. Probability Theory and Related Fields, 100 (1):77–84, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01204954.

Kesten:1995:SGM

[KS95] Harry Kesten and Roberto H. Schonmann. On some growth models with a small parameter. Probability Theory and Related Fields, 101(4):435–468, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01202779.

Kesten:2000:ABC

[KS00] Harry Kesten and Zhong-Gen Su. Asymptotic behavior of the critical probability for ρ-percolation in high dimensions. Probability Theory and Related Fields, 117(3):419–447, July 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050013.

Kulik:2018:GCC

[KS18] Alexei Kulik and Michael Scheutzow. Generalized couplings and convergence of transition probabilities. Probability Theory and Related Fields, 171(1–2):333–376, June 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0779-8.

Kramli:1986:RWI

[KSS86] András Krámli, Nándor Simányi, and Domokos Szász. Random walks with internal degrees of freedom. Probability Theory and Related Fields, 72(4):603–617, July 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00344723.

Knapik:2016:BPA

[KSvdVvZ16] B. T. Knapik, B. T. Szabó, A. W. van der Vaart, and J. H. van Zanten. Bayes procedures for adaptive inference in inverse problems for the white noise model. Probability Theory and Related Fields, 164(3–4):771–813, April 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
015-0619-7;   http://link.springer.com/content/pdf/
10.1007/s00440-015-0619-7.pdf.

Kenyon:2016:ADT

[KSW16] Richard W. Kenyon, Nike Sun, and David B. Wilson. On the asymptotics of dimers on tori. Probability Theory and Related Fields, 166(3–4):971–1023, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0687-8;   http://link.springer.com/
article/10.1007/s00440-015-0687-8.

Khoshnevisan:2008:HDC

[KSX08] Davar Khoshnevisan, Narn-Rueih Shieh, and Yimin Xiao. Hausdorff dimension of the contours of symmetric additive Lévy processes. Probability Theory and Related Fields, 140(1–2): 129–167, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0060-7.

Khoshnevisan:2009:HDC

[KSX09] Davar Khoshnevisan, Narn-Rueih Shieh, and Yimin Xiao. Hausdorff dimension of the contours of symmetric additive Lévy processes. Probability Theory and Related Fields, 143(3–4):665–666, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00440-008-0184-4.pdf.

Klemela:2004:ECP

[KT04] Jussi Klemelä and Alexandre B. Tsybakov. Exact constants for pointwise adaptive estimation under the Riesz transform. Probability Theory and Related Fields, 129(3):441–467, July 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0348-9.

Katori:2007:ISN

[KT07] Makoto Katori and Hideki Tanemura. Infinite systems of noncolliding generalized meanders and Riemann–Liouville differintegrals. Probability Theory and Related Fields, 138(1–2):113–156, May 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-006-0015-4.

Kawazu:1989:LTO

[KTT89] Kiyoshi Kawazu, Yozo Tamura, and Hiroshi Tanaka. Limit theorems for one-dimensional diffusions and random walks in random environments. Probability Theory and Related Fields, 80(4):501–541, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00318905.

Kuwae:1997:WCS

[KU97] Kazuhiro Kuwae and Toshihiro Uemura. Weak convergence of symmetric diffusion processes. Probability Theory and Related Fields, 109(2):159–182, October 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050129.

Kulske:1998:REM

[Kül98] Christof Külske. A random energy model for size dependence: recurrence vs. transience. Probability Theory and Related Fields, 111(1):57–100, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050162.

Kulske:2001:WGR

[Kül01] Christof Külske. Weakly Gibbsian representations for joint measures of quenched lattice spin models. Probability Theory and Related Fields, 119(1):1–30, January 2001. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/PL00012737.

Kulske:2003:UBS

[Kül03] Christof Külske. Universal bounds on the selfaveraging of random diffraction measures. Probability Theory and Related Fields, 126(1):29–50, May 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0261-7.

Kumagai:1993:ETD

[Kum93] Takashi Kumagai. Estimates of transition densities for Brownian motion on nested fractals. Probability Theory and Related Fields, 96(2):205–224, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192133.

Kuwada:2009:LAS

[Kuw09] Kazumasa Kuwada. Laplace approximation for stochastic line integrals. Probability Theory and Related Fields, 144(1–2):1–51, May 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-008-0140-3.

Kleptsyna:2008:DTE

[KV08] M. L. Kleptsyna and A. Yu. Veretennikov. On discrete time ergodic filters with wrong initial data. Probability Theory and Related Fields, 141(3–4):411–444, July 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0089-7.

Kifer:2014:NLD

[KV14] Yuri Kifer and S. R. S. Varadhan. Nonconventional large deviations theorems. Probability Theory and Related Fields, 158(1–2): 197–224, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0481-4.

Kriecherbauer:2018:ESC

[KV18] Thomas Kriecherbauer and Martin Venker. Edge statistics for a class of repulsive particle systems. Probability Theory and Related Fields, 170(3–4):617–655, April 2018. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0765-1.

Klopp:2019:OGE

[KV19] Olga Klopp and Nicolas Verzelen. Optimal graphon estimation in cut distance. Probability Theory and Related Fields, 174(3–4): 1033–1090, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0878-1.

Kenyon:2004:CRN

[KW04] Richard W. Kenyon and David B. Wilson. Critical resonance in the non-intersecting lattice path model. Probability Theory and Related Fields, 130(3):289–318, November 2004. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0293-z.

Kassel:2015:TCP

[KW15] Adrien Kassel and Wei Wu. Transfer current and pattern fields in spanning trees. Probability Theory and Related Fields, 163(1– 2):89–121, October 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0588-2.

Kemppainen:2016:NSC

[KW16] Antti Kemppainen and Wendelin Werner. The nested simple conformal loop ensembles in the Riemann sphere. Probability Theory and Related Fields, 165(3–4):835–866, August 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0647-3;   http://link.springer.com/
article/10.1007/s00440-015-0647-3.

Kwasnicki:2024:FVC

[Kwa24] Mateusz Kwaśnicki. Fleming–Viot couples live forever. Probability Theory and Related Fields, 188(3–4):1385–1408, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01247-z.

Kwon:1992:SPB

[Kwo92] Youngmee Kwon. The submartingal problem for Brownian motion in a cone with non-constant oblique reflection. Probability Theory and Related Fields, 92(3):351–391, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01300561.

Kwon:1995:RBM

[Kwo95] Youngmee Kwon. Reflected Brownian motion in a cone: Semimartingale property. Probability Theory and Related Fields, 101 (2):211–226, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01375825.

Khoshnevisan:2009:HAA

[KX09] Davar Khoshnevisan and Yimin Xiao. Harmonic analysis of additive Lévy processes. Probability Theory and Related Fields, 145(3–4):459–515, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0175-5.

Kusuoka:1992:DFF

[KY92] Shigeo Kusuoka and Zhou Xian Yin. Dirichlet forms on fractals: Poincaré constant and resistance. Probability Theory and Related Fields, 93(2):169–196, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01195228.

Kusuoka:2000:MCG

[KY00] Shigeo Kusuoka and Nakahiro Yoshida. Malliavin calculus, geometric mixing, and expansion of diffusion functionals. Probability Theory and Related Fields, 116(4):457–484, April 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400070001.

Knowles:2013:EDW

[KY13] Antti Knowles and Jun Yin. Eigenvector distribution of Wigner matrices. Probability Theory and Related Fields, 155(3–4): 543–582, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0407-y.

Knowles:2017:ALL

[KY17] Antti Knowles and Jun Yin. Anisotropic local laws for random matrices. Probability Theory and Related Fields, 169(1–2):

257–352, October 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-016-0730-4.

Kyprianou:2000:MCS

[Kyp00] A. E. Kyprianou. Martingale convergence and the stopped branching random walk. Probability Theory and Related Fields, 116(3):405–419, March 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050256.

Kesten:1997:CLT

[KZ97] Harry Kesten and Yu Zhang. A central limit theorem for “critical” first-passage percolation in two dimensions. Probability Theory and Related Fields, 107(2):137–160, February 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050080.

Kusuoka:1998:WFL

[KZ98] Shigeo Kusuoka and Xian Yin Zhou. Waves on fractal-like manifolds and effective energy propagation. Probability Theory and Related Fields, 110(4):473–495, May 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050156.

Kosygina:2017:ZOL

[KZ17] Elena Kosygina and Martin P. W. Zerner. A zero–one law for recurrence and transience of frog processes. Probability Theory and Related Fields, 168(1–2):317–346, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Lacey:1990:LLL

[Lac90] Michael T. Lacey. Limit laws for local times of the Brownian sheet. Probability Theory and Related Fields, 86(1):63–85, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01207514.

Lacoin:2010:HPM

[Lac10] Hubert Lacoin. Hierarchical pinning model with site disorder: disorder is marginally relevant. Probability Theory and Related Fields, 148(1–2):159–175, September 2010. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-009-0226-6.

Lacoin:2014:NCQ

[Lac14] Hubert Lacoin. Non-coincidence of quenched and annealed connective constants on the supercritical planar percolation cluster. Probability Theory and Related Fields, 159(3–4):777–808, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-013-0520-1.

Lacker:2016:GCM

[Lac16] Daniel Lacker. A general characterization of the mean field limit for stochastic differential games. Probability Theory and Related Fields, 165(3–4):581–648, August 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0641-9;   http://link.springer.com/
article/10.1007/s00440-015-0641-9.

Lahiri:1999:ADE

[Lah99] S. N. Lahiri. Asymptotic distribution of the empirical spatial cumulative distribution function predictor and prediction bands based on a subsampling method. Probability Theory and Related Fields, 114(1):55–84, April 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050221.

Lahiri:2001:EBL

[Lah01] S. N. Lahiri. Effects of block lengths on the validity of block resampling methods. Probability Theory and Related Fields, 121 (1):73–97, September 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008798.

Lalley:1986:RTC

[Lal86] S. P. Lalley. Renewal theorem for a class of stationary sequences. Probability Theory and Related Fields, 72(2):195–213, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00699103.

Lalley:2009:SEC

[Lal09] Steven P. Lalley. Spatial epidemics: critical behavior in one dimension. Probability Theory and Related Fields, 144(3–4): 429–469, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0151-0.

Lambert:2002:GCS

[Lam02] Amaury Lambert. The genealogy of continuous-state branching processes with immigration. Probability Theory and Related Fields, 122(1):42–70, January 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400100155.

Lammers:2022:HFD

[Lam22] Piet Lammers. Height function delocalisation on cubic planar graphs. Probability Theory and Related Fields, 182(1–2): 531–550, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01087-9.

Lapeyre:1990:ATE

[Lap90] Bernard Lapeyre. Une application de la théorie des excursions à une diffusion réfléchie dégénérée. (French) [An application of the theory of excursions with reflected degenerate diffusion]. Probability Theory and Related Fields, 87(2):189–207, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01198429.

Laroche:1995:HPS

[Lar95] Etienne Laroche. Hypercontractivité pour des systèmes de spins de portée infinie. (French) [Hypercontractivity for spin systems with infinite range]. Probability Theory and Related Fields, 101 (1):89–132, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192197.

Last:1994:PPP

[Las94] Günter Last. Predictable projections for point process filtrations. Probability Theory and Related Fields, 99(3):361–388, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01199897.

Laslier:2019:LLL

[Las19] Benoît Laslier. Local limits of lozenge tilings are stable under bounded boundary height perturbations. Probability Theory and Related Fields, 173(3–4):1243–1264, April 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0853-x.

Lawler:1992:EPS

[Law92] Gregory F. Lawler. Escape probabilities for slowly recurrent sets. Probability Theory and Related Fields, 94(1):91–117, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01222512.

Le:1995:ICT

[LB95] Huiling Le and Dennis Barden. Itô correction terms for the radial parts of semimartingales on manifolds. Probability Theory and Related Fields, 101(1):133–146, ???? 1995. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192198.

Le:2008:SGI

[LB08] H. Le and D. Barden. Semimartingales and geometric inequalities on locally symmetric manifolds. Probability Theory and Related Fields, 142(1–2):285–311, September 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0105-y.

Lei:2018:AHD

[LBE18] Lihua Lei, Peter J. Bickel, and Noureddine El Karoui. Asymptotics for high dimensional regression M-estimates: fixed design results. Probability Theory and Related Fields, 172(3–4): 983–1079, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0824-7.

Li:2002:ASS

[LDS02] C. W. Li, Z. Dong, and R. Situ. Almost sure stability of linear stochastic differential equations with jumps. Probability Theory and Related Fields, 123(1):121–155, May 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200198.

LeGall:1987:MBC

[Le 87] Jean-François Le Gall. Mouvement brownien, cônes et processus stables. (French) [Brownian motion, cones, and stables processes]. Probability Theory and Related Fields, 76(4):587–627, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00960076.

LeGall:1988:CMK

[Le 88] J.-F. Le Gall. Sur une conjecture de M. Kac. (French) [On a conjecture of M. Kac]. Probability Theory and Related Fields, 78(3):389–402, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00334202.

LeGall:1993:CPV

[Le 93a] J. F. Le Gall. A class of path-valued Markov processes and its applications to superprocesses. Probability Theory and Related Fields, 95(1):25–46, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01197336.

LeGall:1993:URT

[Le 93b] Jean-François Le Gall. The uniform random tree in a Brownian excursion. Probability Theory and Related Fields, 96(3):369–383, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01292678.

LeGall:1995:BSS

[Le 95a] Jean-François Le Gall. The Brownian snake and solutions of Δu = u^{2} in a domain. Probability Theory and Related Fields, 102(3):393–432, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192468.

LeJan:1995:FEB

[Le 95b] Y. Le Jan. Free energy for Brownian and geodesic homology. Probability Theory and Related Fields, 102(1):57–61, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01295221.

LeGall:2015:BCI

[Le 15] Jean-François Le Gall. The Brownian cactus II: upcrossings and local times of super-Brownian motion. Probability Theory and Related Fields, 162(1–2):199–231, June 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0569-5.

LeGall:2018:STB

[Le 18] Jean-François Le Gall. Subordination of trees and the Brownian map. Probability Theory and Related Fields, 171(3–4):819–864, August 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-017-0794-9.

Le:1996:LAB

[Le96] Huiling Le. Limiting angle of Brownian motion on certain manifolds. Probability Theory and Related Fields, 106(1):137– 149, September 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050060.

Le:1999:LAM

[Le99] Huiling Le. Limiting angles of Γ-martingales. Probability Theory and Related Fields, 114(1):85–96, April 1999. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050222.

Leandre:1987:IFA

[Léa87a] Rémi Léandre. Integration dans la fibre associée à une diffusion dégénérée. (French) [Integration in the fiber associated with a degenerate distribution]. Probability Theory and Related Fields, 76(3):341–358, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01297490.

Leandre:1987:MTP

[Léa87b] Rémi Léandre. Majoration en temps petit de la densité d’une diffusion dégénérée. (French) [Increase in small time density of degenerate diffusion]. Probability Theory and Related Fields, 74 (2):289–294, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00569994.

Leandre:1988:TAS

[Léa88] Rémi Léandre. Sur le théoreme d’Atiyah–Singer. (French) [On the Atiyah–Singer theorem]. Probability Theory and Related Fields, 80(1):119–137, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00348755.

Leandre:2001:SCC

[Léa01] Rémi Léandre. Stochastic cohomology of Chen–Souriau and line bundle over the Brownian bridge. Probability Theory and Related Fields, 120(2):168–182, June 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/PL00008779.

Leble:2017:LMB

[Leb17] Thomas Leblé. Local microscopic behavior for 2D Coulomb gases. Probability Theory and Related Fields, 169(3–4):931– 976, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-016-0744-y.

Leduc:2000:CCG

[Led00] Guillaume Leduc. The complete characterization of a general class of superprocesses. Probability Theory and Related Fields, 116(3):317–358, March 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050252.

Lee:1990:SLT

[Lee90] T.-Y. Lee. Some limit theorems for critical branching Bessel processes, and related semilinear differential equations. Probability Theory and Related Fields, 84(4):505–520, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198317.

Lehner:2003:CNP

[Leh03] Franz Lehner. Cumulants in noncommutative probability theory II. Probability Theory and Related Fields, 127(3):407– 422, November 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0292-0.

Lejay:2001:HDF

[Lej01] Antoine Lejay. Homogenization of divergence-form operators with lower-order terms in random media. Probability Theory and Related Fields, 120(2):255–276, June 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/PL00008783.

Lemanczyk:1988:WIT

[Lem88] M. Lemańczyk. Weakly isomorphic transformations that are not isomorphic. Probability Theory and Related Fields, 78(4): 491–507, August 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00353873.

Leonard:1995:LDP

[Léo95] C. Léonard. On large deviations for particle systems associated with spatially homogeneous Boltzmann type equations. Probability Theory and Related Fields, 101(1):1–44, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192194.

Leonard:2022:FKF

[Léo22] Christian Léonard. Feynman–Kac formula under a finite entropy condition. Probability Theory and Related Fields, 184 (3–4):1029–1091, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-022-01155-8.

Leuridan:2010:PPA

[Leu10] Christophe Leuridan. Un processus ponctuel associé aux maxima locaux du mouvement brownien. (French) [A point process associated with local maxima of Brownian motion]. Probability Theory and Related Fields, 148(3–4):457–477, November 2010.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-009-0236-4.

Levental:1988:ULT

[Lev88] Shlomo Levental. Uniform limit theorems for Harris recurrent Markov chains. Probability Theory and Related Fields, 80(1): 101–118, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00348754.

Li:1992:LIR

[Li92] Wenbo V. Li. Lim inf results for the Wiener process and its increments under the L_{2}-norm. Probability Theory and Related Fields, 92(1):69–90, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01205237.

Li:1994:SDE

[Li94a] Xue-Mei Li. Stochastic differential equations on noncompact manifolds: moment stability and its topological consequences. Probability Theory and Related Fields, 100(4):417–428, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01268987.

Li:1994:SCS

[Li94b] Xue-Mei Li. Strong p-completeness of stochastic differential equations and the existence of smooth flows on noncompact manifolds. Probability Theory and Related Fields, 100 (4):485–511, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01268991.

Li:2003:SSC

[Li03] Xiang Dong Li. Sobolev spaces and capacities theory on path spaces over a compact Riemannian manifold. Probability Theory and Related Fields, 125(1):96–134, January 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200227.

Li:2008:MTP

[Li08] Xiang-Dong Li. Martingale transforms and L^{p}-norm estimates of Riesz transforms on complete Riemannian manifolds. Probability Theory and Related Fields, 141(1–2):247–281, May 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/

s00440-007-0085-y. See erratum [Li14a].

Li:2011:WPH

[Li11] Xiang-Dong Li. On the weak L^{p}-hodge decomposition and Beurling–Ahlfors transforms on complete Riemannian manifolds. Probability Theory and Related Fields, 150(1–2):111–144, June 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-010-0270-2. See erratum [Li14b].

Li:2014:EMT

[Li14a] Xiang-Dong Li. Erratum to: Martingale transforms and L^{p}- norm estimates of Riesz transforms on complete Riemannian manifolds. Probability Theory and Related Fields, 159(1–2):405– 408, June 2014. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s00440-014-0560-1.pdf. See [Li08].

Li:2014:EWL

[Li14b] Xiang-Dong Li. Erratum to: On the weak L^{p} Hodge decomposition and Beurling–Ahlfors transforms on complete Riemannian manifolds. Probability Theory and Related Fields, 159(1–2):409– 411, June 2014. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s00440-014-0561-0.pdf. See [Li11].

Li:2016:LRD

[Li16] Xue-Mei Li. Limits of random differential equations on manifolds. Probability Theory and Related Fields, 166(3–4):659– 712, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0669-x;
http://link.springer.com/article/10.1007/s00440-015-
0669-x.

Li:2019:CIF

[Li19] Jhih-Huang Li. Conformal invariance in the FK-representation of the quantum Ising model and convergence of the interface to the SLE_{16/3}. Probability Theory and Related Fields, 173(1–2): 87–156, February 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0831-3.

Liao:1993:SFB

[Lia93] Ming Liao. Stochastic flows on the boundaries of SL(n, R). Probability Theory and Related Fields, 96(2):261–281, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192136.

Liao:2000:DSF

[Lia00] Ming Liao. Decomposition of stochastic flows and Lyapunov exponents. Probability Theory and Related Fields, 117(4):589–607, August 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/PL00008736.

Liao:2022:MPD

[Lia22] Yuchen Liao. Multi-point distribution of discrete time periodic TASEP. Probability Theory and Related Fields, 182(3–4): 1053–1131, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01107-8.

Liero:1989:SUC

[Lie89] Hannelore Liero. Strong uniform consistency of nonparametric regression function estimates. Probability Theory and Related Fields, 82(4):587–614, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00341285.

Liebscher:1998:DVQ

[Lie98] Volkmar Liebscher. Diagonal versions and quantum stochastic integrals on the symmetric Fock space with nonadapted integrands. Probability Theory and Related Fields, 112(2): 255–295, October 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050190.

Liggett:1987:ADP

[Lig87] Thomas M. Liggett. Applications of the Dirichlet principle to finite reversible nearest particle systems. Probability Theory and Related Fields, 74(4):505–528, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00363513.

Liggett:1996:BRW

[Lig96] Thomas M. Liggett. Branching random walks and contact processes on homogeneous trees. Probability Theory and Related Fields, 106(4):495–519, December 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400050073.

Liming:1999:UND

[Lim99] Wu Liming. Uniqueness of Nelson’s diffusions. Probability Theory and Related Fields, 114(4):549–585, July 1999. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050234.

Lindsay:1986:FM

[Lin86] J. M. Lindsay. Fermion martingales. Probability Theory and Related Fields, 71(2):307–320, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00332314.

Lin:1987:CDM

[Lin87] G. D. Lin. On characterizations of distributions via moments of record values. Probability Theory and Related Fields, 74 (4):479–483, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00363510.

Liorit:2005:LMI

[Lio05] Grégory Liorit. Laplace’s method for iterated complex Brownian integrals. Probability Theory and Related Fields, 133(1): 18–42, September 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0409-0.

Liptser:1996:LDT

[Lip96] Robert Liptser. Large deviations for two scaled diffusions. Probability Theory and Related Fields, 106(1):71–104, September 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050058.

Liu:1996:EHD

[Liu96] Quanshing Liu. The exact Hausdorff dimension of a branching set. Probability Theory and Related Fields, 104(4):515–538, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01198165.

Liu:2017:NCQ

[Liu17] Yuan Liu. A new characterization of quadratic transportationinformation inequalities. Probability Theory and Related Fields, 168(3–4):675–689, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-016-0721-5.

Liu:2022:MMV

[Liu22a] Jingbo Liu. Minoration via mixed volumes and Cover’s problem for general channels. Probability Theory and Related Fields, 183 (1–2):315–357, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01111-6.

Liu:2022:OPD

[Liu22b] Zhipeng Liu. One-point distribution of the geodesic in directed last passage percolation. Probability Theory and Related Fields, 184(1–2):425–491, October 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-022-01123-2.

LeJan:2004:PBM

[LL04] Y. Le Jan and S. Lemaire. Products of beta matrices and sticky flows. Probability Theory and Related Fields, 130(1): 109–134, September 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0358-7.

LeGouic:2017:ECW

[LL17] Thibaut Le Gouic and Jean-Michel Loubes. Existence and consistency of Wasserstein barycenters. Probability Theory and Related Fields, 168(3–4):901–917, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
016-0727-z.

Lacker:2023:SUT

[LL23] Daniel Lacker and Luc Le Flem. Sharp uniform-in-time propagation of chaos. Probability Theory and Related Fields, 187(1– 2):443–480, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01192-x.

Li:2024:ELD

[LL24] Songzi Li and Xiang-Dong Li. W -entropy and Langevin deformation on Wasserstein space over Riemannian manifolds. Probability Theory and Related Fields, 188(3–4):911–955, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01256-y.

Levin:2010:GDM

[LLP10] David A. Levin, Malwina J. Luczak, and Yuval Peres. Glauber dynamics for the mean-field Ising model: cut-off, critical power law, and metastability. Probability Theory and Related Fields, 146(1–2):223–265, January 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0189-z.

Li:2010:NSC

[LLR10] Deli Li, Wei-Dong Liu, and Andrew Rosalsky. Necessary and sufficient conditions for the asymptotic distribution of the largest entry of a sample correlation matrix. Probability Theory and Related Fields, 148(1–2):5–35, September 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
009-0220-z.

Landon:2024:SEF

[LLS24] Benjamin Landon, Patrick Lopatto, and Philippe Sosoe. Single eigenvalue fluctuations of general Wigner-type matrices.

Probability Theory and Related Fields, 188(1–2):1–62, February 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01181-6.

Litvak:2019:SSV

[LLT^{+}19] Alexander E. Litvak, Anna Lytova, Konstantin Tikhomirov, Nicole Tomczak-Jaegermann, and Pierre Youssef. The smallest singular value of a shifted d-regular random square matrix. Probability Theory and Related Fields, 173(3–4):1301–1347, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-018-0852-y.

Lamarre:2023:KES

[LLT23] Pierre Yves Gaudreau Lamarre, Yier Lin, and Li-Cheng Tsai. KPZ equation with a small noise, deep upper tail and limit shape. Probability Theory and Related Fields, 185(3–4):885– 920, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-022-01185-2.

LeBreton:1989:LLN

[LM89] A. Le Breton and M. Musiela. Laws of large numbers for semimartingales with applications to stochastic regression. Probability Theory and Related Fields, 81(2):275–290, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00319555.

LeGall:1992:PCD

[LM92] J. F. Le Gall and T. Meyre. Points cônes du mouvement brownien plan, le cas critique. (French) [Conic points of planar Brownian motion, the critical case]. Probability Theory and Related Fields, 93(2):231–247, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01195230.

Letac:2000:MCA

[LM00] Gérard Letac and Dhafer Malouche. The Markov chain associated to a pick function. Probability Theory and Related Fields, 118(4):439–454, December 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/PL00008750.

Leandre:2001:SFD

[LM01] Rémi Léandre and Salah-Eldin A. Mohammed. Stochastic functional differential equations on manifolds. Probability Theory and Related Fields, 121(1):117–135, September 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008795.

Lecue:2009:AER

[LM09] Guillaume Lecué and Shahar Mendelson. Aggregation via empirical risk minimization. Probability Theory and Related Fields, 145(3–4):591–613, November 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0180-8.

Lewicka:2017:OPL

[LM17] Marta Lewicka and Juan J. Manfredi. The obstacle problem for the p-Laplacian via optimal stopping of tug-of-war games. Probability Theory and Related Fields, 167(1–2):349– 378, February 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0684-y;
http://link.springer.com/article/10.1007/s00440-015-
0684-y.

Lelarge:2019:FLS

[LM19a] Marc Lelarge and Léo Miolane. Fundamental limits of symmetric low-rank matrix estimation. Probability Theory and Related Fields, 173(3–4):859–929, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-018-0845-x.

Lugosi:2019:NOM

[LM19b] Gábor Lugosi and Shahar Mendelson. Near-optimal mean estimators with respect to general norms. Probability Theory and Related Fields, 175(3–4):957–973, December 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00906-4.

Lim:2022:MEP

[LM22] Tongseok Lim and Robert J. McCann. Maximizing expected powers of the angle between pairs of points in projective space.

Probability Theory and Related Fields, 184(3–4):1197–1214, December 2022. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-022-01108-1.

Lichev:2024:NRR

[LM24] Lyuben Lichev and Dieter Mitsche. New results for the random nearest neighbor tree. Probability Theory and Related Fields, 189(1–2):229–279, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-024-01268-2.

Luczak:2003:LRR

[LMU03] Malwina J. Luczak, Colin McDiarmid, and Eli Upfal. On-line routing of random calls in networks. Probability Theory and Related Fields, 125(4):457–482, April 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-002-0242-2.

Lorinczi:1998:TGM

[LMV98] József Lörinczi, Christian Maes, and Koen Vande Velde. Transformations of Gibbs measures. Probability Theory and Related Fields, 112(1):121–147, September 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050185.

Lo:1989:DHR

[LMW89] S. H. Lo, Y. P. Mack, and J. L. Wang. Density and hazard rate estimation for censored data via strong representation of the Kaplan–Meier estimator. Probability Theory and Related Fields, 80(3):461–473, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01794434.

Landim:2005:PLS

[LN05] C. Landim and J. Noronha Neto. Poincaré and logarithmic Sobolev inequality for Ginzburg–Landau processes in random environment. Probability Theory and Related Fields, 131(2): 229–260, February 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0370-y.

Lu:2015:ERT

[LN15a] Jianfeng Lu and James Nolen. Erratum to: “Reactive trajectories and the transition path process”. Probability Theory and Related Fields, 161(1–2):245, February 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-014-0558-8.pdf. See [LN15b].

Lu:2015:RTT

[LN15b] Jianfeng Lu and James Nolen. Reactive trajectories and the transition path process. Probability Theory and Related Fields, 161(1–2):195–244, February 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0547-y. See er-

ratum [LN15a].

Licea:1996:SFP

[LNP96] C. Licea, C. M. Newman, and M. S. T. Piza. Superdiffusivity in first-passage percolation. Probability Theory and Related Fields, 106(4):559–591, December 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400050075.

Lepski:1999:ENR

[LNS99] O. Lepski, A. Nemirovski, and V. Spokoiny. On estimation of the L_{r} norm of a regression function. Probability Theory and Related Fields, 113(2):221–253, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004409970006.

Lo:1989:SRB

[Lo89] Shaw-Hwa Lo. On some representations of the Bootstrap. Probability Theory and Related Fields, 82(3):411–418, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00339995.

Liverani:1996:EIH

[LO96] Carlangelo Liverani and Stefano Olla. Ergodicity in infinite Hamiltonian systems with conservative noise. Probability Theory and Related Fields, 106(3):401–445, November 1996. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050071.

Lammers:2024:DAV

[LO24] Piet Lammers and Sébastien Ott. Delocalisation and absolutevalue-FKG in the solid-on-solid model. Probability Theory and Related Fields, 188(1–2):63–87, February 2024. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-023-01202-y.

Loukianova:1997:LBR

[Lou97] Daria Loukianova. Lower bounds on the restitution error in the Hopfield model. Probability Theory and Related Fields, 107(2): 161–176, February 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050081.

Louf:2022:PNS

[Lou22] Baptiste Louf. Planarity and non-separating cycles in uniform high genus quadrangulations. Probability Theory and Related Fields, 182(3–4):1183–1206, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01050-
8.

Loviscach:1995:GNV

[Lov95] J. Loviscach. A generalized notion of variation applied to Markov chains and Anosov maps. Probability Theory and Related Fields, 103(4):553–570, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01246339.

Landim:1998:CDE

[LOY98] C. Landim, S. Olla, and H. T. Yau. Convection-diffusion equation with space–time ergodic random flow. Probability Theory and Related Fields, 112(2):203–220, October 1998. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050187.

Last:2011:PPF

[LP11a] Günter Last and Mathew D. Penrose. Poisson process Fock space representation, chaos expansion and covariance inequalities. Probability Theory and Related Fields, 150(3–4):663–690, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-010-0288-5.

Lavancier:2011:SCR

[LP11b] Frédéric Lavancier and Anne Philippe. Some convergence results on quadratic forms for random fields and application to empirical covariances. Probability Theory and Related Fields, 149(3– 4):493–514, April 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0262-2.

Ledoit:2011:ESL

[LP11c] Olivier Ledoit and Sandrine Péché. Eigenvectors of some large sample covariance matrix ensembles. Probability Theory and Related Fields, 151(1–2):233–264, October 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0298-3.

Lambert:2019:LLN

[LP19] Gaultier Lambert and Elliot Paquette. The law of large numbers for the maximum of almost Gaussian log-correlated fields coming from random matrices. Probability Theory and Related Fields, 173(1–2):157–209, February 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0832-2.

Lyons:1996:BRW

[LPP96] Russell Lyons, Robin Pemantle, and Yuval Peres. Biased random walks on Galton–Watson trees. Probability Theory and Related Fields, 106(2):249–264, October 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050064.

LePage:1997:SCI

[LPR97] Raoul LePage, Krzysztof Podgórski, and Michal Ryznar. Strong and conditional invariance principles for samples attracted to stable laws. Probability Theory and Related Fields, 108(2): 281–298, June 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050110.

Last:2016:NAP

[LPS16] Günter Last, Giovanni Peccati, and Matthias Schulte. Normal approximation on Poisson spaces: Mehler’s formula, second order Poincaré inequalities and stabilization. Probability Theory and Related Fields, 165(3–4):667–723, August 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0643-7;   http://link.springer.com/
article/10.1007/s00440-015-0643-7.

Le:2013:TUA

[LPW13] V. Le, E. Pardoux, and A. Wakolbinger. “trees under attack”: a Ray–Knight representation of Feller’s branching diffusion with logistic growth. Probability Theory and Related Fields, 155(3– 4):583–619, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0408-x.

Lyons:1997:SJF

[LQ97] Terry Lyons and Zhongmin Qian. Stochastic Jacobi fields and vector fields induced by varying area on path spaces. Probability Theory and Related Fields, 109(4):539–570, November 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050141.

Lyons:1998:FDI

[LQ98] Terry Lyons and Zhongmin Qian. Flow of diffeomorphisms induced by a geometric multiplicative functional. Probability Theory and Related Fields, 112(1):91–119, September 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050184.

LeJan:2013:SSB

[LQ13] Yves Le Jan and Zhongmin Qian. Stratonovich’s signatures of Brownian motion determine Brownian sample paths. Probability Theory and Related Fields, 157(1–2):209–223, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-012-0454-z.

Liu:2019:PTE

[LQ19] Xuan Liu and Zhongmin Qian. Parabolic type equations associated with the Dirichlet form on the Sierpinski gasket. Probability Theory and Related Fields, 175(3–4):1063–1098, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-019-00910-8;   https://link.springer.
com/content/pdf/10.1007/s00440-019-00910-8.pdf.

Lagodowski:1986:RCS

[LR86] Z. A. Lagodowski and Z. Rychlik. Rate of convergence in the strong law of large numbers for martingales. Probability Theory and Related Fields, 71(3):467–476, ???? 1986. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01000217.

Lalley:1988:SSC

[LR88] Steven Lalley and Herbert Robbins. Stochastic search in a convex region. Probability Theory and Related Fields, 77(1):99–116, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01848133.

Leandre:1990:EVP

[LR90] Rémi Léandre and Francesco Russo. Estimation de Varadhan pour des diffusions à deux paramètres. (French) [Varadhan estimation for diffusion with two parameters]. Probability Theory and Related Fields, 84(4):429–451, ???? 1990. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198313.

Lee:1994:OTS

[LR94] Tzong-Yow Lee and Bruno Remillard. Occupation times in systems of null recurrent Markov processes. Probability Theory and Related Fields, 98(2):245–259, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01192516.

LeJan:2004:SFC

[LR04] Yves Le Jan and Olivier Raimond. Sticky flows on the circle and their noises. Probability Theory and Related Fields, 129(1):63–82, May 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0324-9.

LeGall:2021:SRN

[LR21] Jean-François Le Gall and Armand Riera. Spine representations for non-compact models of random geometry. Probability Theory and Related Fields, 181(1-3):571–645, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01069-x.

Lacoin:2023:PAU

[LRV23] Hubert Lacoin, Rémi Rhodes, and Vincent Vargas. A probabilistic approach of ultraviolet renormalization in the boundary Sine–Gordon model. Probability Theory and Related Fields, 185 (1–2):1–40, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01174-5.

Lin:1994:LTR

[LRW94] Michael Lin, Ben-Zion Rubshtein, and Rainer Wittmann. Limit theorems for random walks with dynamical random transitions. Probability Theory and Related Fields, 100(3):285–300, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01193702.

Lacker:2023:MDI

[LRW23] Daniel Lacker, Kavita Ramanan, and Ruoyu Wu. Marginal dynamics of interacting diffusions on unimodular Galton–Watson trees. Probability Theory and Related Fields, 187(3–4):817–884, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01226-4.

Lo:1986:PLE

[LS86] Shaw-Hwa Lo and Kesar Singh. The product-limit estimator and the bootstrap: Some asymptotic representations. Probability Theory and Related Fields, 71(3):455–465, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01000216.

Lemanczyk:1987:CLR

[LS87a] Mariusz Lemańczyk and Andrzej Sikorski. A class of not local rank one automorphisms arising from continuous substitutions. Probability Theory and Related Fields, 76(4):421–428, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00960066.

LePage:1987:AIL

[LS87b] Raoul LePage and Bertram M. Schreiber. Addendum to: An iterated logarithm law for families of Brownian paths. Probability Theory and Related Fields, 76(4):629, ???? 1987. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00960077.

Lebowitz:1988:PFE

[LS88] J. L. Lebowitz and R. H. Schonmann. Pseudo-free energies and large deviations for non-Gibbsian FKG measures. Probability Theory and Related Fields, 77(1):49–64, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01848130.

Luckhaus:1989:MAN

[LS89] S. Luckhaus and W. Sauermann. Multinomial approximations for nonparametric experiments which minimize the maximal loss of Fisher information. Probability Theory and Related Fields, 81 (2):159–184, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319548.

Lalley:1997:HBB

[LS97a] Steven P. Lalley and Tom Sellke. Hyperbolic branching Brownian motion. Probability Theory and Related Fields, 108(2): 171–192, June 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050106.

LeJan:1997:SCD

[LS97b] Y. Le Jan and A. S. Sznitman. Stochastic cascades and 3-dimensional Navier–Stokes equations. Probability Theory and Related Fields, 109(3):343–366, November 1997. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050135.

Lalley:2000:EHT

[LS00] Steven P. Lalley and Thomas Sellke. An extension of Hawkes theorem on the Hausdorff dimension of a Galton–Watson tree. Probability Theory and Related Fields, 116(1):41–56, January 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008722.

Last:2001:SDS

[LS01a] G. Last and R. Schassberger. On the second derivative of the spherical contact distribution function of smooth grain models. Probability Theory and Related Fields, 121(1):49–72, September 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008797.

Li:2001:CTB

[LS01b] Wenbo V. Li and Qi-Man Shao. Capture time of Brownian pursuits. Probability Theory and Related Fields, 121(1): 30–48, September 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008796.

Li:2002:NCI

[LS02] Wenbo V. Li and Qi-Man Shao. A normal comparison inequality and its applications. Probability Theory and Related Fields, 122 (4):494–508, April 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400100176.

Lalley:2015:MDC

[LS15a] Steven P. Lalley and Yuan Shao. On the maximal displacement of critical branching random walk. Probability Theory and Related Fields, 162(1–2):71–96, June 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-014-0566-8.

Li:2015:LDO

[LS15b] Xinyi Li and Alain-Sol Sznitman. Large deviations for occupation time profiles of random interlacements. Probability Theory and Related Fields, 161(1–2):309–350, February 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0550-3.

Lee:2016:EEE

[LS16] Ji Oon Lee and Kevin Schnelli. Extremal eigenvalues and eigenvectors of deformed Wigner matrices. Probability Theory and Related Fields, 164(1–2):165–241, February 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0610-8.

Lee:2018:LLT

[LS18] Ji Oon Lee and Kevin Schnelli. Local law and Tracy–Widom limit for sparse random matrices. Probability Theory and Related Fields, 171(1–2):543–616, June 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0787-8.

Levine:2019:HLD

[LS19] Lionel Levine and Vittoria Silvestri. How long does it take for Internal DLA to forget its initial profile? Probability Theory and Related Fields, 174(3–4):1219–1271, August 2019. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0880-7;   http://link.springer.com/content/pdf/
10.1007/s00440-018-0880-7.pdf.

Lambert:2020:CTE

[LS20] Amaury Lambert and Emmanuel Schertzer. Coagulationtransport equations and the nested coalescents. Probability Theory and Related Fields, 176(1–2):77–147, February 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00914-4.

Lubetzky:2021:FIC

[LS21] Eyal Lubetzky and Allan Sly. Fast initial conditions for Glauber dynamics. Probability Theory and Related Fields, 181(1-3):647– 667, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-020-01015-3.

Lee:2022:NRM

[LS22] Jungkyoung Lee and Insuk Seo. Non-reversible metastable diffusions with Gibbs invariant measure I: Eyring–Kramers formula. Probability Theory and Related Fields, 182(3–4):849–903, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01102-z.

Lelli:2024:MTR

[LS24] Andrea Lelli and Alexandre Stauffer. Mixing time of random walk on dynamical random cluster. Probability Theory and Related Fields, 189(3–4):981–1043, August 2024. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-024-01262-8.

Lupu:2020:FML

[LST20] Titus Lupu, Christophe Sabot, and Pierre Tarrès. Fine mesh limit of the VRJP in dimension one and Bass–Burdzy flow. Probability Theory and Related Fields, 177(1–2):55–90, June 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-019-00944-y.

Losert:1986:UDS

[LT86] V. Losert and R. F. Tichy. On uniform distribution of subsequences. Probability Theory and Related Fields, 72(4):517–528, July 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00344719.

Ledoux:1988:CPB

[LT88] M. Ledoux and M. Talagrand. Un critère sur les petites boules dans le théorème limite central. (French) [A criterion on small balls in the Central Limit Theorem]. Probability Theory and Related Fields, 77(1):29–47, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01848129.

Lepski:2000:AEN

[LT00] O. V. Lepski and A. B. Tsybakov. Asymptotically exact nonparametric hypothesis testing in sup-norm and at a fixed point. Probability Theory and Related Fields, 117(1):17–48, May 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050265.

Laslier:2015:HQC

[LT15] Benoît Laslier and Fabio Lucio Toninelli. How quickly can we sample a uniform domino tiling of the 2L × 2L square via Glauber dynamics? Probability Theory and Related Fields, 161 (3–4):509–559, April 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0553-0.

Lytova:2020:DER

[LT20] Anna Lytova and Konstantin Tikhomirov. On delocalization of eigenvectors of random non-Hermitian matrices. Probability Theory and Related Fields, 177(1–2):465–524, June 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00956-8.

Lees:2021:EDT

[LT21] Benjamin Lees and Lorenzo Taggi. Exponential decay of transverse correlations for O(N) spin systems and related models. Probability Theory and Related Fields, 180(3–4):1099–1133, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01053-5.

Lefevre:2003:ENS

[LU03] Claude Lefèvre and Sergey Utev. Exact norms of a Stein-type operator and associated stochastic orderings. Probability Theory and Related Fields, 127(3):353–366, November 2003. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0288-9.

Luczak:1991:OSO

[Luc91] Andrzej Luczak. Operator-semistable operator Lévy’s measures on finite dimensional vector spaces. Probability Theory and Related Fields, 90(3):317–340, ???? 1991. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01193749.

Lucas:2014:LSD

[Luc14] Cyrille Lucas. The limiting shape for drifted internal diffusion limited aggregation is a true heat ball. Probability Theory and Related Fields, 159(1–2):197–235, June 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0505-0.

Luschgy:1992:CLM

[Lus92a] Harald Luschgy. Comparison of location models for stochastic processes. Probability Theory and Related Fields, 93(1):39–66, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01195387.

Luschgy:1992:LAM

[Lus92b] Harald Luschgy. Local asymptotic mixed normality for semimartingale experiments. Probability Theory and Related Fields, 92(2):151–176, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01194919.

Liu:1986:LDM

[LV86] Regina Y. Liu and John Van Ryzin. The limiting distribution of the maximal deviation of a density estimate and a hazard rate estimate. Probability Theory and Related Fields, 72(4):505–516, July 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00344718.

Landim:1996:EER

[LV96] Claudio Landim and M. Eulália Vares. Exponential estimate for reaction diffusion models. Probability Theory and Related Fields, 106(2):151–186, October 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050061.

Levina:2012:PEC

[LV12] Elizaveta Levina and Roman Vershynin. Partial estimation of covariance matrices. Probability Theory and Related Fields, 153(3–4):405–419, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-011-0349-4.

Legut:1990:OPM

[LW90] Jerzy Legut and Maciej Wilczynski. Optimal partitioning of a measurable space into countably many sets. Probability Theory and Related Fields, 86(4):551–558, ???? 1990. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198174.

Lindsay:2000:EPC

[LW00] J. Martin Lindsay and Stephen J. Wills. Existence, positivity and contractivity for quantum stochastic flows with infinite dimensional noise. Probability Theory and Related Fields, 116 (4):505–543, April 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050261.

Lawler:2004:BLS

[LW04a] Gregory F. Lawler and Wendelin Werner. The Brownian loop soup. Probability Theory and Related Fields, 128(4):565–588, April 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0319-6.

Linusson:2004:PPC

[LW04b] Svante Linusson and Johan Wästlund. A proof of Parisi’s conjecture on the random assignment problem. Probability Theory and Related Fields, 128(3):419–440, March 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0308-9.

Lupu:2018:RPM

[LW18] Titus Lupu and Wendelin Werner. The random pseudo-metric on a graph defined via the zero-set of the Gaussian free field on its metric graph. Probability Theory and Related Fields, 171(3– 4):775–818, August 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0792-y.

Lyons:2020:SPR

[LW20] Russell Lyons and Graham White. A stationary planar random graph with singular stationary dual: dyadic lattice graphs. Probability Theory and Related Fields, 176(3–4):1011–1043, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-019-00934-0.

Li:1992:ASB

[LWR92] Deli Li, Xiangchen Wang, and M. Bhaskara Rao. Almost sure behaviour of f-valued random fields. Probability Theory and Related Fields, 93(3):393–405, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01193058.

Li:2004:DSP

[LWX04] Zenghu Li, Hao Wang, and Jie Xiong. A degenerate stochastic partial differential equation for superprocesses with singular interaction. Probability Theory and Related Fields, 130(1): 1–17, September 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0313-z.

Li:2012:JCS

[LWXZ12] Zenghu Li, Hao Wang, Jie Xiong, and Xiaowen Zhou. Joint continuity of the solutions to a class of nonlinear SPDEs. Probability Theory and Related Fields, 153(3–4):441–469, August 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0351-x.

LeGall:1986:EAC

[LY86] J. F. Le Gall and M. Yor. Étude asymptotique de certains mouvements browniens complexes avec drift. (French) [Asymptotic study of certain complex Brownian motion with drift]. Probability Theory and Related Fields, 71(2):183–229, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00332310.

LeGall:1987:EAE

[LY87] J. F. Le Gall and M. Yor. Étude asymptotique des enlacements du mouvement brownien autour des droites de l’espace. (French)

[Asymptotic study of the embraces of Brownian motion around the rights of space]. Probability Theory and Related Fields, 74 (4):617–635, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00363519.

Ledrappier:1988:EFR

[LY88] F. Ledrappier and L.-S. Young. Entropy formula for random transformations. Probability Theory and Related Fields, 80(2): 217–240, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00356103.

Landim:1997:FDE

[LY97] C. Landim and H. T. Yau. Fluctuation-dissipation equation of asymmetric simple exclusion processes. Probability Theory and Related Fields, 108(3):321–356, July 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050112.

Lyons:2010:KI

[Lyo10] Terence J. Lyons. Kiyoshi Itô (1915–2008). Probability Theory and Related Fields, 148(1–2):1–4, September 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-010-0286-7.pdf.

Lalley:2010:SEL

[LZ10] Steven P. Lalley and Xinghua Zheng. Spatial epidemics and local times for critical branching random walks in dimensions 2 and 3. Probability Theory and Related Fields, 148(3–4):527– 566, November 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0239-1.

Lygkonis:2024:MEE

[LZ24] Dimitris Lygkonis and Nikos Zygouras. A multivariate extension of the Erdős–Taylor theorem. Probability Theory and Related Fields, 189(1–2):179–227, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-024-01267-3.

MacLaurin:2021:EAF

[Mac21] James MacLaurin. An emergent autonomous flow for meanfield spin glasses. Probability Theory and Related Fields, 180 (1–2):365–438, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01040-w.

Madras:1989:RWK

[Mad89] Neal Madras. Random walks with killing. Probability Theory and Related Fields, 80(4):581–600, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00318907.

Maisonneuve:1989:SSP

[Mai89] B. Maisonneuve. Systèmes de sortie F_{Dt} prévisibles. (French) [Predictable output systems F_{Dt}]. Probability Theory and Related Fields, 80(3):395–405, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01794431.

Maillard:2016:SFP

[Mai16] Pascal Maillard. Speed and fluctuations of N-particle branching Brownian motion with spatial selection. Probability Theory and Related Fields, 166(3–4):1061–1173, December 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-016-0701-9;   http://link.springer.com/
article/10.1007/s00440-016-0701-9.

Major:1988:TBD

[Maj88a] P. Major. On the tail behaviour of the distribution function of multiple stochastic integrals. Probability Theory and Related Fields, 78(3):419–435, July 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00334204.

Major:1988:SVO

[Maj88b] Péter Major. On the set visited once by a random walk. Probability Theory and Related Fields, 77(1):117–128, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01848134.

Major:1992:PLN

[Maj92] Péter Major. Poisson law for the number of lattice points in a random strip with finite area. Probability Theory and Related Fields, 92(4):423–464, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01274263.

Major:2006:ESN

[Maj06] Péter Major. An estimate on the supremum of a nice class of stochastic integrals and U-statistics. Probability Theory and Related Fields, 134(3):489–537, March 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-005-0440-9.

Malric:1995:TLZ

[Mal95] M. Malric. Transformation de Lévy et zéros du mouvement Brownien. (French) [Lévy transformation and zeros of Brownian motion]. Probability Theory and Related Fields, 101(2):227–236, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01375826.

Male:2012:NPL

[Mal12] Camille Male. The norm of polynomials in large random and deterministic matrices. Probability Theory and Related Fields, 154(3–4):477–532, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-011-0375-2.

Mammen:1987:OLG

[Mam87] Enno Mammen. Optical local Gaussian approximation of an exponential family. Probability Theory and Related Fields, 76(1): 103–119, September 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00390278.

Mammen:1992:BWB

[Mam92] Enno Mammen. Bootstrap, wild bootstrap, and asymptotic normality. Probability Theory and Related Fields, 93(4):439–455, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01192716.

Mandelbaum:1986:DMA

[Man86] Avi Mandelbaum. Discrete multi-armed bandits and multiparameter processes. Probability Theory and Related Fields, 71 (1):129–147, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00366276.

Mansmann:1991:SCL

[Man91] Ulrich Mansmann. Strong coupling limit for a certain class of polaron models. Probability Theory and Related Fields, 90 (4):427–446, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192137.

Marton:1987:ICC

[Mar87] K. Marton. Isomorphisms of certain correlated probability spaces and of discrete memoryless correlated sources. Probability Theory and Related Fields, 75(1):129–142, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00320086.

Marohn:1991:GSE

[Mar91] Frank Marohn. Global sufficiency of extreme order statistics in location models of Weibull type. Probability Theory and Related Fields, 88(2):261–268, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01212562.

Marton:1998:MCC

[Mar98] Katalin Marton. Measure concentration for a class of random processes. Probability Theory and Related Fields, 110(3): 427–439, March 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050154.

Marinucci:2008:CLT

[Mar08] Domenico Marinucci. A central limit theorem and higher order results for the angular bispectrum. Probability Theory and Related Fields, 141(3–4):389–409, July 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0088-8.

Mariani:2010:LDP

[Mar10] Mauro Mariani. Large deviations principles for stochastic scalar conservation laws. Probability Theory and Related Fields, 147 (3–4):607–648, July 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0218-6.

Massart:1988:APD

[Mas88] P. Massart. About the Prohorov distance between the uniform distribution over the unit cube in R^{d} and its empirical measure. Probability Theory and Related Fields, 79(3):431–450, October 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00342234.

Mase:1992:ULC

[Mas92] Shigeru Mase. Uniform LAN condition of planar Gibbsian point processes and optimality of maximum likelihood estimators of soft-core potential functions. Probability Theory and Related Fields, 92(1):51–67, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01205236.

Mattner:1992:CLF

[Mat92] L. Mattner. Completeness of location families, translated moments, and uniqueness of charges. Probability Theory and Related Fields, 92(2):137–149, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01194918.

Mathieu:1994:ZWN

[Mat94] Pierre Mathieu. Zero white noise limit through Dirichlet forms, with application to diffusions in a random medium. Probability Theory and Related Fields, 99(4):549–580, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01206232.

Mathieu:1995:IBY

[Mat95] P. Mathieu. Inegalités Barlow–Yor pour les temps locaux d’intersection de deux mouvements browniens plans. (French) [Barlow–Yor inequalities for the local intersection times of two planar Brownian motions]. Probability Theory and Related Fields, 102(1):83–103, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01295223.

Mattner:2004:CUH

[Mat04] Lutz Mattner. Cumulants are universal homomorphisms into Hausdorff groups. Probability Theory and Related Fields, 130 (2):151–166, October 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0354-y.

Mazza:1999:GSP

[Maz99] Christian Mazza. Gauge symmetries and percolation in ±J Ising spin glasses. Probability Theory and Related Fields, 113(2): 171–190, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s440-1999-8032-8.

McGill:1986:MCB

[McG86] Paul McGill. A martingale characterisation of the Brownian excursion compensator. Probability Theory and Related Fields, 71 (1):117–128, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00366275.

McGill:1989:WHF

[McG89] Paul McGill. Wiener–Hopf factorisation of Brownian motion. Probability Theory and Related Fields, 83(3):355–389, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00964370.

Meleard:2001:MCA

[Mél01] Sylvie Méléard. Monte–Carlo approximations for 2D Navier– Stokes equations with measure initial data. Probability Theory and Related Fields, 121(3):367–388, November 2001. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100154.

Mentemeier:2016:FPM

[Men16] Sebastian Mentemeier. The fixed points of the multivariate smoothing transform. Probability Theory and Related Fields, 164(1–2):401–458, February 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-015-0615-y.

Mendelson:2017:AHT

[Men17] Shahar Mendelson. On aggregation for heavy-tailed classes. Probability Theory and Related Fields, 168(3–4):641–674, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-016-0720-6.

Mendelson:2018:LCG

[Men18] Shahar Mendelson. Learning without concentration for general loss functions. Probability Theory and Related Fields, 171(1– 2):459–502, June 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0784-y.

Merkl:2003:QAG

[Mer03] Franz Merkl. Quenched asymptotics of the ground state energy of random Schrödinger operators with scaled Gibbsian potentials. Probability Theory and Related Fields, 126(3):307–338, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0266-2.

Metcalfe:2013:UPG

[Met13] Anthony P. Metcalfe. Universality properties of Gelfand–Tsetlin patterns. Probability Theory and Related Fields, 155(1–2):303– 346, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-011-0399-7.

Miclo:1999:REI

[Mic99] Laurent Miclo. Relations entre isopérimétrie et trou spectral pour les chaînes de Markov finies. (French) [Relations between isoperimetry and spectral gap for finite Markov chains]. Probability Theory and Related Fields, 114(4):431–485, July 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050231.

Miclo:2008:EMP

[Mic08] Laurent Miclo. On eigenfunctions of Markov processes on trees. Probability Theory and Related Fields, 142(3–4):561– 594, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0115-9.

Miclo:2022:M

[Mic22] Laurent Miclo. On metastability. Probability Theory and Related Fields, 184(1–2):275–322, October 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-022-01147-
8.

Miermont:2003:SSF

[Mie03] Grégory Miermont. Self-similar fragmentations derived from the stable tree I. Probability Theory and Related Fields, 127(3): 423–454, November 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0295-x.

Miermont:2005:SSF

[Mie05] Grégory Miermont. Self-similar fragmentations derived from the stable tree II: splitting at nodes. Probability Theory and Related Fields, 131(3):341–375, March 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0373-8.

Mijnheer:1988:FFL

[Mij88a] Joop Mijnheer. A functional form for the lower Lipschitz condition for the stable subordinator. Probability Theory and Related Fields, 77(4):515–520, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00959614.

Mijnheer:1988:RCS

[Mij88b] Joop Mijnheer. On the rate of convergence of the sum of the sample extremes. Probability Theory and Related Fields, 79(3): 317–325, October 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00342230.

Mikami:2004:MPQ

[Mik04] Toshio Mikami. Monge’s problem with a quadratic cost by the zero-noise limit of h-path processes. Probability Theory and Related Fields, 129(2):245–260, June 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0340-4.

Milasevic:1989:ESL

[Mil89] Philip Milasevic. Estimation of the slope in a linear functional relationship. Probability Theory and Related Fields, 82(1):19–31, June 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00340010.

Milman:2012:PIF

[Mil12] Emanuel Milman. Properties of isoperimetric, functional and Transport–Entropy inequalities via concentration. Probability Theory and Related Fields, 152(3–4):475–507, April 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0328-1.

Mirek:2011:HTP

[Mir11] Mariusz Mirek. Heavy tail phenomenon and convergence to stable laws for iterated Lipschitz maps. Probability Theory and Related Fields, 151(3–4):705–734, December 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0312-9.

Mirek:2013:FPG

[Mir13] Mariusz Mirek. On fixed points of a generalized multidimensional affine recursion. Probability Theory and Related Fields, 156(3–4):665–705, August 2013. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s00440-
012-0439-y.pdf.

Maejima:2009:LNS

[MiS09] Makoto Maejima and Ken iti Sato. The limits of nested subclasses of several classes of infinitely divisible distributions are identical with the closure of the class of stable distributions. Probability Theory and Related Fields, 145(1–2):119– 142, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0163-9.

Mitro:1986:DTC

[Mit86] Joanna B. Mitro. A discontinuous time change for natural additive functionals which preserves duality. Probability Theory and Related Fields, 72(3):463–470, June 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00334196.

Mitoma:1987:GOU

[Mit87] Itaru Mitoma. Generalized Ornstein–Uhlenbeck process having a characteristic operator with polynomial coefficients. Probability Theory and Related Fields, 76(4):533–555, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00960073.

Mlotkowski:1999:FPA

[Mło99] Wojciech Młotkowski. Free probability on algebras with infinitely many states. Probability Theory and Related Fields, 115(4):579–596, November 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050250.

Makagon:1990:SRS

[MM90a] A. Makagon and V. Mandrekar. The spectral representation of stable processes: Harmonizability and regularity. Probability Theory and Related Fields, 85(1):1–11, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01377623.

Mazziotto:1990:ACL

[MM90b] G. Mazziotto and A. Millet. Absolute continuity of the law of an infinite dimensional Wiener functional with respect to the Wiener probability. Probability Theory and Related Fields, 85 (3):403–411, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193945.

Meckes:2013:CCR

[MM13a] Elizabeth S. Meckes and Mark W. Meckes. Concentration and convergence rates for spectral measures of random matrices. Probability Theory and Related Fields, 156(1–2):145–164, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-012-0423-6.

Mikosch:2013:LDM

[MM13b] Thomas Mikosch and Martin Moser. The limit distribution of the maximum increment of a random walk with dependent regularly varying jump sizes. Probability Theory and Related Fields, 156(1–2):249–272, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0427-2.

Mathieu:2015:AAD

[MM15] Pierre Mathieu and Jean-Christophe Mourrat. Aging of asymmetric dynamics on the random energy model. Probability Theory and Related Fields, 161(1–2):351–427, February 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0551-2.

Meiners:2017:SCS

[MM17] Matthias Meiners and Sebastian Mentemeier. Solutions to complex smoothing equations. Probability Theory and Related Fields, 168(1–2):199–268, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Mammen:1992:SAM

[MMF92] E. Mammen, J. S. Marron, and N. I. Fisher. Some asymptotics for multimodality tests based on kernel density estimates. Probability Theory and Related Fields, 91(1):115–132, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01194493.

Montanari:2012:WLI

[MMS12] Andrea Montanari, Elchanan Mossel, and Allan Sly. The weak limit of Ising models on locally tree-like graphs. Probability Theory and Related Fields, 152(1–2):31–51, February 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0315-6.

Maida:2014:FTE

[MMS14] Mylène Maïda and Édouard Maurel-Segala. Free transportentropy inequalities for non-convex potentials and application to concentration for random matrices. Probability Theory and Related Fields, 159(1–2):329–356, June 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-013-0508-x.

Marciniak:2021:PLB

[MMŚ21] Mikołaj Marciniak, Lukasz Maślanka, and Piotr Śniady. Poisson limit of bumping routes in the Robinson–Schensted correspondence. Probability Theory and Related Fields, 181(4):1053– 1103, December 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01084-y.

Mischler:2015:NAQ

[MMW15] Stéphane Mischler, Clément Mouhot, and Bernt Wennberg. A new approach to quantitative propagation of chaos for drift, diffusion and jump processes. Probability Theory and Related Fields, 161(1–2):1–59, February 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-013-0542-8.

Mikosch:1987:SLL

[MN87] T. Mikosch and R. Norvaisa. Strong laws of large numbers for fields of Banach space valued random variables. Probability Theory and Related Fields, 74(2):241–253, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00569991.

Malliavin:1993:QSA

[MN93] Paul Malliavin and David Nualart. Quasi sure analysis and Stratonovich anticipative stochastic differential equations. Probability Theory and Related Fields, 96(1):45–55, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01195882.

Maejima:1998:SSD

[MN98a] Makoto Maejima and Yoshihiro Naito. Semi-selfdecomposable distributions and a new class of limit theorems. Probability Theory and Related Fields, 112(1):13–31, September 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050181.

Milstein:1998:DAN

[MN98b] Grigori Milstein and Michael Nussbaum. Diffusion approximation for nonparametric autoregression. Probability Theory and Related Fields, 112(4):535–543, December 1998. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050199.

Moret:2002:OMF

[MN02] Sílvia Moret and David Nualart. Onsager–Machlup functional for the fractional Brownian motion. Probability Theory and Related Fields, 124(2):227–260, October 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400200211.

Michoel:2004:CLT

[MN04] Tom Michoel and Bruno Nachtergaele. Central limit theorems for the large-spin asymptotics of quantum spins. Probability Theory and Related Fields, 130(4):493–517, December 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0364-9.

Millet:1989:TRI

[MNS89] Annie Millet, David Nualart, and Marta Sanz. Time reversal for infinite-dimensional diffusions. Probability Theory and Related Fields, 82(3):315–347, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00339991.

Mossel:2015:REP

[MNS15] Elchanan Mossel, Joe Neeman, and Allan Sly. Reconstruction and estimation in the planted partition model. Probability Theory and Related Fields, 162(3–4):431–461, August 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0576-6.

Michaeli:2021:DUS

[MNS21] Peleg Michaeli, Asaf Nachmias, and Matan Shalev. The diameter of uniform spanning trees in high dimensions. Probability Theory and Related Fields, 179(1–2):261–294, February 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00999-2.

Mori:1986:LIL

[MO86] Toshio Mori and Hiroshi Oodaira. The law of the iterated logarithm for self-similar processes represented by multiple Wiener integrals. Probability Theory and Related Fields, 71(3):367–391, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01000212.

Mori:1987:FIL

[MO87] Toshio Mori and Hiroshi Oodaira. The functional iterated logarithm law for stochastic processes represented by multiple Wiener integrals. Probability Theory and Related Fields, 76(3):299–310, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01297487.

Marahrens:2015:AEG

[MO15] Daniel Marahrens and Felix Otto. Annealed estimates on the Green function. Probability Theory and Related Fields, 163 (3–4):527–573, December 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0598-0.

Monmarche:2018:HMS

[Mon18] Pierre Monmarché. Hypocoercivity in metastable settings and kinetic simulated annealing. Probability Theory and Related Fields, 172(3–4):1215–1248, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-0828-y.

Monch:2024:ILR

[Mön24] Christian Mönch. Inhomogeneous long-range percolation in the weak decay regime. Probability Theory and Related Fields, 189(3–4):1129–1160, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-024-01281-5.

Morimoto:1986:NZS

[Mor86] Hiroaki Morimoto. Non-zero-sum discrete parameter stochastic games with stopping times. Probability Theory and Related Fields, 72(1):155–160, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00343901.

Morrow:1987:CLT

[Mor87] Gregory J. Morrow. Central limit theorem for linearly dependent fields of continuous elements. Probability Theory and Related Fields, 75(1):87–95, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00320083.

Mori:1991:WTT

[Mór91a] Tamás F. Móri. On the waiting time till each of some given patterns occurs as a run. Probability Theory and Related Fields, 87(3):313–323, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01312212.

Morimoto:1991:ACS

[Mor91b] Hiroaki Morimoto. On average cost stopping time problems. Probability Theory and Related Fields, 90(4):469–490, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01192139.

Mori:1992:RLA

[Mor92] Toshio Mori. Representation of linearly additive random fields. Probability Theory and Related Fields, 92(1):91–115, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01205238.

Morters:2000:ASK

[Mör00] Peter Mörters. Almost sure Kallianpur–Robbins laws for Brownian motion in the plane. Probability Theory and Related Fields, 118(1):49–64, September 2000. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008742;
 http://link.springer.com/article/10.1007/s004400000077.

Morters:2001:HFP

[Mör01] Peter Mörters. How fast are the particles of super-Brownian motion? Probability Theory and Related Fields, 121(2):171– 197, October 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008801.

Morris:2003:CWS

[Mor03] Ben Morris. The components of the wired spanning forest are recurrent. Probability Theory and Related Fields, 125(2): 259–265, February 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-002-0236-0.

Morris:2011:ZTG

[Mor11] Robert Morris. Zero-temperature Glauber dynamics on Z^{d}. Probability Theory and Related Fields, 149(3–4):417–434, April 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-009-0259-x.

Mossel:2012:QAT

[Mos12] Elchanan Mossel. A quantitative Arrow theorem. Probability Theory and Related Fields, 154(1–2):49–88, October 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0362-7.

Mountford:1990:LBO

[Mou90] T. S. Mountford. Limiting behaviour of the occupation of wedges by complex Brownian motion. Probability Theory and Related Fields, 84(1):55–65, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01288558.

Mountford:1992:CVS

[Mou92a] T. S. Mountford. The critical value for some non-attractive long range nearest particle systems. Probability Theory and Related Fields, 93(1):67–76, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01195388.

Mountford:1992:ECR

[Mou92b] T. S. Mountford. The ergodicity of a class of reversible reactiondiffusion processes. Probability Theory and Related Fields, 92 (2):259–274, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01194924.

Mountford:1992:RPL

[Mou92c] T. S. Mountford. Rates for the probability of large cubes being non-internally spanned in modified bootstrap percolation. Probability Theory and Related Fields, 93(2):159–167, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01195227.

Mourrat:2014:KDM

[Mou14] Jean-Christophe Mourrat. Kantorovich distance in the martingale CLT and quantitative homogenization of parabolic equations with random coefficients. Probability Theory and Related Fields, 160(1–2):279–314, October 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-013-0529-5.

Morrow:1986:IPP

[MP86] Gregory J. Morrow and Walter Philipp. Invariance principles for partial sum processes and empirical processes indexed by sets. Probability Theory and Related Fields, 73(1):11–42, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01845991.

Metivier:1987:TCP

[MP87] M. Métivier and P. Priouret. Théorèmes de convergence presque sure pour une classe d’algorithmes stochastiques à pas décroissant. (French) [Almost sure convergence theorems for a class of stochastic algorithms with decreasing steps]. Probability Theory and Related Fields, 74(3):403–428, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00699098.

Miamee:1988:WDP

[MP88] A. G. Miamee and M. Pourahmadi. Wold decomposition, prediction and parameterization of stationary processes with infinite variance. Probability Theory and Related Fields, 79(1): 145–164, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319110.

Monrad:1991:NVN

[MP91] Ditlev Monrad and Walter Philipp. Nearby variables with nearby conditional laws and a strong approximation theorem for Hilbert space valued martingales. Probability Theory and Related Fields, 88(3):381–404, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01418867.

Mueller:1992:CSP

[MP92] Carl Mueller and Edwin A. Perkins. The compact support property for solutions to the heat equation with noise. Probability Theory and Related Fields, 93(3):325–358, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01193055.

Mazza:2002:PCS

[MP02] Christian Mazza and Didier Piau. Products of correlated symmetric matrices and q-Catalan numbers. Probability Theory and Related Fields, 124(4):574–594, December 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
002-0225-3.

Morris:2005:ESM

[MP05] B. Morris and Yuval Peres. Evolving sets, mixing and heat kernel bounds. Probability Theory and Related Fields, 133(2): 245–266, October 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-005-0434-7.

Mytnik:2011:PUS

[MP11] Leonid Mytnik and Edwin Perkins. Pathwise uniqueness for stochastic heat equations with Hölder continuous coefficients: the white noise case. Probability Theory and Related Fields, 149 (1–2):1–96, February 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0241-7.

Mytnik:2019:DBS

[MP19] Leonid Mytnik and Edwin Perkins. The dimension of the boundary of super-Brownian motion. Probability Theory and Related Fields, 174(3–4):821–885, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0866-5.

Merlevede:2011:BTI

[MPR11] Florence Merlevède, Magda Peligrad, and Emmanuel Rio. A Bernstein type inequality and moderate deviations for weakly dependent sequences. Probability Theory and Related Fields, 151(3–4):435–474, December 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0304-9.

Muirhead:2018:DPA

[MPS18] Stephen Muirhead, Richard Pymar, and Nadia Sidorova. Delocalising the parabolic Anderson model through partial duplication of the potential. Probability Theory and Related Fields, 171(3–4):917–979, August 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0798-5;   http://link.springer.com/content/pdf/
10.1007/s00440-017-0798-5.pdf.

Matoussi:2019:PIS

[MPS19] Anis Matoussi, Dylan Possamaï, and Wissal Sabbagh. Probabilistic interpretation for solutions of fully nonlinear stochastic PDEs. Probability Theory and Related Fields, 174(1–2):177–233, June 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-018-0859-4.

Menshikov:2001:CPC

[MPV01] M. V. Menshikov, S. Yu. Popov, and M. Vachkovskaia. On the connectivity properties of the complementary set in fractal percolation models. Probability Theory and Related Fields, 119(2): 176–186, February 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008757.

Ma:1994:SFB

[MPY94] Jin Ma, Philip Protter, and Jiongmin Yong. Solving forwardbackward stochastic differential equations explicitly — a four step scheme. Probability Theory and Related Fields, 98(3): 339–359, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192258.

Miller:2020:GLQ

[MQ20] Jason Miller and Wei Qian. The geodesics in Liouville quantum gravity are not Schramm-Loewner evolutions. Probability Theory and Related Fields, 177(3–4):677–709, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-019-00949-7.

Monrad:1995:SVG

[MR95] Ditlev Monrad and Holger Rootzén. Small values of Gaussian processes and functional laws of the iterated logarithm. Probability Theory and Related Fields, 101(2):173–192, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01375823.

Matzinger:2003:RRS

[MR03] Heinrich Matzinger and Silke W. W. Rolles. Reconstructing a random scenery observed with random errors along a random walk path. Probability Theory and Related Fields, 125 (4):539–577, April 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0257-3.

Mattner:2007:SPK

[MR07a] Lutz Mattner and Bero Roos. A shorter proof of Kanter’s Bessel function concentration bound. Probability Theory and Related Fields, 139(1–2):191–205, September 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-006-0043-0.

Merkl:2007:REL

[MR07b] Franz Merkl and Silke W. W. Rolles. A random environment for linearly edge-reinforced random walks on infinite graphs. Probability Theory and Related Fields, 138(1–2):157–176, May 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-006-0016-3.

Merkl:2009:ERR

[MR09] Franz Merkl and Silke W. W. Rolles. Edge-reinforced random walk on one-dimensional periodic graphs. Probability Theory and Related Fields, 145(3–4):323–349, November 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0170-x.

Mikulevicius:2012:USN

[MR12] R. Mikulevicius and B. L. Rozovskii. On unbiased stochastic Navier–Stokes equations. Probability Theory and Related Fields, 154(3–4):787–834, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-011-0384-1.

Meister:2013:AEN

[MR13] Alexander Meister and Markus Reiß. Asymptotic equivalence for nonparametric regression with non-regular errors. Probability Theory and Related Fields, 155(1–2):201–229, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-011-0396-x.

Morters:2016:SET

[MR16] Peter Mörters and István Redl. Skorokhod embeddings for twosided Markov chains. Probability Theory and Related Fields, 165 (1–2):483–508, June 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0637-5.

Matetski:2023:TGM

[MR23] Konstantin Matetski and Daniel Remenik. TASEP and generalizations: method for exact solution. Probability Theory and Related Fields, 185(1–2):615–698, February 2023. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-022-01129-w.

Merkl:2019:CVR

[MRT19] Franz Merkl, Silke W. W. Rolles, and Pierre Tarrès. Convergence of vertex-reinforced jump processes to an extension of the supersymmetric hyperbolic nonlinear sigma model. Probability Theory and Related Fields, 173(3–4):1349–1387, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/

s00440-018-0855-8. See correction [MRT20].

Merkl:2020:CCV

[MRT20] Franz Merkl, Silke W. W. Rolles, and Pierre Tarrès. Correction to: Convergence of vertex-reinforced jump processes to an extension of the supersymmetric hyperbolic nonlinear sigma model. Probability Theory and Related Fields, 176(3–4):1445, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-020-00958-x;   https://link.springer.
com/content/pdf/10.1007/s00440-020-00958-x.pdf. See

[MRT19].

Mao:2023:EMR

[MRT23] Cheng Mao, Mark Rudelson, and Konstantin Tikhomirov. Exact matching of random graphs with constant correlation. Probability Theory and Related Fields, 186(1–2):327–389, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01184-3.

March:1987:SCB

[MS87] Peter March and Alain-Sol Sznitman. Some connections between excursion theory and the discrete Schrödinger equation with random potentials. Probability Theory and Related Fields, 75(1):11–53, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00320079.

Mukherjea:1990:IPD

[MS90] Arunava Mukherjea and Richard Stephens. Identification of parameters by the distribution of the maximum random variable: The general multivariate normal case. Probability Theory and Related Fields, 84(3):289–296, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01197886.

Millet:1992:COU

[MS92] A. Millet and W. Smoleński. On the continuity of Ornstein– Uhlenbeck processes in infinite dimensions. Probability Theory and Related Fields, 92(4):529–547, ???? 1992. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01274267.

Mueller:1993:BHE

[MS93] Carl Mueller and Richard Sowers. Blowup for the heat equation with a noise term. Probability Theory and Related Fields, 97 (3):287–320, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01195068.

Miamee:1995:CSD

[MS95] Abol G. Miamee and Bernd S. W. Schröder. On completeness of the spectral domain of harmonizable processes. Probability Theory and Related Fields, 101(3):303–309, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01200498.

Maes:1999:FTI

[MS99] Christian Maes and Senya Shlosman. Freezing transition in the Ising model without internal contours. Probability Theory and Related Fields, 115(4):479–503, November 1999. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050246.

Maslowski:2000:PAS

[MS00] Bohdan Maslowski and Jan Seidler. Probabilistic approach to the strong Feller property. Probability Theory and Related Fields, 118(2):187–210, October 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s440-000-8014-0.

Moynot:2002:LDM

[MS02] Olivier Moynot and Manuel Samuelides. Large deviations and mean-field theory for asymmetric random recurrent neural networks. Probability Theory and Related Fields, 123(1):41–75, May 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400100182.

Morters:2009:EPM

[MS09] Peter Mörters and Narn-Rueih Shieh. The exact packing measure of Brownian double points. Probability Theory and Related Fields, 143(1–2):113–136, January 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0122-x.

Mossel:2010:GRS

[MS10] Elchanan Mossel and Allan Sly. Gibbs rapidly samples colorings of G(n, d/n). Probability Theory and Related Fields, 148(1–2): 37–69, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00440-009-0222-x.pdf.

Miller:2016:IGI

[MS16] Jason Miller and Scott Sheffield. Imaginary geometry I: interacting SLEs. Probability Theory and Related Fields, 164(3–4):553–705, April 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-016-0698-0;
 http://link.springer.com/content/pdf/10.1007/s00440-
016-0698-0.pdf.

Miller:2017:IGI

[MS17a] Jason Miller and Scott Sheffield. Imaginary geometry IV: interior rays, whole-plane reversibility, and space-filling trees. Probability Theory and Related Fields, 169(3–4):729–869, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s00440-017-0780-2.pdf;   http://
link.springer.com/article/10.1007/s00440-017-0780-2.

Miller:2017:ULP

[MS17b] Jason Miller and Perla Sousi. Uniformity of the late points of random walk on Z_{n}^{d} for d ≥ 3. Probability Theory and Related Fields, 167(3–4):1001–1056, April 2017. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Miller:2021:LQG

[MS21] Jason Miller and Scott Sheffield. Liouville quantum gravity and the Brownian map III: the conformal structure is determined. Probability Theory and Related Fields, 179(3–4):1183– 1211, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01026-8.

Mikulincer:2024:BTM

[MS24a] Dan Mikulincer and Yair Shenfeld. The Brownian transport map. Probability Theory and Related Fields, 190(1–2):379–444, October 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-024-01286-0.

Mu:2024:QUV

[MS24b] Yingxin Mu and Artem Sapozhnikov. On questions of uniqueness for the vacant set of Wiener sausages and Brownian interlacements. Probability Theory and Related Fields, 190(3-4): 703–751, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-024-01315-y.

Millet:1994:SSH

[MSS94] Annie Millet and Marta Sanz-Solé. The support of the solution to a hyperbolic SPDE. Probability Theory and Related Fields, 98(3):361–387, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192259.

Marcus:2022:FFC

[MSS22] Adam W. Marcus, Daniel A. Spielman, and Nikhil Srivastava. Finite free convolutions of polynomials. Probability Theory and Related Fields, 182(3–4):807–848, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01105-w.

Mossel:2014:ALB

[MST14] Elchanan Mossel, Allan Sly, and Omer Tamuz. Asymptotic learning on Bayesian social networks. Probability Theory and Related Fields, 158(1–2):127–157, February 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0479-y.

Muller:2009:EID

[MSW09] Ursula U. Müller, Anton Schick, and Wolfgang Wefelmeyer. Estimating the innovation distribution in nonparametric autoregression. Probability Theory and Related Fields, 144(1–2):53–77, May 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-008-0141-2.

Miller:2021:NSC

[MSW21] Jason Miller, Scott Sheffield, and Wendelin Werner. Nonsimple conformal loop ensembles on Liouville quantum gravity and the law of CLE percolation interfaces. Probability Theory and Related Fields, 181(1-3):669–710, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01070-4.

McConnell:1987:DBV

[MT87] Terry R. McConnell and Murad S. Taqqu. Decoupling of Banach-valued multilinear forms in independent symmetric Banach-valued random variables. Probability Theory and Related Fields, 75(4):499–507, August 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00320330.

Mueller:1994:PTS

[MT94] Carl Mueller and Roger Tribe. A phase transition for a stochastic PDE related to the contact process. Probability Theory and Related Fields, 100(2):131–156, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01199262.

Muller:1995:SPD

[MT95] C. Müller and R. Tribe. Stochastic p.d.e.’s arising from the long range contact and long range voter processes. Probability Theory and Related Fields, 102(4):519–545, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01198848.

Mueller:2011:PDS

[MT11] Carl Mueller and Roger Tribe. A phase diagram for a stochastic reaction diffusion system. Probability Theory and Related Fields, 149(3–4):561–637, April 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0265-z.

Menz:2020:VPN

[MT20] Georg Menz and Martin Tassy. A variational principle for a nonintegrable model. Probability Theory and Related Fields, 177(3– 4):747–822, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-020-00959-w.

McKeague:1990:INC

[MU90] Ian W. McKeague and Klaus J. Utikal. Identifying nonlinear covariate effects in semimartingale regression models. Probability Theory and Related Fields, 87(1):1–25, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01217745.

Mijatovic:2012:MPC

[MU12] Aleksandar Mijatović and Mikhail Urusov. On the martingale property of certain local martingales. Probability Theory and Related Fields, 152(1–2):1–30, February 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0314-7.

Muck:1994:LDW

[Müc94] Stefan Mück. Large deviations w.r.t. quasi-every starting point for symmetric right processes on general state spaces. Probability Theory and Related Fields, 99(4):527–548, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01206231.

Mueller:1991:CBS

[Mue91a] Carl Mueller. A connection between Strassen’s and Donsker– Varadhan’s laws of the iterated logarithm. Probability Theory and Related Fields, 87(3):365–388, ???? 1991. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01312216.

Mueller:1991:LTE

[Mue91b] Carl Mueller. Long time existence for the heat equation with a noise term. Probability Theory and Related Fields, 90(4): 505–517, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192141.

Mueller:1998:LTE

[Mue98] Carl Mueller. Long-time existence for signed solutions of the heat equation with a noise term. Probability Theory and Related Fields, 110(1):51–68, January 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050144.

Mukherjea:1991:TPD

[Muk91] Arunava Mukherjea. Tightness of products of i.i.d. random matrices. Probability Theory and Related Fields, 87(3):389–401, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01312217.

Mukhrjea:1993:RRW

[Muk93] Arunava Mukhrjea. Recurrent random walks in nonnegative matrices II. Probability Theory and Related Fields, 96(4):415–434, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01200204.

Muller:1989:MWR

[Mül89] Josef S. Müller. On the mean width of random polytopes. Probability Theory and Related Fields, 82(1):33–37, June 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00340011.

Murugan:2024:HKR

[Mur24] Mathav Murugan. Heat kernel for reflected diffusion and extension property on uniform domains. Probability Theory and Related Fields, 190(1–2):543–599, October 2024. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-024-01266-4.

Macht:1989:LLP

[MW89] Wolfgang Macht and Werner Wolf. The local limit problem and Hölder-continuity. Probability Theory and Related Fields, 82 (2):295–305, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00354765.

Muller:1990:LAH

[MW90] Hans-Georg Müller and Jane-Ling Wang. Locally adaptive hazard smoothing. Probability Theory and Related Fields, 85(4): 523–538, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01203169.

Meyre:1995:OTC

[MW95] Thierry Meyre and Wendelin Werner. On the occupation times of cones by Brownian motion. Probability Theory and Related Fields, 101(3):409–419, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01200504.

Merkl:2001:PTP

[MW01] Franz Merkl and Mario V. Wüthrich. Phase transition of the principal Dirichlet eigenvalue in a scaled Poissonian potential. Probability Theory and Related Fields, 119(4):475–507, April 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008768.

Morvai:2005:PDT

[MW05] Gusztáv Morvai and Benjamin Weiss. Prediction for discrete time series. Probability Theory and Related Fields, 132(1):1–12, May 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-004-0386-3.

Mikosch:2013:PLD

[MW13] Thomas Mikosch and Olivier Wintenberger. Precise large deviations for dependent regularly varying sequences. Probability Theory and Related Fields, 156(3–4):851–887, August 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0445-0.

Mikosch:2014:CIR

[MW14] Thomas Mikosch and Olivier Wintenberger. The cluster index of regularly varying sequences with applications to limit theory for functions of multivariate Markov chains. Probability Theory and Related Fields, 159(1–2):157–196, June 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0504-1.

Mikosch:2016:LDA

[MW16] Thomas Mikosch and Olivier Wintenberger. A large deviations approach to limit theory for heavy-tailed time series. Probability Theory and Related Fields, 166(1–2):233–269, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-015-0654-4.

Miller:2017:ISP

[MW17] Jason Miller and Hao Wu. Intersections of SLE paths: the double and cut point dimension of SLE. Probability Theory and Related Fields, 167(1–2):45–105, February 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
015-0677-x;   http://link.springer.com/content/pdf/
10.1007/s00440-015-0677-x.pdf.

Mossel:2009:HSI

[MWW09] Elchanan Mossel, Dror Weitz, and Nicholas Wormald. On the hardness of sampling independent sets beyond the tree threshold. Probability Theory and Related Fields, 143(3–4):401–439, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-007-0131-9.

Miller:2015:CLE

[MWW15] Jason Miller, Samuel S. Watson, and David B. Wilson. The conformal loop ensemble nesting field. Probability Theory and Related Fields, 163(3–4):769–801, December 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0604-6.

Mayer-Wolf:2005:DBS

[MWZ05] E. Mayer-Wolf and M. Zakai. The divergence of Banach space valued random variables on Wiener space. Probability Theory and Related Fields, 132(2):291–320, June 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0397-0.

Mayer-Wolf:2008:DBS

[MWZ08] E. Mayer-Wolf and M. Zakai. The divergence of Banach space valued random variables on Wiener space. Probability Theory and Related Fields, 140(3–4):631–633, March 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0111-0.

Ma:1999:LDB

[MY99] Jin Ma and Jiongmin Yong. On linear, degenerate backward stochastic partial differential equations. Probability Theory and Related Fields, 113(2):135–170, February 1999. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050205.

Mykland:1995:EAE

[Myk95] P. A. Mykland. Embedding and asymptotic expansions for martingales. Probability Theory and Related Fields, 103(4):475–492, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01246335.

Mytnik:1998:UMC

[Myt98] Leonid Mytnik. Uniqueness for a mutually catalytic branching model. Probability Theory and Related Fields, 112(2): 245–253, October 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050189.

Mytnik:2002:SPD

[Myt02] Leonid Mytnik. Stochastic partial differential equation driven by stable noise. Probability Theory and Related Fields, 123 (2):157–201, June 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400100180.

Merzbach:1987:STP

[MZ87] Ely Merzbach and Moshe Zakai. Stopping a two parameter weak martingale. Probability Theory and Related Fields, 76 (4):499–507, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00960070.

Ma:2002:PRS

[MZ02] Jin Ma and J. Zhang. Path regularity for solutions of backward stochastic differential equations. Probability Theory and Related Fields, 122(2):163–190, February 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400100144.

Ma:2011:WSF

[MZ11] Jin Ma and Jianfeng Zhang. On weak solutions of forwardbackward SDEs. Probability Theory and Related Fields, 151 (3–4):475–507, December 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0305-8.

Montanari:2024:TOE

[MZZ24] Andrea Montanari, Yiqiao Zhong, and Kangjie Zhou. Tractability from overparametrization: the example of the negative perceptron. Probability Theory and Related Fields, 188(3–4):805– 910, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-023-01248-y.

Nacu:2003:GDC

[Nac03] Serban Nacu. Glauber dynamics on the cycle is monotone. Probability Theory and Related Fields, 127(2):177–185, October 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-003-0279-x.

Nagai:1987:NZS

[Nag87] Hideo Nagai. Non zero-sum stopping games of symmetric Markov processes. Probability Theory and Related Fields, 75 (4):487–497, August 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00320329.

Nagasawa:1989:TDS

[Nag89] Masao Nagasawa. Transformations of diffusion and Schrödinger processes. Probability Theory and Related Fields, 82(1):109–136, June 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00340014.

Najnudel:2018:EVR

[Naj18] Joseph Najnudel. On the extreme values of the Riemann zeta function on random intervals of the critical line. Probability Theory and Related Fields, 172(1–2):387–452, October 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0812-y.

Nandori:2011:NDS

[Nán11] Péter Nándori. Number of distinct sites visited by a random walk with internal states. Probability Theory and Related Fields, 150(3–4):373–403, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0277-8.

Ndumu:1991:HKF

[Ndu91] Martin Ngu Ndumu. The heat kernel formula in a geodesic chart and some applications to the eigenvalue problem of the 3-sphere. Probability Theory and Related Fields, 88(3):343–361, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01418865.

Nestoridi:2019:OSS

[Nes19] Evita Nestoridi. Optimal strong stationary times for random walks on the chambers of a hyperplane arrangement. Probability Theory and Related Fields, 174(3–4):929–943, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-018-0872-7.

Neuhauser:1990:ETS

[Neu90] Claudia Neuhauser. An ergodic theorem for Schlögl models with small migration. Probability Theory and Related Fields, 85(1):27–32, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01377625.

Neufeld:2024:WSB

[Neu24] Leonie Neufeld. Weighted sums and Berry–Esseen type estimates in free probability theory. Probability Theory and Related Fields, 190(3-4):803–879, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-024-01294-
0.

Nguyen:1988:FTP

[Ngu88] Bao Gia Nguyen. Fourier transform of the percolation free energy. Probability Theory and Related Fields, 78(2):165–168, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00322016.

Nickl:2007:DTT

[Nic07] Richard Nickl. Donsker-type theorems for nonparametric maximum likelihood estimators. Probability Theory and Related Fields, 138(3–4):411–449, July 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-006-0031-4.

Nickl:2008:DTT

[Nic08] Richard Nickl. Donsker type theorems for nonparametric maximum likelihood estimators. Probability Theory and Related Fields, 141(1–2):331–332, May 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s00440-
007-0136-4.

Nicolaescu:2016:SGB

[Nic16] Liviu I. Nicolaescu. A stochastic Gauss–Bonnet–Chern formula. Probability Theory and Related Fields, 165(1–2):235–265, June 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-015-0630-z.

Nieuwenhuis:1989:EFL

[Nie89] Gert Nieuwenhuis. Equivalence of functional limit theorems for stationary point processes and their Palm distributions. Probability Theory and Related Fields, 81(4):593–608, May 1989.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00367306.

Nishiyama:1997:SCL

[Nis97] Yoichi Nishiyama. Some central limit theorems for l^{∞}- valued semimartingales and their applications. Probability Theory and Related Fields, 108(4):459–494, August 1997. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050117.

Nishikawa:2003:HLG

[Nis03] Takao Nishikawa. Hydrodynamic limit for the Ginzburg– Landau Δϕ interface model with boundary conditions. Probability Theory and Related Fields, 127(2):205–227, October 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-003-0283-1.

Nolan:1989:UCA

[NM89] Deborah Nolan and J. Stephen Marron. Uniform consistency of automatic and location-adaptive delta-sequence estimators. Probability Theory and Related Fields, 80(4):619–632, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00318909.

Ney:1993:RCI

[NN93] P. Ney and E. Nummelin. Regeneration for chains with infinite memory. Probability Theory and Related Fields, 96(4):503–520, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01200207.

Nazarov:2004:ESB

[NN04] A. I. Nazarov and Ya. Yu. Nikitin. Exact L_{2}-small ball behavior of integrated Gaussian processes and spectral asymptotics of boundary value problems. Probability Theory and Related Fields, 129(4):469–494, August 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0337-z.

Nourdin:2020:FBM

[NN20] Ivan Nourdin and David Nualart. The functional Breuer–Major theorem. Probability Theory and Related Fields, 176(1–2):203– 218, February 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-019-00917-1.

Nguyen:2015:CRM

[NO15] Hoi H. Nguyen and Sean O’Rourke. On the concentration of random multilinear forms and the universality of random block matrices. Probability Theory and Related Fields, 162(1–2):97–154, June 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-014-0567-7.

Nolan:1989:LNL

[Nol89] John P. Nolan. Local nondeterminism and local times for stable processes. Probability Theory and Related Fields, 82(3):387–410, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00339994.

Nolen:2014:NAR

[Nol14] James Nolen. Normal approximation for a random elliptic equation. Probability Theory and Related Fields, 159(3–4):661–700, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-013-0517-9.

Norberg:1986:RCT

[Nor86] Tommy Norberg. Random capacities and their distributions. Probability Theory and Related Fields, 73(2):281–297, September 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00339941.

Norris:1993:PIF

[Nor93] J. R. Norris. Path integral formulae for heat kernels and their derivatives. Probability Theory and Related Fields, 94 (4):525–541, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192562.

Nestoridi:2022:LPR

[NOT22] Evita Nestoridi and Sam Olesker-Taylor. Limit profiles for reversible Markov chains. Probability Theory and Related Fields, 182(1–2):157–188, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-021-01061-5.

Nualart:1988:SCA

[NP88] D. Nualart and E. Pardoux. Stochastic calculus with anticipating integrands. Probability Theory and Related Fields, 78 (4):535–581, August 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00353876.

Nualart:1992:WND

[NP92] D. Nualart and E. Pardoux. White noise driven quasilinear SPDEs with reflection. Probability Theory and Related Fields, 93(1):77–89, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01195389.

Nourdin:2009:SMW

[NP09] Ivan Nourdin and Giovanni Peccati. Stein’s method on Wiener chaos. Probability Theory and Related Fields, 145(1–2):75– 118, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0162-x.

Nechita:2012:RRQ

[NP12] Ion Nechita and Clément Pellegrini. Random repeated quantum interactions and random invariant states. Probability Theory and Related Fields, 152(1–2):299–320, February 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0323-6.

Nachmias:2022:LLU

[NP22] Asaf Nachmias and Yuval Peres. The local limit of uniform spanning trees. Probability Theory and Related Fields, 182(3– 4):1133–1161, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01072-2.

Neeman:2023:TLG

[NRS23] Joe Neeman, Charles Radin, and Lorenzo Sadun. Typical large graphs with given edge and triangle densities. Probability Theory and Related Fields, 186(3–4):1167–1223, August 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01187-8.

Nickl:2016:HFD

[NRST16] Richard Nickl, Markus Reiß, Jakob Söhl, and Mathias Trabs. High-frequency Donsker theorems for Lévy measures. Probability Theory and Related Fields, 164(1–2):61–108, February 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-014-0607-3.

Norris:1987:SAR

[NRW87] J. R. Norris, L. C. G. Rogers, and David Williams. Selfavoiding random walk: A Brownian motion model with local time drift. Probability Theory and Related Fields, 74(2):271–287, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00569993.

Nadtochiy:2023:SPS

[NS23a] Sergey Nadtochiy and Mykhaylo Shkolnikov. Stefan problem with surface tension: global existence of physical solutions under radial symmetry. Probability Theory and Related Fields, 187(1– 2):385–422, October 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-023-01206-8.

Nestoridi:2023:RWU

[NS23b] Evita Nestoridi and Allan Sly. The random walk on upper triangular matrices over Z/mZ. Probability Theory and Related Fields, 187(3–4):571–601, December 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-023-01228-
2.

Norris:2023:SLP

[NST23] James Norris, Vittoria Silvestri, and Amanda Turner. Scaling limits for planar aggregation with subcritical fluctuations. Probability Theory and Related Fields, 185(1–2):185–250, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01141-0.

Narayanan:2024:SGM

[NST24] Hariharan Narayanan, Scott Sheffield, and Terence Tao. Sums of GUE matrices and concentration of hives from correlation decay of eigengaps. Probability Theory and Related Fields, 190 (3-4):1121–1165, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01250-4.

Naumov:2019:BCS

[NSU19] Alexey Naumov, Vladimir Spokoiny, and Vladimir Ulyanov. Bootstrap confidence sets for spectral projectors of sample covariance. Probability Theory and Related Fields, 174(3–4): 1091–1132, August 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0877-2.

Nagasawa:1986:PCD

[NT86] Massao Nagasawa and Hiroshi Tanaka. Propagation of chaos for diffusing particles of two types with singular mean field interaction. Probability Theory and Related Fields, 71(1):69–83, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00366273.

Nagasawa:1987:DIC

[NT87] Masao Nagasawa and Hiroshi Tanaka. Diffusion with interactions and collisions between coloured particles and the propagation of chaos. Probability Theory and Related Fields, 74 (2):161–198, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00569988.

Nguyen:2017:RMT

[NTV17] Hoi Nguyen, Terence Tao, and Van Vu. Random matrices: tail bounds for gaps between eigenvalues. Probability Theory and Related Fields, 167(3–4):777–816, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-016-0693-5.

Nualart:1991:GAC

[NU91] D. Nualart and A. S. Ustunel. Geometric analysis of conditional independence on Wiener space. Probability Theory and Related Fields, 89(4):407–422, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199786.

Nummelin:1990:LDF

[Num90] Esa Nummelin. Large deviations for functionals of stationary processes. Probability Theory and Related Fields, 86(3):387–401, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01208257.

Nutz:2012:RAA

[Nut12] Marcel Nutz. Risk aversion asymptotics for power utility maximization. Probability Theory and Related Fields, 152(3–4): 703–749, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0334-3.

Nualart:1990:SRA

[NUZ90] D. Nualart, A. S. Ustunel, and M. Zakai. Some relations among classes of σ-fields on Wiener space. Probability Theory and Related Fields, 85(1):119–129, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01377633.

Najnudel:2021:BPB

[NV21] Joseph Najnudel and Bálint Virág. The bead process for beta ensembles. Probability Theory and Related Fields, 179(3–4):589– 647, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01034-8.

Newman:1990:MFB

[NW90] Charles M. Newman and C. Chris Wu. Markov fields on branching planes. Probability Theory and Related Fields, 85 (4):539–552, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01203170.

Nualart:1991:IPP

[NW91] David Nualart and Mario Wschebor. Intégration par parties dans l’espace de Wiener et approximation du temps local. (French) [Integration by parts in space and Wiener approximation of the local time]. Probability Theory and Related Fields, 90(1):83–109, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01321135.

Nutz:2022:EOT

[NW22] Marcel Nutz and Johannes Wiesel. Entropic optimal transport: convergence of potentials. Probability Theory and Related Fields, 184(1–2):401–424, October 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-021-01096-
8.

Nualart:1986:GSI

[NZ86] David Nualart and Moshe Zakai. Generalized stochastic integrals and the Malliavin calculus. Probability Theory and Related Fields, 73(2):255–280, September 1986. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00339940.

Nualart:1990:MWI

[NZ90] David Nualart and Moshe Zakai. Multiple Wiener–Itô integrals possessing a continuous extension. Probability Theory and Related Fields, 85(1):131–145, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01377634.

Neuman:2017:MDS

[NZ17] Eyal Neuman and Xinghua Zheng. On the maximal displacement of subcritical branching random walks. Probability Theory and Related Fields, 167(3–4):1137–1164, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-016-0702-8.

Neuman:2021:MDN

[NZ21] Eyal Neuman and Xinghua Zheng. On the maximal displacement of near-critical branching random walks. Probability Theory and Related Fields, 180(1–2):199–232, June 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01042-8.

Orsingher:2004:TFT

[OB04] Enzo Orsingher and Luisa Beghin. Time-fractional telegraph equations and telegraph processes with Brownian time. Probability Theory and Related Fields, 128(1):141–160, January 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-003-0309-8.

OConnell:1998:SLD

[O’C98] Neil O’Connell. Some large deviation results for sparse random graphs. Probability Theory and Related Fields, 110(3):277–285, March 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050149.

OConnell:2021:IDP

[O’C21] Neil O’Connell. Interacting diffusions on positive definite matrices. Probability Theory and Related Fields, 180(3–4):679–726, August 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01039-3.

Odasso:2008:EMS

[Oda08] Cyril Odasso. Exponential mixing for stochastic PDEs: the non-additive case. Probability Theory and Related Fields, 140 (1–2):41–82, January 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0057-2.

Oelschlager:1987:FTM

[Oel87] Karl Oelschläger. A fluctuation theorem for moderately interacting diffusion processes. Probability Theory and Related Fields, 74(4):591–616, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00363518.

Oelschlager:1989:DRD

[Oel89] Karl Oelschläger. On the derivation of reaction-diffusion equations as limit dynamics of systems of moderately interacting stochastic processes. Probability Theory and Related Fields, 82 (4):565–586, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00341284.

Offret:2014:IDS

[Off14] Yoann Offret. Invariant distributions and scaling limits for some diffusions in time-varying random environments. Probability Theory and Related Fields, 158(1–2):1–38, February 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0475-7.

Ogura:2001:WCL

[Ogu01] Yukio Ogura. Weak convergence of laws of stochastic processes on Riemannian manifolds. Probability Theory and Related Fields, 119(4):529–557, April 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/PL00008770.

Ohlsson:1989:ANT

[Ohl89] Esbjörn Ohlsson. Asymptotic normality for two-stage sampling from a finite population. Probability Theory and Related Fields, 81(3):341–352, April 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00340058.

Oliveira:2016:LTR

[Oli16] Roberto Imbuzeiro Oliveira. The lower tail of random quadratic forms with applications to ordinary least squares. Probability Theory and Related Fields, 166(3–4):1175–1194, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-016-0738-9;
http://link.springer.com/article/10.1007/s00440-016-
0738-9.

Olla:1987:LDA

[Oll87] Stefano Olla. Large deviations for almost Markovian processes. Probability Theory and Related Fields, 76(3):395–409, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01297493.

Olla:1988:LDG

[Oll88] Stefano Olla. Large deviations for Gibbs random fields. Probability Theory and Related Fields, 77(3):343–357, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00319293.

Ocone:1989:LSD

[OP89] Daniel Ocone and Etienne Pardoux. Linear stochastic differential equations with boundary conditions. Probability Theory and Related Fields, 82(4):489–526, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00341281.

Ortega:1989:UCI

[Ort89] J. Ortega. Upper classes for the increments of fractional Wiener processes. Probability Theory and Related Fields, 80(3):365–379, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01794429.

Osada:1995:IPN

[OS95] Hirofumi Osada and Toshifumi Saitoh. An invariance principle for non-symmetric Markov processes and reflecting diffusions in random domains. Probability Theory and Related Fields, 101 (1):45–63, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192195.

Olla:2013:MED

[OS13] Stefano Olla and Makiko Sasada. Macroscopic energy diffusion for a chain of anharmonic oscillators. Probability Theory and Related Fields, 157(3–4):721–775, December 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0469-5.

Osada:2016:ACS

[OS16] Hirofumi Osada and Tomoyuki Shirai. Absolute continuity and singularity of Palm measures of the Ginibre point process. Probability Theory and Related Fields, 165(3–4):725– 770, August 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0644-6;
http://link.springer.com/article/10.1007/s00440-015-

0644-6. See correction [OS22].

Osada:2022:CAC

[OS22] Hirofumi Osada and Tomoyuki Shirai. Correction to: Absolute continuity and singularity of Palm measures of the Ginibre point process. Probability Theory and Related Fields, 183 (1–2):679–680, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-021-01094-w. See [OS16].

Osada:1990:ICE

[Osa90] Hirofumi Osada. Isoperimetric constants and estimates of heat kernels of pre-Sierpinski carpets. Probability Theory and Related Fields, 86(4):469–490, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01198170.

Osada:1998:PSD

[Osa98] Hirofumi Osada. Positivity of the self-diffusion matrix of interacting Brownian particles with hard core. Probability Theory and Related Fields, 112(1):53–90, September 1998. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050183.

Osada:2001:FDP

[Osa01] Hirofumi Osada. A family of diffusion processes on Sierpinski carpets. Probability Theory and Related Fields, 119(2): 275–310, February 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008761.

Osada:2012:IDS

[Osa12] Hirofumi Osada. Infinite-dimensional stochastic differential equations related to random matrices. Probability Theory and Related Fields, 153(3–4):471–509, August 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0352-9.

Osada:2022:CID

[Osa22] Hirofumi Osada. Correction to: Infinite-dimensional stochastic differential equations related to random matrices. Probability Theory and Related Fields, 183(1–2):677–678, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01124-1. See [OT20].

Osekowski:2008:WTI

[Ose08] Adam Osekowski. Weak type inequality for noncommutative differentially subordinated martingales. Probability Theory and Related Fields, 140(3–4):553–568, March 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0075-0.

Ossiander:2005:PRS

[Oss05] Mina Ossiander. A probabilistic representation of solutions of the incompressible Navier–Stokes equations in R^{3}. Probability Theory and Related Fields, 133(2):267–298, October 2005. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0418-z.

Oh:2017:QIG

[OT17] Tadahiro Oh and Nikolay Tzvetkov. Quasi-invariant Gaussian measures for the cubic fourth order nonlinear Schrödinger equation. Probability Theory and Related Fields, 169(3–4): 1121–1168, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s00440-016-0748-7.
pdf;   http://link.springer.com/article/10.1007/s00440-
016-0748-7.

Osada:2020:IDS

[OT20] Hirofumi Osada and Hideki Tanemura. Infinite-dimensional stochastic differential equations and tail σ-fields. Probability Theory and Related Fields, 177(3–4):1137–1242, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00981-y. See correction [Osa22].

OBrien:1990:SSS

[OTV90] George L. O’Brien, Paul J. J. F. Torfs, and Wim Vervaat. Stationary self-similar extremal processes. Probability Theory and Related Fields, 87(1):97–119, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01217748.

Olesker-Taylor:2024:GBF

[OTZ24] Sam Olesker-Taylor and Luca Zanetti. Geometric bounds on the fastest mixing Markov chain. Probability Theory and Related Fields, 188(3–4):1017–1062, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-023-01257-
x.

Ouaki:2022:SCL

[Oua22] Mehdi Ouaki. Scalar conservation laws with white noise initial data. Probability Theory and Related Fields, 182(3–4):955– 998, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01083-z.

Ott:2019:AEE

[OV19] Sébastien Ott and Yvan Velenik. Asymptotics of even–even correlations in the Ising model. Probability Theory and Related Fields, 175(1–2):309–340, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0890-5.

Overbeck:1993:CSB

[Ove93] L. Overbeck. Conditioned super-Brownian motion. Probability Theory and Related Fields, 96(4):545–570, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01200209.

Overbeck:1994:PCA

[Ove94] L. Overbeck. Pathwise construction of additive h-transforms of super-Brownian motion. Probability Theory and Related Fields, 100(4):429–437, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01268988.

Owhadi:2003:AEC

[Owh03] Houman Owhadi. Approximation of the effective conductivity of ergodic media by periodization. Probability Theory and Related Fields, 125(2):225–258, February 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-002-0240-4.

Pal:2012:CMD

[Pal12] Soumik Pal. Concentration for multidimensional diffusions and their boundary local times. Probability Theory and Related Fields, 154(1–2):225–254, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0368-1.

Pancheva:1990:SDM

[Pan90] E. I. Pancheva. Selfdecomposable distributions for maxima of independent random vectors. Probability Theory and Related Fields, 84(2):267–278, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01197848.

Panchenko:2015:HEP

[Pan15] Dmitry Panchenko. Hierarchical exchangeability of pure states in mean field spin glass models. Probability Theory and Related Fields, 161(3–4):619–650, April 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0555-y.

Papangelou:1987:APT

[Pap87] F. Papangelou. On the absence of phase transition in continuous one-dimensional Gibbs systems with no hard core. Probability Theory and Related Fields, 74(4):485–496, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00363511.

Papanicolaou:1990:PST

[Pap90] Vassilis G. Papanicolaou. The probabilistic solution of the third boundary value problem for second order elliptic equations. Probability Theory and Related Fields, 87(1):27–77, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01217746.

Park:1990:EET

[Par90] Byeong Uk Park. Efficient estimation in the two-sample semiparametric location-scale models. Probability Theory and Related Fields, 86(1):21–39, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01207511.

Park:1991:AME

[Par91] Byeong Uk Park. Asymptotic minimax estimation in semiparametric models. Probability Theory and Related Fields, 88 (1):107–120, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193584.

Parraud:2022:ONN

[Par22] Félix Parraud. On the operator norm of non-commutative polynomials in deterministic matrices and iid Haar unitary matrices. Probability Theory and Related Fields, 182(3–4):751–806, April 2022. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-021-01101-0.

Peche:2006:LESa

[Péc06a] S. Péché. The largest eigenvalue of small rank perturbations of Hermitian random matrices. Probability Theory and Related Fields, 134(1):127–173, January 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-005-0466-z.

Peche:2006:LESb

[Péc06b] S. Péché. The largest eigenvalue of small rank perturbations of Hermitian random matrices. Probability Theory and Related Fields, 134(1):174, January 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-005-0480-1.

Peche:2009:URL

[Péc09] Sandrine Péché. Universality results for the largest eigenvalues of some sample covariance matrix ensembles. Probability Theory and Related Fields, 143(3–4):481–516, March 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0133-7.

Peligrad:2012:CLT

[Pel12] Magda Peligrad. Central limit theorem for triangular arrays of non-homogeneous Markov chains. Probability Theory and Related Fields, 154(3–4):409–428, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0371-6.

Pemantle:1992:VRR

[Pem92] Robin Pemantle. Vertex-reinforced random walk. Probability Theory and Related Fields, 92(1):117–136, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01205239.

Penkov:1990:PEA

[Pen90] Boyan Penkov. Persönliche Erinnerung an Andrei Nikolaevitsch Kolmogorov. (German) [Personal recollections of Andrei Nikolaevitsch Kolmogorov]. Probability Theory and Related Fields, 84(3):427–428, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01197894.

Penrose:1996:TCP

[Pen96] M. D. Penrose. The threshold contact process: A continuum limit. Probability Theory and Related Fields, 104(1):77–95, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01303804.

Peng:1999:MLT

[Pen99] Shige Peng. Monotonic limit theorem of BSDE and nonlinear decomposition theorem of Doob–Meyers type. Probability Theory and Related Fields, 113(4):473–499, March 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050214.

Penrose:2023:REC

[Pen23] Mathew D. Penrose. Random Euclidean coverage from within. Probability Theory and Related Fields, 185(3–4):747–814, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-022-01182-5.

Perkins:1992:MVB

[Per92] Edwin Perkins. Measure-valued branching diffusions with spatial interactions. Probability Theory and Related Fields, 94(2): 189–245, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192444.

Peszat:1994:LDP

[Pes94] Szymon Peszat. Large deviation principle for stochastic evolution equations. Probability Theory and Related Fields, 98 (1):113–136, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01311351.

Peskir:2002:LZB

[Pes02] Goran Peskir. Limit at zero of the Brownian first-passage density. Probability Theory and Related Fields, 124(1):100– 111, September 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400200208.

Petz:1990:FTT

[Pet90] Dénes Petz. A de Finetti-type theorem with m-dependent states. Probability Theory and Related Fields, 85(1):65–72, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01377629.

Petit:1992:TLC

[Pet92] Bernard Petit. Le théorème limite central pour des sommes de Riesz–Raikov. (French) [The Central Limit Theorem for Riesz– Raikov sums]. Probability Theory and Related Fields, 93(4): 407–438, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192715.

Petrov:2014:ARL

[Pet14] Leonid Petrov. Asymptotics of random lozenge tilings via Gelfand–Tsetlin schemes. Probability Theory and Related Fields, 160(3–4):429–487, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-013-0532-x.

Pfeifer:1987:JSA

[Pfe87] D. Pfeifer. On a joint strong approximation theorem for record and inter-record times. Probability Theory and Related Fields, 75(2):213–221, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00354034.

Picard:1989:CPP

[Pic89] Jean Picard. Convergence in probability for perturbed stochastic integral equations. Probability Theory and Related Fields, 81 (3):383–452, April 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00340060.

Picard:1996:ESD

[Pic96] Jean Picard. On the existence of smooth densities for jump processes. Probability Theory and Related Fields, 105(4):481–511, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/

10.1007/BF01191910. See erratum [Pic10].

Picard:2002:GES

[Pic02] Jean Picard. Gradient estimates for some diffusion semigroups. Probability Theory and Related Fields, 122(4):593–612, April 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400100185.

Picard:2010:EES

[Pic10] Jean Picard. Erratum to: On the existence of smooth densities for jump processes. Probability Theory and Related Fields, 147(3–4):711–713, July 2010. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s00440-
010-0267-x.pdf. See [Pic96].

Pinelis:2007:EIS

[Pin07] Iosif Pinelis. Exact inequalities for sums of asymmetric random variables, with applications. Probability Theory and Related Fields, 139(3–4):605–635, November 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0055-4.

Pisztora:1996:SOL

[Pis96] Agoston Pisztora. Surface order large deviations for Ising, Potts and percolation models. Probability Theory and Related Fields, 104(4):427–466, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01198161.

Pitman:1995:EPE

[Pit95] Jim Pitman. Exchangeable and partially exchangeable random partitions. Probability Theory and Related Fields, 102 (2):145–158, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01213386.

Pitman:1996:CSB

[Pit96] Jim Pitman. Cyclically stationary Brownian local time processes. Probability Theory and Related Fields, 106(3):299– 329, November 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050066.

Popescu:2009:GTR

[Pop09] Ionel Popescu. General tridiagonal random matrix models, limiting distributions and fluctuations. Probability Theory and Related Fields, 144(1–2):179–220, May 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-008-0145-y.

Portnoy:1986:CLT

[Por86] Stephen Portnoy. On the central limit theorem in R^{p} when p → ∞. Probability Theory and Related Fields, 73(4):571– 583, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00324853.

Porod:1995:LBS

[Por95] Ursula Porod. L_{2}-lower bounds for a special class of random walks. Probability Theory and Related Fields, 101(2):277–289, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01375829.

Porod:1996:PRR

[Por96] U. Porod. The out-off phenomenon for random reflections II: Complex and quaternionic cases. Probability Theory and Related Fields, 104(2):181–209, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01247837.

Pourahmadi:1988:ARM

[Pou88] Mohsen Pourahmadi. Autoregressive representations of multivariate stationary stochastic processes. Probability Theory and Related Fields, 80(2):315–322, December 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00356109.

Povel:1999:CBM

[Pov99] Tobias Povel. Confinement of Brownian motion among Poissonian obstacles in R^{d}, d ≥ 3. Probability Theory and Related Fields, 114(2):177–205, May 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s440-1999-8036-0.

Powell:2019:IPB

[Pow19] Ellen Powell. An invariance principle for branching diffusions in bounded domains. Probability Theory and Related Fields, 173 (3–4):999–1062, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0847-8.

Pardoux:1987:TSS

[PP87] E. Pardoux and P. Protter. A two-sided stochastic integral and its calculus. Probability Theory and Related Fields, 76(1): 15–49, September 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00390274.

Pietruska-Paluba:1991:LSD

[PP91] Katarzyna Pietruska-Paluba. The Lifschitz singularity for the density of states on the Sierpinski gasket. Probability Theory and Related Fields, 89(1):1–33, ???? 1991. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01225823.

Pardoux:1994:BDS

[PP94] Etienne Pardoux and Shige Peng. Backward doubly stochastic differential equations and systems of quasilinear SPDEs. Probability Theory and Related Fields, 98(2):209–227, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192514.

Pemantle:1996:TSB

[PPS96] Robin Pemantle, Yuval Peres, and Jonathan W. Shapiro. The trace of spatial Brownian motion is capacity-equivalent to the unit square. Probability Theory and Related Fields, 106(3): 379–399, November 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050070.

Perman:1992:SBS

[PPY92] Mihael Perman, Jim Pitman, and Marc Yor. Size-biased sampling of Poisson point processes and excursions. Probability Theory and Related Fields, 92(1):21–39, ???? 1992. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01205234.

Pisztora:1999:PLD

[PPZ99] Agoston Pisztora, Tobias Povel, and Ofer Zeitouni. Precise large deviation estimates for a one-dimensional random walk in a random environment. Probability Theory and Related Fields, 113(2):191–219, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050206.

Paouris:2014:CLT

[PPZ14] Grigoris Paouris, Peter Pivovarov, and Joel Zinn. A central limit theorem for projections of the cube. Probability Theory and Related Fields, 159(3–4):701–719, August 2014. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
013-0518-8.

Petz:2005:LDE

[PR05] Dénes Petz and Júlia Réffy. Large deviation for the empirical eigenvalue density of truncated Haar unitary matrices. Probability Theory and Related Fields, 133(2):175–189, October 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0420-5.

Puchala:2008:EAC

[PR08] Zbigniew Puchala and Tomasz Rolski. The exact asymptotic of the collision time tail distribution for independent Brownian particles with different drifts. Probability Theory and Related Fields, 142(3–4):595–617, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-007-0116-8.

Pitman:2012:GCM

[PR12] Jim Pitman and Nathan Ross. The greatest convex minorant of Brownian motion, meander, and bridge. Probability Theory and Related Fields, 153(3–4):771–807, August 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0385-0.

Prigent:2020:NSE

[PR20] Martin Prigent and Matthew I. Roberts. Noise sensitivity and exceptional times of transience for a simple symmetric random walk in one dimension. Probability Theory and Related Fields, 178(1–2):327–367, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-020-00978-7.

Prat:1993:RES

[Pra93] J. J. Prat. Résolution de l’équation de Schrödinger par intégration stochastique du champ magnétique. (French) [Solving the Schrödinger equation by stochastic integration of the magnetic field]. Probability Theory and Related Fields, 95(2): 217–235, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192271.

Privault:1996:GTA

[Pri96a] N. Privault. Girsanov theorem for anticipative shifts on Poisson space. Probability Theory and Related Fields, 104(1):61–76, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01303803.

Privault:1996:DQS

[Pri96b] Nicolas Privault. A different quantum stochastic calculus for the Poisson process. Probability Theory and Related Fields, 105 (2):255–278, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01203837.

Procaccia:2016:QIP

[PRS16] Eviatar B. Procaccia, Ron Rosenthal, and Artëm Sapozhnikov. Quenched invariance principle for simple random walk on clusters in correlated percolation models. Probability Theory and Related Fields, 166(3–4):619–657, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0668-y;   http://link.springer.com/
article/10.1007/s00440-015-0668-y.

Pfander:2013:RIP

[PRT13] Götz E. Pfander, Holger Rauhut, and Joel A. Tropp. The restricted isometry property for time-frequency structured random matrices. Probability Theory and Related Fields, 156(3–4): 707–737, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0441-4.

Pruss:1998:BTI

[Pru98] Alexander R. Pruss. A bounded N-tuplewise independent and identically distributed counterexample to the CLT. Probability Theory and Related Fields, 111(3):323–332, July 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050170.

Parthasarathy:1987:STF

[PS87] K. R. Parthasarathy and Kalyan B. Sinha. Stop times in Fock space stochastic calculus. Probability Theory and Related Fields, 75(3):317–349, July 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00318706.

Protter:1993:GCV

[PS93] Philip Protter and Jaime San Martín. General change of variable formulas for semimartingales in one and finite dimensions. Probability Theory and Related Fields, 97(3):363–381, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01195071.

Peres:1998:NIC

[PS98] Yuval Peres and Jeffrey E. Steif. The number of infinite clusters in dynamical percolation. Probability Theory and Related Fields, 111(1):141–165, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050165.

Peres:2013:MUT

[PS13] Yuval Peres and Allan Sly. Mixing of the upper triangular matrix walk. Probability Theory and Related Fields, 156(3–4): 581–591, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-012-0436-1.

Profeta:2015:PIS

[PS15] Christophe Profeta and Thomas Simon. Persistence of integrated stable processes. Probability Theory and Related Fields, 162(3–4):463–485, August 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-014-0577-5.

Peres:2016:DFB

[PS16] Yuval Peres and Perla Sousi. Dimension of fractional Brownian motion with variable drift. Probability Theory and Related Fields, 165(3–4):771–794, August 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0645-5;   http://link.springer.com/
article/10.1007/s00440-015-0645-5.

Panagiotou:2018:SLR

[PS18] Konstantinos Panagiotou and Benedikt Stufler. Scaling limits of random Pólya trees. Probability Theory and Related Fields, 170 (3–4):801–820, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-017-0770-4.

Petrov:2022:MTB

[PS22] Leonid Petrov and Axel Saenz. Mapping TASEP back in time. Probability Theory and Related Fields, 182(1–2):481–530, February 2022. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01074-0.

Peres:2015:RWD

[PSS15] Yuval Peres, Alexandre Stauffer, and Jeffrey E. Steif. Random walks on dynamical percolation: mixing times, mean squared displacement and hitting times. Probability Theory and Related Fields, 162(3–4):487–530, August 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-014-0578-4.

Peres:2020:MTR

[PSS20] Yuval Peres, Perla Sousi, and Jeffrey E. Steif. Mixing time for random walk on supercritical dynamical percolation. Probability Theory and Related Fields, 176(3–4):809–849, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-019-00927-z;   https://link.springer.
com/content/pdf/10.1007/s00440-019-00927-z.pdf.

Peres:2013:MGG

[PSSS13] Yuval Peres, Alistair Sinclair, Perla Sousi, and Alexandre Stauffer. Mobile geometric graphs: detection, coverage and percolation. Probability Theory and Related Fields, 156(1–2):273–305, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-012-0428-1.

Perkins:1987:UMR

[PT87] Edwin A. Perkins and S. James Taylor. Uniform measure results for the image of subsets under Brownian motion. Probability Theory and Related Fields, 76(3):257–289, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01297485.

Pardoux:1999:FBS

[PT99] Etienne Pardoux and Shanjian Tang. Forward-backward stochastic differential equations and quasilinear parabolic PDEs. Probability Theory and Related Fields, 114(2):123–150, May 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004409970001.

Pipiras:2000:IQR

[PT00] Vladas Pipiras and Murad S. Taqqu. Integration questions related to fractional Brownian motion. Probability Theory and Related Fields, 118(2):251–291, October 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s440-000-8016-7.

Pipiras:2002:DSS

[PT02] Vladas Pipiras and Murad S. Taqqu. Decomposition of selfsimilar stable mixed moving averages. Probability Theory and Related Fields, 123(3):412–452, July 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400200196.

Panchenko:2004:BDM

[PT04] Dmitry Panchenko and Michel Talagrand. Bounds for diluted mean-fields spin glass models. Probability Theory and Related Fields, 130(3):319–336, November 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0342-2.

Pinsky:2017:TRS

[PT17] Ross G. Pinsky and Nicholas F. Travers. Transience, recurrence and the speed of a random walk in a site-based feedback environment. Probability Theory and Related Fields, 167 (3–4):917–978, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-016-0695-3.

Peres:2020:CPR

[PTZ20] Yuval Peres, Ryokichi Tanaka, and Alex Zhai. Cutoff for product replacement on finite groups. Probability Theory and Related Fields, 177(3–4):823–853, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-020-00962-
1.

Pfister:1997:LDC

[PV97] C.-E. Pfister and Y. Velenik. Large deviations and continuum limit in the 2D Ising model. Probability Theory and Related Fields, 109(4):435–506, November 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050139.

Paindaveine:2020:DDS

[PV20] Davy Paindaveine and Thomas Verdebout. Detecting the direction of a signal on high-dimensional spheres: non-null and Le Cam optimality results. Probability Theory and Related Fields, 176(3–4):1165–1216, April 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-019-00937-x.

Perman:1997:PBM

[PW97] Mihael Perman and Wendelin Werner. Perturbed Brownian motions. Probability Theory and Related Fields, 108(3):357–383, July 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050113.

Pal:2020:MSP

[PW20] Soumik Pal and Ting-Kam Leonard Wong. Multiplicative Schrödinger problem and the Dirichlet transport. Probability Theory and Related Fields, 178(1–2):613–654, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00987-6.

Pfaffelhuber:2011:TLE

[PWW11] P. Pfaffelhuber, A. Wakolbinger, and H. Weisshaupt. The tree length of an evolving coalescent. Probability Theory and Related Fields, 151(3–4):529–557, December 2011. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-010-0307-6.

Pisier:1988:SVM

[PX88] Gilles Pisier and Quanhua Xu. The strong p-variation of martingales and orthogonal series. Probability Theory and Related Fields, 77(4):497–514, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00959613.

Patzschke:1993:SSR

[PZ93] N. Patzschke and M. Zähle. Self-similar random measures are locally scale invariant. Probability Theory and Related Fields, 97(4):559–574, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192964.

Pardoux:1998:GBN

[PZ98] Etienne Pardoux and Shuguang Zhang. Generalized BSDEs and nonlinear Neumann boundary value problems. Probability Theory and Related Fields, 110(4):535–558, May 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050158.

Peszat:2000:NSW

[PZ00] Szymon Peszat and Jerzy Zabczyk. Nonlinear stochastic wave and heat equations. Probability Theory and Related Fields, 116 (3):421–443, March 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050257.

Peres:2008:CLT

[PZ08] Yuval Peres and Ofer Zeitouni. A central limit theorem for biased random walks on Galton–Watson trees. Probability Theory and Related Fields, 140(3–4):595–629, March 2008. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0077-y.

Priola:2011:SPS

[PZ11] Enrico Priola and Jerzy Zabczyk. Structural properties of semilinear SPDEs driven by cylindrical stable processes. Probability Theory and Related Fields, 149(1–2):97–137, February 2011.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-009-0243-5.

Quiroz:1991:SNT

[QD91] Adolfo J. Quiroz and R. M. Dudley. Some new tests for multivariate normality. Probability Theory and Related Fields, 87 (4):521–546, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01304278.

Qian:2018:CRT

[Qia18] Wei Qian. Conformal restriction: the trichordal case. Probability Theory and Related Fields, 171(3–4):709–774, August 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-017-0791-z.

Qian:2021:GDE

[Qia21] Wei Qian. Generalized disconnection exponents. Probability Theory and Related Fields, 179(1–2):117–164, February 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-01005-5.

Quastel:2013:LBH

[QR13] Jeremy Quastel and Daniel Remenik. Local behavior and hitting probabilities of the Airy_{1} process. Probability Theory and Related Fields, 157(3–4):605–634, December 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0466-8.

Quastel:1999:LDS

[QRV99] J. Quastel, F. Rezakhanlou, and S. R. S. Varadhan. Large deviations for the symmetric simple exclusion process in dimensions d ≥ 3. Probability Theory and Related Fields, 113 (1):1–84, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050202.

Qian:2003:CTE

[QRZ03] Zhongmin Qian, Francesco Russo, and Weian Zheng. Comparison theorem and estimates for transition probability densities of diffusion processes. Probability Theory and Related Fields, 127(3):388–406, November 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0291-1.

Qi:2024:LDP

[QSY24] Weiwei Qi, Zhongwei Shen, and Yingfei Yi. Large deviation principle for quasi-stationary distributions and multiscale dynamics of absorbed singular diffusions. Probability Theory and Related Fields, 188(1–2):667–728, February 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01246-0.

Quas:1999:IPL

[Qua99] Anthony N. Quas. Infinite paths in a Lorentz lattice gas model. Probability Theory and Related Fields, 114(2):229–244, May 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s440-1999-8037-y.

Quint:2010:PTD

[Qui10] J.-F. Quint. Percolation on the three dot system. Probability Theory and Related Fields, 146(1–2):49–60, January 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
008-0183-5.

Raimond:1997:SAD

[Rai97a] Olivier Raimond. Self-attracting diffusions: Case of the constant interaction. Probability Theory and Related Fields, 107(2): 177–196, February 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050082.

Rains:1997:HPR

[Rai97b] E. M. Rains. High powers of random elements of compact Lie groups. Probability Theory and Related Fields, 107(2): 219–241, February 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050084.

Rains:1998:NLT

[Rai98] Eric M. Rains. Normal limit theorems for symmetric random matrices. Probability Theory and Related Fields, 112(3):411– 423, November 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050195.

Rains:2003:IED

[Rai03] Eric M. Rains. Images of eigenvalue distributions under power maps. Probability Theory and Related Fields, 125(4):522–538, April 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-002-0250-2.

Raimond:2009:SID

[Rai09] Olivier Raimond. Self-interacting diffusions: a simulated annealing version. Probability Theory and Related Fields, 144(1– 2):247–279, May 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0147-9.

Ramasubramanian:1988:HSD

[Ram88] S. Ramasubramanian. Hitting of submanifolds by diffusions. Probability Theory and Related Fields, 78(1):149–163, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00718043.

Ramasubramanian:1990:HSD

[Ram90] S. Ramasubramanian. Hitting of submanifolds by diffusions. Probability Theory and Related Fields, 84(2):279, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01197849.

Ramirez:1998:REM

[Ram98] A. F. Ramírez. Relative entropy and mixing properties of infinite dimensional diffusions. Probability Theory and Related Fields, 110(3):369–395, March 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050152.

Rassoul-Agha:2005:ASI

[RAS05] Firas Rassoul-Agha and Timo Seppäläinen. An almost sure invariance principle for random walks in a space–time random environment. Probability Theory and Related Fields, 133(3): 299–314, November 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0424-1.

Rassoul-Agha:2014:QPP

[RAS14] Firas Rassoul-Agha and Timo Seppäläinen. Quenched pointto-point free energy for random walks in random potentials. Probability Theory and Related Fields, 158(3–4):711–750, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-013-0494-z.

Raugi:1988:TCD

[Rau88] Albert Raugi. Un théorème de Choquet–Deny pour les groupes moyennables. (French) [A theorem of Choquet–Deny for amenable groups]. Probability Theory and Related Fields, 77 (4):481–496, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00959612.

Reynaud-Bouret:2003:AEI

[RB03] Patricia Reynaud-Bouret. Adaptive estimation of the intensity of inhomogeneous Poisson processes via concentration inequalities. Probability Theory and Related Fields, 126(1):103–153, May 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0259-1.

Rohde:2013:SIO

[RD13] Angelika Rohde and Lutz Dümbgen. Statistical inference for the optimal approximating model. Probability Theory and Related Fields, 155(3–4):839–865, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0414-7.

Regts:2023:AZI

[Reg23] Guus Regts. Absence of zeros implies strong spatial mixing. Probability Theory and Related Fields, 186(1–2):621–641, June 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01190-z.

Reimers:1989:HMA

[Rei89a] Mark Reimers. Hyperfinite methods applied to the critical branching diffusion. Probability Theory and Related Fields, 81 (1):11–27, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00343736.

Reimers:1989:ODS

[Rei89b] Mark Reimers. One dimensional stochastic partial differential equations and the branching measure diffusion. Probability Theory and Related Fields, 81(3):319–340, April 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00340057.

Reitzner:2005:CLT

[Rei05] Matthias Reitzner. Central limit theorems for random polytopes. Probability Theory and Related Fields, 133(4):483– 507, December 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-005-0441-8.

Ressel:1988:IRD

[Res88] Paul Ressel. Integral representations for distributions of symmetric stochastic processes. Probability Theory and Related Fields, 79(3):451–467, October 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00342235.

Revesz:1992:BHP

[Rév92] P. Révész. Black holes on the plane drawn by a Wiener process. Probability Theory and Related Fields, 93(1):21–37, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01195386.

Rezakhanlon:1996:KLC

[Rez96] F. Rezakhanlon. Kinetic limits for a class of interacting particle systems. Probability Theory and Related Fields, 104(1):97–146, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01303805.

Rhodes:2009:DLS

[Rho09] Rémi Rhodes. Diffusion in a locally stationary random environment. Probability Theory and Related Fields, 143(3–4):545–568, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-007-0135-5.

Richier:2018:LBR

[Ric18] Loïc Richier. Limits of the boundary of random planar maps. Probability Theory and Related Fields, 172(3–4):789–827, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-017-0820-y.

Rider:2004:DCL

[Rid04] B. Rider. Deviations from the circular law. Probability Theory and Related Fields, 130(3):337–367, November 2004. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0355-x.

Rio:1994:LIP

[Rio94] Emmanuel Rio. Local invariance principles and their application to density estimation. Probability Theory and Related Fields, 98 (1):21–45, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01311347.

Rio:1996:TBE

[Rio96] Emmanuel Rio. Sur le théorème de Berry–Esseen pour les suites faiblement dépendantes. (French) [On the Berry-Esseen theorem for weakly dependent sequences]. Probability Theory and Related Fields, 104(2):255–282, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01247840.

Rio:1998:PEA

[Rio98] Emmanuel Rio. Processus empiriques absolument réguliers et entropie universelle. (French) [Absolutely regular empirical processes and universal entropy]. Probability Theory and Related Fields, 111(4):585–608, August 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050179.

Rio:2000:LFG

[Rio00] Emmanuel Rio. Lois fortes des grands nombres presque sûres pour les sommes de Riesz–Raikov. (French) [The almost-sure Strong Law of Large Numbers on Riesz–Raikov sums]. Probability Theory and Related Fields, 118(3):342–348, November 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008745.

Rio:2001:ICP

[Rio01] Emmanuel Rio. Inégalités de concentration pour les processus empiriques de classes de parties. (French) [Concentration inequalities for empirical processes of classes of parts]. Probability Theory and Related Fields, 119(2):163–175, February 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
PL00008756.

Robinson:1994:RCO

[Rob94] P. M. Robinson. Rates of convergence and optimal spectral bandwidth for long range dependence. Probability Theory and Related Fields, 99(3):443–473, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01199901.

Roberto:2010:SDG

[Rob10] Cyril Roberto. Slow decay of Gibbs measures with heavy tails. Probability Theory and Related Fields, 148(1–2):247– 268, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-009-0229-3.

Rodriguez:2017:LMG

[Rod17] Pierre-François Rodriguez. A 0–1 law for the massive Gaussian free field. Probability Theory and Related Fields, 169(3–4): 901–930, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-016-0743-z.

Rodriguez:2019:PFI

[Rod19] Pierre-François Rodriguez. On pinned fields, interlacements, and random walk on (Z/NZ)^{2}. Probability Theory and Related Fields, 173(3–4):1265–1299, April 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-018-0851-z.

Rogers:1987:CMB

[Rog87] L. C. G. Rogers. Continuity of martingales in the Brownian excursion filtration. Probability Theory and Related Fields, 76 (3):291–298, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01297486.

Rogers:1993:JLM

[Rog93] L. C. G. Rogers. The joint law of the maximum and terminal value of a martingale. Probability Theory and Related Fields, 95 (4):451–466, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01196729.

Rohde:2011:OCM

[Roh11] Angelika Rohde. Optimal calibration for multiple testing against local inhomogeneity in higher dimension. Probability Theory and Related Fields, 149(3–4):515–559, April 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0263-1.

Rolski:1990:EPP

[Rol90] Tomasz Rolski. Ergodic properties of Poisson processes with almost periodic intensity. Probability Theory and Related Fields, 84(1):27–37, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01288556.

Rolles:2003:HER

[Rol03] Silke W. W. Rolles. How edge-reinforced random walk arises naturally. Probability Theory and Related Fields, 126(2):243– 260, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/s00440-003-0260-8.

Romersperger:1996:NNV

[Röm96] M. Römersperger. A note on nearby variables with nearby conditional laws. Probability Theory and Related Fields, 106(3): 371–377, November 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050069.

Romito:2014:UBS

[Rom14] Marco Romito. Uniqueness and blow-up for a stochastic viscous dyadic model. Probability Theory and Related Fields, 158(3– 4):895–924, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0499-7.

Rosenblatt:1987:RLT

[Ros87] Murray Rosenblatt. Remarks on limit theorems for nonlinear functionals of Gaussian sequences. Probability Theory and Related Fields, 75(1):1–10, May 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00320078.

Rosenthal:1988:FUG

[Ros88] A. Rosenthal. Finite uniform generators for ergodic, finite entropy, free actions of amenable groups. Probability Theory and Related Fields, 77(2):147–166, February 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00334034.

Rosenkrantz:1989:ECT

[Ros89] Walter A. Rosenkrantz. Ergodicity conditions for twodimensional Markov chains on the positive quadrant. Probability Theory and Related Fields, 83(3):309–319, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00964367.

Roy:1991:PPS

[Roy91] Rahul Roy. Percolation of Poisson sticks on the plane. Probability Theory and Related Fields, 89(4):503–517, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01199791.

Rozovsky:1986:AEP

[Roz86] L. V. Rozovsky. Asymptotic expansions for probabilities of large deviations. Probability Theory and Related Fields, 73(2): 299–318, September 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00339942.

Rozkosz:2003:BSC

[Roz03] Andrzej Rozkosz. Backward SDEs and Cauchy problem for semilinear equations in divergence form. Probability Theory and Related Fields, 125(3):393–407, March 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-002-0244-0.

Rajput:1989:SRI

[RR89] Balram S. Rajput and Jan Rosinski. Spectral representations of infinitely divisible processes. Probability Theory and Related Fields, 82(3):451–487, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00339998.

Rachev:1992:NIM

[RR92] S. T. Rachev and L. Rüschendorf. A new ideal metric with applications to multivariate stable limit theorems. Probability Theory and Related Fields, 94(2):163–187, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192443.

Ramachandran:1995:GDT

[RR95] D. Ramachandran and L. Rüschendorf. A general duality theorem for marginal problems. Probability Theory and Related Fields, 101(3):311–319, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01200499.

Ren:2000:RHC

[RR00] Jiagang Ren and Michael Röckner. Ray Hölder-continuity for fractional Sobolev spaces in infinite dimensions and applications. Probability Theory and Related Fields, 117(2):201–220, June 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050004.

Rinott:2003:EED

[RR03] Yosef Rinott and Vladimir Rotar. On Edgeworth expansions for dependency-neighborhoods chain structures and Stein’s method. Probability Theory and Related Fields, 126(4):528–570, August 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0271-5.

Ramirez:2017:SRM

[RR17] José A. Ramírez and Brian Rider. Spiking the random matrix hard edge. Probability Theory and Related Fields, 169(1–2): 425–467, October 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-016-0733-1.

Riera:2024:SLT

[RRO24] Armand Riera and Alejandro Rosales-Ortiz. The structure of the local time of Markov processes indexed by Lévy trees. Probability Theory and Related Fields, 189(1–2):1–99, June 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01258-w.

Ressel:1991:NCL

[RS91] Paul Ressel and Walter Schmidtchen. A new characterization of Laplace functionals and probability generating functionals. Probability Theory and Related Fields, 88(2):195–213, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01212559.

Rogers:1993:IBP

[RS93] L. C. G. Rogers and Z. Shi. Interacting Brownian particles and the Wigner law. Probability Theory and Related Fields, 95 (4):555–570, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01196734.

Rajput:1994:SEP

[RS94] Balram S. Rajput and Carl Sundberg. On some extremal problems in H^{p} and the prediction of L^{p}-harmonizable stochastic processes. Probability Theory and Related Fields, 99(2):197–210, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01199022.

Raimond:2012:EBM

[RS12] Olivier Raimond and Bruno Schapira. Excited Brownian motions as limits of excited random walks. Probability Theory and Related Fields, 154(3–4):875–909, December 2012. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
011-0388-x.

Roch:2017:PTS

[RS17] Sebastien Roch and Allan Sly. Phase transition in the sample complexity of likelihood-based phylogeny inference. Probability Theory and Related Fields, 169(1–2):3–62, October 2017. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0793-x.

Rosenzweig:2023:GSA

[RS23] Matthew Rosenzweig and Gigliola Staffilani. Global solutions of aggregation equations and other flows with random diffusion. Probability Theory and Related Fields, 185(3–4):1219– 1262, April 2023. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-022-01171-8.

Restrepo:2013:IMC

[RST^{+}13] Ricardo Restrepo, Jinwoo Shin, Prasad Tetali, Eric Vigoda, and Linji Yang. Improved mixing condition on the grid for counting and sampling independent sets. Probability Theory and Related Fields, 156(1–2):75–99, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0421-8.

Rakhlin:2015:SCU

[RST15] Alexander Rakhlin, Karthik Sridharan, and Ambuj Tewari. Sequential complexities and uniform martingale laws of large numbers. Probability Theory and Related Fields, 161(1–2): 111–153, February 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0545-5.

Rosenberg:2001:SGR

[RSV01] Dinah Rosenberg, Eilon Solan, and Nicolas Vieille. Stopping games with randomized strategies. Probability Theory and Related Fields, 119(3):433–451, March 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/PL00008766.

Rider:2013:SMC

[RSX13] Brian Rider, Christopher D. Sinclair, and Yuan Xu. A solvable mixed charge ensemble on the line: global results. Probability Theory and Related Fields, 155(1–2):127–164, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-011-0394-z.

Rockner:2023:MSW

[RSZ23] Michael Röckner, Yiming Su, and Deng Zhang. Multi solitary waves to stochastic nonlinear Schrödinger equations. Probability Theory and Related Fields, 186(3–4):813–876, August 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01201-z.

Rhee:1986:UCD

[RT86] WanSoo Rhee and Michel Talagrand. Uniform convexity and the distribution of the norm for a Gaussian measure. Probability Theory and Related Fields, 71(1):59–67, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00366272.

Roelly:2002:CRP

[RT02] Sylvie Roelly and Michèle Thieullen. A characterization of reciprocal processes via an integration by parts formula on the path space. Probability Theory and Related Fields, 123(1):97–120, May 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400100184.

Rucinski:1988:WSS

[Ruc88] Andrzej Ruciński. When are small subgraphs of a random graph normally distributed? Probability Theory and Related Fields, 78(1):1–10, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00718031.

Rukhin:1986:ATS

[Ruk86] Andrew L. Rukhin. Adaptive tests in statistical problems with finite nuisance parameter. Probability Theory and Related Fields, 73(4):529–538, November 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00324850.

Rutkowski:1990:SSD

[Rut90] M. Rutkowski. On solutions of stochastic differential equations with drift. Probability Theory and Related Fields, 85(3):387–402, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01193944.

Russo:1993:FBS

[RV93] Francesco Russo and Pierre Vallois. Forward, backward and symmetric stochastic integration. Probability Theory and Related Fields, 97(3):403–421, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01195073.

Russo:1996:IFF

[RV96] F. Russo and P. Vallois. Ito formula for C^{1}-functions of semimartingales. Probability Theory and Related Fields, 104 (1):27–41, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01303801.

Reiman:1988:BPS

[RW88] M. I. Reiman and R. J. Williams. A boundary property of semimartingale reflecting Brownian motions. Probability Theory and Related Fields, 77(1):87–97, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01848132. See cor-

rection [RW89].

Reiman:1989:CBP

[RW89] M. I. Reiman and R. J. Williams. Correction to: A boundary property of semimartingale reflecting Brownian motions. Probability Theory and Related Fields, 80(4):633, February 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00318910. See [RW88].

Rogers:1991:ILT

[RW91] L. C. G. Rogers and J. B. Walsh. The intrinsic local time sheet of Brownian motion. Probability Theory and Related Fields, 88 (3):363–379, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01418866.

Rempala:2005:ATR

[RW05] Grzegorz A. Rempala and Jacek Wesolowski. Approximation theorems for random permanents and associated stochastic processes. Probability Theory and Related Fields, 131(3):442–458, March 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-004-0380-9.

Rider:2019:USB

[RW19] Brian Rider and Patrick Waters. Universality of the stochastic Bessel operator. Probability Theory and Related Fields, 175(1– 2):97–140, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-018-0888-z.

Rockner:2020:SPN

[RXZ20] Michael Röckner, Longjie Xie, and Xicheng Zhang. Superposition principle for non-local Fokker–Planck–Kolmogorov operators. Probability Theory and Related Fields, 178(3–4):699–733, December 2020. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-020-00985-8.

Rockner:2009:STN

[RZ09] Michael Röckner and Xicheng Zhang. Stochastic tamed 3D Navier–Stokes equations: existence, uniqueness and ergodicity. Probability Theory and Related Fields, 145(1–2):211– 267, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0167-5.

Rohde:2016:BSS

[RZ16] Steffen Rohde and Dapeng Zhan. Backward SLE and the symmetry of the welding. Probability Theory and Related Fields, 164 (3–4):815–863, April 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0620-1.

Rezaei:2018:GFC

[RZ18] Mohammad A. Rezaei and Dapeng Zhan. Green’s functions for chordal SLE curves. Probability Theory and Related Fields, 171(3–4):1093–1155, August 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0802-0.

Sabot:2011:RWR

[Sab11] Christophe Sabot. Random walks in random Dirichlet environment are transient in dimension d ≥ 3. Probability Theory and Related Fields, 151(1–2):297–317, October 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0300-0.

Saisho:1987:SDE

[Sai87] Yasumasa Saisho. Stochastic differential equations for multidimensional domain with reflecting boundary. Probability Theory and Related Fields, 74(3):455–477, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00699100.

Salisbury:1986:CRP

[Sal86a] Thomas S. Salisbury. Construction of right processes from excursions. Probability Theory and Related Fields, 73(3):351–367, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00776238.

Salisbury:1986:IEP

[Sal86b] Thomas S. Salisbury. On the Itô excursion process. Probability Theory and Related Fields, 73(3):319–350, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00776237.

Salminen:1990:MEF

[Sal90] Paavo Salminen. On multiplicative excessive functions of a branching Brownian motion. Probability Theory and Related Fields, 85(1):43–56, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01377627.

Samur:1987:IPS

[Sam87] Jorge D. Samur. On the invariance principle for stationary φmixing triangular arrays with infinitely divisible limits. Probability Theory and Related Fields, 75(2):245–259, June 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00354036.

Samson:2022:ECG

[Sam22] Paul-Marie Samson. Entropic curvature on graphs along Schrödinger bridges at zero temperature. Probability Theory and Related Fields, 184(3–4):859–937, December 2022. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-022-01167-4.

Sanchis:1994:FLTa

[San94a] Gabriela R. Sanchis. A functional limit theorem for Erdős and Rényi’s law of large numbers. Probability Theory and Related Fields, 98(1):1–5, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01311345.

Sanchis:1994:FLTb

[San94b] Gabriela R. Sanchis. A functional limit theorem for Erdős and Rényi’s law of large numbers. Probability Theory and Related Fields, 99(3):475, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01199902.

Sasvari:1986:CDR

[Sas86] Zoltán Sasvári. Characterizing the distributions of the random variables X_{1}, X_{2}, X_{3} by the distribution of ( X_{1}- X_{3}, X_{2}- X_{3} ). Probability Theory and Related Fields, 73(1):43–49, ???? 1986.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01845992.

Savov:2009:STT

[Sav09] Mladen Savov. Small time two-sided LIL behavior for Lévy processes at zero. Probability Theory and Related Fields, 144 (1–2):79–98, May 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0142-1.

Szekely:2003:EPG

[SB03] Gábor J. Székely and Nail K. Bakirov. Extremal probabilities for Gaussian quadratic forms. Probability Theory and Related Fields, 126(2):184–202, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0262-6.

Sur:2019:LRT

[SCC19] Pragya Sur, Yuxin Chen, and Emmanuel J. Candès. The likelihood ratio test in high-dimensional logistic regression is asymptotically a rescaled chi-square. Probability Theory and Related Fields, 175(1–2):487–558, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
018-00896-9.

Scheutzow:1986:PBS

[Sch86] Michael Scheutzow. Periodic behavior of the stochastic Brusselator in the mean-field limit. Probability Theory and Related Fields, 72(3):425–462, June 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00334195.

Schatte:1988:LIL

[Sch88a] Peter Schatte. On a law of the iterated logarithm for sums mod 1 with application to Benford’s law. Probability Theory and Related Fields, 77(2):167–178, February 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00334035.

Schneller:1988:SPM

[Sch88b] W. Schneller. A short proof of Motoo’s combinatorial central limit theorem using Stein’s method. Probability Theory and Related Fields, 78(2):249–252, June 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00322021.

Schief:1989:ASC

[Sch89a] Andreas Schief. Almost surely convergent random variables with given laws. Probability Theory and Related Fields, 81 (4):559–567, May 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00367303.

Schonmann:1989:ECU

[Sch89b] Roberto H. Schonmann. Exponential convergence under mixing. Probability Theory and Related Fields, 81(2):235–238, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00319552.

Schurmann:1990:NSP

[Sch90] Michael Schürmann. Noncommutative stochastic processes with independent and stationary increments satisfy quantum stochastic differential equations. Probability Theory and Related Fields, 84(4):473–490, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01198315.

Schied:1996:SPL

[Sch96] A. Schied. Sample path large deviations for super-Brownian motion. Probability Theory and Related Fields, 104(3):319–347, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01213684.

Schilling:1998:GHC

[Sch98] René L. Schilling. Growth and Hölder conditions for the sample paths of Feller processes. Probability Theory and Related Fields, 112(4):565–611, December 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400050201.

Schonmann:1999:SIC

[Sch99] Roberto H. Schonmann. Stability of infinite clusters in supercritical percolation. Probability Theory and Related Fields, 113(2):287–300, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050209.

Scheffler:2000:LIL

[Sch00] Hans-Peter Scheffler. A law of the iterated logarithm for heavytailed random vectors. Probability Theory and Related Fields, 116(2):257–271, February 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/PL00008729.

Schultz:2005:NCP

[Sch05] Hanne Schultz. Non-commutative polynomials of independent Gaussian random matrices. the real and symplectic cases. Probability Theory and Related Fields, 131(2):261–309, February 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-004-0366-7.

Schapira:2007:HOM

[Sch07] Bruno Schapira. The Heckman–Opdam Markov processes. Probability Theory and Related Fields, 138(3–4):495–519, July 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-006-0034-1.

Schweinsberg:2009:LER

[Sch09] Jason Schweinsberg. The loop-erased random walk and the uniform spanning tree on the four-dimensional discrete torus. Probability Theory and Related Fields, 144(3–4):319–370, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-008-0149-7.

Schnellmann:2015:LIL

[Sch15] Daniel Schnellmann. Law of iterated logarithm and invariance principle for one-parameter families of interval maps. Probability Theory and Related Fields, 162(1–2):365–409, June 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0575-7.

Schoug:2020:MBS

[Sch20] Lukas Schoug. A multifractal boundary spectrum for SLE_{κ}(ρ). Probability Theory and Related Fields, 178(1–2):173–233, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-020-00975-w.

Schonbauer:2023:MCD

[Sch23] Philipp Schönbauer. Malliavin calculus and densities for singular stochastic partial differential equations. Probability Theory and Related Fields, 186(3–4):643–713, August 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01207-7.

Saloff-Coste:2021:IPR

[SCZ21] Laurent Saloff-Coste and Tianyi Zheng. Isoperimetric profiles and random walks on some groups defined by piecewise actions. Probability Theory and Related Fields, 181(1-3):711–756, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-021-01067-z.

Seignourel:2000:DSP

[Sei00] Paul Seignourel. Discrete schemes for processes in random media. Probability Theory and Related Fields, 118(3):293– 322, November 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008743.

Seppalainen:1993:LDL

[Sep93] Timo Seppäläinen. Large deviations for lattice systems I. Parametrized independent fields. Probability Theory and Related Fields, 96(2):241–260, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01192135.

Seppalainen:1995:MEP

[Sep95] Timo Seppäläinen. Maximum entropy principles for disordered spins. Probability Theory and Related Fields, 101(4):547–576, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01202784.

Seppalainen:1998:LDI

[Sep98] Timo Seppäläinen. Large deviations for increasing sequences on the plane. Probability Theory and Related Fields, 112(2): 221–244, October 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050188.

Serlet:1995:SDR

[Ser95] Laurent Serlet. Some dimension results for super-Brownian motion. Probability Theory and Related Fields, 101(3):371–391, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01200502.

Shao:1992:ATG

[Sha92a] Jun Shao. Asymptotic theory in generalized ilinear models with nuisance scale parameters. Probability Theory and Related Fields, 91(1):25–41, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01194488.

Shao:1992:ERT

[Sha92b] Qi-Man Shao. An Erdős–Révész type law of the iterated logarithm for stationary Gaussian processes. Probability Theory and Related Fields, 94(1):119–133, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01222513.

Shcherbina:2013:CFC

[Shc13] T. Shcherbina. On the correlation functions of the characteristic polynomials of the Hermitian sample covariance matrices. Probability Theory and Related Fields, 156(1–2):449–482, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-012-0433-4.

Sheu:1993:HMC

[She93] Yuan-Chung Sheu. A Hausdorff measure classification of Gpolar sets for the superdiffusions. Probability Theory and Related Fields, 95(4):521–533, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01196732.

Sheffield:2007:GFF

[She07] Scott Sheffield. Gaussian free fields for mathematicians. Probability Theory and Related Fields, 139(3–4):521–541, November 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-006-0050-1.

Shiga:1990:RCM

[Shi90] Tokuzo Shiga. A recurrence criterion for Markov processes of Ornstein–Uhlenbeck type. Probability Theory and Related Fields, 85(4):425–447, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01203163.

Shimura:1992:FPT

[Shi92] Michio Shimura. Flat points of two-dimensional Brownian motion. Probability Theory and Related Fields, 93(2):197–229, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01195229.

Shigekawa:1994:SSB

[Shi94a] Ichiro Shigekawa. Sobolev spaces of Banach-valued functions associated with a Markov process. Probability Theory and Related Fields, 99(3):425–441, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199900.

Shih:1994:CMP

[Shi94b] C. T. Shih. Construction of Markov processes and associated multiplicative functionals from given harmonic measures. Probability Theory and Related Fields, 98(2):185–207, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192513.

Shiga:1997:EDR

[Shi97] Tokuzo Shiga. Exponential decay rate of survival probability in a disastrous random environment. Probability Theory and Related Fields, 108(3):417–439, July 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050115.

Shinoda:2003:NEP

[Shi03] Masato Shinoda. Non-existence of phase transition of oriented percolation on Sierpinski carpet lattices. Probability Theory and Related Fields, 125(3):447–456, March 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-002-0247-x.

Shirikyan:2006:LLN

[Shi06] Armen Shirikyan. Law of large numbers and central limit theorem for randomly forced PDE’s. Probability Theory and Related Fields, 134(2):215–247, February 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-005-0427-6.

Shiraishi:2012:EVR

[Shi12] Daisuke Shiraishi. Exact value of the resistance exponent for four dimensional random walk trace. Probability Theory and Related Fields, 153(1–2):191–232, June 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-

011-0343-x. See correction [CS23].

Silvestri:2017:FRH

[Sil17] Vittoria Silvestri. Fluctuation results for Hastings–Levitov planar growth. Probability Theory and Related Fields, 167(1–2): 417–460, February 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0688-7;
http://link.springer.com/article/10.1007/s00440-015-
0688-7.

Simon:1999:SWS

[Sim99] Thomas Simon. Subordination in the wide sense for Lévy processes. Probability Theory and Related Fields, 115(4):445– 477, November 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050245.

Simon:2011:ACM

[Sim11] Thomas Simon. On the absolute continuity of multidimensional Ornstein–Uhlenbeck processes. Probability Theory and Related Fields, 151(1–2):173–190, October 2011. CODEN PTRFEU.

ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-010-0296-5.

Schiopu-Kratina:1986:WCP

[SK86] I. Schiopu-Kratina. Weak convergence of processes in D(R) and compensation of point processes. Probability Theory and Related Fields, 72(1):99–109, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00343898.

Schiopu-Kratina:1987:WCP

[SK87] I. Schiopu-Kratina. Weak convergence of processes and preservation of predictability. Probability Theory and Related Fields, 76(2):231–241, October 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00319985.

Slud:1991:MWI

[Slu91] Eric Slud. Multiple Wiener-itô integral expansions for levelcrossing-count functionals. Probability Theory and Related Fields, 87(3):349–364, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01312215.

Song:1993:PAD

[Son93] Renming Song. Probabilistic approach to the Dirichlet problem of perturbed stable processes. Probability Theory and Related Fields, 95(3):371–389, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192170.

Song:2004:SBD

[Son04] Renming Song. Sharp bounds on the density, Green function and jumping function of subordinate killed BM. Probability Theory and Related Fields, 128(4):606–628, April 2004. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0316-9.

Soo:2017:MIT

[Soo17] Terry Soo. A monotone isomorphism theorem. Probability Theory and Related Fields, 167(3–4):1117–1136, April 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-016-0700-x.

Sowers:1992:LDI

[Sow92] Richard Sowers. Large deviations for the invariant measure of a reaction-diffusion equation with non-Gaussian perturbations. Probability Theory and Related Fields, 92(3):393–421, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01300562.

Speicher:1990:NEI

[Spe90] Roland Speicher. A new example of ‘independence’ and ‘white noise’. Probability Theory and Related Fields, 84(2):141–159, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01197843.

Srimurthy:2000:EML

[Sri00] Vikram K. Srimurthy. On the equivalence of measures on loop space. Probability Theory and Related Fields, 118(4):522– 546, December 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008753.

Surgailis:1993:SMM

[SRMC93] Donatas Surgailis, Jan Rosinski, V. Mandrekar, and Stamatis Cambanis. Stable mixed moving averages. Probability Theory and Related Fields, 97(4):543–558, ???? 1993. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192963.

Sawyer:1987:REA

[SS87a] Stanley Sawyer and Tim Steger. The rate of escape for anisotropic random walks in a tree. Probability Theory and Related Fields, 76(2):207–230, October 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00319984.

Schal:1987:SPM

[SS87b] Manfred Schäl and William Sudderth. Stationary policies and Markov policies in Borel dynamic programming. Probability Theory and Related Fields, 74(1):91–111, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01845641.

Sidoravicius:2004:QIP

[SS04] Vladas Sidoravicius and Alain-Sol Sznitman. Quenched invariance principles for walks on clusters of percolation or among random conductances. Probability Theory and Related Fields, 129(2):219–244, June 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0336-0.

Schramm:2013:CLC

[SS13] Oded Schramm and Scott Sheffield. A contour line of the continuum Gaussian free field. Probability Theory and Related Fields, 157(1–2):47–80, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0449-9.

Sandier:2015:LGR

[SS15] Etienne Sandier and Sylvia Serfaty. 1D log gases and the renormalized energy: crystallization at vanishing temperature. Probability Theory and Related Fields, 162(3–4):795–846, August 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-014-0585-5.

Sapozhnikov:2018:BMS

[SS18] Artem Sapozhnikov and Daisuke Shiraishi. On Brownian motion, simple paths, and loops. Probability Theory and Related Fields, 172(3–4):615–662, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-017-0817-6.

Sah:2024:IDK

[SS24] Ashwin Sah and Mehtaab Sawhney. The intransitive dice kernel: \frac{∞_{x≥y}−∞_{x≤y}}{4} − − − \frac{3(x−y)(1+xy)}{8}. Probability Theory and Related Fields, 189(3–4):1073–1128, August 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-024-01270-
8.

Sinclair:2017:SMC

[SSSY17] Alistair Sinclair, Piyush Srivastava, Daniel Stefankovic, and Yitong Yin. Spatial mixing and the connective constant: optimal bounds. Probability Theory and Related Fields, 168(1–2):153– 197, June 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

Sly:2022:NSR

[SSZ22] Allan Sly, Nike Sun, and Yumeng Zhang. The number of solutions for random regular NAE-SAT. Probability Theory and Related Fields, 182(1–2):1–109, February 2022. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-021-01029-5.

Sherwood:1988:DSM

[ST88] H. Sherwood and M. D. Taylor. Doubly stochastic measures with hairpin support. Probability Theory and Related Fields, 78 (4):617–626, August 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00353879.

Sabot:2016:IRK

[ST16] Christophe Sabot and Pierre Tarres. Inverting Ray–Knight identity. Probability Theory and Related Fields, 165(3–4): 559–580, August 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0640-x;
http://link.springer.com/article/10.1007/s00440-015-
0640-x.

Stacey:2001:CPF

[Sta01] Alan Stacey. The contact process on finite homogeneous trees. Probability Theory and Related Fields, 121(4):551–576, December 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400100149.

Stannat:2002:LTB

[Sta02] Wilhelm Stannat. Long-time behaviour and regularity properties of transition semigroups of Fleming–Viot processes. Probability Theory and Related Fields, 122(3):431–469, March 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400100166.

Stannat:2004:CGA

[Sta04a] Wilhelm Stannat. On the convergence of genetic algorithmsa variational approach. Probability Theory and Related Fields, 129(1):113–132, May 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0330-y.

Stannat:2004:RCR

[Sta04b] Wilhelm Stannat. A remark on the CLT for a random walk in a random environment. Probability Theory and Related Fields, 130(3):377–387, November 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-004-0359-6.

Stauffer:2017:LFG

[Sta17] Alexandre Stauffer. A Lyapunov function for Glauber dynamics on lattice triangulations. Probability Theory and Related Fields, 169(1–2):469–521, October 2017. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s00440-
016-0735-z.pdf;   http://link.springer.com/article/
10.1007/s00440-016-0735-z.

Steinsaltz:1997:ZWR

[Ste97] David Steinsaltz. Zeno’s walk: A random walk with refinements. Probability Theory and Related Fields, 107(1):99–121, January 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050078.

Stoll:1986:NCL

[Sto86] Andreas Stoll. A nonstandard construction of Lévy Brownian motion. Probability Theory and Related Fields, 71(3):321–334, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01000208.

Strassen:1964:IPL

[Str64] V. Strassen. An invariance principle for the law of the iterated logarithm. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 3(3):211–226, ???? 1964. CODEN ZWVGAA.

ISSN 0044-3719. URL http://link.springer.com/article/

10.1007/BF00534910. See remarks [GK96a].

Strasser:1986:MDA

[Str86a] Helmut Strasser. Martingale difference arrays and stochastic integrals. Probability Theory and Related Fields, 72(1):83–98, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00343897.

Stroyan:1986:PSH

[Str86b] K. D. Stroyan. Previsible sets for hyperfinite filtrations. Probability Theory and Related Fields, 73(2):183–195, September 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00339935.

Strauch:2016:EAP

[Str16] Claudia Strauch. Exact adaptive pointwise drift estimation for multidimensional ergodic diffusions. Probability Theory and Related Fields, 164(1–2):361–400, February 2016. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0614-4.

Sturm:1991:GTR

[Stu91] Karl-Theodor Sturm. Gauge theorems for resolvents with application to Markov processes. Probability Theory and Related Fields, 89(4):387–406, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199785.

Stummer:1993:NEC

[Stu93] Wolfgang Stummer. The Novikov and entropy conditions of multidimensional diffusion processes with singular drift. Probability Theory and Related Fields, 97(4):515–542, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01192962.

Sturm:1995:SEC

[Stu95] K.-T. Sturm. Sharp estimates for capacities and applications to symmetric diffusions. Probability Theory and Related Fields, 103(1):73–89, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199032.

Soner:2012:WSO

[STZ12] H. Mete Soner, Nizar Touzi, and Jianfeng Zhang. Wellposedness of second order backward SDEs. Probability Theory and Related Fields, 153(1–2):149–190, June 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-011-0342-y.

Shiga:1986:SST

[SU86] Tokuzo Shiga and Kohei Uchiyama. Stationary states and their stability of the stepping stone model involving mutation and selection. Probability Theory and Related Fields, 73(1):87–117, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01845994.

Suzuki:1993:HLS

[SU93] Yuki Suzuki and Kôhei Uchiyama. Hydrodynamic limit for a spin system on a multidimensional lattice. Probability Theory and Related Fields, 95(1):47–74, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01197337.

Subag:2023:FES

[Sub23] Eliran Subag. The free energy of spherical pure p-spin models: computation from the TAP approach. Probability Theory and Related Fields, 186(3–4):715–734, August 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01200-0.

Sugita:1994:PHW

[Sug94] Hiroshi Sugita. Properties of holomorphic Wiener functionsskeleton, contraction, and local Taylor expansion. Probability Theory and Related Fields, 100(1):117–130, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01204956.

Sun:1998:AEP

[Sun98] Yeneng Sun. The almost equivalence of pairwise and mutual independence and the duality with exchangeability. Probability Theory and Related Fields, 112(3):425–456, November 1998.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050196.

Suzuki:2024:EIP

[Suz24] Kohei Suzuki. On the ergodicity of interacting particle systems under number rigidity. Probability Theory and Related Fields, 188(1–2):583–623, February 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.

springer.com/article/10.1007/s00440-023-01243-3.

Schonmann:1986:SLD

[SV86a] Roberto H. Schonmann and Maria Eulália Vares. The survival of the large dimensional basic contact process. Probability Theory and Related Fields, 72(3):387–393, June 1986. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00334192.

Sznitman:1986:MPI

[SV86b] A. S. Sznitman and S. R. S. Varadhan. A multidimensional process involving local time. Probability Theory and Related Fields, 71(4):553–579, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00699041.

Salisbury:1999:CEM

[SV99] Thomas S. Salisbury and John Verzani. On the conditioned exit measures of super Brownian motion. Probability Theory and Related Fields, 115(2):237–285, September 1999. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050271.

Song:2003:PTS

[SV03] Renming Song and Zoran Vondracek. Potential theory of subordinate killed Brownian motion in a domain. Probability Theory and Related Fields, 125(4):578–592, April 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
002-0251-1.

Schapira:2017:ETC

[SV17] Bruno Schapira and Daniel Valesin. Extinction time for the contact process on general graphs. Probability Theory and Related Fields, 169(3–4):871–899, December 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-016-0742-0.

Shao:1995:SBP

[SW95] Q.-M. Shao and D. Wang. Small ball probabilities of Gaussian fields. Probability Theory and Related Fields, 102(4):511–517, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01198847.

Swart:2000:CLI

[Swa00] J. M. Swart. Clustering of linearly interacting diffusions and universality of their long-time limit distribution. Probability Theory and Related Fields, 118(4):574–594, December 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/PL00008755.

Swart:2004:UID

[Swa04] Jan M. Swart. Uniqueness for isotropic diffusions with a linear drift. Probability Theory and Related Fields, 128(4):517–524, April 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0315-x.

Swanson:2007:WCS

[Swa07] Jason Swanson. Weak convergence of the scaled median of independent Brownian motions. Probability Theory and Related Fields, 138(1–2):269–304, May 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-006-0024-3.

Swart:2018:SPE

[Swa18] Jan M. Swart. A simple proof of exponential decay of subcritical contact processes. Probability Theory and Related Fields, 170 (1–2):1–9, February 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-016-0741-1.

Swindle:1990:MFL

[Swi90] Glen Swindle. A mean field limit of the contact process with large range. Probability Theory and Related Fields, 85(2): 261–282, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01277984.

Seppalainen:2001:LDP

[SY01] Timo Seppäläinen and J. E. Yukich. Large deviation principles for Euclidean functionals and other nearly additive processes. Probability Theory and Related Fields, 120(3):309–345, July 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/PL00008785.

Schramm:2010:BPS

[SZ10] Oded Schramm and Wang Zhou. Boundary proximity of SLE. Probability Theory and Related Fields, 146(3–4):435–450, March 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-008-0195-1.

Subag:2017:EPC

[SZ17] Eliran Subag and Ofer Zeitouni. The extremal process of critical points of the pure p-spin spherical spin glass model. Probability Theory and Related Fields, 168(3–4):773–820, August 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-016-0724-2.

Sznitman:1986:PCR

[Szn86] A. S. Sznitman. A propagation of chaos result for Burgers’ equation. Probability Theory and Related Fields, 71(4):581–613, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00699042.

Sznitman:1989:LRS

[Szn89] Alain Sol Sznitman. A limiting result for the structure of collisions between many independent diffusions. Probability Theory and Related Fields, 81(3):353–381, April 1989. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00340059.

Sznitman:1993:BAP

[Szn93a] Alain-Sol Sznitman. Brownian asymptotics in a Poissonian environment. Probability Theory and Related Fields, 95(2):155–174, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01192268.

Sznitman:1993:BMP

[Szn93b] Alain-Sol Sznitman. Brownian motion in a Poissonian potential. Probability Theory and Related Fields, 97(4):447–477, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01192959.

Sznitman:1996:BCPa

[Szn96a] Alain-Sol Sznitman. Brownian confinement and pinning in a Poissonian potential. I. Probability Theory and Related Fields, 105(1):1–29, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192069.

Sznitman:1996:BCPb

[Szn96b] Alain-Sol Sznitman. Brownian confinement and pinning in a Poissonian potential. II. Probability Theory and Related Fields, 105(1):31–56, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192070.

Sznitman:1999:SNP

[Szn99] Alain-Sol Sznitman. Slowdown and neutral pockets for a random walk in random environment. Probability Theory and Related Fields, 115(3):287–323, October 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050239.

Sznitman:2002:ECB

[Szn02] Alain-Sol Sznitman. An effective criterion for ballistic behavior of random walks in random environment. Probability Theory and Related Fields, 122(4):509–544, April 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400100177.

Sznitman:2009:RWD

[Szn09] Alain-Sol Sznitman. Random walks on discrete cylinders and random interlacements. Probability Theory and Related Fields, 145(1–2):143–174, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-008-0164-8.

Sznitman:2011:LBC

[Szn11] Alain-Sol Sznitman. A lower bound on the critical parameter of interlacement percolation in high dimension. Probability Theory and Related Fields, 150(3–4):575–611, August 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0284-9.

Sznitman:2017:DRW

[Szn17] Alain-Sol Sznitman. Disconnection, random walks, and random interlacements. Probability Theory and Related Fields, 167(1– 2):1–44, February 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0676-y;
http://link.springer.com/article/10.1007/s00440-015-
0676-y.

Szulga:1992:LDS

[Szu92] Jerzy Szulga. Limit distributions of U-statistics resampled by symmetric stable laws. Probability Theory and Related Fields, 94(1):83–90, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01222511.

Tempelman:1996:MTM

[TA96] Arkady A. Tempelman and Michael G. Akritas. Model testing for multivariate censored data. Probability Theory and Related Fields, 106(3):351–369, November 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050068.

Takanobu:1987:ESS

[Tak87] Satoshi Takanobu. On the existence of solutions of stochastic differential equations with singular drifts. Probability Theory and Related Fields, 74(2):295–315, ???? 1987. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00569995.

Takacs:1998:BRW

[Tak98] Christiane Takacs. Biased random walks on directed trees. Probability Theory and Related Fields, 111(1):123–139, June 1998.

CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050164.

Talagrand:1988:LRS

[Tal88a] M. Talagrand. On liftings and the regularization of stochastic processes. Probability Theory and Related Fields, 78(1):127–134, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00718040.

Talagrand:1988:DCS

[Tal88b] Michel Talagrand. Donsker classes of sets. Probability Theory and Related Fields, 78(2):169–191, June 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00322017.

Talagrand:1998:MPT

[Tal98a] Michel Talagrand. Multiple points of trajectories of multiparameter fractional Brownian motion. Probability Theory and Related Fields, 112(4):545–563, December 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400050200.

Talagrand:1998:RRH

[Tal98b] Michel Talagrand. Rigorous results for the Hopfield model with many patterns. Probability Theory and Related Fields, 110(2): 177–275, February 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050148.

Talagrand:1998:SKM

[Tal98c] Michel Talagrand. The Sherrington–Kirkpatrick model: a challenge for mathematicians. Probability Theory and Related Fields, 110(2):109–176, February 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400050147.

Talagrand:2000:MLS

[Tal00a] Michel Talagrand. Multiple levels of symmetry breaking. Probability Theory and Related Fields, 117(4):449–466, August 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400000070.

Talagrand:2000:RLT

[Tal00b] Michel Talagrand. Rigorous low-temperature results for the mean field p-spins interaction model. Probability Theory and Related Fields, 117(3):303–360, July 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050009.

Talagrand:2001:HTC

[Tal01a] Michel Talagrand. The high temperature case for the random K-sat problem. Probability Theory and Related Fields, 119(2): 187–212, February 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008758.

Talagrand:2001:HMC

[Tal01b] Michel Talagrand. On the Hopfield model at the critical temperature. Probability Theory and Related Fields, 121(2): 237–268, October 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008804.

Talagrand:2005:EWS

[Tal05] Michel Talagrand. Expected wasted space of optimal simple rectangle packing. Probability Theory and Related Fields, 131(2):145–153, February 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0299-6.

Talagrand:2006:FES

[Tal06] Michel Talagrand. Free energy of the spherical mean field model. Probability Theory and Related Fields, 134(3):339–382, March 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-005-0433-8.

Talagrand:2009:GFC

[Tal09] Michel Talagrand. A general form of certain mean field models for spin glasses. Probability Theory and Related Fields, 143(1– 2):97–111, January 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0121-y.

Talagrand:2010:CPS

[Tal10] Michel Talagrand. Construction of pure states in mean field models for spin glasses. Probability Theory and Related Fields, 148(3–4):601–643, November 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-009-0242-6.

Talagrand:2016:MFS

[Tal16] Michel Talagrand. A mean-field spin glass model based on diluted V -statistics. Probability Theory and Related Fields, 165 (1–2):401–445, June 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-015-0634-8.

Taniguchi:1986:BET

[Tan86] Masanobu Taniguchi. Berry–Esseen theorems for quadratic forms of Gaussian stationary processes. Probability Theory and Related Fields, 72(2):185–194, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00699102.

Tanemura:1996:SIM

[Tan96] Hideki Tanemura. A system of infinitely many mutually reflecting Brownian balls in R^{d}. Probability Theory and Related Fields, 104(3):399–426, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01213687.

Tanemura:1997:UDF

[Tan97] Hideki Tanemura. Uniqueness of Dirichlet forms associated with systems of infinitely many Brownian balls in r^{d}. Probability Theory and Related Fields, 109(2):275–299, October 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050133.

Taniguchi:1999:SOI

[Tan99] S. Taniguchi. Stochastic oscillatory integrals with quadratic phase function and Jacobi equations. Probability Theory and Related Fields, 114(3):291–308, June 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050227.

Tao:2013:ADS

[Tao13a] Terence Tao. The asymptotic distribution of a single eigenvalue gap of a Wigner matrix. Probability Theory and Related Fields, 157(1–2):81–106, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0450-3.

Tao:2013:EOS

[Tao13b] Terence Tao. Erratum to: Outliers in the spectrum of iid matrices with bounded rank perturbations. Probability Theory and Related Fields, 157(1–2):511–514, October 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-013-0513-0.pdf. See [Tao13c].

Tao:2013:OSI

[Tao13c] Terence Tao. Outliers in the spectrum of iid matrices with bounded rank perturbations. Probability Theory and Related Fields, 155(1–2):231–263, February 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-

011-0397-9. See erratum [Tao13b].

Taylor:1989:DMS

[Tay89] J. C. Taylor. Do minimal solutions of heat equations characterize diffusions? Probability Theory and Related Fields, 83 (3):321–330, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00964368.

Tearne:2008:CAO

[Tea08] Oliver M. Tearne. Collapse of attractors for ODEs under small random perturbations. Probability Theory and Related Fields, 141(1–2):1–18, May 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0051-0.

Teicher:1988:DMC

[Tei88] H. Teicher. Distribution and moment convergence of martingales. Probability Theory and Related Fields, 79(2):303– 316, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00320924.

Teixeira:2011:SFV

[Tei11] Augusto Teixeira. On the size of a finite vacant cluster of random interlacements with small intensity. Probability Theory and Related Fields, 150(3–4):529–574, August 2011. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
010-0283-x.

Teixeira:2016:PLI

[Tei16] Augusto Teixeira. Percolation and local isoperimetric inequalities. Probability Theory and Related Fields, 165(3–4): 963–984, August 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0653-5;
http://link.springer.com/article/10.1007/s00440-015-
0653-5.

Telcs:1989:RWG

[Tel89] A. Telcs. Random walks on graphs, electric networks and fractals. Probability Theory and Related Fields, 82(3):435–449, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00339997.

Telcs:1990:SGF

[Tel90] András Telcs. Spectra of graphs and fractal dimensions. I. Probability Theory and Related Fields, 85(4):489–497, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01203166.

Terhesiu:2016:MRI

[Ter16] Dalia Terhesiu. Mixing rates for intermittent maps of high exponent. Probability Theory and Related Fields, 166(3–4):1025– 1060, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0690-0;
http://link.springer.com/article/10.1007/s00440-015-
0690-0.

Thang:1991:CVR

[Tha91] Dang Hung Thang. On the convergence of vector random measures. Probability Theory and Related Fields, 88(1):1–16, ????

1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01193580.

Thalmaier:1996:BMF

[Tha96] Anton Thalmaier. Brownian motion and the formation of singularities in the heat flow for harmonic maps. Probability Theory and Related Fields, 105(3):335–367, ???? 1996. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192212.

Thieullen:1991:CSN

[Thi91] Michèle Thieullen. Calcul stochastique non adapté pour des processus à deux paramètres: formules de changement de variables de type Stratonovitch et de type Skorohod. (French) [Stochastic calculus unsuitable for processes with two parameters: the change-of-variables formulas of Stratonovitch type and Skorokhod type]. Probability Theory and Related Fields, 89 (4):457–485, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01199789.

Thorisson:1988:FIT

[Tho88] Hermann Thorisson. Future independent times and Markov chains. Probability Theory and Related Fields, 78(1):143–148, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00718042.

Thorisson:1994:SCC

[Tho94] Hermann Thorisson. Shift-coupling in continuous time. Probability Theory and Related Fields, 99(4):477–483, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01206228.

Thomas:2001:PLC

[Tho01] Erik G. F. Thomas. Projective limits of complex measures and martingale convergence. Probability Theory and Related Fields, 119(4):579–588, April 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008772.

Thomine:2014:GCL

[Tho14] Damien Thomine. A generalized central limit theorem in infinite ergodic theory. Probability Theory and Related Fields, 158(3– 4):597–636, April 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0491-2.

Thrum:1987:RAS

[Thr87] Rolf Thrum. A remark on almost sure convergence of weighted sums. Probability Theory and Related Fields, 75(3):425–430, July 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00318709.

Thu:1986:AAM

[Thu86] Nguyen Van Thu. An alternative approach to multiply selfdecomposable probability measures on Banach spaces. Probability Theory and Related Fields, 72(1):35–54, April 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00343895.

Tsaknakis:1986:RPI

[TKPK86] Haralampos Tsaknakis, Dimitri Kazakos, and P. Papantoni-Kazakos. Robust prediction and interpolation for vector stationary processes. Probability Theory and Related Fields, 72 (4):589–602, July 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00344722.

Toth:1986:PRW

[Tót86] Bálint Tóth. Persistent random walks in random environment. Probability Theory and Related Fields, 71(4):615–625, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00699043.

Toubol:1998:HTR

[Tou98] Alain Toubol. High temperature regime for a multidimensional Sherrington–Kirkpatrick model of spin glass. Probability Theory and Related Fields, 110(4):497–534, May 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050157.

Trautner:1990:NPK

[Tra90] Rolf Trautner. A new proof of the Komlós-Révész-theorem. Probability Theory and Related Fields, 84(3):281–287, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01197885.

Trevisan:2015:LFD

[Tre15] Dario Trevisan. Lagrangian flows driven by BV fields in Wiener spaces. Probability Theory and Related Fields, 163(1–2):123– 147, October 2015. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-014-0589-1.

Tribe:1991:CCC

[Tri91] Roger Tribe. The connected components of the closed support of super Brownian motion. Probability Theory and Related Fields, 89(1):75–87, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01225826.

Tribe:1995:LTB

[Tri95] Roger Tribe. Large time behavior of interface solutions to the heat equation with Fisher–Wright white noise. Probability Theory and Related Fields, 102(3):289–311, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192463.

Trouve:1995:ABS

[Tro95] Alain Trouvé. Asymptotical behaviour of several interacting annealing processes. Probability Theory and Related Fields, 102 (1):123–143, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01295225.

Tropp:2012:CPF

[Tro12] Joel A. Tropp. A comparison principle for functions of a uniformly random subspace. Probability Theory and Related Fields, 153(3–4):759–769, August 2012. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s00440-
011-0360-9.pdf.

Troscheit:2021:QEB

[Tro21] Sascha Troscheit. On quasisymmetric embeddings of the Brownian map and continuum trees. Probability Theory and Related Fields, 179(3–4):1023–1046, April 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-020-01024-
2.

Trutnau:2003:SDR

[Tru03] Gerald Trutnau. Skorokhod decomposition of reflected diffusions on bounded Lipschitz domains with singular non-reflection part. Probability Theory and Related Fields, 127(4):455–495, December 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-003-0296-9.

Tsai:2016:IDS

[Tsa16] Li-Cheng Tsai. Infinite dimensional stochastic differential equations for Dyson’s model. Probability Theory and Related Fields, 166(3–4):801–850, December 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s00440-015-0672-2;   http://link.springer.com/
article/10.1007/s00440-015-0672-2.

Terrin:1991:CDS

[TT91] Norma Terrin and Murad S. Taqqu. Convergence in distribution of sums of bivariate Appell polynomials with long-range dependence. Probability Theory and Related Fields, 90(1):57–81, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01321134.

Tindel:2003:SEE

[TTV03] S. Tindel, C. A. Tudor, and F. Viens. Stochastic evolution equations with fractional Brownian motion. Probability Theory and Related Fields, 127(2):186–204, October 2003. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0282-2.

Taylor:1993:EUS

[TW93] L. M. Taylor and R. J. Williams. Existence and uniqueness of semimartingale reflecting Brownian motions in an orthant. Probability Theory and Related Fields, 96(3):283–317, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01292674.

Toth:1998:TSR

[TW98] Bálint Tóth and Wendelin Werner. The true self-repelling motion. Probability Theory and Related Fields, 111(3):375–452, July 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s004400050172.

Tracy:2001:DLL

[TW01] Craig A. Tracy and Harold Widom. On the distributions of the lengths of the longest monotone subsequences in random words. Probability Theory and Related Fields, 119(3):350–380, March 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008763.

Tykesson:2012:PVS

[TW12] Johan Tykesson and David Windisch. Percolation in the vacant set of Poisson cylinders. Probability Theory and Related Fields, 154(1–2):165–191, October 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0366-3.

Tsatsoulis:2020:ELM

[TW20] Pavlos Tsatsoulis and Hendrik Weber. Exponential loss of memory for the 2-dimensional Allen–Cahn equation with small noise. Probability Theory and Related Fields, 177(1–2):257–322, June 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-019-00945-x;   https://link.springer.
com/content/pdf/10.1007/s00440-019-00945-x.pdf.

Tanemura:2003:LTF

[TY03] Hideki Tanemura and Nobuo Yoshida. Localization transition of d-friendly walkers. Probability Theory and Related Fields, 125 (4):593–608, April 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-002-0253-z.

Tikhomirov:2023:SPL

[TY23] Konstantin Tikhomirov and Pierre Youssef. Sharp Poincaré and log-Sobolev inequalities for the switch chain on regular bipartite graphs. Probability Theory and Related Fields, 185(1–2): 89–184, February 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-022-01172-7.

Thieullen:1997:SSC

[TZ97] M. Thieullen and J. C. Zambrini. Symmetries in the stochastic calculus of variations. Probability Theory and Related Fields, 107 (3):401–427, March 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050091.

Tchoumatchenko:2001:AFT

[TZ01] Konstantin Tchoumatchenko and Sergei Zuyev. Aggregate and fractal tessellations. Probability Theory and Related Fields, 121 (2):198–218, October 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008802.

Tzvetkov:2010:CGM

[Tzv10] N. Tzvetkov. Construction of a Gibbs measure associated to the periodic Benjamin–Ono equation. Probability Theory and Related Fields, 146(3–4):481–514, March 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-008-0197-z.

Uzun:2003:LBB

[UA03] Hasan B. Uzun and Kenneth S. Alexander. Lower bounds for boundary roughness for droplets in Bernoulli percolation. Probability Theory and Related Fields, 127(1):62–88, September 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-003-0276-0.

Uchiyama:1988:FMS

[Uch88] Kohei Uchiyama. Fluctuations in a Markovian system of pairwise interacting particles. Probability Theory and Related Fields, 79(2):289–302, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00320923.

Uchiyama:1989:TPP

[Uch89] Kôhei Uchiyama. A tagged particle process in the Boltzmann– Grad limit for the Broadwell modell. Probability Theory and Related Fields, 82(3):419–433, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00339996.

Uchiyama:1994:SLI

[Uch94] Kôhei Uchiyama. Scaling limits of interacting diffusions with arbitrary initial distributions. Probability Theory and Related Fields, 99(1):97–110, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01199591.

Ueki:1989:SAP

[Uek89] Naomasa Ueki. A stochastic approach to the Poincaré-Hopf theorem. Probability Theory and Related Fields, 82(2):271–293, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00354764.

Ueltschi:2002:SAW

[Uel02] Daniel Ueltschi. A self-avoiding walk with attractive interactions. Probability Theory and Related Fields, 124(2):189– 203, October 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400200209.

Umanita:2006:CDQ

[Uma06] Veronica Umanità. Classification and decomposition of Quantum Markov Semigroups. Probability Theory and Related Fields, 134(4):603–623, April 2006. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-005-0450-7.

Urbanski:1990:PDH

[Urb90] M. Urbański. The probability distribution and Hausdorff dimension of self-affine functions. Probability Theory and Related Fields, 84(3):377–391, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01197891.

Ustunel:1988:IFA

[Ust88] A. S. Ustunel. The Itô formula for anticipative processes with nonmonotonous time scale via the Malliavin calculus. Probability Theory and Related Fields, 79(2):249–269, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00320921.

Ustunel:1992:TWM

[ÜZ92] Ali Süleyman Üstünel and Moshe Zakai. Transformation of Wiener measure under anticipative flows. Probability Theory and Related Fields, 93(1):91–136, ???? 1992. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01195390.

Ustunel:1993:ADT

[ÜZ93] A. S. Üstünel and M. Zakai. Applications of the degree theorem to absolute continuity on Wiener space. Probability Theory and Related Fields, 95(4):509–520, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01196731.

Ustunel:1994:CWF

[ÜZ94a] A. S. Üstünel and M. Zakai. The composition of Wiener functionals with non absolutely continuous shifts. Probability Theory and Related Fields, 98(2):163–184, ???? 1994. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192512.

Ustunel:1994:TWM

[ÜZ94b] A. S. Üstünel and M. Zakai. Transformation of the Wiener measure under non-invertible shifts. Probability Theory and Related Fields, 99(4):485–500, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01206229.

Ustunel:1995:RRW

[ÜZ95] A. S. Üstünel and M. Zakai. Random rotations of the Wiener path. Probability Theory and Related Fields, 103(3):409–429, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01195481.

Ustunel:1996:MIW

[ÜZ96] Ali Süleyman Üstünel and Moshe Zakai. Measures induced on Wiener space by monotone shifts. Probability Theory and Related Fields, 105(4):545–563, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01191913.

Ustunel:1997:DTW

[ÜZ97] A. Süleyman Üstünel and Moshe Zakai. Degree theory on Wiener space. Probability Theory and Related Fields, 108(2): 259–279, June 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050109.

Ustunel:2007:SCI

[ÜZ07] Ali Süleyman Üstünel and Moshe Zakai. Sufficient conditions for the invertibility of adapted perturbations of identity on the Wiener space. Probability Theory and Related Fields, 139(1–2): 207–234, September 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0044-z.

Vuolle-Apiala:1992:ETR

[VA92] Juha Vuolle-Apiala. Excursion theory for rotation invariant Markov processes. Probability Theory and Related Fields, 93 (2):153–158, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01195226.

Vallois:1991:ETR

[Val91] P. Vallois. Une extension des théorèmes de Ray et Knight sur les temps locaux Browniens. (French) [An extension of the theorems of Ray and Knight on local Brownian times]. Probability Theory and Related Fields, 88(4):445–482, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192552.

Vargas:2007:SLM

[Var07] Vincent Vargas. Strong localization and macroscopic atoms for directed polymers. Probability Theory and Related Fields, 138 (3–4):391–410, July 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0030-5.

vandenBerg:1994:HCB

[vdB94] M. van den Berg. Heat content and Brownian motion for some regions with a fractal boundary. Probability Theory and Related Fields, 100(4):439–456, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01268989.

vandenBerg:2000:HEA

[vdB00] M. van den Berg. Heat equation on the arithmetic von Koch snowflake. Probability Theory and Related Fields, 118(1):17–36, September 2000. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/PL00008740;   http://link.springer.
com/article/10.1007/s004400000068.

vandenBerg:1994:AGF

[vdBB94] Michiel van den Berg and Erwin Bolthausen. Asymptotics of the generating function for the volume of the Wiener sausage. Probability Theory and Related Fields, 99(3):389–397, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01199898.

vandenBerg:2004:AVC

[vdBB04] M. van den Berg and E. Bolthausen. Area versus capacity and solidification in the crushed ice model. Probability Theory and Related Fields, 130(1):69–108, September 2004. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
003-0332-9.

vandenBerg:2005:BSA

[vdBBdH05] M. van den Berg, E. Bolthausen, and F. den Hollander. Brownian survival among Poissonian traps with random shapes at critical intensity. Probability Theory and Related Fields, 132 (2):163–202, June 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0393-4.

vandeBrug:2016:RWL

[vdBCL16] Tim van de Brug, Federico Camia, and Marcin Lis. Random walk loop soups and conformal loop ensembles. Probability Theory and Related Fields, 166(1–2):553–584, October 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-015-0666-0.pdf.

vandenBerg:2013:BTI

[vdBG13] J. van den Berg and A. Gandolfi. BK-type inequalities and generalized random-cluster representations. Probability Theory and Related Fields, 157(1–2):157–181, October 2013. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
012-0452-1.

vandenBerg:2012:BIR

[vdBJ12] J. van den Berg and J. Jonasson. A BK inequality for randomly drawn subsets of fixed size. Probability Theory and Related Fields, 154(3–4):835–844, December 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-011-0386-z.

vandenBerg:2021:NCP

[vdBN21] Jacob van den Berg and Pierre Nolin. Near-critical 2D percolation with heavy-tailed impurities, forest fires and frozen percolation. Probability Theory and Related Fields, 181(1-3): 211–290, November 2021. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-020-01022-4.

vandenBerg:1990:HFB

[vdBS90] M. van den Berg and S. Srisatkunarajah. Heat flow and Brownian motion for a region in R^{2} with a polygonal boundary. Probability Theory and Related Fields, 86(1):41–52, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01207512.

vandenBerg:1991:EEW

[vdBT91] M. van den Berg and B. Tóth. Exponential estimates for the Wiener sausage. Probability Theory and Related Fields, 88(2): 249–259, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01212561.

vandeGeer:2013:BON

[vdGL13] Sara van de Geer and Johannes Lederer. The Bernstein–Orlicz norm and deviation inequalities. Probability Theory and Related Fields, 157(1–2):225–250, October 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-012-0455-y.

vanderHofstad:2001:LEA

[vdH01] Remco van der Hofstad. The lace expansion approach to ballistic behaviour for one-dimensional weakly self-avoiding walks. Probability Theory and Related Fields, 119(3):311–349, March 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/PL00008762.

vanderHofstad:2003:WIL

[vdHdHK03] Remco van der Hofstad, Frank den Hollander, and Wolfgang König. Weak interaction limits for one-dimensional random polymers. Probability Theory and Related Fields, 125(4): 483–521, April 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-002-0248-9.

vanderHofstad:1998:NIA

[vdHdHS98] Remco van der Hofstad, Frank den Hollander, and Gordon Slade. A new inductive approach to the lace expansion for self-avoiding walks. Probability Theory and Related Fields, 111 (2):253–286, June 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050168.

vanderHofstad:2007:SPC

[vdHdHS07] Remco van der Hofstad, Frank den Hollander, and Gordon Slade. The survival probability for critical spread-out oriented percolation above 4 + 1 dimensions. I. Induction. Probability Theory and Related Fields, 138(3–4):363–389, July 2007. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
006-0028-z.

vanderHofstad:2010:MER

[vdHH10] Remco van der Hofstad and Mark Holmes. Monotonicity for excited random walk in high dimensions. Probability Theory and Related Fields, 147(1–2):333–348, May 2010. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00440-009-0215-9.pdf.

vanderHofstad:2020:RWB

[vdHHN20] Remco van der Hofstad, Tim Hulshof, and Jan Nagel. Random walk on barely supercritical branching random walk. Probability Theory and Related Fields, 177(1–2):1–53, June 2020. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
019-00942-0.

vanderHofstad:2010:RSH

[vdHL10] Remco van der Hofstad and Malwina J. Luczak. Random subgraphs of the 2D Hamming graph: the supercritical phase. Probability Theory and Related Fields, 147(1–2):1–41, May 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-009-0200-3.

vanderHofstad:2002:GIA

[vdHS02] Remco van der Hofstad and Gordon Slade. A generalised inductive approach to the lace expansion. Probability Theory and Related Fields, 122(3):389–430, March 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400100175.

vanderHofstad:2005:CPS

[vdHS05] Remco van der Hofstad and Akira Sakai. Critical points for spread-out self-avoiding walk, percolation and the contact process above the upper critical dimensions. Probability Theory and Related Fields, 132(3):438–470, July 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-004-0405-4.

vanEs:1988:ESP

[vEH88] A. J. van Es and R. Helmers. Elementary symmetric polynomials of increasing order. Probability Theory and Related Fields, 80 (1):21–35, December 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00348750.

Velenik:2004:ERI

[Vel04] Y. Velenik. Entropic repulsion of an interface in an external field. Probability Theory and Related Fields, 129(1):83–112, May 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-003-0328-5.

Velenik:2009:WGF

[Vel09] Yvan Velenik. Wetting of gradient fields: pathwise estimates. Probability Theory and Related Fields, 143(3–4):379–399, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-007-0130-x.

Vervaat:1990:TFI

[Ver90] Wim Vervaat. Transformations in functional iterated logarithm laws and regular variation. Probability Theory and Related Fields, 87(1):121–128, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01217749.

Verzani:1997:CPP

[Ver97] John Verzani. Cone paths for the planar Brownian snake. Probability Theory and Related Fields, 107(4):517–526, April 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050096.

Veretennikov:2003:LDA

[Ver03] A. Yu. Veretennikov. On large deviations for approximations of SDEs. Probability Theory and Related Fields, 125(1): 135–152, January 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-002-0235-1.

Vershynin:2011:SNP

[Ver11] Roman Vershynin. Spectral norm of products of random and deterministic matrices. Probability Theory and Related Fields, 150(3–4):471–509, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0281-z.

vanHandel:2009:ONF

[vH09] Ramon van Handel. Observability and nonlinear filtering. Probability Theory and Related Fields, 145(1–2):35–74, September 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-008-0161-y.

vanHandel:2011:MPM

[vH11] Ramon van Handel. On the minimal penalty for Markov order estimation. Probability Theory and Related Fields, 150(3–4): 709–738, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0290-y.

vanHandel:2013:UGC

[vH13] Ramon van Handel. The universal Glivenko–Cantelli property. Probability Theory and Related Fields, 155(3–4):911–934, April 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-012-0416-5.

vanHandel:2018:BEG

[vH18] Ramon van Handel. The Borell–Ehrhard game. Probability Theory and Related Fields, 170(3–4):555–585, April 2018. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
017-0762-4.

Viennet:1997:IAR

[Vie97] Gabrielle Viennet. Inequalities for absolutely regular sequences: application to density estimation. Probability Theory and Related Fields, 107(4):467–492, April 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050094.

Virag:2002:FGR

[Vir02] Bálint Virág. Fast graphs for the random walker. Probability Theory and Related Fields, 124(1):50–72, September 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200200.

Virag:2003:BB

[Vir03] Bálint Virág. Brownian beads. Probability Theory and Related Fields, 127(3):367–387, November 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0289-8.

Vondracek:1992:CBM

[Von92] Zoran Vondracek. A characterization of h-Brownian motion by its exit distributions. Probability Theory and Related Fields, 92(1):41–50, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01205235.

Vondracek:1995:SMP

[Von95] Zoran Vondracek. Some Markov processes with Brownian exit distributions. Probability Theory and Related Fields, 101(3): 393–407, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01200503.

Viklund:2014:CTS

[VR14] Fredrik Johansson Viklund and Steffen Rohde. On the continuity of SLE_{κ} in κ. Probability Theory and Related Fields, 159(3– 4):413–433, August 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0506-z.

vonVintschger:1989:EIM

[vV89] R. von Vintschger. The existence of invariant measures for C[0, 1]-valued diffusions. Probability Theory and Related Fields, 82(2):307–313, ???? 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00354766.

vonWeizsacker:1997:STM

[vW97] Heinrich von Weizsäcker. Sudakov’s typical marginals, random linear functionals and a conditional central limit theorem. Probability Theory and Related Fields, 107(3):313–324, March 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050087.

Vatutin:2009:LPR

[VW09] Vladimir A. Vatutin and Vitali Wachtel. Local probabilities for random walks conditioned to stay positive. Probability Theory and Related Fields, 143(1–2):177–217, January 2009. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0124-8.

Vidmar:2024:SLR

[VW24] Matija Vidmar and Jon Warren. Stationary local random countable sets over the Wiener noise. Probability Theory and Related Fields, 188(3–4):1063–1129, April 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https:/

/link.springer.com/article/10.1007/s00440-023-01227-
3.

Wachtel:2012:LLT

[Wac12] Vitali Wachtel. Local limit theorem for the maximum of asymptotically stable random walks. Probability Theory and Related Fields, 152(3–4):407–424, April 2012. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0326-3.

Wang:1994:ACM

[Wan94a] Feng-Yu Wang. Application of coupling methods to the Neumann eigenvalue problem. Probability Theory and Related Fields, 98(3):299–306, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192256.

Wang:1994:CFS

[Wan94b] Kongming Wang. Critical fluctuations of sums of weakly dependent random vectors. Probability Theory and Related Fields, 98 (2):229–243, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192515.

Wang:1995:EFD

[Wan95] Feng-Yu Wang. Estimates of the first Dirichlet eigenvalue by using diffusion processes. Probability Theory and Related Fields, 101(3):363–369, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01200501.

Wang:1997:LSI

[Wan97a] Feng-Yu Wang. Logarithmic Sobolev inequalities on noncompact Riemannian manifolds. Probability Theory and Related Fields, 109(3):417–424, November 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050137.

Wang:1997:ELS

[Wan97b] Feng-Yu Wang. On estimation of the logarithmic Sobolev constant and gradient estimates of heat semigroups. Probability Theory and Related Fields, 108(1):87–101, May 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050102.

Wang:1997:SCC

[Wan97c] H. Wang. State classification for a class of measure-valued branching diffusions in a Brownian medium. Probability Theory and Related Fields, 109(1):39–55, September 1997. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050124.

Wang:2001:MFB

[Wan01] Wei-Min Wang. Mean field bounds on Lyapunov exponents in Z^{d} at the critical energy. Probability Theory and Related Fields, 119(4):453–474, April 2001. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/PL00008767.

Wang:2003:WCF

[Wan03] Wensheng Wang. Weak convergence to fractional Brownian motion in Brownian scenery. Probability Theory and Related Fields, 126(2):203–220, June 2003. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-002-0249-8.

Wang:2018:EIP

[Wan18a] Feng-Yu Wang. Estimates for invariant probability measures of degenerate SPDEs with singular and path-dependent drifts. Probability Theory and Related Fields, 172(3–4):1181–1214, December 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-017-0827-4.

Wang:2018:PHI

[Wan18b] Zhenan Wang. A probabilistic Harnack inequality and strict positivity of stochastic partial differential equations. Probability Theory and Related Fields, 171(3–4):653–684, August 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-017-0789-6.

Warren:2005:SFA

[War05] Jon Warren. A stochastic flow arising in the study of local times. Probability Theory and Related Fields, 133(4):559–572, December 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-005-0449-0.

Watkins:1987:FCL

[Wat87] Joseph C. Watkins. Functional central limit theorems and their associated large deviation principles for products of random matrices. Probability Theory and Related Fields, 76(2): 133–166, October 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00319982.

Watanabe:1988:AFP

[Wat88] Hisao Watanabe. Averaging and fluctuations for parabolic equations with rapidly oscillating random coefficients. Probability Theory and Related Fields, 77(3):359–378, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00319294.

Watanabe:1993:FOS

[Wat93] Shinzo Watanabe. Fractional order Sobolev spaces on Wiener space. Probability Theory and Related Fields, 95(2):175–198, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF01192269.

Watanabe:1996:SFB

[Wat96] Toshiro Watanabe. Sample function behavior of increasing processes of class L. Probability Theory and Related Fields, 104 (3):349–374, ???? 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01213685.

Watanabe:2000:ACS

[Wat00] Toshiro Watanabe. Absolute continuity of some semiselfdecomposable distributions and self-similar measures. Probability Theory and Related Fields, 117(3):387–405, July 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050011.

Watanabe:2008:CED

[Wat08] Toshiro Watanabe. Convolution equivalence and distributions of random sums. Probability Theory and Related Fields, 142(3–4): 367–397, November 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-007-0109-7.

Weber:1989:SGP

[Web89] Michel Weber. The supremum of Gaussian processes with a constant variance. Probability Theory and Related Fields, 81 (4):585–591, May 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00367305.

Weba:1991:QSS

[Web91] Michael Weba. Quadrature of smooth stochastic processes. Probability Theory and Related Fields, 87(3):333–347, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01312214.

Wee:1988:LFP

[Wee88] In-Suk Wee. Lower functions for processes with stationary independent increments. Probability Theory and Related Fields, 77(4):551–566, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00959617.

Wee:1990:LFA

[Wee90] In-Suk Wee. Lower functions for asymmetric Lévy processes. Probability Theory and Related Fields, 85(4):469–488, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01203165.

Wee:1992:LIL

[Wee92] In-Suk Wee. The law of the iterated logarithm for local time of a Lévy process. Probability Theory and Related Fields, 93 (3):359–376, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193056.

Weiner:1986:LIL

[Wei86a] Daniel Charles Weiner. A law of the iterated logarithm for distributions in the generalized domain of attraction of a nondegenerate Gaussian law. Probability Theory and Related Fields, 72(3):337–357, June 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00334190.

Weinryb:1986:EAP

[Wei86b] Sophie Weinryb. Étude asymptotique par des mesures de R^{3} de saucisses de Wiener localisées. (French) [Asymptotic study by measurements of R^{3} of localized Wiener sausages]. Probability Theory and Related Fields, 73(1):135–148, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01845997.

Weissman:1988:AJD

[Wei88] Ishay Weissman. On the asymptotic joint distribution of an unbounded number of sample extremes. Probability Theory and Related Fields, 78(1):39–50, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00718033.

Weisz:1990:MHS

[Wei90] F. Weisz. Martingale Hardy spaces for 0 < p ≤ 1. Probability Theory and Related Fields, 84(3):361–376, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01197890.

Wei:1993:MTN

[Wei93] C. Z. Wei. Martingale transforms with non-atomic limits and stochastic approximation. Probability Theory and Related Fields, 95(1):103–114, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01197340.

Wentzell:1996:AEL

[Wen96] A. D. Wentzell. Asymptotic expansions in limit theorems for stochastic processes. I. Probability Theory and Related Fields, 106(3):331–350, November 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050067.

Wentzell:1999:AEL

[Wen99] Alexander D. Wentzell. Asymptotic expansions in limit theorems for stochastic processes. II. Probability Theory and Related Fields, 113(2):255–271, February 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:/

/link.springer.com/article/10.1007/s004400050207.

Wentzell:2004:AEL

[Wen04] Alexander D. Wentzell. Asymptotic expansions in limit theorems for stochastic processes.- III. Probability Theory and Related Fields, 128(1):63–81, January 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-003-0294-y.

Werner:1994:FCC

[Wer94a] Wendelin Werner. Sur la forme des composantes connexes du complémentaire de la courbe brownienne plane. (French) [On the shape of the connected components of the complement of planar Brownian motion]. Probability Theory and Related Fields, 98(3):307–337, ???? 1994. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192257.

Werner:1994:PAD

[Wer94b] Wendelin Werner. Sur les points autour desquels le mouvement brownien plan tourne beaucoup. (French) [On the points about which planar Brownian motion turns a lot]. Probability Theory and Related Fields, 99(1):111–144, ???? 1994. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01199592.

Werner:1997:ABD

[Wer97] Wendelin Werner. Asymptotic behaviour of disconnection and non-intersection exponents. Probability Theory and Related Fields, 108(1):131–152, May 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s004400050104.

Wang:2007:ACG

[WHZY07] Yanhua Wang, Shuyuan He, Lixing Zhu, and Kam C. Yuen. Asymptotics for a censored generalized linear model with unknown link function. Probability Theory and Related Fields, 138 (1–2):235–267, May 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-006-0022-5.

Wieacker:1986:IRH

[Wie86] John André Wieacker. Intersections of random hypersurfaces and visibility. Probability Theory and Related Fields, 71(3): 405–433, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01000214.

Wieand:2002:EDR

[Wie02] K. Wieand. Eigenvalue distributions of random unitary matrices. Probability Theory and Related Fields, 123(2):202–224, June 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400100186.

Williams:1987:RBM

[Wil87] R. J. Williams. Reflected Brownian motion with skew symmetric data in a polyhedral domain. Probability Theory and Related Fields, 75(4):459–485, August 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00320328.

Willekens:1988:SCS

[Wil88] Eric Willekens. The structure of the class of subexponential distributions. Probability Theory and Related Fields, 77(4): 567–581, ???? 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00959618.

Wilson:1997:RRW

[Wil97] David Bruce Wilson. Random random walks on Z_{2}^{d}. Probability Theory and Related Fields, 108(4):441–457, August 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050116.

Winnicki:1991:EVB

[Win91] J. Winnicki. Estimation of the variances in the branching process with immigration. Probability Theory and Related Fields, 88(1):77–106, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01193583.

Wittmann:1986:NDM

[Wit86] Rainer Wittmann. Natural densities of Markov transition probabilities. Probability Theory and Related Fields, 73(1):1–10, ???? 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01845990.

Withers:1987:CLT

[Wit87a] C. S. Withers. Central limit theorems for dependent variables, II. Probability Theory and Related Fields, 76(1):1–13, September 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF00390273.

Wittmann:1987:SMT

[Wit87b] Rainer Wittmann. Sufficient moment and truncated moment conditions for the law of the iterated logarithm. Probability Theory and Related Fields, 75(4):509–530, August 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00320331.

Wong:2020:UTP

[Won20] Mo Dick Wong. Universal tail profile of Gaussian multiplicative chaos. Probability Theory and Related Fields, 177(3–4): 711–746, August 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-020-00960-3.

Wu:2004:RBS

[WT04] Tiee-Jian Wu and Min-Hsiao Tsai. Root n bandwidths selectors in multivariate kernel density estimation. Probability Theory and Related Fields, 129(4):537–558, August 2004. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
004-0357-8.

Wu:2000:NML

[Wu00] Liming Wu. A new modified logarithmic Sobolev inequality for Poisson point processes and several applications. Probability Theory and Related Fields, 118(3):427–438, November 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
PL00008749.

Wu:2004:ESR

[Wu04] Liming Wu. Essential spectral radius for Markov semigroups (i): discrete time case. Probability Theory and Related Fields, 128(2):255–321, February 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-003-0304-0.

Wuthrich:1998:SIC

[Wüt98] Mario V. Wüthrich. Scaling identity for crossing Brownian motion in a Poissonian potential. Probability Theory and Related Fields, 112(3):299–319, November 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050192.

Weinryb:1993:TCL

[WY93] S. Weinryb and M. Yor. Théorème central limite pour l’intersection de deux saucisses de Wiener indépendantes. (French) [Central Limit Theorem for the intersection of two independent Wiener sausages]. Probability Theory and Related Fields, 97(3):383–401, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01195072.

Wong:1987:MMD

[WZ87] Eugene Wong and Moshe Zakai. Multiparameter martingale differential forms. Probability Theory and Related Fields, 74 (3):429–453, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00699099.

Wong:1989:IGM

[WZ89] Eugene Wong and Moshe Zakai. Isotropic Gauss–Markov currents. Probability Theory and Related Fields, 82(1):137–154, June 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00340015.

Wolf:1993:OMF

[WZ93] Eddy Mayer Wolf and Ofer Zeitouni. Onsager Machlup functionals for non trace class SPDE’s. Probability Theory and Related Fields, 95(2):199–216, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01192270.

Wu:1999:SCI

[WZ99] Y. Wu and M. M. Zen. A strongly consistent information criterion for linear model selection based on M-estimation. Probability Theory and Related Fields, 113(4):599–625, March 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050219.

Wu:2013:MSS

[WZ13] Zheyang Wu and Harrison H. Zhou. Model selection and sharp asymptotic minimaxity. Probability Theory and Related Fields, 156(1–2):165–191, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-012-0424-5.

Xiao:1997:HCL

[Xia97] Yimin Xiao. Hölder conditions for the local times and the Hausdorff measure of the level sets of Gaussian random fields. Probability Theory and Related Fields, 109(1):129–157, September 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s004400050128.

Xia:2004:PPC

[Xia04] Aihua Xia. Point processes characterized by their one dimensional distributions. Probability Theory and Related Fields, 128 (3):467–474, March 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-003-0314-y.

Xia:2007:APC

[Xia07] Aihua Xia. Acknowledgement of priority concerning “Point processes characterized by their one dimensional distributions”. Probability Theory and Related Fields, 138(1–2):323, May 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00440-006-0017-2.

Xu:1991:LCB

[Xu91] Jianming Xu. The lifetime of conditioned Brownian motion in planar domains of infinite area. Probability Theory and Related Fields, 87(4):469–487, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF01304276.

Xiao:2002:LTF

[XZ02] Yimin Xiao and Tusheng Zhang. Local times of fractional Brownian sheets. Probability Theory and Related Fields, 124(2): 204–226, October 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400200210.

Xiao:2013:DFD

[XZ13] Yimin Xiao and Xinghua Zheng. Discrete fractal dimensions of the ranges of random walks in Z^{d} associate with random conductances. Probability Theory and Related Fields, 156(1–2):1–26, June 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-012-0418-3.

Yang:2009:LMR

[Yan09] Ming Yang. Lebesgue measure of the range of additive Lévy processes. Probability Theory and Related Fields, 143(3–4):597–613, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-008-0138-x.

Yang:2022:KEN

[Yan22] Kevin Yang. KPZ equation from non-simple variations on open ASEP. Probability Theory and Related Fields, 183(1–2):415– 545, June 2022. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL https://link.springer.com/
article/10.1007/s00440-022-01133-0.

Yan:2024:FOR

[Yan24] Shuo Yan. Fast-oscillating random perturbations of Hamiltonian systems. Probability Theory and Related Fields, 190(3-4): 1199–1260, December 2024. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL https://link.springer.

com/article/10.1007/s00440-024-01302-3.

Yao:2019:ACN

[Yao19] Chang-Long Yao. Asymptotics for 2D critical and near-critical first-passage percolation. Probability Theory and Related Fields, 175(3–4):975–1019, December 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-019-00908-2.

Yasutomi:2002:LTS

[Yas02] Kenji Yasutomi. A limit theorem for sequences generated by Weyl transformation: Disappearance of dependence. Probability Theory and Related Fields, 124(2):178–188, October 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200207.

Yattselev:2013:LDL

[Yat13] Maxim L. Yattselev. Large deviations and linear statistics for potential theoretic ensembles associated with regular closed sets. Probability Theory and Related Fields, 156(3–4):827–850, August 2013. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-012-0444-1.

Yau:1997:LSI

[Yau97] Horng-Tzer Yau. Logarithmic Sobolev inequality for generalized simple exclusion processes. Probability Theory and Related Fields, 109(4):507–538, November 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s004400050140.

Yin:1988:LLE

[YBK88] Y. Q. Yin, Z. D. Bai, and P. R. Krishnaiah. On the limit of the largest eigenvalue of the large dimensional sample covariance matrix. Probability Theory and Related Fields, 78(4):509–521, August 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00353874.

Yilmaz:2011:EAQ

[Yil11] Atilla Yilmaz. Equality of averaged and quenched large deviations for random walks in random environments in dimensions four and higher. Probability Theory and Related Fields, 149(3– 4):463–491, April 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00440-010-0261-3.pdf.

Yin:2014:LCL

[Yin14] Jun Yin. The local circular law III: general case. Probability Theory and Related Fields, 160(3–4):679–732, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-013-0539-3.

Yang:1988:NCD

[YK88] Wei-Shih Yang and D. Klein. A note on the critical dimension for weakly self-avoiding walks. Probability Theory and Related Fields, 79(1):99–114, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/BF00319107.

Yong:1997:FAS

[Yon97] Jiongmin Yong. Finding adapted solutions of forward-backward stochastic differential equations: method of continuation. Probability Theory and Related Fields, 107(4):537–572, April 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050098.

Yong:2008:WPR

[Yon08] Jiongmin Yong. Well-posedness and regularity of backward stochastic Volterra integral equations. Probability Theory and Related Fields, 142(1–2):21–77, September 2008. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
007-0098-6.

Yor:1992:TED

[Yor92] Mare Yor. Tsirel’son’s equation in discrete time. Probability Theory and Related Fields, 91(2):135–152, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01291422.

Yoshida:1992:AEM

[Yos92] Nakahiro Yoshida. Asymptotic expansions of maximum likelihood estimators for small diffusions via the theory of Malliavin– Watanabe. Probability Theory and Related Fields, 92(3):275– 311, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://link.springer.com/
article/10.1007/BF01300558.

Yoshida:1993:AEB

[Yos93a] Nakahiro Yoshida. Asymptotic expansion of Bayes estimators for small diffusions. Probability Theory and Related Fields, 95 (4):429–450, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01196728.

Yoshida:1993:LDP

[Yos93b] Nobuo Yoshida. A large deviation principle for ( r,p )-capacities on the Wiener space. Probability Theory and Related Fields, 94 (4):473–488, ???? 1993. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01192559.

Yoshida:1996:CID

[Yos96] Minoru W. Yoshida. Construction of infinite dimensional interacting diffusion processes through Dirichlet forms. Probability Theory and Related Fields, 106(2):265–297, October 1996. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050065.

Yoshida:1997:MCA

[Yos97] Nakahiro Yoshida. Malliavin calculus and asymptotic expansion for martingales. Probability Theory and Related Fields, 109(3): 301–342, November 1997. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050134.

Yoshida:1999:LSI

[Yos99] Nobuo Yoshida. The log-Sobolev inequality for weakly coupled lattice fields. Probability Theory and Related Fields, 115 (1):1–40, August 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050235.

Yoshida:2004:PME

[Yos04] Nakahiro Yoshida. Partial mixing and Edgeworth expansion. Probability Theory and Related Fields, 129(4):559–624, August 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-003-0325-8.

Younes:1989:PII

[You89] Laurent Younes. Parametric inference for imperfectly observed Gibbsian fields. Probability Theory and Related Fields, 82(4): 625–645, August 1989. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00341287.

Yu:1992:DCH

[YS92] Bin Yu and T. P. Speed. Data compression and histograms. Probability Theory and Related Fields, 92(2):195–229, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01194921.

Yogeshwaran:2017:RGC

[YSA17] D. Yogeshwaran, Eliran Subag, and Robert J. Adler. Random geometric complexes in the thermodynamic regime. Probability Theory and Related Fields, 167(1–2):107–142, February 2017. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00440-015-0678-9;

http://link.springer.com/article/10.1007/s00440-015-
0678-9.

Yu:1993:GCL

[Yu93] Hao Yu. A Glivenko–Cantelli lemma and weak convergence for empirical processes of associated sequences. Probability Theory and Related Fields, 95(3):357–370, ???? 1993. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01192169.

Yu:2023:PSR

[Yu23] Lei Yu. On the Φ-stability and related conjectures. Probability Theory and Related Fields, 186(3–4):1045–1080, August 2023. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-023-01209-5.

Yukich:1987:SLT

[Yuk87] J. E. Yukich. Some limit theorems for the empirical process indexed by functions. Probability Theory and Related Fields, 74(1):71–90, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01845640.

Yukich:1995:AET

[Yuk95] J. E. Yukich. Asymptotics for the Euclidean TSP with power weighted edges. Probability Theory and Related Fields, 102 (2):203–220, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01213389.

Yurinsky:1999:SBL

[Yur99] V. V. Yurinsky. Spectrum bottom and largest vacuity. Probability Theory and Related Fields, 114(2):151–175, May 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s440-1999-8035-2.

Yuri:2011:OIR

[Yur11] Golubev Yuri. On oracle inequalities related to data-driven hard thresholding. Probability Theory and Related Fields, 150(3–4): 435–469, August 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-010-0280-0.

Yang:2021:RBM

[YY21] Fan Yang and Jun Yin. Random band matrices in the delocalized phase, III: averaging fluctuations. Probability Theory and Related Fields, 179(1–2):451–540, February 2021. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-020-01013-5.

Zahle:1986:CMR

[Zäh86] M. Zähle. Curvature measures and random sets II. Probability Theory and Related Fields, 71(1):37–58, January 1986. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00366271.

Zahle:1988:SSR

[Zäh88] U. Zähle. Self-similar random measures. Probability Theory and Related Fields, 80(1):79–100, December 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00348753.

Zahle:1998:IRF

[Zäh98] M. Zähle. Integration with respect to fractal functions and stochastic calculus. I. Probability Theory and Related Fields, 111(3):333–374, July 1998. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050171.

Zaitsev:1987:GAC

[Zai87] A. Yu. Zaitsev. On the Gaussian approximation of convolutions under multidimensional analogues of S. N. Bernstein’s inequality conditions. Probability Theory and Related Fields, 74 (4):535–566, April 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF00363515.

Zaitsev:1988:ECS

[Zai88] A. Yu. Zaitsev. Estimates for the closeness of successive convolutions of multidimensional symmetric distributions. Probability Theory and Related Fields, 79(2):175–200, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00320918.

Zambotti:2000:AAE

[Zam00] Lorenzo Zambotti. An analytic approach to existence and uniqueness for martingale problems in infinite dimensions. Probability Theory and Related Fields, 118(2):147–168, October 2000. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s440-000-8012-6.

Zambotti:2002:IPF

[Zam02] Lorenzo Zambotti. Integration by parts formulae on convex sets of paths and applications to SPDEs with reflection. Probability Theory and Related Fields, 123(4):579–600, August 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200203.

Zambotti:2004:FIM

[Zam04] Lorenzo Zambotti. Fluctuations for a Δϕ interface model with repulsion from a wall. Probability Theory and Related Fields, 129(3):315–339, July 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0335-1.

Zeitouni:1990:CVF

[ZD90] O. Zeitouni and A. Dembo. A change of variables formula for Stratonovich integrals and existence of solutions for two-point stochastic boundary value problems. Probability Theory and Related Fields, 84(3):411–425, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF01197893.

Zerner:2005:MER

[Zer05] Martin P. W. Zerner. Multi-excited random walks on integers. Probability Theory and Related Fields, 133(1):98–122, September 2005. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/s00440-004-0417-0.

Zessin:1988:ABF

[Zes88] Hans Zessin. On the asymptotic behaviour of the free gas and its fluctuations in the hydrodynamical limit. Probability Theory and Related Fields, 77(4):605–622, ???? 1988. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00959620.

Zhan:2004:SLE

[Zha04] Dapeng Zhan. Stochastic Loewner evolution in doubly connected domains. Probability Theory and Related Fields, 129 (3):340–380, July 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-004-0343-1.

Zhang:2007:VAF

[Zha07] Xicheng Zhang. Variational approximation for Fokker–Planck equation on Riemannian manifold. Probability Theory and Related Fields, 137(3–4):519–539, March 2007. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-006-0001-x.

Zhang:2010:SRB

[Zha10] Xicheng Zhang. A stochastic representation for backward incompressible Navier–Stokes equations. Probability Theory and Related Fields, 148(1–2):305–332, September 2010. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
009-0234-6.

Zhang:2014:LDR

[Zha14] Yu Zhang. Large deviations in the reinforced random walk model on trees. Probability Theory and Related Fields, 160(3–4): 655–678, December 2014. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-013-0537-5.

Zhan:2015:RWP

[Zha15] Dapeng Zhan. Reversibility of whole-plane SLE. Probability Theory and Related Fields, 161(3–4):561–618, April 2015. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/article/10.1007/s00440-
014-0554-z.

Zhan:2016:ETS

[Zha16] Dapeng Zhan. Ergodicity of the tip of an SLE curve. Probability Theory and Related Fields, 164(1–2):333–360, February 2016. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-014-0613-5.

Zhai:2018:HDC

[Zha18a] Alex Zhai. A high-dimensional CLT in W_{∈} distance with near optimal convergence rate. Probability Theory and Related Fields, 170(3–4):821–845, April 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-017-0771-3.

Zhang:2018:LTM

[Zha18b] Yu Zhang. Limit theorems for maximum flows on a lattice. Probability Theory and Related Fields, 171(1–2):149–202, June 2018. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-017-0775-z.

Zhan:2019:OHC

[Zha19] Dapeng Zhan. Optimal Hölder continuity and dimension properties for SLE with Minkowski content parametrization. Probability Theory and Related Fields, 175(1–2):447–466, October 2019. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/s00440-018-0895-0.

Zhang:2020:OBC

[Zha20] Deng Zhang. Optimal bilinear control of stochastic nonlinear Schrödinger equations: mass-(sub)critical case. Probability Theory and Related Fields, 178(1–2):69–120, October 2020. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL https://link.springer.com/article/10.
1007/s00440-020-00971-0.

Zhan:2021:SLM

[Zha21] Dapeng Zhan. SLE loop measures. Probability Theory and Related Fields, 179(1–2):345–406, February 2021. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00440-020-01011-7.

Zheng:1995:LDR

[Zhe95] Weian Zheng. A large deviation result for a class of Dirichlet processes. Probability Theory and Related Fields, 101(2):

237–249, ???? 1995. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01375827.

Zhou:1988:TSB

[Zho88] M. Zhou. Two-sided bias bound of the Kaplan–Meier estimator. Probability Theory and Related Fields, 79(2):165–173, September 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-

2064 (electronic). URL http://link.springer.com/article/
10.1007/BF00320917.

Zieba:1988:NCF

[Zie88] Wieslaw Zieba. A note on conditional Fatou Lemma. Probability Theory and Related Fields, 78(1):73–74, ???? 1988. CO-DEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00718036.

Zygouras:2009:LNR

[Zyg09] N. Zygouras. Lyapounov norms for random walks in low disorder and dimension greater than three. Probability Theory and Related Fields, 143(3–4):615–642, March 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/s00440-008-0139-9.